A Complete Bibliography of Publications in the Journal of Cold War Studies

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

26 September 2023 Version 2.07

Title word cross-reference

$100.00 [Her16, Wal12b]. $110.00 [DF12, McM12a]. $120.00 [Yor17b]. $125.00 [KG11]. $130.00 [Onu14]. $131.95 [Cli07]. $135.00 [Sas14b]. $138.00 [Kan12a, Whi13]. $140.00 [Blu15]. $145.00 [Jer16b, Nei14, San16]. $150.00 [Dar16, Kul13, Sar16]. $16.95 [Smi11]. $18.00 [Gru07]. $18.95 [Tei11]. $19.95 [Lak11, Gaf11, Ham15, Oli11, Mar07d, Nor07, Sha07a]. $20.00 [Ket14]. $21.99 [Duk11]. $22.50 [Joh13a, Lin11, Joh07b]. $23.94 [Gun12]. $23.95 [Kro12, Wei12]. $24.95 [Boy11, Ree16, Bal17, Hoc16, Jer13, Jer16a, Kap17a, Lub11, Pod14, Vuj02, Bra14]. $25.00 [Kur11, Lou16, Woo12]. $25.95 [Her07b, Neu11, Usd11]. $25.99 [Bar07]. $26.00 [Fes07, Jer11]. $26.95 [Sch11, Joh07c, Pos07]. $27.00 [Lee21, Cra17, Mal15]. $27.50 [Kim07]. $27.95 [Gra12a, Whi13, Bel11, Gen17, Gra12b, Hou16, Par16, Woo13, Zhu16, Lit11b]. $27.99 [Rus16]. $28.00 [Bru11]. $28.50 [Sen16]. $28.95 [Joh11b, Law12]. $29.00 [Hus13]. $29.95 [Gra12b, Gra12a, Kan15, Dei15, Gal07, Gan16, Haw07, Her11b, Kim11, Moi14, Nam13, Nam16, Ros17, Sny11, Spr07, Sur13, Usd16]. $29.99 [Bal15, Kop16]. $30.00 [Gal16a]. $32.50 [Bar11a]. $32.95 [Ken13, Mat16a]. $34.95 [Sou21, Gar13, Mas11a, Mil17b, Nor15, Pra17, Rad16, Rai11, Ram11b, Sch13, Whi11a, Wie07]. $35.00 [Abe21, Whi11b, Zar21, Ale12, Edw16, Kal12, Lie12, Lit11a, McM12b, Rut11, She12c, She13, BSK^{+}23, Liv07]. $35.00/£ [Bra16b]. $35.99 [Rad11]. $36.00 [His13]. $36.95 [Chu07b, Pre17, Win11]. $36.99 [Har12b]. $37.00 [Har11]. $39.95 [Aft13, Ass14, Cal14, DeM12, Gal16b, Guz13, Hit16, Kre12b, Kre12a, Mil15b, Rab13, Wir07b]. $40.00 [Mil21, Ble16, Bur13, Col15, Lat14, Ran12, Whi15]. $42.94 [Onu14]. $43.95 [Spe14]. $45.00 [Car13a, DeR17, Haa13, Kle11, Kuz16, tL16, Pai16, Roh17, Wea12, Hel07]. $49.95 [Dar15, Kru07a, Lee12, Sol13, Whi12]. $50.00 [Lat16a, Mat16b, Sau12]. $52.95 [Rak15]. $55.00 [Byr12, Kan12b, Key16, Whi16]. $55.95 [Pud12]. $58.00 [Lin11]. $60.00 [Kan15, Lie21, Dra12b, yL12, Thi14]. $65.00 [Car11, Dan16a, Lüt12, Rei16, Sto11]. $67.50 [Joh13a]. $69.95 [Fer07, Jen17, Sil11]. $69.99 [Sco17]. $70.00 [Sco17]. $70.00/£ [Cul13]. $72.00 [VA12]. $74.94 [Sta13]. $75.00 [Whi11b, Kip15]. $75.00/EUR [Hei16]. $80.00 [Boy11, Fin13, Mar12]. $84.95 [Gun12, Sha11]. $85.00 [Luc07]. $89.00 [Bla15, Rut12a]. $90.00 [Kuk13]. $90.00/£ [Ram11a]. $95.00 [Lee17, Whi14a]. $99.00 [Car13b, Dom15a, Dou15, Sim16, Tro16b]. $99.50 [Pet11a]. $99.99 [Lat16b].

/the [Mas23].

0Warriors [Gaf07].

1 [Her04b, Kra99b, Mor00b]. 109 [Kru07b]. 12.99 [Wir13, Chu07a]. 129 [Kas16]. 16.00 [Nut11]. 17.95 [Zab16]. 1917 [Nei00]. 1920s [Whi04]. 1920s-1950s [Whi04]. 1921 [Kur05b]. 1928 [Kur15]. 1930-er [Kur21]. 1930s [Rut12a, Ert10a, Ert10b, Kri04, Vai05]. 1934 [Hal07b, Myl15, Pip15]. 1939 [Cea06, Mou02, Nat04, Sch00b]. 1940 [Ing04a]. 1940-kh [Rob08]. 1941 [Gel12a, Pon13, She12c, She13, Del07, Pol10, Sny09, Ava04, Her05]. 1942 [Osg04]. 1942-62 [Osg04]. 1943 [Cha21]. 1945 [Rep12, Rug16b, Spr07, Bis10, Git20, Kor10, Low08, McK15a, Sza10a, Bra05, Mas00, Wil03]. 1946 [Ris15, Val08, Mar06a]. 1947 [Ken13, Jol07]. 1948 [Bor09, Cur07, Mar07a, Gat03, Gra05, Ken03, Kus03, Lef03, Nai02, Sny03a]. 1949 [Mar12, Din23b, Nai08b, Dre05, Jon03, Nai04]. 1950 [Dom15a, Haj12, Mas21b, Thi14, Lee06a, Rea06, Ros09, Lev04]. 1950-er [Kur21]. 1950s [Ass14, Bor99, Cho10, Cly13, Dei13, Jos23, Kaź09, Olš18, Sha02, Zha06, Zin11, Wat18, Hus00, Whi04]. 1951 [Mil21]. 1953 [Hes13, VA12, Yor17b, Akt10, Bon07, Bru09, Lev07, Mar06b, McD06, Pec08, Por09, Wen09, Din02, Mad04, Mat04b, Nor00, Sil04]. 1954 [Wan17, Lar07a, Cog02a, Pru04]. 1955 [Dan16a, Kir16, Bra02a, Van02a, vD02]. 1956 [Ano08i, Nag18, Joh06, Jon06, McC20, Ste20, Sto06, Bor05a, Byr01]. 1957 [Dei07, Mor06]. 1958 [Guz13, Sha11, Zha23b, Zhu16, Ash10, Gar01a, Mil04a]. 1960 [Haa13, Whi11a, Ban07, Luň10, Sha07b, Kru04b, Sfi02, Zha02]. 1960s [Bus14, Din07, Jen17, Lut22b, Olš18, Ric22, McM16, Hen09, Ler15, Gar04b]. 1961 [Ado08, Jon07, Lak11, SS18a, Cha07a]. 1962 [Car13a, Col19, Hes12, Wys12, Aro06, Coo04, Sel02]. 1963 [Lüt06, Mar20, Pen06, Gre05a, Hus05, Mil04b, Sep01, Wil02, Wis00]. 1964 [Goo15, Goo17, Kan12b, Kul13, LX14, SF11, Sie06, Tau02]. 1965 [Chu07b, Kan15, McM12a, Ale09, Sav14, Smi08, Ste10, Tha06, Vei14a, War08, Fis03]. 1966 [Her03b, Sti06, Law08, Yor21, Cat08, Gan06, Sch05, Son02]. 1967 [Col15, Fre07, Nam13, Owe07, Tal11, Doc08, Kal18, Din04]. 1968 [McA07, Mil17a, Pre11b, Cea09, Eas10, Lar17, Lew07, McA08, Rya08a, Cha00]. 1969 [Wal12b, Cal09, Dum20, Joh20a, Sch10, Sar05b]. 1970 [Fer07, Tho11]. 1970s [Blu14, De 20, Hus13, JE19, QG12, Żuk19, Bur16, Qua21]. 1971 [Gan16, Wid05]. 1972 [Gol11, Joh14b, Xia06, Cha14, Coh08b]. 1973 [Cog12, Din14a, Sar16, Yos15, Cha04, Nie03, Wes00]. 1974 [Hat11, Pap23, She14, Soa20b, Van02b]. 1975 [Dav11b, Joh21, Lat16b, Rai11, Mul23, MC23, Law01]. 1976 [Cra13, Dav11a, Nie10a, Dom03]. 1979 [Etc23, Qua23, Wet17, Zha15]. 1980 [Mar07b, Dap14, Lee23, Moo23]. 1980s [Boz20a, Boz20b, JE19, Nun22, Pet20, Żuk19]. 1983 [Kuk11]. 1985 [Mas12, Nie12, Kuz19]. 1986 [Sim23]. 1988 [Win11, Din18, Kaź19, Ros05]. 1989 [Chi14, Gra12b, Gra12a, Kop16, Rug16a, Sim16, Sny11, Cri10, Eva07, Hut19, Sni22, Sto07, Wal21a, Zie08, Eva00, Ged99, Pik02]. 1989/90 [Rug16a]. 1990 [Ble16, Bru08, Alp09, Git07, Kub20, Bru05, Rus05]. 1991 [Rad16, Rei16, Wet11a, Wet11b, Dom15c, Eva09, Vaj20, Cox02, Gen04, Gil05, Hut01, Kar04]. 1992 [Pas03, Wet01]. 1995 [Lar02]. 1996 [Eng02]. 1999 [Val14, Die01, Per04, Ste05].

2 [Her04c, Kra99a, Kra03f, Kra04b, Mor00c, Mue10]. 20 [Pen18]. 2001 [Dom07, Har11]. 2005 [Ser13]. 2006 [Hor07]. 2007 [Pre11a, Duk10]. 2009 [Moy11, Par13, Sar11, Ser13]. 2010 [Myl12]. 2011 [Kol15, Rad12a, Rug16a]. 2013 [Mas16]. 2014 [Tum16]. 2022 [Mas23]. 20th [Cea17, Har23, Pet11a]. 20th-Century [Pet11a]. 21.15 [Lew12]. 21st [Kal14, LX23]. 22.00 [SF11]. 24.95 [Bra16b]. 25.00 [Bar11b, Gra13b]. 28.00 [Rom17]. 29.90 [Mey11, Mil17a, Mol16]. 2nd [Haw07, Nei10b, Ros06, Vuj02].

3 [Kra99c, Kra05a]. 30-Year [Ass23]. 32.00 [Vil16]. 36.00 [Mar16]. 39.90 [She12a]. 3rd [Sny17].

40 [Mar13b]. 40-kh [Odo02]. 40.00 [Cul13]. ’44 [Cie07]. 45.00 [Ram11a]. 47 [Ald07]. 47.00 [Hei16]. 49 [Sch02b].

5 [Gar02]. 50.00 [Tho11]. 50.00/$85.00 [Has07]. 52.25 [Lin11]. 53 [Van99a]. 535pp [BSK^{+}23]. 55.00 [Neh11]. 57.00/£ [Hei16].

60-kh [Odo02]. 60.00 [Goo15, Goo17]. 61 [Ald08]. 62 [Osg04, Smi03]. 63 [Nue06]. 64 [Cou04]. ’68 [Nie10a]. 68.00 [Imm16]. 68er [Mil17a]. 69 [And20, Spe18].

73 [Qua09]. 74 [Ben21]. 75 [Moi14, Dav10]. 78.00 [Cra07]. 79 [Rey23].

83 [Mil20]. 89 [Wei04a].

90 [Rug16a]. 91 [Tho08, Zub99]. 98 [Cla01].

în [Sni22].

A&M [Col15, Gaf11, McA07]. A. [Byr12, Cri20a, Fin13]. Aaron [Hof20]. Abandoning [Ale07]. Abduction [Ess15]. Abingdon [Jon14]. Able [Mas09, Mil20]. Abolition [Gly06]. Abolitionism [Eva23]. Abrams [Har11]. Abroad [DF12, God11, Kuz19, Zhu16, Kru08, Mil15a, Bra02b, Bre08]. Absher [Aft13]. Absolute [Ric09]. Abstract [Ano04a, Ano04b]. Abstracts [Ano99a, Ano99b, Ano99c, Ano00a, Ano00b, Ano00c, Ano01a, Ano01b, Ano01c, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano04c, Ano04d, Ano04e, Ano05a, Ano05b, Ano05c, Ano05d]. Abwanderung [Wet09]. Abyss [Her07b]. Abzug [Zar21]. Academia [Arg04]. Academic [Arg04, SS14, Wys12]. Academy [Lun04]. Accent [Kau03]. Acceptable [Col07]. Accepting [Sto15]. Accession [McK08]. Accidental [Rus16]. Accommodate [PS20]. Accords [Dav11b, Mas23, Mar21]. Account [Sza10b, Wet01, Sto11]. Accused [Mil14]. Acheson [Har08, Kau03]. Achievement [Pec13b]. Achille [Joh07c]. Acquisition [Eng07]. Act [You23]. Action [Cob18, Cor14, Elm04, Gar01a, Joh07d, Yor23a]. Active [Neh21, Sel21a]. Activism [Boe21, Kom18, Mic17]. Activists [Li18, Pet20, Sch19]. Activities [Alo05, Cur23, HK09, Łuk23, Ris15, Tro16a]. Actors [Sch08]. Ad [Win15]. Adam [Kos15, Tis15, Kra00]. Adamic [Lic20]. Adams [KH23, Ric15]. Adapting [Hop03]. Adaptive [Dob10b]. Addicted [Sch11]. Adelman [Dan15]. Adenauer [Sch05]. Adinolfi [Gun12]. Administration [Ces21, Lat08, Mar23b, Mit07, All14, Bar19, Bro02b, Bur05, Bur21, Gue17, Han17, Joh07a, Jon19, KW09, Lin09, Lud13, Lum04, Mat15, McD02, Njø02, PL22, PR21, Kau09, Sch17, Søn19]. Administrations [Mar05]. Adrian [Rey23, Rog15]. Advance [Zei08]. Advanced [Gra21a]. Advancing [Mas23, Tan05]. Advent [Bis05]. Adventure [Pit23]. Adventures [Dan14]. Adversaries [Bro02a]. Adversary [Hat09, Luk07]. Advertising [Bro10]. Advice [Ran10]. Adviser [GSP^{+}19]. Advisers [Gaw21, Kap16]. Advisory [Aft13]. Affair [Mak17]. Affairs [Dan16a, Dav11b, Hor07, Jos20, Kuk11, Ros05]. Afgantsy [Gra12b, Gra12a]. Afghanistan [Bla05, Bra16a, Gra12b, Gra12a, Sai02, DF19, Gra12b, Gra12a, Jer11, Kal09, Kli22, Pra99, Win17, Bea06, Van99b]. Afraid [Sci20]. Africa [Bri15, Cog12, DeR17, Dom03, KG11, Kan12b, Kan15, Lar17, Mul23, Nie10b, Yor15, DFK^{+}20, Gle06a, Gle06b, Gru07, Mil13, Rei19, Sch22a, Tel18]. Africa/L’Amérique [Cog12]. African [Hut20, Mue17, Par16, Smi03, The15, Whi16, Li18, MC23]. Africans [Par16, Whi16]. Afrique [Cog12]. After [Ada03, Fed17, Leb03, Mid00, Sim23, AS22, Bal05, BS02, Boy23, Con07, DF12, Dei14, Din14b, Dou15, Gre08, Gut23, Ham18, Hei06, Her23, Jer08, JMG08, Mad08a, Mas21b, MSPN21, Mic08, Ree16, Sta13, Sza10a, Tho21, Wol13, McM12a, Oli11, Dei02, Sfi02]. Afterlife [Chi14]. Aftermath [Bud22, Pos07, Pol05]. Aga [Dra12b]. Aga-Rossi [Dra12b]. against [Bac02, Blu14, Bus14, Eva13, Fet20, FF22, Ing04b, Kru05, Mac06, McC20, Pen06, Ric07, Sur13, Zak10, Mad08b, Lee21]. Age [Ans17, Cap08, Her03a, Jen03, Sha18, TJ21, Wat18, Bar22, Boy11, Pol07a, Her99]. Agencies [Ric09]. Agency [Kel19, PS20, Kru07a, Pud12]. Agenda [Bro02b, Dar12, Mas19b, Mor00a, Tan05, Ula99]. Agenda-Setting [Dar12]. Agendas [Kir16]. Agent [Bra14, Kle23, Tsy08]. Agents [Eck07b, Kap16, Wal12a, Cap08, Tra09, Wem21]. Aggression [Che20]. Agitators [Hol06]. Agius [Eng09]. Agnes [Eck07b, Mer13]. Ago [Ger13]. Agreement [Ass07, Law05]. Agyeman [Woo12]. ägyptischer [Ste20]. Ahab [Eth05]. Ahern [Bur13]. Ahlberg [Man10]. Aid [Bod08, Fer11, Gan06, Lou23, Mid11, Sto14, Zha15]. AIDS [Sel19, Sel21b]. Air [Her11a, Poc03, Ric07, Sha18, Wys18, Zal04]. Air-Atomic [Sha18]. Aircraft [Usd22]. Airlift [Gra13a, Kuk13]. Airman [Pfa02]. Akteure [Huh20, Sch08]. al [Cap08, Din14b, Pet18, Rug18, Sta13, Lei08]. al-Qaeda [Lei08]. al. [Yor20a]. Alamos [Nor09, Kru07b]. Alan [Rai11, Low08, Sch09]. Albania [Boy23, Mar22, Mëh11a, Neu23]. Albanien [Las19]. Albaniya [Boy23]. Albert [Abe21, Huh20]. Albertson [HK20]. Alchemy [Cra21, Len12]. Aldershot [Eng07, Lin11]. Aldo [Dra06]. Alecu [Bis12]. Aleksandr [Mue10]. Alenius [McC20]. Alessandro [Dom20, Her12]. Alex [Sny09]. Alexander [Kir09, Rad15, HK09, Kno20, Sza10a]. Alexei [Nei10a]. Alexievich [Bra16a]. Alfred [Mar23a, Pos07, Gal16b, Nai16]. Alger [Kle16, Fet20, KLHK14, Mar08a, Mar09]. Algeria [Gif05, Mor06, Yor18]. Algerian [Gos08]. Algiers [CS07]. Ali [Ric12]. Alienation [Bak09]. Aligned [Jak16, Lüt16]. Alignment [Raj14, ŽČ16, Geh05]. Alignments [HF14]. Alike [Hel22]. Alison [Mac20b, Jos20]. Alive [ST17]. Allan [Suh08, Wea12]. Allcock [McP19]. alle [Lar07a]. Allen [Kho14, Lee21, Pra14, Wal21b, vD21, Bar11a, Kho14]. Allende [Gus10, HT19, Lov22]. Alliance [Bra05, Bro02a, Coh12, Kno20, Kuz16, Lou23, Lüt22a, Nau04, Pen06, Pud12, Rea11, Sch05, Wis00, Xia23, All14, Cur15, Dar12, FF22, Hug08, McK08, Svi16, Tro16a, Zha12b, Kap04, Kli05, Von03]. Alliances [Mar22, Mas02, Sön22, Nue05]. Allianz [Nue06]. Allied [Bre08, Fei10, PL22, Bor12b, Kou12]. Allies [Bre21, Del22, God11, Hel22, Bro02a, Sim04]. Allowed [Fei10]. Alltag [War06]. Ally [Sae07]. Almost [Ole18]. Alone [Koc10]. Alpo [Mas19a]. Altered [Sch02c]. Alternative [Tum16, TSW^{+}22]. Alternatives [Sen16]. Aluminum [Mar15]. Alvandi [Key16]. Alyosha [Cob18]. Am [Sci20, Bis11, Lar07a]. Ambassador [Kan10, Taa13, Wei04a]. Ambassadors [Tei11]. Ambition [McM22, Hes21]. Ambitions [Bla15, Vu19]. Ambivalent [Cho10, Sch05]. Amendment [Mar21]. Amerasians [Sch22b]. America [Ale07, Bar03b, Bel09, Bel13, Bro10, Bro04, Byr01, Dar15, Dar16, Dav09, Dei14, Din02, Din23a, Dom20, Dom08, Dom15b, Eva00, Gra13a, Hay99, Her99, Jer13, Jer16a, Law10, Lee06b, Lov10, Low08, Mar20, Mat16b, Nau04, Nie10a, Nor07, Par13, Pas03, Pen06, Poc03, Pru04, Rob20, Rus04b, Sau12, Sch02a, Sha08a, Shm23, Sur13, War08, Wat17, Wat18, Zei08, All14, Bal15, Bel11, Car99, Cog12, Din13, FF22, Har19, Har11, Joh13a, Lov22, Mar07c, McC14, Moy11, Rem08, Sae07, Sho15, Soa06, Sön22, Whi12, Yar09, You12a, Zim09, Bar20, Boc06, Gil06, Iac22, Kau02, Lee21, Wee18, Zie21, Duk02, Fin04, Lun04, McM05, War01]. American [Alo05, Ans17, Bac09, Bai18, Bak09, Bel02, Blu15, Boc10, Bos15, Bra02b, Bur16, Bur08, Cal06, Cal09, Che19, Cli07, Cob18, Cra17, Cri10, Dan10, Dav12, Dei02, Die01, Din04, Din07, Din18, Dom07, Dom09, Dum17, Fer15, Fes07, Gar02, Gar01b, Gas06, Gem18, Gen17, Gle05, Gre10, Haa20, Haf08b, Her17, Her04a, Hes21, Hoc16, Hor07, Hut20, Iri14, Jen17, Jer11, Joh03, Joh14b, Kal12, Kap14, Kau03, Ken05, Kir16, Kle22, Kra04a, Kuk21, Lan10, Laz07, Lem14, Lin15, Lom09, Low04, Low06, Mad07, Mar07a, Mar08b, Mar07d, Mat04b, Mav18, McE03, McP19, Mil06, Mur02, Nic12, Nin03, Nor10, Nor20, Nut02a, Osg04, Per04, Pra21, Pre17, Pud12, Rak15, Ram01, Rec17, Rep12, Rog15, Ros04, Rus04a, Sap03, Sch02b, Sha18, Sha08a]. American [She07, Sib11, Sin18, Smi07, Smi11, Sny17, Soa20a, Sto11, Sto06, Sza10a, Sza10b, The15, Usd11, Val08, Vau15, Vei14b, Vit04, Wal07, Wal17b, Whi03a, Whi04, Whi11a, Whi15, Wil08b, Win15, Wys12, Aro06, Bor05b, Cho10, CKP05b, CKP05a, Dar12, Her13, Ing04b, Joh11a, Kam13, Kur13, Kus02, Li18, tL16, Luk07, McB15, Mic11, Mid00, Ole11, Ria17, Rin12, Sha07a, Tod17, Xia06, You07, Zha06, Cap14, Her03a, Hes02, Jon03, Wet01]. americana [For02]. Americanism [Pay21, Sza10a, Kom18]. Americanization [His13, Sza10a]. Americans [Boy11, Haa13, Hat10, Kle08, Qua21, Roz11, Wel06, Wen09, Lon11, Sci20, Kur14]. Americas [Rya08b, Rya08b]. Amerika [Lar07a]. amerikanischen [Haf08b]. Amérique [Cog12]. Amherst [Boy11, Gen17, Law12, Sch11, Whi11a]. amid [Nag17, Nag18, RTYD19]. among [Chu19, Dan12b, Ric09]. Amsterdam [Taa13]. Amt [Rug18]. Amts [Rug16a]. Amy [Buh15, Rut12a]. Ana [Wei02]. Analogies [Gar04c, Tie07]. Analysen [vD05]. Analysis [Alp09, Gra17, Haa05, SY17]. Analyzing [Ber13]. Anastasakis [Pap23]. Anatomy [Puc21, Her12, Lib01]. Ancel [Del07]. Ancestries [Law07]. Andalucía [Sti06]. Anders [Mol16].

Anderson [Ass10, Kau02, Mat10b, Pre11a, Sza12a]. Andrea [Rut10, Gol11]. Andreas [Mas19b, Dav10, Dra22, Hen09, Nue07, Rug16a]. Andrei [McM11, Sni22, Gra13a, Zub99]. Andrew [Cal09, Ger14, Lat08, Wem21, Zha15, Bac09, BMLG11, Bal08, Bla15, Dal10, DeM12, Hol15, Kee00, Mil00, Nor20, Sha08a, Sto06, Str12, Usd11, Whi11a, You12b]. Andrews [Hal07b, Soa09]. Andropov [Roz03]. Andrzej [Kor06, Sim16]. Ang [Tay19]. Angelo [Kap14]. Angleton [Ric10, Joh13b]. Anglo [Dei02, Kir16, Nor10, Pec13a, Pra21, Smi03, Gol11, Mid00, You07, Gas06, Mur02]. Anglo-American [Kir16, Nor10]. Anglo-Australian [Gol11]. Anglo-Soviet [Pec13a]. Angola [Dom15c, Rei19, Sob19, dS19]. Angolan [Mil13]. Anguelov [Nei07]. Animal [Ros10b]. Anita [Cur07]. Anja [Nai10b]. Ann [Dan11, Ing17]. Ann-Marie [Ing17]. Anna [Hen09, Clo09, Eng20a, Szp11]. Annals [Fie03]. Annapolis [Gal16b, Kas16, Rak15, Ram11b, Rub11, Win11]. Anne [Iac22, DF12, Kur11]. Anne-Emanuelle [Iac22]. Annegret [Cal14]. Annelise [Mat10b, Kau02]. Annette [Whi14a]. anni [Nut10, Nut11]. Anniversary [Sza12a]. Anspruch [Rug16b]. Answer [Mor00a, Nie11]. Anthology [Kur11]. Anthony [Moi15, Rem08, Sha20, Pra21, Rem08]. Anthropology [Ros17, Ros17]. Anti [AE02, Bra04, Buc17, Cor14, Cos22, Fei10, Fou14, Gil18, Hay00, Jos20, Kom18, Lew07, Lyn16, Mar06a, Nun22, Rob14, Rub06, Rya10, Sha08a, Sza10a, Zha23a, Whi10]. Anti- [Jos20]. Anti-Americanism [Sza10a, Kom18]. Anti-Communism [Cos22, Fou14, Lew07, Rob14, Rya10, Hay00, Lyn16, Whi10]. Anti-Communist [Buc17, Gil18, Sha08a, Cor14]. Anti-Cosmopolitan [AE02]. Anti-Fascist [Bra04, Rub06]. Anti-Imperialism [Nun22]. Anti-Nazi [Mar06a]. Anti-Peaceful [Zha23a]. Anti-Semitism [Fei10, AE02]. Anti-Stalinism [Lyn16]. Anti-Westernism [AE02]. Antiaircraft [Her16]. Anticommunism [Hay13, Her03a, Kle22, TJ21, Wee18]. Antikolonialismus [Ste20]. Anxiety [Ans17, DX18, Her03a]. Any [Fet20, Kli05, Poc03]. Anyway [Ger13]. ANZUS [RG15]. Apartheid [Mue17, vW19]. apertura [For02]. Apolitical [Whi03a]. Apor [Nai10a]. Apostles [Hol06]. Apparat [Gil05]. Apple [Hay15b]. Applebaum [Kur11]. Appleby [Mag14a]. Applying [Din13]. Appraising [Alk05]. Appreciation [Her11a]. Approach [Cor14, Gil00, Hof06, Jon19]. Approaches [Wea15]. Appropriation [Pos19]. April [Mas21b, Mil21, Moi14, Thi14]. Ära [Mas19b]. Arab [Din04, Din14a, Fre07, Joh07b, Kal18, Lak11, Lar10, Qua13, Qua23, Rad12a, Din22, Hug08, Kob06, She20]. Arab-Israeli [Fre07, Lak11, Qua13]. Arabia [Mat20b]. Arabists [Bal15]. Arabs [Qua21, Kar04]. Arbeit [O’S17]. Arbor [Dan11]. Arch [Kul10]. Archduke [Wan09]. Archer [Mas09, Mil20]. Archie [Mat08, Rad11, Rut06, Sav09]. Archipelago [Bar03a]. Architect [Cal06, Hes21]. Architects [Gav05]. Archival [Mag10, Sny03b]. Archive [Rad12a, Wir13]. Archives [Hut20, Ler04, Mar09, Kar15c, Kip15, Pon01a]. ARCI [Oye14]. Arctic [Gol12, Far15]. Arena [Ros04]. Arévalo [FF22].

Arguing [Pay21]. Arguments [Fet20]. Arie [Dom07]. Arkhangelsk [Gol12]. arkhivov [Kar15c]. Arktike [Gol12]. Armageddon [Fer07, Wel06]. Armaments [CK13]. Armas [Sie06]. armata [Dra08]. Armchair [Bai18]. Armed [Dra08, Sew23, Kuk21]. Armee [Bis10]. armi [Eva09]. Armies [Gru07, McA14, Spr07]. Armin [Wen09]. Arming [Fre01]. Arms [Dan12a, Her16, Kal22, McD02, Mit07, Pra15a, Pra15b, She12c, She13, Tra22, Wis00, Zha15, Che10]. Armstrong [Lee06a, Sch09]. Army [Bak09, Don05, Sch11, Bak09, Car99, Hit16, Wei01]. Arne [Hol21, Nai22, Nie12, Nie08]. Arno [Rub06]. Arnold [Dav09, Lie21]. Arras [SF11]. Arrested [Myl12]. Art [BS02, Bro22, Kan10, Kus02, Mar08b, Rot21, Whi12, Whi15]. Artan [Neu23]. Artemy [Bit17, Gra12b, Gra12a, Kan12a]. Arthur [KH23]. Article [Mil00]. articles [Kar15c]. Artificial [Kan10]. Artist [Sta07]. Artists [Chu19, Sto06, Sto06]. Artois [SF11]. Arts [Kan10, Mar07b]. Arvid [Wet17, Zie08]. as-Ser [Yos15]. Asaf [Qua14, Gar09]. asasinat [Del07]. Ascendance [Joh11a]. Ashes [Chi14]. Ashgate [Cro07, Eng07, Lin11, Stu15]. Ashton [Ban06, Qua09, Sas14b]. Asia [BJ10, Ben21, Cha07a, Cli15a, Gar04a, Gra07, Hac03, Hes12, Hus00, Jon07, Kho14, Kna02, Lef21, McM12a, Rad14, Sch13, Soa20b, Spr07, Tay19, BJ11, Coh12, Edw16, Geh22, Hal16, Han17, Iri11, Jon08, KSR18, Kra06f, Lüt12, McM06, Rec17, Rub11, Tho21, Wol13, Edg08, Pai14, Ros08, Wil08b]. Asian [Lin15, Mat04b, BJ19, Dee07, Soa20b]. Aspirations [Her21]. Assassination [Iat18, Pos07]. Asselin [Vei14a]. Assessing [Ber13]. Assessment [Kur13, May13, Nei14]. Assessments [Hat09]. Assets [Bre08]. Assignment [Dan22]. Assimilation [Rea06]. Assistance [Čav19, Vám22, Vit04, Dra12a]. Assuming [Lom09]. Assurance [Kno20]. Assured [Sha18]. Astronauts [Har18]. Asymmetry [Dar12]. Athan [Ric09]. Athena [Bac02]. Athens [Bra14]. Atkinson [Kuk11]. Atlantic [Hus05, Mil04b, Hat17]. Atlanticist [Bro02b]. Atom [Bar11a, Mar01]. atomaren [Ban07]. Atomic [Boy11, Her99, Mad15, Mal15, Nor09, Sha18, Kno20, Pan01, Wal21b]. atomiche [Eva09]. Atoms [Cur23, Doc04, Šmi23, Cur23, Her23, Jos23]. Attached [Gan06]. Attachés [Sie06]. Attack [Gal09, Gal16b, Gut23, Sch09]. Attempt [PS20, Pra17, Rad06, Usd22]. Attempts [Mun12]. Attitude [Röh18]. Attitudes [Con03, Jar16, Lan20, Sza18]. Attwood [Dom15b]. Atwood [Che08a, Ler15, Lom09, McM10]. Audra [Sch20]. Aufbau [Ban07]. Aufstieg [Kur21]. Aufzeichnungen [Nai10b]. August [Taa13, Dun03]. Augustine [Ble16]. Aunesluoma [Sal06]. Aurélie [She14]. ausgewählten [Jon06]. Ausland [O’S17]. Aussaresses [Mor06]. Aussen [Sza07]. Aussen- [Sza07]. Außenministerium [Rug18]. Aussenpolitik [Bru06, Rug16b, Ste20]. Austin [Lüt22a, Xia23]. Australia [Ben21, BJ10, BJ19, Cur15, Hal16]. Australian [Gar02, BJ11, Gol11]. Austria [Bis11, Wil03, Wil08a, Geh05, Röh18, Sch16, SM12, Wil07, Bis04]. Austrian [Bis12, Dav11b, MBR^{+}21, SM12, Wil00]. Auswärtige [Rug18]. Auswärtigen [Rug16a, Kan10]. Authoritarian [Dra09, Eas10, PL22, Tor22]. Authorities [Shi13]. Authors [Dan16c]. Automobile [Joh14a]. Avant [She03]. Avant-garde [She03]. Averting [Gre05b]. Avi [Qua13]. Aviation [Con23]. Aviva [Whi11b]. Avoiding [Fer07]. Avoids [Woh99]. Awaiting [Wel06]. Away [Mun12]. Ayatollah [Fri18]. azerbaidzhanskaya [Min19]. Azerbaijan [Val08]. Azorian [Kas16].

B [Bis14, Buh15, Dei14, Dum20, Gaf07, God14, Gru07, Her19, Hes21, Hof20, Imm16, Lei08, McM21b, Moï10, Nei10b, Pos07, Ran12, Sch19, Sha11, Taa13, Tha06, The15, Wal12b, Wie07, You13b]. Bacevich [Bal08, BMLG11]. Bach [Bri15]. Bachmann [Ste20]. Back [Lar10, RT13, Zub13, Lou16]. Back-Channel [RT13, Zub13]. Backdrop [Mac06]. Backyard [Pas03, Lee17]. Bad [Bar20]. Badel [Mar13b]. Baden [Dav11b]. Baden-Baden [Dav11b]. Baghdad [SJ23]. Bagley [Cla09]. Bailey [Wil08b]. Baines [Nie10b]. Balance [Mat16a, Ole18, Owe07, Sap03]. Balancing [Cos22, Jer00, Mei15, You23]. Balázs [Lee12, Nai10a]. Balint [Bur23]. Balkan [Hay17]. Balkanite [Boy23]. Balkans [Boy23, Mar23a, Pet18, Yor20a]. Ballantine [Her07b]. Ballet [Zhu20]. Baltic [Ber20]. Bamboo [Lüt12, Ros08]. Ban [Mas08, Pie16]. Band [How14]. Bandung [Jak16]. Bange [Mas10a]. Bangladesh [Gan16]. Bank [Sto08]. Bao [Sza18]. Barack [Sim14]. Barak [Pai16]. Bardgett [Din14b, Sta13]. Bargain [Pet08]. Bargaining [Raj14]. Barlow [Dan16a]. Barnes [Ale12]. Barnett [Die07, Mic17]. Barnhisel [Whi15]. Barnsley [Duk11]. Barrass [Gar10b]. Barrett [Whi13]. Barry [Win15, Cli15a]. Base [Fer11]. Based [KH23]. Basic [Bal15, Dan12b, Hay16, Hou16, Rus16, Rut11, She12c, She13]. Basketball [And23]. Basosi [Nut10, Nut11]. Bass [Sza11, Buh15]. Bastiansen [Vil16]. Bateman [Den06]. Ba’thist [DS14, Mat15]. Baton [Dan14, Dei15]. Battle [Buh15, Cie07, Kor10, Kru08, Mor06, Rya08a, Vei13, CS07, Din14b, Rus07, Sta13]. Battlefield [Kob06]. Battlefields [Gol07]. Battleground [Blu14, Kan15, Cro00]. Battles [Din02]. Battling [Che99, Zar21]. Baumann [Whi13]. Bay [Bar19]. Bayard [Joh11b]. Bayly [Edw16, Spr07]. Be [Dan12b, Kel19, Kle23, Kor06, Moy11]. be.bra [Rug16c]. Bear [Nor10]. Beatles [Ric22]. Beatriz [Lov22]. Became [Lee20]. Becoming [Dan22]. Bedingungen [Ban07]. before [Fox09, Her23, Lüt08, Mas21b, Rob04, Thi14, Leb03]. Begin [Bar22]. Beginning [Ale01, Cos22, Mas21a]. Beginnings [Dei14, Mit07]. Begins [Gro02]. Behind [Haf08a, Mic17, Fro23, Haa13, Kis06, Lüt12, Mar15, Rad12a, Ris14, Ros08]. Behrends [Shl06]. Beijing [DeM12, Wal21a, Lüt09, Rei16]. Being [Joh07a]. Beisner [Har08]. Beitrage [Bru05, Nai08a]. Békés [Vaj20]. Belarus [Bud22]. Belated [Mac06]. Belgian [Van11]. Belgrade [Čav19, Jak16, ŽČ16]. Belief [McD02]. Belief-Systems [McD02]. Beliefs [Bla15, Leb06]. Believe [Mas23]. Believer [Cra17]. Belknap [Edw16, Gal07, McM12b, Spr07]. Bell [Mad07, Usd22]. Bella [Zar21].

Belonger [Bro06]. Below [Pet20]. Ben [Spa14]. Bend [Dom07]. Beneath [Col05]. Bengal [Tal11]. Benjamin [Joh22, Kol15, Mey11]. Bennett [Bre21]. Bentley [Hay06, Hay06]. Benvenuti [Gol11]. Bereitschaft [vD02]. Berets [Wil02]. Berger [Dob10b]. Berghahn [Cul13, Han13, McM16, Rai11, Ram11a, Sar11, Ser13, Sza12a, Whi14a, vD12]. Bergholz [Hay17]. Bergman [McM11]. Berkeley [Jon07, Kim07, Lie21, Oli11, Sou21]. Berkhoff [Kes07]. Berkner [Sap03]. Berlin [Ado08, Cro00, Gra13a, Lau00, Mar12, Nue05, Nue07, Rug16c, Rus16, She12a, Sza12a, Tra02, Wet11a, Wet11b, Col00, Har12b, Kuk13, Lar11, Kar15b, Col05]. Berlin-Krise [Ado08]. Berliner [Nue06]. Berlinguer [Bro11]. Bern [Pud12]. Bernd [Bis14, Sza21, BMLG11, BM12, Mas16]. Bernhard [GB10, Mol16]. Bernstein [Her07a, Ran10]. Bert [Har18]. Berthold [Bis11]. Bertjan [Cro07]. Besatzung [Bis10, Wil08a]. beschwerliche [Kra01]. Besier [Pen18]. Bespiel [Lar07a]. Besporyadki [Gor01]. Bestandsaufnahme [Bru09]. Beth [Kur14, Rub06]. Bethe [Aft02]. Betrayal [Lüt09, Dan16b, yL01a]. Better [Kru04a, Moi03, ZRS^{+}21]. Betting [Par16, Whi16]. Between [Bal10, Fav18, Hur10, McA22, Mor00b, Mor00c, Pet08, Rub02, Shl06, Del22, Dom20, Dra08, Fal20, Fri02, Gra07, Hay13, Hof06, Ken13, Kom18, Lou16, Lun04, Mis23a, Neh21, Ped19, Pud12, Rad16, Rua04, Sha11, SJ23, Tis15, Eng20a, Jol07, Rea06, Dap14, Nau04, Won04]. Beyond [Cur15, Law22, Lüt09, Mas23, Mor00a, Hen09, Joh13a, Lüt12, Ros08, Byr12, Nie10b, Bri15, Ric15, Ber04b, Dom15b, Ler15, Tal99]. Beziehungen [Mas19b, Rug16a, Haf08b, Mue10]. Beziehungen/Soobshcheniya [Mue10]. Bibliothek [Bis11]. Bid [Wei01]. Biddle [Neu11, Rus07]. Bidlack [Kip15]. Bien [Nor08a]. Big [Bre21]. Bijedić [Kar15a]. Bilateral [Dei14]. bildenden [Kan10]. Bilder [Sch08]. Bill [Joh11b]. Billion [Dan16b]. Bingham [Bla15]. Binkiewicz [Mar07b]. Biografie [Joh17, vD03]. biografii [Szp09]. biografija [Kar15a]. Biographical [Cap08]. Biography [Ber13, Hay06, Kra13b, May13, Sny07, Szp09, Ale08, Zha09, Gol03]. Biological [Sha20]. Biomedical [Gel12b, Kre07]. Bird [Sha12]. Birdnow [Rya10]. Birn [Iac22]. Birth [Kal14, Šmi23]. bis [Ado08, Ban07, Bru05, Gri10, Her08, Kur21, Lar07a, Las19, Ste20, Wet09, Wet11a, Wet11b]. Bischof [Mol16]. Biser [Boy23]. Bishops [McB15]. Bispinck [Wet09]. Bissau [Mac06]. Bitter [Bea06, Law05, Low08]. Black [And23, Cor22, Joh13a, KG11, Li18, Wet10, BH17, Joh13a, Lew07, Mar12, Moi14]. Blackburn [Mar07a]. Blacklist [Hol08, Sny08, Whi14b, Fal05]. Blackwell [Ash10, Cli07, Bos15]. Blair [Par10]. Blaive [Bis11]. Blanket [Puc21]. Blind [Eng08]. Blitzkrieg [Wir07b]. Bloc [Byr01, Imm16, Ler04, Mar02, Nie11, Sel02, Sel21a, Bék16, Dra12a, HP17, Nie16a, Yor23b]. Blockade [Col19, Kip15, Kuk13]. Blöcken [Bis11]. blokk [Vaj20]. Blomstedt [Mat16b]. Bloody [Coo04]. Bloomington [Luc07]. Bloomsbury [Cea17]. Blows [Tei11]. Blowtorch [Rak15]. Bludeau [Rub06]. Blue [Sha12]. Bo [Tum14]. Bożena [Pol13]. Board [Aft13]. Boat [Gal09]. Boats [Hou15]. Bob [Pav19]. Boden [Wet17]. Bodies [Gar04b]. Böhlau [Bis11, Mil17a]. Bohn [Joh07c]. Bois [Buh15]. Bojkot [Kaź19]. Bolivia [Kuz16, Sie06]. Bollinger [Pol07b]. Bolshevik [Hal07b, Kip03, Nei08]. Bolsheviks [Mou02, Sch00b]. Bolton [Bug15]. Bomb [Aft02, Gei12, Joc07, Kre07, Mal15, Mas00, Nor07, Nor08b, Que02, Sti06, vW19, Bar11a, Ker15, Her20a, Mad09, Nor10]. Bombers [Gra13a]. Bombing [Bar22, Clo06, Her03b]. Bombs [Mad15]. Bonds [CK11]. Bonn [Sza12a, Spo12]. Bonnie [Spe14, Spe15]. Book [And23, Ano01d, Ano01e, Att23, Boy23, Bro22, Bur23, Con23, Cul22, Del23, Din23a, Din23b, Etc23, Eva23, Gil23, Har23, How22, Jac23, Joh22, Kal22, Kle23, Lar22, Lee23, Lüt22a, Mar23a, Mas23, Mis23a, Mis23b, Moo23, Mul23, MC23, Neu23, Pap23, Pee23, Pit23, Qua23, Rey23, Rub06, Sch22b, Sew23, Shm23, Soa22, Soa23, Sul22, Xia23, Yor23a]. Books [Bal15, Bor12a, Cul13, Dan11, Dan12b, Dei13, Fre07, Gel12a, Gra12b, Gra12a, Han13, Har12b, Hay16, Her07b, Hou16, Joh11b, Joh14b, Kol15, Lee17, Lub11, Moi14, Rai11, Ram11a, Rus16, Rut11, Ser13, She12c, She13, Sza12a, VA12, Whi14a, Yor17b, vD12, Joh20b]. Boone [Ste08]. Boose [Stu15]. bor’be [Min19]. Border [Che20, Hof06, Nie10b, SX14, Tag22]. Borderland [Sny07]. Borghese [Mar23b]. Borhi [Bal10, Hut19]. Boris [Gol10, Pip15]. Borisova [Mag14b]. Born [Har15]. Bose [Gan07]. Boskovska [Jak16]. Bostdorff [Che10]. Botschafter [Kan10]. Bottled [Bur17]. Bottoni [Gor09]. Bouchet [Sim14]. Boulder [Gru07, Kru07a, Lin11, Sil11]. Bounties [Her23]. Bowers [Mil21]. Boy [Hay15a, Nor20]. Boyadjieva [Val14, Yor17a]. Boycott [Soa13]. Boyd [Whi13]. Boyd-Barrett [Whi13]. Boydell [Gol11, Nor15]. Boyle [Dei07]. Boyne [Pol06]. Boys [Bra16a]. Bozo [Sar11, Say17]. Bradley [Her11b, Moï10, Eas10, Hay15a]. Bradsher [Dan14]. Brainwashing [Gen17, Sha08b]. Braithwaite [Gra12b, Gra12a]. Branch [Bur04]. Brandeis [Din07]. Brandenburg [Nai02]. Brandt [Nie08]. Brassey [Joh07c]. braucht [Kar14a]. Braun [Neu11]. Brazil [Her04b, Her04c, Lou23, PS20]. Brazilian [Her20b]. Brazinsky [Hat10]. Break [Cal02]. Break-up [Cal02]. Breakdown [HS15, Mav18, Smi03]. Breaking [Kho10, Kot17]. Breaks [Mun10]. Breakthrough [Pie16]. Breakup [Hal07a, Tum03]. Breisgau [Mas16]. Brendon [Haa20]. Brent [Nic12]. Brewer [Bel13]. Brewing [Lan20]. Brezhnev [Bit17, Dom05, Ell03, Roz03, BL10a, BL10b, Gor01, Mue11]. Brezhneve [Gor01]. Brezinová [Huh20]. Brian [Chu07a, Ram11b, Rya10]. Bridge [Kle23, Li18, Lum04, Nor20, Tol10]. Bridger [Roh17]. Bridoux [Cap14]. Brief [Jen03, Ste08]. Briefings [Bar19]. Brigades [Her12]. Bright [Her16]. Bringing [Log01]. Brink [Dei15, Kuk13, Sco09]. Britain [Cat05, Ces21, Cro07, Den21, Doc08, Fre01, Gar04a, Goo15, Goo17, Kir08, Mur07, Bro08, Cea09, Dee07, Dei14, Doc08, Hug08, Kra18, Mac06, Owe07, Sal06, Sha08a, Wys18, Cox02, Hac03, Kna05, Rua04, Sta00, Wil03, Wym03]. Britanskata [Boy23]. British [Ald07, Boy23, Cea06, Cla01, Dav11a, Dum17, Gar02, Joh11a, Kau03, Rab13, Ram07, Sch02b, Sha08a, Spr07, Zha06, Ash10, Cea17, Chu07b, Cor14, Cor22, Dei13, Eng07, Goo10, Jon19, Mid00, Vic08, Mur04, Whi02a]. Broadcasting [Cri99, Cri10, Cri20b, Cri20a, Cul22, Fal05, Haa13, Joh21, Kom18, Pos19, Cri02]. Broadcasts [Din23a, Gue17, Shm23]. Broadening [Bri13]. Brocheux [Zha09]. Brogi [Dom20]. Brookfield [Ber15]. Brookings [Din13]. Brotherhood [CK11, Fel05]. Brothers [Pra14, Wis00, Zha15]. Brown [Din13, Mat08, McM21b, Rad11, Rut06, Rya08a, Sav09, Sny07, Tum16, Wal21a]. Bruce [Eck19, Has07, Nam16, Ped07, WF12, Aro06, Eng08]. Bruckmüller [Bis08]. Bruno [Hay16, Kra00, Ban07]. Bruns [Den21]. Brunswick [Nam13, Pét13]. Brussels [Cog12]. Bryan [Sas14b]. Bryant [Wet10]. Bryn [Sco17]. Brysac [Par10]. Brzezinski [Ste15]. Buchverlag [Zab16]. Buddhist [Cat08, Law08]. Build [Pra17, ZRS^{+}21, Mit07]. Build-Up [Mit07]. Building [Gav05, Hat10, Li18, Lum04, Nut02a, Pre11b, Svi16, Tra22, Uel20, vH09, Tol10, Fri04, Pik02, She14]. Buildup [Kom14, Wys18]. Bulgaria [DS14, Neh22]. Bulgarian [Bae16, Neh21]. Bullitt [Mar07a]. Bullough [Dan12b]. Bundesrepublik [Moe03, Pen18, Rie08, Mil17a]. Bundeswehr [Ban07]. Bündnis [Kra01]. Bündnispolitk [Sza07]. Bündnisses [Jon06]. Bunker [Wie07]. Bunkers [Duk11]. Burden [Lom09]. Bureau [Goo15, Goo17]. Bureaucratic [Mar02]. Burgess [Har18, Smi08]. Buried [Yor15]. Burke [Gil18, Kle16, Nor15, Wal21b]. Burlington [Cro07]. Burnett [Dom15b]. Burning [Suh08]. Burns [Koc10]. Burt [Lub11]. Bury [Pra15a, Pra15b]. Bush [Koc10, Lin11, Sul22]. Business [Fav18, Kru04b]. Buttino [Gra07]. Byrne [Yor18]. Byung [Cli15b]. Byung-Kook [Cli15b].

C [Aft13, Coh06, Cul13, Gaf11, Haf08c, Her21, Her07b, Joh06, Kau09, Kes07, Kir16, Kuk11, Kuz16, Lan10, Lat08, Lub11, Mar07a, McA07, McA08, McM21b, McP19, Mic08, Nei08, Par13, Ros10a, Rub11, Shl06, Soa20a, Spe18, Xia08c]. C.S.C [vDO08]. CA [Dan16a, DeR17, Fre07, Hei16, Kap17a, Lub11, Lüt12]. Cables [Cha21]. Cadres [KM22]. Calico [Lie21]. California [Jon07, Kim07, Lie21, Lub11, Oli11, Sou21]. Call [Che10]. Callaghan [Haa20]. Callahan [Bes07]. Calori [Eng20a]. Cambodia [Cly01, Tur01, Yor23a, Ray20]. Cambodian [Etc23, McA14]. Cambridge [Abe21, Bal17, Bar07, Bel11, Bla15, Ble16, Bru11, Bru16, Cli15a, Dav11a, DeM12, Dom15a, Dou15, Edw16, Fer15, Fin13, Gal07, Gan16, Gra12b, Gra12a, Har11, Has07, Hay11, Hes21, Hes13, His13, Hol15, Hol10, Hor07, Hus13, Iri14, Kal16, Kle16, Kop16, Kuk11, Lat16b, Lef12, Mad15, Mar12, McM12a, McM12b, Mil15b, Nai16, Nam16, Neh11, Pai16, Pre11b, Qua13, Rat21, Roh17, Sae07, Sib11, Sny17, Soa13, Sol13, Spr07, Sza21, Tal11, Tei11, Tro16b, Vuj02, Woo12, Zar21]. Came [Nor15, Wea12]. Campaign [Joh14c, Mar02, Rid16, Sar05b, Sel19, Sel21b]. Campaigns [AE02, Kap17b, Zak10]. Campbell [Mal15]. Camps [Git20, Tod17, Ber07]. Campus [Rom17]. Can [Koc10]. Canada [Her11a, Mas11a, Joh03].

Canadian [Ber15, Don05, Her17, Joc07, McK15b, Ste14]. Canal [Cit14]. Cancer [Fie03]. Candace [Fre07]. Candy [Gra13a]. Cannot [Mil14, Sci20]. Canon [Spr17]. Can’t [Tay15]. Canyon [Ale04]. Capabilities [Jer13, Jer16a]. Capital [SM12]. Capitalism [Mag14a]. Capitalist [Zak10]. capovolta [Gra07]. Captain [Eth05, Rob20]. Capture [Usd22]. Card [Bac02, Mar08c]. Cards [Kle16]. Cardwell [Lef12]. Carl [Wal17b]. Carlos [Sil11]. Carmel [Jer13, Jer16a]. Carnival [Kri05]. Carola [Jos20]. Carole [Sza21, Myl12]. Carolina [Bre21, Dei07, Gar13, Hat11, Hay13, Hel07, Joh07b, Joh14c, Lak11, Lou16, Mar07b, Mar07d, Rab13, Rad16, Ram07, Sha07a, Str12, Wie07]. Carrier [Kar15b]. Carroll [Ram08]. Carrot [Mar07a]. Carrying [Bal05]. Cars [Joh14a]. Carter [DeR17, Gre01, Haf08b, Pre11b, Soa13, Yon08, BL10a, BL10b, Bur21, Gue17, Mil22, Njø02, Soa06]. Cartner [Mas23]. Cartographic [Che20]. Caryl [Kal14]. Casbah [Mor06]. Case [Col19, Dav11a, Dra06, Gil00, Goo05, Iat18, Kra22, Luk01, Lyn16, Ped19, Sch16, Tho08, Woh05, Zak18, Man21, Ped07, Gar00b]. Casey [Wea15]. Cash [Wir13]. Cass [Chu07b, Owe07, Rob14]. Cassella [Mar07a]. Cassella-Blackburn [Mar07a]. Castro [Att23, Pod14, Sco09, Tau02, Che07]. Casualty [Mar02]. Cataclysm [Son02]. Catastrophe [Chi04]. Catherine [McK15a]. Catholic [Szp10]. Catholicism [Leu13]. Caucasus [Dan12b]. Caught [Kom18, Kra04a, Lee12, Lak11]. Cauldron [Ass14]. Causes [Tur01]. Caution [Röh18]. Cautionary [Bur10]. Cave [Kor06]. CCP [Rad17, ZP23]. CDU [Sch05]. CDU/CSU [Sch05]. Ceausescu [Wei04a, Bot17]. Cecil [Kle23]. Cede [Rug16b]. Celia [Yor21]. Cenckiewicz [Szp09]. Center [Byr12, Car11, Cli15a, Dra12b, Hes12, Kan15, Kuk11, Pod14, Rei16]. Central [Edg08, Gat03, Gra07, Hut19, Joh06, Ken03, Kra99b, Kra99a, Kra99c, KCG^{+}17, Kus03, Lef03, Luň10, McC14, Odo06, Pas03, Pud12, Shl09, Sny03a, Soa06]. centrale [Gra07]. Centrifuge [Bur17, Hel22]. Century [Aki02, Dou06, Gor05, Har23, Her17, Lib07, Mil06, Pet11a, Rut10, Spr17, Zei08, AS22, Bor12a, Cea17, Dav11a, LX23, Roz11, Bur08, Cob18, Kal14, Rad14, Kag05, Sch02c]. Certain [Car17, Woh99]. Cervin [Soa23]. Cesari [SF11]. Ceskoslovenské [Gáš13b]. Ch [She12a, Wet11a, Wet11b]. Chad [Wet10]. Chair [Par19]. Challenge [Gho00, Goo10, McB15, O’D09, Tol03, Tor22, Wys18]. Challenges [Eng09, Hen09, Iri14, Ler15, Ser13]. Chambers [Whi10]. Chance [Bru08, Div00, Rad17]. Chang [Mat20a]. Change [Bro08, Dom09, Eng05, Har11, Kap16, Kra22, Mar05, Nie16b, PL22, Ric22, Sto07, Won04, Sco17, Mil02]. Changed [Bra14, Con23, Mas23, Mat08, Win15]. Changing [Che06, Dob10a, Joh07a, Leb06, Luk01, Pet20, Yar09, Ald07, Pra99, Van99b]. Channel [Lou16, RT13, Zub13]. Chaos [Dre21]. Chapel [Bre21, Dei07, Gar13, Hat11, Hay13, Hel07, Joh07b, Joh14c, Lak11, Lou16, Mar07b, Mar07d, Rab13, Rad16, Ram07, Sha07a, Str12, Wie07]. Chapman [Ass14, Pay21]. Charap [Ste18]. charismatiques [Mar16]. Charles [How22, Hus13, Mil15a, Ste15, Bru16, GP11, Kol15, Lee06a, Mor12]. Charter [Bug15]. Chase [Joh14b]. Chasing [Hay03]. Chast [Val14]. Chatterjee [Kur14]. Chatterji [Tal11]. Chatto [Bra16a]. Chaufour [Sze19]. Chautauqua [Eva07]. che [Rom17]. Cheddi [Rab13]. Chef [Kro12]. Chemical [Lei08, Wal12a]. Chen [Iri11]. Cheng [yL12, Mat20a, Tay19]. Cherny [Gra13a]. Chess [Ken13]. Cheung [Dre13]. Chi [Zha09]. Chicago [Lit11a, Rad12a]. Chief [Aft20]. Chiefs [Bro02a]. Child [Ber15, Kro12]. Children [Att23, Bal05, Lee21, McM12b, Rub10]. Chile [Ped19, QG12, Yor19b, Dra09, Gus10]. Chilean [HT19, US21]. Chimäre [Bru08]. Chimera [Bru08]. Chin [Eng07]. China [Cha04, Cha14, Cle04, Coh06, Coh08a, Coh08b, Dre21, Ell03, Etc23, Hus00, Kra01, Lea01, Lev07, yL01a, yL07, Lin11, Lüt12, Pen06, Rad16, Rea06, Rec17, Ros08, Roz11, Sch13, Sha11, Sza14, Wem21, Xia08c, Xia14, Zal04, Zha02, Bar18, Che06, DX18, Dre13, FTRP17, Gra17, Hof06, Jac16, Jon08, Kap16, KM22, yL01b, Li18, LX23, Lin09, Lin11, Lum04, Lüt08, Lüt14, Mar08c, Mei15, Pol07a, RLV^{+}10, Vám22, Wan17, Wem21, WLX^{+}08, Xia06, Xia08b, XL23, Zha10, Zha23a, Zha23b, Kru05, Law01, Won04]. China-India [Hof06]. Chinese [Dan10, Dit09, Lev04, Mat20a, Pai16, Pai17, Ros09, Zha15, ZP23, Ber14, Haj12, JM23, Kho10, Knü20, yL09, Rad07, RTYD19, She20, SX14, Zha06, ZP23, Dre05]. Chinese-British-American [Zha06]. Chips [Raj14]. Chodakiewicz [Jol07]. Choi [Kur14]. Choice [LSvdV04, Mid11, Rua04]. Choices [McD02, Mil04a]. Choosing [Ass07, Div00]. Chris [Cha14, Che08b, Lee12]. Christensen [Coh12]. Christian [Dav10, Gil23, Hen09, Hes12, Kal14, Rug16b, US21, Wet11a, Wet11b, Kan10, Neh11, O’S17, Ped19]. Christianity [Kir08]. Christiansen [McM16]. Christina [Zhu20, Cha07a, Jon07]. Christine [Dom15c, Sim16, Eng09]. Christof [Nue06]. Christoph [Mil17a, Law07]. Christophe [Hut20]. Christopher [Dav09, Edw16, Gre10, Her16, Hes12, Kal12, Mar07d, Pra08a, Spr07, Wal07]. Chruschtschow [Gri10]. Chruschtschows [Ado08]. Chull [Pai14]. Chun [Dit09]. Chung [Cli15b]. Church [Moe07, Joh13b]. Churches [Bur19]. Churchill [Dei15, Kir16, Bes07, Ram11b, Rob14]. Chushi [McG06]. CIA [Aft20, Bal15, Bel11, Bur13, Byr12, Cro00, Cul22, Gar01a, Imm16, Joh11b, Kas16, Lea01, Low06, McA14, Myl12, Pra08b, Ric10, Ros17, Ros10b, Sew23, Whi03a, Whi13, Whi03b, Dav04, Gas13a, Kna03, Lit11a, Lon11, McG06, Meh11b, Mor13]. Cienciala [Clo09]. Cihat [Chu07b]. Cinema [Fro23, Whi02a, Whi13, You13b, Wei12]. Cinematic [Sco17, Sha02]. Cinerama [Mar15]. Circle [McD06]. Ciscel [Mic21]. Citino [Kal18, Wir07b]. Citizen [Dra08]. Citizens [Bai18, yL09, Sch01]. cittadino [Dra08]. City [Clo06, Har12b, Kru07b]. Civic [Lub11]. Civil [Ass23, Bel02, Dan10, Hay13, Hor07, Iat05, Kar14b, Lev04, Mar13a, McM08, Mil13, Pay21, Rin12, Ros10a, Ros05, ZP23, Cur07, Gro02]. Civilian [Wee18]. Civilians [Mat04b]. Clandestine [Cul22, Eck07b, Kna02, Pra08b, Ris14]. Clap [Tay15]. Claremont [Fre07, Lub11]. Clarence [dL13, KH23].

Clashing [McA14]. Class [Ano08i, Göt08, ST17]. Classes [Jan08]. Classics [Duk11]. Classification [Hel22]. Classroom [Low04]. Claude [Dim08]. Claudio [Dra21]. Cleansing [Dou15, Gat03, Ken03, Kus03, Lef03, Sny99, Sny03a]. Cleanup [Nor05]. Clearance [Aft03]. Clever [Hay06]. Cloak [Low06]. Close [Hay16, Sim23]. Closer [Bor12a, DF07]. cloth [Boy11, Gun12, Kan15, Whi11b, Joh13a, VA12]. Clouds [Win15]. Cloutier [Hut20]. Club [Mar07a]. Clune [Usd16]. CMEA [De 20, Sto08]. CO [Gru07, Kru07a, Lin11, Sil11, Haa13, Nor15]. Co-op [Nor15]. Coast [Lea01, Her09]. Cocktail [Gun12]. Coder [Wir13]. Coercive [Coh12, JLBK17]. Cohen [Sen16, Iri14]. Colby [Low06]. Cold [Abe21, Ano08i, Aro06, Bak09, Bal10, Ban06, Bar04, Bar22, Bar07, Bel09, Ben21, Bis04, Boc06, Bod08, Bro04, Car13a, Car11, Cha08, Cha04, Cha14, Cha07b, Che08a, Che10, Chu07a, Cle04, Cog02a, Coh02, Coh08a, Coh08b, Con23, Cox02, Cri99, Cri02, Cro00, Dan16a, Dan10, Dan12a, Dan15, Dan16b, Dan22, Dar15, Dar16, DeR17, Dee07, Dei02, Dei14, Dei15, Del23, Diz10, Dob10a, Doc05, Dom15a, Dom20, Dom08, Dom09, Dom15b, Dra12b, Duk11, Dum17, Eck14, Eng20a, Eng20b, Eva23, Fel05, Fie03, Fin04, Fou14, Fri10, Gaf11, Gal09, Gal16b, Gar00a, Gar10b, Gas06, Gem18, Gen04, Gen17, Ger14, Gif05, Gil18, Gil06, Gle05, Gol07, Gol12, Goo15, Goo17, Gre08, Gro07, Guz13, Haa13, Haf08a, Haf08c, Ham15, Har08, Har12b, Hay03, Hay15b]. Cold [Hay15a, Hei16, Hen09, Her16, Her21, Her03a, Her07b, Hes12, Hit16, Hol21, Hut01, Iac22, Ing17, Jak16, Jer08, Jer16b, Jos20, Jos23, Kal22, Kan12a, Kan12b, Kan15, Kap04, Kap10, Key16, Kir08, Kir09, Kle04, Kle23, Kna02, Kre12b, Kre12a, Kru08, Kuk13, Kuk21, Kur05a, Lar22, Lat08, Lat14, Lat16a, Lau00, Leb03, Lee21, Lee06b, Lee23, Lef12, Ler15, Lev10a, Lev10b, Lew12, Lib01, tL16, Lov10, Low04, Luk17, Lun04, Mad02, Mad04, Mad07, Mad08a, Mag14b, Mal15, Man13, Mar18, Mar06a, Mar22, Mar08b, Mar12, Mar23b, Mas11b, Mas21a, Mas23, Mav18, McM14, McP19, Mey11, Mik12, Mil20, Mil22, Mil04a, Mit07, Moe03, Moo23, Nai22, Nau04, Nie12, Nie10b, Nin03, Nor15, Nue05, OB20, Osg04, Pai17, PW05, Par09, Pec08, Pee23, Pen06, Pet18]. Cold [Pfa02, Pit23, Poc03, Pol06, Por09, Pra14, Pra15a, Pra15b, Pra20, Pud12, Qua09, Rak15, Ram07, Ram01, Ram08, Ram11a, Rat21, Ree16, Rei16, Rey23, Ric15, Rie15, Rob10, Roh17, Ros10a, Ros00, Ros04, Rot21, Rub20, Sal06, Sap03, Sar05a, Sar16, Sar05b, Sar11, Sau12, Sch08, Sch03, Sch20, Sch17, Sco17, Sel02, Sen16, Sha11, Sha06, Sha07b, Sib11, Sin18, Sny11, Soa09, Soa13, Soa20a, Soa22, Spe01, Spe14, Spe15, Sti10, Str12, Str09, Tal99, Tay19, Tei11, Tho11, Tol10, Usd16, Val08, VA12, Vám10, Van99a, Vau15, Ver18, Vil16, Wal21b, Wei01, Wen09, Wet01, Whe06, Whi02a, Whi02b, Whi06, Whi08b, Whi11a, Whi12, Whi13, Whi14b, Win07, Wir13, Woo13, Yar09, Yor15, Yor19a, Yor20a, Yor20b, You12a, You22, Zhu16]. Cold [Zie21, vDO08, vD21, ARN16, Blu14, Bri13, Bro08, Buc17, Car99, Car17, Cha07a, Chu19, Cly13, Cos22, Cri04, Cri10, Cri20b, Cri20a, Cur15, Dan11, DFK^{+}20, Duk11, Dum14, Eng03, Eng05, Ess15, Eva01, Fal20, Fav18, Fed17, FAM16, FM16, FP21, For20, For05, Fox09, FF22, Fro23, Gal22, Gar04d, Geh22, Gel12b, Gho00, GK02, GSP^{+}19, Gol06b, Goo10, Göt08, GD99, Gue17, Hal16, Har19, Hat21, Hay00, HK20, Hel22, HF14, HGP^{+}23, His13, Hop03, HT19, Iat18, Ing04b, Jen21, JE19, Jer01, JMG08, JM23, Joh01, Jon07, Kac02, KG11, Kap17b, Kap20, Kar14b, Kel19, KSR18, Knü20, Kou17, Kra18, KNS18, Kra06f, Kra13b, KCG^{+}17, Kre07, Lei18, Li18, LX23, Lit10, Lou23, Lov22, Łuk23, Luk01]. Cold [Luk07, Luk16, Lüt16, Lyn16, Mac06, Mac20b, Mad13, MBR^{+}21, Mak17, Mak12, Man21, Mar02, Mar07c, Mar07d, Mar15, Mas02, Mat20b, May13, McA22, McD06, McG06, McK08, McM06, Mei15, Mic17, Mil15b, Mod17, Mue13, Mue16, Mur07, Nam13, Nut02b, O’M21, Odo06, Osg02, Par22, Pem16, Pip11, Pon01b, Pow16, QMLL16, Rad06, Raj14, Ray20, Rei19, Ria17, Ris14, RG15, Ros17, Rub20, Sac18, SS18a, Sar23, Sch16, Sch22a, Sha01, Sha02, Sha12, SY17, She20, Shi13, Sny13, Sur02, SJ23, TW05, Tho05, Tra09, Tro16a, TGS^{+}23, Ula99, Wal07, Wal13b, Wei12, Whi14a, Whi15, Woh99, Woh05, Wys18, Xia08b, XL23, Zak10, ZRS^{+}21, Zie08, Zim09, Żuk19, vW19, Alo05, Bel02, Ber15, Bra02a, Cul22, Din23a, Eva00, Fri02, Gal15, Gas00, Ger14, Gil23]. Cold [Haa05, Hei06, Hus00, Lef21, Mas00, Mul23, Nin03, Shm23, Van02a]. Cold-War [Ram11a]. Colhoun [Yor21]. Colin [Smi08, Har18, Kle16, Rab13]. Collaboration [Sha12, Spa14]. Collaborator [Ryd14]. Collapse [BL10a, BL10b, Cal03, Ged99, Gra07, KWKS07, KW07, Kni03, Kor07, Kra03a, Kra03g, Kra03f, Kra04b, Kra05a, Mat04a, Mul23, Rus16, Rus05, Sim17, Szp07a, Tay00, BSK^{+}23, Jen17]. Collapsed [Mou03]. Collateral [Zha12b]. Collection [Cri20b, Cri20a, Kar15c]. Collective [Elm04, McG99]. Collectivization [Maw08]. Colleen [TJ21]. College [Col15, Gaf11, McA07, Gem18]. Colman [Wal12b, Dum20]. Colonel [Mav18]. Colonial [Che08a, Dom08, Smi03]. Colonialism [Hei15]. Color [Ros04, Yor19a]. Colton [Rut11, Ste18]. Columbia [Ham15, Kre12b, Kre12a, Lat16a, Mad09, Rep12, Roz11, Sen16, Whi15, Zha12b]. Combat [Bur19, Mil21, Wal12a]. Combatants [Mat04b]. Comfortable [Boz15]. Comforts [Ber15]. Coming [Kna05, Son02, Zub14]. Comintern [Cha21, Nat04, Tra09]. Command [Dav09, Gal15, Her05, Odo06]. Commander [Gal22, Pfa02]. Commemoration [Oli11]. Comment [Hof00, Mil00, Pas10, She03]. Commentaries [BMLG11, BL10b, Gad13, Gre11, Lüt10, MS08, Nai12, NLM12, Rad12b, WLX^{+}08]. Commentary [Apt08, Ell12, Esh08, Gar10a, Gar10c, Gra12c, Har12a, Hol12, Kam13, Kap12, Kol12, Lev08, Mat10a, Ros12, Rub12, She12b, Szp12, Wal12c, Whi08a, Xia08a, Zha12a]. Commission [Pos07, Wal17a]. Commitment [Eck07a, Lar11, Lom09, Zim09]. Committee [Bra06, Bri13, Cor14, Hei16, Lar07a, Hei16, Joh13b, Rub06, SÃņ23, Bra04]. Common [Cob18, Her11a, Ket14]. Commonsense [Hay13]. Commonwealth [Ben21]. Communication [Nie16b]. Communing [Kir08]. Communism [Bru16, Cal02, Cos22, Dim08, Dou15, Fou14, Gor03, Hay15b, Hol08, Imm16, Jos20, KWKS07, KW07, Kle08, Kor07, Law07, Lew07, Mur05, Rob14, Rya10, Szp07a, Vár09, Whi11b, Ber14, Cho10, Dra04a, Dra04b, Dra10, Hay00, Kra03a, Kra04b, Kra05a, Kub07, Lyn16, Nei07, PA23, Pet11a, Rad11, Rya10, Szp07b, Tho05, Van04, Zin11, Bug15, Ell09, Joh20b, McM08, Whi10, Cra03, Hei06, Tol03]. Communist [And23, Ass23, Bor09, Buc17, Chi14, DS14, Etc23, Fri10, Ger13, Gil18, Jan08, Joh14b, Kap10, Ken07, Lev10b, Neu23, Rob10, Sch02a, Sha08a, Sil04, Sti10, Sto06, Vám10, vD12, Cor14, Del22, Gas19, Haj12, Iat05, Kli22, Lan01, Ler04, yL09, Łuk22, Mar13a, Mar08c, Mat20b, Nie21, Puc21, Rad07, Sfi01, Szp10, Tra09, Wil07, Fis03, Wei02]. Communist-Bloc [Ler04]. Communist-Era [Szp10]. Communist-Inspired [Wil07]. Communists [Dom20, Ran12, For20, Hai16, Njø02, Vu19, You23]. Communities [Bak09, Leu13]. Community [Cob18, Dei02, Doc08, Hur05, Kal16, Kaź09, Kaź09, Lie12, Hay17, Nor09]. Compagno [Dra08]. Companion [Bos15, Cli07, Hol15, Hol10, Law05, Stu15]. Companions [Bos15]. Comparative [Boo03, Gra17, Ket14, McM05, SY17]. Compared [Sai02]. Comparing [Bis12, Cap14]. Comparison [DS14]. Compensation [FM16]. Competition [Mar07c, Pif15, Pol07a, Sin18, Won04, Zhu16]. Complex [Ber01, Gho00, Nor07, Pra15a, Pra15b, Whi02b]. Complexity [Gol20, KNS18]. Complications [Dim08]. Composing [Spr17]. Compound [Elm04]. Comprehensive [Bur04]. Compromise [Par19]. Comrade [Dra08, ST17, Vár09]. Comrades [Joh14a, Ell09, Yor19a]. comunismo [Gor09, Bro11]. Conant [Her19, Kru07b, Her19]. Conboy [McA14]. Concentration [Ber07]. Concept [Bru08, Yor17b]. Concerns [Whe06]. Concise [McM10]. Condemned [Eng02]. Condor [Wee18]. Conduct [BMLG11, BM12]. Conducting [Bur04]. Conference [Dad19, SX09, Mid00, Röh18, Wan17]. Conferences [Eva07, Jos20, Rob07, Bar18, KNS18]. Confidential [Aft13]. Confirm [Ole11]. Conflict [Bru01, Che08a, Din04, Din14a, Din22, Fre07, Gan07, Her17, Her20a, Hof06, Kho14, Kna05, Lak11, Lüt06, Mas11b, Mel01, Mic17, Mil15b, Por09, Qua09, Rad16, Rai11, Rea11, She20, Ste02, Tof05, You12a, Dal10, Ger14, Kan12a, Gar04a]. Conflicting [BH17, Mei15, Dom03]. Conflicts [Nie10b]. conflitto [Mas11b]. Confront [Joc07, Mas00, Moe07]. Confrontation [Fri02, Gar02, Gar04a, Kli02, McK15b, Nam13, Zha23b, Zlo03]. Confronting [Lüt06, Dom20, Gar02]. Confronts [Meh11b, Soa06]. Congo [Kan12b, Kan15, Nam13, Nam16, Ian14, O’M21, You12a]. Congress [Joh05, Mat16b, Whi03a, Bar07, Joh01, Sci20]. Conley [Gal15]. Connecting [Hes12]. Connection [Hol99]. Conquest [Far15]. Consensus [All15, Cog02b, Sur02, Whi02a]. Consequences [Car11, Clo06, FP21, Sou21, Qua13]. Conservatism [Kle22, Rec17, Vei14b]. Conservative [Whi10]. Consolation [Oli11]. Consolidation [Ros06]. Conspirator [Fed17]. Constitution [DX18]. Constitutionalism [yL01b]. Constraints [Mat16a]. Constructing [Dav09, PA23]. Construction [Boc10, Duk10, Eng09, Lie08, Ste05]. Constructions [Nin03]. Constructivist [Adl03]. Consultation [Hat17]. Consumerism [Cal07a, Cal07b, Por08]. Consumption [Gun02, Rut12a]. Contact [Mas16]. Contained [Wal13a]. Containing [Hel22, Joh07b, Fre01]. Containment [Aro06, Bor99, Cho10, Jer06, McA22]. Contemporary [Har12b, Hur05, Ket14]. Contending [MBR^{+}21, MSPN21, Kri04, Zha02]. Contest [Kel19, Ste18]. Contested [Tum14, Pol05]. Contesting [Aro06]. Context [Con03, Sta16, Ken07]. Continental [Her11a, Her16]. Continues [Hay00]. Continuity [Cog02b, Mar05]. Continuum [Kur05b]. Contradiction [Bak09]. Contradictions [Del22, Zha02]. Contrarevolutia [Sni22]. contrats [Mar13b]. Contre [Lef02]. Contre-espionnage [Lef02]. Contribution [Szp09]. Contributors [Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00n, Ano00o, Ano00p, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01o, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11d, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano13d, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a]. Contributors [Ano15b, Ano15c, Ano15d, Ano16a, Ano16b, Ano16c, Ano16d, Ano17a, Ano18a, Ano18b, Ano18c, Ano18d, Ano19a, Ano19b, Ano19c, Ano19d, Ano20a, Ano20b, Ano20c, Ano20d, Ano21a, Ano21b, Ano21c, Ano21d, Ano22a]. Control [Bot17, Cra13, Dav09, McD02, Olš18, Sha08b, Wal12a, Aft20]. Controversial [Mad15]. Controversy [Kra18]. Convention [Car17, Den21]. Conventional [Car17, Usd09]. Conversation [Nor05]. Conversations [Pos07, Tis15]. Conversion [Kli22]. Convicts [Pip15]. Cooper [Bro22]. Cooperation [Dom15c, Gel12b, Has13, HF14, Mid00, Neh21, Olš18, Pec13a, Rea11, Ste02, Tho21]. Coordinating [Bae16, God11]. Coordination [Mar02]. Copeland [Dre15]. Copenhagen [Nie12]. Coping [Rai11]. Corber [Wei12]. Corina [Mac20b]. Corinna [Mac19]. Corkin [Win07]. Cornell [Ass14, Dav12, Die07, Geo13, Her12, Joh14a, Kal12, Kuz16, Lie12, Lit11b, Man13, McM11, Qua21, Ree16, Wan16, Whi13, Zhu16]. Cornerstone [Den21]. Corney [Nei08]. Corporate [Fav18, Man21]. Corporation [Gho00]. Corps [Cly13, Hof20]. Correspondence [Dei07, KLHK14, Pip03, Spe18]. Correspondent [Dan10, Dan22]. Cosgrove [Tho08]. Cosmopolitan [AE02, Aki02]. Costigliola [Sny17, Mis23a]. Costs [Moi03]. Cotton [Fer11]. Could [Cle00]. Couldn’t [Moy11]. Council [Bur19, Lar07a, McC20, Nei14, Dra12a]. Countdown [Mis23b]. Counter [Kau09, Knü20, Mor06, Sni22, Shl06]. Counter-Cultures [Shl06]. Counter-Revolution [Sni22]. Counter-Terrorism [Kau09, Mor06].

Countering [Imm16]. Counterinsurgency [Cra13, Kap17b, Lon11, Bur13, Rem08]. Counterintelligence [Fox09, Gro07, Hay03, Lef02, Ric10]. Counternarcotic [Pem16]. Counterrevolution [Sim16, Kos15]. Counterspy [Usd22, Gro07]. Countrie [Yor19b]. Countries [Nil19]. Country [Buc17, Kar15c, Vuj02, Pac06]. Countryside [Neu23]. Coup [Dun03, FF22, Gas13a, Get15, Gun15, Gus10, Hol05a, Mar23b, Mav18, Rad06]. Couple [Sha11]. Courage [Pra20]. Course [Con23, Eck19, Pra99, Van99b]. Court [Ale08, Bel09]. Court-Martial [Bel09]. Courting [Par16, Whi16]. Courts [Rya10]. Cousins [Pie16]. Covert [Cor14, Gar01a, Gut23, Joh07d, Kir08, Kna03, Lea01, Mar08a, McA14, Pac06, Ric07, Sew23, Yor23a]. Cowboys [Win07]. Cox [Sim14]. Craft [tL16, Ric10]. Craig [Mal15, Ped07, Rob14, Din14a]. Crash [Eck19]. Crawford [Kno06]. Create [Sny11]. Creating [yL12]. Creation [Gan16, Gar04a, Kis06]. Creative [Hay17]. Crescent [Joh13a]. Creswell [Hit03, Cog03, She03]. Creuzberger [Kur10, Nai08b, Sch10]. Crime [Clo09, Ree16]. Crimes [Bar08, Pai16, War01]. Cripps [Mar10]. Crises [Han17, Wes00, Ser13, Ler06]. crisi [Rom17]. Crisis [Åse16, Ash10, Bor05a, Cou04, Dan10, Dav04, De 20, Dei02, Gas06, Gre05b, Gut23, Haf08b, Hat17, Hou15, Ian14, Kur05b, Ler04, Mas12, Nue05, Par19, Pod14, Smi03, Ste18, Ste14, Sto08, Tag22, Val08, VA12, Wen04, Rom17, Col07, Col19, Cro07, Her04b, Her04c, Her20b, HSB^{+}23, Hol99, JKM12, Jon19, Koz14, Tie07, ZP23, Zha23b, Che08a, Gar09, Mas21b, Pra21, Bru01, Gre01, Rey02, Wel06]. Critchfield [Kis06]. Critical [BL10a, BL10b, Gar10c, Luc07, Mil02, Sch01]. Critics [Mor00a, Joc07]. Critique [Kur18, Nei14]. Croatia [Git20]. crollo [Gra07]. Cross [HF14, Ket14, SX14]. Cross-Border [SX14]. Cross-Cultural [Ket14]. Cross-Curtain [HF14]. Crossed [Von03]. Crouse [Joh14b]. Crown [Chi14]. Crucible [Dan22, Mil15b]. Crump [And20, Joh20a]. Crusade [Haa13, Sch02a]. Crusader [Low06]. Cry [She12c, She13]. Cryptology [Ste08]. Csaba [Vaj20]. CSCE [Dav11b, Bék16, Jar16, Nie16a]. CSU [Sch05]. CT [Fer07]. Cualann [Chu07a]. Cuba [Lou16, Rat21, Rom17, Sie06, Tra02, Yor21, Col19, Gle06a, Gle06b, Kel19, Lat16b, McK15b, Pow16, Yor23b, Att23]. Cuban [Dan10, Gre05b, Nam13, Pod14, Rom17, Smi07, Wel06, Yor21, Che07, Col07, Har19, Her04b, Her04c, Her20b, HSB^{+}23, Hol99, Koz14, Tie07]. Cubans [Dom15c]. Cullinane [SS18b]. Cult [Len12]. Cultural [BS02, Bri13, Cho10, Cul13, Din18, Fro23, Geh22, Hoc16, Ket14, Kle04, Kus02, Nin03, Pon13, RTYD19, Sci20, Sha07b, Shl09, Son02, Wem21, WLX^{+}08, Whi03a, Whi11a, Whi15, Guo10, Lüt08]. Culture [And05, Aro06, AS22, AE02, Bac09, Bak09, Bar22, Bel09, Bug15, Cea17, Cha07b, DF07, Gem18, GK02, Her99, Ing17, Kan10, Loe10, Nin03, PA23, Ram01, Ray20, Rog15, Sar05a, Sha01, Soa20a, Sol13, Sön22, Sza10a, Tha06, Tur01, Whi02b, Whi03a, Whi04, Whi12, Rut10, Whi14b, McE03, Ric04, Sha06]. Cultures [Kar15b, Lat16b, Whi14a, Del23, Shl06]. Cummings [Cri10, Haa13, Cul22]. Curb [Sar19]. Cure [Fie03, PR21]. Curious [Sha12]. Currency [Loe10]. Currents [SX09]. Curt [Pra07, Lef12]. Curtain [HF14, Lüt12, Nie12, Ros08, Shl09, Fro23, Mar15]. Curtis [Ric07]. Curtiss [Gho00]. Cutler [Gro07]. Cyberspeak [Cam05]. Cyprus [Cou04]. Czech [Bug15, Wet10, Wil17]. Czechoslovak [JKM12, Luk07, Šmi23, Svi16]. Czechoslovakia [Lar17, Mun10, Puc21, Sme15, Gar01b]. CZK [Gáš13b].

D [Alp09, Che08b, Cli07, Han13, How22, McM21b, Moe07, Mur07, vDO08]. D. [Dan07, Wal12a]. Daddis [Pre17, Pit23]. D’Agati [Vau15]. Dai [Sza18]. Daigle [Din14a]. Daily [Har23]. Daisy [Win15]. dal [Dra21]. Dalai [Dan14]. Dale [Dre15]. Dalsjö [Eng09]. Damage [Zha12b]. Damascus [Mar15]. Dame [Dom07]. Damian [Wet09]. Damnable [Bur21]. Dan [Ben21, Gal15]. Dana [Kro12]. Dance [Bac02]. Dancer [Kle04]. Danger [Win11]. Dangerous [Cri10, Dan12a, Ler04, Sza11]. Dangerously [Bra16b]. Daniel [Bos15, Bri15, Dar15, Dar16, Hus13, Bur16, Cra13, Kal16, Kir16, Kuk13, Ros10b, Yan08]. Daniela [Lov10, Pét13]. Danish [Mar18, Ole11]. dans [Lie08]. Daoud [Win17]. Daring [Sew23]. Dark [Neu11]. Darren [Dra09]. Dashed [Vit04]. Data [Jer22b]. Daugherty [Joh07d]. Daughter [yL01a]. Daulatzai [Joh13a]. Daum [Nue07]. Davenport [Sau12]. David [Foo21, Pon13, SS18b, Sew23, Soa09, Soa22, Sza14, Usd20, Wal21b, Wen09, Whi13, Ass10, Bar11b, Boc10, Bra14, Cro09, Dan12a, Dan16b, Dav09, Dei15, Ger13, Hol08, Kau09, Kuk11, Law12, Laz07, Mat20a, Nic12, Nor15, Par13, Pre11a, Ran06, Ros17, Sha08b, Str09, Tsi08]. David-Fox [Pon13]. Davidson [Bai18]. Davies [Cie07]. Davis [Jer13, Jer16a, Ros10a, Roz11]. Davoliūtė [Kot17]. Davy [Mas23]. Dawn [Kag05, Tra09, Val08]. Dawsey [Mol16]. Day [Gil23, Gol06a, Gol10, Kap17a, Kna05]. Days [Len18, Qua04, Rom17]. DC [Byr12, Car11, Din13, Dra12b, Hes12, Joh07c, Kan12b, Kan15, Pod14, Rei16, Smi11]. DDR [Rug17, Rug18, Wet09, Kan10, Bru06, Las19]. DDR-Außenministerium [Rug18]. De-Stalinization [Mëh11a]. Dead [Dan12a, Ert10a, Ert10b]. Deadly [Cla09]. Dealing [Alp09, Pec13a, Hut19, Nie03]. Dean [Sny08, Har08, Bra02b]. Death [Ber07, Git20, Gre08, Jer08, JMG08, Kes07, yL09, Mad08a, New07, ST17, SM12, Usd16, Ale12, Lic20, Nam16]. Debate [Cog02b, Hay00, Mar21, Moi03, Rug11]. Debates [McK08, Rai11, Haa05]. Debating [QG12, Gle05]. Decade [Eva23, Mei15, Wem21]. Decades [Din13, HSB^{+}23, Mas23]. December [Haj12, Sni22, Col07, Her03b]. decembrie [Sni22]. Decentering [Rei19]. Decides [Mad02]. Decimals [Kle16]. Decision [Bor05a, Hou15, Kal09, Mad15, Mad08b, Rea11, Wes00, Zha10, Cro07]. Decision-Making [Hou15, Kal09, Mad08b, Cro07]. Decisive [Lev04]. Declaration [Hat17]. Declassified [Bur13, Pra20]. Decline [Dra10, Van04]. Declining [Cea09]. Decolonisation [Cea09, Mur04, You12a]. Decolonization [Ces21, Del23, Gos08, Hes12, Mac06, Rei19, Yor18, dS19, Yor20b, TJ21].

Deconstructing [Onu14]. Dee [Lit11a]. Deep [Gal22]. Deery [Hay15b]. DEFCON [Hat17]. Defeat [Mad15, Nor08a, Rus07, Sch00a]. Defeated [Kar14b, Sai02]. Defects [Kle04]. Defence [Ald07, Ber01, Cla01, Dei02, Duk11, Lar07b, Goo15, Goo17]. Defence-Industry [Ber01]. Defenders [Joc07]. Defending [Bra02b, Bel09, Soa20a]. Defense [Bri13, Cog02b, Dre13, Gho00, Gil06, Gro02, Has13, Her11a, Her16, Iat05, Kaź09, Kis06, Kom14, Lüt08, Mei15, Rin12, Whi13, SÃņ23, Haf08a, Ros10a]. Defiant [Dan12b]. Defined [Moe18]. Defrosting [Mas23]. Defty [Sha08a]. Defying [Mëh11a]. Degrees [Soa23]. DeGroot [Nor08b]. Deimperialization [Iri11]. Dej [Fis03, ST17]. DeKalb [Ran12]. Dekolonisierung [Mas19b]. Del [Dav12, Bro11, Nut10, Nut11]. Delayed [Bur10, Moi03]. Delhi [Jak16]. della [For02, Mas11b]. dell’impero [Gra07]. dell’Urss [Gra07]. Delusion [Dal00, Szo10]. Delusions [Jer11]. Demand [Cal07a, Cal07b, Por08]. Demands [McM11]. Demise [BSK^{+}23, Bud22, Cra03, Ken07, Lan01, Sav09, Tel18, Tho05, Wal03, Yor15]. Democracies [Wir06]. Democracy [Ber04a, Boo03, Elm04, Her02, Ken13, Lit11a, Mav18, McC14, Neh11, PL22, Pee23, Pud12, Rub11, Sim14, Whi04, SÃņ23, Dim08, GB10, Hat10, Lic20, Bel02, Søn19]. Democratic [Bru04, Dal10, Dom11, Kli22, Pik02, Spr17, US21, Wet09, Ped19]. Demonstrating [Ste10]. Denazification [Wen09, Nai02]. Deng [Del22, FTRP17, Rad16, Zha10]. Denial [Mur05]. Denis [Zub14]. Denise [Che10]. Denmark [Mid11, Ole18]. Dennis [vD12]. Denuclearization [Kou12]. Denver [Sel21b, Sel19]. Deny [Ole11]. Department [All15, Her21, Kna03, Rid16, Sha08a]. Dependence [Sza14]. Depictions [Lor17]. Deportees [JK07]. Deporting [Pol10]. Derek [Jer11]. Desch [Aft13]. Desert [Rad12a, Wir07b]. Design [Gra21a, Sil99]. Designs [Hus05, Mil04b]. Despair [Kes07]. Destination [Bur23]. Destroy [Clo06, Sba05]. Destruction [Der02, Mou02, Sch00b, Sch01, Sha18]. d’État [Hol05a]. Detective [Lor17]. Detente [Geh05, Gen04, Nie03, Wen04, All15, BJ19, BL10b, Din14a, Din22, Lee23, Mar21, Mas08, Mas12, Nie08, Nie12, Pet20, Rei19, Sel21a, Ban16, BL10a, Mue11, Yor15, Fin05]. determination [Hez02]. Deterrence [Hen09, Kal12, Kno06, Kur18, QMLL16]. deutsch [Haf08b, Mue10, Rug16a]. deutsch-amerikanischen [Haf08b]. deutsch-russischen [Mue10]. deutsch-sowjetischen [Rug16a]. deutsche [Gri10, Rie08, Rug16a, vD03]. Deutschen [Nai08a, Rug17, Bis13]. deutscher [Bru06]. Deutschland [Ban14, Bru08, Moe03, vD02, Nai10b]. Deutschland-Politik [vD02]. Deutschland-Tagebuch [Nai10b]. Deutschlandpolitik [Gri10]. Developed [Sto08]. Development [Chu09, Cit14, Eas10, Han17, Her20a, Kap20, Kuz16, Mac19, Man13, McK08, PW05, Por09, SL23, Kal16, Laz07]. Devil [Mil01, Nie03, Pec13a]. Devils [Pai16]. Devine [Joh14c]. Dewey [Kle16]. Dexter [Ped07]. d’Hiv [Kar15b]. Dialogue [Kam17]. Diaries [Mar10]. Dictator [ST17]. Dictators [Alp09, Hut19, Whi04]. Dictatorship [Ble16, Kuz16, Mar06b, Nie21, Pap23, Ran10, Cal07a, Cal07b, Por08, Har23]. Dictatorships [Bru16, Cro09]. Dictionary [Pet11a]. Did [Fox09, Jer16b, Mat04a]. Didn’t [Ben05]. Died [Har18]. Diedrich [Wet11a, Wet11b]. Diem [Ass14, Spe13]. Dien [Nor08a]. Dienesch [Pra17]. Dieter [Zab16]. Difference [Jer16b, Kal12]. Different [Ste13]. Difficulty [Soa23]. Dilemma [Bur21, Jer01, Han13, Xia14]. Dilemmas [Adl03, Alp09, Dim08, Elm04, Kou17, Ole11, Sza14, Yan08]. Dillon [Onu14]. Dimension [Dra04b, Mad13, Mag14b, Man21]. Dimensions [Dra10, FP21]. Dimitrov [Pon01a]. Dina [Bit17]. Dinh [Ass14]. Diplomacy [Alp09, Bar18, Bor05b, Boy23, Bur17, Cha04, Che08b, Cog02a, Coh12, CKP05b, Cri04, Dav11a, Den21, Din04, Diz10, Fro23, Gar02, Geh22, GP11, Gut23, Han13, Her03b, JLBK17, Jos20, Kru07a, Kus02, Leu13, Mas00, Pai17, Pie16, Pon13, Qua14, RG15, Sau12, Sch17, Sny13, Spo12, Ste02, Szo10, Usd16, Wet01, Whi11a, Yor23a, Zub13, Cul13, Dei15, HTP^{+}15, Ric12, Whi15, Lun04, Myl15]. Diplomat [Mav18, Sto11, Taa13, Pfa02]. Diplomatic [Ass11, Bae16, Bro22, Den21, Fre07, Gif05, O’D09, Rob02b, Rob14, Tag22, Van11]. Diplomatie [Mar13b, Rug16a]. Diplomatiya [Boy23]. dir [Sze19]. Direction [Ald07, dS19]. Director [Low06]. Directorate [Whi03b]. Dirk [Fin13, Har12b]. Disarmament [Gly06]. Disaster [Mat10b]. Discontents [Con12]. Discontinuous [Kra22]. Discord [Haf08b]. Discourse [Bos04]. Disenchanted [Sim04]. Disinformation [Sel19, Sel21b]. Disintegration [Tay03]. Diskontinuitäten [Pen18]. Disorder [Wem21]. Dispersal [JK07]. Displaced [Bur23]. Displacement [Myl15]. Dispute [Che20, Qua23]. Disruption [Kli02]. Dissent [Bug15, Lor17, Mad08b, Whi11a]. Dissenting [Kra18]. Dissolution [Bék23, Kra22, Hal07a]. Dissonant [Koz14]. Distance [Sim10]. Distant [Gal16a, Kan12b, Roz03]. Distorted [Gel12b, Roz11]. District [Nor05]. Disturbances [Gor01]. diversiya [Boy23]. Divide [SS18a]. Divided [Ald08, Har12b, Hay16, Moe07, Ram01, Sav14]. Division [Mai05, Sme13]. Dizard [Kru07a]. Dmitar [Lee17]. DMZ [Rya08a]. do [Szp09]. Dobbs [Sco09]. Dober [Lee21]. Dobson [Ree16]. Dobzhansky [dL13]. Doctrinal [Kal12, Rob04]. Doctrine [Bur05, Bur10, Che10, Dom05, Fra09, Joh07b, Kom14, Ole11, Tho21, Win17]. Documentary [Kip15]. Documentation [HK09]. Documents [Cri20b, Rad07, Zub13, Kar15c, Kip15, Git07, Cri20a, Yon08, Jen03, Ran02]. Does [Bra16a]. dokumentacní [Gáš13b]. dokumentakh [Kar15c]. Dokumente [Rug16a]. dokumenty [Sny06]. Dollar [Dan16b, Low06]. dollaro [Nut10, Nut11]. Dollars [Hel07, Wid05]. Dolores [Ble16]. Domestic [All15, Din02, Dra10, FP21, Hay00, Jac17, Kus02, Ser01, Smi07, Uel20, Vei14b, You12b]. Domesticity [Sny13]. Dominance [Kim07, Wat17]. Dominik [Bru08]. Domino [Bor05a]. Domnitz [O’S17]. Donal [Pec13a]. Donald [Stu15, Con07, Hes13, Roz11]. Donfried [Cul13]. Dong [Wem21]. Donkey [Mar07a]. Donna [Mar07b]. Don’t [Tor22]. Doolan [Har18]. Doomed [Nun22]. Doomsday [Bur10]. Door [Fer15]. Doppelter [Wet17]. Dorado [Ale04]. Doran [Wat17]. Doslidzhennya [Sny06]. Double [Mun21]. Doubleday [Dan12a, Dan16b]. Doubt [Ped07]. Doucette [Joh20b]. Douglas [Edg08, Sha07a]. Down [Gra07]. Dr. [Her99]. Draft [DX18]. Dragon [Boy11, JM23]. Drake [Hol06]. Drama [Fal05]. Dramatic [Gal15]. Dream [Hoc16, Rut12a]. Dreamboat [Con23]. Dreams [Eva23]. Dreamworld [Chi04]. Dribble [Wal21b]. Dritte [War06]. Dritten [Bru08]. Driving [Ado08]. Drohpolitik [Ado08]. Dropping [Soa13]. Drove [She12c, She13]. Drugs [Cra13, Cra21, Pem16, Sch11]. Dry [Ste14]. DST [Lef02]. Dual [Rea11, Ros17]. Dual-Track [Rea11]. Duan [Sho15]. Dubonosov [Tsy08]. Duc [Law07]. Duccio [Nut10, Nut11]. Dueling [Luc02]. Duke [Gun12, Iri11, Kro12, Myl12, Ros17, Wei12]. Dulles [Joh11b, Pra14, Yaq02, Tud05, Che08b]. d’un [Lef02]. Duncan [Hay15a, Rub20]. Dunfermline [Chu07a]. Dunn [dL13]. Dunne [Gen17]. Durham [Duk13, Gun12, Iri11, Kro12, Myl12, Ros17, Wei12]. Duric [Lar07b]. During [GD99, Her03b, Lea01, Lib01, Mil01, Ros00, Bor05a, Boy23, Cal14, Car17, Che07, Chu19, Cob18, Coh08b, Col19, Cri04, Cro07, Dei13, Dom11, DFK^{+}20, Duk11, Dum14, Fal20, Fes07, FAM16, Gas19, Gav05, Geh05, Has07, HK09, HF14, Her21, Hou15, Ing04b, JM23, Kap04, Kau09, Kle04, Knü20, Kuk21, Lei18, Li18, Lie08, Lud13, Luk07, Lyn16, Mar23a, McM06, Mey11, Mod17, Nau04, Pec13a, Pem16, Pol05, Rei16, RG15, Sac18, Sha12, SX09, Sin18, Sny13, SM12, SJ23, Szo10, Tra11, Tra22, US21, Wal12a, Wir13, Wys18, Zak10, ZP23, Zhu16, Zim09]. DVD [Rut12b]. Dwight [Che08b, Cli15a]. Dworok [Mil17a]. Dylan [Goo15, Goo17]. Dynamics [Guo10, Kra13a]. Dzemal [Kar15a].

E. [Has07, Mar06b]. Each [Lüt14, Pas10]. Eagle [Dom15b]. Earl [Cha21, Mur05, Sib11]. Early [Che07, Cur23, Dum17, Goo10, Guz13, HTP^{+}15, Joh06, Jos20, Kra18, Kra99b, Kra99a, Kra99c, Kre07, Kuk13, Lef12, Lut22b, McM06, Nei14, Pem16, Pud12, Ria17, Sny13, Yar09, Zha06, Zhu16, Her21, Jos23, Mar22, OB20, Sib11, Ver18, You22]. Ears [JM23]. Earth [BH17]. East [Ado08, Ash10, Ber04b, Ble16, Bru01, Bru10, Cal02, Cal07a, Cal07b, Chi04, Eng20a, Gar04a, Gat03, Ged99, Hut19, Ing17, Joh06, Jon03, Ken13, Ken03, Ken07, Kli05, Kna05, Kop16, Kru07b, Kub20, Kus03, Lak11, Lef03, Lef21, Loe10, MC23, Nie03, Owe07, Oye14, Pai14, Pik02, Por08, Por09, Qua09, Qua21, Ram11a, Rua04, Sar05b, Sel02, Sha07a, Sha08a, Shl09, Sny03a, Ste02, Tho11, Wes00, Yor19a, Zie08, vD12, Bal15, Ban16, Blu15, Bru03, Bus14, Cit14, Con12, Cos22, Din10, Fri18, Geh05, Geh22, Hal16, Han17, Her14, JE19, Joh07b, Jon08, Kra99b, Kra99a, Kra99c, Kra03a, Kra04b, Kra05a, KCG^{+}17, Lor17, QMLL16, Rad12a, Rub02, Sas14a, Sel19, Sel21b, Soa09, Sto14, Zhi00, Par10, Wat17, Qua04, Sze19]. East-Central [Joh06, Shl09]. East-West [Ing17, Por09, Tho11]. Eastern [And20, Arg04, Cal02, Chi14, Cri20a, Cul22, Ger14, Gor03, Gor05, Jac23, Joh20a, Kuz19, Luc02, Łuk22, Mac20b, Nai10a, Whi14a, Zha02, Bor99, Cri20b, Nie16a, Olš18, Tra08]. eBook [Sco17]. EC [Ler06, Mor00b, Mor00c]. EC-121 [Ler06]. Eccentric [Dav12]. Echternkamp [Mac20a]. Ecocide [Bra14]. Ecology [Zie08]. Economic [Cal22, Dra12a, For05, Han13, Jer00, Lev07, Lou23, Mad13, McK15a, Per04, Rei16, SS18a, Sny09, Yor23b, Pen06, Dre15]. economica [Nut10, Nut11]. Economics [Bod08, Dav07]. Economies [Sak06]. Economists [Eas10]. Economy [Büt05, Eng20a, Eng20b, Ham18, Her23, Lef12, Lib09, Mad13, Mor00b, Mor00c, Cli15a]. ed [Bal08, Bis08, Bis11, Bis13, Bla03, Car11, Cha21, Chi14, Chu07a, Cli07, Cra07, Din07, Dre13, Duk10, Eng20a, Git07, Gol12, Haw07, Hei16, Hol15, Hol10, Iac22, JK07, KG11, Kol15, Kur11, Las19, Lew12, Lib07, Lie08, Liv07, Lüt12, Mar06b, Mas19a, Mas12, Mas16, Mas21b, McM21b, Mey11, Mil21, Nai08a, Nai10b, Par19, Pen18, Pét13, Pod14, Qua09, Ran06, Ros08, Rug16a, Rug17, Rug18, Sar16, Sav09, Sch08, Sch13, Ser13, Shl09, Sny06, Sny17, Soa20a, Spe18, Sza10a, Sza21, Woo13, Yan08, Yon08, Yor19a, Yor20a, Zab16]. ed. [Coh10, Nei10b, Ros06]. Eden [Dei07, Dei15, Pra21]. Edgar [The15, Whi13]. Edge [Luk16, Sme15]. Edgerton [Gru07]. Edinburgh [Wal12b]. edited [Bis14, Bos15, Bri15, Cli15b, Dap14, Dom15b, Ger14, Kar14a, Kur14, Ler15, McM14, Pai14, Qua14, Sim14, Spe15, Ste15, Stu15, Sul22, Tum14, Wea15]. Edition [Cra07, Vuj02]. Editor [Ano00j, Ano00k, Ano00l, Ano01l, Ano01m, Ano01n, Ano02n, Ano02o, Ano02p, Ano02q, Ano04o, Ano08d, Ano08e, Ano08f, Ano09e, Ano09f, Ano09g, Ano09h, Ano10e, Ano10f, Ano10g, Ano10h, Ano11e, Ano11f, Ano11g, Ano11h, Ano12e, Ano12f, Ano12g, Ano12h, Ano13e, Ano13f, Ano13g, Ano13h, Ano14e, Ano14f, Ano14g, Ano14h, Ano15e, Ano15f, Ano15g, Ano15h, Ano16e, Ano16f, Ano16g, Ano16h, Ano17b, Ano17c, Ano17d, Ano17e, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano19e, Ano19f, Ano19g, Ano19h, Ano20e, Ano20f, Ano20g, Ano20h, Ano21e, Ano21f, Ano21g, Ano21h, Ano22b, Ano22c, Ano22d, Ano22e, Ano23a, Ano23b, Ano23c, Ano23d, Kra99d, Kra99e, Kra99f, Kra03b, Kra03c, Kra03d, Kra03e, Kra04c, Kra04d, Kra04e, Kra05b, Kra05c, Kra05d, Kra05e]. Editor [Kra06b, Kra06c, Kra06d, Kra06e, Kra07a, Kra07b, Kra07c, Kra07d, Kra08a]. Editore [Bro11]. Edizioni [Nut11]. Edmond [Ket14]. eds [Alp09, Ass10, Bis10, Bit17, Bru08, Bru09, Cal09, Cap08, Che08a, Cli15a, Clo09, Cog12, Coh06, Cri20a, Cul13, Dar15, Dar16, Dav07, Dav10, Din14b, Dob10b, Duk13, Gaf07, Gol10, Hay11, Hen09, Hes12, His13, Huh20, Hus13, Ing17, Jak16, Jan08, Jer16b, Jos20, Kan12a, Kau02, Kho14, Kip15, Kul13, Lat08, Lov10, Mac20a, Mac20b, Mag10, Mag14b, Mas10a, McA07, McA08, McM16, Mil17a, Mil06, Moï10, Mol16, Mue10, Mur07, Nai08b, Nai10a, Nei10b, Nie12, Nie10b, Pet11a, Pet18, Pre11a, Qua13, Rai11, Ric12, Rut10, Rut06, Sas14b, SS18b, Sha07b, She07, Shl06, Sny17, Soa09, Spe14, Sta13, Sto07, Sza07, Sza12a, Van99a, Vil16, Wal08, Wet09, Wet11a, Wet11b, Whi14a, Woo12, Xia08c, Yor20b]. eds. [Del23, Gil23, Pap23]. Education [Cho10, Dro18, War08, Low04]. Edward [Bor12a, Gem18, Hat11, Mar08a, Cha08, Hal07a, Sha18]. Edwin [McM21a, BMLG11]. EEC [Mue11]. Effect [Cra03]. Effectiveness [Kar04]. Effects [Pos19]. Efficacy [Bla15, Mic11]. Effort [Gaw21]. Efforts [Din10, Neu03, Wan17]. Egypt [Fer08, Fer11, Vit04]. Ehrenburg [Rub02]. Eichengreen [Cli15a]. Eigensinn [War06]. Eight [Dan15]. Einaudi [Bro11, Mas11b]. Einheit [Rug16a, Rug18, Sch10, vD02]. Einstein [Tsi08]. Eisenhower [Bos04, Ces21, Che08b, Dei07, Dei15, Her16, Joh07b, Jon04, Kir16, Kru08, Lat08, Pra17, Wat18, All14, Bro02b, Col00, Han17, Her11a, Mar05, Tud05, Wan16, Doc04, Luc02, Mad02, Pra15a, Pra15b, Yor23a]. Eisenmann [War08]. Eisernen [Bis11]. Eisman [Smi11]. Ekbladh [Boc10]. Ekengren [Ing17]. Elaine [Spr17]. Election [HT19]. Elena [Dra12b]. Elisa [She14]. Elisabeth [Yor20b]. Elise [Rus16, Sny11]. Elite [Cea06, Van11, Xia06]. Elites [Eng05]. Elizabeth [Hay06, Yor15, Bel09]. Elke [Nai10b, Bis13]. Elliot [Lib07]. Ellison [Wal12a]. Ellsworth [Wie07]. Elsewhere [Bur23]. Elzbieta [Tis15]. emancipáció [Vaj20]. Emanuel [Sar11]. Emanuelle [Iac22]. Embargo [Dob10a, Pen06, Smi07]. Embassy [Gar01b, Mav18]. Embracing [Sch00a]. Emergence [Bri13, Car13b, Dre13, Etc23, Nie08, Por09, Rad07, Soa20b]. Emergency [Gal22, Ros10a]. Emigration [Git07, Pip15, Wet09]. É [?]igréMoll:2016:GBJ. Émigrés [Mik12, Tro16a]. Emily [Fes07]. Emma [Hur10]. Emmanuel [Nam16]. Emotion [Tof05]. Emotionen [Nue07]. Empathy [BL10a, Gar10a, Gar10c, Mat10a, Pas10]. Empire [Cea09, Duk13, Edg08, Gra07, Kau03, Łuk22, Mas21a, Mul23, Nau04, Par09, Rey23, Sny03b, Sol13, Zei08, vH09, Cra21, Duk02]. Empire-Building [vH09]. Empty [Par19]. Emre [Bar20]. Enbergs [Las19]. Encirclement [Kho10]. Encounter [Ket14, Lov10, Moi14, Rea06, SS18a, Sza10a]. Encountering [Kur14]. Encounters [Hol09, Lev04]. End [Ber20, Bro15, Con03, Eck07a, Eng05, Ess15, Eva01, Eva23, For05, Gaf11, Hop03, Jon14, Kac02, Kal22, Kan12a, KSR18, Kub07, Mas21a, McM22, Mis23b, Sar11, Sco17, Spe01, Sur02, TW05, Tho05, TGS^{+}23, Wei04a, Wet01, Woh05, Ale01]. Ended [Dan15]. Endgame [Haa05]. Ending [Nor15, ZRS^{+}21]. Endowment [Mar07b]. Ends [Ass10, Pre11a, Soa22]. Enduring [DF19, Rog15]. Endy [Mar07d, Wal07]. Enemies [Rus04b, Nie21, Hol05b, Nat04]. Enemy [Bot17, Coh08b, Fri18, Gal07, Kir08, Par13, Ria17, Van11, Whi02b, Whi04, Sae07]. Energiemacht [Kur21]. Energy [De 20, Han13]. Engagement [BJ11, Bro08, Hal16, Ric12]. Engaging [BJ10, Mas21a]. Engel [Car11]. Engelmann [Bru09]. Engerman [Laz07, Par13]. Engineering [Ble16, Kle08]. England [Bar11b]. Enigma [Dra22, Haf08a, Kur14, Van11, Her07a]. Enlargement [Gol20, Kap04, Ros06, Spo12]. Enlightenment [yL12]. Enormoz [Her09]. Enrollment [Fri04]. Entangled [Ros09]. Entanglement [Gre01]. Entanglements [Gil23]. Entering [ZRS^{+}21]. Enterprise [Kap20]. Entry [Shi13]. Entwicklung [Nai04, Nai08b]. Environment [Bra14, Kru05]. Environmental [Sch01, Ste02, Woo12]. Envisioning [Whi04, Kal18]. Envoys [Geh22]. Enyeart [Lic20]. Enyhülés [Vaj20]. Enzyklopädie [Bru06]. Epic [Smi11]. Epigenetics [Bal17]. Epilogue [Ray20]. Equilibrium [Sta16]. Era [Ass11, Bit17, Cli15b, Fal05, Gas19, Gho00, Her16, Hut01, QMLL16, Smi08, Szp10, Szp14, Tis15, Wei04a, Wir07a, Zak18, ZRS^{+}21, HK09, McM22, Sau12, Bac09, Gem18, Soa22, Wal21b, Cog02b, Gif05, Hay99, Nut02a, Tom04]. Erbe [Bis14]. Erdölzeitalter [Kur21]. Ereignisse [Mil17a]. Erez [Hus13]. Erfahrungen [Ban14, Bis11]. Erforschung [Mue10]. Eric [Mag10, Sza12a, Cog12, Joh14b]. Erika [Joc07]. Erinnerungskultur [Rie08]. Erkenntnisse [Ban14]. Erlebnisberichte [Mer13]. Ernest [Alp09]. Ernst [Ass10, Pre11a, Bis08]. Errata [Ano04p, Ano14i]. Erratum [Ano00m, Ano02r, Ano03k, Ano04q, Ano04r, Ano04s, Ano08i, Ano08g, Ano08h, Ano17f, Ano18j]. Escalation [Div00, Lin09, McM21a]. Eschen [Tei11]. Esparza [Dar15, Dar16]. Espionage [Dan16b, Gar00b, Hay99, Her09, Knü20, Sib11, Usd09, Nor15, Hay06, Mur05]. espionnage [Lef02]. Essay [Con12, Eck07b, Moi03, Neu03, Wil12]. Essays [Arg04, Kag05]. Essen [Mey11]. Essential [Ran06, Sul22]. Establishing [Din23b]. Establishment [Bae16, Bor09, yL01b]. Establishments [Kis06]. Esteem [Mil04a]. estera [Eva09]. état [Mar13b]. Ethan [And10]. Ethics [Roh17]. Ethnic [Dou15, Gat03, Gor05, Ken03, Kus03, Lef03, Pol10, Sny99, Sny03a, Sny07, Tum03, Zak18]. Ethnicity [Woo13]. Ethnicized [Bot17]. Etiudy [Sok99]. EU [Bis12]. Eugene [Cri20a, Roz11, Shm23]. Euphoria [Cur15]. EUR [Cra07, Mar13b, Mar16, Mil17a, Mol16, Nut11, Rom17, She12a, SF11, Vil16, Zab16]. Eurasia [Nai16, Ste18]. Euromissiles [Gaf14]. Europa [Huh20, Joh17, Joh21, Nai04, Nai08b]. Europe [And20, Byr12, Chi14, Cog02a, Cog12, Coh02, Cri02, Cri10, Cri20a, Dap14, Din23a, Din14b, Duk10, Eva00, Fer07, Gat03, Ged99, Gol11, Gor03, Gor05, Gra05, Haa13, Hei16, Hus05, Hut19, Ing17, Joh06, Joh20a, Joh21, Ken03, Ken07, Kus03, Kuz19, Lef03, Lie08, Lom09, Luc02, Lun04, Mag10, Mas11b, Mas12, Mey11, Mil04b, Nai10a, Nau04, Nie08, Nie10a, Por09, Ris15, Ros06, Rua04, Ser01, Sha07b, Shm23, Sny03a, Spe14, Spe15, Sta13, Sto07, Van99a, Wet01, Whi14b, Zha02, AS22, Bor99, Car99, Chi14, Cri20b, Gaf14, Imm16, Kra99b, Kra99a, Kra99c, KCG^{+}17, LSvdV04, Lie12, Lud13, Łuk22, Mai05, Mak12, MSPN21, Mor12, Nil19, Olš18, Pon01b, Pos19, Rea11, SS14, Sel21a, Sme13, Sny11, Tra00, Tra08, Ula99, Zim09, Bur23]. Europe [Cul22, Har23, Lai08, Lef21, Lie21, Mac20b, Mas10a, Mas23, Mil06, Ram08, Shl09, Wal17a, Wal08, Zei08, Arg04, Ber04a, Cal02, Fin04, Gor05, Loc05, Mid02]. Europe/Free [Hei16]. Europe/Radio [Din23a, Shm23]. European [Ber04b, Coh10, Dav10, Dei02, Doc08, Duk10, Fei10, His13, Kra03a, Kra04b, Kra05a, Lew12, Lie12, Luň10, Mac20a, Mat04b, Neh11, Nie12, Ser13, She14, Spe14, Spe15, Sza21, Tum14, Wal17a, Büt05, Gil00, Kaź09, Leu13, Lud13, Sar16, Sar19, US21]. Europeanization [Bis12]. Européenne [Lie08]. Eva [Kar14a]. Evaluating [BSK^{+}23, Mic11]. Evan [Mis23b, Soa13, Pol07a]. Evangelism [Etc23]. Everitt [Hol08]. Everyday [Gro02, Jan08, Kap17a, Mar12, Pét13]. Everyone [Ste18]. Everything [Nei10a]. Evidence [Ber13, CK11, Ler04, LSvdV04, Mic11, Per07]. Evil [Mas21a, McC20, Yan08]. Evolution [Cor14, Ell03, Gaf14, Haw07, Rin12, Rob20, Tel18, Wir07b, Zha23a, ŽČ16]. Ex [Mic21]. Ex-Soviet [Mic21]. Examination [Hou15, Sta15]. examining [Cha04]. Exchange [BS02, KW09, KLHK14, Myl15, QG12, Tol10, Bit17]. Exchanged [Kle23]. Exchanges [CK11]. Executing [Usd16]. Executive [Bur04, Pif15, Joh07d]. Exhibiting [Kus02]. Exhibition [Kus02, Mic11]. Exhibits [Tol10]. Exile [Jac16, Tol03]. Exiled [Pip15]. Exiles [Łuk23, Mik12, Wee18, Whi14b]. Exit [Pru04]. Exotica [Gun12]. Expedition [Vám22]. Experience [Dom07, Dre05, Kur14, PA23, Zak10, Ken05]. Experiencing [Duk10]. Experiment [Pon13, Mar06a]. Experimentation [Lee21]. Expertise [Kuk21]. Experts [yL07, Par13]. Explaining [CK13, Law10, Rus07, Sur02, Sny17]. Explanation [Adl03, Tan05]. Exploit [Mar02]. Exploitation [Sny09]. Exploiting [Mik12, Mil15a]. Exploring [She07]. Explosive [O’D09]. Explosives [Zie21]. Exports [Mei15, Sar19]. Exposing [Hei15]. Exposition [Mar15]. Expressive [Dob10a]. extérieurs [Mar13b]. Exterminate [Fei10]. External [Kra99a, Kra99c]. Extraordinary [Čav19, Pet08]. Extrenal [Kra99b]. Eye [Sza10b]. Eyeing [Pra17]. Eyes [JM23, The15, Wil23]. Eyewitness [Sto11]. Ezra [Cli15b]. Ezrahi [Zhu20].

F [Che19, Cli15b, Cro09, Gre05b, Gre05a, Her21, Her13, Hes12, Hol09, Hol15, Kam13, Kra13b, Kuk13, Kur13, Mag10, Mid02, Mur07, Nic12, Par16, Pra08a, Sen16, Sme13, Wal13a, Wal08, Wet11a, Wet11b, Whi16]. F. [The15]. Fabio [Dra18]. Face [Boy11, Bra16a, For19]. Faced [Wel06]. Faces [Jer22a]. Facilities [Nor05]. Factor [Boz20b, Has13, Svi16, Sza18, Tra11, Zub13]. Facts [Geo13]. Failed [BL10a, Gar10a, Hay03, Koc10, Mar20, Par09, Pas10, PS20, RT13, Ric09, Sim10, Tud05, Sch02a]. Fails [Lit11b]. Failure [Gra05, Gre05b, Gun15, Her04a, Kal12, Kap17b, Sha08a]. Failures [Fre07]. Fain [Joh11a]. Fainberg [Bit17]. Falk [Whi11a]. Fall [BSK^{+}23, Bru04, Che18, Dei02, Dom05, Eng09, Eng02, Hei06, Joh06, Joh11b, Kip03, Moi14, Par13, Pud12, Rad11, Spr07, Wei02, Wis00, Mar08b]. Fall-Out [Mar08b]. Fallen [Har18]. Fallout [Kra18, McE03]. Fame [Dan12b]. Family [Sfi02, Pra08b]. Famine [Lan20]. Famines [Gra17]. Fan [Sza14]. Far [Hor07, Kop16, Mar23b, Mun12, Tal08, Oye14, Zhi00]. Far-Right [Mar23b]. Farm [Ros10b]. Farms [Zie08]. Farrar [Gal16a]. Fascism [Lic20]. Fascist [Bra04, Rub06, Ryd14]. Fast [Gal09]. Fast-Attack [Gal09]. Fate [GB10, MSPN21, Ree16, Sol13]. Fates [Sen16]. Father [Ker15]. Fauser [Cal14]. FBI [Gar00b, Hay03, McM21b, Whi13, Goo05, Usd22]. FBI’s [HK20]. FDR [Moe18]. Fear [Gor05, Kap17a]. Feathers [Wee18]. February [Xia06, Col07]. Federal [Fre01, Rya10, Str09, Boz20a, Boz20b, FP21]. Federalizing [Mar07b]. Federation [Mat15, Pud12]. Federico [Mas11b]. Feel [Boz15]. Feet [Gar01a]. Feierstein [Dar15, Dar16]. Feigenbaum [Pol07a]. Felix [Kur21]. Fell [San16, Soa23, Moi15]. Femme [Wei12]. Fences [Ell03]. Fenemore [Ram11a]. Fenner [Mil14]. Ferguson [Hus13]. Fernando [Wee18]. Ferrell [Ham15]. Festivals [Fro23]. Fetz [Mol16]. Feudal [Del22]. Few [Che18, Lou23, Ula99]. Fiasco [Bar19]. Fiat [Cur23, Fav18]. Fictions [Sha08b]. Field [Kuz16, LX23, Mag14b, Alk05]. Fiend [Ess15]. Fifties [Ans17, Gro02]. Fifty [Ada03, Bro04]. Fifty-Year [Bro04]. Fight [Gar01b, Gif05, Lic20, Mat10b, Sur13]. Fighter [Die08]. Fighters [Kar14b]. Fighting [Gal09, Wal17b, How22]. Fights [Bel13]. figures [Mar16]. Files [Wir07a, McM21b]. Film [Eck14, Fro23, Rin12, Rut12b, Sha12, SY17, Whi14b]. Filming [Ros10b]. Films [Dum17, Ria17, Sha08b]. Filtzer [Con07, Hes13]. Final [Gre05b, Mar09, Pol10, Sny08]. finale [Dra18]. Finding [Bot17]. fine [Bro11]. Finest [Gra13a]. Fink [Sza21]. Finland [Ano08i, Rob14, ARN16, Göt08, JE19]. Finnemore [Die07]. Finnish [Ing04a]. Fire [Bac02, Gar01a, Dan07]. Firenze [Nut11]. Firm [WF12]. Firms [Won04]. Firsov [Cha21]. First [All14, Bor05a, Che08a, Fed17, Geh05, Guo10, Hor07, Iri02, Jon14, Kau09, KH23, Kle23, Kur05a, Lin09, Mat16b, McD02, Pra17, Rec17, Usd22, Wil08b, AS22, Mar15]. Fischer [Dav11b, Jak16]. Fischer-Tiné [Jak16]. Fisher [Ric12]. Fist [Kul10]. Fitzpatrick [Bra16b]. Five [Bar22, Wil23, Rya10]. Fix [Koc10]. Fixed [CS07]. Fixer [McK15b]. Fixing [Geo13]. Flame [BS02]. Flanks [Mas16]. Flash [Ler06]. Flashpoint [Wil07]. Flawed [Cal06, Joh11b]. Fleet [Pol07b]. Fleming [Mor13]. Flers [Bis12]. Flew [Pra20]. Flight [Wet09]. Flights [Con23, Pra20]. Florence [Rom17]. Flucht [Wet09]. Folgen [Kan14]. Food [Man10]. Foong [Luc07]. Football [Gem18]. Force [Ben05, Ber20, Cle00, Cro15, Gar04c, Hay17, Kro08, Lut22b, Rog15, Wys18, Leb06]. Forced [Dav07]. Forces [Gaf14, Nam13, Spe01, Szp07b, Tag22, Wil02, Wil08a, Rep03]. Ford [Lud13]. Fordham [Ket14]. Forecasts [Leb03]. Foreign [Ava04, Bea06, BJ11, Bis12, Bla15, Bur16, Cal06, Cal09, Cap14, Chu07b, Cli07, Cly01, CKP05a, Dan22, Dav12, Dob10b, Dom05, Dum20, Fel05, Gan06, Gar00a, Gar09, Gar04d, Haa20, Her04a, Hol05a, Iri14, Jen17, Jer11, JM23, Jon03, Kan10, Kau03, Lar02, Ler02, Lin11, tL16, Man13, McP19, Mil22, Nic12, Osg04, Pay21, Pem16, Rob14, Ros00, Ros05, Sar05a, Say17, Sch19, SS18b, She07, Sim14, Smi07, Sny17, Søn19, Ste10, Ste05, Tra11, Vit04, Wal12b, Yaq02, Yor17b, Zak18, Joh05, Yor15]. Foreshadowing [Gaf11]. Forests [Zie08]. Forever [Con23, Eck19, Nei10a, ST17]. Forgeries [Cor22]. Forget [Mil14]. Forging [Kur05b, Ber04a]. Forgotten [Myl12, Edw16, Spr07]. Form [Del23]. Formation [Dra12a, Gra07, Nie08, Wil23]. formazione [Gra07]. Former [Woo12, Lef02]. Forsaken [Ale09]. Forschung [Nai08a]. Fortnightly [Yon08]. Forty [Dan15]. Forty-Eight [Dan15]. Forum [BL10a, GRS^{+}12, Kra13b, RLV^{+}10, WLX^{+}08, vDO08]. Foster [Che08b, Pra14, Yaq02]. Fought [Pol06]. Foundation [Mue13]. Founded [Kle08]. Four [Han17]. Fourth [Wal21a]. Fox [Pon13]. Fractured [DeM12]. Frage [Gri10]. Fragile [Sta16]. Framed [The15]. Framework [Bae16]. Framing [HK20]. francais [Mar13b]. France [Dom20, Gra13b, Hit03, Kar15b, Lai08, Lan10, Mar07d, Mar16, Mar20, Mil04a, She14, SF11, Wal08, vC13, Boz15, Boz20a, Boz20b, CT03a, Lüt14, Mar08c, Pow16, Sci20, She03, Foo21, Gos08, Wal07, Cog02a]. Frances [Ran10]. Francesca [Del23, Van99a]. Francesco [Gun12]. Francis [Cro15, Ler15, Hel07, Pra07, Smi08]. Franco [Pay21, Spe13]. Frank [Chu07b, Kar14a, Mis23a, Owe07, Rob14, Akt10, Hay16, Mic17, Nam13, Rak15, Sny17]. Franklin [Eck19]. Franz [Rug16b]. Fred [Kle22]. fredda [Mas11b]. Frédéric [For19, Sar11, Say17]. Frederick [Nei08]. Fredrik [Che08a, Cal09]. Free [Byr12, Cri02, Din23a, Fes07, Haa13, Has07, Hei16, Joh21, Nin03, Pos19, Rot21, SS14, Shm23, Sto15, SÃņ23]. Freedman [Nei14]. Freedom [Cri02, Eck14, Edw16, Gar01b, Haa13, Joh13a, Kor06, McC14, Ros05, Whi03a, Sch20, Sci20, Rog15, Pee23, TJ21]. Freeman [Eva23]. FREEZE [Kal22]. Frei [Yor19b]. Freiburg [Mas16]. Freiheit [vD02]. Freire [Str12]. French [Cog02b, Dap14, Foo21, Kro12, Mur02, Nor08a, Smi03, Smi08, Hai16, Jon19, Mor00b, Mor00c, O’D09, Say17, Tra11, Van04]. Frequencies [Cul22]. Fridrikh [Cha21]. Friede [Kar14a]. Friedman [Sin18]. Friedrich [Rub06]. Friend [Ess15, Kli05]. Friendly [Vám22]. Friends [BJ19, Luk17]. Friis [Las19]. Front [Cor22, Dan10, Fes07, Kan12b, Kou12, You12b, JE19]. Frontier [Eck14]. Frontline [Kir09]. Fronts [Law11, Moï11, BMLG11]. FSB [Kar15c]. Fuchs [Goo05]. Fulbright [Gar00a]. Fulbrook [Bru10]. Fulda [Zab16]. Function [Kar15b]. Functioned [Mou03]. Fundamentalism [Wal17b]. Fundamentalist [Wal17b]. Funktionsweise [Jon06]. Fürst [Car13b]. Further [CKP05a]. Fusion [Sti06]. Futility [Eng07]. Future [Boc06, Joh14c, Kal18, Lee06b, Mor00a, Ric12, Leb03].

G [Chu09, Dan16a, DeM12, Din23a, Dra09, Gra13a, Lin11, tL16, McM08, New07, Ram07, She07, Spe14, Spe15, Sul22, Vil16]. G. [Mar08a]. G.Mahnken [Rep12]. GA [Bra14]. Gábor [Shl06]. Gabriel [Mar10]. Gaddis [Ber13, Kam13, Kra13b, Kur13, May13, Pec13b, Ste13]. Gaiduk [Lüt06, Sav14]. Galen [Qua23]. Galili [Pip15]. Gallagher [GP11]. Galway [Etc23]. Gamble [Tau02, Wat17]. Game [Bal15, Ken13, Sza11, Mun21]. Games [Jac17, Vau15, Rid16, Cla09]. Ganesan [Pai14]. Gangdrug [McG06]. Gangsterismo [Yor21]. Gannon [Gal09]. Gap [Gei12, Zab16, Pra08a]. garde [She03]. Gardner [McA07, McA08]. Gareth [Kim07]. Garnet [Lee12]. Garrard [Dom15b]. Garrard-Burnett [Dom15b]. Garrison [Lin11]. Garrisons [Duk02]. Gary [Pra07, Law10, New07, Nie10b, Pol06, WF12]. Gas [Bur17]. Gassert [Rai11]. Gatekeeper [Moe18]. Gates [Nat04]. Gathering [Mad13]. Gati [Ste15]. GATT [McK08]. Gaulle [Ass15, Kol15, Loc05, Mar16, Mar20, Kol12, LSvdV04, Lüt14, Mor00b, Mor00c, Mor12, Tra00, Tra12, Van00, Wal12c]. Gave [Bar11a]. Gavin [Hel07, Ler15]. Gawlikowski [Joh17]. Gays [Str09].

Gazibo [Bri15]. GDR [Har15, Kan10, Mad13, Kir08, Cal03]. Gedächtnis [Bis13]. gegen [Bru09, Lar07a]. Gegenwelten [Shl06]. Geisha [Sae07]. Gelfand [Nai10b]. Gemeinsamen [Mue10]. Gender [Åse16, Dum14, Edg08, Fel05, Ing04b, Pit23, Spe14, Spe15, Tod17, Rut10]. Gendered [Lyn16]. Gendering [Ria17]. Gene [Smi11]. General [Die08, GK02, Nut02b, Bos04]. Generals [Bes07, Mil17b]. Generation [Car13b, Gun12, Nei10a]. Generational [Rai11]. Genesis [Rea11]. Geneva [Bra02a, Van02a, Wan17, Ass07]. Genie [Bur17]. Genocide [Dar15, Dar16, Dou06]. Genocides [Ell12, Gra12c, Har12a, Hol12, Ros12, Rub12, Szp12, GRS^{+}12]. Genter [Whi12]. Gentry [Pra07, Mil14]. Geoffrey [Rie15, Pec08]. Geographic [Smi11]. Geometry [Tag22]. Geophysical [Vám22]. Geopolitics [Ver18, Rob14]. George [Hol10, Mit04, Her21, Her13, Kam13, Kra13b, Kur13, Moi14, Sme13, Wal13a]. Georges [Wal08, Van04]. Georges-Henri [Wal08]. Georgetown [Luk17]. Georgia [Bra14, Soa23]. Geppert [Bru08]. Geraint [Tho11]. Gerald [Sny08]. Gerard [Nam16, Nor08b]. Gerd [Nie10a, Sto07]. Gerd-Rainer [Nie10a, Sto07]. Gerhard [Ado08, Pen18, Por09, Rug16c]. German [Ado08, Bru01, Bru04, Bru08, Bru10, Bru16, Dal10, Dal00, Haf08b, Hit03, Kan10, Kop16, Lar02, Mac20a, Moe07, Pik02, Pru04, Rai11, Sar11, Sas14a, Sel02, Sel19, Sel21b, Sny09, Spr17, Ste10, Sza10a, Wen09, Wet09, Wil08a, Zie08, Bae16, Bék16, Boz15, Bus14, CT03a, Fin13, For05, Geh22, Gut23, Jar16, Kom18, Lor17, Nie16a, Rob04, Rug11, She03, Sto14, Ser01]. Germans [Dou15]. Germany [Ald08, Ble16, Bru08, Cal02, Fre01, Ged99, Haf08c, Kar15b, Kis06, Kop16, Kub20, MC23, Nai02, Nie03, Sar05b, Ste99, Sza12a, Wal08, Wen09, Yor19a, Zab16, Ban16, Bor12b, Boz20a, Boz20b, Bru03, Cal22, CK13, Fri18, Her14, Lut22b, McA22, Mue11, Ram11a, Rob02b, Uel20, vD12, Bak09, Cal07a, Cal07b, Loe10, Por08, vD21, Gas00, Lau00, Moe03, Ram01]. Gerrard [Rob14, Wir13]. Gerrit [Mil17a]. gesamtdeutsche [Sch10]. Geschichte [Bru06, Mue10]. Gesellschaften [Shl06]. Getting [Par22]. Getty [Kul10]. Gfeller [She14]. gg [Rob08]. Ghana [Tel18]. Gheorghe [ST17]. Gheorghiu [Fis03, ST17]. Gheorghiu-Dej [Fis03, ST17]. Ghor [Cit14]. Ghost [Bal17]. Ghostreaders [The15]. Ghosts [Cit14]. Giai [Zin11]. Gibson [Sas14b]. Giebel [Law07]. Gienow [Cul13]. Gienow-Hecht [Cul13]. Gifted [Dra22]. Gil [Sur13]. Gilbert [Lov10]. Giles [Fou14, Sha07b]. Gill [Rya08b]. Gillespie [Chu09]. Gillian [Mur07]. Gilman [Boc06]. Gilpatric [Bra06]. giorni [Rom17]. Girl [Hay06]. Giroux [Gal16a]. Gittinger [McA07, McA08]. Glacis [Boz20a, Boz20b]. Glajar [Mac20b]. Glamour [Gun02]. Glass [Mar08a]. Gleichschaltung [Nai08b, Nai04]. Global [Ber15, Car11, Con23, Die07, Etc23, Fin05, Gan16, Ger14, Ham18, HT19, Hus13, Jon14, KW22, KNS18, Lee23, Ler15, Mas23, McM16, PW05, Pem16, Ros04, SL23, Sar05b, Sto15, Sza21, Wie07, Woo13, ŽČ16, Hur05]. Globalism [Gav05, Coh08a]. Globalization [BD22, Eng20b, Lat08, Der02].

Globalizing [Eng20a, Kol12, Wal12c]. Glory [Gra13b, vC13]. go [Zha10]. Goals [Dob10a, Iat05, Mic11, Zhi00]. God [Dum17]. Godless [Pow01]. godov [Odo02]. gody [Odo02]. Goedde [McM14]. Goes [Ram11b, Whi13]. Göktepe [Chu07b]. Gold [Bar11a, Hel07, Wid05]. Goldstein [Cob18]. Goldwater [Win15]. Gong [Coh06, Xia08c]. Gontarczyk [Szp09]. Good [Eck19, Lee20, Nun22, Vár09]. Goodall [Pra08b]. Goodbye [Gra12b, Gra12a]. Goodman [Nor10]. Gorbachev [Gil05, Her02, Mas12, Roz03, Bro08, Con03, For05, RT13, Rut06, Sim23, Zlo03, Par09, Pra19, Ste99, Wei04b]. Gordon [Haw07, Kle11, Gar10b]. Gorham [And05]. Gori [Van99a]. Gorlizki [McD06]. Gorodetsky [Mar10]. Gorsuch [DF12]. Görtemaker [Nai08b]. Goscha [Hes12]. Goscilo [Rut10]. Gothenburg [Vil16]. Gottfried [Mas10a]. Gottlieb [Aft20]. Gough [Vár09]. Gouvernance [Lie08]. Governance [Gel12a, Lie08, SS18a]. Governing [Mic08]. Government [Gaw21, Goo10, Kru04a, Kul13, Mur04, Ped19, Sto14, Sza18, Tho11, US21, Str09]. governo [Nut10, Nut11]. Governor [Yon08]. Gower [Kre12b, Kre12a]. Grace [Cea06, Soa23]. Graf [Rug17]. Graham [Bal17]. Grain [Mor00a, Mor00b, Mor00c]. Grand [Bro02a, Dre21, Gre10, Lib09, Gel12a, Dal00, Hus05, Mil04b]. grandes [SF11]. Grandeur [Mor00a, Mor00b, Mor00c]. Grandin [Dom08, Nei10b]. grands [Mar13b]. granits [Min19]. Grant [Ham18]. Grassroots [Kal22]. Grazia [Zei08]. Great [Bal15, Boc10, Cro07, Dan12b, Ert10a, Ert10b, Gar10b, Hix05, Joh22, Kob06, Kra06f, Man10, Maw08, McC20, Pon13, Qua23, Ros09, Shl06, RTYD19]. Great-Power [Kob06, Kra06f]. Greatest [Bar08, Sew23]. Greece [Hat11, Sfi02, Iat18, Kou17, Sfi01, Mav18]. Greek [Myl15, Pap23, Dra22, Iat05, Kar14b, Mar13a, Kap14]. Green [Wil02]. Greenland [Pet11b]. Greenwood [Mil21]. Greg [Nei10b, Dom08, Whi15]. Gregg [Hat10, Luk17]. Gregory [Dav07, Lee21, Lib07, Pit23, Pre17, Sch17]. Greiner [Bis14, BMLG11, BM12]. Grenzfälle [Bis11]. Gretchen [Fer15]. Grimal [Cro15]. Grimstead [Mag10]. Gronbeck [Lat16b, Att23]. Gronbeck-Tedesco [Lat16b, Att23]. Grossbritannien [Las19]. Ground [Cly13]. Group [Mar02]. Grouping [Wil23]. Groups [Cor22]. Grow [Dom09]. Growth [Cli15a, Nor09, PW05, Ros17, SS18a]. Grünbacher [Wen09]. Gründung [Wet11a, Wet11b]. Grüner [Akt10]. Guan [Tay19]. Guang [Pen06, Rei16]. Guard [DeM12]. Guatemala [Sie06, FF22, Get15, Hol05a]. Guatemalan [Tod17]. guerra [Mas11b]. Guerre [CS07]. Guerrilla [Kra06f, Rem08]. Guerrillas [Joh22]. Guiana [Rab13, Ram07]. Guide [Mad09, Bur04]. Guided [Chu09]. Guidelines [Ano99i, Ano00n, Ano00o, Ano00p, Ano01o]. Guinea [Mac06]. Guinea-Bissau [Mac06]. Gulag [Ale12, Bar03a, Maw08, Pol07b, Ree16, Ber07, Kur11, Dav07, Hol05b]. Gulf [Fri05, Joh11a, McM21a, Mue05, Win11]. Gun [Den06]. Guns [Ben05, Eas10, Fer11, Mun21, Joh22]. Günter [Mol16]. Günther [Mol16]. Guoqiang [Wem21]. Gusterson [Nor07]. Guyana [Tay15]. Gwilym [Jer16b]. Gymnastics [Soa23]. György [Jan08, Shl09].

H [Ale07, Bur10, Cla09, Cli15a, Cri10, Cul22, Eck19, Haa13, Ham15, Joh14a, Kis06, Lie21, McM14, Mic21, Ros17, Sew23]. Haas [Usd22]. Habielski [Joh21]. Habsburg [Wan09]. Hacker [Sto11]. Haddad [Hur10]. Hadzisehović [Woo07]. Haftendorn [Wal08]. Hagopian [Lem14]. Hahn [Lak11]. Haiti [Sep01]. Hajimu [Mil15b]. Half [Hay16]. Half-Life [Hay16]. Hall [Gar10b]. Halt [Kal22]. Halvorson [Ben21]. Hamburg [Cra07]. Hamerow [Fei10]. Hammer [Eva07]. Hammerich [Lar07a]. Hammond [Ger14, Ger14]. Hand [Dan12a, Kau02, Tay15]. Handbook [McM14]. Handshakes [Jen21]. Haney [Smi07]. Hanford [Nor05, Nor05]. Hang [Gar13]. Hangmen [Kra06a]. Hanhimäki [Cal06]. Hank [Wal12a]. Hannfried [Ste10]. Hanoi [Ass07, Law05, Ass11, Gar13, You23, Vei14a]. Hans [Sha07b, Gol03]. Hansen [Ris15]. Happened [Ger13, Kra04a]. Harald [Jak16]. Harbor [Bar03b, Tie07, Sch09]. Hard [McK15b, Woh05]. hardcover [Gra12a, Lin11, Sco17, Gra12a, Onu14]. Hardesty [Smi11]. Hargittai [Bor12a]. Harmut [Nai10b]. Harold [Doc08, Vic08, Din13, Pra08b, Tho11, Win11]. Harper [Edw16, Spr07]. HarperCollins [Rad11]. Harrington [Kuk13]. Harrison [Ado08, Kap17a]. Harry [Bar11a, Bos15, Ped07, Sch13, Ham15]. Harvard [Abe21, Bal17, Bel11, Bru16, Cli15a, DeM12, Edw16, Fer15, Gan16, Gra12b, Gra12a, Hes21, Hor07, Hus13, Kal16, Kuk11, McM12b, Mil15b, Nam16, Pai16, Roh17, Sae07, Sol13, Spr07, Tei11, Zar21]. Harvest [Bea06, Kes07]. Harvey [Cha21, Gre20, Mur05, Sib11, Joh11b, Usd20]. Hasanli [Val08, VA12, Jac23, Min19]. Hasegawa [Gal07]. Haslam [Gal16a]. Hastings [Kor10]. Hatred [Gor05]. Hatzky [Dom15c]. Haunted [Hay99, Dre21]. Havana [Dom03, Kam17, Lou16]. Havel [Gáš13b, Tis15]. Haven [Bar11a, BSK^{+}23, Din14a, Ger13, Gra13b, GP11, Hay11, Kip15, Kur11, Lat14, Len12, Mal15, vC13]. Havoc [She12c, She13]. Hawai’i [yL12]. Hawk [Wie07]. Hawkins [Bak09, Dra09]. Hay [Ste14]. Haynes [Mur05, Sib11, Cha21]. Hazards [Hes13, Lei18]. head [Lef02]. Healing [Lem14]. Health [Hes13, Iac22]. Hearing [Aft03]. Hearings [Boe21]. Heart [Bal05, ST17, Wal17a]. Heartland [Fer15, Sny07]. Hearts [Osg02]. Heavens [Pow01, SZK12]. Heavy [SL23]. Hecht [Cul13]. Hee [Cli15b]. Heefner [Fer15]. Hegemony [Ros09, Wys12, Whi03a]. Heinemann [Jon06, Sza07, Wet11a, Wet11b]. Helen [Pra15a, Pra15b, Doc08]. Helena [Hei16, Rut10]. Helga [Wal08]. Hell [Tau02]. Helma [Kub20]. Helmut [Haf08b, Las19, Lar07a]. Help [Čav19, Eng09]. Helpful [McK15b]. Helsinki [Cra03, Dav10, Dav11b, Mas12, Mas23, Mar21, Mas10a, Sel21a]. Hemingway [Sto06]. Hemispheric [Get15]. Henig [Ros06]. Henri [Wal08, Van11]. Henrik [Ris15, Wet09, Vil16]. Henry [Cal06, Dar15, Dar16, Dav12, Kal22, Rub11, All15, Bur08, Dan14, Joh14c, PS20, Iri02]. Heonik [Lat16a, Oli11]. Her [Moo23]. Herbert [Moy11]. Heresthetics [Eva01]. Herf [Kop16, Din23b]. Heritage [Sny03b]. Herken [Luk17]. Herman [Che18, Jer16b]. Hermann [Clo06, dL13]. Hermanos [Sie06]. Heroes [Har18]. Herrera [Sak06, Whi13]. Hess [Law10]. Hessman [Rut12b]. Hester [Har15]. Heurtebize [For19]. Hey [Haf08a]. Hidden [Lou16, SX09]. High [Fri02, Gra21a, Her05, Hus00, Jan08, Kap16]. High-Seas [Fri02]. High-Yield [Gra21a]. Higher [War08]. Hijacked [Mat20a]. Hijacking [Joh07c]. Hilger [Rug16a, Mas19b]. Hill [Bre21, Dei07, Gar13, Hat11, Hay13, Hel07, Joh07b, Joh14c, Lak11, Lou16, Mar07b, Mar07d, Rab13, Rad16, Ram07, Sha07a, Str12, Wie07]. Hillman [Bel09]. Hills [Mun12]. Him [Buh15, Kra06a]. Himalayas [Hus00]. Himmelberger [Woo12]. Hindenburg [Ste10]. Hiroaki [Ert10a, Ert10b]. Hiroshima [Mad09, McM12a, vDO08]. Hiskey [KH23]. Hiss [Mar08a, Fet20, Kle16, KLHK14, Mar09]. Historian [Bar04]. Historians [Mur05]. Historic [Nor05]. Historical [Din10, Gar04c, Hay00, Kub07, Mad08b, Pik02, Sme15, Sur02, Tho11]. Histories [Dom15b, Eng03, Hes12, KNS18, Myl12, Tum14]. Historiography [Gar04d, Mar18, Rob02b, Sar23]. Historische [Rug16c, Rie08]. History [Aft13, Ale01, Arg04, Bal08, Bar08, Ber04a, BL10a, BL10b, Bos15, Bri13, Car13a, Col19, Cri10, Den06, Ell09, Gal16a, Gan16, Gar13, Gar10c, Gen04, Gly06, Gru07, Haa05, Hat11, Hei16, Hei06, Her17, Her07b, Her05, Iri14, Jen03, Jer22b, Kag05, Kip15, Kos15, Law07, Lee21, LX23, Liv07, Lou16, Low06, Mag14a, Mas02, Mas11b, Maw08, Nor05, Pra08b, Rad14, Rem08, Rog15, Ros09, Sch03, She07, Sil05, Sny17, Ste08, Szp07b, Tho11, Tum16, Vuj02, Wil17, Wir07a, Xia08b, Kar15c, Coh10, Hoc16, Jon14, Luc07, Mas11a, Tum16, Coh06, Hol21, Lüt22a, Mac19, McM10, Nai22, Tay19, Win07, Xia08c, Xia23, Bal05, Kel02, Ler02, Ric04]. Hitchcock [Wat18]. Hitler [Bru10, Fei10, Spa14]. Hoberek [Hol15]. Hockenos [Moe07]. Hodgson [Wir13]. Hoffman [Dan12a, Dan16b, Ran06]. Hofmann [Nie08]. Hogan [She07, Sny17]. Hogg [Cea17]. Höhn [Duk13]. Holidays [Mar07d, Wal07]. Holland [Pos07]. Hollander [Eck07a]. Holly [Mas23]. Hollywood [Bre21, Sny08, Spa14, Wei12, Whi13, Gun02, Sco17, Whi08b, Whi13, Whi14b]. Holmes [Gel12a]. Holmgren [Kur14]. Holocaust [Kar15b, Pol05, vH09, Bor12b, Kar15b]. Holy [Bra05]. Holzman [Ric10]. Home [Chu19, DF12, Fes07, Gro02, Kru08, Mil15a, Mil14]. Honecker [Ban16, Bru10]. Honeymoon [SX09]. Hong [Cho10, Oye14]. Hongwei [Sza14]. Honolulu [yL12, Dad19]. Honor [Gru07]. Hoogewoud [Mag10]. Hoover [Kap17a, The15, Whi13]. Hope [Bak09, Ado08]. Hopes [Con12, Nag17, Nag18, Vit04]. Hopf [Guz13]. Hopkins [Mur07]. Horn [Nie10a, Sto07, Yor15, DFK^{+}20]. Hornblum [Bar11a, Lee21]. Horne [Sny08]. Hornsby [Tro16b]. Horror [Bur05]. Horst [Rug18, Mue10]. Hosni [Yos15]. Hostage [Hou15]. Hostages [O’M21]. Hot [Dan16a]. Hour [Her19, Gra13a]. Hours [Dan15]. House [Car13a, Mad08b, Pos07, Suh08]. Housing [Ran12]. Howard [Coh10, Sny08, Wie07]. Hoxha [Neu23]. Hsing [Iri11]. Hsu [Lee20]. Hua [Lev07]. Hua-yu [Lev07]. Huber [Kar14a]. Hubert [Kle22]. Huebner [Bac09, Soa20b]. Hugh [Bal15, Bel11, Nor07].

Hughes [Jer16b, Pud12, Tho11, Eck14]. Hugo [Dou15]. hui [Rub11]. Human [Boe21, Bri13, Clo06, Cra03, Dra09, Haw07, KW22, Mar08b, Mar21, Mas23, Nei14, Pee23, Sch19, Soa06, Tho05, Jen17, Luc07, Nei10b]. Humanitarian [Mid11]. Humanities [Mue13]. Humphrey [Fed17]. Huneeus [Sil11]. Hungarian [Bal05, Bor05a, Nag18, Tol03, Her03b, Nag17, PA23, Szo10, Vám22]. Hungary [Arg04, Bor09, Hut19, Ken13, Bal10, Bék16, Bék23, Bor09, Bor12b, Man21, Mat12, Jan08, Vár09]. Hunner [Nor09]. Hunt [Law12, Nei10b]. Hunter [Gar00b]. Hunting [Eth05]. Hurley [GP11]. Husnija [Kar15a]. Hussein [Ash10]. Huttenbach [Dar15, Dar16]. Huw [Goo15, Goo17]. Hyam [Cea09]. Hydrogen [Her20a]. Hygiene [Hes13].

I. [Hol09]. IAEA [Röh18]. Ian [Kho14, Mor13]. Iasi [Del07]. Ibidem [Tum16]. Ice [Wei01]. Icelandic [Ing04b]. Icon [Rob20]. Iconoclast [Mol16]. Idea [Lai08, Woh99]. Ideals [yL12, Sap03]. Ideas [Ben05, CK13, For05, Jan08, Pra99, Tan05, TW05, Tho05, Van99b, Woh05, Kol15]. ideelle [Moe03]. Identität [Mas19b]. Identities [Ste05]. Identity [ARN16, Eng09, Eng05, Git07, Hay17, Mic21, Rut10, She14, Spe13, Wei12, Wil23, Rub11]. Ideologe [Kur10]. Ideological [Fav18, Xia14, Zak10]. Ideologies [Haa20]. Ideology [Bit17, Bos04, Del23, GD99, McC14, Nut02a, Ram08, Sil04, Soa06]. igrzysk [Kaź19]. ihre [Kan14, Wil08a]. II [Bur04, Dei14, Fal05, Gro07, JK07, Koc10, Rog15, Sny03b, Bor12b, BS02, Cal14, Fes07, Has07, McB15, Pie16, Rob04, Wol13, Bal08, Bre21, Clo06, Con07, Ham18, Mis23b, Sha08b, Bro02a, Gav05, Sch00a]. III [Hat17, Kal22]. Ike [Bro02b, Wat17]. Il-Sung [Del22]. Ilic [Kul13]. Ilko [Bru09]. Ilko-Sascha [Bru09]. Illegal [Tsy08]. Illinois [Ran12]. Illusions [Eva00, Gre10]. Ilya [Sav14, Lüt06, Rub02]. im [Akt10, Bis13, Huh20, Kan14, Mas16, Nue07, O’S17, Rug16c, Rug17, Sch08]. Image [Bac09, Bel02, Kaź09, Sar05a]. Images [Hat09, Hat21, Lee12, Sch08]. Imaginary [Jen21]. Imagination [Cha07a, Hal07b, Jon07]. Imagined [Law07, Sak06]. Imagining [Pem16, Sch02b, Sil05]. Imaginings [Tum16]. Imbalance [Ian14, Kim07]. Immerman [McM14]. Immerwahr [Kal16]. Immigrants [Lee20]. Immigration [Ver18]. Immunizing [Ing04b]. Impact [AE02, Cri20b, Cri20a, Dom07, Dun03, Jar16, Mun10, Que02, Żuk19, Rya08a]. Imperfect [Qua21]. Imperial [Zhu16, Fel05, Ste99]. Imperiale [Mas19b]. Imperialism [Che99, Nin03, Nun22]. Implications [Cri04, TGS^{+}23]. Importanza [For02]. Impossible [FF22]. Imprisoned [Mil14]. Imprisoning [Nie21]. Improbable [Zim09]. Inaction [Bor99, McG99]. Inauguration [Hei16]. Incentives [For05]. inchiesta [Dra18]. Incident [Her04a, Sti06, Den06, Pra07]. Including [Fet20]. Incorporated [TJ21]. Incredible [Pra20]. Independence [Gif05, Mac06, Rea06, Ros09, Sob19, Sto11, TJ21, Rab13, How22, Yor20b]. Indes [Mar16]. Index [Kle16]. India [Gan06, Tal11, Che06, Hof06, Mas10b, Que02]. Indian [Che20]. Indiana [Luc07]. Indies [Jon03]. Indies/Indonesia [Jon03]. indische [Mas19b]. Indochina [Jon14, Lüt06, Pra21, Yor23a, Vu19, Wan17]. Indochinese [Ass23]. Indonesia [Gar01a, Gar04a, Jon03, Jon04, Bod08]. Indonesian [Eas10, Gar02, Jon03]. Industrial [Gho00, Pra15a, Pra15b, Usd09]. Industrial-Scale [Usd09]. Industries [HP17]. Industry [Ber01, Cur23, Fal05, Her23]. Inernal [Kra99a]. Inernal-External [Kra99a]. Inevitable [Mel01]. INF [Adl03]. Infamy [Gru07]. Influence [CS07, Gre01, Kou17, Kur18, Wal13a]. Influences [BS02]. Influential [Bor12a]. Informal [Geh22]. Informants [SM12]. Information [Bro10, Kel19, Kru07a, Pol07a, Sha08a, Sti06, DS14, Kle16]. Ingo [Hit16]. Initial [Fet20]. Initiative [Hel22, Lar07b]. Innocent [Nei07]. Innovation [Ava04, Rob04]. Innsbruck [Mol16, Rug16b]. Inquisition [Bra04]. ins [Kur21]. Insecurity [Ber15]. Insider [Her07b, Sza10b, Wet01]. Inspiration [Pos19]. Inspired [Wil07]. Instance [Sha12]. Institute [Dav11b, Gal16b, Kas16, Pol13, Rak15, Ram11b, Rub11, Win11]. Institution [Din13, Kap17a]. Institutional [McK08, Won04]. Institutions [Lar02, GB10, Pét13, Sil99]. instrument [Kaź19]. Instrumental [Dob10a]. Insurrection [Law12]. Integration [Büt05, Gil00, Lew12, Sar16, Ser13, Ser01, Uel20]. Intellectual [Fin04, Gol03, Kro08, Lun04, Whi02b]. Intellectuals [Buc17, Eck07a, Eng08, Fin13, Gil06, Laz07, Tol03]. Intellectuels [Mar16]. Intelligence [Aft13, Ald08, Alp09, Ava04, Gar04d, Goo15, Goo17, Gun15, Gut23, Haf08c, HK09, Hol09, JM23, Kel02, Kis06, Kna02, Lit11b, Mad13, Mas09, Mat04a, Nor10, Pec13a, Pud12, Ric09, Ric15, Ris15, Rus05, SM12, Tro16a, Wil23, Żuk19, Knü20, Gal16a, Geo13, Jer16b, Low06]. Intelligence-Gathering [Mad13]. Intelligentsia [McM12b, Rub10]. Interaction [Eng20a, Li18]. Interactions [Ing17, Mag14b]. Interdependence [Dre15, Jer00, Mid00, Ban06, Doc05]. Interdipendenza [Nut10, Nut11]. Interdoc [Fou14]. Interest [Pud12, Raj14]. Interests [Fav18, Kou17, Mei15]. Interfering [Njø02]. Intergovernmentalist [Gil00]. Interloper [Gre20]. Internal [Kra99b, Kra99c, yL09]. Internal-External [Kra99c]. International [Alk05, Alo05, And20, Ber20, Bla15, Blu14, Boe21, Bor05a, Coh06, Cra03, Cro15, Dav11b, Dei13, Die01, Die07, Din23b, Dob10b, Dom07, Dra10, Elm01, Eva13, Eva23, FP21, Fro23, Gar13, GD99, Hat21, Haw07, Hel07, HF14, Hop03, Hur05, Hur10, HT19, Joh05, Joh13a, Joh20a, Kap20, Kuk11, Luň10, Lüt14, Lüt22a, Mar13a, Mar15, Mat10a, McM05, McM10, Mid11, Olš18, Onu14, QMLL16, Ric04, SL23, Sas14b, Sny03b, Sob19, Ste05, Sto08, Tra22, Wel00, Wid05, Woh99, Xia08b, Xia08c, Xia23, Yor17a, Rom17, Cos22, Dra09, Mat15, Mac19, McK15a]. Internationale [Fou14]. Internationalism [Gil18, Ped19, Sch02b, Sob19, Nic12]. internazionali [Rom17]. Interpretation [Guo10, Zin11]. Interpretations [Ber13, Law22]. Interpreting [Nei00]. Interpretive [Tay19]. Interrogation [Stu19, vD21]. Intersections [AS22]. Intervention [Ash10, Blu15, Fer08, Gun15, Kal09, Lan10, Leb06, Mil13, Ram07, Sob19, Yor15]. Interventionen [Kan14].

Interventionism [Eng02, Pee23]. Interventions [Dom09]. Interviews [Mit04]. Intimate [Low08]. Intra [Hel22, Rea11]. Intra-Alliance [Rea11]. Intra-NATO [Hel22]. Intrigue [Kle16]. Introduction [FAM16, GK02, KNS18, Kra03g, Kra03f, Kra06f, Kra08b, Mas23, Nie16a, Nut02b, TW05, XL23, Say17]. Inturist [Zak18]. Invaded [Tur01]. Invasion [Dal00, Mun10]. Inventing [Kru07a, Nor09, Pre11b]. Invention [Bra14, Gos08, Par10]. Investigation [Gar00b, Gra21a]. Investment [Sto08]. Invisible [Bar11a, Ber04b]. Involvement [Dra12a, Gus10, Kob06, QG12, Yor19b]. Involving [Lei18]. Ioanid [Pet08]. ioho [Sny06]. Iran [Key16, Sas14b, Gas13a, Gas19, Gav99, Joh07a, Mar05, Nun22, SJ23, Gas06]. Iranian [Lit11b, Val08, Fri18, Joh07a]. Iraq [Lit11b, Sas14b, Bel13, Sas14a, SJ23, Cap14, DS14]. Iraqi [Mat15]. Ireland [Whe06]. Irina [Rad14]. Irish [Bis12]. Iron [Kul10, Nie12, Fro23]. Irony [Ban06, Doc05, Lum04]. Irregular [Kro08]. Irresistible [Zei08]. Irresolute [Wes00]. IRS [Sba05]. Isaevych [Sny06]. Islam [Gue17, Kli22, Wei04b]. Islamic [Nun22]. Island [Jac16]. Isolate [Sar05b, Wan17]. Isolation [You22]. Isotopes [Her23]. Israel [Kli05, Kop16, Ste10, Wei04a, Din18, Her14, Pol13, Pet08, Din23b]. Israeli [Din04, Din14a, Din18, Fre07, Lak11, Qua13, Qua23, Din22, Git20, Gut23, Hug08, Kob06, She20, Kli02, Wei04a]. Israelyan [Gol07]. Issue [CT03b, GK02, KNS18, Kra03g, Kra03f]. Issues [Ber13, Kre12b, Kre12a, Mag10]. Istanbul [Pem16]. istorii [Mue10]. Istoriya [Kar15c]. István [Bor12a]. Italia [For02]. Italian [Her12, Dra04a, Dra04b, Dra10, Fal20, Hai16, Mar23b, Pif15]. italiana [Dra21, Eva09]. Italy [Bro10, Dom15a, Dra12b, Hol06, Mil04a, Bro02b, Cur23, For20, Gun02, Njø02, Nut02b, Nut02c, Sci20, Dom20]. Ithaca [Ass14, Dav12, Die07, Geo13, Her12, Joh14a, Kal12, Kuz16, Lie12, Lit11b, Man13, McM11, Qua21, Ree16, Wan16, Whi13, Zhu16]. Itself [Ano08i, Göt08]. ITT [Man21]. iunie [Del07]. Ivan [Lit11a, Lee17]. ix [Dou15, Liv07, Nor15]. izucheniyu [Mue10].

J [Aft03, Bac09, BMLG11, Bal08, Bur16, Che19, Coh12, Cur07, Dei13, Gol03, Her16, Her12, Her08, Hus13, Joh07d, Kal18, Law08, Lee21, Ler15, Mad08b, Mag10, Mar23a, Moi14, Nai16, Nor08b, Pol06, Pol07b, Qua09, Ros10b, Rut11, Sch20, She07, Sim14, Smi07, Sny09, Sny17, Ste18, Str12, Sza12a, The15, Wei12, Whi11a, Yor23a, dL13]. J. [Kul10, Rob20, Ste08, Whi13, Gar00a]. ja [Mas19a]. Jacco [Dap14]. Jack [Yor21, Żuk19]. Jackson [Sau12, Mar21, Qua23]. Jacob [Ket14]. Jacobs [Boy11]. Jaeger [Mac20a]. Jagan [Rab13]. Jahre [Pen18, Kur21]. jako [Kaź19]. jalanjäljet [Mas19a]. James [Bro22, Mas21b, Moi15, Rem08, Ryd14, Sha08a, Stu15, Yor18, Yor20b, Ale07, Ans17, Coh08a, Hal07b, Hat11, Joh13b, Jon14, Kis06, Koc10, Pre11b, Ric10, Sch13, Sza10b]. Jamie [Mue17, You13b]. Jamil [Jac23, Min19, Val08, VA12]. Jamison [Chu07a]. Jan [Jol07, Shl06, Che18, Len12]. Jane [Kor06]. János [Vár09, Jan08, Kro08]. January [Her03b, Yon08, Ber20, Col07, Xia06]. Japan [Cap14, Kho14, Kor10, Mey11, Sch00a, Sch13, Dob10b, Ess15, Gal07, KM22, Kom14, Mad15, OB20, SS18a, Shi13, Wol13, Zub13]. Japanese [Pai16, Roz03, Sae07, Has13, KM22, Kur18]. Jarmo [Mas19a]. Jason [Mol16]. Jay [Sol13, McM11]. Jazz [Kub20, Tei11, Sau12, Ram01]. Jean [Par19, Del07, Hut20, Lin11]. Jean-Christophe [Hut20]. Jean-Marie [Par19]. Jeder [Lar07a]. Jeff [Cap14, Lew07]. Jefferson [Haa13, Ric15]. Jeffrey [Din23b, Car11, Dan16a, Kop16, Nei10b, Sza12a, Wir07a, Yor18]. Jeffreys [Low06]. Jeffreys-Jones [Low06]. Jenkins [Dan11, Wal17a]. Jennet [Her19, Kru07b]. Jennifer [Gil06, Hay13]. Jeremi [Bur08]. Jeremy [Wal21a, Her07a, Kul13, Sch11, Sin18]. Jersild [Lüt22a, Xia23]. Jervis [Lit11b]. Jessica [Ass14, Cul13]. Jesus [Ric10]. Jewish [Bra04, Din07, Din23b, Git07, Rub06, San16, Wei02, Whi11b]. Jewry [Fei10]. Jews [Pet08, Pol05, Rub06]. JFK [Cat05, Koc10, Mar20]. Jim [Con23, Whi13, Rya08a]. Jimmy [Haf08b, Soa13, DeR17, Soa06]. Jinping [Dre21]. Jistuo [Dob10b]. Joachim [Dap14, Bru06]. Joanna [Spe15, Spe14]. Jockel [Mas11a]. Jockeying [LX14]. Joe [Jen17]. Joel [Bai18]. Joes [Rem08, Moi15]. Joey [Ces21]. Johann [Kar14a]. John [Ass10, Att23, Cha21, Che19, Che08b, Gaf11, Gre05b, Gre05a, Haa20, Hol15, Mil21, Nor20, Par16, Pre11a, Sny08, Whi16, Bak09, Bur10, Dav11a, Gra13b, Hol10, Joh06, Kam13, Kra13b, Kur13, Lat16b, Lic20, Lit11a, Low06, McB15, Mur05, Pra08a, Sib11, Whi13, You12a, vC13, Mid02, Yaq02]. Johnny [Mil14, Kle11]. Johns [Lat08, You12b]. Johnson [Bar07, Byr12, Cri20a, Dum20, God14, Hes02, Ler15, Man10, McP19, Pos07, Soa22, Spe18, Str09, Wal12b, Cur15, Joh07a, Lin09, Lum04]. Join [Neh22]. Joint [Bro02a, Goo15, Goo17]. Jon [Nor09]. Jonathan [Bug15, Dum20, Car13a, Cea17, Cha08, Gal16a, Hor07, Lei08, Loe10, Mad07, Wal12b]. Jones [Gaf11, Hou16, Low06, McM12a, Pra21, Rak15]. Jordan [Ash10, Cit14]. Jörg [Mac20a]. Joseph [GP11, Lov10, Mas11a, She12c, She13]. Josephson [Far15]. Joshua [Sou21, Geo13, Len18]. Journal [Dan10]. Journalism [Mic08]. Journalists [Bar19]. Journey [Gar10b, Kro08]. Journeys [Dan12b, Ram11b]. Journoud [Ass15]. Joy [Kuk21, Lie21]. Joya [Tal11]. Joyce [Din13, Mag14a, Rec17]. Jozef [Ryd14]. JPat [McM21b]. Jr [Bur13, Kru07a, Kuz16, Pra08b, Wal08]. Jr. [Stu15]. Juden [Akt10]. Judging [Bor12a]. Judgment [Col19]. Judicial [Iat18]. Judith [Lee21, Mil14]. Juhana [Sal06]. Julia [Bra16a, Die08, Kro12]. Julian [Ald07, Woo12]. Juliane [Car13b]. Julie [Lai08]. Juliet [Lyn16]. Julio [Dom15b]. July [LX14]. June [Gol10, Mas21b, Mil21, Qua04, Thi14, Gol06a, Wal21a]. jüngeren [Mue10]. Juni [Bru09]. Jürgen [Nai08a]. Jussi [Cal06]. Justice [Bur19, tL16, McC20, Woo12, Pai16, Sta13, Yan08].

K-129 [Kas16]. Kacowicz [Dom07]. Kádár [Hei16, Vár09]. Kaeten [Dom15a]. Kagan [Rub11]. Kai [Den21]. Kaiser [Neh11]. Kaji [Ess15]. Kalb [Dan22]. Kaldewey [Kub20]. Kalemat [Yos15]. Kalic [Col15]. Kalinovsky [Bit17, Gra12b, Gra12a, Kan12a]. Kaltem [Mas19b]. Kalten [Bis14, Huh20, Kan14, Nue07, Sch08, Sch10]. Kalter [Bru05, Moe03]. kam [Boy23]. Kamberović [Kar15a]. Kampf [Moe03, Sch10]. Kandiah [Mur07]. Kansas [Bes07, Hit16, Mil17b, Wea12, Wir07b]. Kaplan [Mas21b, Sha18, Thi14]. Karabakh [Mel01]. Karel [Kes07]. Kari [McC20]. Karin [His13]. Karl [Par10]. Karla [Kre12b, Kre12a]. Karlsch [Luň10]. Karner [Bis10]. Karp [Fre07]. Karr [Usd20]. karta [Min19]. Kashmir [Gan07]. Katalin [Hei16]. Kate [Har18, Sny07]. Katerina [Pap23, Huh20]. Kathleen [Dum17]. Kathryn [Hay06, Lan10, Lat08, Moe18]. Kati [Cra17]. Katyn [Clo09]. Kaufman [Zie21]. Kay [Sny09, Usd22]. Kazakhstan [Bud22]. Keating [Hol99]. Keeley [Mav18]. Keep [Bur17, Goo05]. Keepers [BS02]. Keeping [Njø02, Cou04]. Keith [vD21]. Keller [Rat21]. Kelly [Spr17]. Kemper [Gem18]. Kempton [Dan11]. Ken [Her20a]. Kenez [Bor09]. Kennan [Eng08, Ber13, Fra09, Her21, Her13, Kam13, Kra13b, Kur13, May13, Pec13b, Sme13, Ste13, Wal13a, Mis23a, Mit04]. Kennedy [Bla15, Che19, Gre05b, Gre05a, Hes02, Kli05, Mag10, Mid02, Mun21, Nie08, Nut02a, Par16, Pod14, Pos07, Pra08a, Sba05, Sch17, Sco09, Tau02, Whi16, Ban06, Bar19, FF22, Hol15, Jon19, Mat15, Nue06, Nue07, Bay01, Doc05, God14, Loc05, Tra02]. Kenneth [Dan15, McA14, Aft13, Kru08, Lub11]. Kenney [Sto07]. Kent [And23, Cra13, Mas21b, Thi14, You12a, Chu19]. Kentomania [And23]. Kentucky [Aft13, Bur13, Kuk13, tL16, Liv07, Mat16b, Mil21, Pre11a, Vei13, You12b]. Kern [New07]. Kessler [Hay06]. Ketelaar [Mag10]. Kevin [Soa20a, Cap08, Lib09, Pra21]. Key [Her14, Yon08]. KGB [Cro00, Kni03, Mar09, Neh21, Sel21b, Kle23]. KGB’s [Sel19]. kh [Odo02, Rob08]. Khan [Dre21]. Khanin [Git07]. Khlevniuk [Maw08, McD06]. Khmer [Zha15]. Kholmshchyna [Sny06]. Kholodnaya [Gol12]. kholodnoi [Odo02]. Khomeini [Fri18]. Khong [Luc07]. Khoo [Zha12b]. Khristoforov [Kar15c]. Khrushchev [Mun21, Pod14, Sco09, Tau02, Tro16b, Gor01, Kul13, Ran12, Ree16, RLV^{+}10, SX09, Tor22, Eng20b, Ber01, Tom04]. Khrushcheve [Gor01]. Kill [Sha18]. Killed [Kra06a]. Kim [Cli15b, Del22, Pai14, Stu19, Tha06]. Kimball [Wir07a]. Kimmage [Whi10]. Kind [Man13]. King [Ash10]. Kingmakers [Par10]. Kingston [Wir13]. Kingston-upon-Thames [Wir13]. Kinzer [Aft20, Pra14]. Kirby [Coh06, Xia08c]. kirchlichen [Shl06]. Kiril [Zhu16]. Kirk [Che18]. Kirkendall [Str12]. Kirkland [Sie06]. Kirksville [Sch13]. Kissinger [Bur08, Cal06, Dav12, Gar09, Gus10, Key16, Lin11, All15, PS20, Tra22, Eng08]. Kitchen [His13]. Kittel [Rie08]. Kizny [Ber07]. Klaas [Bis14, Kan14]. Klartext [Mey11]. Klaus [Goo05, Haf08b]. Klehr [Cha21, Mur05, Sib11, Usd20]. Klein [Cha07a, Jon07]. Kleinstaaten [Huh20]. Klimke [Dap14]. KMT [ZP23]. Knell [Clo06]. Knopf [Pos07].

Know [Pra08b, Tor22, Par13]. Knowledge [Dom15c, Fri04]. Kokoshin [Zub99]. Koleva [Pét13]. Kolyma [Pol07b]. Komandirovka [Nai21]. Komárek [Luk01]. Komer [Rak15]. Komine [Lin11]. komissii [Mue10]. Kommission [Mue10]. kommunistische [Kra01]. kompleks [Odo02]. Komsomolets [Dan07]. Konfrontation [Bru05]. Kong [Cho10, Oye14]. Konrad [Sch05]. Konstantin [Tum16]. Kontinuitäten [Pen18]. Kontrolle [O’S17]. Kontroverse [Rug16c]. Kontroversen [Mil17a]. Kontum [Vei13]. Konzepte [Bru08]. Kook [Cli15b]. Kooperation [War06, O’S17]. Korea [Bon07, Eck14, Gre01, Hat10, Jer11, Joh22, Lee12, Ler06, Mat04b, Mat16b, Mil14, Mil21, Suh08, Sza10b, Wea12, Kim16, Lan01, Lan20, Ler04, Sta15, Sza12b, SJ23, Tay15, Cli15b, Don05, Zal04]. Korean [Boo03, Cli15a, Den06, Lea01, Lee06a, Liv07, Mas21b, Mat20a, Mil01, Mil15b, Mil15a, Mil17b, Stu15, Stu19, Sza10b, Thi14, Wea15, CK11, Haj12, Lan20, Mid11, SX14, She06, Zhi00]. Korean-American [Sza10b]. Koreans [Hat10]. Koreas [Sch13]. Korespondence [Gáš13b]. Korhonen [Mas19a]. Kornai [Kro08]. Kornbluh [Lou16]. Kort [Mad09]. Kosmopoliten [Akt10]. Kotkin [Kur15]. Kowalczuk [Bru09]. Kozlov [Zub14]. Krabbendam [Sha07b]. Kraft [Jos20]. Kraushaar [Cra07]. Kreditanstalt [Wen09]. Kremlin [Byr01, Kul13, Rad12a, Wes00, Zhu20, Mas19a]. Krenn [Mar08b]. Krieg [Bon07, Bru05, Huh20, Kan14, Moe03, Sch08, Mas19b, Nue07]. Krieges [Sch10, Bis14]. Kriegsgefangenen [Bis13]. Kriegsschauplatz [Ban14]. Krise [Ado08, Haf08b]. Krisenfällen [Jon06]. Kristie [Bru11]. Kristin [Man10]. Krivitsky [New07]. Krona [Wys12]. Krüger [Zab16]. KS [Bes07, Hit16, Mil17b, Wea12, Wir07b]. Kuan [Iri11]. Kuan-Hsing [Iri11]. Kubakrise [Nue06]. Kuhne [Kle23]. Kuhnhardt [Ser13]. Kuklick [Eng08, Nam16]. Kultur [Nue07, Sch10]. kulturelle [Huh20]. Kulturpolitik [Kan10]. Kunst [Kan10, Kan10]. künstlichen [Kan10]. Kuomintang [Rad17]. Kurds [Lit10]. Kuromiya [Ert10a, Ert10b]. Kurt [Gem18]. Kurzbach [Die08]. Kushner [Pai16]. Kuzmarov [Sch11]. Kwame [Tel18]. Kwan [Lat16a]. Kwanho [Cli15a]. Kwon [Oli11]. KY [Bur13, Kuk13, tL16, Mat16b, Mil21, Pre11a, Vei13, You12b]. Kyle [Gil18].

L [Ass10, Ble16, Bur13, Den06, Eva23, Har08, Iri02, Lak11, Lat08, Lee21, Man10, Mar08b, Mas23, Moo23, Pre11a, Ran06, Soa09, You12b]. Labor [Dav07, Hay13, Mat15, Pud12, Sch02b, Tro16a, BJ10]. Laboratory [Sch20]. Labour [Con07, Tho11, Vic08]. Laetitia [Del23]. Lagos [Pap23]. Lai [Oli11]. Laković [Lee17]. Lama [Dan14]. Land [Hor07, Kim16]. Landlocked [Aki02]. Landscape [Zie08]. Landscapes [Din14b, Sch02c, Sta13]. Landsman [Cal07a, Cal07b, Por08]. Lane [Gaf07]. Lang [Cog12, Pec13a, Pud12]. Langenbacher [Sza12a]. Langley [Whi03b]. Language [And05, Mic21, Nei00]. Lanham [Dei13, Gel12a, Har12b, Jer13, Jer16a, Joh14b, Kol15, Lee17, Moy11, Nic12, VA12, Yor17b, vD21]. Lanoszka [Kno20]. Lanoux [Rut10]. Lansdale [Cha08]. Lansford [Bea06]. Lansing [Bru16]. Laos [SF11, Tra02, Che19]. Lapidus [San16]. LaPorte [vD12].

Larry [Usd22, Gel12a, Mat16b]. Last [Car13b, Cra17, Dom08, Len18, Mas11b, McM12b, Mei15, Nei10a, Nor08a, Rub10, Taa13, WLX^{+}08, Xia08b, Yon08, Pav19]. Lastenteilung [Lar07a]. László [Hut19, Bal10]. latach [Joh21, Kaź19]. Late [Con07, Hes13, Moi03, Nie10b, Pee23, Bar22, Whi12]. Late-Cold [Nie10b]. Lateinamerika [Huh20]. Later [HSB^{+}23]. Latest [Bra16a]. Latham [Man13]. Latin [Dar12, Dar15, Dar16, Dom07, Dom08, Dom09, Dom15b, Iac22, Lov10, Lov22, Low08, McP19, All14, FF22, Har19]. Latina [Kel19]. Latino [Lub11]. Latvia [Sil05]. Latvian [Zak10, Zak18]. Lauren [Haw07, Hay06]. Laurence [Mil06, Mar13b]. Laurent [SF11]. Laurien [Joh20a, And20]. Lauro [Joh07c]. Lautsch [Ban14]. Lavagno [Dra18]. Lavender [Str09]. Lavery [Ram11b]. Law [Cro15, Eva13]. Lawrence [Bes07, Che08a, Dom15b, Hit16, Lom09, Mas21b, McM10, Wea12, Wir07b, BMLG11, Gly06, Hoc16, Ler15, Thi14, Wil08b]. Lawson [Sny08]. Layne [Gre10]. Lazarev [Dav07]. Lazic [Lee23]. LBJ [Win15, Bra06, Div01, Hes21]. Leab [Ros10b]. Lead [Nin03, Pip11]. Lead-up [Pip11]. Leader [Joh22, Tor22]. Leaders [Bla03, Par16, Whi16]. Leadership [Bru01, Cog02a, Koc10, LX14, Rut06]. League [Pow01]. Leake [Yor20b]. Leandra [Zar21]. Learned [Din13]. Learning [Wir13]. Leasehold [Duk02]. Leave [Bur23]. Leaving [DF19]. Lebanon [Blu15]. Lebedeva [Clo09]. Lech [Szp09, Szp14]. Lechosław [Joh17]. Lee [Gre20, Hay15a, Ran10, Rub11, Rub11, Soa22]. Leebaert [Jer11]. Leeke [Con23]. LeeWise [Win11]. Leffler [Lev10a]. Lefort [Dim08]. Left [Ale07, Ber04a, Fal20, Kop16, Lat16b, Nut02c, Soa06, Sto06]. Legacies [Cha07b, Har12b]. Legacy [Abe21, Ber04b, Kol15, Kur05a, Rat21, Sch13, Szp11, Yor21, Dan12a, Qua14]. Legal [Car17, Col19, Mag14b, Sza18]. Legend [Joh11b]. Legion [Bus14]. Legitimacy [Chi14]. LeHand [Moe18]. Leiden [Mag14b, Ric12]. Leith [Rak15]. Lemke [Mas16]. Lenin [Cap08]. Leningrad [Kip15, Hou16]. Leninism [Sav09]. Leninist [ST17]. Lenoe [DF07]. Lens [Kal12]. Lentz [Kre12b, Kre12a]. LeoGrande [Lou16]. Leonardo [Rom17]. Leonid [Tsy08]. Leopoldo [Eva09, Mas12]. Lesbianism [Wei12]. Lesbians [Str09]. Lesley [Rya08b]. Leslie [Yor20b, dL13]. L’Espagne [Foo21]. Lessons [Din13, DF19, Gor05, Joh07c, Jos23, Kap17b, Lee06b, Lit11b, Par22, Whi10]. Let [Dan12b]. Letters [Pon01a, Mol16]. l’eurocommunisme [For19]. l’Europa [Mas11b]. Levee [Mun10]. Level [Jac17]. Lewin [Lib07]. Lewis [Ald07, Kam13, Kur13, Mac20b, Joh14a, Rog15]. Lexington [Aft13, Bur13, Dei13, Gel12a, Har12b, Joh14b, Kol15, Kuk13, Lee17, tL16, Mat16b, Mil21, Pre11a, VA12, Vei13, Yor17b, You12b]. Li [Coh06, Lev07, Xia08c, Lin15, Xia14]. Liaison [Sar05b]. Liberal [Gil00, Mad07]. Liberalism [Joh14c]. Liberate [Sta07]. Liberated [Foo21]. Liberation [Bor99, Mul23, Ros09, Tud05, vW19, KG11]. Liberties [Hay13]. Liberty [Byr12, Din23a, Shm23, Cri02]. Libya [Blu15]. Licensed [Sar05b]. Lien [Gar13]. Lien-Hang [Gar13]. Life [Arg04, Ber07, Gre05a, Gro02, Har08, Har23, Hay06, Hay15a, Hay16, Her13, Hes13, Jan08, Joh14a, Kam13, Kes07, Kur13, Lov22, Mar08a, Mar12, McM11, Mis23a, Mol16, Pac06, Pra19, Ran02, Shl09, Soa20a, Usd20, Wal13a, Yor21, Fed17, Eng09, Hay16, Kle11, Rut11, Nor08b, Div01]. Life-Line [Eng09]. Lifecycles [Hut20]. Lifestyle [Ran10]. Light [CT03b, Gil06, Per07, Rad12a]. Like [BJ19, Joh07a]. Lilia [Rut06]. Limitation [Tra22]. Limited [Bur05, Pie16]. limiti [For02]. Limits [Din14a, Din22, FM16, Kam17, Ole11, Olš18, Sny13, Zha23b]. Lin [Dit09]. Lincoln [Pra17]. Linda [War08]. Lindenberger [Sch08, Whi14a]. Line [Lüt08, Ros04, Eng09]. Lines [Dan10]. Linkage [Hof06]. Linkages [Kra99b, Kra99a, Kra99c]. linke [Cra07]. Links [She12a, Wet11a, Wet11b]. Lionel [Whi10, Yon08]. Lippert [Han13]. Lipton [McM21b]. Lisa [Har23, Sau12]. Lise [Kan15]. Listeners [Shm23]. litaratura [Bra16a]. Literacy [Str12]. Literary [Rub20]. Literature [Ger14, Tho08, Whi15, Ket14, Hut20, The15, Eck14]. Lithuania [Kot17]. litso [Bra16a]. Little [Sha07a]. Littlefield [Kir16, Moy11, Nic12, Sco17, vD21]. Litvinov [Rob02a]. Liu [Ros09]. Live [Kap17a]. Liverpool [Bar11b]. Lives [Kap17a, Pai17, Pét13, Wan09]. Living [Bra16b, Duk13, Har15, Hes13, Pan01, Her17]. Llewellyn [GSP^{+}19]. Lloyd [Pra08b, Sap03, McA07, McA08]. Lobbying [God11]. Lobbyists [Pay21]. Local [Moï10, SL23, Car11]. Locher [Hen09]. Logevall [Cal09, Che08a]. Logic [Wir03]. Lomagin [Kip15]. London [Bis12, Bra16a, Cea17, Chu07b, Din14b, Gra12b, Gra12a, Imm16, Jak16, Kul13, Lew12, Mas12, Onu14, Owe07, Rob14, Sco17, Sha11, Sta13, Tho11, Whi11b, Whi13, You12a, You13b]. Long [Bal08, Cea17, Ger13, Gra12b, Gra12a, Imm16, Jan08, Rad16, Wei01, Yor23b, Ces21]. Long-Range [Wei01]. Look [Bor12a, Sza12b]. Looking [Bon05, Mar08a]. Looking-Glass [Mar08a]. López [Wee18, Iac22]. Loren [Bal17]. Lorenz [Lef21]. Lori [Usd16]. Lose [Che19, Wir06]. Loses [Ste18]. Loss [Sme15]. Lost [Ale07, Div00, Eng09, Gar01b, HTP^{+}15, Lat14, Law10, Qua23, Rad17, Rob02a, Sen16, Usd11, Zub13, Low06]. Loth [Duk10, Gri10, Lie08]. lotta [Dra08]. Lotus [Cat08, Law08]. Louis [Kho14, Lic20, Qua13, Rot21, Rya10, Sim17]. Louise [Yor15, Cea06]. Louisiana [Dan14, Dei15]. Love [Bon05, yL01a]. Loved [Sza11]. Lower [Jan08, vH09]. Lubna [Gus10]. Lucas [Ric12]. Luce [Hol99]. Lucile [Sze19]. Lucky [Che18]. Ludger [Ser13]. Ludlow [Lew12, Sar16, Wal17a]. Luff [Hay13]. Lukewarm [Sch16]. Lumumba [Nam16]. Lure [Kal16]. Lustiger [Rub06]. Lutes [Bel09]. Lüthi [Lef21]. Lutter [Nai10b]. Luxembourg [Par19]. Lviv [Sol13]. Lynch [Sim14]. Lyndon [Dum20, Ler15, Man10, Pos07, Wal12b, Cur15]. Lynn [Hei16, Nei10b]. Lynne [Kru07a, Lin11, Sil11]. Lysenko [Bal17]. Lysenkoism [dL13].

M [Ado08, Ass14, Bar11a, Bit17, Bre21, BSK^{+}23, Car13a, Clo09, Dei15, Dob10b, Dom07, Gaf11, Gra12b, Gra12a, Har11, Kim11, Kol15, Lee21, Lef21, Lou16, Mar07b, Par09, Pra17, Pre11b, Sak06, Sch17, Whi14a, Wir07b]. MA [Abe21, Bal17, Bel11, Ble16, Bru16, Cli07, Cli15a, DeM12, Din07, Edw16, Fer15, Gal07, Gan16, Gra12b, Gra12a, Hes21, His13, Hor07, Hus13, Kal16, Kle16, Kuk11, McM12b, Mil15b, Nam16, Pai16, Roh17, Sae07, Sol13, Spr07, Tei11, Woo12, Zar21]. Maar [Kal22]. MacArthur [Mil17b]. Mace [Ryd14]. MacFarlane [Luc07]. MacGregor [Koc10]. Machcewicz [Joh21]. Machine [BD22]. Machinery [Gar09]. Machines [Fri04]. Machtpolitiker [Kur10]. Mackenzie [Eva07]. MacMillan [Joh11a, Ban06, Bay01, Doc05, Her16, Lee21, Rub11, Rut12a]. Macrakis [Bru11]. Maddrell [Ald08]. Made [Abe21]. Madeline [Lee20]. Madsen [Ham18]. Mafia [Yor21]. Magazines [Pit23]. Magic [Jer11]. Magyarország [Vaj20]. Maier [Hus13]. Maine [Dan16c]. Mainland [KM22]. Maiolo [She12c, She13]. Make [Jer16b, Neu03]. Making [Bar20, Bla15, Bor05a, Dav11b, Fel05, FF22, Gar09, GSP^{+}19, Hat10, Hat11, HGP^{+}23, Hof20, Hou15, Hur05, Kal09, Kli05, Kot17, Kra99b, Kra99a, Kra99c, Law05, Lie12, Lou23, Mad08b, Nei08, Pai17, Qua04, Rec17, Rus04b, Ryd14, Sön22, Tro16a, Vei14b, Wes00, Whi02b, Cro07, Lin11, Moï10, Kip03, Kur05b]. Malarkey [All15]. Malaya [Dee07]. Malaysia [BJ10, Gar04a]. Malaysian [Gar02]. Malden [Cli07]. Małgorzata [Tum14]. Malte [Shl06]. Mamoudou [Bri15]. Man [Bar11a, Kle23, yL12, Mol16, Tom04, Her19, Iri02]. Management [Kap16, Tag22, ZP23]. Managing [Szp10, Kir09]. Mandarins [Boc06]. Mandler [Lat14]. Manduley [Yor21]. Manela [Hus13]. Manetti [Sim16]. Maneuvering [SJ23]. Manfred [Nai08b, Rie08, She12a]. Manfried [Bis11]. Manhattan [Bar08, Mil17b, Pan01]. Manke [Huh20]. Mankind [Kre12b, Kre12a, Lev10a]. Mann [McP19, Win15]. Many [Jer22a, War01]. Mao [Ber14, Bla15, Che99, Dre21, LX14, Rec17, SX09, She06, Won04, Lev07, Lüt14, RTYD19, RLV^{+}10, WLX^{+}08, Zha23a, Cle04, Kru05, Rad15, Ter05, Xia14]. Maoism [Etc23]. Marble [Nor20]. Marc [Hit03, Cog03]. Marcello [Gaf07]. March [Yon08]. Marchais [Van04]. marchés [Mar13b]. Marci [Chi14]. Marcia [Dar15, Dar16]. Marco [Gra07]. Marcus [Whi14a]. Marek [Jol07]. Margaret [Joh14b, Lat14]. Marginal [Gre01]. Margolies [Bos15]. Maria [Duk13, Ken13]. Mariager [Ing17]. Marian [Kip15]. Marie [Ing17, Par19, Tei11]. Marilyn [Moï10, Nei10b]. Marine [Hof20]. Marines [ZP23]. Mario [Dav12]. Mark [Cul13, Din23a, Dom15b, Gil23, Hay15a, Kho14, Ler15, Shm23, Ale09, BMLG11, Cal07a, Cal07b, Che08a, Her11b, Kap17a, Lom09, Mad08b, McM10, Moï10, Por08, Ram08, Ram11a, Ran12]. Market [Büt05, Mar12]. Marketing [Fav18]. Markets [Won04]. Markku [Rob14, Wal17b]. Marriage [Kru04a]. Mars [Sel21a]. Marshall [CKP05b, Her21, Mai05, Tra05, Whe06]. Martha [Die07]. Martial [Bel09]. Martians [Sha02]. Martin [Chi14, Dap14, Kau02, Lar22, Bla03, Mat10b, Nor08a, Pol07b]. Martinus [Mag14b, Ric12]. Marton [Cra17]. Martyrs [Sha02]. Martz [Wil08b]. Marvin [Dan22]. Marxism [Sav09, Sil99]. Marxist [Hol06]. Mary [Mar16, Rub06, Bru10, Rus16, Sny11]. Marz [vD05]. Masculinities [Dum17]. Masculinity [Rob20]. Mass [Chi04, Gor01, GRS^{+}12, Gun02, Sch08, Sny09]. Massachusetts [Boy11, Gen17, Law12, Sch11, Whi11a]. Massacre [Dom08, Oli11, Wal21a]. Massenmedien [Sch08]. Masses [Cal03, DF07, Hal07b, Sch01, Ran10]. Massie [Dan16c]. Massive [Jon08, Gre01]. massiven [Ban07]. Massovye [Gor01]. Mastatskaya [Bra16a]. Mastny [Dav10]. Masuda [Mil15b]. Match [Ste14]. Material [Čav19]. materialov [Kar15c]. materials [Kar15c]. Materski [Clo09]. Matray [Sch13, Stu15]. Matter [Aft03]. Matters [Luň10]. Matthew [Etc23, Mic21, Pra21, DF07, Gen17, McM12a, Moe07]. Matthews [Joh06]. Mattia [Blu15]. Mature [Car13b]. Maturity [Cli15a]. Matynia [Tis15]. Mauer [Nue06, She12a]. Mauerbau [Ado08]. Maurizio [Mil06]. Max [Hay17, Kor10, Pos07, Sac18]. Maximilian [Rug17]. Maxwell [The15]. May [Alp09, Dap14]. Mayhem [Jer11]. Mazov [Kan12b]. mbH [Cra07]. McCamley [Duk11]. McCarthy [Whi11b]. McCarthyism [Hay03, Hay15b, Jen03, War01, Goo10]. McConachie [Aro06]. McDonnell [Eva13]. McFarland [Haa13]. McGill [Ken13, Mas11a]. McGranahan [Myl12]. McHale [Gal09]. McIntire [Wal17b]. McKenna [Vei13]. McLaren [Gal16b]. McLaughlin [Mar20]. McNamara [Owe07]. McPherson [Low08]. MD [Dei13, Gel12a, Har12b, Jer13, Jer16a, Joh14b, Kol15, Lee17, Moy11, Nic12, Ram11b, Rub11, VA12, Yor17b, vD21]. ME [Dan16c, Dan16c]. Mead [Lat14]. Means [Poc03]. Measure [Pec13b]. Measures [Neh21, Sel21a]. Mecca [Yor18]. Media [Che20, Sch08, Vil16]. Mediate [Mun12]. Mediation [McK15b]. Medical [Lee21]. Medicine [Iac22]. Medvedev [Mar21]. Meeting [Bro22, McM11]. Meetings [Gre05b]. Meets [Gro02]. Meier [Usd11]. Melanie [Kul13]. Melinda [How14]. Melis [Wet09]. Melita [Nor15]. Melson [How22]. Melville [Eth05]. Melvyn [Lev10a]. Members [Fet20]. Memoir [Pra07]. Mémoires [Lef02]. Memoirs [Bro06, Gro07, Kro08, Lef02]. Memorial [Kra00]. Memorials [Lem14, Mac20a]. Memories [Har12b, Myl12, Pol05, Sze19]. Memory [Hay17, Kar15b, Kot17, Kub07, Lem14, Nei08, Sta13, ST17, Tum14, Her17, Law07]. Men [Bac02, Kis06, Mis23b, Pai16, Pit23, Pra20, Pai16, Pav19]. Menace [Joh14b]. Menand [Rot21]. Mendelson [Pra99, Van99b]. Mendès [Mar16]. Mending [Ell03]. Menzies [BJ11]. Meredith [Pai17]. Merrilyn [Kir08]. Mertha [Zha15]. Merve [Bar20]. Messages [Die01, Yar09]. Messenger [Foo21]. Method [Iri11]. Mevius [Chi14]. Mexican [Rat21]. Mexico [Tod17, Rat21]. Meyer [Par10]. Mezhdunarodni [Yor17a]. MI5 [Goo05]. Miami [Kam17]. Michael [Aft13, Eva13, Hit03, Kas16, McM21b, Mur07, Pay21, Pee23, Sim14, Sny17, And05, Cog03, Coh10, Dei13, Die07, Din07, Dom09, GB10, Hou16, Jer16b, Joh07c, Kir09, Mad09, Man13, Mar07a, Mar08b, Mur07, Nor10, Onu14, Pon13, Ric10, SS18b, Sco09, She07, Wat17, Whi10]. Michał [Kaź19]. Michnik [Kos15, Tis15]. Micronesia [Hez02]. Middle [Ash10, Bal15, Blu15, Joh07b, Kli05, Kna05, Lak11, Lef21, Owe07, Par10, Qua04, Qua09, Qua21, Rad12a, Sha07a, Sha08a, Sim10, Wat17, Wes00, Win17, Din10, Her14, QMLL16]. Middlebrow [Cha07a, Jon07]. Midnight [Sco09]. Mieczkowski [Wan16]. miedzynarodowej [Kaź19]. Mighty [Bel11]. Migrants [Ver18]. Migration [Att23, Cha07b, Pai17, SX14]. Mikael [Wys12]. Mike [Dob10b, Wir13, vD12]. Mikhail [Gol12]. Mikoyan [Pod14, SS18a]. Miles [Mas21a]. Militant [Cos22, Pow01, Sza12b]. Militants [Fal20]. militärische [Huh20, Kar14a]. Militarization [Col15, Gro02, Kuk21]. Military [Ald07, Ash10, Ava04, Bel09, BJ10, Car13a, Den06, Duk11, Duk13, Eng02, Gho00, Gra13b, Ham18, HP17, Her05, How22, Ing04b, Jer13, Jer16a, Kal12, Kar04, Kim16, Lee17, Moo23, Nam13, Pap23, Pra15a, Pra15b, Rad16, Rob04, Rog15, Rya08b, Sar05b, Sch03, Sie06, Sto14, Tay03, US21, Whi02b, Win11, Wys12, vC13, Rus07, Kel02, Sou21, Tay00]. Miller [Hat11, Mue17, You13b]. Millett [Suh08, Wea12]. Millionaire [Usd20]. Milorad [Lee23]. min. [Rut12b]. Mind [Aft20, Din18, Gen17, Her12, Sha08b, Lar22, Mur04]. Mind-Set [Her12]. Minds [Osg02, Sha08a]. Mineshaft [Gei12]. Ming [Dre13]. Mingjiang [Xia14]. Minh [Spe13, Zha09]. Minister [Mil13]. Ministerium [Sch10]. Ministries [Yon08]. Ministry [Sel19, Sel21b, Sas14a]. Minneapolis [Joh13a, Nor07]. Minnesota [Joh13a, Nor07]. Minorities [vD12]. Minority [Lee20]. Minsk [Bra16a]. Minute [Sco09]. Minuteman [Fer15]. Minutes [Bar22]. Minzhurenko [Nai21]. Mir [Zub14]. Mira [Lar07b]. Miracle [Cal22, Cli15a]. Miracles [Sha02]. Miriam [Ree16]. Miroslav [Wil17]. Mirror [Gel12b]. Mirrors [Roz11, Gar10b]. Mis [Gel12b]. Miscalculation [Ler04]. Miscamble [Mad15, vDO08]. Misconduct [Iat18]. Misković [Jak16]. Misreading [DX18]. Misrepresentation [BJ11]. Missed [Gre08, Jer08, JMG08, Mad08a, Mas08, Fre07]. Missile [Col07, Dan10, Fer15, Gre05b, Haf08a, Her04b, Her04c, Her20b, HSB^{+}23, Hol99, Koz14, Pod14, Pra08a, Tie07, Wel06, Rom17, Col19]. Missiles [Col07, Pod14]. missili [Rom17]. Missing [Pas10]. Mission [Boc10, Bra02b, Che18, Mat12, Pac06, Sar05b, Sep01, Sew23, Sta07, Wil08a, Żuk19, Wil08a]. Missions [Dom03, Gal22]. Mississippi [Sau12, Wal21b]. Missouri [Ham15, Kre12b, Kre12a, Roz11, Rya10]. Missy [Moe18]. Mistificatori [Sni22]. Mistrust [Lib01]. Mistry [Dom15a]. Misunderstanding [Yor23b]. Misuse [Hei06]. Mitchell [DeR17, Din14b, Haf08c, Sta13, Whi11b]. Mitelpunkt [Din18]. Mitgliedstaaten [Sza07]. Mitteilungen [Mue10]. Mitterrand [Cog02b, Mar16, Sar11]. Mixed [Die01]. MO [Kre12a, Sch13]. Mobilization [Koz14]. Mobilizing [Car99]. Mobley [Ler06]. Mochizuki [Dob10b]. Model [Lee20]. Modern [Bal15, Den21, Gaf07, Hat11, Her12, Hof20, Par10, Qua04, Rec17, Rus07, Sch11, Soa20b, Sön22, Vei14b, Sza14]. Modernism [Whi12]. Modernists [Whi15]. Modernity [Kot17, Sil05]. Modernization [Boc06, Boc10, Cra13, Kal18, Man13, PW05, Soa22, Laz07, Nut02a]. Modest [Ula99]. Mohamed [Yos15]. Moïse [BMLG11, McM21a]. Molden [Bis11].

Moldovans [Mic21]. Mole [Usd20, Usd22]. Moles [Cla09]. Möller [Mue10, Rug18]. Molotov [Rie15]. Moment [Din23b, HT19, Sur13, Wan16]. Momentous [Kra13a]. Moments [Sto07]. mondo [Rom17, Gun12]. Monetary [Hel07, Wid05]. Money [Loe10]. Mongolia [Rad14, Aki02]. Mongolian [Ros09, Rad06]. Monica [Stu19]. Monnier [Rom17]. Monod [Wen09]. Monolith [Coh12]. Monroe [Kap14]. Montefiore [Ale08]. Montreal [Ken13, Mas11a]. Monument [Kar15b]. Monuments [Kar15b]. Moon [Duk13, Joh13a, Neu11, Har18]. Moore [Mil14, Mil06]. Moral [Aft02, Tie07]. Morality [Kos15, Eck07a]. Moravcsik [Gil00, Hof00, Kee00, LSvdV04, Mil00, Tra00]. Moreno [Dom15b]. Morgan [Cap08]. Morgenthau [Gol03]. Morisy [McM21b]. Moro [Dra06, Dra18]. Morozov [Gol10, Pip15]. Morozova [Rad14]. Mortal [Bar08]. Moscow [Ell03, Mar10, Mas21a, Sch19, Sim17, Ste05, Gle06a, Gle06b, Kus02, Lüt09, Mic11, Mue11, SX09, Spo12, Tra09]. Moses [Fin13]. Moshe [Lib07]. Mossadegh [Gas19]. Most [Bor12a, Gra21a, Kle23, Mad15, You07]. Mother [Fes07]. Mountbatten [Rey23]. Movement [Cat08, Etc23, Gly06, Hor07, Jac23, Jak16, Kal22, KW22, Law08, Lüt16, Spe01, Sto06, Vei14b, DeM12, How14]. Movements [Mey11]. Movies [Whi13]. Moyar [Ale09]. Moynihan [Sur13]. Mozambique [Rei19]. MPLA [Čav19]. Mubarak [Yos15, Yos15]. Much [Che19]. Mücke [Wil17]. Muddling [Jan08]. Mudhakkirat [Yos15]. Muehlenbeck [Lar17, Par16, Whi16, Woo13]. Mueller [Wil08a]. Müller [Bis14, Las19, dL13]. Müller-Enbergs [Las19]. Multilateral [Wil23]. Multilateralization [Wen04]. Munevera [Woo07]. Münger [Nue06]. Munich [Bar03b, Bis13, Mer13, Rug16a, Gut23]. Munitions [Chu09]. Murder [Dra06, Sny09]. Murdering [Nam16]. Muriel [Bis11]. Murrow [Sch17]. Muse [Mar07b]. Museum [Nie12]. Museums [Mac20a]. Mushroom [Win15]. Music [Ans17, Cal14, Kub20, Pos19, Wen09, Zhu16, Ans17, Ram08, Spr17]. Muslim [Joh13a, Woo07]. Muslims [Kli22]. Mutual [Dra12a, Sza14]. Mutually [Sha18]. My [Bal05, Fri18, Oli11, Rut12b]. Myanmar [Sza14]. Mysteries [Cla09, Ula99]. Mysterious [Luk01, Żuk19]. Mysticists [Sni22]. Myth [BJ11, Chu19, Pik02, Sch11, Szp07b, Szp14, Tro16a]. Mythology [BJ10, Pik02]. Myths [Mil20].

N [Col15, Dav11b, Nei10b, Pai14, Tha06, Ram11a]. N. [Lew12, Wal17a]. nach [Mer13]. nachalo [Odo02]. Nada [Jak16]. Nadia [Val14, Yor17a]. Nagorno [Mel01]. Nahe [Ste20]. Näherungen [Kar14a]. Naimark [GRS^{+}12]. Name [Mil15a]. Namikas [Kan15]. Nancy [DeR17]. Nanjing [Guo10]. Naoko [Sae07]. Narcotics [Cra13]. Narizny [Lib09]. Narrative [Ran02]. Narratives [Åse16, Bar03a, Cea17, Lyn16, Spr17]. Nash [Tho08]. Nashel [Cha08]. Nashville [Woo13]. Nassau [Mid00]. Nasser [Ash10, Owe07]. Natalia [Clo09, Mul23]. Natasa [Jak16]. Nathan [Kal18]. Nation [Åse16, Kan10, McE03, Nut02a, Sfi02, Sil05, vD21, Hat10, Dou06].

National [Bal08, Bla15, Bra06, Bur04, Chi14, Coh08a, Dan16a, Geo13, Hei16, Jac23, Kar15b, Kus02, Mar07b, Mic11, Mid11, Moï10, Pol07a, Pol13, Rob20, Rut10, Smi11, Søn19, Wei12, Wir13, Tro16a, Hez02]. nationale [Ban07, Mas19b, Sza07]. Nationalism [Bra05, Dou15, Hay17, Ing04b, Joh07b, Jon03, Nin03, Spe13, Tho08, Tum03, Wet10, Gor03]. Nationalist [Jac16, Par16, Rea06, Tum16, Whi16]. Nations [Ano03l, Ano08i, Fei10, Gat03, Ken03, Kus03, Lef03, Nei14, Sav14, Sha18, Sny03a, Get15, Göt08, Hei15, Nag17, Nag18, O’D09]. Nationwide [Bru03]. Natives [Lat14]. NATO [Bis11, Han13, Hen09, Jon06, Lar07a, Lee17, Mas16, Pfa02, Spo12, Val14, Ban07, Gol20, Hat09, Hat17, Hel22, Joc07, Kou12, Lut22b, Mas21b, McG99, Ole18, Rea11, Thi14, Uel20, Wen04, Sza07, Ada03, Kap04, Mas21b]. NATO-Mitgliedstaaten [Sza07]. NATO-Strategie [Ban07]. Nature [Guo10, Kom18, Kru05, Mar06b, RTYD19]. Naumov [Kul10]. Naval [Gal16b, Kas16, Rak15, Ram11b, Rub11, Sep01, Win11]. Navy [Dan16a, Fri05, Wir13]. Nazi [Bre08, Din14b, Fin13, Kes07, Mag10, Mar06a, Moe07, Pol10, Rai11, Wet10, Rob02b, vH09]. nazionale [Gor09]. Nazis [Bor09, Jol07]. Nazism [Bru16, Rai11]. NC [Duk13, Gun12, Haa13, Iri11, Kro12, Myl12, Ros17, Wei12]. Near [Kuz19, Gal16a]. Nebraska [Pra17]. Necessary [Poc03]. Necochea [Iac22]. Need [Pra08b]. Neelam [Del23]. neft [Min19]. negli [Nut10, Nut11]. Negotiating [Yor20b, Coh08b, Pét13]. Negotiation [Den21]. Negotiations [Lou16, Lüt09, Mun12, RT13]. Nehru [Bla15]. Neighbors [Gal16a, Roz03]. Neil [Luc07]. Neither [Ole11]. Nelegal [Tsy08]. Nelson [Zie08]. nemzetközi [Vaj20]. Neo [Bis05, Bro02b]. Neo-Atlanticist [Bro02b]. Neo-Revisionism [Bis05]. Neopatrimonialism [Bri15]. Nerve [Iac22]. Nerves [Lar22, Lei08]. Netherlands [Jon03, Sny13]. Network [Fou14, Fri04]. Networks [KNS18, Puc21, SS14]. Neubeginn [Bis08]. Neue [Nai08a, vD05, Pen18]. Neutral [Bis12, Buc17, Eng09, Nil19, Dav11b]. Neutralism [Wys18]. Neutralität [Bru08]. Neutrality [ARN16, AS22, Bru08, Eng09, FAM16, FM16, Geh05, Mue16, Sch16, Yor17b, Sal06]. Neutralizing [Wan17]. Neutrals [ŽČ16]. Never [Ass10, Ger13, Kaź09, Moi03, Pre11a]. Nevsky [Sha12]. Newman [Lee21]. News [Kel19]. Newspapers [DF07]. Newspeak [Cam05]. Newton [Lai08]. Next [Fer15]. Ngo [Ass14]. NGOs [Cho10]. Nguyen [Gar13]. Nhân [Zin11]. Niall [Hus13]. Nicaragua [Soa22, HS15, Sto14]. Nicholas [Soa13, Zha12b]. Nick [Duk11]. Nicolas [Sim14, Mar16]. Nicole [Bis12]. Niedhart [Mas10a]. niezwykłe [Joh17]. Nigel [Ban06, Bar08, Haf08a, Hay11, Qua09, Sas14b]. Nigerian [Wys18]. Nightmare [Sha12]. Nijhoff [Mag14b, Ric12]. Nikita [Kip15]. Nils [Boc06]. Nilsson [Wys12]. Nimble [Dan11]. Nineteenth [Spr17]. Nineteenth-Century [Spr17]. Ninh [Tha06]. Nish [Kho14]. Nixon [Cha14, JLBK17, Lin11, Nut10, Nut11, Wir07a, Ass11, Bur05, Cal09, Gar09, Gus10, Key16, KW09, Kom14, Mar23b, Tho21, Tra11, Tra22, Win17, Sba05].

Nixon-Era [Wir07a]. NJ [Ale12, Bra16b, Coh12, Mat16a, Nam13, Pét13, Pet11a, Pra99, Rus07, Sny11]. Nkrumah [Tel18]. No [Nei10a, Raj14, Sny07, CS07, Den06, Gan06, Pru04]. Noble [How14]. Nodes [Her12]. Nomos [Dav11b]. Non [Ben05, Bud22, Geh05, Jak16, Lan01, Lüt16, Raj14, ŽČ16, Bro06]. Non-Aligned [Jak16, Lüt16]. Non-Alignment [Raj14, ŽČ16, Geh05]. Non-Belonger [Bro06]. Non-Communist [Lan01]. Non-Proliferation [Bud22]. Non-Use [Ben05]. Nonaligned [God14]. Nonalignment [FAM16, Nie11]. None [Ris15]. Nonproliferation [Bra06, Fer07, PS20, PR21, Sto15]. nor [Ole11]. NORAD [Mas11a]. Norbert [Sza07]. Nordic [Vil16]. NORDICOM [Vil16]. Normal [Mou03]. Normality [Pét13]. Normalization [Coh06, Xia08c]. Norman [Car13a, vD12, Cie07, GRS^{+}12, Kas16, Pie16]. Norms [Cra03, Dom07, Eva01]. Norstad [Tal08, Pfa02]. North [Bre21, Dei07, Gar13, Hat11, Hay13, Hel07, Joh07b, Joh14c, Joh22, Lak11, Lee06a, Ler06, Lou16, Mar07b, Mar07d, Mun12, Nie10a, Rab13, Rad16, Ram07, Sha07a, SX14, Smi03, Str12, Wie07, CK11, Dro18, Lan01, Lan20, Lee12, Ler04, Sza12b, SJ23, Tay15, Wea12]. Northeast [Sch13, Wol13]. Northern [Ran12, Ris15, Ing17, JE19, Joh03]. Northrop [Edg08]. Norton [Mag14a, Neu11, Usd11]. Norwood [Nor15]. Note [Ano00j, Ano00k, Ano00l, Ano01l, Ano01m, Ano01n, Ano02n, Ano02o, Ano02p, Ano02q, Ano04o, Ano08d, Ano08e, Ano08f, Ano09e, Ano09f, Ano09g, Ano09h, Ano10e, Ano10f, Ano10g, Ano10h, Ano11e, Ano11f, Ano11g, Ano11h, Ano12e, Ano12f, Ano12g, Ano12h, Ano13e, Ano13f, Ano13g, Ano13h, Ano14e, Ano14f, Ano14g, Ano14h, Ano15e, Ano15f, Ano15g, Ano15h, Ano16e, Ano16f, Ano16g, Ano16h, Ano17b, Ano17c, Ano17d, Ano17e, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano19e, Ano19f, Ano19g, Ano19h, Ano20e, Ano20f, Ano20g, Ano20h, Ano21e, Ano21f, Ano21g, Ano21h, Ano22b, Ano22c, Ano22d, Ano22e, Ano23a, Ano23b, Ano23c, Ano23d, Kra99d, Kra99e, Kra99f, Kra03b, Kra03c, Kra03d, Kra03e, Kra04c, Kra04d, Kra04e, Kra05b, Kra05c, Kra05d, Kra05e]. Note [Kra06b, Kra06c, Kra06d, Kra06e, Kra07a, Kra07b, Kra07c, Kra07d, Kra08a, Rug11, Rug16c, vD05]. Notebooks [HK09, KLHK14]. Notes [Mag14b]. Notice [Kra00]. Notorious [Kle23]. Notre [Dom07]. noveishei [Mue10]. Novel [Fed17]. Novels [Sha08b]. November [Col19, Col07, Pod14]. Novocherkassk [Coo04]. Novyi [Zub14]. NSA [Lon11]. NSC [Lef12, You13a]. NSZZ [God11]. Nuclear [Bae02, Bay01, Boz20a, Boz20b, Bud22, Bur05, Bur17, Bur21, Cea17, Cur23, Duk11, Eva23, Fer07, Gaf14, Gal15, Gly06, Han17, Her16, Her09, JLBK17, Jer22b, Jon08, Jon19, Jos23, KH23, Kno20, KNS18, Kur18, Lei18, Lut22b, Mar01, Mas00, Mas08, Mas09, Mat10b, McM12a, Mid00, Nor05, Nor07, Nor10, O’D09, PA23, PS20, Pol07a, Que02, Rea11, Rep03, Ros10a, SL23, Sar19, Sch01, Sco09, Sew23, Šmi23, Ste14, Sto15, Usd09, Wal21b, Wei01, You07, Zha23b, Zie21, Bar11b, JLBK17]. Nuclear-Armed [Sew23]. nucleare [Eva09]. Nuenlist [Dav10, Hen09]. Number [Mil15a]. Nuti [Eva09, Mas12]. NVA [Ban14]. NVA-Offiziers [Ban14]. NY [Ass14, Dav12, Geo13, Her12, Hoc16, Joh14a, Kal12, Kuz16, Lie12, Lit11b, Man13, Qua21, Ree16, Sim16, Wan16, Whi13, Zhu16]. Nylon [Shl09].

O [Sni22, Sou21, Gáš13b, Sok99]. Obama [Sim14]. Observed [Wei04a]. Observer [Cly13]. Observing [Sie06]. Obsessions [Gun12]. Occupation [Jol07, Kim16, Kor06, Mar06a, Ray20, Shi13, Sny09, SM12, Wil03, Ess15]. Occupied [Bis04, Nai02]. Occupying [Wil08a]. O’Connell [Hof20]. O’Connor [Haa20]. October [Nei08, Qua14, Yos15, Col19, Haj12, LX14, Sim23]. Octobers [Mun21]. Odd [Hol21, Nai22, Nie12, Sha11]. Odyssey [Diz10]. Off [Bar19]. Off-the-Record [Bar19]. Offensive [Pee23]. Offentlichkeit [Shl06]. Office [Rob14, God11, Pif15]. Officer [Gro07, Pac06]. Official [Cea17, Kna03, Mur04]. Offiziers [Ban14]. Ogaden [Yor15]. oggi [Dra21]. Ogneva [Woo12]. Ogneva-Himmelberger [Woo12]. OH [Cra13, Mas21b, Thi14]. Oil [Gas06, She14]. Ojserkis [Mit07]. okeanom [Tsy08]. Okey [Ken07]. Oklahoma [Car13a]. Old [CT03b, Rad07]. Oldenbourg [Bis13, Rug16a]. Oldenburg [Mer13]. Oldenziel [His13]. Oleg [Hay11, Kul10, McD06, Maw08]. olimpijskich [Kaź19]. Oliver [Whi13, Dan12b, Mas10a]. Olmsted [Hay06]. Olsted [Pud12]. Olympic [Soa13, Vau15, Rid16]. Olympics [Gut23, Sar23]. Once [Sny99]. One [Bor12a, Jac16, yL01b, Tau02, Kap17a, Sco09, Tay15, Bre21, McE03]. One-Party [yL01b]. Ongoing [Rug11]. Onset [Hac03, Mak17]. Onslow [KG11]. Onto [Ban16]. Onur [Myl15]. op [Nor15]. Open [JM23, Pra15a, Pra15b, Mar07c, Tal08]. Open-Source [JM23]. Opening [Lee17, Nut02c, Rus16]. Operation [Gas13a, HK20, Rog15, Sel19, Sew23, Pra07, Sel21b, Att23]. Operational [Wir07b]. Operations [Bru03, Don05, Kir08, Lea01, Lei18, Mil01, Pem16, Pif15, Ric07, McA14, Kna02]. Operatives [Mar23b]. Operativgruppen [O’S17]. Opinions [Kre12b, Kre12a, Gan08]. Oppenheimer [Aft02, Aft03, Kru07b, Her07a]. Opportunities [Fre07]. Opportunity [Gre08, HTP^{+}15, Jer08, JMG08, Mad08a, Mas08, Wen04]. Opposition [Dap14, Din23b, Her21, Joh20b]. Oppression [Dum14]. Option [Eng09]. Options [Bur05]. Oracles [Eng08]. Oral [Bal05, BL10a, BL10b, Gar10c, Haa05, Wil17]. Orange [Bra14]. Orbits [Von03]. Order [For20, Gav05, Boc10, Yor18]. Organisation [Whi11b]. Organization [Die01]. Organizational [Her20a]. Organizations [Die07, Hei16, Hur05, Alo05]. Organized [Hei16]. Orientalism [Cha07a, Jon07, Sha07a]. Original [Sar05a]. Origins [Ass23, Bar03b, Blu15, Bro08, Coh08a, Dar12, Dav10, Dom15a, Dra12b, Fed17, Ged99, Gif05, Gle05, Haf08c, Han13, Hes02, Imm16, Kar15b, Kra18, Lan10, Lar11, Mal15, McC14, Neh11, Odo06, Pem16, Pon01b, Qua13, Raj14, Rob02b, Sha20, Sho15, Son02, Whi13, Yor23a, You07, Zha02, Zhi00]. Orsini [Del23, Her12]. Orwell [Ros10b, Hol10]. Oscar [Eng20b]. Osgood [Kru08].

Osirak [PR21]. Osnos [Mas23]. OSS [Has07]. Ossolineum [Joh21]. Ost [Bru05]. Osten [Rie08, Ste20]. Ostermann [Hes12, Wet11a, Wet11b]. Österreich [Bis08, Bis11, Rug17, Wil08a]. Österreichische [Bis11]. österreichischen [Kar14a]. Österreichs [Rug16b]. ostlichen [Nai04, Nai08b]. Ostpolitik [Han13, Lew12, Nie03, Sar16, Nie16b, Nie08, Sza21]. Ostpolitik-Westpolitik [Lew12, Sar16]. O’Sullivan [Pec13a]. Oswald [Gre20]. Other [Blu14, Dan14, Gan08, Ing04b, Lat16a, Laz07, Log01, Lüt14, Pas10, Yon08, SS18b]. Othmar [Bis10]. Othon [Pap23]. Otis [Haf08c]. otnoshenii [Mue10]. otnosheniya [Yor17a]. Ounce [PR21]. Our [Che08b, Dan12b, Her23, Sie06, Fes07, Pas03, Rus04b]. Ours [Kor06]. Ouster [Tor22]. Outbreak [Gol06a]. Outreach [Gue17]. Overcome [Gut23]. Overcoming [Gen04]. Overflight [Pra07]. Overseas [Duk02, Tsy08]. Overstretch [Ste99]. Overthrow [Joh20b]. Overtures [Bor12b]. Overview [Haf08c]. Owen [Bar11b]. Own [Kau02, Pas03]. Oxford [Bel13, Bur16, Cal14, Car13b, DF12, Goo15, Goo17, Guz13, Her11b, Key16, McM14, Par13, Par16, Pon13, Pre17, Ram11a, Sur13, Usd16, WF12, Whi16]. Oyen [Pai17]. Ozawa [RT13].

P [Bak09, Cog12, Gra13a, GP11, Joh06, Kal12, Kru07a, Lev10a, Lic20, Mit07, Roz11, Vei13, Gáš13b]. PA [Kle11, Sto11]. Pacific [Hac03, Kho14, RG15]. Pact [And20, Joh20a, Rob02b, Spa14, Yor19b, Bae16, Bék23]. Paczkowski [Kor06, Sim16]. Padraic [Sto07]. Paid [Pac06]. Pakier [Tum14]. Pakistan [Sim04, Bur21]. Pakt [Bis11, Wet11a, Wet11b]. Palace [Kru07b]. Palasik [Ken13]. Palayret [Par19]. Palestine [Pip15]. Palestinian [Fal20, Kou17]. Palgrave [Her16, Joh11a, Lee21, Rub11, Rut12a]. Palmer [Rab13]. Palomares [Sti06]. Pan [Att23, Soa20b]. Pan-Asian [Soa20b]. Pantsov [Rad15]. Papacosma [Mas21b]. Papandreou [Dra22]. paper [Ham15, Lin11, Onu14, Boy11, Gru07, Gun12, Kan15, Lak11, Ree16, Sch11, Whi11b]. paperback [Joh13a]. Papers [Bar11b, Mar10]. Paradigm [You22]. Paradise [Dad19]. Paradox [Tod17]. Paradoxes [Kur15]. Paraliterary [Bar20]. Paramilitary [Gil18]. Paris [Kar15b, Law05, Mar13b, Mar16, Lüt09]. Park [Cli15b, Kle11, Sto11]. parlamentare [Dra08]. Parliament [Dra08]. Parr [Doc08]. Part [Boz20a, Boz20b, Lut22b, Nag17, Nag18, Sel19, Sel21b, Her04b, Her04c, Kra99b, Kra99a, Kra99c, Kra03a, Kra03g, Kra03f, Kra04b, Kra05a, Mor00b, Mor00c]. Parta [Cri20a, Shm23]. partei [Shl06]. partei-staatlicher [Shl06]. Parteien [Nai04, Nai08b]. Participation [Fro23]. Parties [Lan01]. Partition [Gan08, Tal11]. Partners [Kis06]. Partnership [Mas10b, Moe18, Sha11, Mur02]. Party [Gil05, Kli22, Kno06, yL01b, Łuk23, Sfi01, Shl06, ST17, Tra09, Vic08, You12b, Ass23, Fri18, yL09, Mar23b, Mat20b, US21]. Party-State [Shl06]. Parzellers [Zab16]. Pascal [Sul22]. Pascaline [Cog12, Par19]. Passage [Joh03]. Passing [Chi04, Mil21]. Passive [Duk11]. Past [Gra21b, Leb03, Pap23, Pol13, Rai11, TSW^{+}22, Yan08, Fin13, Ger13, Moe07, Zub14]. Paternity [Sch22b]. Paterson [She07]. Paths [Gan07]. Patients [Ber04b]. Patrice [Nam16]. Patricia [Mag10]. Patrick [Lem14, SS18b, Smi07]. Patriot [Joh11b]. Patrioten [Akt10]. Patriotism [Bel13, Kuz19]. Patrol [Gal16b]. Patrols [Don05]. patron [Lef02]. Pattern [Joh07a, ZP23]. Patterns [Luk01]. Pauker [Wei02]. Paul [Far15, McB15, Ald08, Chu09, Dav07, Eck07a, Haw07, Mar12, Mor06, Sto11, Van11]. Paul-Henri [Van11]. Paulo [Str12]. Pause [Her03b]. Pavel [Wil17, Rad12a]. Paweł [Joh21]. Pawley [Hol99, Hol05a]. Payk [Whi14a]. Payne [McM08]. PCI [Dra08, Dra21]. Peace [Ass07, Bor12b, Cat08, Cou04, Div00, Gar13, Gra05, Gre10, Her03b, How14, Jac17, JMG08, Kac02, Law05, Law08, McA07, McA08, McB15, McC20, McD06, Pet20, Qua23, Rad17, Rob02a, Rus04a, Sch02c, Sfi01, Cur23, Her23, HTP^{+}15, Jos23, Kim11, Mey11, Gan07, Kno06, Man10, Din04, Doc04, Gre08, Jer08, Leb00, Mad08a, Wir03]. Peaceful [FF22, Nie16b, Zie21, Zha23a]. Peacemaking [Din10]. Pearl [Bar03b, Sch09, Tie07]. Peasant [Law12]. Pechatnov [Rob08]. Peck [Coh08a]. Pedro [Att23]. Pee [Søn19]. Peebles [Ric07]. Peer [Ris15]. Pekelder [Dap14]. Peloponnesian [Boo03]. Pen [Duk11]. Penetration [Bar08]. Pennsylvania [Haw07, Jen17, Kle11, Sto11, Whi12]. Penny [Tei11]. Pentagon [Ros17]. People [Bar18, Bru10, Bug15, Dan12b, Fri04, Kub20, ST17, Zie08, vDG07, Kli22, Nor07]. Perception [OB20, Rom17]. Perceptions [Bru03, Gas19, KW07, Kor07, Szp07a]. percezioni [Rom17]. Perestroika [Ber20, Kac02, Kni03, Mat08, Rut12b]. Perforating [Nie12]. Peril [Lee20, For19]. Perilous [Gra13b, vC13]. Perils [Con03, Kim07]. Period [Tra11, Tra22, Nai10a]. Peripheral [Iac22]. Periphery [Mas16]. Perkins [Cli15a]. Permanence [Zim09]. Permanent [Mue16, Sou21]. Pero [Dav12]. Persecution [Din14b, Str09]. Persian [Fri05, Joh11a, Win11]. Person [Mic08]. Personal [Bar04, Hol05b]. Persons [Bur23]. Perspective [Jer06, Mas09, Sar05a, Hus13, Woo13, Mad08b, Mat08, Gar01b, Joh05]. Perspectives [Ano03l, Ano10i, Ass10, BH17, Din10, DFK^{+}20, GRS^{+}12, HP17, JMG08, JLBK17, Kan12a, KWKS07, KSR18, Kra13b, MBR^{+}21, MSPN21, Nai10a, Nie10b, Nue05, Pre11a, RLV^{+}10, SZK12, TSW^{+}22, Whi14a, ZRS^{+}21, vDG07, vDO08, vW19, Ger14, Sas14b, Moï10, Mur07, Gle05, McM05]. Persuasion [Sil04, You13b]. Petals [Win15]. Peter [Cog12, Kar15b, Lou16, Mas23, Nai10a, Nie10b, Pec13a, Pud12, Bor09, Dei07, Gaf07, Gre20, Lak11, Lat14, Yor17b]. Péteri [Jan08, Shl09]. Petra [McM14]. Petrescu [Mac20b]. Petroleum [Mat15]. Petrov [Boy23]. Petruccioli [Dra21]. Ph´ âm [Zin11]. Pham [Sza11]. Phat [Spe13]. Phenomenon [Ris14]. Philadelphia [Haw07, Jen17, Whi12]. Philanthropy [Lun04]. Philatelic [Mod17]. Philip [Alp09, Her11b, Moï10, Vau15, Lar17, Par16, Whi16, Woo13]. Philipp [Rai11]. Philippine [TJ21]. Philippines [Bel13]. Phillip [Hay15b, Pol10]. Phillips [Rea06]. Philosopher [Mol16]. Philpott [Yan08]. Phu [Nor08a].

Physicist [Hay16]. Phythian [Haa20]. Pickl [Bis10]. Pickut [Usd22]. Pierre [Vei14a, Zha09]. Piers [Lew12, Wal17a, Sar16]. Piety [Yaq02]. Pigs [Bar19]. Pine [Har23]. Pinochet [Sil11, US21, Yor19b]. Pinprick [Cor14]. Piotr [Szp09]. Piotrowski [JK07]. Pittsburgh [Smi07]. Pius [GP11]. Pivotal [Kno06]. PLA [Dre05]. Place [McK15b, Sny07]. Places [Bon05]. Plamper [Len12]. Plan [Car99, Sch09, Tal08, CKP05b, Mai05, Tal08, Tra05, Whe06]. Planck [Sac18]. Planes [Hou15]. Planning [Ald07, Dad19, Jon08, Lüt08, Mar02, Sny09, You07]. Planung [Ban07]. Plastas [How14]. Plastic [Bug15]. Play [Mun21, Tei11]. Played [Bel11]. Playing [Mar08c]. Plays [Soa09]. Plokhy [Kim11]. Plot [Mar23b]. Plowshare [Zie21]. Plunder [Mag10]. plus [Rug18]. Plutonium [Nor05]. Poetry [Ket14]. Poets [San16]. Pogrom [Bra04]. Pogromul [Del07]. Point [Cha14, Ler06]. Poisoner [Aft20]. Polan [Kro12]. Poland [Cur07, Jol07, KW07, Kor06, Kor07, Sim16, Szp07a, Dom11, KW22, Pol13, Puc21, Sny99, Szp10, Szp11, And23]. Poland-Israel [Pol13]. Polar [Pet11b]. Poles [Dou15, Kor06, Pol05, Pol10]. Police [Bru04, Fis03, Kap17a, Łuk22, Puc21, Szp10]. Policies [Buc17, Dob10b, Gan06, Jon19, Kna03, Mar05, Ste05]. Policy [All14, Bal08, Bea06, BJ11, Bis12, Bor99, Boy23, Bra06, Bro08, Cap14, Che19, Chu07b, Cly01, CKP05a, Din10, Doc08, Dom11, Dum20, Ell03, Eng07, FM16, Foo21, Gar09, Gav99, GSP^{+}19, Her14, Hez02, Hol05a, Jac16, Jac23, Jen17, Jer11, JM23, Joh05, Jon03, Jon04, Kan10, Kau09, Kho10, Kra99b, Kra99a, Kra99c, Kul13, Lak11, Lar10, Lar02, Lar07b, Ler02, Lin11, Lüt06, Man13, Mar07b, Mar23b, McC14, McD02, McM06, Mid00, Mil22, Mor00b, Mor00c, Mur04, Osg04, Pay21, PR21, Rei19, Ros00, Sar05a, Sar19, Say17, SS18b, She20, SX14, Shi13, Sim10, Sim14, Sny09, Søn19, Ste10, Ste99, Sti06, Sto15, Tel18, Tho21, Tra11, Vai05, Ver18, Wal12b, Wal03, Wet01, Wol13, Yor17b, Zha02, Zha23a, Bis12, Bla15, Dav12, tL16, Nic12]. Policy [Smi07, Cal06, Haa20, McP19, Sch19, Dom05, Fel05, Gar00a, Her04a, Kau03, Ser01, Yaq02]. Polish [JK07, Joh20b, Joh21, Pac06, Sel02, Sta07, Bri13, Dav04, Her03b, Jar16, Kaź09, Łuk23, Pip11, Szp07b]. Polistampa [Nut11]. Politburo [Mar06b, DF19]. politica [Eva09]. Political [Arg04, Čav19, Cea06, Dom15a, Dum17, Eck07a, Eng20b, Gra17, Hat17, Hei16, Her05, Imm16, Kar15b, Kar14b, Lef12, yL01b, Lib09, Loe10, Łuk23, Mat16a, Mor00b, Mor00c, Nin03, Rus04b, Rut06, Rya08b, Sba05, Sny09, Sön22, Tur01, Wil08a, Wel00]. Politician [Ryd14]. Politicization [Wal17b]. Politicka [Kar15a]. Politics [And05, Ano08i, AS22, Bra05, Cal07a, Cal07b, Cal22, Car17, Che20, Cla01, Cob18, Coh12, Con12, Cur15, Die01, Din18, Dob10b, Dro18, Dun03, Fie03, FAM16, FP21, Gas00, Gav99, Geo13, Göt08, GD99, Gue17, Hay03, Hel07, How14, Jac17, Jan08, Joh07c, Jol07, KW22, Kir08, Kno20, Kru05, Kus02, Lar02, Law07, Lem14, Lie12, Luk01, Luň10, Mac06, Mad02, Mad07, Mat16b, McC14, McG99, Mil22, Moe03, Nie11, Por08, Pra99, Qua14, Qua23, Rab13, Rad06, Ram01, Rea11, Rem08, Sch22b, Ser01, Sha01, Smi07, Sny03b, Sta13, Ste05, Str12, Tha06, Tho08, Tor22, Tum14, US21, Van99b, Whi02b, Whi03a, Whi12, Wid05, Won04, Xia06, Yan08, Yon08, You12b, You13b, Zar21, Zlo03, Bra16b, Die07, Lub11, Onu14, Sib11, Tsi08, Win15]. Politics [Elm01, Hix05, Ste02]. Politik [Nue07, Rug17, vD02]. politika [Boy23, Min19, Vaj20]. politiques [Mar16]. Politische [Huh20, Sch10, Wil08a]. Politischer [War06]. polityki [Kaź19]. Polk [Iat18]. Polk/Staktopoulos [Iat18]. Pollock [And10]. Polmar [Kas16]. polovina [Odo02]. Polska [Joh21]. Pols’ko [Sny06]. Pols’ko-ukrains’ke [Sny06]. Pomar [Shm23, Din23a]. Pomor [Gol12]. Pompidou [Van02b, Tra11]. Pons [Bro11, Pet11a, Van99a]. Pontecorvo [Hay16]. Pop [Sar05a]. Popadiuk [Aft13]. Pope [GP11]. Popular [Bai18, Cha07b, Hal07b, Kri04]. Popular/Unpopular [Cha07b]. Population [Bot17]. Populations [Myl15]. Port [Dal10]. Porter [Kim07]. Portrait [Her07a]. Portraits [Nor07]. Portrayal [For20]. Portugal [Mac06]. Portuguese [Mul23, Rei19]. Position [Ano08i, Cho10, Dei13, Göt08]. Poslednie [Bra16a]. Post [Ald07, Arg04, Bur04, Car13b, Gif05, Gra21b, Kra99b, Kra99a, Kra99c, Nic12, Sch02a, Sny11, Ste18, TSW^{+}22, Whi03a, Won04, Zak18]. Post-1945 [Arg04]. Post-1991 [Gra21b, TSW^{+}22]. Post-Cold [Gif05, Sny11]. Post-Communist [Sch02a]. Post-Mao [Won04]. Post-Soviet [Ste18]. Post-Stalin [Kra99b, Kra99a, Kra99c, Zak18]. Post-Vietnam [Nic12]. Post-War [Car13b, Whi03a, Ald07]. Post-World [Bur04]. Postage [Mod17]. Postcolonial [Ger14]. Posture [Lut22b]. Postwar [Bar20, Bra04, Bur23, Cap14, Fal05, Gas00, Gen17, Gun02, Joh14c, Kis06, Kle22, Lie21, Mac19, Mil15b, Moe03, Nai10a, Rob02b, Spe14, Spe15, War08]. potere [Nut10, Nut11]. Potomac [Joh11b]. Potsdam [vDO08, Rob07]. Poul [Ing17, Nie12]. Pound [PR21]. Poverty [Cob18]. POW [Wil08b]. Power [Čav19, CK13, Dav11b, Dei13, Edg08, Elm01, Eva00, Gav99, Hat11, Hel07, Hix05, Ian14, KG11, KW07, Kim07, Kob06, Kor07, Kra06f, Kri04, Kur15, Lie12, Luk01, Mar01, Njø02, Owe07, Qua23, Rab13, Ros09, Rus07, Sch03, Szp07a, Vai05, Wid05, Yaq02, Zhu20, Dre13, Gra13b, Jer13, Jer16a, Len12, Mat16a, Wil23, vC13, Bai18, Fin05, Jer00, Rus04a, Sba05, Sil04]. Powers [Hus05, McC20, Mil04b, Pra07]. POWs [Mat20a, Mil15a]. Poyntz [Lyn16]. Poznań [Hai16]. pp [Abe21, Aft13, Ale12, Ass14, Bal15, Bal17, Bar11a, Bar11b, Bar07, Bel11, Bis11, Bla15, Ble16, Blu15, Boy11, Bra14, Bra16b, Bre21, Bru11, Bur13, Byr12, Cal14, Car13a, Car13b, Car11, Chi14, Chu07a, Chu07b, Cli07, Col15, Cra17, Cra07, Cul13, Dan16a, Dar15, Dar16, DF12, DeM12, DeR17, Dei15, Dom15a, Dou15, Dra12b, Duk11, Edw16, Fer07, Fes07, Fin13, Gaf11, Gal16a, Gal07, Gal16b, Gan16, Gar13, Gáš13b, Gen17, Goo15, Goo17, Gra12b, Gra13b, Gru07, Gun12, Guz13, Haa13, Ham15, Har12b, Har11, Has07, Haw07, Hei16, Hel07, Her16, Her07b, Her11b, Hes21, His13, Hit16, Hoc16, Hou16, Hus13, Imm16, Jen17, Jer11, Jer13, Jer16a, Jer16b, Joh07b, Joh07c, Joh11b, Joh13a, Joh21, KG11, Kal12, Kan12a, Kan12b, Kan15, Kap17a]. pp [Ken13, Ket14, Key16, Kim07, Kim11, Kip15, Kle11, Kop16, Kre12b, Kre12a, Kro12, Kru07a, Kuk13, Kul13, Kur11, Kuz16, Lak11, Lat14, Lat16a, Lat16b, Law12, Lee12, Lee21, Lee17, Lew12, yL12, Lie21, Lie12, Lin11, tL16, Lit11a, Lit11b, Liv07, Lou16, Lub11, Luc07, Lüt12, Mal15, Mar12, Mar07d, Mar13b, Mar16, Mas11a, Mas21b, Mat16a, Mat16b, McM12a, Mey11, Mil17a, Mil15b, Mil17b, Mil21, Moi14, Mol16, Nam13, Nam16, Neh11, Nei14, Neu11, Nor07, Nor15, Nut11, Oli11, Onu14, Pai16, Par16, Pet11a, Pod14, Pos07, Pra17, Pre17, Pud12, Qua21, Rab13, Rad11, Rad16, Rai11, Rak15, Ram11a, Ram11b, Ran12, Ree16, Rei16, Rob14, Roh17, Rom17, Ros17, Rus16, Rut12a, Rut11, San16, Sar16, Sas14b, Sau12, Sch13, Sch11, Sco17, Sen16, Sha07a, Sha11]. pp [She12a, She12c, She13, Sil11, Sim16, Smi11, Sny11, Sol13, Sou21, Spe14, Spr07, Sta13, Sto11, SF11, Sur13, Sza21, Tei11, Thi14, Tho11, Tro16b, Usd11, Usd16, VA12, Vil16, Vuj02, Wal12b, Wea12, Wei12, Whi16, Whi11a, Whi11b, Whi12, Whi13, Whi14a, Whi15, Wie07, Win11, Wir07b, Wir13, Woo12, Woo13, Wys12, Yor17b, Zab16, Zar21, Zhu16, vD21]. pp. [Bel13, Bes07, Bis11, Bis12, Bis13, Bor12a, Bra16a, Bro11, Bru16, Bur16, Cea17, Chi14, Cli15a, Cog12, Coh12, Cra13, Cro07, Dan11, Dan12a, Dan12b, Dan14, Dan16b, Dan16c, Dav11a, Dav11b, Dav12, Dei07, Dei13, Dei14, Die07, Din07, Din13, Din14a, Din14b, Dom07, Dre13, Duk13, Eng07, Eva13, Fer15, Fre07, Gel12a, Geo13, Ger13, Gol11, Gol12, Gra13a, GP11, Han13, Hat11, Hay11, Hay13, Hay16, Her12, Hes12, Hes13, Hol15, Iri11, Jak16, Joh11a, Joh14b, Joh14a, Joh14c, Jon07, Jon14, Kal16, Kas16, Kir16, Kle16, Kot17, Kru07b, Kuk11, Lef12, Len12, Mad15, Mag14a, Mag14b, Man13, Mar07b, Mas11b, Mas12, Mas16, McA07, McM16, McM11, Mer13, Nai16, Nic12, Nie12, Owe07, Pec13a, Pét13, Pol13, Pon13, Pra99, Pre11b, Qua13]. pp. [Ram07, Rep12, Ric12, Rob14, Roz11, Rub11, Rug16b, Rug16c, Rus07, Sae07, Sib11, Smi07, Sny17, Soa13, Str12, Sza12a, Sza11, Taa13, Tal11, Tum16, Vei13, Wal12a, Wan16, WF12, Wet11a, Wet11b, You12b, You12a, You13b, Zha12b, vC13, vD12]. Pracownicy [Joh17]. Practice [Dav11a, Kuk11, Vai05]. Practicing [Diz10]. Prados [Lit11a, Low06]. Praeger [Fer07]. Pragmatism [Yaq02]. Prague [Dap14, Gáš13b, McG99, Wet10, Wil12]. Praying [Bur19]. Prazmowska [Cur07]. Preble [Pra08a]. Precan [Gáš13b]. Precedent [Spo12]. Precedent-Setting [Spo12]. Precision [Chu09]. Precluded [Spo12]. Predict [Leb03]. Preemptive [Sch09]. Preferences [CK13]. Prejudice [Joh07c]. Prelude [Bru03, Bur05, Jon19, Pol10]. Preludiu [Del07]. Prensa [Kel19]. Presence [Zim09]. Present [Duk13, Eng07, Gru07, Man13, Gly06, Gre10, Iri14, Kur14]. presenza [For02]. Presidency [Wal17a, Joh07d, Moe18]. President [Aft13, Mad04, Mad08b, McP19, Pav19, HGP^{+}23]. Presidential [Joh14c, Koc10, Mad02, HT19]. Presidents [Col15, Dom09]. Press [Abe21, Aft13, Ale12, Ass14, Bai18, Bal17, Bar11a, Bar11b, Bar07, Bel11, Bel13, Bes07, Bla15, Ble16, Boy11, Bra14, Bra16b, Bre21, BSK^{+}23, Bru11, Bru16, Bur16, Bur13, Byr12, Cal14, Car13a, Car13b, Car11, Chu07a, Cli15a, Coh12, Col15, Cra13, Dan16a, Dan14, DF12, Dav11a, Dav12, DeM12, DeR17, Dei07, Dei15, Die07, Din07, Din13, Din14a, Dom15a, Dom07, Dou15, Dra12b, Dre13, Duk13, Edw16, Fer15, Fes07, Fin13, Gaf11, Gal07, Gal16b, Gan08, Gan16, Gar13, Gen17, Geo13, Ger13, Gol11, Gol12, Goo15, Goo17, Gra12b, Gra12a, Gra13b, GP11, Gru07, Gun12, Guz13, Ham15, Har11, Has07, Hat11, Haw07, Hay11, Hay13, Hei16, Hel07, Her12, Her11b, Hes12, Hes21, Hes13, His13, Hit16, Hoc16, Hol15, Hor07, Hus13, Iri11]. Press [Jen17, Jer13, Jer16a, Joh07b, Joh13a, Joh14a, Joh14c, Joh20b, Jon07, Kal12, Kal16, Kan12b, Kan15, Kap17a, Kas16, Ken13, Ket14, Key16, Kim07, Kip15, Kle11, Kle16, Kop16, Kre12b, Kre12a, Kro12, Kuk11, Kuk13, Kur11, Kuz16, Lak11, Lat14, Lat16a, Lat16b, Law12, Lef12, Len12, yL12, Lie21, Lie12, tL16, Lit11b, Liv07, Lou16, Luc07, Lüt12, Mad15, Mal15, Man13, Mar07b, Mar12, Mar07d, Mas11a, Mas21b, Mat16a, Mat16b, McA07, McM12a, McM11, McM12b, Mic08, Mil15b, Mil17b, Mil21, Myl12, Nai16, Nam16, Neh11, Nie12, Nor07, Nor15, Oli11, Pai16, Par13, Par16, Pet11a, Pod14, Pon13, Pra17, Pra99, Pre11a, Pre11b, Pre17, Qua13, Qua21, Rab13, Rad16, Rak15, Ram07, Ram11b, Ran12, Rat21, Ree16, Rei16, Rep12, Roh17, Ros17, Roz11, Rub11]. Press [Rus07, Sae07, Sau12, Sch13, Sch11, Sen16, Sha07a, Sib11, Sim16, Smi07, Sny11, Sny17, Soa13, Sol13, Sou21, Spr07, Sto11, Str12, Sur13, Sza21, Tal11, Tei11, Thi14, Tro16b, Usd16, Vei13, Vuj02, Wal12b, Wan16, Wea12, WF12, Wei12, Whi16, Whi11a, Whi12, Whi13, Whi15, Wie07, Win11, Wir07b, Wir13, Woo12, Woo13, Wys12, You12b, Zar21, Zha12b, Zhu16, vC13]. Presses [SF11]. Pressures [Tum03]. Prestige [Wan16]. Preston [Cal09]. Prevented [Sch09]. Prevention [PR21]. pri [Gor01]. Price [Bro04, Kim11, Pac06, Rad12a, Ros17]. Priest [Ryd14]. Priests [Szp10]. Primacy [Iri14]. Prime [Whi14b]. Prince [Wan09]. Princes [Wes00]. Princeton [Ale12, Bra16b, Coh12, Mat16a, Pet11a, Pra99, Rus07, Sny11]. Principled [Hay15a]. Print [Del23, Yar09]. Priscilla [Lüt12, Ros08]. Prism [Haj12]. Private [Bai18, Hol05a, Kap20, SS14]. Privileged [Aft13]. Pro [Vei14b]. Pro-War [Vei14b]. Probes [Her03b]. Problem [Col00, Col07, Del22, Gol20, Lar02, Mur02, O’D09, Pru04, Tra12]. Problems [Coh12, Gil06]. Process [Dav10, Din04, Mas23, Bék16]. Proclaiming [Che10]. Proclamation [Mac06]. Procurement [Gho00]. Producing [Aro06]. Production [Nor05, Sch01]. Productions [Cor22, Rut12b]. Professional [Sap03]. Profile [Gra12b, Gra12a, Mor13]. Program [Bur19, Bur21, Cur23, Mar01, Pol10, Ran12, Oye14]. Programs [Kna03, SL23]. Progress [Lou23, All14, Dar12, FF22, Kuz16, Alk05]. Project [Wal21b, Kas16, Bar08, Pan01, Zie21]. Proletarian [RTYD19]. Proliferation [Bud22, Bur17, Kno20, Que02]. Prometheus [Ble16, Bor12a]. Prominence [Soa23]. Promised [Hor07]. Promoted [Hay03]. Promoting [Sau12]. Promotion [McC14, PL22, Pee23, Sim14, Søn19]. promyshlennyi [Odo02]. Propaganda [Bel13, Bus14, Che20, Cor22, Din02, Gue17, Hei15, Kaź09, Kna02, Kre12b, Kre12a, Kru08, Mar15, Mod17, Moe03, Pik02, Rid16, Sha02, Sha08a, SY17, Sta00, Whi02a, Yar09]. Property [Ran12]. Proposal [Tal08]. Proposals [HTP^{+}15, Kou12, Neh22]. Prosl [Rug16b]. Prospekt [Sha12]. Protection [Kom18]. Protest [Fin05, Zar21, Tro16b]. Protestants [Moe07]. Protests [Wal21a]. Protracted [Mic17]. protystoyannya [Sny06]. Provinz [Bis11]. Provoking [Sho15]. Proxy [Gle06a, Gle06b]. Prozess [Rug18]. Przyczynek [Szp09]. Psychological [Ber04b, Bus14, Eth05]. Public [BD22, Cly13, Con03, Cri04, Diz10, Gan08, Haj12, Hal07b, Kru07a, Mor13, Ric12, Sch17, Shl06, Sny13, Whi04]. PublicAffairs [Sza11]. Publications [Mar13b]. Publishers [Chi14, Kru07a, Lin11, Lit11a, Mag14b, Nam13, Pét13, Ric12, Sil11]. Publishing [Cli07, Dan16c, Eng07, Lee12, Lin11, Pec13a, Rad12a]. Pueblo [Her04a, Ler06, Hou15]. Pugach [MC23]. Pugwash [Jos20, Bar18, Kra18, KNS18, Kur18, Olš18, Röh18, Sac18]. puissances [SF11]. Pulp [Pit23]. Punishment [SM12, Clo09]. Punjab [Gan08, Yon08]. Punk [Sze19]. Purpose [Leb06]. Pursuing [Dom09]. Pursuit [Bre08, Kno06]. Put [Gar04b]. Putin [Rut06]. Putnam [Gra13a].

Qaeda [Lei08]. Qingdao [ZP23]. Quaddafi [Ale04]. Quagmire [Law22]. Queen [Hay06, Ken13]. Queens [Mas11a]. Quellen [vD05, Rug16c]. Quenby [Pud12]. Quest [ARN16, Bur23, Chi14, Cog02a, Fer11, Nie16b, Ric09, Rub11, Wal21b]. Question [Bék16, Cly01, CT03a, Har19, Hit03, Jar16, Kou17, Mil04a, She03, Sny99, Nie16a, Spo12]. Questionable [Ber13]. questione [Gor09]. Quiet [Lon11]. Quinn [Yor19a]. Qureshi [Gus10].

R [Alp09, Bai18, Ces21, Cri20a, Dar15, Dar16, Dav07, Eas10, Far15, GP11, Her20a, Hoc16, Joh14b, Joh22, Lar07a, Law10, Loe10, Mac19, McK15a, Mic17, Mil17b, Nam13, Rog15, Sha08a, Shm23, Suh08, Wea12, Wil08b, vD21]. R. [Kle11, Mil06, Ped07]. Rabe [Ram07]. Race [Bel02, Dan12a, DeR17, Dou06, Haf08a, Kal22, McM12a, Mit07, Pra15a, Pra15b, Ros04, Ros05, She12c, She13, Tod17, Wan16, Neu11, Smi11, Woo13, Ros00]. Racial [How14, Kre12b, Kre12a]. Racing [Gal07]. Racism [Bur19, Ing04b, Sur13]. Radar [Shm23]. Radchenko [Kan12a, Mal15, SZK12]. Radia [Joh17, Joh21]. Radical [Vu19]. Radio [Byr12, Cri02, Cri10, Cul22, Din23a, Fal05, HF14, Hol08, Joh21, Shm23, Byr12, Haa13, Pos19]. Radu [Pet08]. RAF [Cra07]. Rafał [Joh21]. Raghavan [Gan16]. Raghuvendra [Gan08]. Raid [PR21]. Raiders [Lea01]. Rain [Oye14]. Rainer [Jan08, Luň10, Nie10a, Sto07]. Raising [Kas16]. Rajak [Pet18, Yor20a]. Rakove [God14]. Rallying [Haa13]. Randall [Hes21, Rut12a, Mat16a]. Range [Wei01]. Rank [Mil15a]. Ransom [Pet08]. Raoul [Mat12]. Raphael [Git20]. Rapprochement [Lin11, Xia06]. Rasmus [Ing17]. rasshireniyu [Min19]. Rauchensteiner [Bis11]. Raúl [Iac22]. Rauschenberg [Kac02]. Raymond [Bes07, Mit07]. Re [Mar09, Cha04]. Re-examining [Cha04]. Reaching [Har18]. Reaction [dL13]. Reactions [Ber20, Haj12, HT19, Kou12, yL09]. Reader [Gaf07]. Readers [Bar20, Zub14]. Readiness [Sou21]. Reading [DX18, Lee12, Bac02].

Readings [Ran06]. Reagan [Ale04, Bro22, Dan16c, Kau02, Kau09, Søn19, Bar22, Gaf11, Kau02, Mat10b, McB15, McD02, Mil22, PL22, PR21, Sim23, Dan15]. Real [Eck07b, Gei12, Lud13, Rad15]. Realignment [Tag22]. Realism [All15]. Realist [Dav12, Hof06]. Realities [yL12, Cly01]. Reality [Haj12, Szo10]. Really [Bor05b]. Reappraisal [Mak12]. Reappraising [GRS^{+}12]. Rearguard [She03]. Rearmament [Sny13]. Rearranging [Lüt14]. Reason [McM11, Tof05]. Reassess [BL10a, BL10b]. Reassessing [JKM12, Kra13a, Mil13, Pre17, TGS^{+}23]. Reassessment [Fet20, Per07, Tra08]. Rebecca [Whi14b]. Rebellion [DeM12, Gor05, Nie10a, Rem08, Che06, O’M21]. Rebels [Kal14, Ram01, Ram11a]. Reborn [Ged99]. Reception [Ric22, Tol10]. Reclaiming [Szp07b]. Recognition [Mar08c, Mue11]. Recognized [Lüt14]. Recollections [JK07]. Reconciliation [Ste10, Tof05, Yan08, Ass15]. Reconnaissance [Pra20]. Reconsidered [And20, Joh20a, Man21, Mue05]. Reconsidering [Myl15, Bit17]. Reconstructing [Sfi02, Van00, Guz13]. Reconstruction [Cal22, Cap14, Low04, Lee12, Wen09]. Record [Bar19, Cri04, Nie21]. Records [Bur04]. Recover [Sew23]. Red [Ale08, DeM12, Etc23, Hay15a, Hei15, Her12, Hol08, Lew07, Pra17, Rub06, Rut12b, Tol03, Wan09, Ble16, Hay06, Eng20b, Hay15b, Mar01, Zal04]. Redemption [Ale12]. Redrawing [Ano03l, Gat03, Ken03, Kus03, Lef03, Sny03a]. Reed [Her07b]. Reenacting [Tud05]. Rees [Mar06b, Nai10a]. Reevaluating [Dom11, PL22]. Reexamining [HSB^{+}23, OB20]. Reflections [CKP05a, Her13, Mil02, Bar04]. Reform [Eng07, Gil05, Kim16, Ree16, Tro16b, Zin11]. Reforms [Con03]. Refugee [Hur10, Oye14, Tod17]. Refugees [Jac16, Kar14b, Oye14, vD21]. Regarding [Pos07]. Regime [Bor09, Dom09, GRS^{+}12, Mat15, Mue17, PL22, Rad07, US21, Che07, Sil11]. Regina [Fre07, Lub11]. Region [Joh11a]. Regional [BJ10, Gel12a, Kan12a, Kou17, Qua09, Sto15, Tho21, vW19]. Regionalism [Sak06]. Regulska [Spe14, Spe15]. Rehschuh [Kur21]. Reich [Neu11, War06]. Reimagining [Sae07]. Reins [Ros09]. Reinterpretation [CS07, RG15, Ter05]. Reinterpreting [Hal16, Mar05]. Reiss [Cra21]. Rekonstruktion [Bis08]. Relations [Ado08, Alk05, And20, Bae16, Bur16, Cal09, Che06, Che07, Cly13, Coh06, Dar12, Dei02, Dei13, Dei14, Din18, Eas10, Ess15, Eva07, Gas06, Gol11, Goo05, Haf08b, HS15, Hal16, Hel07, Hel22, Her20b, Hop03, Ing04b, Iri14, Joh07a, Joh05, Joh20a, Kal18, Kap14, Kir16, Kli02, Kre07, Lib01, Lud13, Lüt14, Mar07a, Mat10a, McK15a, Mun10, Mur02, Nie12, Nun22, Ole11, Pai17, Ped19, Per04, Pra21, Qua21, QMLL16, Rad07, Ric09, Ros04, Roz03, Roz11, Sel02, SX14, Sim17, Smi03, Ste14, Sza14, SJ23, Tho11, Wei04a, Wid05, Win17, Woh99, Wys12, Xia08c, Yor17a, Zha06, Zha23b, Zhi00, Cli07, Kal12, Sny17, Den21, She07, Sza10b, Kur05a, Wel00]. Relationship [Bro08, Dei02, Sac18, You07]. Relationships [Fal20].

Relentless [Mag14a]. Reliance [Vám22]. Religion [Bra05, Sha02, Spe13, Szp10, Wal13b, Yan08, Whi06]. Religious [Her12, Shl06, Wal13a, Wal13b]. Remacle [Cog12]. Remained [Buh15]. Remaining [KM22]. Remaking [Bur16, Gos08]. Remembrance [Pol13, Tum14]. Remorse [Nei07]. Removal [JK07]. Renata [Rat21]. Rendiconto [Dra21]. Reno [Sou21]. Renouard [Jen17]. Reorganization [Hus05, Mil04b]. Repatriation [Shi13]. Repercussions [Kra03a, Kra04b, Kra05a]. Repetition [Eng02]. Replacing [Lan10]. Reply [BL10b, Gad13, Gre11, Luk16, Lüt10, Nai12, NLM12, Rad12b]. Reporting [Dan14, Gan08]. Reports [yL09, Yon08]. Repousis [Kap14]. Representation [Tod17]. Representing [Mac20a]. Repression [DS14, GRS^{+}12, Nag17, Nag18, Nat04, Tro16b]. Reprisals [Bal05]. Reprise [Eng02]. Republic [Boz20a, Boz20b, Neh22, Nun22, Pik02, Sha11, Spr17, Sza12a, Wet09, Lar10, Win17, Dal10, Mic21, Bru04, Fre01]. Républicaine [Foo21]. Republican [You12b, Mar23b]. Republicanism [Foo21]. Republics [Ber20, Fet20]. Republikflucht [Wet09]. Rescue [Che18]. Research [BD22, Bur04, Kuk21, Mor00a, Sha08a, Stu15, Ula99, Roh17]. Resentment [Gor05]. Reserve [Eng09]. Resettlement [Sny09]. Reshaping [All14]. Resistance [Ass14, Die08, Kra06f, Kri04, Pif15, vDG07, Kna03, McG06, Gor05]. Resisters [Joh03]. Resisting [Rem08]. Resolution [Tof05]. Resolve [Sny99]. Resolving [Adl03]. Resonances [Sch08]. Resonanzen [Sch08]. Respite [Bra02a, Van02a]. Respond [Bur21]. Response [Cog03, Hit03, Kee00, MS08, McG99, Oye14, Ray20, Whi11b]. Responses [BM12, Ser13, Sza21]. Responsibility [Aft02, Ben21, GRS^{+}12, Sch22b]. Restitution [Mag10, Sny03b]. Restoration [Con07]. Restored [Cog02a]. Restraints [Car17]. Restricted [Jer22b]. Restructuring [McP19]. Resurgent [Kom14]. Retail [Rut12a]. Retaliation [Jon08]. Rethinking [Bra06, Dar12, Eng03, GD99, Hof06, CKP05b, Sch03, Sha06]. Retreat [Joh11a, Rua04]. Retribution [Kor10]. Retrospect [Dra06, Iat18]. Return [Mue11, Lat14, Bar03a]. Returned [Mag10]. Returnees [Ree16]. Reunification [Boz15]. Reveal [Kau02]. Revealed [Usd09]. Reverse [Tie07]. Review [And23, Att23, Boy23, Bro22, Bur23, Con12, Con23, Cul22, Del23, Din23a, Din23b, Eck07b, Etc23, Eva23, Gil23, Har23, How22, Jac23, Joh22, Kal22, Kle23, Lar22, Lee23, Lüt22a, Mar23a, Mis23a, Mis23b, Moi03, Moo23, Mul23, MC23, Neu03, Neu23, Pap23, Pee23, Pit23, Qua23, Rey23, Sch22b, Sew23, Shm23, Soa22, Soa23, Sul22, Wil12, Xia08b, Xia23, Yor23a, Abe21, Ada03, Adl03, Aft02, Aft03, Aft20, Aki02, Ale04, Ale01, Alk05, Alo05, Arg03, Arg04, Ass15, Ava04, Bac02, Bae02, Bal05, Ban14, Bar03a, Bar04, Bar03b, Bay01, Bel02, Ben21, Ber15, Ber04a, Ber01, Ber04b, Bis04, Bis14, Bla05, Boo03, Bor05a, Bos04, Bos15, Bra04, Bra05, Bra02a, Bra02b, Bre08, Bre21, Bri15, Bro02a, Bro06, Bro04, Bru01, Bru04]. review [Bru05, Bug15, Buh15, Bur04, Byr01, Cal02, Cal03, Cam05, Cap14, Cat05, Cat08, Ces21, Cha04, Cha14, Cha00, Cha21, Che99, Che19, Chi04, Cla01, Cle04, Cli15b, Cly01, Cog02b, Cog02a, Coh02, Col05, Coo04, Cou04, Cox02, Cra21, Cra03, Cri02, Cro00, Cro15, Dal00, Dan15, Dan22, Dap14, Dei02, Den21, Der02, Die01, Din02, Din04, Div00, Div01, Doc04, Doc05, Dom05, Dom20, Dom03, Dom15b, Dom15c, Don05, Dou06, Dra21, Dre05, Dre21, Dre15, Duk02, Dum20, Eck14, Eck19, Ell03, Elm01, Elm04, Eng20b, Eng02, Eth05, Eva00, Fal05, Far15, Fel05, Fie03, Fin04, Fin05, Fis03, Foo21, For19, For02, Fou14, Fre01, Fri02, Fri04, Fri05, Fri10, Gal15, Gan06, Gar00a, Gar01a, Gar02, Gar04a, Gar04b, Gar01b, Gar00b, Gas06, Gas00]. review [Gat03, Gav05, Ged99, Gen04, Ger14, Gif05, Gil05, Git20, Gle05, God14, Gol03, Gor03, Gor05, Gra05, Gre05b, Gre05a, Gre01, Gre08, Gro02, Haa05, Haa20, Hac03, Har15, Hay99, Hay03, Hay15b, Hay15a, Hei06, Her03a, Her02, Her04a, Her99, Her19, Her05, Hes02, Hes21, Hez02, Hix05, Hof20, Hol05b, Hol21, How14, Hur05, Hus00, Hus05, Hut01, Hut19, Iac22, Ing04a, Iri02, Iri14, Jen03, Jer00, Jer08, Joh03, Joh05, Joh20a, Joh21, Jon03, Jon04, Kag05, Kal14, Kan14, Kap04, Kap14, Kap10, Kar15a, Kar15b, Kar14a, Kar04, Kau02, Kau03, Kaź19, Kel02, KW07, Ken05, Ken03, Kip03, Kle04, Kle22, Kli02, Kli05, Kna05, Kna02, Kor07, Kos15, Kra01, Kra04a, Kra06a, Kri04, Kri05, Kru04a, Kru04b, Kru05, Kuk21, Kur05a, Kur14]. review [Kur05b, Kur15, Kur21, Kus03, Kuz19, Lar02, Las19, Lau00, Law01, Law05, Law11, Lea01, Leb00, Leb03, Leb06, Lee21, Lee06b, Lef03, Lef21, Lem14, Ler02, Ler15, Lev04, Lev10b, yL01a, Lib01, Lie21, Lin15, Loc05, Low04, Luc02, Lun04, Mac19, Mad02, Mad04, Mad08a, Mar06a, Mar01, Mas19a, Mas00, Mas19b, Mas21a, Mas21b, Mat04a, Mat04b, McA14, McE03, McK15a, McM05, McM14, McM21a, McM21b, McP19, Mic21, Mid02, Mil02, Mil01, Mil14, Mil15a, Mil21, Mil04b, Mil04a, Min19, Mit04, Moe03, Moï11, Moi15, Mou02, Mou03, Mue05, Mun21, Mur05, Mur04, Mur02, Myl15, Nai02, Nai04, Nai21, Nai22, Nat04, Nau04, Nei00, Nie03, Nin03, Nor00, Nor05, Nue05, Nut02a, Odo02, Osg04, Pac06, Pai14, PW05, Pan01, Par19, Pas03, Pav19, Pay21]. review [Per04, Pfa02, Pik02, Pip15, Poc03, Pol05, Pon01a, Pow01, Pra14, Pra15a, Pra15b, Pra20, Pra19, Pru04, Qua04, Qua14, Qua21, Que02, Rad14, Rad15, Ram01, Ran02, Rat21, Rep03, Rey02, Ric15, Ric04, Rie15, Ris15, Rob10, Rog15, Ros00, Ros04, Ros05, Rot21, Rua04, Rug18, Rus04a, Rus05, Rus04b, Ryd14, Sai02, Sap03, Sar05b, Sav14, Sba05, Sch00a, Sch02a, Sch19, Sch00b, Sch02b, Sch03, Sch01, Sch05, Sch02c, Sel02, Sep01, Ser01, Sfi02, Sha20, Sha06, She14, Sil05, Sil99, Sil04, Sim04, Sim14, Smi03, Sny03a, Sny03b, Soa20a, Soa20b, Sok99, Søn19, Son02, Sou21, Spa14, Spe18, Spe01, Spe15, Sta00, Ste02, Ste20, Ste99, Ste05, Ste15, Sti10, Stu15, Stu19, Sza21, Sza14, Sze19, Szp07a, Tal99, Tau02, Tay00]. review [Tay19, Ter05, The15, Tis15, Tof05, Tol03, Tom04, Tra02, Tum14, TJ21, Tur01, Vai05, Val14, Vám10, Van02a, Van02b, Vau15, Vei14a, Vei14b, Vit04, Von03, Vuj02, Wal21a, Wal21b, War06, War01, Wea15, Wei02, Wei04a, Wei04b, Wel00, Wel06, Wem21, Wes00, Wet01, Whi02a, Whi02b, Whi03a, Whi04, Whi06, Whi14b, Whi03b, Wid05, Wil02, Wil03, Win15, Wir03, Wir06, Wis00, Won04, Wym03, Xia14, Yaq02, Yor17a, Yor19a, Yor20a, Yos15, Zal04, Zar21, Zha02, Zha15, Zhu20, Zie21, Zub14, vD02, vD03, vD05, vD21]. Reviewing [Woh99]. Reviews [Ano01d, Ano01e, Mas23, Roz03, Yor15, Yor21]. Revisionism [Ass11, Bis05, Mor12]. Revisionist [Zin11]. Revisionists [Ham15]. Revisions [Git20]. Revisited [Dav10, Mel01, Rad17, You22]. Revisiting [Get15, Mar08c, Pap23, You13a]. Revival [Eth05]. Revived [Szp14]. Revolt [Dap14, Gar04b]. Revolution [Ass23, Bal05, Bru04, Cap08, DF07, Edw16, Fin05, FF22, Ged99, Gif05, Gra07, Iat05, Kel19, Kos15, Kur05b, Law12, Lee06a, Lit11b, Mag14a, Man13, Nei00, Ric22, Sil05, Sil99, Sni22, Son02, Tel18, Vu19, Wem21, WLX^{+}08, Yor18, Guo10, Lüt08, Nag17, Nag18, RTYD19, Sim16, WLX^{+}08, Nei08, Rat21, Wem21, Arg03, Kip03, Sch03]. Revolutionaries [Eck07a, Gil18]. Revolutionary [And05, Fal20, Har11, Hol06, Kau02, Kru05, Lov22, Sch22a, Tha06, Yor21]. Révolutionnaire [CS07]. Revolutions [Rad12a, Rad14, Wil17, Nei10b]. Reykjavik [Dan15, Sim23]. Reynolds [Has07]. RFE [Kom18]. RFE/RL [Kom18]. Rhetoric [BD22, McC20, Rin12, Sil04, Tud05]. Rhetorical [Che08b]. Rhodesia [How22]. Rhodri [Low06]. Ri [Den06]. Richard [Cul22, Gal15, Kal22, Mas23, McM14, Nut10, Nut11, Rob20, Rub11, Cri10, Dei14, Gro07, Haa13, Har11, Hol06, Kip15, Kre12b, Kre12a, Ler06]. Richmond [Diz10]. Rider [Soa20a]. Rieber [Nai16, Mar23a]. Rienner [Kru07a, Lin11, Sil11]. Rift [Kho10]. Right [Cro09, Gil18, Man13, Mar23b]. Right-Wing [Cro09]. Rights [Bel02, Boe21, Bri13, Cra03, Dra09, Fer11, Haw07, Hor07, Jen17, KW22, Mar21, Mas23, Nei10b, Nei14, Pee23, Ros05, Sch19, Soa06, Tho05]. Rimmington [Sha20]. Rina [San16]. Ring [Kho10, Usd09]. Riot [Wal12a]. Riots [tL16]. Ripple [Gra21a]. Risch [Sol13]. Rise [Bai18, Bri13, Čav19, Dei02, Dom05, Eng09, Eng02, Fin05, Gra13b, Joh11b, Kul10, Par13, Pud12, Rad11, Rus04a, Sha18, Wat17, Wei02, Wis00, vC13]. Rising [Cie07, Pol06]. Risk [Col07]. Rittersporn [Shl06]. Rivalries [Kra06f, Rus04b]. Rivalry [Sin18, Smi11, Zha12b, Elm04]. Rivals [Luk17]. rivoluzione [Gra07]. RL [Kom18]. Road [Cea09, Jac17, Kac02, Kim07, Mou02, Sch00b, Sto11, Vei14a, Wil00, ŽČ16, Mis23b]. Robert [Aft03, Ale08, Coh06, Eng09, God14, Kru07b, Pet11a, Søn19, Tsi08, Win15, Xia08c, Bar07, Boy11, Cli07, Den06, Ell09, Gaf11, Gru07, Ham15, Har08, Her17, Kac02, Law08, Lit11b, Mav18, Owe07, Pra17, Sie06, Tro16b, Wei12, Whi12, Wir07b]. Roberta [Abe21]. Roberts [Lüt12, Pec08, Ros08, Rie15]. Robin [Abe21, Bra16a, Rut12b, Ken07]. Rochester [Sim16]. ROCI [Kac02]. Rock [Ram01, Ram11a]. Rockefeller [Mue13]. Rockland [Dan16c]. Rockwell [Chu19]. Rodden [Hol10]. Rodopi [Taa13]. Rodric [Gra12b, Gra12a]. Roger [Qua13, Bru09, Vár09]. Roham [Key16]. Rohde [Kuk21]. Ro’i [Gol10]. Roithner [Kar14a]. Roland [Sni22]. Role [Cal03, Doc08, GD99, Her14, Hur05, Kan10, Pie16, Pol07b, TW05, Wei01]. Rolf [Shl06, Vil16, Bon07]. Roll [Ram11a]. Rollback [Bor99, McA22, SS14]. Rolle [Kan10]. Roman [Aft13, Leu13]. Romance [Laz07]. Romania [Fis03, Mic21, Pet08, Wei04a, Bot17, Dra12a, Mun12, Sta16]. Romanian [Sni22, Wei04a, Mun10]. Romanov [Dan07]. Rombach [Mas16]. romeno [Gor09]. Romero [Mas11b]. Ron [Abe21, Wil08b]. Ronald [Gaf07, Kau02, Cea09, McB15]. Rooms [Stu19]. Roosevelt [Bar03b, Pay21, Sim14, vDO08]. Roots [Gan07, Nin03, RTYD19]. Rosa [Nei14]. Rosato [Lie12]. Rose [Kle11, Soa23]. Rosenberg [Hor07, Usd09]. Rosenbergs [Usd16]. Ross [Byr12, Coh06, Cri20a, Xia08c, Eva07]. rossa [Gor09]. Rossi [Dra12b]. Rossii [Kar15c]. rossiiskogermanskikh [Mue10]. Rotarmisten [Nai10b]. Rote [Bis10]. Roth [Moo23]. Rouge [Dan14, Dei15, Zha15, For19]. Routledge [Bis12, Blu15, Chi14, Dar15, Dar16, Dre13, Eva13, Jak16, Jer16b, Jon14, KG11, Kan12a, Kot17, Kul13, Lew12, Mas12, Nei14, Onu14, Rob14, San16, Sar16, Sas14b, Spe14, Wal12a, Whi13, You12a]. Rovner [Geo13]. Rowe [Tsi08]. Rowland [Gaf11, Kol15]. Rowman [Kir16, Moy11, Nic12, Sco17, vD21]. Roy [Wal17a]. Royal [Tho11]. Rozgłósnia [Joh21]. rr [Sny06]. Ruane [Pra21]. Rubel [Bur10]. Rubenstein [Len18]. Rückkehr [Mer13]. Rudolf [Her08]. Ruggenthaler [Yor17b]. Ruinous [Ste18]. Rule [Dra09, Wet10, Kes07]. Rules [Die07]. Ruling [McD06]. Run [Mad02]. Running [Koc10]. Ruotsila [Rob14, Wal17b]. Rural [Gaw21, Wem21]. Rusi [Mas19a]. Rusk [Bra02b]. Russia [Dan16c, Dan22, Gel12a, Ger13, Hal07b, Hes13, Kar15c, Kur14, Kur05b, Mag10, Mar01, Mic21, Roz11, Rut06, Shl09, Tum16, Bal17, Gra21b, Lai08, TSW^{+}22, And05, Sav09, Zhu20, Dal00, Ged99, Her02, Kri05, Sch02a, Sch01, Wym03]. Russian [Bae02, Din23a, Eng02, Far15, Gle05, Laz07, Lin15, McM12b, Nei00, Pol06, Roz03, Rub10, Rut10, Sak06, Shm23, Von03, Wir13, Tho08, Tro16a, Wil12, Per04, Rep03]. Russians [Gra12b, Gra12a]. russischen [Mue10]. Russlandheimkehrer [Bis13]. Rust [Che19, Yor23a]. Rusya [Val14]. Ruth [Bur23, His13]. Rutherford [Pol10]. Ruzvel’t [Rob08]. Ryan [SS18b].

S [And05, Bos15, Ces21, Clo09, Coh06, Dan14, Gar10b, Gly06, Ham15, Hay06, Hus13, Kar15c, Kle11, Mas21b, Mil15a, Nor10, Sch13, Xia08c, Gáš13b, Wil12]. S. [Kim11, Luc07]. S.J [GP11]. Sabrina [Sch22b]. SAC [Pet11b]. Sachse [Jos20]. Saehrendt [Kan10]. Safe [Lit11a, Ces21]. Safeguarding [Din13]. Safehaven [Bre08]. Safety [Ber15]. Saga [KH23, Dan11]. Saigon [Che18]. Saint [Hei16]. Sakharov [Boe21, Mar21, McM11]. Salim [Qua21, Joh07b]. SALT [Wet17, Jer22a]. SALT-Verhandlungen [Wet17]. Salvador [HT19]. Salvatore [Dra08]. Samantha [McM16]. Samuel [Hoc16, Wal08, Pra20, Ste18]. Sánchez [Yor21, Eng20b]. Sanchez-Sibony [Eng20b]. Sanctuary [Boz20a, Boz20b]. Sanders [Man21]. Sandinista [Sto14]. Sandra [Vei14b]. Sands [Yor15]. Sang [Sha11]. Sankara [Sch22a]. Santérus [Wys12]. Sara [MC23]. Sarah [Her12, Sch19, Pra99, Roh17, Van99b]. Sarantakes [Soa13]. Sargent [Bur16, Hus13]. Sarotte [Rus16, Sny11]. Sascha [Bru09]. Satchmo [Tei11].

Satellite [Dav09, Pra17]. Satta [Dra18]. Satter [Ger13]. Saturday [Coo04]. Saudi [Mat20b]. Savage [Rus04a]. Savantes [Mar16]. Save [Mat10b, Pac06, Vei13]. Saved [Cle00]. Saving [Tel18]. Savodnik [Gre20]. Savranskaya [Pod14]. SB [Szp09, Szp09]. Sbardellati [Whi13]. Sbornik [Kar15c]. SBZ [Wet09]. SBZ/DDR [Wet09]. Scale [Usd09]. Scanlon [Vei14b]. Scare [Lew07, Mil20, Str09]. Scarlett [McM16]. Scars [Sch22b]. Scenes [Haf08a, Mic17]. Schaffer [Sza21, Wie07]. Schandler [Moy11]. Scharloth [Dap14]. Schatten [Rug17]. Schenk [McK15a]. Scherstjanoi [Bis13, Nai10b]. Schilling [Her20a]. Schlachtfeld [Zab16]. Schlaffer [Her08]. Schlesinger [Bur05]. Schmidli [Pee23]. Schmidt [Haf08b, Yor15]. Schmitz [Cro09, Nic12]. Schoenberg [Lie21]. Scholars [Gaf07]. Scholarship [Log01, Tor22]. Scholtyseck [Bru06]. School [Dro18, Rya08b]. Schoolboy [Joh07a]. Schools [Vai05]. Schoolteachers [Bru16]. Schors [Wet17]. Schrafstetter [Fer07]. Schröder [Nai10b]. Schulmann [Tsi08]. Schulzinger [Cli07]. Schuster [Cra17, Jer11, Kru07b]. Schwartz [Kip15]. Schwarz [Kle22]. Schweller [Mat16a]. Science [Ald08, And10, AS22, Gil23, Hal07b, Her03a, Jos20, Kru04a, Low04, Nut02a, Sön22, Whi03b, Woh99, Hal07b, Jos20, Sch20, Rus04b, Sap03]. Sciences [Mue13]. Scientist [Kur18, Her19, Aft02]. Scientists [Bor12a, Bra14, Her03a, Kra18, KNS18, Roh17, Kru04b, Low04, McM21b]. sconvolsero [Rom17]. Scores [Wen09]. Scotland [Chu07a, Chu07a]. Scott [Gal16b, Kle11, Ric12, Sha07b, Zie21, Fou14]. Scott-Smith [Sha07b, Fou14]. Scowcroft [Nic12]. Screen [Bre21]. Screening [Ver18]. Scrutinising [Mur04]. Sea [Dan07, Fri02, Smi08]. Sean [Col15, Mar20]. Search [Aft20, Bur05, For20, Gar00a, Lub11, McA07, McA08, Mue17]. Searching [Cul13]. Seas [Fri02]. Sebag [Ale08]. Sebastian [Lie12]. Sechi [Dra08]. Second [Bac09, Bar03b, Boy23, Din14b, Dou15, Kip03, Lat14, Mac20a, Mar23a, Pec13a, San16, Sta13, Wei04b, Yor23a, You12b, Mil20, Mil22, Tho21, Ris15]. Secrecy [Jer22b, Lin11]. Secret [Aft13, Bal15, Bal05, Bor12b, Bra04, Bru04, Bur21, Cor14, Dan14, Duk11, Eng09, Goo05, Gre05b, GP11, Has07, Kru08, Kru07b, Lee21, Lit11a, Low06, Łuk22, Mat10b, Pac06, Pet08, Poc03, Pra14, Rad12a, Sar05b, Sch09, Sha20, Sta00, Sta07, Szo10, Szp10, Usd11, Wan09, Wir07a, Żuk19, Cha21]. Secrets [Bru11, Hay11, Joh07d]. Securita [Sni22]. Security [Adl03, Aft03, ARN16, Bal08, Bra06, Bur04, Coh08a, Dan16a, Dav10, DS14, Din13, Geo13, Hez02, Jer01, Ken05, Kho14, Luc07, Łuk22, Mar07c, Mas23, McC20, Mei15, Mid11, Mil02, Pol07a, Ric09, Sas14a, Sni22, Søn19, Sto15, Szp07b, Tal99, Wys12, Neh21, Sel19, Sel21b, Ing17]. SED [Bru09]. SED-Staat [Bru09]. Seduced [Bru11]. Seed [Sha08b]. Seeing [Cra13, Tol03]. Seen [Haw07]. Segregation [Lew07]. sei [Rom17]. seit [Rug16b]. sekundäre [Huh20]. Selbstinszenierung [Shl06]. Selected [Kag05]. Self [Hez02, Iat05, Mil04a, Mou02, Sch00b, Vám22]. Self-Defense [Iat05]. Self-Destruction [Mou02, Sch00b]. Self-determination [Hez02].

Self-Esteem [Mil04a]. Self-Reliance [Vám22]. Sell [Cra21, Sim17]. Selling [Cal22, Yar09]. Selma [Sch19]. Semitism [Fei10, AE02]. Senator [Joh14b, Whi11b, Hol99]. Sense [Kip03, Moï10, Neu03]. senza [Dra18]. Separate [Jac17, Kir16]. Ser [Yos15]. Sergey [Kan12a, Mal15, Kan12b, SZK12]. Sergo [Pod14]. Serial [Mil15a]. Series [Lor17, Sar05a]. Servant [Her23]. Serve [Lou23]. Service [Ale08, Bro10, Dou15, Ell09, Pet11a, Usd11, Vu19]. Servicemen [Wir13]. Services [Pif15, SM12, Luk07]. Set [Her12, Luk17]. Setting [Dar12, Spo12]. Settlement [Col07, HTP^{+}15]. Settling [Wen09]. Seungsook [Duk13]. Seven [Mat08]. Sex [Ran10, Ram11a]. Sexuality [Dum14]. Seydlitz [Die08]. Seydlitz-Kurzbach [Die08]. Seymour [Dan10]. sfida [Eva09]. Shadow [Aft02, Dom15b, Elm01, Har15, Hol08, Kre07, Łuk23, Mil01, Pol13, Rad06, Sin18, Sza12b, Hut20]. Shah [Key16, Joh07a]. Shake [Mat04a]. Shalikashvili [Nor20]. Shame [Åse16]. Shan [Rub11]. Shape [Dad19]. Shaped [Sib11]. Shaping [Ale12, Bal15, Dav12, Haj12, Wal08]. Shareen [Par10]. Sharp [Neh21, Sew23]. Shattered [Nag17, Nag18]. Shaul [Din18]. Shaw [Cea06, Whi08b]. Sheds [Rad12a]. Sheiko [Tum16]. Sheila [Bra16b]. Shelter [McE03]. Shelters [Gei12, Mar08b]. Shen [yL07]. Shepard [Gos08, Lun04]. Shevtsova [Rut06]. Shibusawa [Sae07]. Shield [Bae02, Fri05]. Shifting [Mar22]. Shih [Rub11]. Shih-Shan [Rub11]. Shin [Cli15a]. Ship [Åse16, Her04a]. Ships [Pol07b]. Shlaim [Qua13]. Shock [Hus13, She14]. shook [Rom17]. Shooters [Sni22]. Shore [Chi14, Laz07]. Show [Shl06]. Showcasing [Pon13]. Shu [Pen06, Rei16]. si [Sni22]. Sibony [Eng20b]. sich [Lar07a]. Sicherheit [Kar14a]. Sicherheitsstrategie [Kar14a]. Sickle [Eva07]. Side [Log01, Neu11]. Sideshow [Rob14]. Sidney [Aft20]. siècle [Mar13b]. Siege [Hou16]. Siegelbaum [Joh14a]. Siegfried [Ban14]. Sights [Kri05]. Signs [Sco17]. sigurnost [Val14]. Silent [Buh15, Mun21, Smi08, Gal16b, Pra20]. Silicon [Kle08]. Silvio [Van99a, Bro11, Pet11a]. Simba [O’M21]. Simon [Cra17, Jer11, Kru07b, Mas21a, Spe18, Ale08, Tho08]. Simona [Bro10]. Simons [Mag14b]. Simpson [Eas10, Joc07]. Since [McK15a, Bal08, Din04, Dre05, Kar15b, Low08, Mas00, Rep12, Say17, Sha07a, Sha08b]. Singapore [BJ10, Ces21]. sinistra [Dra21, For02]. Siniver [Gar09, Qua14]. Sino [CK11, Fri10, Kal12, Kap10, Lev10b, Lüt22a, Pen06, Rob10, Sin18, Sti10, Vám10, Wis00, Xia14, Xia23, Zha02, Zha12b, Ano10i, Che07, Che20, Kho10, LX14, Mun10, Rad06, Rad07, RLV^{+}10, Sta16, Tag22, Xia06, You23, Zhi00]. Sino-American [Sin18]. Sino-Soviet [Lüt22a, Pen06, Sin18, Xia23, Zha12b, You23]. Sino-Vietnamese [Zha12b]. Siobhan [Joh20b]. SIOP [Bur10]. SIOP-62 [Bur10]. Sippial [Yor21]. Site [Nor05]. Six [Din13, Gol06a, Gol10, HSB^{+}23, Kna05, Rom17, Qua04]. Six-Day [Gol06a, Kna05]. Sixsmith [Lar22]. Sixteenth [Neh22]. Sixties [Jan08]. Sixty [Har11, Mas21b, Wea15]. Skepticism [Dom11]. Sketches [Sok99, Sta07]. Skies [Pra15a, Pra15b, Tal08, Mar07c]. Skybolt [Jon19]. Slater [Fed17]. Slave [Pol07b]. Sławomir [Szp09]. sled [Boy23, Val14].

Slippery [Ban16]. Slobodian [Yor19a]. Słoniewski [Kaź19]. Slope [Ban16]. Slovakia [Sto11, Ryd14]. Slowakei [War06]. Small [Kal16, Rus04a, Wir06]. Smedley [Eck07b]. Smells [Kri05]. Smith [Joh14b, Kul13, Moe18, Ran12, Sha07b, Spe18, Spe14, Fou14, Rey23, Spe15]. Smoke [Ben05]. Smokestack [Ste02]. Snyder [Sta07, Wan09, BH17, Sch19]. Social [Con03, DF07, Eng09, Gil23, Kuk21, Lor17, Mue13, Nut02a, Rad14, Ric22, Sch03, Ste05]. Socialism [Arg04, Loe10, Šmi23, Sto08, Car13b, Dit09]. Socialist [yL12, Mic08, Pét13, Shl09, Cos22, Rad14]. Societies [Shl06, Whi14a]. Society [Bru10, Con03, Dom07, Haj12, Her99, Hur10, Kom18, Man10, Mou03, RTYD19, Ste10, Tho11, Ale12, Gen17, Sac18, Wil17]. Sociology [Eng05]. SOE [Has07]. Soft [LSvdV04]. Sohail [Joh13a]. Soldaten [Her08]. Soldier [Taa13]. Soldiers [Bac09, Ham18, San16]. Söldner [Kan14]. Sole [Sza18]. Solidarity [Ben21, Get15, Hal16, KW22, Szp11, Szp14, Tod17]. Solidarność [God11]. Solitary [Wil08b]. Solo [HK20]. Solovey [Gil23]. Solution [Pol10]. Some [Eng03, Mag14b]. Son [Ker15]. Sons [Gra13a]. Soobshcheniya [Mue10]. Sorbonne [Mar13b]. Sorge [Her08]. Soul [Lev10a, Sch20]. Sounds [Gun12, Kri05, Cal14]. Source [JM23]. Sourcebook [Din07]. Sources [Hol05a, LSvdV04, Sak06]. South [Ale07, Cat08, Cli15b, Dom07, Dom15c, Dro18, Duk11, Eck14, Eng20a, Gar04a, Hat10, Hus00, Kim16, Law08, Lew07, Moi14, Moi15, Vei13, Wee18, Gaw21, Kra06f, McM06, Mil13]. Southeast [BJ10, Edw16, Hac03, Hes12, Wil08b, BJ19, Tho21, Tay19]. Southern [Ass14, KG11, Law12, Nie10b, Kou12, Rei19]. Southwest [RG15]. Soutou [Wal08]. Sovereign [Ham18]. Sovereigns [Hur10]. Sovereignty [Hal07a, Pem16, Eng09]. Sovetskaya [Min19]. Soviet [Ado08, Ale01, Ale12, And05, And10, Bal10, Bar08, Ber01, Ber04b, Ber07, BL10a, Bra16a, Bra16b, BSK^{+}23, Byr01, Car13b, Cea06, Coo04, Cri20a, Cul22, Dan07, DF07, DF12, Dav07, Der02, Din14a, Dom05, DFK^{+}20, Eng20b, Eng02, Eva00, Eva07, Fri02, Fri10, Gal16a, Ged99, Gol10, Gor03, Gra12b, Gra12a, Gre20, Hal07a, Hal07b, Hay99, Her05, Hut01, Imm16, Ing04a, JK07, Jen17, Joh14a, Kap10, Kho10, Kip15, Kle08, Kle23, Kli02, Kot17, Kra04a, Kri04, Kul13, Kur05a, Kuz19, Lar07b, Lev10a, Lev10b, yL07, Lib01, Lib07, Lüt06, Lüt22a, Mar07a, Mar08a, Mat04a, McD06, McM08, Mic21, Mic08, Mou03, Mul23, Nai02, Nei10a, Nor00, Nor10, Par09, Par13, Pec13a, Pen06, Pip15, Pod14, Pol13, Pon01a, Pon13, Pow01, Pra99, Ran10]. Soviet [Rob10, Rub06, Rus05, Rut12a, San16, Sav14, Sel02, Sew23, Sha20, Shl06, Shm23, Sil99, Sim17, Sin18, Smi11, Sny07, Sny09, Sol13, Sta07, Ste99, Ste18, Sti10, Tag22, Tay00, Tho08, Tro16b, Usd20, Vai05, Val08, Vám10, Van99a, Van99b, Vau15, Wei04b, Wet09, Wil08a, Wis00, Woo12, Wym03, Xia14, Xia23, Zal04, Zha02, Zha12b, Zhu16, Zhu20, Åse16, Ano10i, AE02, Bék16, Ben05, Ber20, BL10b, Bod08, Boz15, Bro15, BSK^{+}23, Che06, Chu19, Cle00, Con03, Con07, Cri20b, De 20, Din22, Dra12a, Dra04a, Dra04b, DF19, Dun03, Fer08, Fer11, Fet20, FP21, Fox09, Fri18, Gel12b, Get15, GK02, Gol06a, GSP^{+}19, Gra21b, HP17, Has13, Hat09, HK09, Her14, Her09, Her20b, Hol09, Ian14, Jen21, JKM12, Kal09, Kap16, KH23, Kni03]. Soviet [Kou12, Koz14, Kra99b, Kra99a, Kra99c, Kra03a, Kra03g, Kra03f, Kra04b, Kra05a, Kra22, Kre07, Lan20, Lar10, LX14, Mar02, Mas10b, Mik12, Mun10, Neh21, Neh22, Olš18, PA23, Ped19, Rad06, Rad07, Ria17, Rob02b, Rob04, RLV^{+}10, Sel21a, Sen16, Sha12, Shi13, Šmi23, Sta16, SM12, Sza18, Tay03, Tel18, Tol10, Tor22, TSW^{+}22, Tum03, Usd22, Van11, Wal03, Yor23b, You23, You13b, Zak10, Zak18, Zhi00, Zub99, Zub13, Jac23]. Soviet-American [Mar07a, Val08]. Soviet-Bloc [Mar02, Sel21a]. Soviet-Brazilian [Her20b]. Soviet-East [Ado08]. Soviet-Occupied [Nai02]. Soviet-Style [PA23]. Soviet-Type [Shl06]. Sovietization [Nai10a]. Sovietologists [Kra04a]. Sovietology [Pip11]. Soviets [Ado08, Bar11a, Bor09, Jol07]. Sovmestnoi [Mue10]. sovremennik [Tho08]. sowietischen [Shl06]. Sowjetisch [Mas19b]. Sowjetisch-indische [Mas19b]. Sowjetische [Bis10, Kur21, Ste20, Wil08a, vD02]. sowjetischen [Bis13, Gri10, Rug16a]. Sowjetstaat [Akt10]. Sowjetunion [Gri10, Kra01]. sozialistischen [O’S17]. Spaak [Van11]. Space [Col15, Dav09, Jen21, Neu11, Smi08, Smi11, Von03, Wan16]. Spaces [Eng20a, Pan01]. Spain [You22]. Spangled [Din13]. Spanish [Foo21, McM08, Pay21]. Speaking [And05]. Special [Dei02, GK02, Kna02, KNS18, Kra03g, Kra03f, Luk07, Mil01, Pif15, Wil02, Wir13, You07]. Spectacle [Mar15]. Spector [JLBK17]. Spectrums [Dum14]. Spenser [Lov10]. Sperling [Mas21b]. Spharen [Shl06]. Spheres [Raj14, Shl06]. Spiders [Fox09]. Spied [Kle08]. Spiegel [Rug16c]. Spies [Hay03, Hou15, Mad13, Sib11, SM12, vD21, Hay11, Kle23, Col05]. Spionage [Las19]. Spirit [Mar08b, Nie10a]. Split [Ano10i, Fri10, Kap10, Lee17, Lev10b, LX14, Meh11b, Mun10, Per07, Rad06, RLV^{+}10, Rob10, Sta16, Sti10, Vám10, Xia14, Zha02, You23]. spohady [Sny06]. Spoils [Tal11]. Sponsored [Hei16]. Spontaneous [Ged99]. Sport [SY17, Soa09, Soa20a]. Sports [Soa20b]. Spread [Jos23]. Spring [Dap14, Joh06, Kor06, McG99, Wil12]. Springer [Lee12]. Sputnik [Wan16]. Spy [Bru11, Dan16b, Fed17, Hay06, Hay15a, Her04a, Kle11, KH23, Mac20b, Nor15, Ped07, Pra17, Sar05b, Sza11, Usd11, Cla09, Cra17, Hay16, Taa13, Mar08a, Gar00b]. Spy-Tech [Bru11]. Spying [Ald08, Dav09, Nor10]. Square [Rut12b]. sredata [Val14]. Srinath [Gan16]. Srivastava [Del23]. SShA [Rob08]. SSR [Zak18]. SSSR [Odo02, Rob08, Gor01]. St [Rya10]. Staat [Bru09]. staatlicher [Shl06]. Stability [Dal10, Dom11]. Stable [Mas21b, Tra22]. Staerck [Mur07]. Staff [Bro02a]. Stafford [Mar10]. Stages [Ros10a]. Staging [PW05]. Stagnation [Bit17]. Staktopoulos [Iat18]. Stalin [Bae02, Ber01, Bra16b, Che99, Cra17, Dal00, Eng20b, Ert10a, Ert10b, Gal07, Gel12a, Gre08, Gri10, Hay99, Jac23, Jer08, Kle08, Kul10, Lee17, Len12, Mad08a, Mar01, Mar06b, McD06, Min19, Mou02, Nai04, Nai08b, New07, Nor00, Par09, Pon01a, Ran12, Rie15, Rug16c, Sch00b, Ste99, Usd11, Yor17b, vD05, Ale08, And10, AE02, Car13b, DF12, Dra12b, Ell12, Gra12c, GRS^{+}12, Har12a, HK09, Hol12, JMG08, Kra99b, Kra99a, Kra99c, Kur10, yL09, MSPN21, Meh11b, Nai08a, Nai16, Pec08, Per07, Pol07b, Pon01b, Por09, Ree16, Rob02b, Rob07, Rob08, Ros12, Rub06, Rub12, Rug11, Szp12, Ula99, VA12, Wol13, You13b, Zak18, Zhi00, Cap08, Len18, Mic08, Bra04, Bra05, Ing04a, Kra06a, Kur15, Sha20]. Stalin-Note [Rug16c, vD05]. Stalinism [Con07, Kri04, Sen16, Vai05, Lyn16, Ran06, Ran02]. Stalinist [Con07, DF07, Edg08, Hes13, Nat04, Zub14, Kap16, Puc21]. Stalinization [Lev07, yL01b, Mëh11a]. Stamps [Mod17]. Stand [Joh14b]. Standards [Hes13]. Stanford [Dan16a, DeR17, Kap17a, Lüt12]. Stanley [McM08, Ros06, Win07]. Stanleyville [O’M21]. Star [Haf08a, Joh13a, Sar05a, Din13, Hat21, Von03]. Star-Crossed [Von03]. Starck [Dum17]. Stars [Eva07]. Start [Mun12]. Stasi [Bru11, O’S17, Bru03, WF12]. Stasi-Operativgruppen [O’S17]. State [Arg04, Bot17, Bru10, Bud22, Cra13, Dan14, Dei15, Din23b, Fis03, Gen17, Gol12, Haj12, Hal07b, Her21, Hol05b, Hou16, Kle11, Kna03, Kru04b, LX23, Mad07, MBR^{+}21, Mas21b, Neh21, Nie21, Pre11b, Sap03, Sas14a, Sch13, Sch03, SS14, Sel19, Sel21b, Sfi02, Shl06, Shl09, Ste10, Sto11, Thi14, Wee18, Whi02a, Wil00, All15, Dar15, Dar16, DS14, Kap17a, Rid16, vD12, Dob10b, Pai14]. State-Private [SS14]. State-Socialist [Shl09]. Statecraft [Kno06, Mar07c, Rei16, Ste15, Tal99]. stat’ei [Kar15c]. Stateless [Car99]. States [Ass14, Bal10, Bis12, Bro10, Cal14, Cly01, Cro09, Dav11b, Din14a, Dom15a, Dom20, Dum20, Elm01, Eva13, Fer07, Fri02, Fri05, Gar04a, Hat11, Hez02, Hut01, Hut19, Jon04, Kal16, Key16, Lat16b, Lee23, Lev10a, Low08, Mar07b, McM12a, Mil04a, Nin03, Pas03, PL22, Pre11b, Puc21, Rat21, Sav14, Sch01, Sha07a, She14, Sim04, Soa22, Wal12b, Wal08, Yor15, Yor21, BJ19, Cit14, Din22, Gol06b, HS15, HK09, Her09, Her04b, Her04c, Jer22b, Lit10, Mat12, McA22, McM22, Nil19, Nut02c, Rea11, Tra08, Tra22, Wan17, Zha23a, dL13, dS19, Che99, Kru04a]. Statesman [Rub11]. Statesmen [Mas00]. Stati [For02]. Station [Col15, Gaf11, Joh21, McA07]. Stationen [She12a]. Statler [Lan10, Lat08]. statunitense [Nut10, Nut11]. Staub [Din07]. Stay [Ris14]. Stayed [Ole18]. Stealth [Gal09]. Steege [Mar12]. Stefan [Soa20b, Bis10, Kur10, Nai08b, Sch10]. Stefano [Gor09]. Steiermark [Bis10]. Steinberg [Sza14]. Steininger [Bon07]. Steinweis [Rai11]. Stephan [Mac20a, Sza10a]. Stephanie [Eva23]. Stephen [Aft20, Fer07, Kur15, Pra14, Tum16, Wal08, Ash10, Imm16, tL16, Mil17b, Ram07, Rus07, Sen16, Soa09]. Stepping [Lar10]. Steve [Sha11]. Steven [Wea15, Ale12, Kle08, Rea06]. Stevens [Rob20]. Still [KM22]. Stillborn [Her02]. Stimson [Iri02]. Stockholm [Wys12]. Stockton [Joh11b]. Stolen [Rad12a]. Stone [Hol09, Lun04]. Storia [Mas11b]. Stories [Hol05b, Lin15, Mac20b, Pra20]. Storm [Pra17, Rad12a, Wir07b]. Storming [Pow01]. Storms [Mar23a]. Story [Dan12a, Dan16b, Fie03, Gal15, Gra13a, Kru07a, yL01a, Lit10, Mat10b, Mat20a, Moe18, Nor20, Pet08, Pol06, Rad07, Smi11, Tud05, WF12, Wir13, Ram07, Rad15, Stu19]. Strait [Zha23b]. Strange [Kal14]. Strangelove [Her99]. Strangeness [Ket14]. Strangers [Chu19, Qua21]. strany [Kar15c]. Strasbourg [SS14]. Strategic [Abe21, Ald07, Bur10, Clo06, Dob10a, Jon19, Lar07b, Mur02, Pif15, Pol07a, Ray20, Rep03, Sch02c, Tra22, Wal08, Zha23a, Zhi00, Zub99, Ric15]. Strategie [Ban07]. Strategies [Elm01, Jer06]. Strategist [Pfa02]. Strategy [Ass11, Bor12b, Bro02a, Bur05, Byr01, Cor14, CK13, Dre21, Fav18, Gre10, Iat05, Jac17, Kap04, Luc02, Mik12, Pet11b, Pre17, Sfi01, Sha18, Soa06, Søn19, Ste15, Sza12b, Zha23a, Rak15, Cro15, Lib09, Wir07a, Mad02, Wir03]. Strath [Tum14]. Straus [Gal16a]. stredisko [Gáš13b]. Strike [Sch09, You07]. Strikes [Wil07]. Strings [Gan06]. Stripes [Eva07]. Stroilov [Rad12a]. Strong [Żuk19]. Structures [FP21]. Struggle [Ash10, Bis04, Blu14, Dra08, Eva13, Fer07, Kle04, Kra99b, Kra99a, Kra99c, Lau00, Lew07, Nai16, Pet20, Rab13, Sch20, Sel21a, Sny11]. Struggles [Low08]. Stuck [Win17]. Studenata [Val14]. Student [Gar04b]. Students [MC23]. Studien [Gri10, Kan10]. StudienVerlag [Mol16, Rug16b]. Studies [Cap08, Cri20b, Cri20a, Kan10, LX23, Tho11]. Study [Boo03, Dav11a, Der02, Fri04, Kra22, Len12, Rob14, Sha08b, Wea15, Xia08b]. Stueck [Liv07]. Stuttgart [Tum16]. Style [PA23]. Sub [Sew23]. Submarine [Åse16, Dan07, Gal22, Gal09]. Submariner [Gal15]. Submarines [Pol06, Gal16b]. Subversion [Boy23, Pos19, Whi13]. Subversives [Zak10]. Subvert [Byr01]. Subverted [Ros10b]. Subverting [Gil05]. Success [Kap17b, Kob06, Nor20]. Successful [Gun15]. Succession [Bud22, Del22, Kra99b, Kra99a, Kra99c]. Such [Mad04]. sue [Rom17, KG11]. Suez [Cro07, Rey02, Rua04]. Sugarland [Neu23]. Sukarno [Gar02]. Sulian [yL07]. Sulmaan [Dre21]. Sumantra [Gan07]. Summer [Joh06, Ree16, SS18a]. Summit [Bra02a, Van02a]. Summitry [Bro22]. Sung [Pai14, Del22]. Suns [SZK12]. Suomettuminen [Mas19a]. Superbomb [Her20a]. Superpower [Bur16, Elm04, Kou17]. Superpowers [Qua09, SY17, Sai02]. Superstate [Büt05]. Support [Col19, Din23b, Fer08, Fri18, Uel20, Vei14b, Kli05]. Supranational [Lie08]. Supranationale [Lie08]. Supremacy [Kle04, Nai16]. Supreme [Pfa02]. Suprun [Gol12]. Suri [Bur08]. Surprise [Ken05, Kra04a]. Surrender [Gal07, Mis23b]. Surrogate [Vau15]. Surveillance [Bru03, McM21b, Puc21]. Survival [Mue17]. Survivor [Lie21, Wil08b]. Survivors [Din14b, Bar03a]. Susan [Sar11, Bel13]. Susanna [Fer07]. Susanne [Din14b]. Sustainability [Woo12]. Suzanna [Cra21]. Suzanne [Dan16c, Sta13]. Svetlana [Bra16a, Pod14, Bra16a]. svetovna [Boy23]. Svetozar [Yor20a, Pet18]. svideteli [Bra16a]. Swans [Zhu20]. Sweden [Eng09, Sal06, Åse16, Mak12, Mat12]. Swedish [Eng09, Wys12]. Swedish-American [Wys12]. Swiss [Buc17, FM16, Knü20]. Sword [Duk11, Fri05]. Symbols [Nei00]. Syndrome [Sta15]. Syracuse [Hoc16]. System [Cle00, Con07, Dav10, Gil05, Hat21, Kap04, Łuk23, Mar13a, Sob19, Tra22]. Systemic [Shl09]. Systems [Dro18, McD02, Dav09]. Szabo [Wal08]. Szalontai [Lee12]. Szaynok [Pol13]. szovjet [Vaj20].

T [Gar13, Hal07b, Mas11a, Mil21, Pol10, Shl06]. Taaffe [Mil17b]. Taboo [Han17]. Tactics [Iat05]. Tadeusz [JK07]. Tagebuch [Nai10b]. Tai [Dre13]. Tail [Boy11]. Taiwan [tL16, Rub11, Zha23b, Jac16, Rub11]. Taiwanese [Rea06]. Taking [Pec13b]. Tales [Bur10]. Talks [Coh08b, Rad17, Tag22]. Tamman [Taa13]. Tango [Ole11]. Tantalus [Tud05]. Tanwar [Gan08]. Tanya [Moo23]. Tapes [Pos07]. Tarah [Ber15]. Target [Her09]. Targeting [Jon08]. Tasić [Lee17]. Taste [Chi14]. Tatiana [Mag14b]. Taubman [Pra19]. Tauris [Dei14, Imm16, Sha11, You13b]. tausta [Mas19a]. Taylor [Joh11a]. Teachers [Vai05]. Teaching [BD22, Kle22]. Team [Bra16b]. Tech [Bru11]. Techno [Pol07a]. Techno-Warriors [Pol07a]. Technological [Dre13, JE19]. Technology [Bur17, Hel22, His13, Jos23, Wys12, Rep12, Whi03b]. Ted [McA07, McA08, Guz13]. Tedesco [Lat16b, Att23]. Teenage [Ram11a]. Tehran [Rob07, SJ23]. Teigrob [Her17]. Teilung [Rug17]. Teilungsgeschichte [She12a]. Telepneva [Mul23]. Teletyped [Kel19]. Television [Lor17, Hol08]. Teller [Bor12a]. Telling [Nei08]. Temporary [Lar07a]. Ten [Mat04a, Sny08]. Tenable [Dei13]. Tendencies [Shl09]. Teng [Rub11]. Teng-hui [Rub11]. Tennent [Cla09]. Tentative [Lum04]. Termination [Zha12b]. Terms [Zub14]. Territoriality [Ros09]. Terror [Blu15, Dra04a, Ert10a, Ert10b, Fis03, GRS^{+}12, Mou02, Sch00b, Maw08, New07, Yor15]. Terrorism [Blu14, Dra10, Kau09, Mor06, Wee18, Eva13, Joh07c]. Terrorismus [Cra07]. Terrorists [Her12]. Test [Gil00, Mas08, Mil21, Pie16]. Testament [Arg03]. Testicles [Dum17]. Testing [O’D09, Wal21b]. Texas [Col15, Gaf11, McA07]. Textual [DX18]. Thai [Has07]. Thailand [Has07, Ray20]. Thames [Wir13]. Thang [Law07]. Thatcher [Bro22, Bro08]. Theater [Gaf14, Gel12a]. Theatre [Aro06]. Theft [Bar08]. Their [Ber04b, Bur23, Lee21, Luk07, Pra14, Roz11, Tol10, Wil08a, dS19, Mat04b]. Them [Pra20]. Theodore [Eck14, Mun21, Sim14, Fei10]. Theodosius [dL13]. Theoharis [Ric09]. Theoretical [Tan05]. Theory [Alk05, Boc06, CS07, Gho00, Hop03, Kuk11, Kur18, Sch03, Sta15, Wel00, Woh99]. Therapists [Ber04b]. Therapy [Tum16]. There [Duk13, Gei12, Vuj02]. Thermonuclear [Gra21a]. These [BJ19]. Think [Bra14]. Thinking [Car17, Eng05, Kal16]. Third [Bru08, Cra13, Del23, Joh22, Kan12a, Kno06, Lat08, Lüt08, McM16, Neu11, Sin18, Yor18, Lar10, McM22, ŽČ16]. Third-Line [Lüt08]. Third-Party [Kno06]. Thirty [Par19]. Thomas [Kar14a, Kir08, Las19, McP19, She07, Whi14a, Bur13, Coh12, Dav11b, Dob10b, Eva13, Her07b, Joh14c, Kuz16, Mic08, Mis23b, Rep12, Sch08, Sch22a, Sch22b, Sza11, Vei13, McP19]. Thomasson [Sni22]. Thompson [GSP^{+}19]. Those [Kra06a, Buh15]. Thoß [Ban07]. Thought [Kro08, McM11, Rot21, Zub99]. Threat [Bru03, Gas19, Jer13, Jer16a, OB20]. Threats [Mat16a, Cro15]. Three [Gal16b, Mis23b, Vu19]. throughout [JK07]. Thugs [Ram11a]. Thule [Ole11, Pet11b]. Thunder [Oye14]. Tiananmen [Wal21a]. Tibet [Myl12, Gol06b, Hof06, McG06, McM06, She06, Zha06]. Tibetan [Che06, Kna03, Kra06f]. Tide [McC14, Pol06]. Ties [Jer00, Low08, Yor23b]. Tiffany [Yor21]. tii [Sni22]. Tim [Bis14, Edw16, Spr07]. Time [Ger13, How22, Lee12, Sok99, Sil99]. Times [Pra19]. Timothy [Sim14, Ste18, BH17, Kno06, Rut11, Sta07, Wan09]. Tina [Taa13]. Tinderbox [Joh06]. Tiné [Jak16]. Tired [Joh07a]. Tiso [Ryd14]. Tito [Lee17, Woo07, Meh11b, Nie21, Per07]. Tito-Stalin [Lee17]. Toaldo [Blu15]. Tobia [Bro10]. Toby [Soa20a]. Today [Mar01]. Todd [Bre21, Rub06, Gos08]. Todor [Neh22]. Together [Sav14]. Togliatti [Dra12b, Pon01b]. Tom [Bea06]. Tomasz [Ber07]. Tomlin [Sch17]. Tomoff [Zhu16]. Ton [Law07]. Tongues [And05]. Tonkin [McM21a]. Tony [Whi08b, Wir13]. Too [Tal08]. Tools [Wys12]. Top [Cor14, Sar05b]. Top-Secret [Cor14]. Topmiller [Law08]. Topping [Dan10]. Torch [Soa13]. Torma [Taa13]. Torsten [Wet11a, Wet11b]. Tossing [Ste14]. Total [Law12, Kru08]. Totalitarian [Mou03, Whi04]. Totalitarianism [Chi14]. Toth [Mer13]. Tourism [DF12, Mar07d, Wal07, Zak18]. TPBEDAMN [Gas13a]. tra [Dra08, Gra07]. Trachtenberg [Cog03, Hit03, She03]. Tracing [Han17]. Track [Pie16, Rea11]. Tracking [Shm23]. Tracy [Ros10a]. Trade [Fer11, Ing17, JE19, Rut12a]. Tradition [Hol06]. Traditionalist [Hay00]. Tragatori [Sni22]. Tragedies [CKP05a]. Tragedy [Bon05, CKP05b, Dan07, Hes02, Hix05, RTYD19, Rub06, Sch02a, Tra05]. Tragic [Bor05b]. Trailblazers [Smi08]. Training [Rya08b]. Traitor [Die08]. Trani [Roz11]. Trans [Shl09, Tod17, Bra16a, Her12, Kip15, Kor06, Nai10b, Rub06, Sar11, Sim16]. Trans-American [Tod17]. Trans-Systemic [Shl09]. Transaction [Nam13, Pét13]. Transactional [Hal16]. Transatlantic [Nie12, Hug08]. transedited [Tis15]. Transfer [Dom15c]. Transformation [Cha07b, Cli15b, Dit09, Dom11, Geh05, Mas10a, Neu23, Nie16b, Wen04]. Transformed [Bur16, Ham18, Har11, Lar02, Sch19, Tha06, Cha00]. Transforming [Hen09]. Transilvania [Gor09]. Transition [Bla03, Rut06, KW07, Kor07, Kri05, Szp07a]. Transitional [Yan08]. Transnational [Boe21, Fal20, Gil23, KNS18, Lat16b, Moï10, Pai17, Pet20, Shl09, Spe01, Wee18, Sto07]. Transparency [Mar07c]. Transplanting [Man10]. Trauma [Ber04b]. Trauschweizer [Hit16]. Travel [Zak18]. Treason [Dum17]. Treasonable [Ped07]. Treated [Joh07a]. Treaty [Mas08, Mue11, Pie16, RG15, Wal21b, Wil00, MBR^{+}21, Adl03]. Trek [Sar05a, Hat21]. Trial [Mad07, Rya10]. Trials [Sib11, Ric12]. Triangle [Wal08, Zub13]. Triangles [O’D09]. Triangular [Spo12]. Trident [Fri04]. Trigger [Mas09]. Triggering [KWKS07, KW07, Kor07, Szp07a]. Trilling [Whi10]. Tripartite [Zub13]. Triplex [Hay11]. Triumph [Cri02, Ale09]. Triumphalism [Hei06]. Triumphant [Ass11]. Troop [Zim09]. Trophies [Sny03b]. Trouble [Kos15, Lan20]. Troubled [Mur02, Pap23]. Troubleshooter [Wie07]. Troy [Sur13]. True [Bro04, Dan16b, yL01a, Cra17]. Truman [Che10, Gal07, Ham15, Jon04, Mad15, Mad04, Mad07, Mit07, Sch13, vDO08, Fra09, Lar11, Mar05, Mat16b, Bos15, Kap14]. Trumen [Rob08]. Trust [Dan16c]. Truth [Che08b, Rid16, Sme15]. Trying [Goo05]. Tsai [Rub11]. Tsang [Sha11]. Tsar [Ale08]. Tsarev [Hay11]. Tsarist [Gra07]. tschechische [Bis11]. Tschubarjan [Mue10]. Tsuchiyama [Dob10b]. Tsuyoshi [Gal07]. Tucker [Lei08]. Tudda [Che08b, Cha14]. Tudeh [Fri18]. Tunstall [McP19]. Turco [Myl15]. Turin [Bro11, Mas11b]. Turkey [Chu07b, Gun15, Sön22]. Turkish [VA12]. Turkistan [Jac23]. Turn [Gol11, Whi10]. Turnabout [Mar06a]. Turned [Gra07]. Turner [Dei13]. Turning [Cha14, dS19]. Turtle [Con23]. Tusculum [Nie12]. Twentieth [Aki02, Bor12a, Gor05, Her17, Kag05, Rad14, Roz11, Rut10, Sch02c, Sza12a, Zei08, AS22, Dav11a]. Twentieth-Century [Gor05, Her17, Rut10, Zei08, AS22, Dav11a]. Twenty [Xia08b]. Twice [Vuj02]. Twigge [Fer07]. Twilight [Ric07]. Twisted [Usd20]. Two [Bru16, Bur10, Duk13, Jac17, Kle08, Mun21, Sch13, SZK12]. Two-Level [Jac17]. Twomey [Kal12]. TX [Col15, Gaf11]. Type [Shl06]. Typs [Shl06]. Tyrant [Kra06a].

U [Dob10b, Pra07, Bra16a]. U-2 [Pra07]. U.N [Mil01]. U.S [Alp09, Ava04, Bal08, Bea06, Blu15, Bro02a, Cat05, Cha04, Coh06, Coh08b, Cra13, Dan16a, Duk13, Eas10, Fre07, Gal16b, Gar09, Ham18, Hit16, Jon03, Kal18, Kli05, Kru07a, Kur05a, Lak11, Lib01, Lin11, Luc02, Man13, Mit07, Moo23, Pai17, Qua21, Ric09, Ric12, Ris15, Rog15, Ros00, Sch19, Sep01, Sie06, Von03, Wet01, Win07, Win11, Xia08c, Yaq02, Tho21]. U.S. [BMLG11, BM12, BD22, BL10a, BL10b, Bor99, Bra06, Bur17, Chu19, Col15, Col19, Dar12, Din10, Dob10a, Dom11, Dom09, Ess15, Fox09, Fro23, Gas19, Gav99, Gaw21, Gel12b, GSP^{+}19, Gun15, Has13, Hei15, Hel22, Her21, Hol05a, Hou15, Ing04b, Jac17, Jen21, Joh07a, Jon08, Kim16, KH23, Kna03, Kre07, Lou23, Lud13, Mar23b, McC14, McK15b, McM06, Mei15, Mic11, Mik12, Mod17, Mun12, Odo06, Oye14, Pem16, Pet11b, Pet20, QG12, Ram07, RG15, Sar05a, Sar19, SS14, SS18b, Sha12, Shi13, Sme15, Sny13, Ste14, Sti06, Sto15, Tol10, Tra11, US21, Ver18, Whe06, Win17, Yar09, ZP23, Zha23b, Zub13]. U.S.-Based [KH23]. U.S.-China [Coh06, Coh08b]. U.S.-Chinese [Pai17]. Udo [Bru08]. UdSSR [Ste20]. UK [Eng07, Gol11, Goo15, Goo17, Has07, Her11b, Jon14, Key16, Lee12, Lin11, Nor15, Sha11, Tal11, Tro16b, Wir13]. Ukraine [Kes07, Kuz19, Sny03b, Ste18, Bud22, Sta07, vH09]. Ukrainian [Git07, Sol13, Sny99]. Ukrainians [Sny99]. ukrains’ke [Sny06]. Ulam [Kra00]. Ulbricht [vD03, Ban16]. Ultimate [Col19, Gaf14]. ultimo [Mas11b]. Unaccompanied [Att23]. Unanswered [Mat16a]. Unarmed [Spe01]. Uncanny [Tis15]. Unconquerable [Sha08a]. Unconventional [Osg02]. Uncovering [Wir07a, Ava04]. Undeclared [Ale04, Mat16b, Kop16]. Underdogs [Hof20]. Underground [Has07, McE03]. Undermining [Byr01]. Understand [Sci20]. Understanding [Gor05, HS15, Sal06, Gel12b, Bar04]. Uneasy [Sac18]. Unexpected [Sou21]. Unfinished [Ass23, Gre05a]. Ungarn [Mer13]. ungarndeutscher [Mer13].

Unger [Mac19]. Unification [For05, Ged99, Rug11, Ser01, Sar11, Sza12a]. Unified [HP17]. Unintended [Sha11]. Union [Ale01, BSK^{+}23, Cea06, Coo04, Cri20a, Cul22, Der02, Din14a, Gol10, Hut01, JK07, Kra04a, Kul13, Lev10a, Mat04a, McM08, Mou03, Mul23, Nor00, Par09, Pip15, Pon13, Rus05, Sav14, Ser01, Sny09, Van99a, Wei04b, Wym03, Zal04, Boz15, Bro15, BSK^{+}23, Che06, Chu19, Cri20b, De 20, Din22, DFK^{+}20, Fri18, Get15, GK02, Gol06a, GSP^{+}19, Hat09, Ian14, Kni03, Kra03a, Kra03g, Kra03f, Kra04b, Kra05a, Kra22, Mas10b, Neh11, Pol13, Shm23, Spe14, Tol10, Tro16b, Tum03, Van11, Wal03, Woo12, Bal10, Gre20, Hal07a, Lar07b, Spe15, Fri02, Gor03]. United [Ano08i, Ass14, Bal10, Bro10, Cal14, Che99, Cro09, Din14a, Dom15a, Dom20, Dum20, Eva13, Fer07, Fri02, Fri05, Gar04a, Hat11, Hez02, Hut01, Hut19, Kal16, Key16, Kru04a, Lat16b, Lee23, Lev10a, Lie12, Low08, Mar07b, McM12a, Mil04a, Nei14, Nin03, Pas03, Pre11b, Rat21, Sav14, Sch01, Sha07a, She14, Sim04, Soa22, Wal12b, Wal08, Yor15, Yor21, BJ19, Cit14, Din22, Get15, Gol06b, Göt08, HS15, Har12b, HK09, Hei15, Her09, Her04b, Her04c, Jer22b, Lar10, Lit10, Mat12, McA22, McM22, Nag17, Nag18, Nil19, Nut02c, O’D09, Rea11, Tra08, Tra22, Wan17, Zha23a, dL13, dS19, Cly01, Jon04]. Uniti [For02]. Uniting [Ros06]. Unity [Hus05, Mil04b]. Universal [Kag05]. Universalism [Fra09]. Universe [Bug15]. Université [SF11]. Universities [Eng03]. University [Abe21, Aft13, Ale12, Ass14, Bal17, Bar11a, Bar11b, Bar07, Bel11, Bel13, BD22, Bes07, Bla15, Boy11, Bra14, Bra16b, Bre21, BSK^{+}23, Bru11, Bru16, Bur16, Bur13, Cal14, Car13a, Car13b, Cli15a, Coh12, Col15, Cra13, Dan16a, Dan14, DF12, Dav11a, Dav12, DeM12, DeR17, Dei07, Dei15, Die07, Din07, Din14a, Dom15a, Dom07, Dou15, Duk13, Edw16, Fer15, Fin13, Gaf11, Gan16, Gar13, Gen17, Geo13, Ger13, Gol12, Goo15, Goo17, Gra12b, Gra12a, Gra13b, GP11, Gun12, Guz13, Ham15, Har11, Has07, Hat11, Haw07, Hay11, Hay13, Hel07, Her12, Her11b, Hes21, Hes13, Hit16, Hoc16, Hol15, Hor07, Hus13, Iri11, Jen17, Jer13, Jer16a, Joh07b, Joh13a, Joh14a, Joh14c, Jon07, Kal12, Kal16, Ken13, Ket14, Key16, Kim07, Kip15, Kle11]. University [Kop16, Kre12b, Kre12a, Kro12, Kuk11, Kuk13, Kur11, Kuz16, Lak11, Lat14, Lat16a, Lat16b, Law12, Lef12, Len12, yL12, Lie21, Lie12, tL16, Lit11b, Liv07, Lou16, Luc07, Lüt12, Mad15, Mal15, Man13, Mar07b, Mar12, Mar07d, Mas11a, Mas21b, Mat16a, Mat16b, McA07, McM12a, McM11, McM12b, Mil15b, Mil17b, Mil21, Myl12, Nai16, Nam16, Neh11, Nor07, Oli11, Pai16, Par13, Par16, Pet11a, Pon13, Pra17, Pra99, Pre11a, Pre11b, Pre17, Qua13, Qua21, Rab13, Rad16, Ram07, Ran12, Rat21, Ree16, Rep12, Roh17, Ros17, Roz11, Rus07, Sae07, Sau12, Sch13, Sch11, Sen16, Sha07a, Sib11, Sim16, Smi07, Sny11, Sny17, Soa13, Sol13, Sou21, Spr07, Sto11, Str12, Sur13, Sza21, Tal11, Tei11, Thi14, Tro16b, Usd16, Vei13, Vuj02, Wal12b, Wan16, Wea12]. University [WF12, Wei12, Whi16, Whi11a, Whi12, Whi13, Whi15, Wie07, Wir07b, Woo13, You12b, Zar21, Zha12b, Zhu16, vC13, Guo10, SS14]. Unleashed [Cat08, Law08]. Unmaking [Lee23]. Unofficial [Cea17]. Unpopular [Cha07b]. Unpursued [Ber13]. Unresolved [Ula99]. Unseen [Gal16b]. Unselfishly [McC20]. Until [Nei10a]. Untold [Dan12a, Gra13a, Mat10b, Moe18, Pol06, Stu19, Wir13]. Unwanted [KSR18]. updated [Coh10]. Upheaval [Wem21]. Upheavals [Kra99b, Kra99a, Kra99c]. upon [Gar04b, Wir13]. Uprising [Guo10, Hai16]. Upside [Gra07]. Upstaging [Whi11a]. Upton [Sco17]. Uranium [Luň10, Luň10]. Urban [Gil06, Hes13, Ran12]. Ursu [Sni22]. Urwand [Spa14]. US-Interventionen [Kan14]. US-Soviet [Eva07]. USA [Hus00, Law11, Mey11, Moe03, Moï11]. USAF [Bur10]. Usdin [Kle08]. Use [Ben05, Ber20, Gar04c, Leb06, Ros17, Zie21]. Users [His13]. Uses [Sba05]. Using [BL10a, BL10b, Leb03, Wir13]. USS [Che18]. USSR [Gra07, Kan12b, Ber14, BS02, Bud22, Con03, Fav18, Gei12, Gor01, Gra17, KSR18, KM22, Kli22, Mue16, Neh22, Nun22, Ped19, RLV^{+}10, Tay03, Ric07, Sim17]. Utopia [Chi04, Šmi23]. Utopias [Dou06].

V [Kul10, Lüt06, Mav18, Maw08, Nai21, Rad15, Sap03, Sav14, Ste08, Sza10b, Taa13, Gol12, Gor01, Kar15c, Min19, Odo02, Rob08]. VA [Joh11b]. Václav [Tis15, Gáš13b]. Vaizey [Har15]. vakoilukohun [Mas19a]. Vale [Nie10b]. Valentina [Mac20b]. Valentine [Din14b]. Valery [Dav07]. Vallentine [Sta13, Whi11b]. Valley [Nor08a, Kle08]. Value [Gol20]. Values [CS07]. Vanderbilt [Woo13]. Vanderbush [Smi07]. Vanek [Wil17]. Vanguard [Wil02]. Vanik [Mar21]. Vannevar [Sul22]. Vasilii [Kar15c]. Vassiliev [HK09, KLHK14]. Vatican [GP11]. Vaudagna [Mil06]. Vaughan [Sha08a]. Veiled [Edg08]. Veith [Moi14]. Vél [Kar15b]. Velvet [Wil17]. Verbeek [Cro07]. Verdeckte [Kan14]. Verdict [Mar09]. Vergeltungsstrategie [Ban07]. vergessene [Bon07]. Verhandlungen [Wet17]. Verheyen [Har12b]. Verification [Wal21b]. Verify [Dan16c]. Verlag [Bis11, Bis13, Mer13, Mey11, Mil17a, She12a, Tum16, Wet11a, Wet11b]. Verlages [Cra07]. Verräter [Die08]. versus [Pif15, Szp10]. Verteidigungsplanung [Ban07]. Vertreibung [Rie08]. Vertriebenen [Rie08]. Vertriebener [Mer13]. Very [Hay15a]. Veterans [Lem14, Lin15]. Vetoes [Par19]. via [Dra08]. Victim [Sny08]. Victimizer [Nei07]. Victims [Bal05]. Victor [Dra12b, Mas21b, Gol07]. Victoria [Zei08]. Victory [HT19, Mad04, Rus07, Bro04, Coh02]. vidlunnya [Sny06]. Vienna [Bis11, Den21, Mil17a]. Vier [Rug18]. Viet [Spe13]. Vietnam [Ale07, Ale09, Ass10, Ass15, Bac09, Bra14, Cat05, Cat08, Che08a, Gar00a, HTP^{+}15, Hes02, Hor07, JLBK17, Joh03, Kim07, Law01, Law05, Law08, Lem14, Lin15, Lüt06, Lüt12, Mar20, McA07, McA08, McM05, McM10, McM21a, Moï10, Moi14, Moi15, Moy11, Nic12, Pit23, Pre11a, Pre11b, Rad16, Rak15, Ros08, Rya08a, Sai02, Sch11, Sza11, Tha06, Tur01, Vei14a, Vei14b, Wal12a, Wie07, Wir07a, Ass07, Ass14, Ass23, BMLG11, BM12, Bur13, Bus14, Dad19, Dro18, Gar13, Gaw21, Her11b, Her03b, Kho10, KW09, Law10, Law12, Law22, Lin09, Log01, Lum04, Lüt08, McM22, Moi03, Neu03, Pre17, Ray20, Sho15, Sta16, Sza12b, Sza18, Szo10, Vei13, Vic08, You12b, Zha10, Dan10, Lan10, Lom09, Nor08a, Sch22b, Bac02, Div00, Law11, Moï11, Tra02].

Vietnamese [Law07, Zha12b, Mun12, Vu19, You23, Zin11]. View [Hay00, Kir09, Mav18, Sim10, Wil12]. Viewed [Hai16]. Views [Mac20a]. Vigilance [Zha23a]. vii [Ken13]. viii [Dre13, Rai11, Ree16]. Viking [Kim11]. Vilém [Gáš13b]. Villafana [Nam13]. Villaume [Ing17, Nie12]. Villeneuve [Kle22]. Violence [Bot17, Dar15, Dar16, Gil18, Gor05, Mac20a, Pai14, Rob20, Rya08b, Hay17]. Violent [Nag17, Nag18]. Violeta [Kot17]. Virginia [Dom15b]. Virtuosi [Zhu16]. Virtuoso [And23]. Vis [Har18, Ano08i, Bor12b, Göt08]. vis-à-vis [Ano08i, Bor12b, Göt08]. Vision [Kau02, Yar09]. Visionaries [KSR18]. Visions [Git20, Gun12, Haw07, Hus05, Jon03, Luc02, Mil04b, Par19]. Visitors [Pon13]. Visual [Kan10, Kaź09]. Vixi [Bro06]. Vladimir [Git07, Rob08, Luk01]. Vladimiro [Dra18]. Vladislav [BSK^{+}23, McM12b, Par09, Rub10]. Vlast’ [Nai21]. vo [Nai21]. Voenno [Odo02]. Voenno-promyshlennyi [Odo02]. Vogel [Cli15b]. Vogeler [Man21]. Vogeler/Sanders [Man21]. Voice [And05, Din02, Din23a, Lub11, Shm23]. Voices [Bra16a, Ert10a, Ert10b, Koz14, Kur11, Lin15]. voina [Boy23, Gol12, Val14]. voiny [Bra16a, Odo02]. Vojtech [Dav10]. Vol [Bru06, Din14b, Iri14, Kur15, Mue10, Roz03, Son02, Sta13]. Volk [Mue17]. Volkserhebung [Bru09]. vols [Cra07]. Volyn’ [Sny06]. vom [vD05, Jon06]. Voorhees [Mun21]. Vorhang [Bis11]. Vortex [Bor12b]. Voß [Bis14, Kan14]. Votes [Par19]. Vowinckel [Whi14a]. vreme [Boy23]. vremeni [Sok99]. vs [Cro00]. VT [Cro07]. Vtorata [Boy23]. Vtoraya [Odo02]. Vulnerabilities [Mar02]. vuoden [Mas19a].

W [Buh15, Dav11a, Gen17, Gro07, Hal07a, Joh14c, Kno06, Lin11, Mag14a, Mas23, Neu11, Nue07, Pra20, Stu15, Usd11, Usd22, Joh21, Kaź19]. W. [Joh11a]. Waged [Rub20]. Wagg [Soa09]. Waging [Dom15a, Mic17]. Waite [Jon14]. Wake [Kot17, Sch00a]. Walder [DeM12, Wem21, Wem21]. Walentynowicz [Szp11]. Wałesa [Szp09, Szp14]. Walker [Gar00b, Hal07a]. Wall [Ado08, Nue05, Rus16, Har15]. Wallace [Joh14c, Par19]. Wallenberg [Mak17, Mat12]. Walls [Ber04b]. Walt [Smi07]. Walter [New07, Pol06, Usd22, vD03]. Waltham [Din07]. Walther [Die08]. Waltzing [Bis04]. War [Ale04, Ale09, Alo05, Ano08i, Aro06, Ass10, Ass23, Ava04, Bac09, Bak09, Bal08, Bal10, Ban06, Bar04, Bel09, Bel02, Ben21, Ber15, Bis04, Blu15, Boc06, Bod08, Boo03, Bor12b, BS02, Boy23, Bra16a, Bra05, Bra02a, Bre21, Bro02a, Bro04, Bur04, Cal14, Car99, Car13a, Car13b, Cha04, Cha07a, Cha14, Cha07b, Che08a, Che10, Chu07a, Clo06, Cog02a, Coh02, Coh08a, Coh08b, Con07, Con23, Cox02, Cri02, Cri10, Cri20a, Cul22, Cur15, Cur07, Dad19, Dan16a, Dan10, Dan11, Dan12a, Dan16b, Dar15, Dar16, Dei02, Dei14, Del23, Den06, Din23a, Din14b, Div00, Diz10, Doc05, Dom15a, Dom20, Don05, Duk11, Duk13, Eck14, Eck19, Eng08, Eng20b, Eva00, Fal05, Fel05, Fes07, For20, Fou14, Fri02, Fri10, Gaf07, Gal09, Gal15, Gal16b, Gar00a]. War [Gar13, Gar10b, Gas06, Gas00, Gav05, Gaw21, Gem18, Gen04, Gen17, Ger14, Gif05, Gil06, Gil23, Gle05, Gol06a, Gol12, Goo15, Goo17, Gos08, Gre08, Gro07, Guz13, Haa05, Haa13, Haf08a, Haf08c, Haj12, Ham15, Ham18, Har12b, Has07, Hay03, Hay15b, Hei16, Hei06, Hen09, Her17, Her14, HTP^{+}15, HGP^{+}23, Hes12, His13, Hit16, Hol21, Hor07, Hug08, Hus00, Hut01, Iac22, Iat18, Ing04a, Ing17, Jak16, JK07, Jer08, Jer16b, JLBK17, Joh03, Jon07, Jon14, KG11, Kal09, Kan12a, Kan12b, Kan15, Kap04, Kap10, Kar14b, Kar04, Kau09, Kim07, Kip03, Kir08, Kir09, Kle23, Kna02, Kra18, KNS18, Kru08, Kru05, Kur05a, Kur05b, Lar22, Lat14, Lau00, Leb00, Leb03, Lee12, Lee21, Lee23, Lei08, Lem14, Ler15, Lev04, Lev10b, Lew12, Lin15]. War [tL16, Lit11b, Liv07, Lom09, Lou23, Lov22, Low04, Łuk23, Luk17, Mac20a, Mac20b, Mad02, Mad04, Mad07, Mad08a, Mag14b, Man13, Mar18, Mar06a, Mar08a, Mar12, Mar07d, Mas00, Mas11b, Mas21b, Mas23, Mat10b, Mat04b, Mat20a, Mav18, McM05, McM10, McM08, McP19, Mel01, Mid11, Mil20, Mil15b, Mil15a, Mil17b, Mil04a, Mis23b, Mit07, Moe03, Moo23, Moy11, Mue05, Mul23, Mur07, Nai22, Nam13, Nic12, Nie12, Nie10b, Nin03, Nor15, Nue05, OB20, Osg04, Pai16, Par09, Pec08, Pem16, Pen06, Pfa02, Pit23, Poc03, Por09, Pra15a, Pra15b, Pra20, Pre11a, Pre17, Pud12, Qua04, Qua09, Qua13, Qua14, Rad16, Rak15, Ram07, Ram01, Ram08, Ram11a, Ram11b, Rat21, Rei16, Rep12, Rey23, Ric15, Rob04, Rob10, Rog15, Roh17, Ros10a, Ros17, Ros04]. War [Sal06, Sap03, Sar05a, Sar16, Sar05b, Sar11, Sas14b, Sau12, Sch08, Sch00a, Sch22a, Sch11, Sch22b, Sch03, Sch20, Sch17, Sch02c, Sco17, Sel02, Sfi01, Sfi02, Sha11, Sha06, Sha07b, Sha08b, She06, She12c, She13, Shm23, Sib11, Sin18, Sny03b, Sny11, Soa22, Sou21, Spe14, Spe15, Sta13, Sta15, Sta07, Sti10, Str12, Str09, Stu19, Suh08, Sza11, Tal99, Tay19, Thi14, Tho11, Tol10, Usd16, Val08, VA12, Vám10, Van99a, Van02a, Vau15, Vei14a, Vei14b, Vic08, Wal12a, Wal21b, Wal07, Wal13b, Wea12, Wea15, Wei04b, Wei01, Wei12, Wen09, Wet01, Whi02a, Whi02b, Whi03a, Whi11a, Whi12, Whi14a, Whi14b, Whi15, Wir03, Wir07a, Wol13, Woo13, Yor15, Yor19a, Yor20b, You12a, Zal04, Zha10, Zhi00, Zhu16, Zie21, Zie08, Żuk19, vD21]. war [Ald07, ARN16, AS22, BMLG11, Bar22, Bar07, Blu14, Bra16a, Bri13, Bro08, Buc17, Car11, Car17, Chu19, Cly13, Coh10, Cos22, Cri99, Cri04, Cri20b, Dad19, DeR17, Dee07, Dei14, Dei15, Dob10a, DFK^{+}20, Dou15, Dra12b, Duk11, Dum14, Eng03, Eng05, Ess15, Eva01, Fal20, Fav18, Fed17, FAM16, FM16, FP21, For05, Fox09, FF22, Fro23, Gaf11, Gal22, Gar04d, Geh22, Gel12b, Gho00, GK02, GSP^{+}19, Gol06b, Goo10, Göt08, GD99, Gue17, Hal16, Har19, Hat21, Hay00, HK20, Hel22, HF14, Her16, Her21, Her07b, Her11b, Her03b, Hop03, HT19, Iat05, Ing04b, Jen21, JE19, Jer01, JMG08, JM23, Joh01, Jos23, Kac02, Kap17b, Kap20, Kar14b, Kel19, Key16, Kho14, KSR18, KW09, Knü20, Kot17, Kou17, Kra06f]. War [Kra13b, KCG^{+}17, Kre07, Kre12b, Kre12a, Kuk13, Lat14, Lat16a, Law12, Law22, Lef12, Lei08, Lei18, Li18, LX23, Lin09, Lit10, Luk01, Luk07, Luk16, Lüt08, Lüt16, Lyn16, Mac06, Mad13, MBR^{+}21, Mak17, Mak12, Mal15, Man21, Mar13a, Mar02, Mar22, Mar07c, Mar23b, Mar15, Mas02, Mat16b, Mat20b, May13, McA22, McG06, McK08, McM06, Mei15, Mey11, Mic17, Mik12, Mil20, Mil22, Mil13, Mod17, Mue13, Mue16, Myl12, Neu03, Nut02b, O’M21, Odo06, Osg02, Par22, Pec13a, Pem16, Pip11, Pon01b, Pow16, Pre11a, QMLL16, Rad06, Raj14, Ray20, Rei19, Ria17, Ris14, RG15, Sac18, SS18a, San16, Sar23, Sch16, Sen16, Sha01, Sha02, Sha12, SY17, She20, Shi13, Sho15, Sny13, Soa13, Spe13, Sta16, Sur02, SJ23, Szo10, TW05, Tei11, Tho05]. War [Tho21, Tra09, Tro16a, TGS^{+}23, Ula99, Ver18, Vil16, Wal13b, Whe06, Whi13, Wir13, Woh99, Woh05, Wys18, Xia08b, XL23, Yar09, You12b, You22, Zak10, ZRS^{+}21, ZP23, Zim09, Żuk19, vW19, Ale07, Ass10, Cha08, Dan15, Dan22, Dom08, Dom09, Dom15b, Dre15, Dum17, Eck19, Eng20a, Eva23, Gil18, Gol07, Gol10, Gro07, Har08, Jos20, Kal22, Kuk21, Lat08, Law10, Lev10a, Lov10, Mar23a, Mar08b, Mas21a, Mat20a, McM14, McM21a, Pai17, Pay21, Pee23, Pet18, Pol06, Pra14, Rot21, Rub20, Sco09, Soa09, Soa20a, Stu15, Vau15, Wea15, Whi08b, Yor20a, Yor23a, vDO08, BMLG11, Bar03b, Boo03, Cat05, Cle04, Coh02, Cro00, Fie03, Fri05, Hac03, Her03a, Hes02, Jer00, Kle04, Kna05, Law05, Law11, Lea01, Lee06b, Lib01, Mil01, Moï11, Nau04, Nin03]. War [Osg04, PW05, Poc03, Ros00, Spe01, Sta00, Tof05, Tur01, Whi06]. War-Era [Gho00, QMLL16]. War-Weariness [Sta15]. Ward [Ryd14]. Warfare [Bus14, Dom15a, Eth05, Gil06, Hei16, Imm16, Lei08, Wal12a, Rem08, Wir07b, Sha20]. Warfighting [Dre05]. Warner [Her20a]. Warren [Iri14, Pos07, Eng07]. Warrior [Bac09, Her19, Hay15a, Rie15]. Warriors [Bai18, Pol07a, Pra20, Ric07, Rub20, Win07]. Wars [Cla09, Edw16, Fin04, Haf08a, Kob06, Kop16, Law01, Lef21, Lit11a, Low06, Lun04, McA14, Moï10, Pec08, Rus04a, Tra02, Vu19, Hay13, Log01, And10, Haf08c, Bla05, Wir06]. Warsaw [And20, Joh20a, Lie21, Pol13, Bae16, Bék23, Yor19b, Cie07]. Warschauer [Bis11, Wet11a, Wet11b]. Warth [Die08]. Wartime [Dro18, Eng09, Ess15, Ram11b]. Was [Bor05b, Čav19, Gei12, Ger13, HTP^{+}15, Jer01, Kaź09, Mas09, Mel01, Nei10a, Sfi02, Usd20, Vuj02, dS19]. Washington [Byr12, Car11, Din13, Dom03, Dra12b, For19, Hes12, Joh07c, Kan12b, Kan15, Law05, Lou16, Mas21a, New07, Pod14, Rei16, Sim17, Smi11, Spo12, And23, Coh08a, Kam17, Luk17]. Washingtons [Kan14]. Wasif [Dre21]. Wasn’t [Mil20]. Wasserstrom [Nei10b]. Waste [Nor05, Sou21]. Wataru [Ess15]. Watched [Fei10]. Water [SL23]. Watry [Dei15]. Way [Bra14, Bru08, Cat05, Ran02, Rep12, Soa20a]. Wayne [Neu11]. Ways [Ava04, Leb00]. Weak [LSvdV04]. Weapon [Che08b, Gra21a, Sha20, Sto15]. Weapons [Boz20a, Bur10, Bur21, Car17, Eng07, JLBK17, Joh20b, Lei18, McM12a, Mid00, Roh17, Sch01, Ste14, Wal21b, Zha23b, Bay01]. Weariness [Sta15]. Wechselbeziehungen [Huh20]. Weg [Kra01, Kur21, She12a]. Wegener [Las19]. Weges [Bru08]. Wehrbeauftragte [Her08]. Wehrmacht [Die08]. Wehrmachtgeneral [Die08]. Weight [Gra21b, TSW^{+}22]. Weimer [Cra13]. Weingarten [Whi11b]. Weir [Pol06]. Weissmann [Mil17a]. Weitra [Bis11]. Welfare [Gil06]. Well [Tor22]. Wells [Wal08]. Welt [Bru08]. Weltmeinung [Moe03]. Wendy [vH09]. Wenger [Dav10, Hen09]. Wengst [Bru08]. Were [Kra04a]. Werenskjold [Vil16]. Werner [Han13].

Wernher [Neu11]. West [Bae16, Ban16, Bar08, Bru05, Hay11, Ing17, Kan12b, Ken13, Kop16, Lee17, Lud13, Por09, Rai11, Sch05, Soa09, Sol13, Ste02, Ste10, Tho11, Bus14, Cal22, Con12, Cos22, CK13, Geh05, Geh22, Gut23, Her09, JE19, Kom18, Lar11, Lut22b, Mad13, Mue11, Pow16, Rub02, Sar19, Uel20, Eng09, Pol07b, Chi04]. West-Konfrontation [Bru05]. Westad [Nie12, Hol21, Nai22]. Westbindung [Moe03]. Western [Ald08, Che99, Duk11, Ger14, Gra13b, Hus05, Kap04, Kna02, Loc05, Mag10, Mey11, Mil04b, Nie10a, Pon13, Sha07b, Shm23, Whi14a, Win07, vC13, BS02, Cri99, McK08, Rea11, SM12, Svi16, Wal03, Wys18, Fou14, Mat04a, Rus05]. Westernism [AE02]. Westlake [Bla03]. westliche [Nue06]. Westminster [Gaf11]. Westmoreland [Pre17]. Westpolitik [Lew12, Sar16]. Westport [Fer07]. Westview [Gru07]. Wettig [Ado08, Por09, Rug16c]. Wevill [Dei14]. Wheels [Gar04b]. Whitby [Bra16a]. White [KG11, Kas16, Mad08b, Mar08a, Ped07, Pos07, Rub20]. Whitehall [Cor14, Rey02]. Whither [Kom14]. Whittaker [Whi10]. Who [Bar11a, Bra14, Buh15, Goo05, Har18, Kra06a, Nor15, Pra20, Rub20, San16, Sza11]. Whom [Goo05]. Widerstandskämpfer [Die08]. Wiebke [Ste20]. Wiederaufbau [Wen09, Bis08]. Wiegrefe [Haf08b]. Wierzbicki [Ans17]. Wieviel [Kar14a]. Wiggershaus [Sza07]. Wilford [Bal15, Bel11]. Wilfried [Duk10, Gri10, Lie08]. Wilke [She12a]. Will [Kap17a, Kel19, Kor06, Lee21]. Willbanks [Ale07]. William [Che19, Gar00a, Low06, Mag14b, Mar07a, Mil21, Pee23, Pra19, The15, Yor23a, Coh06, HK20, Hol99, Hol05a, Joh07d, Liv07, Lou16, Qua13, Sol13, Wat18, Xia08c]. Williamson [Kir16]. Wills [Kau09]. Wilson [Byr12, Car11, Doc08, Dra12b, Hes12, Kan12b, Kan15, Kur05a, Pod14, Rei16, Spe18, Tho11, vDO08, Kru07a, Mad15, Vic08]. Winand [Cog12, Par19]. Window [Ros05]. Windrow [Nor08a]. Winfried [Sza07, Wet11a, Wet11b, Jon06]. Wing [Cro09]. Wings [Zal04]. Winning [Sin18, Lee06b]. Winslow [Din13]. Winston [Ram11b]. Wirklichkeit [Rug16b]. Wirkungen [Mil17a]. Wirtschaft [Rug17]. wirtschaftliche [Huh20]. Wise [Iri02]. Withdrawal [BJ10, Gra12b, Gra12a, Kal09, Pra99, Van99b]. Witherspoon [Soa20a]. within [Bae16, Kra03a, Kra04b, Kra05a, Nat04]. Without [Njø02, BMLG11, Clo09, Kap17a, Law11, Moï11, Oye14]. Witness [Kag05, Mar06a]. Witnesses [Bra16a]. Wittner [Gly06]. Wizards [Whi03b]. Wladimir [Nai10b]. Wohlstetter [Abe21]. Wojciech [Clo09]. Wojnowski [Kuz19]. Wolfe [Mic08, Sch20]. Wolfgang [Pra20, Cra07, Wil08a]. Wolfram [Neh11]. Wolna [Joh17, Joh21]. Woman [Bra16a, Woo07]. Women [Alo05, Ber15, Fes07, How14, Moo23, Soa23, War08, Spe14, Spe15]. Won [Lat14, Moy11]. Wood [Hay99]. Woodbridge [Gol11, Nor15]. Woodman [Gal15]. Woodroofe [Yor15]. Woodrow [Car11, Dra12b, Hes12, Kan12b, Kan15, Kur05a, Pod14, Rei16]. Woods [Hes21, Lew07, TJ21]. Woodward [Pav19]. Words [Fal05]. Work [Mol16]. Worker [Bri13]. Workers [Con07, Mat15]. Working [ST17]. World [Abe21, AS22, Bac09, Bal08, Bar03b, Boc10, Boy23, Bre21, Bro02a, Bru08, Bur04, Cal09, Cal14, Cha00, Clo06, Coh02, Coh08a, Col19, Con07, Cra13, DF12, Dei14, Del23, Die07, Din14b, Doc08, Dou15, Duk11, Duk13, Dum20, Ell09, Fal05, Fes07, FAM16, Gav05, Gly06, Gra21a, Gro07, Ham18, Has07, Hay13, Hol21, Hor07, JK07, Jos20, Kan12a, Kip03, Lat08, Lat14, Lee06b, Lei08, Ler02, Liv07, Lüt12, Mac20a, Mar23a, Mas23, Mat08, Mat10b, McM05, McM16, McM22, Mis23b, Nai22, Nin03, Pec08, Pec13a, Pet20, Rad12a, Rog15, Ros08, Rot21, Rut12a, San16, Sch00a, Sha08b, She12c, She13, Sin18, Sny03b, Sta13, Tei11, Tha06, Tho21, Vu19, Wal12b, Wan16, Wat18, Wei04b, Yar09, Yor18, dS19, Rom17, Bor12b, BS02, Bru11, Bur19]. World [FTRP17, Lar10, Mil15b, Rob04, Usd16, Wol13, ŽČ16, SÃņ23, God14, JK07, Joh22, Sin18, Yor19a, Fri10, Hur05, Kap10, Lev10b, Mat04a, Rob10, Rua04, Sti10, Vám10, ZRS^{+}21]. Worlds [Mis23a, Bug15]. Worse [Kru04a, Coh12]. Would [Sch09, Mas23]. Wound [Bro04]. Wright [Gho00]. Writers [Rub20]. Writing [Ger14, Hay00]. Writings [Kau02, Kho14, Sul22]. Wrocław [Joh21]. Wrong [Bon05, Par22]. Wurlitzer [Bel11].

x [Jon14, Kre12b, Kre12a, You12b, Zha12b]. Xi [Dre21, Pec13a, Rut12a, Zhu16]. Xia [Coh08b]. Xiaobing [Lin15]. Xiaoming [Rad16]. Xiaoping [Rad16, Del22, FTRP17, Zha10]. Xiaoyuan [Ros09]. xii [Boy11, Hay13, Kan12a, Lew12, Pon13, Ran12, GP11]. xiii [Ass14, Pra99]. Xinjiang [Jac16, Jac23]. xiv [Car13b, Gra13a, Jon07, Kan12b, Kan15, Sen16]. xix [Hay16, Neu11]. Xuan [Sza11]. xv [Din14b, Ket14, You13b]. xvi [Dra12b, Goo15, Goo17, Hat11, Mat16b, Soa13]. xvii [Owe07, Taa13, vD12]. xviii [Han13]. XXe [Mar13b]. xxiv [Bro11, Gal16a, Mag14b]. xxix [Kip15]. xxv [Pre17]. xxvi [Moi14]. xxx [Hes13]. xxxii [vD21].

Yaacov [Gol10]. Yafeng [Coh08b]. Yale [Bar11a, BSK^{+}23, Din14a, Ger13, Gra13b, GP11, Hay11, Kip15, Kur11, Lat14, Len12, Mal15, vC13, Diz10]. Yalta [Lau00, Fet20, Kim11, Rob07]. Yalu [Zal04]. Yanek [Wan16]. Yaqub [Joh07b, Qua21]. Yaroslav [Sny06]. Year [Ass23, Bro04, Kra13a]. Years [BL10a, BL10b, Bra16b, Byr12, Che07, Duk10, Guz13, Har11, Lin09, Lud13, Mas21b, Mat04a, Mat08, Nor00, Par19, Van02b, Xia08b, Yon08, Dar15, Dar16, Mad07, Ada03, Hay03, Jon04]. Yelena [Woo12]. Yellin [Fes07]. Yellow [Lee20]. Yeltsin [Rut06, Rut11, Zlo03, Bae02, Ell03, Her02, Roz03]. Yemen [Fer08]. Yes [Mil13]. Yezhov [Kul10]. Yield [Gra21a]. Yildirim [Myl15]. Yinghong [yL12]. Yoram [McD06]. York [Bal15, Bar07, Bel13, Bla15, Blu15, Bor12a, Bru11, Bur16, Cal14, Car13b, Chi14, Cra17, Cul13, Dan12a, Dan12b, Dan16b, Dar15, Dar16, Dav11a, Dei14, Dom15a, Dou15, Dre13, Eva13, Fes07, Fin13, Gal16a, Gra13a, Guz13, Han13, Har11, Hay16, Her16, Her07b, Hes13, Hol15, Hou16, Jer11, Jer16b, Joh11a, KG11, Kan12a, Ket14, Kim11, Kir16, Kop16, Kot17, Kru07b, Lat16a, Lat16b, Lee21, Lef12, Mad15, Mag14a, Mar12, McM12a, McM16, Moi14, Nai16, Neh11, Nei14, Neu11, Par13, Par16, Pec13a, Pon13, Pos07, Pre11b, Pre17, Qua13, Rad11, Rai11, Ram11a, Rat21, Rep12, Rob14, Rub11, Rus16, Rut12a, Rut11, San16, Sar16, Sar11, Sas14b, Sen16, Ser13, She12c, She13, Sib11, Sny17, Soa13, Spe14, Sur13, Sza12a, Sza21, Sza11, Usd11, Usd16, Vuj02]. York [Wal12a, WF12, Whi16, Whi14a, Whi15, Zha12b, vD12, Hay15b]. Yorkshire [Duk11]. Yoshiko [Sak06]. Young [Dav11a, Her20a, Moï10, Nei10b, Sza10b, Joh22, Mil15a, San16]. Youth [Car13b, Koz14, Pik02]. yu [Lev07]. Yuen [Luc07]. Yugoslav [Raj14]. Yugoslavia [Lee17, Lee23, Sil04, Čav19, Nie11, Nie21, ŽČ16, Woo07, Vuj02]. Yui [Sha11]. Yui-Sang [Sha11]. Yukinori [Lin11]. Yuniu [Yos15]. Yurchak [Nei10a]. yuzhnykh [Min19].

Z [Gus10]. za [Min19, Tsy08]. Zachary [McM16, Sul22]. Zachmann [His13]. zai [yL07]. Zaïre [Pow16]. zarista [Gra07]. Zarnow [Zar21]. Zarusky [Nai08a]. Zaslavsky [Dra12b]. Zatlin [Loe10]. Zbig [Ste15]. Zbigniew [Kuz19, Ste15]. Zbynek [Luň10]. Zealand [Gar02, Tra09]. Zecchini [Del23]. Zedong [Dre21]. Zelikow [Alp09]. Zeman [Luň10]. Zero [Sto15]. Zerreissprobe [Nue06]. Zerwürfnis [Haf08b]. Zhang [Pen06, Rei16, Rad16]. zhenskoe [Bra16a]. Zhihua [yL07]. Zhivago [McM12b, Rub10]. Zhivkov [Neh22]. Zhongguo [yL07]. zhuanjia [yL07]. Ziemann [Mey11]. Zierler [Bra14]. Zimbabwe [How22]. Zinky [Bra16a]. Zionism [Sur13]. Zionismus [Ste20]. Zionist [Pip15, Her21]. Ziva [Pip15]. Zivile [Kar14a]. Zlatko [Nei07]. złoty [Joh21]. Zone [Mas16, Rya08a, Wet09, Win11]. Zone/German [Wet09]. Zones [Sto15]. Zubok [McM12b, Par09, Rub10]. zum [Bru09, Kra01, Wet11a, Wet11b]. zur [Bru05, Gri10, Kan10, Kar14a, Kur21, She12a]. Zusammenbruch [Wet11a, Wet11b]. Zusammenwachsen [Jon06]. Zwei [Rug18]. Zwei-plus-Vier-Prozess [Rug18]. Zwischen [Shl06, Huh20, War06, Bis11]. zwykłe [Joh17].

References

Abella:2021:CWT

[Abe21] Alex Abella. The Cold World They Made: The Strategic Legacy of Roberta and Albert Wohlstetter by Ron Robin, Cambridge, MA: Harvard University Press, 2016. 365 pp. $35.00. (review). Journal of Cold War Studies, 23(1):228–229, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

        edu/jcws/article-pdf/23/1/228/1911812/jcws_r_00986.pdf;
        https://muse.jhu.edu/article/788394.

Adamthwaite:2003:NAF

[Ada03] Anthony P. Adamthwaite. NATO After Fifty Years (review). Journal of Cold War Studies, 5(2):115–117, Spring 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40752;   http://muse.jhu.edu/
article/40752/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/115/700408/jcws.2003.5.2.115.pdf.

Adler:2003:RSD

[Adl03] Emanuel Adler. Resolving Security Dilemmas: a Constructivist Explanation of the INF Treaty (review). Journal of Cold War Studies, 5(2): 106–109, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/40753;
 http://muse.jhu.edu/article/40753/pdf;   https://direct.
mit.edu/jcws/article-pdf/5/2/106/700406/jcws.2003.5.2.106.
pdf.

Adomeit:2008:HMH

[Ado08] Hannes Adomeit. Hope M. Harrison, Driving the Soviets up the Wall: Soviet-East German Relations, 1953–1961; and Gerhard Wettig, Chruschtschows Berlin-Krise 1958 bis 1963: Drohpolitik und Mauerbau. Journal of Cold War Studies, 10(1):129– 131, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/231567;
http://muse.jhu.edu/article/231567/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/1/129/700073/jcws.2008.10.1.
129.pdf.

Azadovskii:2002:AWA

[AE02] Konstantin Azadovskii and Boris Egorov. From anti-Westernism to anti-Semitism: Stalin and the impact of the “anti-cosmopolitan” campaigns on Soviet culture. Journal of Cold War Studies, 4(1):66–80, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9219;   http:/
/muse.jhu.edu/article/9219/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/1/66/695508/152039702753344834.pdf.

Aftergood:2002:SBO

[Aft02] Steven Aftergood. In the Shadow of the Bomb: Oppenheimer, Bethe, and the Moral Responsibility of the Scientist (review). Journal of Cold War Studies, 4(1):95–96, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9218;   http://muse.jhu.edu/article/
9218/pdf;   http://muse.jhu.edu/journals/journal_of_cold_
war_studies/v004/4.1aftergood.html;   https://direct.mit.
edu/jcws/article-pdf/4/1/95/695583/jcws.2002.4.1.95.pdf.

Aftergood:2003:MJR

[Aft03] Steven Aftergood. In the Matter of J. Robert Oppenheimer: The Security Clearance Hearing (review). Journal of Cold War Studies, 5 (3):122–124, Summer 2003. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
45179;   http://muse.jhu.edu/article/45179/pdf;   http:/
/muse.jhu.edu/journals/journal_of_cold_war_studies/v005/
5.3aftergood.html;   https://direct.mit.edu/jcws/article-
pdf/5/3/122/695883/jcws.2003.5.3.122.pdf.

Aftergood:2013:KMA

[Aft13] Steven Aftergood. Kenneth Michael Absher, Michael C. Desch, and Roman Popadiuk, Privileged and Confidential: The Secret History of the President’s Intelligence Advisory Board. Lexington: University Press of Kentucky, 2012. 515 pp. $39.95. Journal of Cold War Studies, 15(2):143–145, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/509711;   http://muse.jhu.edu/article/509711/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/2/143/698173/
jcws_r_00346.pdf.

Aftergood:2020:PCS

[Aft20] Steven Aftergood. Poisoner in Chief: Sidney Gottlieb and the CIA Search for Mind Control by Stephen Kinzer (review). Journal of Cold War Studies, 22(1):243–245, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/22/1/243/699852/jcws_r_
00932.pdf;   https://muse.jhu.edu/article/749649.

Akiner:2002:MTC

[Aki02] Shirin Akiner. Mongolia in the Twentieth Century: Landlocked Cosmopolitan (review). Journal of Cold War Studies, 4(3):162–163, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9266;   http:/
/muse.jhu.edu/article/9266/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/162/700270/jcws.2002.4.3.162.pdf.

Akturk:2010:FGP

[Akt10] Sener Akturk. Frank Grüner, Patrioten und Kosmopoliten: Juden im Sowjetstaat 1941–1953. Journal of Cold War Studies, 12(4):205–207, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/403813;   http:

//muse.jhu.edu/article/403813/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/4/205/697585/jcws_r_00065.pdf.

Aldrich:2007:JLC

[Ald07] Richard J. Aldrich. Julian Lewis, Changing Direction: British Military Planning for Post-war Strategic Defence, 1942–47. Journal of Cold War Studies, 9(1):134–135, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210068;   http://muse.jhu.edu/article/210068/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/134/696552/
jcws.2007.9.1.134.pdf.

Aldrich:2008:PMS

[Ald08] Richard J. Aldrich. Paul Maddrell, Spying on Science: Western Intelligence in Divided Germany, 1945–61. Journal of Cold War Studies, 10(4):165–167, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/252669;   http://muse.jhu.edu/article/252669/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/4/165/697059/
jcws.2008.10.4.165.pdf.

Alexopoulos:2001:HSU

[Ale01] Golfo Alexopoulos. A History of the Soviet Union from the Beginning to the End (review). Journal of Cold War Studies, 3(2):111– 113, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9184;   http:/
/muse.jhu.edu/article/9184/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/2/111/695386/jcws.2001.3.2.111.pdf.

Alexander:2004:DCR

[Ale04] Nathan Alexander. El Dorado Canyon: Reagan’s Undeclared War with Quaddafi (review). Journal of Cold War Studies, 6 (3):139–141, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/171146;   http://muse.jhu.edu/article/171146/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/3/139/696055/
jcws.2004.6.3.139.pdf.

Alexander:2007:JHW

[Ale07] Nathan Alexander. James H. Willbanks, Abandoning Vietnam: How America Left and South Vietnam Lost Its War. Journal of Cold War Studies, 9(1):115–117, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210069;   http://muse.jhu.edu/article/210069/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/115/696479/
jcws.2007.9.1.115.pdf.

Alexopoulos:2008:SSM

[Ale08] Golfo Alexopoulos. Simon Sebag Montefiore, Stalin: The Court of the Red Tsar; and Robert Service, Stalin: a Biography. Journal of Cold War Studies, 10(1):132–136, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231568;   http://muse.jhu.edu/article/231568/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/132/700037/
jcws.2008.10.1.132.pdf.

Alexander:2009:MMT

[Ale09] Nathan Alexander. Mark Moyar, Triumph Forsaken: The Vietnam War 1954–1965. Journal of Cold War Studies, 11(4):167– 170, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/366250;
http://muse.jhu.edu/article/366250/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/4/167/1627641/jcws.2009.11.4.
167.pdf.

Alexopoulos:2012:SBD

[Ale12] Golfo Alexopoulos. Steven A. Barnes, Death and Redemption: The Gulag and the Shaping of Soviet Society. Princeton, NJ: Princeton University Press, 2011. 352 pp. $35.00. Journal of Cold War Studies, 14(2):173–175, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/478789;   http://muse.jhu.edu/article/478789/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/2/173/698007/
jcws_r_00241.pdf.

Alker:2005:PIR

[Alk05] Hayward R. Alker. Progress in International Relations Theory: Appraising the Field (review). Journal of Cold War Studies, 7(3):147–149, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185093;   http://muse.jhu.edu/article/185093/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/147/1102549/
1520397054377197.pdf.

Allcock:2014:FAP

[All14] Thomas Tunstall Allcock. The first Alliance for Progress?: Reshaping the Eisenhower Administration’s policy toward Latin America. Journal of Cold War Studies, 16(1):85–110, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543726;   http:
//muse.jhu.edu/article/543726/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/1/85/698908/jcws_a_00432.pdf.

Allen:2015:RMH

[All15] David Allen. Realism and malarkey: Henry Kissinger’s State Department, détente, and domestic consensus. Journal of Cold War Studies, 17(3):184–219, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595069;   http://muse.jhu.edu/article/595069/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/184/
1102407/jcws_a_00548.pdf.

Alonso:2005:CWW

[Alo05] Harriet Hyman Alonso. Cold War Women: The International Activities of American Women’s Organizations (review). Journal of Cold War Studies, 7(3):165–167, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185094;   http://muse.jhu.edu/article/185094/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/165/1102573/
jcws.2005.7.3.165.pdf.

Alpers:2009:ERM

[Alp09] Benjamin L. Alpers. Ernest R. May and Philip D. Zelikow, eds., Dealing with Dictators: Dilemmas of U.S. Diplomacy and Intelligence Analysis, 1945–1990. Journal of Cold War Studies, 11(4): 155–157, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/366251;
http://muse.jhu.edu/article/366251/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/4/155/1627625/jcws.2009.11.4.
155.pdf.

Anderson:2005:MGS

[And05] Richard D. Anderson, Jr. Michael S. Gorham, Speaking in Soviet Tongues: Language Culture and the Politics of Voice in Revolutionary Russia. Journal of Cold War Studies, 7(1):203–206, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/179550;   http:
//muse.jhu.edu/article/179550/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/1/203a/700490/jcws.2005.7.1.203a.
pdf.

Andrews:2010:EPS

[And10] James T. Andrews. Ethan Pollock, Stalin and the Soviet Science Wars. Journal of Cold War Studies, 12(1):195–198, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382490;
http://muse.jhu.edu/article/382490/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/1/195/1627455/jcws.2010.12.1.
195.pdf.

Anderson:2020:LCW

[And20] Sheldon Anderson. Laurien Crump, The Warsaw Pact Reconsidered: International Relations in Eastern Europe, 1955–69. Journal of Cold War Studies, 22(3):262–264, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/3/262/1860632/jcws_r_00959.pdf;
https://muse.jhu.edu/article/762903.

Anderson:2023:BRK

[And23] Sheldon Anderson. Book review: Kentomania: a Black Basketball Virtuoso in Communist Poland by Kent Washington. Journal of Cold War Studies, 25(2):255–257, Spring 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/2/255/2140952/jcws_r_01150.pdf.

Anonymous:1999:Aa

[Ano99a] Anonymous. Abstracts. Journal of Cold War Studies, 1(1):iii–iv, Winter 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/1/
1/iii/695117/jcws.1999.1.1.iii.pdf.

Anonymous:1999:Ab

[Ano99b] Anonymous. Abstracts. Journal of Cold War Studies, 1(2):??, Spring 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/1/
2/iii/695166/jcws.1999.1.2.iii.pdf.

Anonymous:1999:Ac

[Ano99c] Anonymous. Abstracts. Journal of Cold War Studies, 1(3):??, Fall 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/1/3/
iv/695180/jcws.1999.1.3.iv.pdf.

Anonymous:1999:Ca

[Ano99d] Anonymous. Contributors. Journal of Cold War Studies, 1(1):ii, Winter 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9090;   https:/
/direct.mit.edu/jcws/article-pdf/1/1/ii/695115/jcws.1999.
1.1.ii.pdf.

Anonymous:1999:Cb

[Ano99e] Anonymous. Contributors. Journal of Cold War Studies, 1(2):ii, Spring 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9099.

Anonymous:1999:Cc

[Ano99f] Anonymous. Contributors. Journal of Cold War Studies, 1(2):??, Spring 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/1/
2/ii/695162/jcws.1999.1.2.ii.pdf.

Anonymous:1999:Cd

[Ano99g] Anonymous. Contributors. Journal of Cold War Studies, 1(3):iii, Fall 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/9111.

Anonymous:1999:Ce

[Ano99h] Anonymous. Contributors. Journal of Cold War Studies, 1(3): ??, Fall 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/1/
3/iii/695174/jcws.1999.1.3.iii.pdf.

Anonymous:1999:GC

[Ano99i] Anonymous. Guidelines for contributors. Journal of Cold War Studies, 1(3):185, Fall 1999. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/1/3/185/695176/jcws.1999.1.3.185.pdf.

Anonymous:2000:Aa

[Ano00a] Anonymous. Abstracts. Journal of Cold War Studies, 2(1):??, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/2/1/
iv/695224/jcws.2000.2.1.iv.pdf.

Anonymous:2000:Ab

[Ano00b] Anonymous. Abstracts. Journal of Cold War Studies, 2(2):??, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/2/
2/iv/695253/jcws.2000.2.2.iv.pdf.

Anonymous:2000:Ac

[Ano00c] Anonymous. Abstracts. Journal of Cold War Studies, 2(3):??, Fall 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/2/3/
iv/695315/jcws.2000.2.3.iv.pdf.

Anonymous:2000:Ca

[Ano00d] Anonymous. Contributors. Journal of Cold War Studies, 2(1):iii, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9123;   http:
//muse.jhu.edu/article/9123/pdf.

Anonymous:2000:Cb

[Ano00e] Anonymous. Contributors. Journal of Cold War Studies, 2(1):??, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/2/
1/iii/695220/jcws.2000.2.1.iii.pdf.

Anonymous:2000:Cc

[Ano00f] Anonymous. Contributors. Journal of Cold War Studies, 2(2):iii, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9136;   http:
//muse.jhu.edu/article/9136/pdf.

Anonymous:2000:Cd

[Ano00g] Anonymous. Contributors. Journal of Cold War Studies, 2(2):??, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/2/
2/iii/695247/jcws.2000.2.2.iii.pdf.

Anonymous:2000:Ce

[Ano00h] Anonymous. Contributors. Journal of Cold War Studies, 2(3): iii, Fall 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9153;   http:
//muse.jhu.edu/article/9153/pdf.

Anonymous:2000:Cf

[Ano00i] Anonymous. Contributors. Journal of Cold War Studies, 2(3):??, Fall 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/2/3/
iii/695313/jcws.2000.2.3.iii.pdf.

Anonymous:2000:ENa

[Ano00j] Anonymous. Editor’s note. Journal of Cold War Studies, 2(1):1– 2, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9130;   http:/
/muse.jhu.edu/article/9130/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/1/1/695183/15203970051032354.pdf.

Anonymous:2000:ENb

[Ano00k] Anonymous. Editor’s note. Journal of Cold War Studies, 2(2):1– 2, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9141;   http:/
/muse.jhu.edu/article/9141/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/2/1/695210/15203970051032282.pdf.

Anonymous:2000:ENc

[Ano00l] Anonymous. Editor’s note. Journal of Cold War Studies, 2(3):1– 3, Fall 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9158;   http:/
/muse.jhu.edu/article/9158/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/3/1/695272/15203970051032183.pdf.

Anonymous:2000:E

[Ano00m] Anonymous. Erratum. Journal of Cold War Studies, 2(3):149, Fall 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/2/3/
149/695303/jcws.2000.2.3.149.pdf.

Anonymous:2000:GCa

[Ano00n] Anonymous. Guidelines for contributors. Journal of Cold War Studies, 2(1):146, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/2/1/146/695218/jcws.2000.2.1.146.pdf.

Anonymous:2000:GCb

[Ano00o] Anonymous. Guidelines for contributors. Journal of Cold War Studies, 2(2):133, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/2/2/133/695226/jcws.2000.2.2.133.pdf.

Anonymous:2000:GCc

[Ano00p] Anonymous. Guidelines for contributors. Journal of Cold War Studies, 2(3):150, Fall 2000. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/2/3/150/695308/jcws.2000.2.3.150.pdf.

Anonymous:2001:Aa

[Ano01a] Anonymous. Abstracts. Journal of Cold War Studies, 3(1):iv–??, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/3/
1/iv/695454/jcws.2001.3.1.iv.pdf.

Anonymous:2001:Ab

[Ano01b] Anonymous. Abstracts. Journal of Cold War Studies, 3(2):??, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/3/
2/iv/695394/jcws.2001.3.2.iv.pdf.

Anonymous:2001:Ac

[Ano01c] Anonymous. Abstracts. Journal of Cold War Studies, 3(3):??, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL https://direct.mit.edu/jcws/article-pdf/3/3/v/695482/
jcws.2001.3.3.v.pdf.

Anonymous:2001:BRa

[Ano01d] Anonymous. Book reviews. Journal of Cold War Studies, 3(1):126– 148, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/3/
1/126/695414/15203970151032173.pdf.

Anonymous:2001:BRb

[Ano01e] Anonymous. Book reviews. Journal of Cold War Studies, 3(2):101– 116, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/3/
2/101/695380/152039701300373907.pdf.

Anonymous:2001:Ca

[Ano01f] Anonymous. Contributors. Journal of Cold War Studies, 3(1):iii, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9170;   http:
//muse.jhu.edu/article/9170/pdf.

Anonymous:2001:Cb

[Ano01g] Anonymous. Contributors. Journal of Cold War Studies, 3(1):iii, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/3/
1/iii/695450/jcws.2001.3.1.iii.pdf.

Anonymous:2001:Cc

[Ano01h] Anonymous. Contributors. Journal of Cold War Studies, 3(2):iii, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9186;   http:
//muse.jhu.edu/article/9186/pdf.

Anonymous:2001:Cd

[Ano01i] Anonymous. Contributors. Journal of Cold War Studies, 3(2):??, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/3/
2/iii/695396/jcws.2001.3.2.iii.pdf.

Anonymous:2001:Ce

[Ano01j] Anonymous. Contributors. Journal of Cold War Studies, 3(3): iv, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9198;   http:
//muse.jhu.edu/article/9198/pdf.

Anonymous:2001:Cf

[Ano01k] Anonymous. Contributors. Journal of Cold War Studies, 3(3):??, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/3/3/
iv/695478/jcws.2001.3.3.iv.pdf.

Anonymous:2001:ENa

[Ano01l] Anonymous. Editor’s note. Journal of Cold War Studies, 3(1):1– 3, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9176;   http:/
/muse.jhu.edu/article/9176/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/1/1/695400/15203970151032128.pdf.

Anonymous:2001:ENb

[Ano01m] Anonymous. Editor’s note. Journal of Cold War Studies, 3(2):1– 2, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9189;   http:/
/muse.jhu.edu/article/9189/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/2/1/695370/152039701300373853.pdf.

Anonymous:2001:ENc

[Ano01n] Anonymous. Editor’s note. Journal of Cold War Studies, 3(3):1– 3, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9204;   http:/
/muse.jhu.edu/article/9204/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/1/695428/152039701750419484.pdf.

Anonymous:2001:GC

[Ano01o] Anonymous. Guidelines for contributors. Journal of Cold War Studies, 3(1):149, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/3/1/149/695446/jcws.2001.3.1.149.pdf.

Anonymous:2002:Aa

[Ano02a] Anonymous. Abstracts. Journal of Cold War Studies, 4(2):??, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL https://direct.mit.edu/jcws/article-pdf/4/2/v/695615/
jcws.2002.4.2.v.pdf.

Anonymous:2002:Ab

[Ano02b] Anonymous. Abstracts. Journal of Cold War Studies, 4(3):??, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/4/3/v/
700272/jcws.2002.4.3.v.pdf.

Anonymous:2002:Ac

[Ano02c] Anonymous. Abstracts. Journal of Cold War Studies, 4(4):??, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL https://direct.mit.edu/jcws/article-pdf/4/4/v/695738/
jcws.2002.4.4.v.pdf.

Anonymous:2002:Ad

[Ano02d] Anonymous. Abstracts. Journal of Cold War Studies, 4(4):??, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/5/1/
iii/695794/jcws.2003.5.1.iii.pdf.

Anonymous:2002:Ca

[Ano02e] Anonymous. Contributors. Journal of Cold War Studies, 4(1):iv, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9224;   http:
//muse.jhu.edu/article/9224/pdf.

Anonymous:2002:Cb

[Ano02f] Anonymous. Contributors. Journal of Cold War Studies, 4(2):iv, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9245;   http:
//muse.jhu.edu/article/9245/pdf.

Anonymous:2002:Cc

[Ano02g] Anonymous. Contributors. Journal of Cold War Studies, 4(2):??, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/4/
2/iv/695605/jcws.2002.4.2.iv.pdf.

Anonymous:2002:Cd

[Ano02h] Anonymous. Contributors. Journal of Cold War Studies, 4(3):iv, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9274;   http:/
/muse.jhu.edu/article/9274/pdf.

Anonymous:2002:Ce

[Ano02i] Anonymous. Contributors. Journal of Cold War Studies, 4(3):iv, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9270;   http:/
/muse.jhu.edu/article/9270/pdf.

Anonymous:2002:Cf

[Ano02j] Anonymous. Contributors. Journal of Cold War Studies, 4(3):??, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/4/
3/iv/700264/jcws.2002.4.3.iva.pdf.

Anonymous:2002:Cg

[Ano02k] Anonymous. Contributors. Journal of Cold War Studies, 4(4):iv– vi, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9294;   http:
//muse.jhu.edu/article/9294/pdf.

Anonymous:2002:Ch

[Ano02l] Anonymous. Contributors. Journal of Cold War Studies, 4(4):??, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/4/4/
iv/695722/jcws.2002.4.4.iv.pdf.

Anonymous:2002:Ci

[Ano02m] Anonymous. Contributors. Journal of Cold War Studies, 4(4): ??, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/5/
1/ii/695781/jcws.2003.5.1.ii.pdf.

Anonymous:2002:ENa

[Ano02n] Anonymous. Editor’s note. Journal of Cold War Studies, 4(1):1– 2, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9231;   http:/
/muse.jhu.edu/article/9231/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/1/1/695506/152039702753344780.pdf.

Anonymous:2002:ENb

[Ano02o] Anonymous. Editor’s note. Journal of Cold War Studies, 4(2):1– 2, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9248;   http:/
/muse.jhu.edu/article/9248/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/2/1/695538/152039702753649610.pdf.

Anonymous:2002:ENc

[Ano02p] Anonymous. Editor’s note. Journal of Cold War Studies, 4(3):1–2, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9279;   http:/
/muse.jhu.edu/article/9279/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/1/700218/152039702320201030.pdf.

Anonymous:2002:ENd

[Ano02q] Anonymous. Editor’s note. Journal of Cold War Studies, 4(4):1– 2, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9301;   http:/
/muse.jhu.edu/article/9301/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/4/1/695660/15203970260209482.pdf.

Anonymous:2002:E

[Ano02r] Anonymous. Erratum. Journal of Cold War Studies, 4(3):??, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/4/3/
iv/700268/jcws.2002.4.3.ivb.pdf.

Anonymous:2003:Aa

[Ano03a] Anonymous. Abstracts. Journal of Cold War Studies, 5(1):??, Winter 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL https://direct.mit.edu/jcws/article-pdf/5/2/v/700470/
jcws.2003.5.2.v.pdf.

Anonymous:2003:Ab

[Ano03b] Anonymous. Abstracts. Journal of Cold War Studies, 5(3):??, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/5/3/
vi/695945/jcws.2003.5.3.vi.pdf.

Anonymous:2003:Ac

[Ano03c] Anonymous. Abstracts. Journal of Cold War Studies, 5(3):??, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/5/4/
iii/700388/jcws.2003.5.4.iii.pdf.

Anonymous:2003:Ca

[Ano03d] Anonymous. Contributors. Journal of Cold War Studies, 5(1):ii, Winter 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/38249;   http:
//muse.jhu.edu/article/38249/pdf.

Anonymous:2003:Cb

[Ano03e] Anonymous. Contributors. Journal of Cold War Studies, 5(1):??, Winter 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/5/
2/iv/700462/jcws.2003.5.2.iv.pdf.

Anonymous:2003:Cc

[Ano03f] Anonymous. Contributors. Journal of Cold War Studies, 5(2):iv, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/40757;   http:
//muse.jhu.edu/article/40757/pdf.

Anonymous:2003:Cd

[Ano03g] Anonymous. Contributors. Journal of Cold War Studies, 5(3):v, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/45186;   http:
//muse.jhu.edu/article/45186/pdf.

Anonymous:2003:Ce

[Ano03h] Anonymous. Contributors. Journal of Cold War Studies, 5(3):??, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/5/
3/v/695955/jcws.2003.5.3.v.pdf.

Anonymous:2003:Cf

[Ano03i] Anonymous. Contributors. Journal of Cold War Studies, 5(3):??, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/5/
4/ii/700390/jcws.2003.5.4.ii.pdf.

Anonymous:2003:Cg

[Ano03j] Anonymous. Contributors. Journal of Cold War Studies, 5(4):ii, Fall 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/47928;   http://muse.jhu.
edu/article/47928/pdf.

Anonymous:2003:E

[Ano03k] Anonymous. Erratum. Journal of Cold War Studies, 5(3):161, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/45191;   http://
muse.jhu.edu/article/45191/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/3/161/695949/jcws.2003.5.3.161b.pdf.

Anonymous:2003:PRN

[Ano03l] Anonymous. Perspectives on redrawing nations. Journal of Cold War Studies, 5(3):102–114, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/5/3/102/695816/152039703322286809.pdf.

Anonymous:2004:Ac

[Ano04a] Anonymous. Abstract. Journal of Cold War Studies, 6(3):??, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL https://direct.mit.edu/jcws/article-pdf/6/3/v/696137/
jcws.2004.6.3.v.pdf.

Anonymous:2004:Ad

[Ano04b] Anonymous. Abstract. Journal of Cold War Studies, 6(4):??, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL https://direct.mit.edu/jcws/article-pdf/6/4/v/696135/
jcws.2004.6.4.v.pdf.

Anonymous:2004:Aa

[Ano04c] Anonymous. Abstracts. Journal of Cold War Studies, 6(1):??, Winter 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/6/1/
iv/695913/jcws.2004.6.1.iv.pdf.

Anonymous:2004:Ab

[Ano04d] Anonymous. Abstracts. Journal of Cold War Studies, 6(2):??, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL https://direct.mit.edu/jcws/article-pdf/6/2/v/696013/
jcws.2004.6.2.v.pdf.

Anonymous:2004:Ae

[Ano04e] Anonymous. Abstracts. Journal of Cold War Studies, 6(4):??, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL https://direct.mit.edu/jcws/article-pdf/7/1/v/700450/
1520397053326130.pdf.

Anonymous:2004:Ca

[Ano04f] Anonymous. Contributors. Journal of Cold War Studies, 6(1):iii, Winter 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/51861;   http:
//muse.jhu.edu/article/51861/pdf.

Anonymous:2004:Cb

[Ano04g] Anonymous. Contributors. Journal of Cold War Studies, 6(1):??, Winter 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/6/
1/iii/695917/jcws.2004.6.1.iii.pdf.

Anonymous:2004:Cc

[Ano04h] Anonymous. Contributors. Journal of Cold War Studies, 6(2):iv, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/53663;   http:
//muse.jhu.edu/article/53663/pdf.

Anonymous:2004:Cd

[Ano04i] Anonymous. Contributors. Journal of Cold War Studies, 6(2):??, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/6/
2/iv/696005/jcws.2004.6.2.iv.pdf.

Anonymous:2004:Ce

[Ano04j] Anonymous. Contributors. Journal of Cold War Studies, 6(3):iv, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171154;   http://
muse.jhu.edu/article/171154/pdf.

Anonymous:2004:Cf

[Ano04k] Anonymous. Contributors. Journal of Cold War Studies, 6(3):??, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/6/
3/iv/696117/jcws.2004.6.3.iv.pdf.

Anonymous:2004:Cg

[Ano04l] Anonymous. Contributors. Journal of Cold War Studies, 6(4):iv, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/174706;   http://
muse.jhu.edu/article/174706/pdf.

Anonymous:2004:Ch

[Ano04m] Anonymous. Contributors. Journal of Cold War Studies, 6(4): ??, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/6/
4/iv/696129/jcws.2004.6.4.iv.pdf.

Anonymous:2004:Ci

[Ano04n] Anonymous. Contributors. Journal of Cold War Studies, 6(4): ??, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/7/
1/iv/700494/jcws.2005.7.1.iv.pdf.

Anonymous:2004:EN

[Ano04o] Anonymous. Editor’s note. Journal of Cold War Studies, 6(1):1–2, Winter 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/51863;   http://
muse.jhu.edu/article/51863/pdf;   https://direct.mit.edu/
jcws/article-pdf/6/1/1/695889/152039704772741560.pdf.

Anonymous:2004:Ea

[Ano04p] Anonymous. Errata. Journal of Cold War Studies, 6(1):103, Winter 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/51865;   http://
muse.jhu.edu/article/51865/pdf;   https://direct.mit.edu/
jcws/article-pdf/6/1/103/695909/jcws.2004.6.1.103.pdf.

Anonymous:2004:Eb

[Ano04q] Anonymous. Erratum. Journal of Cold War Studies, 6(2):iii, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/53667;   http://muse.jhu.
edu/article/53667/pdf.

Anonymous:2004:Ec

[Ano04r] Anonymous. Erratum. Journal of Cold War Studies, 6(3):181, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/171160;
http://muse.jhu.edu/article/171160/pdf;   https://direct.
mit.edu/jcws/article-pdf/6/3/181/696017/1520397041447391.
pdf.

Anonymous:2004:Ed

[Ano04s] Anonymous. Erratum. Journal of Cold War Studies, 6(4):179, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/174707;   http://
muse.jhu.edu/article/174707/pdf.

Anonymous:2005:Aa

[Ano05a] Anonymous. Abstracts. Journal of Cold War Studies, 7(1):??, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/7/2/
iii/696222/jcws.2005.7.2.iii.pdf.

Anonymous:2005:Ab

[Ano05b] Anonymous. Abstracts. Journal of Cold War Studies, 7(2):??, Spring 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/7/
3/vi/1102605/jcws.2005.7.3.vi.pdf.

Anonymous:2005:Ac

[Ano05c] Anonymous. Abstracts. Journal of Cold War Studies, 7(4):??, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/7/4/
vi/696229/1520397055012415.pdf.

Anonymous:2005:Ad

[Ano05d] Anonymous. Abstracts. Journal of Cold War Studies, 7(4):??, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL https://direct.mit.edu/jcws/article-pdf/8/1/v/696279/
152039706775212049.pdf.

Anonymous:2005:Ca

[Ano05e] Anonymous. Contributors. Journal of Cold War Studies, 7(1):iv, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/179549;   http://
muse.jhu.edu/article/179549/pdf.

Anonymous:2005:Cb

[Ano05f] Anonymous. Contributors. Journal of Cold War Studies, 7(1):??, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/7/
2/ii/696216/jcws.2005.7.2.ii.pdf.

Anonymous:2005:Cc

[Ano05g] Anonymous. Contributors. Journal of Cold War Studies, 7(2):ii, Spring 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/181898;   http://
muse.jhu.edu/article/181898/pdf.

Anonymous:2005:Cd

[Ano05h] Anonymous. Contributors. Journal of Cold War Studies, 7(2):??, Spring 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/7/3/v/
1102601/jcws.2005.7.3.v.pdf.

Anonymous:2005:Ce

[Ano05i] Anonymous. Contributors. Journal of Cold War Studies, 7(3):v, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/185097;   http://
muse.jhu.edu/article/185097/pdf.

Anonymous:2005:Cf

[Ano05j] Anonymous. Contributors. Journal of Cold War Studies, 7(4):v, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/188255;   http://
muse.jhu.edu/article/188255/pdf.

Anonymous:2005:Cg

[Ano05k] Anonymous. Contributors. Journal of Cold War Studies, 7(4):??, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL https://direct.mit.edu/jcws/article-pdf/7/4/v/696281/
1520397055012460.pdf.

Anonymous:2005:Ch

[Ano05l] Anonymous. Contributors. Journal of Cold War Studies, 7(4):??, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/8/1/
iv/696248/152039706775211996.pdf.

Anonymous:2006:Ca

[Ano06a] Anonymous. Contributors. Journal of Cold War Studies, 8(1):iv, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/192204;   http://
muse.jhu.edu/article/192204/pdf.

Anonymous:2006:Cb

[Ano06b] Anonymous. Contributors. Journal of Cold War Studies, 8(2):v, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/196273;   http://
muse.jhu.edu/article/196273/pdf.

Anonymous:2006:Cc

[Ano06c] Anonymous. Contributors. Journal of Cold War Studies, 8(3):ii, Summer 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/201233;   http://
muse.jhu.edu/article/201233/pdf.

Anonymous:2006:Cd

[Ano06d] Anonymous. Contributors. Journal of Cold War Studies, 8(4):iv, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/204618;   http://
muse.jhu.edu/article/204618/pdf.

Anonymous:2007:Ca

[Ano07a] Anonymous. Contributors. Journal of Cold War Studies, 9(1):vi, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/210074;   http://
muse.jhu.edu/article/210074/pdf.

Anonymous:2007:Cb

[Ano07b] Anonymous. Contributors. Journal of Cold War Studies, 9(2):iv, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/214732;   http://
muse.jhu.edu/article/214732/pdf.

Anonymous:2007:Cc

[Ano07c] Anonymous. Contributors. Journal of Cold War Studies, 9(3):v, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/219446;   http://
muse.jhu.edu/article/219446/pdf.

Anonymous:2007:Cd

[Ano07d] Anonymous. Contributors. Journal of Cold War Studies, 9(4):iv, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/223533;   http://
muse.jhu.edu/article/223533/pdf.

Anonymous:2008:Ca

[Ano08a] Anonymous. Contributors. Journal of Cold War Studies, 10(1):v, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/231572;   http://
muse.jhu.edu/article/231572/pdf.

Anonymous:2008:Cb

[Ano08b] Anonymous. Contributors. Journal of Cold War Studies, 10(2):iii, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/237147;   http://
muse.jhu.edu/article/237147/pdf.

Anonymous:2008:Cc

[Ano08c] Anonymous. Contributors. Journal of Cold War Studies, 10(3):vi, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/241494;   http://
muse.jhu.edu/article/241494/pdf.

Anonymous:2008:ENa

[Ano08d] Anonymous. Editor’s note. Journal of Cold War Studies, 10(1): 1–2, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/231573;
http://muse.jhu.edu/article/231573/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/1/1/700051/jcws.2008.10.1.1.
pdf.

Anonymous:2008:ENb

[Ano08e] Anonymous. Editor’s note. Journal of Cold War Studies, 10(2): 1–2, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/237149;
http://muse.jhu.edu/article/237149/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/2/1/696983/jcws.2008.10.2.1.
pdf.

Anonymous:2008:ENc

[Ano08f] Anonymous. Editor’s note. Journal of Cold War Studies, 10(4): 1–4, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/252654;
http://muse.jhu.edu/article/252654/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/4/1/697035/jcws.2008.10.4.1.
pdf.

Anonymous:2008:Ea

[Ano08g] Anonymous. Erratum. Journal of Cold War Studies, 10(1):1, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/231574;   http://
muse.jhu.edu/article/231574/pdf.

Anonymous:2008:Eb

[Ano08h] Anonymous. Erratum. Journal of Cold War Studies, 10(2):162, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/238052;   http://
muse.jhu.edu/article/238052/pdf.

Anonymous:2008:ECI

[Ano08i] Anonymous. Erratum: “In a Class by Itself”: Cold War Politics and Finland’s Position vis-à-vis the United Nations, 1945–1956. Journal of Cold War Studies, 10(4):??, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/252673;   http://muse.jhu.edu/article/252673/

pdf. See [Göt08].

Anonymous:2009:Ca

[Ano09a] Anonymous. Contributors. Journal of Cold War Studies, 11(1):iv, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/259181;   http://
muse.jhu.edu/article/259181/pdf.

Anonymous:2009:Cb

[Ano09b] Anonymous. Contributors. Journal of Cold War Studies, 11(2):iv, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/265889;   http://
muse.jhu.edu/article/265889/pdf.

Anonymous:2009:Cc

[Ano09c] Anonymous. Contributors. Journal of Cold War Studies, 11(3):ii, Summer 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/268387;   http://
muse.jhu.edu/article/268387/pdf.

Anonymous:2009:Cd

[Ano09d] Anonymous. Contributors. Journal of Cold War Studies, 11(4):iii, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/366243;   http://
muse.jhu.edu/article/366243/pdf.

Anonymous:2009:ENa

[Ano09e] Anonymous. Editor’s note. Journal of Cold War Studies, 11(1): 1–2, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/259182;
http://muse.jhu.edu/article/259182/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/1/1/1468508/jcws.2009.11.1.1.
pdf.

Anonymous:2009:ENb

[Ano09f] Anonymous. Editor’s note. Journal of Cold War Studies, 11(2): 1–2, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/265888;
http://muse.jhu.edu/article/265888/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/2/1/697153/jcws.2009.11.2.1.
pdf.

Anonymous:2009:ENc

[Ano09g] Anonymous. Editor’s note. Journal of Cold War Studies, 11(3): 1–5, Summer 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/268388;
http://muse.jhu.edu/article/268388/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/3/1/697246/jcws.2009.11.3.1.
pdf.

Anonymous:2009:ENd

[Ano09h] Anonymous. Editor’s note. Journal of Cold War Studies, 11(4): 1–2, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/366244;
http://muse.jhu.edu/article/366244/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/4/1/1627611/jcws.2009.11.4.1.
pdf.

Anonymous:2010:Ca

[Ano10a] Anonymous. Contributors. Journal of Cold War Studies, 12(1):iv, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/382483;   http://
muse.jhu.edu/article/382483/pdf.

Anonymous:2010:Cb

[Ano10b] Anonymous. Contributors. Journal of Cold War Studies, 12(2):iii, Spring 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/382196;   http://
muse.jhu.edu/article/382196/pdf.

Anonymous:2010:Cc

[Ano10c] Anonymous. Contributors. Journal of Cold War Studies, 12(3):v, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/392774;   http://
muse.jhu.edu/article/392774/pdf.

Anonymous:2010:Cd

[Ano10d] Anonymous. Contributors. Journal of Cold War Studies, 12(4):iv, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/403793;   http://
muse.jhu.edu/article/403793/pdf.

Anonymous:2010:ENa

[Ano10e] Anonymous. Editor’s note. Journal of Cold War Studies, 12(1): 1–2, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382484;

http://muse.jhu.edu/article/382484/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/1/1/1627419/jcws.2010.12.1.1.
pdf.

Anonymous:2010:ENb

[Ano10f] Anonymous. Editor’s note. Journal of Cold War Studies, 12(2): 1–2, Spring 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382197;
http://muse.jhu.edu/article/382197/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/2/1/697488/jcws.2010.12.2.1.
pdf.

Anonymous:2010:ENc

[Ano10g] Anonymous. Editor’s note. Journal of Cold War Studies, 12(3): 1–2, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/392775;
http://muse.jhu.edu/article/392775/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/3/1/697621/jcws_e_00029.pdf.

Anonymous:2010:ENd

[Ano10h] Anonymous. Editor’s note. Journal of Cold War Studies, 12(4): 1–2, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/403794;
http://muse.jhu.edu/article/403794/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/4/1/697592/jcws_r_00069.pdf.

Anonymous:2010:PSS

[Ano10i] Anonymous. Perspectives on the Sino–Soviet split. Journal of Cold War Studies, 12(1):120, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/382489;   http://muse.jhu.edu/article/382489/pdf.

Anonymous:2011:Ca

[Ano11a] Anonymous. Contributors. Journal of Cold War Studies, 13(1):iv, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/424238;   http://
muse.jhu.edu/article/424238/pdf.

Anonymous:2011:Cb

[Ano11b] Anonymous. Contributors. Journal of Cold War Studies, 13(2):ii, Spring 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/435663;   http://
muse.jhu.edu/article/435663/pdf.

Anonymous:2011:Cc

[Ano11c] Anonymous. Contributors. Journal of Cold War Studies, 13(3):iv, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/447460;   http://
muse.jhu.edu/article/447460/pdf.

Anonymous:2011:Cd

[Ano11d] Anonymous. Contributors. Journal of Cold War Studies, 13(4):v, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457190;   http://
muse.jhu.edu/article/457190/pdf.

Anonymous:2011:ENa

[Ano11e] Anonymous. Editor’s note. Journal of Cold War Studies, 13(1): 1–3, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/424239;
http://muse.jhu.edu/article/424239/pdf;   https://direct.
mit.edu/jcws/article-pdf/13/1/1/697746/jcws_e_00099.pdf.

Anonymous:2011:ENb

[Ano11f] Anonymous. Editor’s note. Journal of Cold War Studies, 13(2): 1–3, Spring 2011. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/435664;
http://muse.jhu.edu/article/435664/pdf;   https://direct.
mit.edu/jcws/article-pdf/13/2/1/700674/jcws_e_00139.pdf.

Anonymous:2011:ENc

[Ano11g] Anonymous. Editor’s note. Journal of Cold War Studies, 13(3): 1–3, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/447466;
http://muse.jhu.edu/article/447466/pdf;   https://direct.
mit.edu/jcws/article-pdf/13/3/1/700698/jcws_e_00165.pdf.

Anonymous:2011:ENd

[Ano11h] Anonymous. Editor’s note. Journal of Cold War Studies, 13(4): 1–3, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/457191;
http://muse.jhu.edu/article/457191/pdf;   https://direct.
mit.edu/jcws/article-pdf/13/4/1/697794/jcws_e_00175.pdf.

Anonymous:2012:Ca

[Ano12a] Anonymous. Contributors. Journal of Cold War Studies, 14(1):iv, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/467674;   http://
muse.jhu.edu/article/467674/pdf.

Anonymous:2012:Cb

[Ano12b] Anonymous. Contributors. Journal of Cold War Studies, 14(2):iii, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/478772;   http://
muse.jhu.edu/article/478772/pdf.

Anonymous:2012:Cc

[Ano12c] Anonymous. Contributors. Journal of Cold War Studies, 14(3):iv, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/490198;   http://
muse.jhu.edu/article/490198/pdf.

Anonymous:2012:Cd

[Ano12d] Anonymous. Contributors. Journal of Cold War Studies, 14(4):iv, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/500326;   http://
muse.jhu.edu/article/500326/pdf.

Anonymous:2012:ENa

[Ano12e] Anonymous. Editor’s note. Journal of Cold War Studies, 14(1): 1–2, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/467675;
http://muse.jhu.edu/article/467675/pdf;   https://direct.
mit.edu/jcws/article-pdf/14/1/1/697847/jcws_e_00215.pdf.

Anonymous:2012:ENb

[Ano12f] Anonymous. Editor’s note. Journal of Cold War Studies, 14(2): 1–2, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/478773;
http://muse.jhu.edu/article/478773/pdf;   https://direct.
mit.edu/jcws/article-pdf/14/2/1/697983/jcws_e_00218.pdf.

Anonymous:2012:ENc

[Ano12g] Anonymous. Editor’s note. Journal of Cold War Studies, 14(3): 1–2, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/490172;
http://muse.jhu.edu/article/490172/pdf;   https://direct.
mit.edu/jcws/article-pdf/14/3/1/698043/jcws_e_00245.pdf.

Anonymous:2012:ENd

[Ano12h] Anonymous. Editor’s note. Journal of Cold War Studies, 14(4): 1–3, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/500327;
http://muse.jhu.edu/article/500327/pdf;   https://direct.
mit.edu/jcws/article-pdf/14/4/1/698122/jcws_e_00274.pdf.

Anonymous:2013:Ca

[Ano13a] Anonymous. Contributors. Journal of Cold War Studies, 15(1):iv, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/505774;   http://
muse.jhu.edu/article/505774/pdf.

Anonymous:2013:Cb

[Ano13b] Anonymous. Contributors. Journal of Cold War Studies, 15(2):iv, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/509723;   http://
muse.jhu.edu/article/509723/pdf.

Anonymous:2013:Cc

[Ano13c] Anonymous. Contributors. Journal of Cold War Studies, 15(3):iv, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/522705;   http://
muse.jhu.edu/article/522705/pdf.

Anonymous:2013:Cd

[Ano13d] Anonymous. Contributors. Journal of Cold War Studies, 15(4):iii–iv, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/534410;   http://
muse.jhu.edu/article/534410/pdf.

Anonymous:2013:ENa

[Ano13e] Anonymous. Editor’s note. Journal of Cold War Studies, 15(1):1–2, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/15/1/
1/699958/jcws_e_00305.pdf.

Anonymous:2013:ENb

[Ano13f] Anonymous. Editor’s note. Journal of Cold War Studies, 15(2): 1–3, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/509700;

http://muse.jhu.edu/article/509700/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/2/1/698135/jcws_e_00334.pdf.

Anonymous:2013:ENc

[Ano13g] Anonymous. Editor’s note. Journal of Cold War Studies, 15(3): 1–3, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/522706;
http://muse.jhu.edu/article/522706/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/3/1/698288/jcws_e_00367.pdf.

Anonymous:2013:ENd

[Ano13h] Anonymous. Editor’s note. Journal of Cold War Studies, 15(4): 1–3, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/534411;
http://muse.jhu.edu/article/534411/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/4/1/698365/jcws_e_00392.pdf.

Anonymous:2014:Ca

[Ano14a] Anonymous. Contributors. Journal of Cold War Studies, 16(1):iv, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543720;   http://
muse.jhu.edu/article/543720/pdf.

Anonymous:2014:Cb

[Ano14b] Anonymous. Contributors. Journal of Cold War Studies, 16(2):iii, Spring 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/547320;   http://
muse.jhu.edu/article/547320/pdf.

Anonymous:2014:Cc

[Ano14c] Anonymous. Contributors. Journal of Cold War Studies, 16(3):iii, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/556072;   http://
muse.jhu.edu/article/556072/pdf.

Anonymous:2014:Cd

[Ano14d] Anonymous. Contributors. Journal of Cold War Studies, 16(4):v, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/566649;   http://
muse.jhu.edu/article/566649/pdf.

Anonymous:2014:ENa

[Ano14e] Anonymous. Editor’s note. Journal of Cold War Studies, 16(1): 1–3, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/543721;
http://muse.jhu.edu/article/543721/pdf;   https://direct.
mit.edu/jcws/article-pdf/16/1/1/698892/jcws_e_00428.pdf.

Anonymous:2014:ENb

[Ano14f] Anonymous. Editor’s note. Journal of Cold War Studies, 16(2): 1–2, Spring 2014. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/547321;
http://muse.jhu.edu/article/547321/pdf;   https://direct.
mit.edu/jcws/article-pdf/16/2/1/698423/jcws_e_00446.pdf.

Anonymous:2014:ENc

[Ano14g] Anonymous. Editor’s note. Journal of Cold War Studies, 16(3): 1–4, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/556073;
http://muse.jhu.edu/article/556073/pdf;   https://direct.
mit.edu/jcws/article-pdf/16/3/1/698500/jcws_e_00468.pdf.

Anonymous:2014:ENd

[Ano14h] Anonymous. Editor’s note. Journal of Cold War Studies, 16(4): 1–4, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/566650;
http://muse.jhu.edu/article/566650/pdf;   https://direct.
mit.edu/jcws/article-pdf/16/4/1/698552/jcws_e_00513.pdf.

Anonymous:2014:E

[Ano14i] Anonymous. Errata. Journal of Cold War Studies, 16(1):3, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543722;   http://
muse.jhu.edu/article/543722/pdf.

Anonymous:2015:Ca

[Ano15a] Anonymous. Contributors. Journal of Cold War Studies, 17(1):iii, Winter 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/581227;   http://
muse.jhu.edu/article/581227/pdf.

Anonymous:2015:Cb

[Ano15b] Anonymous. Contributors. Journal of Cold War Studies, 17(2):iii, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/582431;   http://
muse.jhu.edu/article/582431/pdf.

Anonymous:2015:Cc

[Ano15c] Anonymous. Contributors. Journal of Cold War Studies, 17(3):vii, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595114;   http://
muse.jhu.edu/article/595114/pdf.

Anonymous:2015:Cd

[Ano15d] Anonymous. Contributors. Journal of Cold War Studies, 17(4):v, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/606017;   http://
muse.jhu.edu/article/606017/pdf.

Anonymous:2015:ENa

[Ano15e] Anonymous. Editor’s note. Journal of Cold War Studies, 17(1): 1–3, Winter 2015. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/581228;
http://muse.jhu.edu/article/581228/pdf;   https://direct.
mit.edu/jcws/article-pdf/17/1/1/700784/jcws_e_00528.pdf.

Anonymous:2015:ENb

[Ano15f] Anonymous. Editor’s note. Journal of Cold War Studies, 17(2): 1–2, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/582432;
http://muse.jhu.edu/article/582432/pdf;   https://direct.
mit.edu/jcws/article-pdf/17/2/1/698666/jcws_e_00545.pdf.

Anonymous:2015:ENc

[Ano15g] Anonymous. Editor’s note. Journal of Cold War Studies, 17(3): 1–3, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/595076;
http://muse.jhu.edu/article/595076/pdf;   https://direct.
mit.edu/jcws/article-pdf/17/3/1/1102401/jcws_e_00560.pdf.

Anonymous:2015:ENd

[Ano15h] Anonymous. Editor’s note. Journal of Cold War Studies, 17(4): 1–3, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/606018;
http://muse.jhu.edu/article/606018/pdf;   https://direct.
mit.edu/jcws/article-pdf/17/4/1/698753/jcws_e_00594.pdf.

Anonymous:2016:Ca

[Ano16a] Anonymous. Contributors. Journal of Cold War Studies, 18(1):iii, Winter 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/612002;   http://
muse.jhu.edu/article/612002/pdf.

Anonymous:2016:Cb

[Ano16b] Anonymous. Contributors. Journal of Cold War Studies, 18(2):iii, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/620543;   http://
muse.jhu.edu/article/620543/pdf.

Anonymous:2016:Cc

[Ano16c] Anonymous. Contributors. Journal of Cold War Studies, 18(3):v, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/632275;   http://
muse.jhu.edu/article/632275/pdf.

Anonymous:2016:Cd

[Ano16d] Anonymous. Contributors. Journal of Cold War Studies, 18(4):v–vi, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/645903;   http://
muse.jhu.edu/article/645903/pdf.

Anonymous:2016:ENa

[Ano16e] Anonymous. Editor’s note. Journal of Cold War Studies, 18(1): 1–4, Winter 2016. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/612003;
http://muse.jhu.edu/article/612003/pdf;   https://direct.
mit.edu/jcws/article-pdf/18/1/1/698781/jcws_e_00616.pdf.

Anonymous:2016:ENb

[Ano16f] Anonymous. Editor’s note. Journal of Cold War Studies, 18(2): 1–3, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/620544;
http://muse.jhu.edu/article/620544/pdf;   https://direct.
mit.edu/jcws/article-pdf/18/2/1/698850/jcws_e_00635.pdf.

Anonymous:2016:ENc

[Ano16g] Anonymous. Editor’s note. Journal of Cold War Studies, 18(3): 1–2, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/632276;

http://muse.jhu.edu/article/632276/pdf;   https://direct.
mit.edu/jcws/article-pdf/18/3/1/698996/jcws_e_00650.pdf.

Anonymous:2016:ENd

[Ano16h] Anonymous. Editor’s note. Journal of Cold War Studies, 18(4): 1–3, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/645904;
http://muse.jhu.edu/article/645904/pdf;   https://direct.
mit.edu/jcws/article-pdf/18/4/1/700031/jcws_e_00676.pdf.

Anonymous:2017:C

[Ano17a] Anonymous. Contributors. Journal of Cold War Studies, 19(3):iv, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/668155;   http://
muse.jhu.edu/article/668155/pdf.

Anonymous:2017:ENa

[Ano17b] Anonymous. Editor’s note. Journal of Cold War Studies, 19(1):1–3, Winter 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://www.mitpressjournals.org/doi/full/10.
1162/JCWS_e_00715;   https://direct.mit.edu/jcws/article-
pdf/19/1/1/699106/jcws_e_00715.pdf.

Anonymous:2017:ENb

[Ano17c] Anonymous. Editor’s note. Journal of Cold War Studies, 19(2):1–2, Spring 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://www.mitpressjournals.org/doi/full/10.
1162/JCWS_e_00738;   https://direct.mit.edu/jcws/article-
pdf/19/2/1/699214/jcws_e_00738.pdf.

Anonymous:2017:ENc

[Ano17d] Anonymous. Editor’s note. Journal of Cold War Studies, 19(3): 1–3, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/668156;
http://muse.jhu.edu/article/668156/pdf;   https://direct.
mit.edu/jcws/article-pdf/19/3/1/699226/jcws_e_00752.pdf.

Anonymous:2017:ENd

[Ano17e] Anonymous. Editor’s note. Journal of Cold War Studies, 19(4): 1–3, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/683663;
http://muse.jhu.edu/article/683663/pdf;   https://direct.
mit.edu/jcws/article-pdf/19/4/1/699360/jcws_e_00762.pdf.

Anonymous:2017:E

[Ano17f] Anonymous. Erratum. Journal of Cold War Studies, 19(2):222, Spring 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://www.mitpressjournals.org/doi/full/10.
1162/JCWS_e_00751;   https://direct.mit.edu/jcws/article-
pdf/19/2/222/699216/jcws_e_00751.pdf.

Anonymous:2018:Ca

[Ano18a] Anonymous. Contributors. Journal of Cold War Studies, 20(1):ii, Winter 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/691014;   http://
muse.jhu.edu/article/691014/pdf.

Anonymous:2018:Cb

[Ano18b] Anonymous. Contributors. Journal of Cold War Studies, 20(2):??, Spring 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/697057;   http://
muse.jhu.edu/article/697058.

Anonymous:2018:Cc

[Ano18c] Anonymous. Contributors. Journal of Cold War Studies, 20(3):ii, Summer 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/704576.

Anonymous:2018:Cd

[Ano18d] Anonymous. Contributors. Journal of Cold War Studies, 20(4):v, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/718196.

Anonymous:2018:ENa

[Ano18e] Anonymous. Editor’s note. Journal of Cold War Studies, 20(1): 1–3, Winter 2018. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/691015;
http://muse.jhu.edu/article/691015/pdf;   https://direct.
mit.edu/jcws/article-pdf/20/1/1/699333/jcws_e_00798.pdf.

Anonymous:2018:ENb

[Ano18f] Anonymous. Editor’s note. Journal of Cold War Studies, 20(2):1–2, Spring 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/20/2/
1/699421/jcws_e_00812.pdf.

Anonymous:2018:ENc

[Ano18g] Anonymous. Editor’s note. Journal of Cold War Studies, 20(2):??, Spring 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/697058.

Anonymous:2018:ENd

[Ano18h] Anonymous. Editor’s note. Journal of Cold War Studies, 20 (3):1–2, Summer 2018. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
704569;   https://direct.mit.edu/jcws/article-pdf/20/3/1/
699629/jcws_e_00821.pdf.

Anonymous:2018:ENe

[Ano18i] Anonymous. Editor’s note. Journal of Cold War Studies, 20 (4):1–3, Fall 2018. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
718197;   https://direct.mit.edu/jcws/article-pdf/20/4/1/
1969614/jcws_e_00853.pdf.

Anonymous:2018:E

[Ano18j] Anonymous. Erratum. Journal of Cold War Studies, 20(2):159, Spring 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/697066;   https:
//direct.mit.edu/jcws/article-pdf/20/2/159/699425/jcws_e_

00820.pdf. See [Hay17].

Anonymous:2019:Ca

[Ano19a] Anonymous. Contributors. Journal of Cold War Studies, 21(1):iv, Winter 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/723549.

Anonymous:2019:Cb

[Ano19b] Anonymous. Contributors. Journal of Cold War Studies, 21(2):iii, Spring 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/726146.

Anonymous:2019:Cc

[Ano19c] Anonymous. Contributors. Journal of Cold War Studies, 21(3):iii, Summer 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/731279.

Anonymous:2019:Cd

[Ano19d] Anonymous. Contributors. Journal of Cold War Studies, 21(4):ii, Fall 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL https://muse.jhu.edu/article/736851.

Anonymous:2019:ENa

[Ano19e] Anonymous. Editor’s note. Journal of Cold War Studies, 21 (1):1–2, Winter 2019. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
723550;   https://direct.mit.edu/jcws/article-pdf/21/1/1/
699704/jcws_e_00870.pdf.

Anonymous:2019:ENb

[Ano19f] Anonymous. Editor’s note. Journal of Cold War Studies, 21 (2):1–2, Spring 2019. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
726147;   https://direct.mit.edu/jcws/article-pdf/21/2/1/
699547/jcws_e_00874.pdf.

Anonymous:2019:ENc

[Ano19g] Anonymous. Editor’s note. Journal of Cold War Studies, 21 (3):1–3, Summer 2019. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
731280;   https://direct.mit.edu/jcws/article-pdf/21/3/1/
699662/jcws_e_00891.pdf.

Anonymous:2019:ENd

[Ano19h] Anonymous. Editor’s note. Journal of Cold War Studies, 21(4):1–3, Fall 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/21/4/
1/699752/jcws_e_00904.pdf;   https://muse.jhu.edu/article/
736850.

Anonymous:2020:Ca

[Ano20a] Anonymous. Contributors. Journal of Cold War Studies, 22(1):iv, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://muse.jhu.edu/article/749660.

Anonymous:2020:Cb

[Ano20b] Anonymous. Contributors. Journal of Cold War Studies, 22(2):iii, Spring 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://muse.jhu.edu/article/756821.

Anonymous:2020:Cc

[Ano20c] Anonymous. Contributors. Journal of Cold War Studies, 22(3):iii, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://muse.jhu.edu/article/762937.

Anonymous:2020:Cd

[Ano20d] Anonymous. Contributors. Journal of Cold War Studies, 22(4):iii, Fall 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL https://muse.jhu.edu/article/776972.

Anonymous:2020:ENa

[Ano20e] Anonymous. Editor’s note. Journal of Cold War Studies, 22(1):1–3, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/22/1/1/699810/jcws_e_00923.pdf;   https:
//muse.jhu.edu/article/749651.

Anonymous:2020:ENb

[Ano20f] Anonymous. Editor’s note. Journal of Cold War Studies, 22(2):1– 2, Spring 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/2/1/1860552/jcws_e_00937.pdf;   https://muse.jhu.edu/
article/756822.

Anonymous:2020:ENc

[Ano20g] Anonymous. Editor’s note. Journal of Cold War Studies, 22(3):1–3, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/3/1/1860600/jcws_e_00947.pdf;   https://muse.jhu.edu/
article/762889.

Anonymous:2020:ENd

[Ano20h] Anonymous. Editor’s note. Journal of Cold War Studies, 22(4): 1–3, Fall 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/4/1/1860618/jcws_e_00963.pdf;   https://muse.jhu.edu/
article/776973.

Anonymous:2021:Ca

[Ano21a] Anonymous. Contributors. Journal of Cold War Studies, 23(1):iv, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://muse.jhu.edu/article/788377.

Anonymous:2021:Cb

[Ano21b] Anonymous. Contributors. Journal of Cold War Studies, 23(2):iii, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://muse.jhu.edu/article/794730.

Anonymous:2021:Cc

[Ano21c] Anonymous. Contributors. Journal of Cold War Studies, 23(3):iii, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://muse.jhu.edu/article/801903.

Anonymous:2021:Cd

[Ano21d] Anonymous. Contributors. Journal of Cold War Studies, 23(4):iii–iv, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://muse.jhu.edu/article/835874.

Anonymous:2021:ENa

[Ano21e] Anonymous. Editor’s note. Journal of Cold War Studies, 23(1):1– 3, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
23/1/1/1911802/jcws_e_00979.pdf;   https://muse.jhu.edu/
article/788378.

Anonymous:2021:ENb

[Ano21f] Anonymous. Editor’s note. Journal of Cold War Studies, 23(2):1– 3, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
23/2/1/1919927/jcws_e_01006.pdf;   https://muse.jhu.edu/
article/794724.

Anonymous:2021:ENc

[Ano21g] Anonymous. Editor’s note. Journal of Cold War Studies, 23(3):1–3, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
23/3/1/1955761/jcws_e_01025.pdf;   https://muse.jhu.edu/
article/801904.

Anonymous:2021:ENd

[Ano21h] Anonymous. Editor’s note. Journal of Cold War Studies, 23(4): 1–2, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
23/4/1/1970085/jcws_e_01037.pdf;   https://muse.jhu.edu/
article/835875.

Anonymous:2022:C

[Ano22a] Anonymous. Contributors. Journal of Cold War Studies, 24(1):iii, Winter 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://muse.jhu.edu/article/843074.

Anonymous:2022:ENa

[Ano22b] Anonymous. Editor’s note. Journal of Cold War Studies, 24(1):1– 3, Winter 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
24/1/1/1980867/jcws_e_01054.pdf;   https://muse.jhu.edu/
article/843075.

Anonymous:2022:ENb

[Ano22c] Anonymous. Editor’s note. Journal of Cold War Studies, 24(2):1–3, Spring 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/2/
1/2013016/jcws_e_01073.pdf.

Anonymous:2022:ENc

[Ano22d] Anonymous. Editor’s note. Journal of Cold War Studies, 24(3):1–3, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/3/
1/2042007/jcws_e_01086.pdf.

Anonymous:2022:ENd

[Ano22e] Anonymous. Editor’s note. Journal of Cold War Studies, 24(4):1–3, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/4/
1/2062817/jcws_e_01103.pdf.

Anonymous:2023:ENa

[Ano23a] Anonymous. Editor’s note. Journal of Cold War Studies, 25(1):1–3, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/1/
1/2073483/jcws_e_01118.pdf.

Anonymous:2023:ENb

[Ano23b] Anonymous. Editor’s note. Journal of Cold War Studies, 25(2):1–4, Spring 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/2/
1/2140968/jcws_e_01138.pdf.

Anonymous:2023:ENc

[Ano23c] Anonymous. Editor’s note. Journal of Cold War Studies, 25(3):1–5, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/3/
1/2158680/jcws_e_01155.pdf.

Anonymous:2023:ENd

[Ano23d] Anonymous. Editor’s note. Journal of Cold War Studies, 25(4):1–3, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/4/
1/2205010/jcws_e_01167.pdf.

Ansari:2017:JWM

[Ans17] Emily Abrams Ansari. James Wierzbicki, Music in the Age of Anxiety: American Music in the Fifties. Journal of Cold War Studies, 19(4):268–270, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683689;   http://muse.jhu.edu/article/683689/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/268/699455/
jcws_r_00789.pdf.

Apter:2008:C

[Apt08] David E. Apter. Commentary. Journal of Cold War Studies, 10(2): 119–125, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/237153;
http://muse.jhu.edu/article/237153/pdf.

Argentieri:2003:TRR

[Arg03] Federigo Argentieri. A Testament of Revolution (review). Journal of Cold War Studies, 5(3):143–145, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45180;   http://muse.jhu.edu/
article/45180/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/143/695919/jcws.2003.5.3.143.pdf.

Argentieri:2004:ASS

[Arg04] Federigo Argentieri. Academia and State Socialism–Essays on the Political History of Academic Life in Post-1945 Hungary and Eastern Europe (review). Journal of Cold War Studies, 6 (3):160–162, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/171147;   http://muse.jhu.edu/article/171147/pdf;

 https://direct.mit.edu/jcws/article-pdf/6/3/160/696101/
jcws.2004.6.3.160.pdf.

Aunesluoma:2016:NIF

[ARN16] Juhana Aunesluoma and Johanna Rainio-Niemi. Neutrality as identity?: Finland’s quest for security in the Cold War. Journal of Cold War Studies, 18(4):51–78, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645908;   http://muse.jhu.edu/article/645908/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/51/700049/
jcws_a_00680.pdf.

Aronson:2006:BMA

[Aro06] Arnold Aronson. Bruce McConachie, American Theatre in the Culture of the Cold War: Producing and Contesting Containment, 1947–1962. Journal of Cold War Studies, 8(2):152–154, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/196265;   http:
//muse.jhu.edu/article/196265/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/2/152/700534/jcws.2006.8.2.152.pdf.

Autio-Sarasmo:2022:NTC

[AS22] Sari Autio-Sarasmo. Neutrality in Twentieth-Century Europe: Intersections of science, culture, and politics after the First World War. Journal of Cold War Studies, 24(3):248–250, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/3/
248/2041993/jcws_r_01085.pdf.

Aase:2016:SSG

[Åse16] Cecilia Åse. Ship of shame: Gender and nation in narratives of the 1981 Soviet submarine crisis in Sweden. Journal of Cold War Studies, 18(1):112–132, Winter 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/612008;   http://muse.jhu.edu/article/612008/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/1/112/698765/
jcws_a_00621.pdf.

Ashton:2010:SBB

[Ash10] Nigel Ashton. Stephen Blackwell, British Military Intervention and the Struggle for Jordan: King Hussein, Nasser and the Middle East Crisis, 1955–1958. Journal of Cold War Studies, 12(1):169– 171, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382491;
http://muse.jhu.edu/article/382491/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/1/169/1627429/jcws.2010.12.1.
169.pdf.

Asselin:2007:CPH

[Ass07] Pierre Asselin. Choosing peace: Hanoi and the Geneva agreement on Vietnam, 1954–1955. Journal of Cold War Studies, 9(2):95– 126, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/214729;
http://muse.jhu.edu/article/214729/pdf;   https://direct.
mit.edu/jcws/article-pdf/9/2/95/696491/jcws.2007.9.2.95.
pdf.

Asselin:2010:DAJ

[Ass10] Pierre Asselin. David L. Anderson and John Ernst, eds., The War That Never Ends: New Perspectives on the Vietnam War. Journal of Cold War Studies, 12(4):195–197, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/403808;   http://muse.jhu.edu/article/403808/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/4/195/697562/
jcws_r_00060.pdf.

Asselin:2011:RTH

[Ass11] Pierre Asselin. Revisionism triumphant: Hanoi’s diplomatic strategy in the Nixon era. Journal of Cold War Studies, 13(4):101–137, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457195;   http:
//muse.jhu.edu/article/457195/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/4/101/697796/jcws_a_00170.pdf.

Asselin:2014:JMC

[Ass14] Pierre Asselin. Jessica M. Chapman, Cauldron of Resistance: Ngo Dinh Diem, the United States, and 1950s Southern Vietnam. Ithaca, NY: Cornell University Press, 2013. xiii + 276 pp. $39.95. Journal of Cold War Studies, 16(2):129–130, Spring 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/547328;   http://muse.jhu.edu/article/547328/
pdf;   https://direct.mit.edu/jcws/article-pdf/16/2/129/
698435/jcws_r_00455.pdf.

Asselin:2015:GVR

[Ass15] Pierre Asselin. De Gaulle et le Vietnam, 1945–1969: La réconciliation by Pierre Journoud (review). Journal of Cold War Studies, 17(2):154–156, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/582444;   http://muse.jhu.edu/article/582444/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/2/154/698702/
jcws_r_00542.pdf.

Asselin:2023:ICP

[Ass23] Pierre Asselin. The Indochinese Communist Party’s unfinished revolution of 1945 and the origins of Vietnam’s 30-year civil war. Journal of Cold War Studies, 25(1):4–45, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/1/4/2073507/jcws_a_01120.pdf.

Attebery:2023:BRO

[Att23] Clint Attebery. Book review: Operation Pedro Pan: The Migration of Unaccompanied Children from Castro’s Cuba by John A. Gronbeck-Tedesco. Journal of Cold War Studies, 25(4):265–267, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/4/
265/2204999/jcws_r_01184.pdf.

Avant:2004:UWW

[Ava04] Deborah D. Avant. Uncovering Ways of War: U.S. Intelligence and Foreign Military Innovation, 1918–1941 (review). Journal of Cold War Studies, 6(4):159–161, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/174704;   http://muse.jhu.edu/article/174704/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/4/159/696111/
jcws.2004.6.4.159.pdf.

Bacevich:2002:RAD

[Bac02] Andrew J. Bacevich. Reading Athena’s Dance Card: Men against Fire in Vietnam (review). Journal of Cold War Studies, 4(2):135– 137, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9243;   http:/
/muse.jhu.edu/article/9243/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/2/135/695589/jcws.2002.4.2.135.pdf.

Bacevich:2009:AJH

[Bac09] Andrew J. Bacevich. Andrew J. Huebner, The Warrior Image: Soldiers in American Culture from the Second World War to the Vietnam Era. Journal of Cold War Studies, 11(1):150– 151, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/259188;
http://muse.jhu.edu/article/259188/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/1/150/1468494/jcws.2009.11.1.
150.pdf.

Baev:2002:RNS

[Bae02] Pavel K. Baev. The Russian Nuclear Shield from Stalin to Yeltsin (review). Journal of Cold War Studies, 4(3):151–152, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9267;   http:/
/muse.jhu.edu/article/9267/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/151/700248/jcws.2002.4.3.151.pdf.

Baev:2016:EBW

[Bae16] Jordan Baev. The establishment of Bulgarian–West German diplomatic relations within the coordinating framework of the Warsaw Pact. Journal of Cold War Studies, 18(3):158–180, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/632282;   http:
//muse.jhu.edu/article/632282/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/3/158/699047/jcws_a_00656.pdf.

Bailey:2018:JRD

[Bai18] Greg Bailey. Joel R. Davidson, Armchair Warriors: Private Citizens, Popular Press and the Rise of American Power. Journal of Cold War Studies, 20(4):250–251, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718216;   https://direct.mit.
edu/jcws/article-pdf/20/4/250/1969652/jcws_r_00837.pdf.

Baker:2009:JPH

[Bak09] Anni Baker. John P. Hawkins, Army of Hope, Army of Alienation: Culture and Contradiction in the American Army Communities of Cold War Germany. Journal of Cold War Studies, 11(2):149–150, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/265869;   http://muse.jhu.edu/article/265869/pdf;

 https://direct.mit.edu/jcws/article-pdf/11/2/149/697193/
jcws.2009.11.2.149.pdf.

Balogh:2005:CSM

[Bal05] Anna Balogh. Carrying a Secret in My Heart: Children of the Victims of the Reprisals after the Hungarian Revolution in 1956 — An Oral History (review). Journal of Cold War Studies, 7(3):178–179, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185095;   http://muse.jhu.edu/article/185095/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/178/1102579/
jcws.2005.7.3.178.pdf.

Baldwin:2008:AJB

[Bal08] David A. Baldwin. Andrew J. Bacevich, ed., The Long War: a New History of U.S. National Security Policy since World War II. Journal of Cold War Studies, 10(4):149–151, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/252661;   http://muse.jhu.edu/article/252661/
pdf;   https://direct.mit.edu/jcws/article-pdf/10/4/149/
697045/jcws.2008.10.4.149.pdf.

Balogh:2010:LBH

[Bal10] Anna Balogh. László Borhi, Hungary in the Cold War, 1945–1956: Between the United States and the Soviet Union. Journal of Cold War Studies, 12(1):189–191, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/382492;   http://muse.jhu.edu/article/382492/
pdf;   https://direct.mit.edu/jcws/article-pdf/12/1/189/
1627449/jcws.2010.12.1.189.pdf.

Bale:2015:HWA

[Bal15] Jeffrey M. Bale. Hugh Wilford, America’s Great Game: The CIA’s Secret Arabists and the Shaping of the Modern Middle East. New York: Basic Books, 2013. 299 pp. $29.99. Journal of Cold War Studies, 17(4):179–181, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606028;   http://muse.jhu.edu/article/606028/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/179/698751/
jcws_r_00583.pdf.

Balzer:2017:LGL

[Bal17] Harley Balzer. Loren Graham, Lysenko’s Ghost: Epigenetics and Russia. Cambridge, MA: Harvard University Press, 2016. 209 pp. $24.95. Journal of Cold War Studies, 19(2):220–222, Spring 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http:

//www.mitpressjournals.org/doi/full/10.1162/JCWS_r_00748;
 https://direct.mit.edu/jcws/article-pdf/19/2/220/699224/
jcws_r_00748.pdf.

Bange:2006:NAK

[Ban06] Oliver Bange. Nigel Ashton, Kennedy, Macmillan and the Cold War: The Irony of Interdependence. Journal of Cold War Studies, 8(2):161–163, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196266;   http://muse.jhu.edu/article/196266/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/161/700550/
jcws.2006.8.2.161.pdf.

Bange:2007:BTN

[Ban07] Oliver Bange. Bruno Thoß, NATO-Strategie und nationale Verteidigungsplanung: Planung und Aufbau der Bundeswehr unter den Bedingungen einer massiven atomaren Vergeltungsstrategie 1952 bis 1960. Journal of Cold War Studies, 9(3):157–160, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/219439;   http:
//muse.jhu.edu/article/219439/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/3/157/696634/jcws.2007.9.3.157.pdf.

Bange:2014:KDE

[Ban14] Oliver Bange. Kriegsschauplatz Deutschland: Erfahrungen und Erkenntnisse eines NVA-Offiziers by Siegfried Lautsch (review). Journal of Cold War Studies, 16(4):268–271, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566679;   http://muse.jhu.edu/article/566679/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/268/698649/
jcws_r_00498.pdf.

Bange:2016:SSE

[Ban16] Oliver Bange. Onto the slippery slope: East Germany and East–West Détente under Ulbricht and Honecker, 1965–1975. Journal of Cold War Studies, 18(3):60–94, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632279;   http://muse.jhu.edu/article/632279/pdf;

 https://direct.mit.edu/jcws/article-pdf/18/3/60/699039/
jcws_a_00653.pdf.

Barnes:2003:RAN

[Bar03a] Steven Anthony Barnes. Return from the Archipelago: Narratives of Gulag Survivors (review). Journal of Cold War Studies, 5(3):159–160, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/45181;   http://
muse.jhu.edu/article/45181/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/3/159/695929/jcws.2003.5.3.159.pdf.

Barnhart:2003:MPH

[Bar03b] Michael A. Barnhart. From Munich to Pearl Harbor: Roosevelt’s America and the Origins of the Second World War (review). Journal of Cold War Studies, 5(2):84–86, Spring 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40754;   http://muse.jhu.edu/
article/40754/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/84/700426/jcws.2003.5.2.84.pdf.

Barnett:2004:UCW

[Bar04] Thomas P. M. Barnett. Understanding the Cold War: a Historian’s Personal Reflections (review). Journal of Cold War Studies, 6(3):170–172, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/171148;   http://muse.jhu.edu/article/171148/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/3/170/696113/
jcws.2004.6.3.170.pdf.

Barrett:2007:RDJ

[Bar07] David M. Barrett. Robert David Johnson, Congress and the Cold War. New York: Cambridge University Press, 2005. 346 pp. $25.99. Journal of Cold War Studies, 9(4):130–132, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/223530;   http:
//muse.jhu.edu/article/223530/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/4/130/696733/jcws.2007.9.4.130.pdf.

Barbier:2008:NWM

[Bar08] Mary Kathryn Barbier. Nigel West, Mortal Crimes: The Greatest Theft in History — The Soviet Penetration of the Manhattan Project. Journal of Cold War Studies, 10(1):136–138, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/231569;
http://muse.jhu.edu/article/231569/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/1/136/700053/jcws.2008.10.1.
136.pdf.

Barbier:2011:AMH

[Bar11a] Mary Kathryn Barbier. Allen M. Hornblum, The Invisible Harry Gold: The Man Who Gave the Soviets the Atom Bomb. New Haven: Yale University Press, 2010. 464 pp. $32.50. Journal of Cold War Studies, 13(3):210–212, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447471;   http://muse.jhu.edu/article/447471/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/210/700720/
jcws_r_00150.pdf.

Barrass:2011:DON

[Bar11b] Gordon S. Barrass. David Owen, Nuclear Papers. Liverpool, England: Liverpool University Press, 2009. 296 pp. £25.00. Journal of Cold War Studies, 13(3):218–220, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447475;   http://muse.jhu.edu/article/447475/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/218/700728/
jcws_r_00155.pdf.

Barrett:2018:CPD

[Bar18] Gordon Barrett. China’s “people’s diplomacy” and the Pugwash Conferences, 1957–1964. Journal of Cold War Studies, 20(1):140–169, Winter 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/691020;   http:
//muse.jhu.edu/article/691020/pdf;   https://direct.mit.
edu/jcws/article-pdf/20/1/140/699300/jcws_a_00803.pdf.

Barrett:2019:BPF

[Bar19] David M. Barrett. The Bay of Pigs Fiasco and the Kennedy Administration’s off-the-record briefings for journalists. Journal of Cold War Studies, 21(2):3–26, Spring 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/726148;   https://direct.mit.
edu/jcws/article-pdf/21/2/3/699593/jcws_a_00875.pdf.

Barnhisel:2020:MEP

[Bar20] Greg Barnhisel. Merve Emre, Paraliterary: The Making of Bad Readers in Postwar America. Journal of Cold War Studies, 22 (2):168–170, Spring 2020. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL https://direct.mit.edu/jcws/
article-pdf/22/2/168/1860558/jcws_r_00884.pdf;   https://
muse.jhu.edu/article/756830.

Barnhisel:2022:WBB

[Bar22] Greg Barnhisel. We begin bombing in five minutes: Late Cold War culture in the Age of Reagan. Journal of Cold War Studies, 24(3): 250–252, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/24/3/250/2042022/jcws_r_01098.pdf.

Baylis:2001:KMN

[Bay01] John Baylis. Kennedy, Macmillan and Nuclear Weapons (review). Journal of Cold War Studies, 3(1):126–128, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9166;   http://muse.jhu.edu/article/
9166/pdf;   https://direct.mit.edu/jcws/article-pdf/3/1/
126/695412/jcws.2001.3.1.126.pdf.

Bernard-Donals:2022:ITM

[BD22] Michael Bernard-Donals. Inside the teaching machine: Rhetoric and the globalization of the U.S. public research university. Journal of Cold War Studies, 24(3):246–248, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/24/3/246/2041989/jcws_r_01072.pdf.

Bearden:2006:TLB

[Bea06] Milt Bearden. Tom Lansford, A Bitter Harvest: U.S. Foreign Policy and Afghanistan. Journal of Cold War Studies, 8 (2):146–148, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196267;   http://muse.jhu.edu/article/196267/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/146/700528/
jcws.2006.8.2.146.pdf.

Bekes:2016:HSB

[Bék16] Csaba Békés. Hungary, the Soviet Bloc, the German question, and the CSCE Process, 1965–1975. Journal of Cold War Studies, 18(3):

95–138, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/632280;
http://muse.jhu.edu/article/632280/pdf;   https://direct.
mit.edu/jcws/article-pdf/18/3/95/698986/jcws_a_00654.pdf.

Bekes:2023:HDW

[Bék23] Csaba Békés. Hungary and the dissolution of the Warsaw Pact (1988–1991). Journal of Cold War Studies, 25(4):4–23, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/4/
4/2204989/jcws_a_01168.pdf.

Belmonte:2002:CWC

[Bel02] Laura A. Belmonte. Cold War Civil Rights: Race and the Image of American Democracy (review). Journal of Cold War Studies, 4(4): 108–110, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9295;   http:/
/muse.jhu.edu/article/9295/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/4/108/695666/jcws.2002.4.4.108.pdf.

Belknap:2009:ELH

[Bel09] Michal R. Belknap. Elizabeth Lutes Hillman, Defending America: Military Culture and the Cold War Court-Martial. Journal of Cold War Studies, 11(4):145–147, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/366252;   http://muse.jhu.edu/article/366252/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/4/145/
1627621/jcws.2009.11.4.145.pdf.

Belmonte:2011:HWM

[Bel11] Laura A. Belmonte. Hugh Wilford, The Mighty Wurlitzer: How the CIA Played America. Cambridge, MA: Harvard University Press, 2008. 342 pp. $27.95. Journal of Cold War Studies, 13(1):246–248, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/424258;   http:
//muse.jhu.edu/article/424258/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/1/246/697763/jcws_r_00089.pdf.

Belmonte:2013:SBW

[Bel13] Laura A. Belmonte. Susan A. Brewer, Why America Fights: Patriotism and Propaganda from the Philippines to Iraq. New York: Oxford University Press, 2009. 342 pp. Journal of Cold War Studies, 15(3):193–194, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522720;   http://muse.jhu.edu/article/522720/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/193/698297/
jcws_r_00347.pdf.

Bennett:2005:GDS

[Ben05] Andrew Bennett. The guns that didn’t smoke: Ideas and the Soviet non-use of force in 1989. Journal of Cold War Studies, 7(2): 81–109, Spring 2005. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/181897;
http://muse.jhu.edu/article/181897/pdf;   https://direct.
mit.edu/jcws/article-pdf/7/2/81/696209/1520397053630565.
pdf.

Benvenuti:2021:CRC

[Ben21] Andrea Benvenuti. Commonwealth Responsibility and Cold War Solidarity: Australia in Asia, 1944–74 by Dan Halvorson (review). Journal of Cold War Studies, 23(3):243–245, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/23/3/243/1955799/jcws_r_
01034.pdf;   https://muse.jhu.edu/article/801914.

Berliner:2001:SDI

[Ber01] Joseph S. Berliner. The Soviet Defence-Industry Complex from Stalin to Khrushchev (review). Journal of Cold War Studies, 3(3):112– 113, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9197;   http:/
/muse.jhu.edu/article/9197/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/112/695466/jcws.2001.3.3.112.pdf.

Berger:2004:FDH

[Ber04a] Stefan Berger. Forging Democracy: The History of the Left in Europe (review). Journal of Cold War Studies, 6(3):158–160, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171149;   http:
//muse.jhu.edu/article/171149/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/3/158/696085/jcws.2004.6.3.158.pdf.

Bernhard:2004:BIW

[Ber04b] Michael H. Bernhard. Beyond Invisible Walls: The Psychological Legacy of Soviet Trauma, East European Therapists and Their Patients (review). Journal of Cold War Studies, 6(2):109–111, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/53658;   http://
muse.jhu.edu/article/53658/pdf;   https://direct.mit.edu/
jcws/article-pdf/6/2/109/695971/jcws.2004.6.2.109.pdf.

Bernhard:2007:TKG

[Ber07] Michael H. Bernhard. Tomasz Kizny, Gulag: Life and Death Inside the Soviet Concentration Camps. Journal of Cold War Studies, 9(3):191–195, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219440;   http://muse.jhu.edu/article/219440/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/191/696658/
jcws.2007.9.3.191.pdf.

Bernstein:2013:AAG

[Ber13] Barton J. Bernstein. Analyzing and assessing Gaddis’s Kennan biography: Questionable interpretations and unpursued evidence and issues. Journal of Cold War Studies, 15(4):170–182, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/534415;   http:
//muse.jhu.edu/article/534415/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/4/170/698327/jcws_a_00401.pdf.

Bernstein:2014:MCC

[Ber14] Thomas P. Bernstein. Mao, Chinese Communism, and the USSR. Journal of Cold War Studies, 16(2):108–127, Spring 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/547326;   http:
//muse.jhu.edu/article/547326/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/2/108/698427/jcws_a_00453.pdf.

Berend:2015:CWC

[Ber15] Ivan T. Berend. Cold War Comforts: Canadian Women, Child Safety, and Global Insecurity by Tarah Brookfield (review). Journal of Cold War Studies, 17(3):282–283, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595105;   http://muse.jhu.edu/article/595105/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/282/
1102422/jcws_r_00525.pdf.

Bergmane:2020:EPI

[Ber20] Una Bergmane. “Is this the end of perestroika?”: International reactions to the Soviet use of force in the Baltic republics in January 1991. Journal of Cold War Studies, 22(2):26–57, Spring 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/22/2/26/1860548/jcws_a_
00939.pdf;   https://muse.jhu.edu/article/756824.

Best:2007:RCC

[Bes07] Geoffrey Francis Andrew Best. Raymond Callahan, Churchill and His Generals. Lawrence, KS: University of Kansas Press, 2007. 310 pp. Journal of Cold War Studies, 9(4):158–160, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/223531;   http:
//muse.jhu.edu/article/223531/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/4/158/696757/jcws.2007.9.4.158.pdf.

Berenbaum:2017:CPT

[BH17] Michael Berenbaum and Jeffrey Herf. Conflicting perspectives on Timothy Snyder’s Black Earth. Journal of Cold War Studies, 19(4):226–233, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683672;   http://muse.jhu.edu/article/683672/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/226/699358/
jcws_c_00772.pdf.

Bischof:2004:WCW

[Bis04] Günter Bischof. Waltzing into the Cold War: The Struggle for Occupied Austria (review). Journal of Cold War Studies, 6(3):141–144, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/171150;   http://muse.jhu.edu/article/171150/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/3/141/696075/
jcws.2004.6.3.141.pdf.

Bischof:2005:ANR

[Bis05] Günter Bischof. The advent of neo-revisionism? Journal of Cold War Studies, 7(1):141–151, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/179551;   http://muse.jhu.edu/article/179551/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/1/141/700468/
1520397053326176.pdf.

Bischof:2008:EBE

[Bis08] Günter Bischof. Ernst Bruckmüller, ed., Wiederaufbau in Österreich 1945–1955: Rekonstruktion oder Neubeginn? Journal of Cold War Studies, 10(1):163–165, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231570;   http://muse.jhu.edu/article/231570/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/163/700124/
jcws.2008.10.1.163.pdf.

Bischof:2010:SKO

[Bis10] Günter Bischof. Stefan Karner and Othmar Pickl, eds., Die Rote Armee in der Steiermark: Sowjetische Besatzung 1945. Journal of Cold War Studies, 12(1):175–177, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/382493;   http://muse.jhu.edu/article/382493/
pdf;   https://direct.mit.edu/jcws/article-pdf/12/1/175/
1627431/jcws.2010.12.1.175.pdf.

Bischof:2011:MRE

[Bis11] Günter Bischof. Manfried Rauchensteiner, ed., Zwischen den Blöcken: NATO, Warschauer Pakt und Österreich. Vienna: Böhlau Verlag, 2010. 817 pp. and Muriel Blaive and Berthold Molden, Grenzfälle: Österreichische und tschechische Erfahrungen am Eisernen Vorhang. Weitra, Austria: Verlag Bibliothek der Provinz, 2009. 271 pp. Journal of Cold War Studies, 13(4):261–265, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/457221;   http://muse.jhu.edu/article/457221/
pdf;   https://direct.mit.edu/jcws/article-pdf/13/4/261/
697806/jcws_r_00133.pdf.

Bischof:2012:NAF

[Bis12] Günter Bischof. Nicole Alecu de Flers, EU Foreign Policy and the Europeanization of Neutral States: Comparing Irish and Austrian Foreign Policy. London: Routledge, 2012. 167 pp. Journal of Cold War Studies, 14(3):206–207, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/490186;   http://muse.jhu.edu/article/490186/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/3/206/698065/
jcws_r_00259.pdf.

Bischof:2013:ESE

[Bis13] Günter Bischof. Elke Scherstjanoi, ed., Russlandheimkehrer: Die sowjetischen Kriegsgefangenen im Gedächtnis der Deutschen. Munich: Oldenbourg Verlag, 2012. 264 pp. Journal of Cold War Studies, 15(2):145–148, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/509712;   http://muse.jhu.edu/article/509712/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/2/145/698189/
jcws_r_00353.pdf.

Bischof:2014:EKK

[Bis14] Günter Bischof. Erbe des Kalten Krieges edited by Bernd Greiner, Tim B. Müller, and Klaas Voß (review). Journal of Cold War Studies, 16(4):237–239, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566664;   http://muse.jhu.edu/article/566664/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/237/698612/
jcws_r_00484.pdf.

Bittner:2017:DFA

[Bit17] Stephen V. Bittner. Dina Fainberg and Artemy M. Kalinovsky, eds., Reconsidering Stagnation in the Brezhnev Era: Ideology and Exchange. Journal of Cold War Studies, 19(4):247–249, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/683679;   http:
//muse.jhu.edu/article/683679/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/4/247/699445/jcws_r_00779.pdf.

Benvenuti:2010:ESA

[BJ10] Andrea Benvenuti and David Martin Jones. Engaging southeast asia?: Labor’s regional mythology and Australia’s military withdrawal from Singapore and Malaysia, 1972–1973. Journal of Cold War Studies, 12 (4):32–62, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/403796;
http://muse.jhu.edu/article/403796/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/4/32/697546/jcws_a_00047.pdf.

Benvenuti:2011:MMA

[BJ11] Andrea Benvenuti and David Martin Jones. Myth and misrepresentation in Australian foreign policy: Menzies and engagement with Asia. Journal of Cold War Studies, 13(4):57– 78, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/457193;
http://muse.jhu.edu/article/457193/pdf;   https://direct.
mit.edu/jcws/article-pdf/13/4/57/697776/jcws_a_00168.pdf.

Benvenuti:2019:FLA

[BJ19] Andrea Benvenuti and David Martin Jones. With friends like these: Australia, the United States, and Southeast Asian détente.

Journal of Cold War Studies, 21(2):27–57, Spring 2019. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/726149;   https://direct.
mit.edu/jcws/article-pdf/21/2/27/699543/jcws_a_00876.pdf.

Blight:2010:FWE

[BL10a] James G. Blight and Janet M. Lang. Forum: When empathy failed: Using critical oral history to reassess the collapse of U.S.– Soviet Détente in the Carter–Brezhnev years. Journal of Cold War Studies, 12(2):29–74, Spring 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/382199;   http://muse.jhu.edu/article/382199/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/2/29/697495/

jcws.2010.12.2.29.pdf. See commentary [Gar10a, Gar10c].

Blight:2010:RCU

[BL10b] James G. Blight and Janet M. Lang. Reply to the commentaries: Using critical oral history to reassess the collapse of U.S.– Soviet détente in the Carter–Brezhnev years. Journal of Cold War Studies, 12(2):102–109, Spring 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/382205;   http://muse.jhu.edu/article/382205/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/2/102/697501/
jcws.2010.12.2.102.pdf.

Blackwood:2003:MWE

[Bla03] Lee Blackwood. Martin Westlake, ed., Leaders of Transition. Journal of Cold War Studies, 5(3):141–143, Summer 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45183;   http://muse.jhu.edu/
article/45183/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/141/695895/jcws.2003.5.3.141.pdf.

Blank:2005:AWR

[Bla05] Stephen J. Blank. The Afghanistan Wars (review). Journal of Cold War Studies, 7(4):185–186, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188249;   http://muse.jhu.edu/article/188249/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/185/696331/
jcws.2005.7.4.185.pdf.

Blarel:2015:ABK

[Bla15] Nicolas Blarel. Andrew Bingham Kennedy, The International Ambitions of Mao and Nehru: National Efficacy Beliefs and the Making of Foreign Policy. New York: Cambridge University Press, 2012. 272 pp. $89.00. Journal of Cold War Studies, 17(4):181–184, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/606029;   http:
//muse.jhu.edu/article/606029/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/4/181/698769/jcws_r_00584.pdf.

Blessing:2016:DAR

[Ble16] Benita Blessing. Dolores L. Augustine, Red Prometheus: Engineering and Dictatorship in East Germany, 1945–1990. Cambridge, MA: MIT Press, 2007. 381 pp. $40.00. Journal of Cold War Studies, 18(4):247–249, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645930;   http://muse.jhu.edu/article/645930/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/247/700145/
jcws_r_00703.pdf.

Blumenau:2014:OBC

[Blu14] Bernhard Blumenau. The other battleground of the Cold War: The UN and the struggle against international terrorism in the 1970s. Journal of Cold War Studies, 16(1):61–84, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543725;   http:
//muse.jhu.edu/article/543725/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/1/61/698886/jcws_a_00431.pdf.

Blumenau:2015:MTO

[Blu15] Bernhard Blumenau. Mattia Toaldo, The Origins of the U.S. War on Terror: Lebanon, Libya and American Intervention in the Middle East. New York: Routledge, 2013. 214 pp. $140.00. Journal of Cold War Studies, 17(4):186–188, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606031;   http://muse.jhu.edu/article/606031/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/186/698794/
jcws_r_00586.pdf.

Bacevich:2012:RBG

[BM12] Andrew J. Bacevich and Edwin Moïse. Responses to Bernd Greiner on U.S. conduct in Vietnam. Journal of Cold War Studies, 14(1):111–113, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467685;   http://muse.jhu.edu/article/467685/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/111/697843/
jcws_c_00198.pdf.

Bacevich:2011:UCV

[BMLG11] Andrew J. Bacevich, Edwin Moïse, Mark Atwood Lawrence, and Bernd Greiner. U.S. conduct in the Vietnam War: Commentaries by Andrew J. Bacevich, Edwin Moïse, and Mark Atwood Lawrence on Bernd Greiner’s War without Fronts. Journal of Cold War Studies, 13(3):185–190, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447462;   http://muse.jhu.edu/article/447462/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/185/700690/
jcws_c_00147.pdf.

Bockman:2006:NGM

[Boc06] Johanna Bockman. Nils Gilman, Mandarins of the Future: Modernization Theory in Cold War America. Journal of Cold War Studies, 8(2):140–141, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196268;   http://muse.jhu.edu/article/196268/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/140/700532/
jcws.2006.8.2.140.pdf.

Bockman:2010:DEG

[Boc10] Johanna Bockman. David Ekbladh, The Great American Mission: Modernization and the Construction of an American World Order. Journal of Cold War Studies, 12(4):184–186, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.

jhu.edu/article/403802;   http://muse.jhu.edu/article/403802/
pdf;   https://direct.mit.edu/jcws/article-pdf/12/4/184/
697527/jcws_r_00054.pdf.

Boden:2008:CWE

[Bod08] Ragna Boden. Cold war economics: Soviet aid to Indonesia. Journal of Cold War Studies, 10(3):110–128, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/241499;   http://muse.jhu.edu/article/241499/
pdf;   https://direct.mit.edu/jcws/article-pdf/10/3/110/
697089/jcws.2008.10.3.110.pdf.

Boel:2021:ISH

[Boe21] Bent Boel. The international Sakharov hearings and transnational human rights activism, 1975–1985. Journal of Cold War Studies, 23(3): 81–137, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/23/3/81/1955785/jcws_a_01008.pdf;   https://muse.jhu.
edu/article/801906.

Bonds:2005:LLT

[Bon05] John Bledsoe Bonds. Looking for love (or tragedy) in all the wrong places. Journal of Cold War Studies, 7(1):152–158, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/179552;
http://muse.jhu.edu/article/179552/pdf;   https://direct.
mit.edu/jcws/article-pdf/7/1/152/700420/1520397053326121.
pdf.

Bonwetsch:2007:RSV

[Bon07] Bernd Bonwetsch. Rolf Steininger, Der vergessene Krieg: Korea 1950–1953. Journal of Cold War Studies, 9(1):136–137, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/210070;   http:
//muse.jhu.edu/article/210070/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/1/136/696527/jcws.2007.9.1.136.pdf.

Booth:2003:WDC

[Boo03] Ken Booth. War and Democracy: a Comparative Study of the Korean War and the Peloponnesian War (review). Journal of Cold War Studies, 5(2):104–106, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40755;   http://muse.jhu.edu/
article/40755/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/104/700380/jcws.2003.5.2.104.pdf.

Borhi:1999:RLC

[Bor99] László Borhi. Rollback, liberation, containment, or inaction?: U.S. policy and Eastern Europe in the 1950s. Journal of Cold War Studies, 1(3):67–110, Fall 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/9109;   https://direct.mit.edu/jcws/article-pdf/
1/3/67/695148/152039799316976814.pdf.

Borhi:2005:FDI

[Bor05a] László Borhi. The First Domino — International Decision Making during the Hungarian Crisis of 1956 (review). Journal of Cold War Studies, 7(4):181–185, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188250;   http://muse.jhu.edu/article/188250/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/181/696357/
jcws.2005.7.4.181.pdf.

Borhi:2005:WAD

[Bor05b] László Borhi. Was American diplomacy really tragic? Journal of Cold War Studies, 7(1):159–167, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/179553;   http://muse.jhu.edu/article/179553/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/1/159/700443/
1520397053326194.pdf.

Borhi:2009:PKH

[Bor09] László Borhi. Peter Kenez, Hungary from the Nazis to the Soviets: The Establishment of the Communist Regime in Hungary, 1944–1948. Journal of Cold War Studies, 11(2):133–137, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/265876;
http://muse.jhu.edu/article/265876/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/2/133/697185/jcws.2009.11.2.
133.pdf.

Borhi:2012:IHJ

[Bor12a] László Borhi. István Hargittai, Judging Edward Teller: a Closer Look at One of the Most Influential Scientists of the Twentieth Century. New York: Prometheus Books, 2010. 575 pp. Journal of Cold War Studies, 14(4):263–266, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/500353;   http://muse.jhu.edu/article/500353/
pdf;   http://muse.jhu.edu/journals/journal_of_cold_war_
studies/v014/14.4.borhi.pdf;   https://direct.mit.edu/jcws/
article-pdf/14/4/263/698191/jcws_r_00300.pdf.

Borhi:2012:SPO

[Bor12b] László Borhi. Secret peace overtures, the Holocaust, and Allied strategy vis-à-vis Germany: Hungary in the vortex of World War II. Journal of Cold War Studies, 14(2):29–67, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/478775;   http:
//muse.jhu.edu/article/478775/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/2/29/697973/jcws_a_00220.pdf.

Bose:2004:GEI

[Bos04] Meenekshi Bose. General Eisenhower: Ideology and Discourse (review). Journal of Cold War Studies, 6(3):136–138, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171151;   http:
//muse.jhu.edu/article/171151/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/3/136/696035/jcws.2004.6.3.136.pdf.

Bostdorff:2015:BCA

[Bos15] Denise M. Bostdorff. Blackwell Companions to American History: a Companion to Harry S. Truman edited by Daniel S. Margolies (review). Journal of Cold War Studies, 17(3):277–280, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595103;   http:
//muse.jhu.edu/article/595103/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/277/1102414/jcws_r_00523.pdf.

Bottoni:2017:FEE

[Bot17] Stefano Bottoni. Finding the enemy: Ethnicized state violence and population control in Ceauşescu’s Romania. Journal of Cold War Studies, 19(4):113–136, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683667;   http://muse.jhu.edu/article/683667/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/113/699391/
jcws_a_00766.pdf.

Boyer:2011:RJD

[Boy11] Paul Boyer. Robert A. Jacobs, The Dragon’s Tail: Americans Face the Atomic Age. Amherst: University of Massachusetts Press, 2010. 143 + xii pp. $80.00 cloth, $24.95 paper. Journal of Cold War Studies, 13(4):213–215, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457198;   http://muse.jhu.edu/article/457198/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/213/697823/
jcws_r_00176.pdf.

Boyadjieva:2023:BRD

[Boy23] Nadia G. Boyadjieva. Book review: Diplomatiya i diversiya na Balkanite: Britanskata politika kam Albaniya po vreme na i sled Vtorata svetovna voina [Diplomacy and Subversion in the Balkans: British Policy toward Albania during and after the Second World War] by Biser Petrov. Journal of Cold War Studies, 25(4):269–271, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/4/
269/2205008/jcws_r_01186.pdf.

Bozo:2015:FMC

[Boz15] Frédéric Bozo. “I feel more comfortable with you”: France, the Soviet Union, and German reunification. Journal of Cold War Studies, 17(3):116–158, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595073;   http://muse.jhu.edu/article/595073/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/116/
1102420/jcws_a_00563.pdf.

Bozo:2020:SGFa

[Boz20a] Frédéric Bozo. The sanctuary and the Glacis: France, the Federal Republic of Germany, and nuclear weapons in the 1980s (part 1). Journal of Cold War Studies, 22(3):119–179, Summer 2020. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/22/3/119/1860583/
jcws_a_00929.pdf;   https://muse.jhu.edu/article/762894.

Bozo:2020:SGFb

[Boz20b] Frédéric Bozo. The sanctuary and the Glacis: France, the Federal Republic of Germany, and the nuclear factor in the 1980s (part 2). Journal of Cold War Studies, 22(4):175–228, Fall 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/22/4/175/1860634/jcws_a_
00970.pdf;   https://muse.jhu.edu/article/776963.

Brands:2002:CWR

[Bra02a] H. W. Brands. Cold War Respite: The Geneva Summit of 1955 (review). Journal of Cold War Studies, 4(1):88–89, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9221;   http:/
/muse.jhu.edu/article/9221/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/1/88/695510/152039702753344852.pdf.

Brands:2002:DRD

[Bra02b] H. W. Brands. Dean Rusk: Defending the American Mission Abroad (review). Journal of Cold War Studies, 4(2):150–152, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9244;   http:/
/muse.jhu.edu/article/9244/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/2/150/695613/jcws.2002.4.2.150.pdf.

Brandenberger:2004:SSP

[Bra04] David Brandenberger. Stalin’s Secret Pogrom: The Postwar Inquisition of the Jewish Anti-Fascist Committee (review). Journal of Cold War Studies, 6(3):172–174, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/171152;   http://muse.jhu.edu/article/171152/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/3/172/696131/
jcws.2004.6.3.172.pdf.

Brandenberger:2005:SHW

[Bra05] David Brandenberger. Stalin’s Holy War: Religion, Nationalism, and Alliance Politics, 1941–1945 (review). Journal of Cold War Studies, 7(3):196–197, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185092;   http://muse.jhu.edu/article/185092/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/196/1102589/
jcws.2005.7.3.196.pdf.

Brands:2006:RNL

[Bra06] Hal Brands. Rethinking nonproliferation: LBJ, the gilpatric committee, and U.S. national security policy. Journal of Cold War Studies, 8 (2):83–113, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/196269;
http://muse.jhu.edu/article/196269/pdf;   https://direct.
mit.edu/jcws/article-pdf/8/2/83/700572/jcws.2006.8.2.83.
pdf.

Brady:2014:DZI

[Bra14] Lisa M. Brady. David Zierler, The Invention of Ecocide: Agent Orange, Vietnam, and the Scientists Who Changed the Way We Think about the Environment. Athens, GA: University of Georgia Press, 2011. 252 pp. $24.95. Journal of Cold War Studies, 16(1):224–226, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/543738;   http://muse.jhu.edu/article/543738/pdf;

 https://direct.mit.edu/jcws/article-pdf/16/1/224/698950/
jcws_r_00416.pdf.

Braithwaite:2016:SAU

[Bra16a] Sir Rodric Braithwaite. Svetlana Alexievich, U voiny — ne zhenskoe litso: Poslednie svideteli [War Does Not Have a Woman’s Face: The Latest Witnesses]. Minsk: Mastatskaya litaratura, 1985. Svetlana Alexievich, Zinky Boys: Soviet Voices from the Afghanistan War, trans. by Julia Whitby and Robin Whitby. London: Chatto, 1992. 197 pp. Journal of Cold War Studies, 18(3):231–233, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632303;   http://muse.jhu.edu/article/632303/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/231/699069/
jcws_r_00675.pdf.

Brandenberger:2016:SFS

[Bra16b] David Brandenberger. Sheila Fitzpatrick, On Stalin’s Team: The Years of Living Dangerously in Soviet Politics. Princeton, NJ: Princeton University Press, 2015. 384 pp. $35.00/£24.95. Journal of Cold War Studies, 18(4):230–233, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645923;   http://muse.jhu.edu/article/645923/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/230/700136/
jcws_r_00695.pdf.

Breitman:2008:SAP

[Bre08] Richard Breitman. Safehaven: The Allied Pursuit of Nazi Assets Abroad (review). Journal of Cold War Studies, 10(3):164– 166, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/241514;
http://muse.jhu.edu/article/241514/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/3/164/697101/jcws.2008.10.3.
164.pdf.

Brewer:2021:OWB

[Bre21] Susan A. Brewer. One World, Big Screen: Hollywood, the Allies, and World War II, by M. Todd Bennett, Chapel Hill: University of North Carolina Press, 2012. 362 pp. (review). Journal of Cold War Studies, 23(1):248–250, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/1/248/1911796/jcws_r_00997.pdf;
https://muse.jhu.edu/article/788386.

Brier:2013:BCH

[Bri13] Robert Brier. Broadening the cultural history of the Cold War: The emergence of the Polish workers’ defense committee and the rise of human rights. Journal of Cold War Studies, 15(4):104–127, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/534425;   http:
//muse.jhu.edu/article/534425/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/4/104/698301/jcws_a_00396.pdf.

Briggs:2015:NAB

[Bri15] Ryan C. Briggs. Neopatrimonialism in Africa and Beyond edited by Daniel Bach and Mamoudou Gazibo (review). Journal of Cold War Studies, 17(3):264–266, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595096;   http://muse.jhu.edu/article/595096/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/264/
1102428/jcws_r_00526.pdf.

Brogi:2002:AAJ

[Bro02a] Alessandro Brogi. Allies and Adversaries: The Joint Chiefs of Staff, The Grand Alliance, and U.S. Strategy in World War II (review). Journal of Cold War Studies, 4(3):129–132, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9269;   http:/
/muse.jhu.edu/article/9269/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/129/700236/jcws.2002.4.3.129.pdf.

Brogi:2002:IIE

[Bro02b] Alessandro Brogi. Ike and Italy: The Eisenhower Administration and Italy’s “Neo-Atlanticist” agenda. Journal of Cold War Studies, 4(3): 5–35, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9268;   http:/
/muse.jhu.edu/article/9268/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/5/700234/152039702320201058.pdf.

Brown:2004:FYW

[Bro04] Seyom Brown. The Fifty-Year Wound: The True Price of America’s Cold War Victory (review). Journal of Cold War Studies, 6(4): 145–146, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/174705;
http://muse.jhu.edu/article/174705/pdf;   https://direct.
mit.edu/jcws/article-pdf/6/4/145/696050/1520397042350919.
pdf.

Brovkin:2006:VMN

[Bro06] Vladimir N. Brovkin. Vixi: Memoirs of a Non-Belonger (review). Journal of Cold War Studies, 8(1):127–132, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/192205;   http:
//muse.jhu.edu/article/192205/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/1/127/696305/jcws.2006.8.1.127.pdf.

Brown:2008:CEB

[Bro08] Archie Brown. The change to engagement in Britain’s Cold War policy: The origins of the Thatcher–Gorbachev relationship. Journal of Cold War Studies, 10(3):3–47, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/241496;   http://muse.jhu.edu/article/241496/
pdf;   https://direct.mit.edu/jcws/article-pdf/10/3/3/697167/
jcws.2008.10.3.3.pdf.

Brogi:2010:STA

[Bro10] Alessandro Brogi. Simona Tobia, Advertising America: The United States Information Service in Italy (1945–1956). Journal of Cold War Studies, 12(3):133–137, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392785;   http://muse.jhu.edu/article/392785/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/133/697625/
jcws_r_00012.pdf.

Brogi:2011:SPB

[Bro11] Alessandro Brogi. Silvio Pons, Berlinguer e la fine del comunismo. Turin: Einaudi Editore, 2006. xxiv + 265 pp. Journal of Cold War Studies, 13(4):274–278, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457226;   http://muse.jhu.edu/article/457226/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/274/697808/
jcws_r_00158.pdf.

Brown:2015:ESU

[Bro15] Archie Brown. The end of the Soviet Union. Journal of Cold War Studies, 17(4):158–165, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606024;   http://muse.jhu.edu/article/606024/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/158/698721/
jcws_a_00600.pdf.

Brown:2022:BRD

[Bro22] Archie Brown. Book review: A Diplomatic Meeting: Reagan, Thatcher, and the Art of Summitry by James Cooper. Journal of Cold War Studies, 24(4):223–226, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/24/4/223/2062880/jcws_r_01110.pdf.

Bruce:2001:EGL

[Bru01] Gary Bruce. The East German Leadership, 1946–73: Conflict and Crisis (review). Journal of Cold War Studies, 3(1):135–137, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9167;   http:/
/muse.jhu.edu/article/9167/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/1/135/695420/jcws.2001.3.1.135.pdf.

Bruce:2003:PNS

[Bru03] Gary S. Bruce. The prelude to nationwide surveillance in East Germany: Stasi operations and threat perceptions, 1945–1953. Journal of Cold War Studies, 5(2):3–31, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40756;   http://muse.jhu.edu/
article/40756/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/3/700402/152039703763336453.pdf.

Bruce:2004:SPR

[Bru04] Gary Bruce. The Secret Police and the Revolution: The Fall of the German Democratic Republic (review). Journal of Cold War Studies, 6(2):106–109, Spring 2004. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
53659;   http://muse.jhu.edu/article/53659/pdf;   https:/
/direct.mit.edu/jcws/article-pdf/6/2/106/695961/jcws.2004.
6.2.106.pdf.

Bruce:2005:KKB

[Bru05] Gary Bruce. Kalter Krieg: Beitrage zur Ost–West-Konfrontation 1945 bis 1990 (review). Journal of Cold War Studies, 7(4):179–181, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/188251;   http:
//muse.jhu.edu/article/188251/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/4/179/696347/jcws.2005.7.4.179.pdf.

Bruce:2006:JSE

[Bru06] Gary Bruce. Joachim Scholtyseck, Enzyklopädie deutscher Geschichte, Vol. 69, Die Aussenpolitik der DDR. Journal of Cold War Studies, 8(4):155–157, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/204616;   http://muse.jhu.edu/article/204616/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/4/155/700797/
jcws.2006.8.4.155.pdf.

Bruce:2008:DGU

[Bru08] Gary Bruce. Dominik Geppert and Udo Wengst, eds., Neutralität — Chance oder Chimäre? Konzepte des Dritten Weges für Deutschland und die Welt 1945–1990 (German) [Neutrality — A Chance or a Chimera: The Concept of the Third Way for Germany and the World, 1945–1990]. Journal of Cold War Studies, 10(3):183–185, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241523;   http://muse.jhu.edu/article/241523/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/183/697133/
jcws.2008.10.3.183.pdf.

Bruce:2009:REI

[Bru09] Gary Bruce. Roger Engelmann and Ilko-Sascha Kowalczuk, eds., Volkserhebung gegen den SED-Staat: Eine Bestandsaufnahme zum 17. Juni 1953. Journal of Cold War Studies, 11(2):124–125, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/265880;
http://muse.jhu.edu/article/265880/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/2/124/697161/jcws.2009.11.2.
124.pdf.

Bruce:2010:MFP

[Bru10] Gary Bruce. Mary Fulbrook, The People’s State: East German Society from Hitler to Honecker. Journal of Cold War Studies, 12(3):137–140, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392786;   http://muse.jhu.edu/article/392786/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/137/697574/
jcws_r_00013.pdf.

Bruce:2011:KMS

[Bru11] Gary Bruce. Kristie Macrakis, Seduced by Secrets: Inside the Stasi’s Spy-Tech World. New York: Cambridge University Press, 2008.

370 pp. $28.00. Journal of Cold War Studies, 13(4):259–260, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457220;   http:
//muse.jhu.edu/article/457220/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/4/259/697845/jcws_r_00132.pdf.

Bruce:2016:CLN

[Bru16] Gary Bruce. Charles Lansing, From Nazism to Communism: German Schoolteachers under Two Dictatorships. Cambridge, MA: Harvard University Press, 2010. 340 pp. Journal of Cold War Studies, 18(3):211–212, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632294;   http://muse.jhu.edu/article/632294/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/211/698981/
jcws_r_00646.pdf.

Bowlt:2002:KFE

[BS02] John E. Bowlt and Dmitrii Sarab’yanov. Keepers of the flame: an exchange on art and Western cultural influences in the USSR after World War II. Journal of Cold War Studies, 4(1):81–87, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9220;   http:/
/muse.jhu.edu/article/9220/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/1/81/695491/152039702753344843.pdf.

Brown:2023:EDS

[BSK^{+}23] Archie Brown, Thomas W. Simons, Jr., Ivan Kurilla, Andrea Graziosi, Louis D. Sell, and Vladislav Zubok. Evaluating the demise of the Soviet Union: Vladislav M. Zubok, Collapse: The Fall of the Soviet Union. New Haven: Yale University Press, 2021. 535pp. $35.00. Journal of Cold War Studies, 25(3):142–187, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/3/
142/2158668/jcws_c_01162.pdf.

Buclin:2017:SIC

[Buc17] Hadrien Buclin. Swiss intellectuals and the Cold War: Anticommunist policies in a neutral country. Journal of Cold War Studies, 19(4):137–167, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683668;   http://muse.jhu.edu/article/683668/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/137/699413/
jcws_a_00767.pdf.

Budjeryn:2022:NPS

[Bud22] Mariana Budjeryn. Non-proliferation and state succession: The demise of the USSR and the nuclear aftermath in Belarus, Kazakhstan, and Ukraine. Journal of Cold War Studies, 24(2):46–94, Spring 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/2/
46/2013030/jcws_a_01077.pdf.

Bugge:2015:WDC

[Bug15] Peter Bugge. Worlds of Dissent: Charter 77, the Plastic People of the Universe, and Czech Culture under Communism by Jonathan Bolton (review). Journal of Cold War Studies, 17(3):294–295, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595111;   http:
//muse.jhu.edu/article/595111/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/294/1102438/jcws_r_00557.pdf.

Buhle:2015:AHW

[Buh15] Paul Buhle. Those About Him Who Remained Silent: The Battle over W. E. B. Du Bois by Amy Bass (review). Journal of Cold War Studies, 17(3):266–267, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595097;   http://muse.jhu.edu/article/595097/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/266/
1102436/jcws_r_00527.pdf.

Burr:2004:CPW

[Bur04] William Burr. Conducting Post-World War II National Security Research in Executive Branch Records: a Comprehensive Guide (review). Journal of Cold War Studies, 6(3):128–130, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171153;   http:
//muse.jhu.edu/article/171153/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/3/128/696019/jcws.2004.6.3.128.pdf.

Burr:2005:NAH

[Bur05] William Burr. The Nixon Administration, the “horror strategy,” and the search for limited nuclear options, 1969–1972: Prelude to the Schlesinger doctrine. Journal of Cold War Studies, 7(3):34– 78, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/185096;

http://muse.jhu.edu/article/185096/pdf;   https://direct.
mit.edu/jcws/article-pdf/7/3/34/1102571/1520397054377188.
pdf.

Burr:2008:JSH

[Bur08] William Burr. Jeremi Suri, Henry Kissinger and the American Century. Journal of Cold War Studies, 10(3):153–156, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/241509;
http://muse.jhu.edu/article/241509/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/3/153/697085/jcws.2008.10.3.
153.pdf.

Burr:2010:JHR

[Bur10] William Burr. John H. Rubel, Doomsday Delayed: USAF Strategic Weapons Doctrine and SIOP-62, 1959–1962: Two Cautionary Tales. Journal of Cold War Studies, 12(4):188–190, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/403804;   http:
//muse.jhu.edu/article/403804/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/4/188/697556/jcws_r_00056.pdf.

Burton:2013:TAJ

[Bur13] Brian M. Burton. Thomas L. Ahern, Jr., Vietnam Declassified: The CIA and Counterinsurgency. Lexington, KY: The University Press of Kentucky, 2010. 450 pp. $40.00. Journal of Cold War Studies, 15(2):168–169, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/509722;   http://muse.jhu.edu/article/509722/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/2/168/698221/
jcws_r_00363.pdf.

Burke:2016:DJS

[Bur16] Roland Burke. Daniel J. Sargent, A Superpower Transformed: The Remaking of American Foreign Relations in the 1970s. New York: Oxford University Press, 2015. 287 pp. Journal of Cold War Studies, 18(4):228–230, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645922;   http://muse.jhu.edu/article/645922/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/228/700107/
jcws_r_00694.pdf.

Burr:2017:KGB

[Bur17] William Burr. To “keep the genie bottled up”: U.S. diplomacy, nuclear proliferation, and gas centrifuge technology, 1962–1972. Journal of Cold War Studies, 19(2):115–157, Spring 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http:

//www.mitpressjournals.org/doi/full/10.1162/JCWS_a_00743;
 https://direct.mit.edu/jcws/article-pdf/19/2/115/699199/
jcws_a_00743.pdf.

Burlingham:2019:PJW

[Bur19] Kate Burlingham. Praying for justice: The World Council of Churches and the program to combat racism. Journal of Cold War Studies, 21 (1):66–96, Winter 2019. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/723553;
 https://direct.mit.edu/jcws/article-pdf/21/1/66/699773/
jcws_a_00856.pdf.

Burr:2021:CAD

[Bur21] William Burr. The Carter Administration’s “Damnable dilemma”: How to respond to Pakistan’s secret nuclear weapons program, 1978– 1979. Journal of Cold War Studies, 23(1):4–54, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/23/1/4/1911794/jcws_a_
00980.pdf;   https://muse.jhu.edu/article/788379.

Burgard:2023:BRD

[Bur23] Antoine Burgard. Book review: Destination Elsewhere: Displaced Persons and Their Quest to Leave Postwar Europe by Ruth Balint. Journal of Cold War Studies, 25(1):208–210, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/1/
208/2073767/jcws_r_01126.pdf.

Busch:2014:VLW

[Bus14] Peter Busch. The “Vietnam Legion”: West German psychological warfare against East German propaganda in the 1960s. Journal of Cold War Studies, 16(3):164–189, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/556079;   http://muse.jhu.edu/article/556079/
pdf;   https://direct.mit.edu/jcws/article-pdf/16/3/164/
698486/jcws_a_00472.pdf.

Buthe:2005:EIS

[Büt05] Tim Büthe. European integration 1950–2003: Superstate or new market economy? Journal of Cold War Studies, 7(4):166–168, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/188252;   http:
//muse.jhu.edu/article/188252/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/4/166/696311/jcws.2005.7.4.166.pdf.

Byrne:2001:UKA

[Byr01] Malcolm Byrne. Undermining the Kremlin: America’s Strategy to Subvert the Soviet Bloc, 1947–1956 (review). Journal of Cold War Studies, 3(2):103–105, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9185;   http://muse.jhu.edu/article/9185/pdf;
 https://direct.mit.edu/jcws/article-pdf/3/2/103/695384/
jcws.2001.3.2.103.pdf.

Byrne:2012:RJR

[Byr12] Malcolm Byrne. A. Ross Johnson, Radio Free Europe and Radio Liberty: The CIA Years and Beyond. Washington, DC: Wilson Center Press, 2010. 304 pp. $55.00. Journal of Cold War Studies, 14(3):213–215, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/490189;   http://muse.jhu.edu/article/490189/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/3/213/698063/
jcws_r_00262.pdf.

Caldwell:2002:BCE

[Cal02] Peter C. Caldwell. The Break-up of Communism in East Germany and Eastern Europe (review). Journal of Cold War Studies, 4(4): 130–132, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9296;   http:/
/muse.jhu.edu/article/9296/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/4/130/695734/jcws.2002.4.4.130.pdf.

Caldwell:2003:RMC

[Cal03] Peter C. Caldwell. The Role of the Masses in the Collapse of the GDR (review). Journal of Cold War Studies, 5(3):150–152, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45184;   http://muse.jhu.edu/
article/45184/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/150/695941/jcws.2003.5.3.150.pdf.

Caldwell:2006:JHF

[Cal06] Dan Caldwell. Jussi Hanhimäki, The Flawed Architect: Henry Kissinger and American Foreign Policy. Journal of Cold War Studies, 8(2):138–139, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196270;   http://muse.jhu.edu/article/196270/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/138/700522/
jcws.2006.8.2.138.pdf.

Caldwell:2007:MLDa

[Cal07a] Peter C. Caldwell. Mark Landsman, Dictatorship and Demand: The Politics of Consumerism in East Germany. Journal of Cold War Studies, 9(1):147–148, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210071;   http://muse.jhu.edu/article/210071/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/147/696538/
jcws.2007.9.1.147.pdf.

Caldwell:2007:MLDb

[Cal07b] Peter C. Caldwell. Mark Landsman, Dictatorship and Demand: The Politics of Consumerism in East Germany. Journal of Cold War Studies, 9(3):195–197, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219441;   http://muse.jhu.edu/article/219441/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/195/696672/
jcws.2007.9.3.195.pdf.

Caldwell:2009:FLA

[Cal09] Dan Caldwell. Fredrik Logevall and Andrew Preston, eds., Nixon in the World: American Foreign Relations, 1969. Journal of Cold War Studies, 11(4):164–165, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/366253;   http://muse.jhu.edu/article/366253/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/4/164/
1627635/jcws.2009.11.4.164.pdf.

Calico:2014:AFS

[Cal14] Joy H. Calico. Annegret Fauser, Sounds of War: Music in the United States during World War II. New York: Oxford University Press, 2013. 366 pp. $39.95. Journal of Cold War Studies, 16(2):132–134, Spring 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/547330;   http://muse.jhu.edu/article/547330/pdf;

 https://direct.mit.edu/jcws/article-pdf/16/2/132/698437/
jcws_r_00457.pdf.

Caldwell:2022:SEM

[Cal22] Peter C. Caldwell. Selling the economic miracle: Economic reconstruction and politics in West Germany, 1949–1957. Journal of Cold War Studies, 24(3):244–246, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/24/3/244/2042003/jcws_r_01071.pdf.

Campbell:2005:NCR

[Cam05] Robert Wellington Campbell. From Newspeak to Cyberspeak (review). Journal of Cold War Studies, 7(4):189–191, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/188253;   http:
//muse.jhu.edu/article/188253/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/4/189/696355/jcws.2005.7.4.189.pdf.

Capshaw:2008:KMA

[Cap08] Ron Capshaw. Kevin Morgan et al., eds., Agents of Revolution: Biographical Studies in the Age of Lenin and Stalin. Journal of Cold War Studies, 10(1):143–144, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231571;   http://muse.jhu.edu/article/231571/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/143/700061/
jcws.2008.10.1.143.pdf.

Caprio:2014:AFP

[Cap14] Mark E. Caprio. American Foreign Policy and Postwar Reconstruction: Comparing Japan and Iraq by Jeff Bridoux (review). Journal of Cold War Studies, 16(4):264–265, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566677;   http://muse.jhu.edu/article/566677/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/264/698674/
jcws_r_00496.pdf.

Carafano:1999:MES

[Car99] James Jay Carafano. Mobilizing Europe’s stateless: America’s plan for a Cold War Army. Journal of Cold War Studies, 1(2):61–85, Spring 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/9098;   https://direct.
mit.edu/jcws/article-pdf/1/2/61/695131/15203979952559522.
pdf.

Carson:2011:JEE

[Car11] Mark Carson. Jeffrey A. Engel, ed., Local Consequences of the Global Cold War. Washington, DC: Woodrow Wilson Center Press; 2007. 224 pp. $65.00. Journal of Cold War Studies, 13(3):225–227, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447479;   http://muse.jhu.edu/article/447479/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/225/700696/
jcws_r_00112.pdf.

Carafano:2013:JMH

[Car13a] James Jay Carafano. Jonathan M. House, A Military History of the Cold War, 1944–1962. Norman: University of Oklahoma Press, 2012. 546 pp. $45.00. Journal of Cold War Studies, 15(3):185–187, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522716;   http://muse.jhu.edu/article/522716/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/185/698329/
jcws_r_00377.pdf.

Carlback:2013:JFS

[Car13b] Helene Carlbäck. Juliane Fürst, Stalin’s Last Generation: Soviet Post-War Youth and the Emergence of Mature Socialism. New York: Oxford University Press, 2010. xiv + 391 pp. $99.00. Journal of Cold War Studies, 15(3):216–217, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522732;   http://muse.jhu.edu/article/522732/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/216/698363/
jcws_r_00387.pdf.

Carvin:2017:CTC

[Car17] Stephanie Carvin. Conventional thinking? The 1980 convention on certain conventional weapons and the politics of legal restraints on weapons during the Cold War. Journal of Cold War Studies, 19(1): 38–69, Winter 2017. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://www.mitpressjournals.org/doi/
full/10.1162/JCWS_a_00717;   https://direct.mit.edu/jcws/
article-pdf/19/1/38/699098/jcws_a_00717.pdf.

Catton:2005:AWJ

[Cat05] Philip E. Catton. All the Way with JFK?: Britain, the U.S., and the Vietnam War (review). Journal of Cold War Studies, 7(4):148–150, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188254;   http://muse.jhu.edu/article/188254/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/148/696315/
jcws.2005.7.4.148.pdf.

Catton:2008:LUB

[Cat08] Philip E. Catton. The Lotus Unleashed: The Buddhist Peace Movement in South Vietnam, 1964–1966 (review). Journal of Cold War Studies, 10(2):147–149, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/237156;   http://muse.jhu.edu/article/237156/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/2/147/696993/
jcws.2008.10.2.147.pdf.

Cavoski:2019:YHW

[Čav19] Jovan Čavoški. Yugoslavia’s help was extraordinary: Political and material assistance from Belgrade to the MPLA in its rise to power, 1961–1975. Journal of Cold War Studies, 21(1):125–150, Winter 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/723555;   https://direct.mit.
edu/jcws/article-pdf/21/1/125/699725/jcws_a_00857.pdf.

Ceadel:2006:LGS

[Cea06] Martin Ceadel. Louise Grace Shaw, The British Political Elite and the Soviet Union 1937–1939. Journal of Cold War Studies, 8(2):163–164, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196271;   http://muse.jhu.edu/article/196271/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/163/700560/
jcws.2006.8.2.163.pdf.

Ceadel:2009:RHB

[Cea09] Martin Ceadel. Ronald Hyam, Britain’s Declining Empire: The Road to Decolonisation, 1918–1968. Journal of Cold War Studies, 11(2):155–156, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/265866;   http://muse.jhu.edu/article/265866/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/2/155/697203/
jcws.2009.11.2.155.pdf.

Ceadel:2017:JHB

[Cea17] Martin Ceadel. Jonathan Hogg, British Nuclear Culture: Official and Unofficial Narratives in the Long 20th Century. London: Bloomsbury, 2016. 231 pp. Journal of Cold War Studies, 19(3):234–236, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668168;   http://muse.jhu.edu/article/668168/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/234/699243/
jcws_r_00728.pdf.

Cesari:2021:SDE

[Ces21] Laurent Cesari. Safe for Decolonization: The Eisenhower Administration, Britain, and Singapore by S. R. Joey Long (review). Journal of Cold War Studies, 23(3):250–252, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/23/3/250/1955771/jcws_r_
01015.pdf;   https://muse.jhu.edu/article/801917.

Chapman:2000:WTR

[Cha00] James Chapman. 1968: The World Transformed (review). Journal of Cold War Studies, 2(1):130–132, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9120;   http://muse.jhu.edu/article/9120/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/1/130/695199/
jcws.2000.2.1.130.pdf.

Chang:2004:REC

[Cha04] Gordon H. Chang. Re-examining the Cold War: U.S.–China Diplomacy, 1954–1973 (review). Journal of Cold War Studies, 6(2):80–82, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/53660;   http://
muse.jhu.edu/article/53660/pdf;   https://direct.mit.edu/
jcws/article-pdf/6/2/80/695986/jcws.2004.6.2.80.pdf.

Chang:2007:CKC

[Cha07a] Gordon H. Chang. Christina Klein, Cold War Orientalism: Asia in the Middlebrow Imagination, 1945–1961. Journal of Cold War Studies, 9(1):137–139, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210072;   http://muse.jhu.edu/article/210072/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/137/696544/
jcws.2007.9.1.137.pdf.

Chapman:2007:CWL

[Cha07b] Roger Chapman. Cold war legacies: The migration and transformation of popular/unpopular culture. Journal of Cold War Studies, 9(3):137–143, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219442;   http://muse.jhu.edu/article/219442/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/137/696626/
jcws.2007.9.3.137.pdf.

Chandler:2008:JNE

[Cha08] David Chandler. Jonathan Nashel, Edward Lansdale’s Cold War. Journal of Cold War Studies, 10(3):187–189, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/241525;   http:
//muse.jhu.edu/article/241525/pdf;   https://direct.mit.
edu/jcws/article-pdf/10/3/187/697123/jcws.2008.10.3.187.
pdf.

Chang:2014:CWT

[Cha14] Gordon G. Chang. A Cold War Turning Point: Nixon and China, 1969–1972 by Chris Tudda (review). Journal of Cold War Studies, 16(3):218–219, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/556084;   http://muse.jhu.edu/article/556084/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/3/218/698511/
jcws_r_00460.pdf.

Chase:2021:SCC

[Cha21] William J. Chase. Secret Cables of the Comintern, 1933–1943 ed. by Fridrikh I. Firsov, Harvey Klehr and John Earl Haynes (review). Journal of Cold War Studies, 23(2):168–170, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/23/2/168/1919919/jcws_r_
01018.pdf;   https://muse.jhu.edu/article/794718.

Chen:1999:BWI

[Che99] Jian Chen. Battling Western Imperialism: Mao, Stalin, and the United States (review). Journal of Cold War Studies, 1(3):176–177, Fall 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9110;   https://
direct.mit.edu/jcws/article-pdf/1/3/176/695172/jcws.1999.
1.3.176.pdf.

Chen:2006:TRC

[Che06] Jian Chen. The Tibetan Rebellion of 1959 and China’s changing relations with India and the Soviet Union. Journal of Cold War Studies, 8(3):54–101, Summer 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/201236;   http://muse.jhu.edu/article/201236/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/3/54/696419/jcws.
2006.8.3.54.pdf.

Cheng:2007:SCR

[Che07] Yinghong Cheng. Sino–Cuban relations during the early years of the Castro Regime, 1959–1966. Journal of Cold War Studies, 9(3): 78–114, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/219443;
http://muse.jhu.edu/article/219443/pdf;   https://direct.
mit.edu/jcws/article-pdf/9/3/78/696688/jcws.2007.9.3.78.
pdf.

Cheng:2008:MAL

[Che08a] Grace Cheng. Mark Atwood Lawrence and Fredrik Logevall, eds., The First Vietnam War: Colonial Conflict and Cold War Crisis. Journal of Cold War Studies, 10(4):158–160, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/252670;   http://muse.jhu.edu/article/252670/
pdf;   https://direct.mit.edu/jcws/article-pdf/10/4/158/
697065/jcws.2008.10.4.158.pdf.

Chernus:2008:CTT

[Che08b] Ira Chernus. Chris Tudda, The Truth Is Our Weapon: The Rhetorical Diplomacy of Dwight D. Eisenhower and John Foster Dulles. Journal of Cold War Studies, 10(3):133–134, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/241501;
http://muse.jhu.edu/article/241501/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/3/133/697079/jcws.2008.10.3.
133.pdf.

Chernus:2010:DBP

[Che10] Ira Chernus. Denise Bostdorff, Proclaiming the Truman Doctrine: The Cold War Call to Arms. Journal of Cold War Studies, 12(3):121–123, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392781;   http://muse.jhu.edu/article/392781/pdf;

 https://direct.mit.edu/jcws/article-pdf/12/3/121/697643/
jcws_r_00008.pdf.

Chen:2018:JKH

[Che18] Kai Chen. Jan K. Herman, The Lucky Few: The Fall of Saigon and the Rescue Mission of the USS Kirk. Journal of Cold War Studies, 20 (4):237–239, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/718210;
https://direct.mit.edu/jcws/article-pdf/20/4/237/1969626/
jcws_r_00828.pdf.

Chen:2019:MLJ

[Che19] Kai Chen. So Much to Lose: John F. Kennedy and American Policy in Laos by William J. Rust (review). Journal of Cold War Studies, 21(1): 206–207, Winter 2019. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/723544;
 https://direct.mit.edu/jcws/article-pdf/21/1/206/699793/
jcws_r_00865.pdf.

Chervin:2020:CAM

[Che20] Reed Chervin. “Cartographic aggression”: Media politics, propaganda, and the Sino–Indian border dispute. Journal of Cold War Studies, 22(3):225–247, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/3/225/1860581/jcws_a_00911.pdf;
https://muse.jhu.edu/article/762896.

Chirot:2004:DCP

[Chi04] Daniel Chirot. Dreamworld and Catastrophe: The Passing of Mass Utopia in East and West (review). Journal of Cold War Studies, 6(2):84–86, Spring 2004. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
53661;   http://muse.jhu.edu/article/53661/pdf;   https:/
/direct.mit.edu/jcws/article-pdf/6/2/84/695980/jcws.2004.
6.2.84.pdf.

Chirot:2014:MST

[Chi14] Daniel Chirot. Marci Shore, A Taste of Ashes: The Afterlife of Totalitarianism in Eastern Europe. New York: Crown Publishers, 2013. 370 pp. Martin Mevius, ed., The Communist Quest for National Legitimacy in Europe, 1918–1989. New York: Routledge, 2013. 174 pp. Journal of Cold War Studies, 16(1):245–249, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543748;   http:
//muse.jhu.edu/article/543748/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/1/245/698989/jcws_r_00439.pdf.

Chou:2010:CEC

[Cho10] Grace Ai-Ling Chou. Cultural education as containment of Communism: The ambivalent position of American NGOs in Hong Kong in the 1950s. Journal of Cold War Studies, 12(2):3–28, Spring 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/382198;   http:
//muse.jhu.edu/article/382198/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/2/3/697503/jcws.2010.12.2.3.pdf.

Church:2007:BJE

[Chu07a] Matthew Church. Brian Jamison, ed., Scotland and the Cold War. Dunfermline, Scotland: Cualann Press, 2003. 190 pp. £12.99. Journal of Cold War Studies, 9(4):138–140, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/223532;   http://muse.jhu.edu/article/223532/
pdf;   https://direct.mit.edu/jcws/article-pdf/9/4/138/696719/
jcws.2007.9.4.138.pdf.

Church:2007:CGB

[Chu07b] Matthew Church. Cihat Göktepe, British Foreign Policy towards Turkey, 1959–1965. London: Frank Cass, 2003. 236 pp. $36.95. Journal of Cold War Studies, 9(2):181–183, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/214730;   http:
//muse.jhu.edu/article/214730/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/2/181/696467/jcws.2007.9.2.181.pdf.

Chun:2009:PGG

[Chu09] Clayton K. S. Chun. Paul G. Gillespie, The Development of Precision Guided Munitions. Journal of Cold War Studies, 11(4):160– 162, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/366254;
http://muse.jhu.edu/article/366254/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/4/160/1627631/jcws.2009.11.4.
160.pdf.

Chunikhin:2019:HAS

[Chu19] Kirill Chunikhin. At home among strangers: U.S. artists, the Soviet Union, and the myth of Rockwell Kent during the Cold War.

Journal of Cold War Studies, 21(4):175–207, Fall 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/21/4/175/699759/jcws_a_
00910.pdf;   https://muse.jhu.edu/article/736844.

Cienciala:2007:NDR

[Cie07] Anna M. Cienciala. Norman Davies, Rising ’44: The Battle for Warsaw. Journal of Cold War Studies, 9(3):161–163, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/219444;   http:
//muse.jhu.edu/article/219444/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/3/161/696652/jcws.2007.9.3.161.pdf.

Citino:2014:GDU

[Cit14] Nathan J. Citino. The ghosts of development: The United States and Jordan’s East ghor canal. Journal of Cold War Studies, 16(4):159–188, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566656;   http://muse.jhu.edu/article/566656/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/159/698623/
jcws_a_00519.pdf.

Cathcart:2011:BBN

[CK11] Adam Cathcart and Charles Kraus. The bonds of brotherhood: New evidence on sino–North Korean exchanges, 1950– 1954. Journal of Cold War Studies, 13(3):27–51, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/447468;   http:
//muse.jhu.edu/article/447468/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/3/27/700702/jcws_a_00141.pdf.

Creswell:2013:PPI

[CK13] Michael H. Creswell and Dieter H. Kollmer. Power, preferences, or ideas?: Explaining West Germany’s armaments strategy, 1955–1972. Journal of Cold War Studies, 15(4):55–103, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/534426;
http://muse.jhu.edu/article/534426/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/4/55/698311/jcws_a_00395.pdf.

Cox:2005:TAF

[CKP05a] Michael Cox and Caroline Kennedy-Pipe. The tragedies of American foreign policy: Further reflections. Journal of Cold War Studies, 7(1):

175–181, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/179557;
http://muse.jhu.edu/article/179557/pdf;   https://direct.
mit.edu/jcws/article-pdf/7/1/175/700458/1520397053326158.
pdf.

Cox:2005:TAD

[CKP05b] Michael Cox and Caroline Kennedy-Pipe. The tragedy of American diplomacy? Rethinking the Marshall Plan. Journal of Cold War Studies, 7(1):97–134, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/179556;   http://muse.jhu.edu/article/179556/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/1/97/700428/1520397053326202.
pdf.

Clark:2001:PBD

[Cla01] Ian Clark. The Politics of British Defence, 1979–98 (review). Journal of Cold War Studies, 3(1):128–130, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9168;   http://muse.jhu.edu/article/9168/pdf;
 https://direct.mit.edu/jcws/article-pdf/3/1/128/695398/
jcws.2001.3.1.128.pdf.

Clark:2009:THB

[Cla09] J. Ransom Clark. Tennent H. Bagley, Spy Wars: Moles, Mysteries, and Deadly Games. Journal of Cold War Studies, 11(2):137– 139, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/265875;
http://muse.jhu.edu/article/265875/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/2/137/697181/jcws.2009.11.2.
137.pdf.

Clemens:2000:CMF

[Cle00] Walter C. Clemens, Jr. Could more force have saved the Soviet system? Journal of Cold War Studies, 2(1):116–123, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9121;   http:/
/muse.jhu.edu/article/9121/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/1/116/695191/15203970051032390.pdf.

Clemens:2004:MCC

[Cle04] Walter C. Clemens, Jr. Mao’s China and the Cold War (review). Journal of Cold War Studies, 6(2):78–80, Spring 2004. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/53662;   http://muse.jhu.edu/
article/53662/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/2/78/695993/jcws.2004.6.2.78.pdf.

Clifford:2007:RDS

[Cli07] J. Garry Clifford. Robert D. Schulzinger, ed., A Companion to American Foreign Relations. Malden, MA: Blackwell Publishing, 2003. 562 pp. $131.95. Journal of Cold War Studies, 9(2):144–146, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214731;   http://muse.jhu.edu/article/214731/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/144/696425/
jcws.2007.9.2.144.pdf.

Clifford:2015:BED

[Cli15a] Mark L. Clifford. Barry Eichengreen, Dwight H. Perkins, and Kwanho Shin, eds., From Miracle to Maturity: The Growth of the Korean Economy. Cambridge, MA: Harvard University Press for the Harvard Asia Center, 2012. 366 pp. Journal of Cold War Studies, 17(4):217–219, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606047;   http://muse.jhu.edu/article/606047/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/217/698811/
jcws_r_00615.pdf.

Clifford:2015:PCH

[Cli15b] Mark L. Clifford. The Park Chung Hee Era: The Transformation of South Korea edited by Byung-Kook Kim and Ezra F. Vogel (review). Journal of Cold War Studies, 17(3):273–275, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595101;   http:
//muse.jhu.edu/article/595101/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/273/1102403/jcws_r_00521.pdf.

Clodfelter:2006:HKD

[Clo06] Mark Clodfelter. Hermann Knell, To Destroy a City: Strategic Bombing and Its Human Consequences in World War II. Journal of Cold War Studies, 8(2):156–159, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196272;   http://muse.jhu.edu/article/196272/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/156/700558/
jcws.2006.8.2.156.pdf.

Cloud:2009:AMC

[Clo09] Stanley Cloud. Anna M. Cienciala, Natalia S. Lebedeva, and Wojciech Materski, eds., Katyn: a Crime without Punishment. Journal of Cold War Studies, 11(2):122–124, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/265881;   http://muse.jhu.edu/article/265881/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/2/122/697145/
jcws.2009.11.2.122.pdf.

Clymer:2001:USF

[Cly01] Kenton J. Clymer. United States Foreign Policy Towards Cambodia, 1977–92: a Question of Realities (review). Journal of Cold War Studies, 3(1):146–148, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9169;   http://muse.jhu.edu/article/9169/pdf;
 https://direct.mit.edu/jcws/article-pdf/3/1/146/695426/
jcws.2001.3.1.146.pdf.

Clymer:2013:GOC

[Cly13] Kenton Clymer. The ground observer Corps: Public relations and the Cold War in the 1950s. Journal of Cold War Studies, 15(1): 34–52, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/505772;
http://muse.jhu.edu/article/505772/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/1/34/699964/jcws_a_00307.pdf.

Cobb:2018:AGP

[Cob18] Daniel M. Cobb. Alyosha Goldstein, Poverty in Common: The Politics of Community Action during the American Century. Journal of Cold War Studies, 20(4):248–250, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718215;   https://direct.mit.
edu/jcws/article-pdf/20/4/248/1969667/jcws_r_00835.pdf.

Cogan:2002:FRC

[Cog02a] Charles Cogan. France Restored: Cold War Diplomacy and the Quest for Leadership in Europe, 1944–1954 (review). Journal of Cold War Studies, 4(4):114–118, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9297;   http://muse.jhu.edu/article/9297/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/4/114/695710/
jcws.2002.4.4.114.pdf.

Cogan:2002:FDD

[Cog02b] Charles G. Cogan. The French Defense Debate: Consensus and Continuity in the Mitterrand Era (review). Journal of Cold War Studies, 4(1):115–117, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9222;   http://muse.jhu.edu/article/9222/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/1/115/695528/
jcws.2002.4.1.115.pdf.

Cogan:2003:RMC

[Cog03] Charles Cogan. Response to Michael Creswell and Marc Trachtenberg. Journal of Cold War Studies, 5(3):29–32, Summer 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45185;   http://muse.jhu.edu/
article/45185/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/29/695833/152039703322286755.pdf. See [CT03a].

Cogan:2012:ERP

[Cog12] Charles Cogan. Eric Remacle and Pascaline Winand, eds., America, Europe, Africa/L’Amérique, l’Europe, l’Afrique: 1945–1973. Brussels: P. I. E. Peter Lang, 2009. 328 pp. Journal of Cold War Studies, 14(4):237–239, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500342;   http://muse.jhu.edu/article/500342/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/237/698141/
jcws_r_00289.pdf.

Cohen:2002:VEW

[Coh02] Warren I. Cohen. Victory in Europe 1945: From World War to Cold War (review). Journal of Cold War Studies, 4(1):89–91, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9223;   http:/
/muse.jhu.edu/article/9223/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/1/89/695587/jcws.2002.4.1.89.pdf.

Cohen:2006:WCK

[Coh06] Warren I. Cohen. William C. Kirby, Robert S. Ross, and Gong Li, eds., Normalization of U.S.-China Relations: an International History. Journal of Cold War Studies, 8(4):149–150, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/204617;   http:
//muse.jhu.edu/article/204617/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/4/149/700805/jcws.2006.8.4.149.pdf.

Cohen:2008:JPW

[Coh08a] Warren I. Cohen. James Peck, Washington’s China: The National Security World, the Cold War, and the Origins of Globalism. Journal of Cold War Studies, 10(4):163–165, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/252666;   http://muse.jhu.edu/article/252666/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/4/163/697063/
jcws.2008.10.4.163.pdf.

Cohen:2008:YXN

[Coh08b] Warren I. Cohen. Yafeng Xia, Negotiating with the Enemy: U.S.- China Talks during the Cold War, 1949–1972. Journal of Cold War Studies, 10(3):176–178, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241519;   http://muse.jhu.edu/article/241519/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/176/697137/
jcws.2008.10.3.176.pdf.

Cohen:2010:MHW

[Coh10] Eliot A. Cohen. Michael Howard, War in European History, updated ed. Journal of Cold War Studies, 12(4):181–182, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/403800;   http:
//muse.jhu.edu/article/403800/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/4/181/697522/jcws_r_00052.pdf.

Cohen:2012:TJC

[Coh12] Warren I. Cohen. Thomas J. Christensen, Worse than a Monolith: Alliance Politics and Problems of Coercive Diplomacy in Asia. Princeton, NJ: Princeton University Press, 2011. 306 pp. Journal of Cold War Studies, 14(2):156–158, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/478781;   http://muse.jhu.edu/article/478781/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/2/156/697985/
jcws_r_00226.pdf.

Coleman:2000:EBP

[Col00] David G. Coleman. Eisenhower and the Berlin problem, 1953–1954. Journal of Cold War Studies, 2(1):3–34, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9122;   http://muse.jhu.edu/article/
9122/pdf;   https://direct.mit.edu/jcws/article-pdf/2/1/
3/695189/15203970051032363.pdf.

Coleman:2005:SBB

[Col05] David G. Coleman. Spies Beneath Berlin (review). Journal of Cold War Studies, 7(1):200–202, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/179554;   http://muse.jhu.edu/article/179554/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/1/200/700492/
jcws.2005.7.1.200.pdf.

Coleman:2007:MND

[Col07] David G. Coleman. The missiles of November, December, January, February, …: The problem of acceptable risk in the Cuban Missile Crisis settlement. Journal of Cold War Studies, 9(3):5–48, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/219445;   http:
//muse.jhu.edu/article/219445/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/3/5/696686/jcws.2007.9.3.5.pdf.

Coletta:2015:SKU

[Col15] Damon Coletta. Sean N. Kalic, U.S. Presidents and the Militarization of Space, 1946–1967. College Station, TX: Texas A&M University Press, 2012. 182 pp. $40.00. Journal of Cold War Studies, 17(4):184–186, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606030;   http://muse.jhu.edu/article/606030/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/184/698779/
jcws_r_00585.pdf.

Colman:2019:TWS

[Col19] Jonathan Colman. Toward “world support” and “the ultimate judgment of history”: The U.S. legal case for the blockade of Cuba during the Missile Crisis, October–November 1962. Journal of Cold War Studies, 21(2):150–173, Spring 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/726152;   https://direct.mit.
edu/jcws/article-pdf/21/2/150/699563/jcws_a_00879.pdf.

Connor:2003:SSP

[Con03] Walter D. Connor. Soviet society, public attitudes, and the perils of Gorbachev’s reforms: The social context of the end of the USSR. Journal of Cold War Studies, 5(4):43–80, Fall 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/47927;   http://muse.jhu.edu/
article/47927/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/4/43/700372/152039703322483756.pdf.

Connor:2007:DFS

[Con07] Walter D. Connor. Donald Filtzer, Soviet Workers and Late Stalinism: Labour and the Restoration of the Stalinist System after World War II. Journal of Cold War Studies, 9 (1):167–169, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210073;   http://muse.jhu.edu/article/210073/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/167/696606/
jcws.2007.9.1.167.pdf.

Connor:2012:REP

[Con12] Walter D. Connor. Review essay: Politics, discontents, hopes: 1968 East and West. Journal of Cold War Studies, 14(2):142–153, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/478779;   http:
//muse.jhu.edu/article/478779/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/2/142/697987/jcws_a_00224.pdf.

Connor:2023:BRT

[Con23] Roger Connor. Book review: The Turtle and the Dreamboat: The Cold War Flights That Forever Changed the Course of Global Aviation by Jim Leeke. Journal of Cold War Studies, 25(2):253–255, Spring 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/2/
253/2140948/jcws_r_01149.pdf.

Cook:2004:BSS

[Coo04] Linda J. Cook. Bloody Saturday in the Soviet Union: Novocherkassk, 1962 (review). Journal of Cold War Studies, 6(3):175–176, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171155;   http:
//muse.jhu.edu/article/171155/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/3/175/696105/jcws.2004.6.3.175.pdf.

Cormac:2014:PAW

[Cor14] Rory Cormac. The pinprick approach: Whitehall’s top-secret anti-Communist committee and the evolution of British covert action strategy. Journal of Cold War Studies, 16(3):5–28, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/556074;
http://muse.jhu.edu/article/556074/pdf;   https://direct.
mit.edu/jcws/article-pdf/16/3/5/698477/jcws_a_00469.pdf.

Cormac:2022:BBP

[Cor22] Rory Cormac. British “Black” productions: Forgeries, front groups, and propaganda, 1951–1977. Journal of Cold War Studies, 24(3): 4–42, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/24/3/4/2042017/jcws_a_01087.pdf.

Costa:2022:EWB

[Cos22] Ettore Costa. From East–West balancing to militant anti-communism: The Socialist International and the beginning of the Cold War, 1947–1949. Journal of Cold War Studies, 24(2):95–131, Spring 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/2/
95/2013026/jcws_a_01044.pdf.

Coufoudakis:2004:KPC

[Cou04] Van Coufoudakis. Keeping the Peace in the Cyprus Crisis of 1963– 64 (review). Journal of Cold War Studies, 6(2):92–94, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/53664;   http://muse.jhu.edu/
article/53664/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/2/92/696001/jcws.2004.6.2.92.pdf.

Cox:2002:BCW

[Cox02] Michael Cox. Britain and the Cold War, 1945–1991 (review). Journal of Cold War Studies, 4(2):110–113, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9242;   http://muse.jhu.edu/article/9242/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/2/110/695558/
jcws.2002.4.2.110.pdf.

Crawford:2003:HEI

[Cra03] Neta C. Crawford. The Helsinki Effect: International Norms, Human Rights, and the Demise of Communism (review). Journal of Cold War Studies, 5(2):109–111, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40758;   http://muse.jhu.edu/
article/40758/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/109/700418/jcws.2003.5.2.109.pdf.

Crawford:2007:WKE

[Cra07] David Crawford. Wolfgang Kraushaar, ed., Die RAF und der linke Terrorismus, 2 vols. Hamburg: Hamburg Edition HIS Verlages mbH, 2006, 1,415 pp. EUR 78.00. Journal of Cold War Studies, 9(4):160–164, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223535;   http://muse.jhu.edu/article/223535/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/160/696769/
jcws.2007.9.4.160.pdf.

Crandall:2013:DWS

[Cra13] Russell Crandall. Daniel Weimer, Seeing Drugs: Modernization, Counterinsurgency, and U.S. Narcotics Control in the Third World, 1969–1976. Kent, OH: Kent State University Press, 2011. 328 pp. Journal of Cold War Studies, 15(1):151–153, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/505766;   http:
//muse.jhu.edu/article/505766/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/1/151/699996/jcws_r_00313.pdf.

Craig:2017:KMT

[Cra17] R. Bruce Craig. Kati Marton, True Believer: Stalin’s Last American Spy. New York: Simon & Schuster, 2016. 289 pp. $27.00. Journal of Cold War Studies, 19(2):215–217, Spring 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http:

//www.mitpressjournals.org/doi/full/10.1162/JCWS_r_00724;
 https://direct.mit.edu/jcws/article-pdf/19/2/215/699220/
jcws_r_00724.pdf.

Crandall:2021:WSD

[Cra21] Russell Crandall. We Sell Drugs: The Alchemy of US Empire by Suzanna Reiss (review). Journal of Cold War Studies, 23(3):255–256, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/3/255/1955765/jcws_r_01017.pdf;
https://muse.jhu.edu/article/801919.

Critchlow:1999:WCW

[Cri99] James Critchlow. Western Cold War broadcasting. Journal of Cold War Studies, 1(3):168–175, Fall 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9112;   https://direct.mit.edu/jcws/
article-pdf/1/3/168/695160/152039799316976841.pdf.

Critchlow:2002:BFC

[Cri02] James Critchlow. Broadcasting Freedom: The Cold War Triumph of Radio Free Europe and Radio Liberty (review). Journal of Cold War Studies, 4(1):97–99, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9225;   http://muse.jhu.edu/article/
9225/pdf;   https://direct.mit.edu/jcws/article-pdf/4/1/
97/695579/jcws.2002.4.1.97.pdf.

Critchlow:2004:PDD

[Cri04] James Critchlow. Public diplomacy during the Cold War: The record and its implications. Journal of Cold War Studies, 6(1):75–89, Winter 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/51862;   http://
muse.jhu.edu/article/51862/pdf;   https://direct.mit.edu/
jcws/article-pdf/6/1/75/695894/152039704772741597.pdf.

Critchlow:2010:RHC

[Cri10] James Critchlow. Richard H. Cummings, Cold War Radio: The Dangerous History of American Broadcasting in Europe, 1950– 1989. Journal of Cold War Studies, 12(3):175–179, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/392803;   http:
//muse.jhu.edu/article/392803/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/3/175/697674/jcws_r_00031.pdf.

Critchlow:2020:RJR

[Cri20a] James Critchlow. A. Ross Johnson and R. Eugene Parta, eds., Cold War Broadcasting: Impact on the Soviet Union and Eastern Europe — A Collection of Studies and Documents. Journal of Cold War Studies, 22(2):178–180, Spring 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://muse.jhu.edu/

article/756819.

Critchlow:2020:CWB

[Cri20b] James Critchlow. Cold War broadcasting: Impact on the Soviet Union and Eastern Europe — a collection of studies and documents. Journal of Cold War Studies, 22(2):175–180, Spring 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/22/2/
175/1860562/jcws_r_00945.pdf.

Croan:2000:BBC

[Cro00] Melvin Croan. Battleground Berlin: CIA vs. KGB in the Cold War (review). Journal of Cold War Studies, 2(1):139–141, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9124;   http:/
/muse.jhu.edu/article/9124/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/1/139/695208/jcws.2000.2.1.139.pdf.

Cronin:2007:BVD

[Cro07] James E. Cronin. Bertjan Verbeek, Decision-Making in Great Britain during the Suez Crisis. Burlington, VT: Ashgate, 2003. 192 pp. Journal of Cold War Studies, 9(4):164–166, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/223536;   http://muse.jhu.edu/article/223536/
pdf;   https://direct.mit.edu/jcws/article-pdf/9/4/164/696749/
jcws.2007.9.4.164.pdf.

Crotty:2009:DFS

[Cro09] William Crotty. David F. Schmitz, The United States and Right-Wing Dictatorships. Journal of Cold War Studies, 11(4):157– 160, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/366255;
http://muse.jhu.edu/article/366255/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/4/157/1627627/jcws.2009.11.4.
157.pdf.

Cronin:2015:TFI

[Cro15] Audrey Kurth Cronin. Threats of Force: International Law and Strategy by Francis Grimal (review). Journal of Cold War Studies, 17(3):227–228, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595077;   http://muse.jhu.edu/article/595077/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/227/
1102474/jcws_r_00566.pdf.

Cradock:2007:NFV

[CS07] Christopher Cradock and M. L. R. Smith. “No fixed values”: a reinterpretation of the influence of the theory of guerre révolutionnaire and the Battle of Algiers, 1956–1957. Journal of Cold War Studies, 9(4):68–105, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/223534;

http://muse.jhu.edu/article/223534/pdf;   https://direct.
mit.edu/jcws/article-pdf/9/4/68/696775/jcws.2007.9.4.68.
pdf.

Creswell:2003:FGQ

[CT03a] Michael Creswell and Marc Trachtenberg. France and the German question, 1945–1955. Journal of Cold War Studies, 5(3):5–28, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/45187;   http://
muse.jhu.edu/article/45187/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/3/5/695820/152039703322286746.pdf. See

comment [She03] and reply [CT03b].

Creswell:2003:NLO

[CT03b] Michael Creswell and Marc Trachtenberg. New light on an old issue? Journal of Cold War Studies, 5(3):46–53, Summer 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45188;   http://muse.jhu.edu/
article/45188/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/46/695810/152039703322286764.pdf. See [CT03a].

Cull:2013:JCG

[Cul13] Nicholas J. Cull. Jessica C. E. Gienow-Hecht and Mark C. Donfried, eds., Searching for a Cultural Diplomacy. New York: Berghahn Books, 2010. 265 pp. $70.00/£40.00. Journal of Cold War Studies, 15(3):196–198, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522722;   http://muse.jhu.edu/article/522722/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/196/698305/
jcws_r_00349.pdf.

Cull:2022:BRC

[Cul22] Nicholas J. Cull. Book review: Cold War Frequencies: CIA Clandestine Radio Broadcasting to the Soviet Union and Eastern Europe by Richard H. Cummings. Journal of Cold War Studies, 24(4):231–233, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/4/
231/2062878/jcws_r_01113.pdf.

Curry:2007:AJP

[Cur07] Jane Leftwich Curry. Anita J. Prazmowska, Civil War in Poland, 1942–1948. Journal of Cold War Studies, 9(3):163–165, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/219447;   http:
//muse.jhu.edu/article/219447/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/3/163/696630/jcws.2007.9.3.163.pdf.

Curran:2015:BEL

[Cur15] James Curran. Beyond the euphoria: Lyndon Johnson in Australia and the politics of the Cold War Alliance. Journal of Cold War Studies, 17(1):64–96, Winter 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/581231;   http://muse.jhu.edu/article/581231/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/1/64/700774/
jcws_a_00531.pdf.

Curli:2023:AIE

[Cur23] Barbara Curli. Atoms for industry: The early nuclear activities of Fiat and the Atoms for Peace program in Italy, 1956–1959. Journal of Cold War Studies, 25(3):68–88, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/3/68/2158676/jcws_a_01159.pdf.

Daddis:2019:PWP

[Dad19] Gregory A. Daddis. Planning for a war in paradise: The 1966 Honolulu conference and the shape of the Vietnam War. Journal of Cold War Studies, 21(3):152–184, Summer 2019. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/731286;   https://direct.mit.
edu/jcws/article-pdf/21/3/152/699611/jcws_a_00897.pdf.

Dallin:2000:GDS

[Dal00] Alexander Dallin. Grand Delusion: Stalin and the German Invasion of Russia (review). Journal of Cold War Studies, 2(2):120– 122, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9137;   http:/
/muse.jhu.edu/article/9137/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/2/120/695259/jwcs.2000.2.2.120.pdf.

Dale:2010:APC

[Dal10] Gareth Dale. Andrew Port, Conflict and Stability in the German Democratic Republic. Journal of Cold War Studies, 12(3):147–149, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/392790;   http:

//muse.jhu.edu/article/392790/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/3/147/697581/jcws_r_00017.pdf.

Daniel:2007:DRF

[Dan07] D. C. F. Daniel. D. A. Romanov, Fire at Sea: The Tragedy of the Soviet Submarine Komsomolets. Journal of Cold War Studies, 9(3):178–180, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219448;   http://muse.jhu.edu/article/219448/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/178/696668/
jcws.2007.9.3.178.pdf.

Daniloff:2010:STF

[Dan10] Nicholas Daniloff. Seymour Topping, On the Front Lines of the Cold War: an American Correspondent’s Journal from the Chinese Civil War to the Cuban Missile Crisis and Vietnam. Journal of Cold War Studies, 12(3):165–168, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392799;   http://muse.jhu.edu/article/392799/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/165/697694/
jcws_r_00028.pdf.

Daniloff:2011:KJC

[Dan11] Nicholas Daniloff. Kempton Jenkins, Cold War Saga. Ann Arbor: Nimble Books, 2010. 425 pp. Journal of Cold War Studies, 13(4):281–282, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457228;   http://muse.jhu.edu/article/457228/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/281/697863/
jcws_r_00161.pdf.

Daniloff:2012:DHD

[Dan12a] Nicholas Daniloff. David E. Hoffman, The Dead Hand: The Untold Story of the Cold War Arms Race and Its Dangerous Legacy. New York: Doubleday, 2009. 577 pp. Journal of Cold War Studies, 14(1):129–132, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467694;   http://muse.jhu.edu/article/467694/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/129/697919/
jcws_r_00205.pdf.

Daniloff:2012:OBL

[Dan12b] Nicholas Daniloff. Oliver Bullough, Let Our Fame Be Great: Journeys among the Defiant People of the Caucasus. New York: Basic Books, 2010. 496 pp. Journal of Cold War Studies, 14(3):220–222, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/490193;   http:
//muse.jhu.edu/article/490193/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/3/220/698084/jcws_r_00266.pdf.

Daniloff:2014:HBD

[Dan14] Nicholas Daniloff. Henry S. Bradsher, The Dalai Lama’s Secret and Other Reporting Adventures. Baton Rouge: Louisiana State University Press, 2013. 312 pp. Journal of Cold War Studies, 16(1):218–220, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/543735;   http://muse.jhu.edu/article/543735/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/1/218/698943/
jcws_r_00413.pdf.

Daniloff:2015:RRF

[Dan15] Nicholas Daniloff. Reagan at Reykjavik: Forty-Eight Hours That Ended the Cold War by Kenneth Adelman (review). Journal of Cold War Studies, 17(2):145–148, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/582440;   http://muse.jhu.edu/article/582440/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/2/145/698694/
jcws_r_00538.pdf.

Daniel:2016:JGB

[Dan16a] Donald C. F. Daniel. Jeffrey G. Barlow, From Hot War to Cold: The U.S. Navy and National Security Affairs, 1945–1955. Stanford, CA: Stanford University Press, 2009. 710 pp. $65.00. Journal of Cold War Studies, 18(4):249–251, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645931;   http://muse.jhu.edu/article/645931/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/249/700168/
jcws_r_00704.pdf.

Daniloff:2016:DHB

[Dan16b] Nicholas Daniloff. David E. Hoffman, The Billion Dollar Spy: a True Story of Cold War Espionage and Betrayal. New York: Doubleday, 2015. 312 pp. Journal of Cold War Studies, 18(3):225–229, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/632301;   http:
//muse.jhu.edu/article/632301/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/3/225/699072/jcws_r_00668.pdf.

Daniloff:2016:SMT

[Dan16c] Nicholas Daniloff. Suzanne Massie, Trust but Verify: Reagan, Russia, and Me. Rockland, ME: Maine Authors Publishing, 2013. 380 pp. Journal of Cold War Studies, 18(4):225–228, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/645921;   http:
//muse.jhu.edu/article/645921/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/4/225/700089/jcws_r_00693.pdf.

Daniloff:2022:ARB

[Dan22] Nicholas Daniloff. Assignment Russia: Becoming a Foreign Correspondent in the Crucible of the Cold War by Marvin Kalb (review). Journal of Cold War Studies, 24(1):250–252, Winter 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/24/1/250/1980865/jcws_r_
01061.pdf;   https://muse.jhu.edu/article/843083.

Dapprich:2014:BPS

[Dap14] Matthias Dapprich. Between Prague Spring and French May: Opposition and Revolt in Europe, 1960–1980 edited by Martin Klimke, Jacco Pekelder, and Joachim Scharloth (review). Journal of Cold War Studies, 16(4):266–268, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566678;   http://muse.jhu.edu/article/566678/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/266/698639/
jcws_r_00497.pdf.

Darnton:2012:AAS

[Dar12] Christopher Darnton. Asymmetry and agenda-setting in U.S.– Latin American relations: Rethinking the origins of the Alliance for Progress. Journal of Cold War Studies, 14(4):55– 92, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/500329;
http://muse.jhu.edu/article/500329/pdf;   https://direct.
mit.edu/jcws/article-pdf/14/4/55/698112/jcws_a_00276.pdf.

Darnton:2015:MEH

[Dar15] Christopher Darnton. Marcia Esparza, Henry R. Huttenbach, and Daniel Feierstein, eds., State Violence and Genocide in Latin America: The Cold War Years. New York: Routledge, 2011. 303 pp. $49.95. Journal of Cold War Studies, 17(4):191–193, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/606033;   http:
//muse.jhu.edu/article/606033/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/4/191/698777/jcws_r_00589.pdf.

Darnton:2016:MEH

[Dar16] Christopher Darnton. Marcia Esparza, Henry R. Huttenbach, and Daniel Feierstein, eds., State Violence and Genocide in Latin America: The Cold War Years. New York: Routledge, 2010. 272 pp. $150.00. Journal of Cold War Studies, 18(2):199–201, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/620560;   http:
//muse.jhu.edu/article/620560/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/2/199/698898/jcws_r_00631.pdf.

Davies:2004:CPC

[Dav04] Richard T. Davies. The CIA and the Polish crisis of 1980– 1981. Journal of Cold War Studies, 6(3):120–123, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171156;   http:
//muse.jhu.edu/article/171156/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/3/120/696027/1520397041447346.pdf.

Davies:2007:PRG

[Dav07] R. W. Davies. Paul R. Gregory and Valery Lazarev, eds., The Economics of Forced Labor: The Soviet Gulag. Journal of Cold War Studies, 9(1):165–167, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210075;   http://muse.jhu.edu/article/210075/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/165/696587/
jcws.2007.9.1.165.pdf.

David:2009:DCA

[Dav09] James David. David Christopher Arnold, Spying from Space: Constructing America’s Satellite Command and Control Systems. Journal of Cold War Studies, 11(2):151–152, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/265868;   http://muse.jhu.edu/article/265868/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/2/151/
697201/jcws.2009.11.2.151.pdf.

Davy:2010:AWV

[Dav10] Richard Davy. Andreas Wenger, Vojtech Mastny, and Christian Nuenlist, eds., Origins of the European Security System: The Helsinki Process Revisited, 1965–75. Journal of Cold War Studies, 12(4):210–212, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/403815;   http://muse.jhu.edu/article/403815/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/4/210/697600/
jcws_r_00067.pdf.

Davy:2011:JWY

[Dav11a] Richard Davy. John W. Young, Twentieth-Century Diplomacy: a Case Study of British Practice 1963–1976. New York: Cambridge University Press, 2008. 244 pp. Journal of Cold War Studies, 13(3):234–236, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447484;   http://muse.jhu.edu/article/447484/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/234/700712/
jcws_r_00117.pdf.

Davy:2011:TFN

[Dav11b] Richard Davy. Thomas Fischer, Neutral Power in the CSCE: The N + N States and the Making of the Helsinki Accords 1975. Baden-Baden: Nomos for the Austrian Institute for International Affairs, 2009. 392 pp. Journal of Cold War Studies, 13(1):252–253, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/424261;   http://muse.jhu.edu/article/424261/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/1/252/697771/
jcws_r_00092.pdf.

Davy:2012:MPE

[Dav12] Richard Davy. Mario Del Pero, The Eccentric Realist: Henry Kissinger and the Shaping of American Foreign Policy. Ithaca, NY: Cornell University Press, 2010. 193 pp. Journal of Cold War Studies, 14(2):166–168, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/478786;   http://muse.jhu.edu/article/478786/pdf;

 https://direct.mit.edu/jcws/article-pdf/14/2/166/697999/
jcws_r_00231.pdf.

DeGroot:2020:SUC

[De 20] Michael De Groot. The Soviet Union, CMEA, and the energy crisis of the 1970s. Journal of Cold War Studies, 22(4):4–30, Fall 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/4/4/1860608/jcws_a_00964.pdf;   https://muse.jhu.edu/
article/776958.

Deery:2007:MBA

[Dee07] Phillip Deery. Malaya, 1948: Britain’s Asian Cold War? Journal of Cold War Studies, 9(1):29–54, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210076;   http://muse.jhu.edu/article/210076/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/29/696582/jcws.
2007.9.1.29.pdf.

Deighton:2002:RFE

[Dei02] Anne Deighton. The Rise and Fall of the European Defence Community: Anglo–American Relations and the Crisis of European Defence, 1950–55, and: a Special Relationship: Anglo–American Relations in the Cold War and After (review). Journal of Cold War Studies, 4(4): 118–120, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9298;   http:/
/muse.jhu.edu/article/9298/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/4/118/695694/jcws.2002.4.4.118.pdf.

Deighton:2007:PBE

[Dei07] Anne Deighton. Peter Boyle, The Eden–Eisenhower Correspondence, 1955–1957. Chapel Hill: University of North Carolina Press, 2005. 230 pp. Journal of Cold War Studies, 9(4):127–128, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223537;   http://muse.jhu.edu/article/223537/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/127/696741/
jcws.2007.9.4.127.pdf.

Deighton:2013:MJT

[Dei13] Anne Deighton. Michael J. Turner, British Power and International Relations during the 1950s: a Tenable Position? Lanham, MD: Lexington Books, 2009. 353 pp. Journal of Cold War Studies, 15(1):169–170, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/505757;   http://muse.jhu.edu/article/505757/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/1/169/700006/
jcws_r_00322.pdf.

Deighton:2014:RWB

[Dei14] Anne Deighton. Richard Wevill, Britain and America after World War II: Bilateral Relations and the Beginnings of the Cold War. New York: I. B. Tauris, 2012. 343 pp. Journal of Cold War Studies, 16(1):226–227, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/543739;   http://muse.jhu.edu/article/543739/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/1/226/698958/
jcws_r_00417.pdf.

Deighton:2015:DMW

[Dei15] Anne Deighton. David M. Watry, Diplomacy at the Brink: Eisenhower, Churchill and Eden in the Cold War. Baton Rouge: Louisiana State University Press, 2014. 240 pp. $29.95. Journal of Cold War Studies, 17(4):212–214, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606044;   http://muse.jhu.edu/article/606044/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/212/698803/
jcws_r_00612.pdf.

Deletant:2007:JAP

[Del07] Dennis Deletant. Jean Ancel, Preludiu la asasinat: Pogromul de la Iaşi, 29 iunie 1941. Journal of Cold War Studies, 9(3):205–207, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219449;   http://muse.jhu.edu/article/219449/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/205/696666/
jcws.2007.9.3.205.pdf.

Delury:2022:FCB

[Del22] John Delury. Feudal contradictions between Communist allies: Deng Xiaoping, Kim Il-Sung, and the problem of succession, 1976–1984. Journal of Cold War Studies, 24(2):4–28, Spring 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/2/
4/2013032/jcws_a_01057.pdf.

Delbos:2023:BRF

[Del23] Stephan Delbos. Book review: The Form of Ideology and the Ideology of Form: Cold War, Decolonization and Third World Print Cultures by Francesca Orsini, Neelam Srivastava, and Laetitia Zecchini, eds. Journal of Cold War Studies, 25(1):210–212, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/1/
210/2073763/jcws_r_01127.pdf.

DeMare:2012:AGW

[DeM12] Brian J. DeMare. Andrew G. Walder, Fractured Rebellion: The Beijing Red Guard Movement. Cambridge, MA: Harvard University Press, 2009. 416 pp. $39.95. Journal of Cold War Studies, 14(3):230–231, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/490197;   http://muse.jhu.edu/article/490197/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/3/230/698077/
jcws_r_00270.pdf.

Dennis:2006:RBN

[Den06] Peter Dennis. Robert L. Bateman, No Gun Ri: a Military History of the Korean War Incident. Journal of Cold War Studies, 8(2):176–177, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196274;   http://muse.jhu.edu/article/196274/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/176/700554/
jcws.2006.8.2.176.pdf.

Denza:2021:CMD

[Den21] Eileen Denza. A Cornerstone of Modern Diplomacy: Britain and the Negotiation of the 1961 Vienna Convention on Diplomatic Relations by Kai Bruns (review). Journal of Cold War Studies, 23(3):235–237, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/3/235/1955797/jcws_r_01031.pdf;
https://muse.jhu.edu/article/801911.

Derlugian:2002:DSU

[Der02] Georgi Derlugian. The Destruction of the Soviet Union: a Study in Globalization (review). Journal of Cold War Studies, 4(3):157–160, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9271;   http:/
/muse.jhu.edu/article/9271/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/157/700260/jcws.2002.4.3.157.pdf.

DeRoche:2017:NMJ

[DeR17] Andy DeRoche. Nancy Mitchell, Jimmy Carter in Africa: Race and the Cold War. Stanford, CA: Stanford University Press, 2016. 883 pp. $45.00. Journal of Cold War Studies, 19(3):258–260, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/668179;   http:
//muse.jhu.edu/article/668179/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/3/258/699336/jcws_r_00746.pdf.

David-Fox:2007:MLC

[DF07] Michael David-Fox. Matthew Lenoe, Closer to the Masses: Stalinist Culture, Social Revolution, and Soviet Newspapers. Journal of Cold War Studies, 9(1):162–164, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/211266;   http://muse.jhu.edu/article/211266/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/162/696573/
jcws.2007.9.1.162.pdf.

David-Fox:2012:AGA

[DF12] Michael David-Fox. Anne E. Gorsuch, All This is Your World: Soviet Tourism at Home and Abroad after Stalin. Oxford: Oxford University Press, 2011. 222 pp. $110.00. Journal of Cold War Studies, 14(4):257–259, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500351;   http://muse.jhu.edu/article/500351/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/257/698179/
jcws_r_00298.pdf.

Drozdova:2019:LAE

[DF19] Katya Drozdova and Joseph H. Felter. Leaving Afghanistan: Enduring lessons from the Soviet Politburo. Journal of Cold War Studies, 21(4):31–70, Fall 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.

mit.edu/jcws/article-pdf/21/4/31/699696/jcws_a_00906.pdf;
 https://muse.jhu.edu/article/736848.

Donaldson:2020:PSU

[DFK^{+}20] Robert H. Donaldson, Jeremy Friedman, Edward A. Kolodziej, Margot Light, Robert G. Patman, Sergey Radchenko, and Radoslav A.

Yordanov. Perspectives on The Soviet Union and the Horn of Africa during the Cold War. Journal of Cold War Studies, 22(1):210–242, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/1/210/699798/jcws_c_00931.pdf;   https://muse.jhu.edu/
article/749653.

Diehl:2001:MMA

[Die01] Paul F. Diehl. Mixed Messages: American Politics and International Organization, 1919–1999 (review). Journal of Cold War Studies, 3(3): 96–97, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9199;   http:/
/muse.jhu.edu/article/9199/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/96/695470/jcws.2001.3.3.96.pdf.

Diehl:2007:MBM

[Die07] Paul F. Diehl. Michael Barnett and Martha Finnemore, Rules for the World: International Organizations in Global Politics. Ithaca: Cornell University Press, 2004. 226 pp. Journal of Cold War Studies, 9(4):129–130, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223538;   http://muse.jhu.edu/article/223538/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/129/696723/
jcws.2007.9.4.129.pdf.

Diedrich:2008:JWV

[Die08] Torsten Diedrich. Julia Warth, Verräter oder Widerstandskämpfer? Wehrmachtgeneral Walther von Seydlitz-Kurzbach [Traitor or Resistance Fighter? Wehrmacht General Walther von Seydlitz-Kurzbach]. Journal of Cold War Studies, 10(3):169–172, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/241516;
http://muse.jhu.edu/article/241516/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/3/169/697111/jcws.2008.10.3.
169.pdf.

Dimitrov:2008:CLC

[Dim08] Martin Dimitrov. Claude Lefort, Complications: Communism and the Dilemmas of Democracy. Journal of Cold War Studies, 10(4):172–173, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/252671;   http://muse.jhu.edu/article/252671/pdf;

 https://direct.mit.edu/jcws/article-pdf/10/4/172/697069/
jcws.2008.10.4.172.pdf.

Dine:2002:VAD

[Din02] Thomas A. Dine. The Voice of America and the Domestic Propaganda Battles, 1945–1953 (review). Journal of Cold War Studies, 4(3):125– 127, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9272;   http:/
/muse.jhu.edu/article/9272/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/125/700232/jcws.2002.4.3.125.pdf.

Dine:2004:PPA

[Din04] Thomas A. Dine. Peace Process: American Diplomacy and the Arab– Israeli Conflict since 1967 (review). Journal of Cold War Studies, 6(2):89–92, Spring 2004. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
53665;   http://muse.jhu.edu/article/53665/pdf;   https:/
/direct.mit.edu/jcws/article-pdf/6/2/89/696007/jcws.2004.
6.2.89.pdf.

Dine:2007:MSE

[Din07] Thomas A. Dine. Michael E. Staub, ed., The Jewish 1960s: an American Sourcebook. Waltham, MA: Brandeis University Press, 2004. 371 pp. Journal of Cold War Studies, 9(4):132–138, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/223539;   http:
//muse.jhu.edu/article/223539/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/4/132/696743/jcws.2007.9.4.132.pdf.

Dine:2010:UPP

[Din10] Thomas A. Dine. U.S. policy and peacemaking efforts in the Middle East: Historical perspectives. Journal of Cold War Studies, 12(2):117–125, Spring 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/382207;   http://muse.jhu.edu/article/382207/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/2/117/697491/
jcws.2010.12.2.117.pdf.

Dine:2013:HBJ

[Din13] Thomas A. Dine. Harold Brown with Joyce Winslow, Star Spangled Security: Applying Lessons Learned over Six Decades Safeguarding America. Washington, DC: Brookings Institution Press, 2012. 249 pp. Journal of Cold War Studies, 15(3):188–189, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/522717;   http:
//muse.jhu.edu/article/522717/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/3/188/698321/jcws_r_00378.pdf.

Dine:2014:CDL

[Din14a] Thomas A. Dine. Craig Daigle, The Limits of Détente: The United States, the Soviet Union, and the Arab–Israeli Conflict, 1969–1973. New Haven: Yale University Press, 2012. 423 pp. Journal of Cold War Studies, 16(1):227–231, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/543740;   http://muse.jhu.edu/article/543740/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/1/227/698976/
jcws_r_00418.pdf.

Dinnerstein:2014:SBA

[Din14b] Leonard Dinnerstein. Susanne Bardgett et al., eds., Survivors of Nazi Persecution in Europe after the Second World War, Vol. 1, Landscapes after Battle. London: Valentine Mitchell, 2010. xv + 236 pp. Journal of Cold War Studies, 16(1):222–224, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543737;   http:
//muse.jhu.edu/article/543737/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/1/222/698960/jcws_r_00415.pdf.

Dine:2018:SMI

[Din18] Thomas A. Dine. Shaul Mitelpunkt, Israel in the American Mind: The Cultural Politics of US–Israeli Relations, 1958–1988. Journal of Cold War Studies, 20(4):254–256, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718218;   https://direct.mit.
edu/jcws/article-pdf/20/4/254/1969610/jcws_r_00844.pdf.

Dine:2022:LDU

[Din22] Thomas A. Dine. The limits of détente: The United States, the Soviet Union, and the Arab–Israeli conflict, 1969–1973. Journal of Cold War Studies, 24(3):255–258, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/24/3/255/2042011/jcws_r_01101.pdf.

Dine:2023:BRC

[Din23a] Thomas A. Dine. Book review: Cold War Radio: The Russian Broadcasts of the Voice of America and Radio Free Europe/Radio Liberty by Mark G. Pomar. Journal of Cold War Studies, 25(3):215–218, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/3/
215/2158660/jcws_r_01152.pdf.

Dine:2023:BRI

[Din23b] Thomas A. Dine. Book review: Israel’s Moment: International Support for and Opposition to Establishing the Jewish State, 1945–1949 by Jeffrey Herf. Journal of Cold War Studies, 25(1):227–230, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/1/
227/2073761/jcws_r_01134.pdf.

Dittmer:2009:LCT

[Dit09] Lowell Dittmer. Lin Chun, The Transformation of Chinese Socialism. Journal of Cold War Studies, 11(2):141–142, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/265873;   http:
//muse.jhu.edu/article/265873/pdf;   https://direct.mit.
edu/jcws/article-pdf/11/2/141/697173/jcws.2009.11.2.141.
pdf.

Divine:2000:CWL

[Div00] Robert A. Divine. Choosing War: The Lost Chance for Peace and the Escalation of War in Vietnam (review). Journal of Cold War Studies, 2(2):109–111, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9138;   http://muse.jhu.edu/article/9138/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/2/109/695263/
jwcs.2000.2.2.109.pdf.

Divine:2001:LLR

[Div01] Robert A. Divine. LBJ: a Life (review). Journal of Cold War Studies, 3(3):106–107, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9200;   http://muse.jhu.edu/article/9200/pdf;
 https://direct.mit.edu/jcws/article-pdf/3/3/106/695452/
jcws.2001.3.3.106.pdf.

Dizard:2010:YRP

[Diz10] Wilson Dizard, Jr. Yale Richmond, Practicing Public Diplomacy: a Cold War Odyssey. Journal of Cold War Studies, 12(3):173–174, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/392802;   http:
//muse.jhu.edu/article/392802/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/3/173/697653/jcws_r_00030.pdf.

deJong-Lambert:2013:HJM

[dL13] William deJong Lambert. Hermann J. Muller, Theodosius Dobzhansky, Leslie Clarence Dunn, and the reaction to Lysenkoism in the United States. Journal of Cold War Studies, 15(1):78–118, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/505770;
http://muse.jhu.edu/article/505770/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/1/78/699960/jcws_a_00309.pdf.

Dobson:2010:IEC

[Dob10a] Alan P. Dobson. From instrumental to expressive: The changing goals of the U.S. Cold War strategic embargo. Journal of Cold War Studies, 12(1):98–119, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/382488;   http://muse.jhu.edu/article/382488/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/1/98/1627469/
jcws.2010.12.1.98.pdf.

Dobson:2010:TUB

[Dob10b] Hugo Dobson. Thomas U. Berger, Mike M. Mochizuki, and Jistuo Tsuchiyama, eds., Japan in International Politics: The Foreign Policies of an Adaptive State. Journal of Cold War Studies, 12(1): 166–167, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382494;
http://muse.jhu.edu/article/382494/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/1/166/1627417/jcws.2010.12.1.
166.pdf.

Dockrill:2004:EAP

[Doc04] Saki Ruth Dockrill. Eisenhower’s Atoms for Peace (review). Journal of Cold War Studies, 6(3):134–136, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/171157;   http://muse.jhu.edu/article/171157/
pdf;   https://direct.mit.edu/jcws/article-pdf/6/3/134/696033/
jcws.2004.6.3.134.pdf.

Dockrill:2005:KMC

[Doc05] Saki Ruth Dockrill. Kennedy, Macmillan and the Cold War: The Irony of Interdependence (review). Journal of Cold War Studies, 7(4):160–162, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188256;   http://muse.jhu.edu/article/188256/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/160/696303/
jcws.2005.7.4.160.pdf.

Dockrill:2008:HPB

[Doc08] Saki Ruth Dockrill. Helen Parr, Britain’s Policy towards the European Community: Harold Wilson and Britain’s World Role, 1964–1967. Journal of Cold War Studies, 10(3):139–140, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/241503;
http://muse.jhu.edu/article/241503/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/3/139/697097/jcws.2008.10.3.
139.pdf.

Dominguez:2003:CMH

[Dom03] Jorge I. Domínguez. Conflicting Missions: Havana, Washington, and Africa, 1959–1976 (review). Journal of Cold War Studies, 5(3):135– 137, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/45189;
 http://muse.jhu.edu/article/45189/pdf;   https://direct.
mit.edu/jcws/article-pdf/5/3/135/695905/jcws.2003.5.3.135.
pdf.

Domber:2005:RFB

[Dom05] Gregory F. Domber. The Rise and Fall of the Brezhnev Doctrine in Soviet Foreign Policy (review). Journal of Cold War Studies, 7(3):186–188, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185099;   http://muse.jhu.edu/article/185099/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/186/1102581/
jcws.2005.7.3.186.pdf.

Dominguez:2007:AMK

[Dom07] Jorge I. Domínguez. Arie M. Kacowicz, The Impact of Norms in International Society: The Latin American Experience, 1881– 2001. South Bend, IN: University of Notre Dame Press, 2005. 228 pp. Journal of Cold War Studies, 9(2):173–175, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/214733;   http:
//muse.jhu.edu/article/214733/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/2/173/696449/jcws.2007.9.2.173.pdf.

Dominguez:2008:GGL

[Dom08] Jorge I. Domínguez. Greg Grandin, The Last Colonial Massacre: Latin America in the Cold War. Journal of Cold War Studies, 10(3):147–149, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241507;   http://muse.jhu.edu/article/241507/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/147/697103/
jcws.2008.10.3.147.pdf.

Dominguez:2009:MGU

[Dom09] Jorge I. Domínguez. Michael Grow, U.S. Presidents and Latin American Interventions: Pursuing Regime Change in the Cold War. Journal of Cold War Studies, 11(2):145–147, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/265871;   http://muse.jhu.edu/article/265871/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/2/145/
697177/jcws.2009.11.2.145.pdf.

Domber:2011:SSR

[Dom11] Gregory F. Domber. Skepticism and stability: Reevaluating U.S. policy during Poland’s democratic transformation in 1989. Journal of Cold War Studies, 13(3):52–82, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/447464;   http://muse.jhu.edu/article/447464/
pdf;   https://direct.mit.edu/jcws/article-pdf/13/3/52/700686/
jcws_a_00142.pdf.

Domenico:2015:KMU

[Dom15a] Roy Domenico. Kaeten Mistry, The United States, Italy and the Origins of the Cold War: Waging Political Warfare, 1945–1950. New York: Cambridge University Press, 2014. 296 pp. $99.00. Journal of Cold War Studies, 17(4):178–179, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606027;   http://muse.jhu.edu/article/606027/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/178/698728/
jcws_r_00582.pdf.

Dominguez:2015:BES

[Dom15b] Jorge I. Domínguez. Beyond the Eagle’s Shadow: New Histories of Latin America’s Cold War edited by Virginia Garrard-Burnett, Mark Attwood Lawrence, and Julio E. Moreno (review). Journal of Cold War Studies, 17(2):148–150, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/582441;   http://muse.jhu.edu/article/582441/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/2/148/698696/
jcws_r_00539.pdf.

Dominguez:2015:CAS

[Dom15c] Jorge I. Domínguez. Cubans in Angola: South–South Cooperation and Transfer of Knowledge, 1976–1991 by Christine Hatzky (review). Journal of Cold War Studies, 17(3):234–236, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595081;   http:
//muse.jhu.edu/article/595081/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/234/1102480/jcws_r_00570.pdf.

Domenico:2020:CAC

[Dom20] Roy Domenico. Confronting America: The Cold War between the United States and the Communists in France and Italy by Alessandro Brogi (review). Journal of Cold War Studies, 22(1):257–259, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/1/257/699821/jcws_r_00915.pdf;   https://muse.jhu.edu/
article/749642.

Donaghy:2005:WPC

[Don05] Greg Donaghy. A War of Patrols: Canadian Army Operations in Korea (review). Journal of Cold War Studies, 7(4):150–152, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/188257;   http:
//muse.jhu.edu/article/188257/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/4/150/696293/jcws.2005.7.4.150.pdf.

Douglas:2006:CGU

[Dou06] Lawrence Douglas. A Century of Genocide: Utopias of Race and Nation (review). Journal of Cold War Studies, 8(1):143–144, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/192206;   http:
//muse.jhu.edu/article/192206/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/1/143/696337/jcws.2006.8.1.143.pdf.

Douglas:2015:HSG

[Dou15] R. M. Douglas. Hugo Service, Germans to Poles: Communism, Nationalism, and Ethnic Cleansing after the Second World War. New York: Cambridge University Press, 2013. ix + 378 pp. $99.00. Journal of Cold War Studies, 17(4):199–201, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/606037;   http://muse.jhu.edu/article/606037/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/4/199/
698796/jcws_r_00605.pdf.

Drake:2004:ICS

[Dra04a] Richard Regis Drake. Italian Communism and Soviet terror. Journal of Cold War Studies, 6(2):57–63, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/53666;   http://muse.jhu.edu/
article/53666/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/2/57/695963/152039704773254768.pdf.

Drake:2004:SDI

[Dra04b] Richard Regis Drake. The Soviet dimension of Italian Communism. Journal of Cold War Studies, 6(3):115–119, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171158;   http:
//muse.jhu.edu/article/171158/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/3/115/696039/1520397041447355.pdf.

Drake:2006:AMM

[Dra06] Richard Regis Drake. The Aldo Moro murder case in retrospect. Journal of Cold War Studies, 8(2):114–125, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/196275;   http:
//muse.jhu.edu/article/196275/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/2/114/700520/jcws.2006.8.2.114.pdf.

Drake:2008:SSC

[Dra08] Richard Drake. Salvatore Sechi, Compagno e cittadino: Il PCI tra via parlamentare e lotta armata [Comrade and Citizen: The PCI between Parliament and Armed Struggle]. Journal of Cold War Studies, 10(3):172–174, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241517;   http://muse.jhu.edu/article/241517/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/172/697117/
jcws.2008.10.3.172.pdf.

Drake:2009:DGH

[Dra09] Paul W. Drake. Darren G. Hawkins, International Human Rights and Authoritarian Rule in Chile. Journal of Cold War Studies, 11(2):147–149, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/265870;   http://muse.jhu.edu/article/265870/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/2/147/697191/
jcws.2009.11.2.147.pdf.

Drake:2010:TDI

[Dra10] Richard Drake. Terrorism and the decline of Italian Communism: Domestic and international dimensions. Journal of Cold War Studies, 12(2):110–116, Spring 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/382206;   http://muse.jhu.edu/article/382206/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/2/110/697499/
jcws.2010.12.2.110.pdf.

Dragomir:2012:FSB

[Dra12a] Elena Dragomir. The formation of the Soviet Bloc’s Council for Mutual Economic Assistance: Romania’s involvement. Journal of Cold War Studies, 14(1):34–47, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467677;   http://muse.jhu.edu/article/467677/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/34/697829/
jcws_a_00190.pdf.

Drake:2012:EAR

[Dra12b] Richard Drake. Elena Aga-Rossi and Victor Zaslavsky, Stalin and Togliatti: Italy and the Origins of the Cold War. Washington, DC: Woodrow Wilson Center Press, 2011. xvi + 340 pp. $60.00. Journal of Cold War Studies, 14(3):203–206, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/490185;   http://muse.jhu.edu/article/490185/
pdf;   https://direct.mit.edu/jcws/article-pdf/14/3/203/
698055/jcws_r_00258.pdf.

Drake:2018:FLV

[Dra18] Richard Drake. Fabio Lavagno, Vladimiro Satta, Moro: L’inchiesta senza finale. Journal of Cold War Studies, 20(4):251–254, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718217;   https://direct.mit.
edu/jcws/article-pdf/20/4/251/1969657/jcws_r_00843.pdf.

Drake:2021:RSI

[Dra21] Richard Drake. Rendiconto: La sinistra italiana dal Pci a oggi by Claudio Petruccioli (review). Journal of Cold War Studies, 23(3):248–250, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/3/248/1955777/jcws_r_01036.pdf;
https://muse.jhu.edu/article/801916.

Draenos:2022:GGE

[Dra22] Stan Draenos. Gifted Greek: The enigma of Andreas Papandreou. Journal of Cold War Studies, 24(2):173–176, Spring 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/24/2/173/2013012/jcws_r_
01065.pdf.

Dreyer:2005:CWP

[Dre05] June Teufel Dreyer. Chinese Warfighting: The PLA Experience since 1949 (review). Journal of Cold War Studies, 7(4):172–174, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188258;   http://muse.jhu.edu/article/188258/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/172/696339/
jcws.2005.7.4.172.pdf.

Dreyer:2013:TMC

[Dre13] June Teufel Dreyer. Tai Ming Cheung, ed., China’s Emergence as a Defense Technological Power. New York: Routledge Press, 2013. viii + 221 pp. Journal of Cold War Studies, 15(3):206–207, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/522727;   http:
//muse.jhu.edu/article/522727/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/3/206/698333/jcws_r_00382.pdf.

Drezner:2015:EIW

[Dre15] Daniel W. Drezner. Economic Interdependence and War by Dale Copeland (review). Journal of Cold War Studies, 17(3):228–231, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595078;   http:
//muse.jhu.edu/article/595078/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/228/1102444/jcws_r_00567.pdf.

Dreyer:2021:HCC

[Dre21] June Teufel Dreyer. Haunted by Chaos: China’s Grand Strategy from Mao Zedong to Xi Jinping by Sulmaan Wasif Khan (review). Journal of Cold War Studies, 23(2):164–166, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/23/2/164/1919923/jcws_r_
01013.pdf;   https://muse.jhu.edu/article/794719.

Dror:2018:EPW

[Dro18] Olga Dror. Education and politics in wartime: School systems in North and south Vietnam, 1965–1975. Journal of Cold War Studies, 20(3):57–113, Summer 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http:

//muse.jhu.edu/article/704571;   https://direct.mit.edu/
jcws/article-pdf/20/3/57/699642/jcws_a_00819.pdf.

Dimitrov:2014:SSI

[DS14] Martin K. Dimitrov and Joseph Sassoon. State Security, Information, and Repression: a Comparison of Communist Bulgaria and Ba’thist Iraq. Journal of Cold War Studies, 16(2):3– 31, Spring 2014. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/547322;
http://muse.jhu.edu/article/547322/pdf;   https://direct.
mit.edu/jcws/article-pdf/16/2/3/698409/jcws_a_00448.pdf.

deSa:2019:WWT

[dS19] Tiago Moreira de Sá. The world was not turning in their direction: The United States and the decolonization of Angola. Journal of Cold War Studies, 21(1):52–65, Winter 2019. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/723552;   https://direct.
mit.edu/jcws/article-pdf/21/1/52/699757/jcws_a_00871.pdf.

Duke:2002:AOG

[Duk02] Simon Duke. America’s Overseas Garrisons: The Leasehold Empire (review). Journal of Cold War Studies, 4(2):146–147, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9247;   http:/
/muse.jhu.edu/article/9247/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/2/146/695611/jcws.2002.4.2.146.pdf.

Duke:2010:WLE

[Duk10] Simon Duke. Wilfried Loth, ed., Experiencing Europe: 50 Years of European Construction 1957–2007. Journal of Cold War Studies, 12(3):151–153, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392792;   http://muse.jhu.edu/article/392792/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/151/697612/
jcws_r_00019.pdf.

Duke:2011:NMC

[Duk11] Simon Duke. Nick McCamley. Cold War Secret Nuclear Bunkers: The Passive Defence of the Western World during the Cold War. Barnsley, South Yorkshire: Pen and Sword Military Classics, 2007. 281 pp. $21.99. Journal of Cold War Studies, 13(3):216–218, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/447474;   http:
//muse.jhu.edu/article/447474/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/3/216/700718/jcws_r_00154.pdf.

Duke:2013:MHS

[Duk13] Simon Duke. Maria Höhn and Seungsook Moon, eds., Over There: Living with the U.S. Military Empire from World War Two to the Present. Durham, NC: Duke University Press, 2010. 303 pp. Journal of Cold War Studies, 15(2):138–140, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/509708;   http://muse.jhu.edu/article/509708/
pdf; http://muse.jhu.edu/article/509709;   http://muse.jhu.
edu/article/509709/pdf;   https://direct.mit.edu/jcws/article-
pdf/15/2/138/698165/jcws_r_00342.pdf.

Dumancic:2014:SOG

[Dum14] Marko Dumančić. Spectrums of oppression: Gender and sexuality during the Cold War. Journal of Cold War Studies, 16(3):190–204, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/556080;   http:
//muse.jhu.edu/article/556080/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/3/190/698481/jcws_a_00473.pdf.

Dumancic:2017:KST

[Dum17] Marko Dumančić. Kathleen Starck, Of Treason, God, and Testicles: Political Masculinities in British and American Films of the Early Cold War. Journal of Cold War Studies, 19(4):275–277, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/683692;   http:
//muse.jhu.edu/article/683692/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/4/275/699461/jcws_r_00792.pdf.

Dumbrell:2020:FPL

[Dum20] John Dumbrell. The Foreign Policy of Lyndon B. Johnson: The United States and the World, 1963–1969 by Jonathan Colman (review). Journal of Cold War Studies, 22(1):271–272, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/1/271/699864/jcws_r_00920.pdf;   https://muse.jhu.edu/
article/749635.

Dunlop:2003:ACI

[Dun03] John B. Dunlop. The August 1991 coup and its impact on Soviet politics. Journal of Cold War Studies, 5(1):94–127, Winter 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/38250;   http://muse.jhu.edu/
article/38250/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/1/94/695771/152039703320996731.pdf.

Diamant:2018:TAR

[DX18] Neil J. Diamant and Feng Xiaocai. Textual anxiety: Reading (and misreading) the draft constitution in China, 1954. Journal of Cold War Studies, 20(3):153–179, Summer 2018. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/704573;   https://direct.mit.
edu/jcws/article-pdf/20/3/153/699618/jcws_a_00769.pdf.

Easter:2010:BRS

[Eas10] David Easter. Bradley R. Simpson, Economists with Guns: Authoritarian Development and U.S.–Indonesian Relations, 1960– 1968. Journal of Cold War Studies, 12(3):157–159, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/392795;   http:
//muse.jhu.edu/article/392795/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/3/157/697667/jcws_r_00022.pdf.

Eckstein:2007:PHE

[Eck07a] Arthur M. Eckstein. Paul Hollander, The End of Commitment: Intellectuals, Revolutionaries, and Political Morality. Journal of Cold War Studies, 9(3):187–191, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219450;   http://muse.jhu.edu/article/219450/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/187/696676/
jcws.2007.9.3.187.pdf.

Eckstein:2007:REC

[Eck07b] Arthur M. Eckstein. Review essay: Clandestine agents: The real Agnes Smedley. Journal of Cold War Studies, 9(4):106–114, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/223540;   http:
//muse.jhu.edu/article/223540/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/4/106/696735/jcws.2007.9.4.106.pdf.

Eckert:2014:LFC

[Eck14] Carter J. Eckert. Literature and Film in Cold War South Korea: Freedom’s Frontier by Theodore Hughes (review). Journal of Cold War Studies, 16(4):241–243, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566666;   http://muse.jhu.edu/article/566666/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/241/698629/
jcws_r_00488.pdf.

Eckstein:2019:CCG

[Eck19] Arthur Eckstein. Crash Course: From the Good War to the Forever War by H. Bruce Franklin (review). Journal of Cold War Studies, 21 (1):189–193, Winter 2019. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
723538;   https://direct.mit.edu/jcws/article-pdf/21/1/189/
699748/jcws_r_00859.pdf.

Edgar:2008:DNV

[Edg08] Adrienne Edgar. Douglas Northrop, Veiled Empire: Gender and Power in Stalinist Central Asia. Journal of Cold War Studies, 10(3):207–209, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241535;   http://muse.jhu.edu/article/241535/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/207/697147/
jcws.2008.10.3.207.pdf.

Edwards:2016:CBT

[Edw16] Peter Edwards. Christopher Bayly and Tim Harper, Forgotten Wars: Freedom and Revolution in Southeast Asia. Cambridge, MA: The Belknap Press of Harvard University Press, 2007. 674 pp. $35.00. Journal of Cold War Studies, 18(3):196–197, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/632287;   http:
//muse.jhu.edu/article/632287/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/3/196/699051/jcws_r_00661.pdf.

Ellison:2003:MFE

[Ell03] Herbert J. Ellison. Mending Fences: The Evolution of Moscow’s China Policy from Brezhnev to Yeltsin (review). Journal of Cold War Studies, 5(2):125–127, Spring 2003. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
40759;   http://muse.jhu.edu/article/40759/pdf;   https:/
/direct.mit.edu/jcws/article-pdf/5/2/125/700422/jcws.2003.
5.2.125.pdf.

Ellison:2009:RSC

[Ell09] Herbert J. Ellison. Robert Service, Comrades! A History of World Communism. Journal of Cold War Studies, 11(1):179– 180, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/259201;
http://muse.jhu.edu/article/259201/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/1/179/1468565/jcws.2009.11.1.
179.pdf.

Ellman:2012:CSG

[Ell12] Michael Ellman. Commentary on Stalin’s genocides. Journal of Cold War Studies, 14(3):172–175, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/490177;   http://muse.jhu.edu/article/490177/

pdf. See [GRS^{+}12].

Elman:2001:SPS

[Elm01] Colin Elman. In the Shadow of Power: States and Strategies in International Politics (review). Journal of Cold War Studies, 3(1):132– 133, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9171;   http:/
/muse.jhu.edu/article/9171/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/1/132/695416/jcws.2001.3.1.132.pdf.

Elman:2004:CDD

[Elm04] Miriam Fendius Elman. Compound Dilemmas: Democracy, Collective Action, and Superpower Rivalry (review). Journal of Cold War Studies, 6(1):95–96, Winter 2004. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
51864;   http://muse.jhu.edu/article/51864/pdf;   https:/
/direct.mit.edu/jcws/article-pdf/6/1/95/695903/jcws.2004.
6.1.95.pdf.

English:2002:CRR

[Eng02] Robert D. English. Condemned to Repetition? The Rise, Fall, and Reprise of Soviet–Russian Military Interventionism, 1973– 1996 (review). Journal of Cold War Studies, 4(3):148–151, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9273;   http:/
/muse.jhu.edu/article/9273/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/148/700254/jcws.2002.4.3.148.pdf.

Engerman:2003:RCW

[Eng03] David C. Engerman. Rethinking Cold War universities: Some recent histories. Journal of Cold War Studies, 5(3):80–95, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45190;   http://muse.jhu.edu/
article/45190/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/80/695824/152039703322286782.pdf.

English:2005:SNT

[Eng05] Robert English. The sociology of new thinking: Elites, identity change, and the end of the Cold War. Journal of Cold War Studies, 7 (2):43–80, Spring 2005. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/181899;
http://muse.jhu.edu/article/181899/pdf;   https://direct.
mit.edu/jcws/article-pdf/7/2/43/696218/1520397053630628.
pdf.

Engel:2007:WCB

[Eng07] Jeffrey A. Engel. Warren A. Chin, British Weapons Acquisition Policy and the Futility of Reform, 1945 to the Present. Aldershot, UK: Ashgate Publishing, 2004. 300 pp. Journal of Cold War Studies, 9(2):179–181, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214734;   http://muse.jhu.edu/article/214734/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/179/696481/
jcws.2007.9.2.179.pdf.

Engerman:2008:BKB

[Eng08] David C. Engerman. Bruce Kuklick, Blind Oracles: Intellectuals and War from Kennan to Kissinger. Journal of Cold War Studies, 10(2):153–155, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/237157;   http://muse.jhu.edu/article/237157/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/2/153/696999/
jcws.2008.10.2.153.pdf.

Engelbrekt:2009:CAS

[Eng09] Kjell Engelbrekt. Christine Agius, The Social Construction of Swedish Neutrality: Challenges to Swedish Identity and Sovereignty; Robert Dalsjö, Life-Line Lost: The Rise and Fall of “Neutral” Sweden’s Secret Reserve Option of Wartime Help from the West. Journal of Cold War Studies, 11(1):154–157, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/259190;   http://muse.jhu.edu/article/259190/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/1/154/
1468506/jcws.2009.11.1.154.pdf.

Engerman:2020:ACE

[Eng20a] David Engerman. Anna Calori, ed., Between East and South: Spaces of Interaction in the Globalizing Economy of the Cold War. Journal of Cold War Studies, 22(3):256–258, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/22/3/256/1860616/jcws_r_
00956.pdf;   https://muse.jhu.edu/article/762900.

Engerman:2020:RGP

[Eng20b] David C. Engerman. Red Globalization: The Political Economy of the Soviet Cold War from Stalin to Khrushchev by Oscar Sanchez-Sibony (review). Journal of Cold War Studies, 22(1):256–257, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/1/256/699836/jcws_r_00914.pdf;   https://muse.jhu.edu/
article/749643.

Ertz:2010:HKVa

[Ert10a] Simon Ertz. Hiroaki Kuromiya, The Voices of the Dead: Stalin’s Great Terror in the 1930s. Journal of Cold War Studies, 12(1):198– 200, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382512;

http://muse.jhu.edu/article/382512/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/1/198/1627437/jcws.2010.12.1.
198.pdf.

Ertz:2010:HKVb

[Ert10b] Simon Ertz. Hiroaki Kuromiya, The Voices of the Dead: Stalin’s Great Terror in the 1930s. Journal of Cold War Studies, 12(4):207–210, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/403814;   http:
//muse.jhu.edu/article/403814/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/4/207/697571/jcws_r_00066.pdf.

Esherick:2008:C

[Esh08] Joseph W. Esherick. Commentary. Journal of Cold War Studies, 10(2):116–119, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/237153;   http://muse.jhu.edu/article/237153/pdf.

Esselstrom:2015:WFC

[Ess15] Erik Esselstrom. From wartime friend to Cold War fiend: The abduction of Kaji Wataru and U.S.–Japan relations at Occupation’s end. Journal of Cold War Studies, 17(3):159–183, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595074;   http:
//muse.jhu.edu/article/595074/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/159/1102448/jcws_a_00564.pdf.

Etcheson:2023:BRE

[Etc23] Craig Etcheson. Book review: The Emergence of Global Maoism: China’s Red Evangelism and the Cambodian Communist Movement, 1949–1979 by Matthew Galway. Journal of Cold War Studies, 25(1): 222–224, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/25/1/222/2073745/jcws_r_01132.pdf.

Etheridge:2005:HCA

[Eth05] Brian Craig Etheridge. Hunting Captain Ahab: Psychological Warfare and the Melville Revival (review). Journal of Cold War Studies, 7(4):143–145, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188259;   http://muse.jhu.edu/article/188259/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/143/696277/
jcws.2005.7.4.143.pdf.

Evangelista:2000:CWI

[Eva00] Matthew Evangelista. Cold War Illusions: America, Europe, and Soviet Power, 1969–1989 (review). Journal of Cold War Studies, 2(1):124–126, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9125;   http://muse.jhu.edu/article/9125/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/1/124/695193/
jcws.2000.2.1.124.pdf.

Evangelista:2001:NHE

[Eva01] Matthew Evangelista. Norms, heresthetics, and the end of the Cold War. Journal of Cold War Studies, 3(1):5–35, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9172;   http:/
/muse.jhu.edu/article/9172/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/1/5/695410/15203970151032137.pdf.

Evangelista:2007:RMW

[Eva07] Matthew Evangelista. Ross Mackenzie, When Stars and Stripes Met Hammer and Sickle: The Chautauqua Conferences on US-Soviet Relations, 1985–1989. Journal of Cold War Studies, 9 (3):173–175, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219452;   http://muse.jhu.edu/article/219452/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/173/696640/
jcws.2007.9.3.173.pdf.

Evangelista:2009:LNS

[Eva09] Matthew Evangelista. Leopoldo Nuti, La sfida nucleare: La politica estera italiana e le armi atomiche 1945–1991. Journal of Cold War Studies, 11(2):152–155, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/265867;   http://muse.jhu.edu/article/265867/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/2/152/697197/
jcws.2009.11.2.152.pdf.

Evangelista:2013:TMM

[Eva13] Matthew Evangelista. Thomas Michael McDonnell, The United States, International Law, and the Struggle against Terrorism. New York: Routledge, 2011. 298 pp. Journal of Cold War Studies, 15(1):149–151, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/505767;   http://muse.jhu.edu/article/505767/pdf;

 https://direct.mit.edu/jcws/article-pdf/15/1/149/699976/
jcws_r_00312.pdf.

Evangelista:2023:BRD

[Eva23] Matthew A. Evangelista. Book review: Dreams for a Decade: International Nuclear Abolitionism and the End of the Cold War by Stephanie L. Freeman. Journal of Cold War Studies, 25(3):212–215, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/3/
212/2158666/jcws_r_01164.pdf.

Falk:2005:WWW

[Fal05] Andrew J. Falk. Words at War: World War II Era Radio Drama and the Postwar Broadcasting Industry Blacklist (review). Journal of Cold War Studies, 7(4):141–143, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188260;   http://muse.jhu.edu/article/188260/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/141/696252/
jcws.2005.7.4.141.pdf.

Falciola:2020:TRB

[Fal20] Luca Falciola. Transnational relationships between the Italian revolutionary left and Palestinian militants during the Cold War. Journal of Cold War Studies, 22(4):31–70, Fall 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/22/4/31/1860602/jcws_a_
00966.pdf;   https://muse.jhu.edu/article/776959.

Fischer:2016:INN

[FAM16] Thomas Fischer, Juhana Aunesluoma, and Aryo Makko. Introduction: Neutrality and nonalignment in world politics during the Cold War. Journal of Cold War Studies, 18(4):4– 11, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/645905;
http://muse.jhu.edu/article/645905/pdf;   https://direct.
mit.edu/jcws/article-pdf/18/4/4/700033/jcws_a_00677.pdf.

Farish:2015:CRA

[Far15] Matthew Farish. The Conquest of the Russian Arctic by Paul R. Josephson (review). Journal of Cold War Studies, 17(3):251–253, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595090;   http:

//muse.jhu.edu/article/595090/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/251/1102476/jcws_r_00579.pdf.

Fava:2018:BBI

[Fav18] Valentina Fava. Between business interests and ideological marketing: The USSR and the Cold War in Fiat corporate strategy, 1957– 1972. Journal of Cold War Studies, 20(4):26–64, Fall 2018. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718199;   https://direct.mit.
edu/jcws/article-pdf/20/4/26/1969630/jcws_a_00822.pdf.

Fedyashin:2017:FCW

[Fed17] Anton Fedyashin. The first Cold War spy novel: The origins and after life of Humphrey Slater’s conspirator. Journal of Cold War Studies, 19(3):134–159, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668160;   http://muse.jhu.edu/article/668160/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/134/699208/
jcws_a_00756.pdf.

Fein:2010:THW

[Fei10] Helen Fein. Theodore Hamerow, Why We Watched: How Anti-Semitism in the Allied Nations Allowed Hitler to Exterminate European Jewry. Journal of Cold War Studies, 12(1):210–211, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382495;
http://muse.jhu.edu/article/382495/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/1/210/1627459/jcws.2010.12.1.
210.pdf.

Feldstein:2005:IBG

[Fel05] Ruth Feldstein. Imperial Brotherhood: Gender and the Making of Cold War Foreign Policy (review). Journal of Cold War Studies, 7(1):193–195, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/179555;   http://muse.jhu.edu/article/179555/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/1/193/700476/
jcws.2005.7.1.193.pdf.

Ferguson:2007:SSS

[Fer07] Charles D. Ferguson. Susanna Schrafstetter and Stephen Twigge, Avoiding Armageddon: Europe, the United States and the Struggle for Nuclear Nonproliferation, 1945–1970. Westport, CT: Praeger, 2004. 256 pp. $69.95. Journal of Cold War Studies, 9(2):177–179, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/214735;   http:
//muse.jhu.edu/article/214735/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/2/177/696459/jcws.2007.9.2.177.pdf.

Ferris:2008:SSE

[Fer08] Jesse Ferris. Soviet support for Egypt’s intervention in Yemen, 1962–1963. Journal of Cold War Studies, 10(4):5–36, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/252652;   http:
//muse.jhu.edu/article/252652/pdf;   https://direct.mit.
edu/jcws/article-pdf/10/4/5/697071/jcws.2008.10.4.5.pdf.

Ferris:2011:GCA

[Fer11] Jesse Ferris. Guns for cotton?: Aid, trade, and the Soviet quest for base rights in Egypt, 1964–1966. Journal of Cold War Studies, 13 (2):4–38, Spring 2011. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/435665;
http://muse.jhu.edu/article/435665/pdf;   https://direct.
mit.edu/jcws/article-pdf/13/2/4/700662/jcws_a_00101.pdf.

Fernlund:2015:GHM

[Fer15] Kevin Jon Fernlund. Gretchen Heefner, The Missile Next Door: The Minuteman in the American Heartland. Cambridge, MA: Harvard University Press, 2012. 294 pp. Journal of Cold War Studies, 17(4):197–199, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606036;   http://muse.jhu.edu/article/606036/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/197/698783/
jcws_r_00604.pdf.

Fessler:2007:EYO

[Fes07] Diane Burke Fessler. Emily Yellin, Our Mothers’ War: American Women at Home and at the Front during World War II. New York: Free Press, 2004. 448 pp. $26.00. Journal of Cold War Studies, 9(4):166–167, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223541;   http://muse.jhu.edu/article/223541/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/166/696759/
jcws.2007.9.4.166.pdf.

Fetter:2020:AHY

[Fet20] Henry D. Fetter. Alger Hiss at Yalta: A reassessment of Hiss’s arguments against including any of the Soviet republics as initial UN members. Journal of Cold War Studies, 22(1):46–88, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/1/46/699804/jcws_a_00925.pdf;   https://muse.jhu.edu/
article/749657.

Friedman:2022:MPR

[FF22] Max Paul Friedman and Roberto García Ferreira. Making peaceful revolution impossible: Kennedy, Arévalo, the 1963 coup in Guatemala, and the Alliance against Progress in Latin America’s Cold War. Journal of Cold War Studies, 24(1):155–187, Winter 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/24/1/155/1980875/
jcws_a_01058.pdf;   https://muse.jhu.edu/article/843080.

Field:2003:CSC

[Fie03] Mark G. Field. The Cure: a Story of Cancer and Politics from the Annals of the Cold War (review). Journal of Cold War Studies, 5(3):157–158, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45192;   http://muse.jhu.edu/
article/45192/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/157/695939/jcws.2003.5.3.157.pdf.

Fink:2004:AIC

[Fin04] Carole K. Fink. America and the Intellectual Cold Wars in Europe (review). Journal of Cold War Studies, 6(2):71–73, Spring 2004. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/53668;   http://muse.jhu.edu/
article/53668/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/2/71/695991/jcws.2004.6.2.71.pdf.

Fink:2005:PPG

[Fin05] Carole K. Fink. Power and Protest: Global Revolution and the Rise of Detente (review). Journal of Cold War Studies, 7 (3):155–157, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185100;   http://muse.jhu.edu/article/185100/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/155/1102555/
jcws.2005.7.3.155.pdf.

Fink:2013:DMG

[Fin13] Carole Fink. A. Dirk Moses, German Intellectuals and the Nazi Past. New York: Cambridge University Press, 2007. 293 pp. $80.00. Journal of Cold War Studies, 15(2):150–151, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/509714;   http://muse.jhu.edu/article/509714/
pdf;   https://direct.mit.edu/jcws/article-pdf/15/2/150/
698197/jcws_r_00355.pdf.

Fischer:2003:CTR

[Fis03] Mary Ellen Fischer. Communist Terror in Romania: Gheorghiu-Dej and the Police State, 1948–1965 (review). Journal of Cold War Studies, 5(3):152–154, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/45193;   http://muse.jhu.edu/article/45193/pdf;
https://direct.mit.edu/jcws/article-pdf/5/3/152/695927/jcws.
2003.5.3.152.pdf.

Fischer:2016:LCS

[FM16] Thomas Fischer and Daniel Möckli. The limits of compensation: Swiss neutrality policy in the Cold War. Journal of Cold War Studies, 18 (4):12–35, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/645906;
http://muse.jhu.edu/article/645906/pdf;   https://direct.
mit.edu/jcws/article-pdf/18/4/12/700043/jcws_a_00678.pdf.

Footitt:2021:LRF

[Foo21] Hilary Footitt. L’Espagne Républicaine: French Policy and Spanish Republicanism in Liberated France by David A. Messenger (review). Journal of Cold War Studies, 23(2):173–174, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/23/2/173/1919925/jcws_r_
01020.pdf;   https://muse.jhu.edu/article/794720.

Forsyth:2002:SUA

[For02] Douglas J. Forsyth. Gli Stati Uniti e l’apertura a sinistra. Importanza e limiti della presenza americana in Italia (review). Journal of Cold War Studies, 4(2):141–143, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9249;   http://muse.jhu.edu/article/9249/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/2/141/695607/
jcws.2002.4.2.141.pdf.

Forsberg:2005:EII

[For05] Tuomas Forsberg. Economic incentives, ideas, and the end of the Cold War: Gorbachev and German unification. Journal of Cold War Studies, 7(2):142–164, Spring 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/181900;   http://muse.jhu.edu/article/181900/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/2/142/696195/
1520397053630556.pdf.

Formigoni:2019:PRW

[For19] Guido Formigoni. Le Péril rouge: Washington face à l’eurocommunisme by Frédéric Heurtebize (review). Journal of Cold War Studies, 21(3): 264–266, Summer 2019. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/731289;
 https://direct.mit.edu/jcws/article-pdf/21/3/264/699685/
jcws_r_00903.pdf.

Forlenza:2020:SOP

[For20] Rosario Forlenza. In search of order: Portrayal of Communists in Cold War Italy. Journal of Cold War Studies, 22(2):94–132, Spring 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/2/94/1860568/jcws_a_00878.pdf;   https://muse.jhu.edu/
article/756826.

Fousek:2014:WAC

[Fou14] John Fousek. Western Anti-Communism and the Interdoc Network: Cold War Internationale by Giles Scott-Smith (review). Journal of Cold War Studies, 16(4):232–235, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566662;   http://muse.jhu.edu/article/566662/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/232/698571/
jcws_r_00482.pdf.

Fox:2009:WSD

[Fox09] John F. Fox, Jr. What the spiders did: U.S. and Soviet counterintelligence before the Cold War. Journal of Cold War Studies, 11(3):206–224, Summer 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/268394;   http://muse.jhu.edu/article/268394/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/3/206/697248/
jcws.2009.11.3.206.pdf.

Forestier-Peyrat:2021:CWP

[FP21] Étienne Forestier-Peyrat. The Cold War politics of Soviet Federal structures, 1945–1965: International dimensions and domestic consequences. Journal of Cold War Studies, 23(3):175–207, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/23/3/175/1955781/
jcws_a_01026.pdf;   https://muse.jhu.edu/article/801908.

Frazier:2009:KUT

[Fra09] Robert Frazier. Kennan, “Universalism,” and the Truman doctrine. Journal of Cold War Studies, 11(2):3–34, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.

jhu.edu/article/265887;   http://muse.jhu.edu/article/265887/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/2/3/697213/
jcws.2009.11.2.3.pdf.

Freedman:2001:CGB

[Fre01] Lawrence Freedman. Containing Germany: Britain and the Arming of the Federal Republic (review). Journal of Cold War Studies, 3(3): 94–95, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9201;   http:/
/muse.jhu.edu/article/9201/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/94/695448/152039701750419538.pdf.

Freiberger:2007:CKM

[Fre07] Steven Z. Freiberger. Candace Karp, Missed Opportunities: U.S. Diplomatic Failures and the Arab-Israeli Conflict 1947–1967. Claremont, CA: Regina Books, 2004. 309 pp. Journal of Cold War Studies, 9(2):183–185, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214736;   http://muse.jhu.edu/article/214736/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/183/696485/
jcws.2007.9.2.183.pdf.

Friedman:2002:CWS

[Fri02] Norman Friedman. Cold War at Sea: High-Seas Confrontation between the United States and the Soviet Union (review). Journal of Cold War Studies, 4(1):101–103, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9226;   http://muse.jhu.edu/article/9226/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/1/101/695524/
jcws.2002.4.1.101.pdf.

Friedman:2004:BTN

[Fri04] Norman Friedman. Building the Trident Network: a Study of the Enrollment of People, Knowledge, and Machines (review). Journal of Cold War Studies, 6(2):94–96, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/53669;   http://muse.jhu.edu/
article/53669/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/2/94/696009/jcws.2004.6.2.94.pdf.

Friedman:2005:SSU

[Fri05] Norman Friedman. Shield and Sword: The United States Navy and the Persian Gulf War (review). Journal of Cold War Studies, 7(4):152–155, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188261;   http://muse.jhu.edu/article/188261/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/152/696307/
jcws.2005.7.4.152.pdf.

Friedman:2010:SSS

[Fri10] Jeremy Friedman. The Sino–Soviet Split: Cold War in the Communist World (review). Journal of Cold War Studies, 12(1):146–151, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/382489;   http://
muse.jhu.edu/article/382489/pdf.

Friedman:2018:EME

[Fri18] Jeremy Friedman. The enemy of my enemy: The Soviet Union, East Germany, and the Iranian Tudeh Party’s support for Ayatollah Khomeini. Journal of Cold War Studies, 20(2):3–37, Spring 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/697059;   https://direct.
mit.edu/jcws/article-pdf/20/2/3/699415/jcws_a_00815.pdf.

Frost:2023:CCD

[Fro23] Jennifer Frost. Cinema as cultural diplomacy and the Cold War: U.S. participation in international film festivals behind the Iron Curtain, 1959–1971. Journal of Cold War Studies, 25(1):75–100, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/1/
75/2073493/jcws_a_01122.pdf.

Fewsmith:2017:DXC

[FTRP17] Joseph Fewsmith, Frederick C. Teiwes, Sergey Radchenko, and Alexander V. Pantsov. Deng Xiaoping, China, and the World. Journal of Cold War Studies, 19(4):211–225, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/683671;   http://muse.jhu.edu/article/683671/
pdf;   https://direct.mit.edu/jcws/article-pdf/19/4/211/
699409/jcws_c_00771.pdf.

Gaddis:2013:RC

[Gad13] John Lewis Gaddis. Reply to the commentaries. Journal of Cold War Studies, 15(4):241–245, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/534423;   http://muse.jhu.edu/article/534423/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/4/241/698359/
jcws_a_00408.pdf. See [Kam13].

Gaffney:2007:PBL

[Gaf07] H. H. Gaffney. Peter B. Lane and Ronald E. Marcello, eds., 0Warriors and Scholars: a Modern War Reader. Journal of Cold War Studies, 9(1):119–120, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210077;   http://muse.jhu.edu/article/210077/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/119/696514/
jcws.2007.9.1.119.pdf.

Gaffney:2011:RCR

[Gaf11] H. H. Gaffney. Robert C. Rowland and John M. Jones, Reagan at Westminster: Foreshadowing the End of the Cold War. College Station, TX: Texas A&M University Press, 2010. 160 pp. $19.95. Journal of Cold War Studies, 13(4):224–226, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/457204;   http://muse.jhu.edu/article/457204/
pdf;   https://direct.mit.edu/jcws/article-pdf/13/4/224/
697868/jcws_r_00184.pdf.

Gaffney:2014:EUE

[Gaf14] Henry H. Gaffney. Euromissiles as the ultimate evolution of theater nuclear forces in Europe. Journal of Cold War Studies, 16(1):180–199, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/543729;   http://muse.jhu.edu/article/543729/pdf;

 https://direct.mit.edu/jcws/article-pdf/16/1/180/698913/
jcws_a_00435.pdf.

Gallicchio:2007:THR

[Gal07] Marc S. Gallicchio. Tsuyoshi Hasegawa, Racing the Enemy: Stalin, Truman, and the Surrender of Japan. Cambridge, MA: Belknap Press, 2005. 382 pp. $29.95. Journal of Cold War Studies, 9(4):168–170, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223542;   http://muse.jhu.edu/article/223542/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/168/696767/
jcws.2007.9.4.168.pdf.

Galpin:2009:GMS

[Gal09] Timothy J. Galpin. Gannon McHale, Stealth Boat: Fighting the Cold War in a Fast-Attack Submarine. Journal of Cold War Studies, 11 (4):162–163, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/366256;
http://muse.jhu.edu/article/366256/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/4/162/1627633/jcws.2009.11.4.
162.pdf.

Galpin:2015:CWC

[Gal15] Timothy J. Galpin. Cold War Command: The Dramatic Story of a Nuclear Submariner by Richard Woodman and Dan Conley (review). Journal of Cold War Studies, 17(3):287–289, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595108;   http:
//muse.jhu.edu/article/595108/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/287/1102440/jcws_r_00554.pdf.

Galeotti:2016:JHN

[Gal16a] Mark Galeotti. Jonathan Haslam, Near and Distant Neighbors: a New History of Soviet Intelligence. New York: Farrar, Straus & Giroux, 2015. xxiv + 367 pp. $30.00. Journal of Cold War Studies, 18(3):217–218, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632297;   http://muse.jhu.edu/article/632297/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/217/699037/
jcws_r_00649.pdf.

Galpin:2016:ASM

[Gal16b] Timothy J. Galpin. Alfred Scott McLaren, Silent and Unseen: On Patrol in Three Cold War Attack Submarines. Annapolis: U.S. Naval Institute Press, 2015. 256 pp. $39.95. Journal of Cold War Studies, 18(2):194–195, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/620555;   http://muse.jhu.edu/article/620555/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/2/194/698871/
jcws_r_00628.pdf.

Galpin:2022:EDC

[Gal22] Timothy J. Galpin. Emergency deep: Cold War missions of a submarine commander. Journal of Cold War Studies, 24(2):162–164, Spring 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/2/
162/2013028/jcws_r_01080.pdf.

Ganguly:2006:NSA

[Gan06] Sumit Ganguly. No Strings Attached?: India’s Policies and Foreign Aid, 1947–1966 (review). Journal of Cold War Studies, 8(1):152–153, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/192207;   http://muse.jhu.edu/article/192207/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/1/152/696323/
jcws.2006.8.1.152.pdf.

Ganguly:2007:SBK

[Gan07] Sumit Ganguly. Sumantra Bose, Kashmir: Roots of Conflict, Paths to Peace. Journal of Cold War Studies, 9(1):144–146, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/210078;   http:
//muse.jhu.edu/article/210078/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/1/144/696532/jcws.2007.9.1.144.pdf.

Ganguly:2008:RTR

[Gan08] Sumit Ganguly. Raghuvendra Tanwar, Reporting the Partition of Punjab, 1947: Press, Public and Other Opinions. Journal of Cold War Studies, 10(3):160–162, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241512;   http://muse.jhu.edu/article/241512/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/160/697107/
jcws.2008.10.3.160.pdf.

Ganguly:2016:SRG

[Gan16] Sumit Ganguly. Srinath Raghavan, 1971: a Global History of the Creation of Bangladesh. Cambridge, MA: Harvard University Press, 2013. 358 pp. $29.95. Journal of Cold War Studies, 18(3):194–195, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632286;   http://muse.jhu.edu/article/632286/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/194/699043/
jcws_r_00660.pdf.

Gardner:2000:JWF

[Gar00a] Lloyd Gardner. J. William Fulbright, Vietnam, and the Search for a Cold War Foreign Policy (review). Journal of Cold War Studies, 2(2):115–117, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9139;   http://muse.jhu.edu/article/9139/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/2/115/695255/
jwcs.2000.2.2.115.pdf.

Garthoff:2000:SHI

[Gar00b] Raymond L. Garthoff. Spy Hunter: Inside the FBI Investigation of the Walker Espionage Case (review). Journal of Cold War Studies, 2(2):127–129, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9140;   http://muse.jhu.edu/article/9140/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/2/127/695267/
jwcs.2000.2.2.127.pdf.

Gardner:2001:FFC

[Gar01a] Paul F. Gardner. Feet to the Fire: CIA Covert Action in Indonesia, 1957–1958 (review). Journal of Cold War Studies, 3(1):145–146, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9173;   http:/
/muse.jhu.edu/article/9173/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/1/145/695442/jcws.2001.3.1.145.pdf.

Garrison:2001:CLF

[Gar01b] Mark Garrison. Czechoslovakia’s Lost Fight for Freedom 1967– 1969: an American Embassy Perspective (review). Journal of Cold War Studies, 3(3):113–115, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9202;   http://muse.jhu.edu/article/9202/pdf;

 https://direct.mit.edu/jcws/article-pdf/3/3/113/695458/
jcws.2001.3.3.113.pdf.

Gardner:2002:CSB

[Gar02] Paul F. Gardner. Confronting Sukarno: British, American, Australian and New Zealand Diplomacy in the Malaysian–Indonesian Confrontation, 1961–5 (review). Journal of Cold War Studies, 4(3):121–123, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9275;   http:/
/muse.jhu.edu/article/9275/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/121/700230/jcws.2002.4.3.121.pdf.

Gardner:2004:CCS

[Gar04a] Paul F. Gardner. Conflict and Confrontation in South East Asia, 1961–1965: Britain, the United States, Indonesia and the Creation of Malaysia (review). Journal of Cold War Studies, 6(2):77–78, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/53670;   http://
muse.jhu.edu/article/53670/pdf;   https://direct.mit.edu/
jcws/article-pdf/6/2/77/695976/jcws.2004.6.2.77.pdf.

Garfinkle:2004:PYB

[Gar04b] Adam M. Garfinkle. Put Your Bodies upon the Wheels: Student Revolt in the 1960s (review). Journal of Cold War Studies, 6(1):99–101, Winter 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/51866;   http://
muse.jhu.edu/article/51866/pdf;   https://direct.mit.edu/
jcws/article-pdf/6/1/99/695911/jcws.2004.6.1.99.pdf.

Garofano:2004:HAU

[Gar04c] John Garofano. Historical analogies and the use of force. Journal of Cold War Studies, 6(2):64–68, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/53671;   http://muse.jhu.edu/
article/53671/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/2/64/695957/152039704773254777.pdf.

Garthoff:2004:FIH

[Gar04d] Raymond L. Garthoff. Foreign intelligence and the historiography of the Cold War. Journal of Cold War Studies, 6(2):21–56, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/53672;   http://muse.jhu.edu/

article/53672/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/2/21/695951/152039704773254759.pdf.

Gardner:2009:ASN

[Gar09] Lloyd C. Gardner. Asaf Siniver, Nixon, Kissinger, and U.S. Foreign Policy Making: The Machinery of Crisis. Journal of Cold War Studies, 11(4):165–167, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/366257;   http://muse.jhu.edu/article/366257/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/4/165/
1627637/jcws.2009.11.4.165.pdf.

Garrison:2010:CWE

[Gar10a] Mark Garrison. Commentary on “when empathy failed”. Journal of Cold War Studies, 12(2):75–78, Spring 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/382200;   http://muse.jhu.edu/article/382200/
pdf;   https://direct.mit.edu/jcws/article-pdf/12/2/75/697505/
jcws.2010.12.2.75.pdf. See [BL10a].

Garthoff:2010:GBG

[Gar10b] Raymond L. Garthoff. Gordon S. Barrass, The Great Cold War: a Journey through the Hall of Mirrors. Journal of Cold War Studies, 12(3):123–129, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392782;   http://muse.jhu.edu/article/392782/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/123/697588/
jcws_r_00009.pdf.

Garthoff:2010:ECO

[Gar10c] Raymond L. Garthoff. On empathy and critical oral history: a commentary. Journal of Cold War Studies, 12(2):79–88, Spring 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/382201;   http:
//muse.jhu.edu/article/382201/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/2/79/697507/jcws.2010.12.2.79.pdf.

See [BL10a].

Gardner:2013:LHN

[Gar13] Lloyd Gardner. Lien-Hang T. Nguyen, Hanoi’s War: an International History of the War for Peace in Vietnam. Chapel Hill: University of North Carolina Press, 2012. 444 pp. $34.95. Journal of Cold War Studies, 15(3):204–206, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522726;   http://muse.jhu.edu/article/522726/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/204/698341/
jcws_r_00381.pdf.

Gassert:2000:CWP

[Gas00] Philipp Gassert. Cold War Politics in Postwar Germany (review). Journal of Cold War Studies, 2(1):141–143, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9126;   http://muse.jhu.edu/article/
9126/pdf;   https://direct.mit.edu/jcws/article-pdf/2/1/
141/695212/jcws.2000.2.1.141.pdf.

Gasiorowski:2006:AAR

[Gas06] Mark J. Gasiorowski. Anglo–American Relations and Cold War Oil: Crisis in Iran (review). Journal of Cold War Studies, 8(1):154–155, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/192208;   http://muse.jhu.edu/article/192208/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/1/154/696343/
jcws.2006.8.1.154.pdf.

Gasiorowski:2013:CTO

[Gas13a] Mark J. Gasiorowski. The CIA’s TPBEDAMN operation and the 1953 coup in Iran. Journal of Cold War Studies, 15(4):4– 24, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/534427;
http://muse.jhu.edu/article/534427/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/4/4/698303/jcws_a_00393.pdf.

Gaskova:2013:VHV

[Gáš13b] Ivana Gášková. Václav Havel — Vilém Prečan: Korespondence [1983–1989]. Prague: Československé dokumentační středisko, o. p. s., 2011. 834 pp. 300 CZK. Journal of Cold War Studies, 15(1):185–187, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/505749;   http://muse.jhu.edu/article/505749/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/1/185/700022/
jcws_r_00330.pdf.

Gasiorowski:2019:UPC

[Gas19] Mark J. Gasiorowski. U.S. perceptions of the Communist threat in Iran during the Mossadegh era. Journal of Cold War Studies, 21(3): 185–221, Summer 2019. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/731287;
 https://direct.mit.edu/jcws/article-pdf/21/3/185/699673/
jcws_a_00898.pdf.

Gati:2003:RNE

[Gat03] Charles Gati. Redrawing Nations: Ethnic Cleansing in East–Central Europe, 1944–1948 (review). Journal of Cold War Studies, 5(3):107– 109, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/45194;   http:
//muse.jhu.edu/article/45194/pdf.

Gavin:1999:PPU

[Gav99] Francis J. Gavin. Politics, power, and U.S. policy in Iran, 1950–1953. Journal of Cold War Studies, 1(1):56–89, Winter 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9092;   https://direct.mit.edu/jcws/
article-pdf/1/1/56/695101/15203970152521890.pdf.

Gavin:2005:AGB

[Gav05] Francis J. Gavin. Architects of Globalism: Building a New World Order during World War II (review). Journal of Cold War Studies, 7(3):167–169, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185101;   http://muse.jhu.edu/article/185101/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/167/1102583/
jcws.2005.7.3.167.pdf.

Gawthorpe:2021:RGA

[Gaw21] Andrew J. Gawthorpe. Rural government advisers in South Vietnam and the U.S. war effort, 1962–1973. Journal of Cold War Studies, 23(1):196–227, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/1/196/1911792/jcws_a_00984.pdf;
https://muse.jhu.edu/article/788384.

Grzymala-Busse:2010:MBI

[GB10] Anna Grzymala-Busse. Michael Bernhard, Institutions and the Fate of Democracy. Journal of Cold War Studies, 12(1):181–183, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382499;
http://muse.jhu.edu/article/382499/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/1/181/1627433/jcws.2010.12.1.
181.pdf.

Gould-Davies:1999:RRI

[GD99] Nigel Gould-Davies. Rethinking the role of ideology in international politics during the Cold War. Journal of Cold War Studies, 1(1):90–109, Winter 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/9097;   https://direct.mit.edu/jcws/article-pdf/
1/1/90/695103/15203970152521908.pdf.

Gedmin:1999:RGR

[Ged99] Jeffrey Gedmin. Russia and Germany Reborn: Unification, the Soviet Collapse, and the New Europe, and: Origins of a Spontaneous Revolution: East Germany, 1989 (review). Journal of Cold War Studies, 1(3):178–180, Fall 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9113;   https://direct.mit.edu/jcws/
article-pdf/1/3/178/695168/jcws.1999.1.3.178.pdf.

Gehler:2005:NAN

[Geh05] Michael Gehler. From non-alignment to neutrality: Austria’s transformation during the first East–West detente, 1953–1958. Journal of Cold War Studies, 7(4):104–136, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188262;   http://muse.jhu.edu/article/188262/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/104/696266/
1520397055012451.pdf.

Gehrig:2022:ICW

[Geh22] Sebastian Gehrig. Informal Cold War envoys: West German and East German cultural diplomacy in East Asia. Journal of Cold War Studies, 24(4):112–156, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/24/4/112/2062843/jcws_a_01092.pdf.

Geist:2012:WTR

[Gei12] Edward Geist. Was there a real “mineshaft gap”?: Bomb shelters in the USSR, 1945–1962. Journal of Cold War Studies, 14(2):

3–28, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/478774;
http://muse.jhu.edu/article/478774/pdf;   https://direct.
mit.edu/jcws/article-pdf/14/2/3/697977/jcws_a_00219.pdf.

Gelman:2012:LHG

[Gel12a] Vladimir Gel’man. Larry E. Holmes, Grand Theater: Regional Governance in Stalin’s Russia, 1931–1941. Lanham, MD: Lexington Books, 2009. 257 pp. Journal of Cold War Studies, 14(4):253–255, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/500349;   http:
//muse.jhu.edu/article/500349/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/4/253/698187/jcws_r_00296.pdf.

Geltzer:2012:DMC

[Gel12b] Anna Geltzer. In a distorted mirror: The Cold War and U.S.–Soviet biomedical cooperation and (mis)understanding, 1956– 1977. Journal of Cold War Studies, 14(3):39–63, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/490174;   http:
//muse.jhu.edu/article/490174/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/3/39/698039/jcws_a_00247.pdf.

Gems:2018:KEK

[Gem18] Gerald R. Gems. Kurt Edward Kemper, College Football and American Culture in the Cold War Era. Journal of Cold War Studies, 20 (4):232–234, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/718207;
https://direct.mit.edu/jcws/article-pdf/20/4/232/1969659/
jcws_r_00809.pdf.

Gendlin:2004:OCW

[Gen04] Gerry Gendlin. Overcoming the Cold War: a History of Detente, 1950–1991 (review). Journal of Cold War Studies, 6(2):97–99, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/53673;   http://
muse.jhu.edu/article/53673/pdf;   https://direct.mit.edu/
jcws/article-pdf/6/2/97/696003/jcws.2004.6.2.97.pdf.

Genter:2017:MWD

[Gen17] Robert Genter. Matthew W. Dunne, A Cold War State of Mind: Brainwashing and Postwar American Society. Amherst: University of Massachusetts Press, 2013. 281 pp. $27.95. Journal of Cold War Studies, 19(3):250–251, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668175;   http://muse.jhu.edu/article/668175/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/250/699308/
jcws_r_00735.pdf.

George:2013:JRF

[Geo13] Roger Zane George. Joshua Rovner, Fixing the Facts: National Security and the Politics of Intelligence. Ithaca, NY: Cornell University Press, 2011. 261 pp. Journal of Cold War Studies, 15(1):153–155, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/505765;   http://muse.jhu.edu/article/505765/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/1/153/699984/
jcws_r_00314.pdf.

Gerber:2013:DSI

[Ger13] Theodore P. Gerber. David Satter, It Was a Long Time Ago, and It Never Happened Anyway: Russia and the Communist Past. New Haven: Yale University Press, 2012. 383 pp. Journal of Cold War Studies, 15(1):180–182, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/505751;   http://muse.jhu.edu/article/505751/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/1/180/700020/
jcws_r_00328.pdf.

Gery:2014:CWL

[Ger14] John Gery. Cold War Literature: Writing the Global Conflict edited by Andrew Hammond, and: Global Cold War Literature: Western, Eastern and Postcolonial Perspectives edited by Andrew Hammond (review). Journal of Cold War Studies, 16(4):228–232, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/566661;   http:
//muse.jhu.edu/article/566661/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/4/228/698554/jcws_r_00481.pdf.

Getchell:2015:RCG

[Get15] Michelle Denise Getchell. Revisiting the 1954 coup in Guatemala: The Soviet Union, the United Nations, and “hemispheric solidarity”. Journal of Cold War Studies, 17(2):73–102, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/582435;   http:

//muse.jhu.edu/article/582435/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/2/73/698678/jcws_a_00549.pdf.

Gholz:2000:CWC

[Gho00] Eugene Gholz. The Curtiss–Wright Corporation and Cold Warera defense procurement: a challenge to military–industrial complex theory. Journal of Cold War Studies, 2(1):35–75, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9127;   http:/
/muse.jhu.edu/article/9127/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/1/35/695187/15203970051032372.pdf.

Gifford:2005:DRA

[Gif05] Prosser Gifford. A Diplomatic Revolution: Algeria’s Fight for Independence and the Origins of the Post-Cold War Era (review). Journal of Cold War Studies, 7(4):170–172, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/188263;   http://muse.jhu.edu/article/188263/
pdf;   https://direct.mit.edu/jcws/article-pdf/7/4/170/696329/
jcws.2005.7.4.170.pdf.

Gillingham:2000:TCM

[Gil00] John Gillingham. A test case of Moravcsik’s “liberal intergovernmentalist” approach to European integration. Journal of Cold War Studies, 2(3):81–86, Fall 2000. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/9154;
 http://muse.jhu.edu/article/9154/pdf;   https://direct.
mit.edu/jcws/article-pdf/2/3/81/695289/15203970051032237.
pdf.

Gill:2005:SSG

[Gil05] Graeme J. Gill. Subverting the System: Gorbachev’s Reform of the Party’s Apparat, 1986–1991 (review). Journal of Cold War Studies, 7(3):197–199, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185102;   http://muse.jhu.edu/article/185102/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/197/1102599/
jcws.2005.7.3.197.pdf.

Gilman:2006:JLW

[Gil06] Nils Gilman. Jennifer Light, From Warfare to Welfare: Defense Intellectuals and Urban Problems in Cold War America. Journal of Cold War Studies, 8(2):141–144, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196276;   http://muse.jhu.edu/article/196276/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/141/700508/
jcws.2006.8.2.141.pdf.

Gill:2018:KBR

[Gil18] Lesley Gill. Kyle Burke, Revolutionaries of the Right: Anti-Communist Internationalism and Paramilitary Violence in the Cold War. Journal of Cold War Studies, 20(4):268–270, Fall 2018. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718224;   https://direct.mit.
edu/jcws/article-pdf/20/4/268/1969638/jcws_r_00850.pdf.

Gilman:2023:BRC

[Gil23] Nils Gilman. Book review: Cold War Social Science: Transnational Entanglements by Mark Solovey and Christian Day, eds. Journal of Cold War Studies, 25(1):217–219, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/25/1/217/2073749/jcws_r_
01130.pdf.

Gitelman:2007:VKE

[Git07] Zvi Y. Gitelman. Vladimir Khanin, ed., Documents on Ukrainian Jewish Identity and Emigration 1944–1990. Journal of Cold War Studies, 9(3):199–201, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219453;   http://muse.jhu.edu/article/219453/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/199/696654/
jcws.2007.9.3.199.pdf.

Gitman:2020:DCC

[Git20] Esther Gitman. The Death Camps of Croatia: Visions and Revisions, 1941–1945 by Raphael Israeli (review). Journal of Cold War Studies, 22(1):266–268, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/22/1/266/699848/jcws_r_
00918.pdf;   https://muse.jhu.edu/article/749637.

Glants:2002:SIC

[GK02] Musya Glants and Pamela Kachurin. Special issue: Culture, the Soviet Union, and the Cold War: General introduction. Journal of Cold War Studies, 4(1):3–5, Winter 2002. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/9227;
 http://muse.jhu.edu/article/9227/pdf;   https://direct.
mit.edu/jcws/article-pdf/4/1/3/695514/152039702753344799.
pdf.

Gleason:2005:DOC

[Gle05] Abbott Gleason. Debating the Origins of the Cold War: American and Russian Perspectives (review). Journal of Cold War Studies, 7(1):198–199, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/179559;   http://muse.jhu.edu/article/179559/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/1/198/700482/
jcws.2005.7.1.198.pdf.

Gleijeses:2006:MPCa

[Gle06a] Piero Gleijeses. Moscow’s proxy? Cuba and Africa 1975–1988. Journal of Cold War Studies, 8(2):3–51, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/196277;   http://muse.jhu.edu/article/196277/
pdf;   https://direct.mit.edu/jcws/article-pdf/8/2/3/700568/
jcws.2006.8.2.3.pdf.

Gleijeses:2006:MPCb

[Gle06b] Piero Gleijeses. Moscow’s proxy? Cuba and Africa 1975–1988. Journal of Cold War Studies, 8(4):98–146, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/204619;   http://muse.jhu.edu/article/204619/
pdf;   https://direct.mit.edu/jcws/article-pdf/8/4/98/700833/
jcws.2006.8.4.98.pdf.

Glynn:2006:LWT

[Gly06] Patrick Glynn. Lawrence S. Wittner, Toward Nuclear Abolition: a History of the World Disarmament Movement 1971 to the Present. Journal of Cold War Studies, 8(2):128–130, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/196278;   http:
//muse.jhu.edu/article/196278/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/2/128/700514/jcws.2006.8.2.128.pdf.

Goddeeris:2011:LAN

[God11] Idesbald Goddeeris. Lobbying allies?: The NSZZ Solidarność Coordinating Office abroad, 1982–1989. Journal of Cold War Studies, 13(3):

83–125, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/447467;
http://muse.jhu.edu/article/447467/pdf;   https://direct.
mit.edu/jcws/article-pdf/13/3/83/700694/jcws_a_00143.pdf.

Goduti:2014:KJN

[God14] Philip A. Goduti, Jr. Kennedy, Johnson and the Nonaligned World by Robert B. Rakove (review). Journal of Cold War Studies, 16(4):271–273, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566680;   http://muse.jhu.edu/article/566680/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/271/698657/
jcws_r_00499.pdf.

Goldmann:2003:HJM

[Gol03] Kjell Goldmann. Hans J. Morgenthau: an Intellectual Biography (review). Journal of Cold War Studies, 5(1):165–167, Winter 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/38252;   http://muse.jhu.edu/
article/38252/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/1/165/695786/jcws.2003.5.1.165.pdf.

Golan:2006:SUO

[Gol06a] Galia Golan. The Soviet Union and the outbreak of the June 1967 six-day war. Journal of Cold War Studies, 8(1):3–19, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/192209;
http://muse.jhu.edu/article/192209/pdf;   https://direct.
mit.edu/jcws/article-pdf/8/1/3/696275/152039706775212003.
pdf.

Goldstein:2006:UST

[Gol06b] Melvyn C. Goldstein. The United States, Tibet, and the Cold War. Journal of Cold War Studies, 8(3):145–164, Summer 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/201234;   http:
//muse.jhu.edu/article/201234/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/3/145/696404/jcws.2006.8.3.145.pdf.

Golan:2007:VIB

[Gol07] Galia Golan. Victor Israelyan, On the Battlefields of the Cold War. Journal of Cold War Studies, 9(3):168–170, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/219454;   http:
//muse.jhu.edu/article/219454/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/3/168/696650/jcws.2007.9.3.168.pdf.

Golan:2010:YRB

[Gol10] Galia Golan. Yaacov Ro’i and Boris Morozov, eds., The Soviet Union and the June 1967 Six Day War. Journal of Cold War Studies, 12(3):195–197, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392811;   http://muse.jhu.edu/article/392811/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/195/697678/
jcws_r_00039.pdf.

Goldsworthy:2011:ABA

[Gol11] David Goldsworthy. Andrea Benvenuti, Anglo-Australian Relations and the “Turn to Europe,” 1961–1972. Woodbridge, UK: The Boydell Press, 2008. 215 pp. Journal of Cold War Studies, 13(3):232–234, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447483;   http://muse.jhu.edu/article/447483/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/232/700706/
jcws_r_00116.pdf.

Golubev:2012:MSE

[Gol12] Alexey Golubev. Mikhail Suprun, ed., Kholodnaya voina v Arktike [The Cold War in the Arctic]. Arkhangelsk: Pomor State University Press, 2009. 380 pp. Journal of Cold War Studies, 14(1):143–147, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467700;   http://muse.jhu.edu/article/467700/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/143/697925/
jcws_r_00211.pdf.

Goldgeier:2020:NEP

[Gol20] James Goldgeier. NATO enlargement and the problem of value complexity. Journal of Cold War Studies, 22(4):146–174, Fall 2020. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/22/4/146/1860598/
jcws_a_00968.pdf;   https://muse.jhu.edu/article/776962.

Goodman:2005:WTK

[Goo05] Michael S. Goodman. Who is trying to keep what secret from whom and why? MI5–FBI relations and the Klaus Fuchs case. Journal of Cold War Studies, 7(3):124–146, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/185103;   http:
//muse.jhu.edu/article/185103/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/3/124/1102551/1520397054377160.pdf.

Goodman:2010:BGC

[Goo10] Giora Goodman. The British government and the challenge of Mc-Carthyism in the early Cold War. Journal of Cold War Studies, 12(1):62–97, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/382487;   http://muse.jhu.edu/article/382487/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/1/62/1627465/
jcws.2010.12.1.62.pdf.

Goodman:2015:HDD

[Goo15] Michael S. Goodman. Huw Dylan, Defence Intelligence and the Cold War: Britain’s Joint Intelligence Bureau 1945–1964. Oxford, UK: Oxford University Press, 2014. xvi + 240 pp. £60.00. Journal of Cold War Studies, 17(4):201–203, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606038;   http://muse.jhu.edu/article/606038/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/201/698798/
jcws_r_00606.pdf.

Goodman:2017:HDD

[Goo17] Michael S. Goodman. Huw Dylan, Defence Intelligence and the Cold War: Britain’s Joint Intelligence Bureau 1945–1964. Oxford, UK: Oxford University Press, 2014. xvi + 240 pp. £60.00. Journal of Cold War Studies, 19(3):244–245, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668172;   http://muse.jhu.edu/article/668172/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/244/699258/
jcws_r_00732.pdf.

Gorenburg:2001:MBV

[Gor01] Dmitry P. Gorenburg. Massovye besporyadki v SSSR pri Khrushcheve i Brezhneve (mass disturbances in the USSR under Khrushchev and Brezhnev). Journal of Cold War Studies, 3(1):138–140, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9174;   http:/
/muse.jhu.edu/article/9174/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/1/138/695418/jcws.2001.3.1.138.pdf.

Gorenburg:2003:NCE

[Gor03] Dmitry P. Gorenburg. Nationalism and Communism in Eastern Europe and the Soviet Union (review). Journal of Cold War Studies, 5(3): 139–141, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/45195;
 http://muse.jhu.edu/article/45195/pdf;   https://direct.
mit.edu/jcws/article-pdf/5/3/139/695870/jcws.2003.5.3.139.
pdf.

Gorenburg:2005:RRL

[Gor05] Dmitry P. Gorenburg. Resistance and Rebellion: Lessons of Eastern Europe, and: Understanding Ethnic Violence: Fear, Hatred, and Resentment in Twentieth-Century Eastern Europe (review). Journal of Cold War Studies, 7(3):182–186, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/185104;   http:
//muse.jhu.edu/article/185104/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/3/182/1102597/jcws.2005.7.3.182.pdf.

Gori:2009:SBT

[Gor09] Francesca Gori. Stefano Bottoni, Transilvania rossa: Il comunismo romeno e la questione nazionale (1944–1965). Journal of Cold War Studies, 11(2):131–133, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/265877;   http://muse.jhu.edu/article/265877/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/2/131/697169/
jcws.2009.11.2.131.pdf.

Gosnell:2008:TSI

[Gos08] Jonathan Gosnell. Todd Shepard, The Invention of Decolonization: The Algerian War and the Remaking of France. Journal of Cold War Studies, 10(2):141–143, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/237158;   http://muse.jhu.edu/article/237158/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/2/141/696989/
jcws.2008.10.2.141.pdf.

Gotz:2008:CIC

[Göt08] Norbert Götz. “In a class by itself”: Cold War politics and Finland’s position vis-à-vis the United Nations, 1945–1956. Journal of Cold War Studies, 10(2):71–96, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/237152;   http://muse.jhu.edu/article/237152/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/2/71/697013/

jcws.2008.10.2.71.pdf. See brief erratum [Ano08i].

Griech-Polelle:2011:CRG

[GP11] Beth A. Griech-Polelle. Charles R. Gallagher, S.J., Vatican Secret Diplomacy: Joseph P. Hurley and Pope Pius XII. New Haven: Yale University Press, 2008. 283 pp. Journal of Cold War Studies, 13(4):239–242, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457211;   http://muse.jhu.edu/article/457211/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/239/697804/
jcws_r_00123.pdf.

Gramer:2005:FPE

[Gra05] Regina U. Gramer. The Failure of Peace in Europe, 1943–1948 (review). Journal of Cold War Studies, 7(1):196–198, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/179560;   http:
//muse.jhu.edu/article/179560/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/1/196/700488/jcws.2005.7.1.196.pdf.

Graziosi:2007:MBR

[Gra07] Andrea Graziosi. Marco Buttino, La rivoluzione capovolta: L’Asia centrale tra il crollo dell’impero zarista e la formazione dell’Urss [The Revolution Turned Upside Down: Central Asia between the Collapse of the Tsarist Empire and the Formation of the USSR]. Journal of Cold War Studies, 9(3):182–186, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219455;   http://muse.jhu.edu/article/219455/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/182/696660/
jcws.2007.9.3.182.pdf.

Grau:2012:RBAb

[Gra12a] Lester W. Grau. Rodric Braithwaite, Afgantsy: The Russians in Afghanistan 1979–1989. London: Profile Books, 2011, $29.95 hardcover. Artemy M. Kalinovsky, A Long Goodbye: The Soviet Withdrawal from Afghanistan. Cambridge, MA: Harvard University Press, 2011, $27.95 hardcover. Journal of Cold War Studies, 14(4):250–253, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500348;   http://muse.jhu.edu/article/500348/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/250/698175/
jcws_r_00295.pdf.

Grau:2012:RBAa

[Gra12b] Lester W. Grau. Rodric Braithwaite, Afgantsy: The Russians in Afghanistan 1979–1989. London: Profile Books, 2011. 432 pp. $29.95. Artemy M. Kalinovsky, A Long Goodbye: The Soviet Withdrawal from Afghanistan. Cambridge, MA: Harvard University Press, 2011. 320 pp. $27.95. Journal of Cold War Studies, 14(2):170–173, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/478788;   http://muse.jhu.edu/article/478788/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/2/170/698005/
jcws_r_00240.pdf.

Graziosi:2012:CSG

[Gra12c] Andrea Graziosi. Commentary: Stalin’s genocides, and …. Journal of Cold War Studies, 14(3):155–169, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/490177;   http://muse.jhu.edu/article/490177/

pdf. See [GRS^{+}12].

Grathwol:2013:ACC

[Gra13a] Robert P. Grathwol. Andrei Cherny, The Candy Bombers: The Untold Story of the Berlin Airlift and America’s Finest Hour. New York: G. P. Putnam’s Sons, 2008, xiv + 625 pp. Journal of Cold War Studies, 15(1):174–176, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/505754;   http://muse.jhu.edu/article/505754/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/1/174/700012/
jcws_r_00325.pdf.

Gray:2013:JFP

[Gra13b] Colin S. Gray. John France, Perilous Glory: The Rise of Western Military Power. New Haven: Yale University Press, 2011. 438 pp. £25.00. Journal of Cold War Studies, 15(3):183–185, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/522715;   http:
//muse.jhu.edu/article/522715/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/3/183/698313/jcws_r_00376.pdf.

Graziosi:2017:PFU

[Gra17] Andrea Graziosi. Political famines in the USSR and China: a comparative analysis. Journal of Cold War Studies, 19(3):42–103, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/668158;
http://muse.jhu.edu/article/668158/pdf;   https://direct.
mit.edu/jcws/article-pdf/19/3/42/699222/jcws_a_00744.pdf.

Grams:2021:RIW

[Gra21a] Jon Grams. Ripple: An investigation of the world’s most advanced high-yield thermonuclear weapon design. Journal of Cold War Studies, 23(2):133–161, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/2/133/1922984/jcws_a_01011.pdf;
https://muse.jhu.edu/article/794729.

Graziosi:2021:WSP

[Gra21b] Andrea Graziosi. The weight of the Soviet past in post-1991 Russia. Journal of Cold War Studies, 23(1):89–125, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/23/1/89/1911828/jcws_a_
00969.pdf;   https://muse.jhu.edu/article/788381.

Gregg:2001:MEM

[Gre01] Donald P. Gregg. Massive Entanglement, Marginal Influence: Carter and Korea in Crisis (review). Journal of Cold War Studies, 3(2):105– 107, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9187;   http:/
/muse.jhu.edu/article/9187/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/2/105/695388/jcws.2001.3.2.105.pdf.

Greenstein:2005:ULJ

[Gre05a] Fred I. Greenstein. An Unfinished Life: John F. Kennedy, 1917– 1963 (review). Journal of Cold War Studies, 7(3):169–171, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/185105;   http:
//muse.jhu.edu/article/185105/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/3/169/1102593/jcws.2005.7.3.169.pdf.

Greenstein:2005:AFF

[Gre05b] Fred I. Greenstein. Averting “The Final Failure”: John F. Kennedy and the Secret Cuban Missile Crisis Meetings (review). Journal of Cold War Studies, 7(4):145–146, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188264;   http://muse.jhu.edu/article/188264/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/145/696285/
jcws.2005.7.4.145.pdf.

Greiner:2008:CWA

[Gre08] Bernd Greiner. The Cold War after Stalin’s Death: a Missed Opportunity for Peace? (review). Journal of Cold War Studies, 10(2): 137–138, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/237155;
http://muse.jhu.edu/article/237155/pdf.

Greiner:2010:CLP

[Gre10] Bernd Greiner. Christopher Layne, The Peace of Illusions: American Grand Strategy from 1940 to the Present. Journal of Cold War Studies, 12(1):211–213, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/382497;   http://muse.jhu.edu/article/382497/
pdf;   https://direct.mit.edu/jcws/article-pdf/12/1/211/
1627463/jcws.2010.12.1.211.pdf.

Greiner:2011:RC

[Gre11] Bernd Greiner. Reply to the commentaries. Journal of Cold War Studies, 13(3):199–204, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447988;   http://muse.jhu.edu/article/447988/pdf.

Gregory:2020:PSI

[Gre20] Paul Gregory. Peter Savodnik, The Interloper: Lee Harvey Oswald Inside the Soviet Union. Journal of Cold War Studies, 22 (2):167–168, Spring 2020. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL https://direct.mit.edu/jcws/
article-pdf/22/2/167/1860566/jcws_r_00943.pdf;   https://
muse.jhu.edu/article/756829.

Grieder:2010:WLS

[Gri10] Peter Grieder. Wilfried Loth, Die Sowjetunion und die deutsche Frage: Studien zur sowjetischen Deutschlandpolitik von Stalin bis Chruschtschow. Journal of Cold War Studies, 12(1):208–210, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382498;
http://muse.jhu.edu/article/382498/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/1/208/1627457/jcws.2010.12.1.
208.pdf.

Grossman:2002:CDB

[Gro02] Andrew D. Grossman. Civil Defense Begins at Home: Militarization Meets Everyday Life in the Fifties (review). Journal of Cold War Studies, 4(2):137–139, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9250;   http://muse.jhu.edu/article/9250/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/2/137/695601/
jcws.2002.4.2.137.pdf.

Grose:2007:RWC

[Gro07] Peter Grose. Richard W. Cutler, Counterspy: Memoirs of a Counterintelligence Officer in World War II and the Cold War. Journal of Cold War Studies, 9(3):207–210, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/219456;   http:
//muse.jhu.edu/article/219456/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/3/207/696674/jcws.2007.9.3.207.pdf.

Graziosi:2012:FRM

[GRS^{+}12] Andrea Graziosi, Joshua Rubenstein, Roman Szporluk, Paul Hollander, Jeffrey Hardy, et al. Forum: Reappraising mass terror, repression, and responsibility in Stalin’s regime: Perspectives on Norman Naimark’s Stalin’s Genocides. Journal of Cold War Studies, 14(3):149–189, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/490177;   http://muse.jhu.edu/article/490177/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/3/149/698045/

jcws_a_00250.pdf. See commentaries [Rub12, Hol12, Gra12c, Har12a, Ell12, Szp12, Ros12] and reply [Nai12].

Grundy:2007:RBE

[Gru07] Kenneth W. Grundy. Robert B. Edgerton, Africa’s Armies: From Honor to Infamy — A History from 1791 to the Present. Boulder, CO: Westview Press, 2002. 328 pp. $18.00 paper. Journal of Cold War Studies, 9(2):172–173, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214737;   http://muse.jhu.edu/article/214737/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/172/696455/
jcws.2007.9.2.172.pdf.

Goldgeier:2019:CWA

[GSP^{+}19] James Goldgeier, Thomas W. Simons, Jr., Vladimir Pechatnov, Vladislav Zubok, Dan Caldwell, Jenny Thompson, and Sherry Thompson. Cold War adviser: Llewellyn Thompson and the making of U.S. policy toward the Soviet Union. Journal of Cold War Studies, 21(3):222–257, Summer 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/731288;   https://direct.mit.
edu/jcws/article-pdf/21/3/222/699679/jcws_c_00899.pdf.

Guerrero:2017:PBC

[Gue17] Javier Gil Guerrero. Propaganda broadcasts and Cold War politics: The Carter Administration’s outreach to Islam. Journal of Cold War Studies, 19(1):4–37, Winter 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://www.mitpressjournals.org/doi/full/10.1162/JCWS_a_
00716;   https://direct.mit.edu/jcws/article-pdf/19/1/4/
699102/jcws_a_00716.pdf.

Gundle:2002:HGM

[Gun02] Stephen Gundle. Hollywood glamour and mass consumption in postwar Italy. Journal of Cold War Studies, 4(3):95–118, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9276;   http:/
/muse.jhu.edu/article/9276/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/95/700228/152039702320201085.pdf.

Gundle:2012:FAM

[Gun12] Stephen Gundle. Francesco Adinolfi, Mondo Exotica: Sounds, Visions, Obsessions of the Cocktail Generation. Durham, NC: Duke University Press, 2008. 376 pp. $84.95 cloth, $23.94 paper. Journal of Cold War Studies, 14(1):139–141, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467698;   http://muse.jhu.edu/article/467698/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/139/697927/
jcws_r_00209.pdf.

Gunn:2015:CTU

[Gun15] Christopher Gunn. The 1960 coup in Turkey: a U.S. intelligence failure or a successful intervention? Journal of Cold War Studies, 17(2):103–139, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/582436;   http://muse.jhu.edu/article/582436/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/2/103/698686/
jcws_a_00550.pdf.

Guoqiang:2010:FUC

[Guo10] Dong Guoqiang. The first uprising of the Cultural Revolution at Nanjing University: Dynamics, nature and interpretation. Journal of Cold War Studies, 12(3):30–49, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392777;   http://muse.jhu.edu/article/392777/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/30/697610/
jcws_a_00002.pdf.

Gustafson:2010:LZQ

[Gus10] Kristian C. Gustafson. Lubna Z. Qureshi, Nixon, Kissinger and Allende: US Involvement in the 1973 Coup in Chile. Journal of Cold War Studies, 12(3):153–155, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392793;   http://muse.jhu.edu/article/392793/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/153/697618/
jcws_r_00020.pdf.

Guttmann:2023:CDO

[Gut23] Aviva Guttmann. Covert diplomacy to overcome a crisis: West German and Israeli intelligence after the Munich Olympics attack. Journal of Cold War Studies, 25(4):101–126, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/25/4/101/2204973/jcws_a_
01172.pdf.

Guzzini:2013:THR

[Guz13] Stefano Guzzini. Ted Hopf, Reconstructing the Cold War: The Early Years, 1945–1958. New York: Oxford University Press, 2012. 305 pp. $39.95. Journal of Cold War Studies, 15(2):131–134, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/509706;   http:
//muse.jhu.edu/article/509706/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/2/131/698153/jcws_r_00340.pdf.

Haas:2005:CWE

[Haa05] Mark L. Haas. Cold War Endgame: Oral History, Analysis, Debates (review). Journal of Cold War Studies, 7(1):189–191, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/179561;   http:
//muse.jhu.edu/article/179561/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/1/189/700474/jcws.2005.7.1.189.pdf.

Haas:2013:RHC

[Haa13] Susan D. Haas. Richard H. Cummings, Radio Free Europe’s “Crusade for Freedom”: Rallying Americans behind Cold War Broadcasting, 1950–1960. Jefferson, NC: McFarland & Co., 2010. 265 pp. $45.00. Journal of Cold War Studies, 15(1):182–185, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/505750;   http:
//muse.jhu.edu/article/505750/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/1/182/700026/jcws_r_00329.pdf.

Haas:2020:IAF

[Haa20] Mark L. Haas. Ideologies of American Foreign Policy by John Callaghan, Brendon O’Connor, and Mark Phythian (review). Journal of Cold War Studies, 22(1):260–262, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/22/1/260/699856/jcws_r_
00922.pdf;   https://muse.jhu.edu/article/749641.

Hack:2003:BSA

[Hac03] Karl Hack. Britain, Southeast Asia and the Onset of the Pacific War (review). Journal of Cold War Studies, 5(3):133–134, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45196;   http://muse.jhu.edu/
article/45196/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/133/695848/jcws.2003.5.3.133.pdf.

Hafemeister:2008:NHS

[Haf08a] David Hafemeister. Nigel Hey, The Star Wars Enigma: Behind the Scenes of the Cold War Race for Missile Defense. Journal of Cold War Studies, 10(2):139–141, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/237159;   http://muse.jhu.edu/article/237159/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/2/139/696991/
jcws.2008.10.2.139.pdf.

Haftendorn:2008:KWZ

[Haf08b] Helga Haftendorn. Klaus Wiegrefe, Das Zerwürfnis: Helmut Schmidt, Jimmy Carter und die Krise der deutsch-amerikanischen Beziehungen. (German) [The Discord: Helmut Schmidt, Jimmy Carter and the Crisis in German–American Relations]. Journal of Cold War Studies, 10(1):165–168, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231575;   http://muse.jhu.edu/article/231575/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/165/700155/
jcws.2008.10.1.165.pdf.

Haftendorn:2008:OCM

[Haf08c] Helga Haftendorn. Otis C. Mitchell, The Cold War in Germany: Overview, Origins, and Intelligence Wars. Journal of Cold War Studies, 10(4):167–168, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/252672;   http://muse.jhu.edu/article/252672/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/4/167/697067/
jcws.2008.10.4.167.pdf.

Haig:2016:PUV

[Hai16] Fiona Haig. The Poznań uprising of 1956 as viewed by French and Italian Communists. Journal of Cold War Studies, 18(2):160–187, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/620550;   http:
//muse.jhu.edu/article/620550/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/2/160/698838/jcws_a_00641.pdf.

Hajimu:2012:KWT

[Haj12] Masuda Hajimu. The Korean war through the prism of Chinese society: Public reactions and the shaping of “reality” in the Communist state, October–December 1950. Journal of Cold War Studies, 14(3): 3–38, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/490173;
http://muse.jhu.edu/article/490173/pdf;   https://direct.
mit.edu/jcws/article-pdf/14/3/3/698049/jcws_a_00246.pdf.

Hale:2007:EWW

[Hal07a] Henry E. Hale. Edward W. Walker, Dissolution: Sovereignty and the Breakup of the Soviet Union. Journal of Cold War Studies, 9(1):169–171, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210079;   http://muse.jhu.edu/article/210079/pdf;

 https://direct.mit.edu/jcws/article-pdf/9/1/169/696569/
jcws.2007.9.1.169.pdf.

Hall:2007:JAS

[Hal07b] Karl Hall. James T. Andrews, Science for the Masses: The Bolshevik State, Public Science, and the Popular Imagination in Soviet Russia, 1917–1934. Journal of Cold War Studies, 9 (1):158–159, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210080;   http://muse.jhu.edu/article/210080/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/158/696564/
jcws.2007.9.1.158.pdf.

Halvorson:2016:CWS

[Hal16] Dan Halvorson. From Cold War solidarity to transactional engagement: Reinterpreting Australia’s relations with East Asia, 1950–1974. Journal of Cold War Studies, 18(2):130–159, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/620549;   http:
//muse.jhu.edu/article/620549/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/2/130/698830/jcws_a_00640.pdf.

Hamby:2015:RHF

[Ham15] Alonzo L. Hamby. Robert H. Ferrell, Harry S. Truman and the Cold War Revisionists. Columbia: University of Missouri Press, 2006, paper 2015. 142 pp. $19.95. Journal of Cold War Studies, 17(4):203–204, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606039;   http://muse.jhu.edu/article/606039/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/203/698807/
jcws_r_00607.pdf.

Hamby:2018:GMS

[Ham18] Alonzo Hamby. Grant Madsen, Sovereign Soldiers: How the U.S. Military Transformed the Global Economy after World War II. Journal of Cold War Studies, 20(4):270–272, Fall 2018. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718225;   https://direct.mit.
edu/jcws/article-pdf/20/4/270/1969665/jcws_r_00851.pdf.

Hanson:2013:WDL

[Han13] Philip Hanson. Werner D. Lippert, The Economic Diplomacy of Ostpolitik: Origins of NATO’s Energy Dilemma. New York:

Berghahn Books, 2011. xviii + 238 pp. Journal of Cold War Studies, 15(2):154–156, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/509716;   http://muse.jhu.edu/article/509716/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/2/154/698199/
jcws_r_00357.pdf.

Hanania:2017:TDN

[Han17] Richard Hanania. Tracing the development of the nuclear taboo: The Eisenhower Administration and four crises in East Asia. Journal of Cold War Studies, 19(2):43–83, Spring 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http:

//www.mitpressjournals.org/doi/full/10.1162/JCWS_a_00740;
 https://direct.mit.edu/jcws/article-pdf/19/2/43/699239/
jcws_a_00740.pdf.

Harper:2008:RBD

[Har08] John L. Harper. Robert L. Beisner, Dean Acheson: a Life in the Cold War. Journal of Cold War Studies, 10(3):156–158, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/241510;
http://muse.jhu.edu/article/241510/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/3/156/697093/jcws.2008.10.3.
156.pdf.

Hartmann:2011:RMA

[Har11] Susan M. Hartmann. Richard M. Abrams, America Transformed: Sixty Years of Revolutionary Change, 1941–2001. New York: Cambridge University Press, 2006. 345 pp. $37.00. Journal of Cold War Studies, 13(3):230–232, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447482;   http://muse.jhu.edu/article/447482/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/230/700700/
jcws_r_00115.pdf.

Hardy:2012:CSG

[Har12a] Jeffrey S. Hardy. Commentary on Stalin’s genocides. Journal of Cold War Studies, 14(3):169–171, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/490177;   http://muse.jhu.edu/article/490177/

pdf. See [GRS^{+}12].

Harrison:2012:DVU

[Har12b] Hope M. Harrison. Dirk Verheyen, United City, Divided Memories? Cold War Legacies in Contemporary Berlin. Lanham, MD: Lexington Books, 2010. 299 pp. $36.99. Journal of Cold War Studies, 14(4):239–242, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500343;   http://muse.jhu.edu/article/500343/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/239/698155/
jcws_r_00290.pdf.

Harrison:2015:BGL

[Har15] Hope M. Harrison. Born in the GDR: Living in the Shadow of the Wall by Hester Vaizey (review). Journal of Cold War Studies, 17(3):283–285, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595106;   http://muse.jhu.edu/article/595106/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/283/
1102430/jcws_r_00552.pdf.

Harland:2018:CBK

[Har18] David M. Harland. Colin Burgess, Kate Doolan, and Bert Vis, Fallen Astronauts: Heroes Who Died Reaching for the Moon. Journal of Cold War Studies, 20(4):241–243, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718212;   https://direct.mit.
edu/jcws/article-pdf/20/4/241/1969634/jcws_r_00831.pdf.

Harmer:2019:CQC

[Har19] Tanya Harmer. The “Cuban question” and the Cold War in Latin America, 1959–1964. Journal of Cold War Studies, 21(3):114–151, Summer 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/731285;   https:
//direct.mit.edu/jcws/article-pdf/21/3/114/699627/jcws_a_
00896.pdf.

Harrison:2023:BRD

[Har23] Mark Harrison. Book review: Dictatorship and Daily Life in 20th Century Europe by Lisa Pine. Journal of Cold War Studies, 25(4): 249–250, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/25/4/249/2204971/jcws_r_01177.pdf.

Haseman:2007:BRT

[Has07] John B. Haseman. E. Bruce Reynolds, Thailand’s Secret War: OSS, SOE, and the Free Thai Underground during World War II. Cambridge, UK: Cambridge University Press, 2005. 462 pp. £50.00/$85.00. Journal of Cold War Studies, 9(4):170–172, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/223543;   http:
//muse.jhu.edu/article/223543/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/4/170/696751/jcws.2007.9.4.170.pdf.

Hasegawa:2013:SFU

[Has13] Tsuyoshi Hasegawa. The Soviet factor in U.S.–Japanese defense cooperation, 1978–1985. Journal of Cold War Studies, 15(2):72– 103, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/509704;
http://muse.jhu.edu/article/509704/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/2/72/698139/jcws_a_00338.pdf.

Hatzivassiliou:2009:IAN

[Hat09] Evanthis Hatzivassiliou. Images of the adversary: NATO assessments of the Soviet Union, 1953–1964. Journal of Cold War Studies, 11(2):89–116, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/265884;   http://muse.jhu.edu/article/265884/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/2/89/697215/
jcws.2009.11.2.89.pdf.

Hathaway:2010:GBN

[Hat10] Robert M. Hathaway. Gregg Brazinsky, Nation Building in South Korea: Koreans, Americans, and the Making of a Democracy. Journal of Cold War Studies, 12(1):168–169, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/382500;   http://muse.jhu.edu/article/382500/
pdf;   https://direct.mit.edu/jcws/article-pdf/12/1/168/
1627425/jcws.2010.12.1.168.pdf.

Hatzivassiliou:2011:JEM

[Hat11] Evanthis Hatzivassiliou. James Edward Miller, The United States and the Making of Modern Greece: History and Power, 1950–1974. Chapel Hill: University of North Carolina Press, 2009. xvi + 301 pp. Journal of Cold War Studies, 13(4):267–269, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/457223;   http://muse.jhu.edu/article/457223/
pdf;   https://direct.mit.edu/jcws/article-pdf/13/4/267/
697877/jcws_r_00135.pdf.

Hatzivassiliou:2017:CNP

[Hat17] Evanthis Hatzivassiliou. The crisis of NATO political consultation, 1973–1974: From DEFCON III to the Atlantic Declaration. Journal of Cold War Studies, 19(3):104–133, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/668159;   http://muse.jhu.edu/article/668159/
pdf;   https://direct.mit.edu/jcws/article-pdf/19/3/104/
699228/jcws_a_00755.pdf.

Hatzivassiliou:2021:IIS

[Hat21] Evanthis Hatzivassiliou. Images of the international system and the Cold War in Star Trek, 1966–1991. Journal of Cold War Studies, 23 (1):55–88, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/23/1/55/1911814/jcws_a_00981.pdf;   https://muse.jhu.
edu/article/788380.

Hawkins:2007:PGL

[Haw07] Darren G. Hawkins. Paul Gordon Lauren, The Evolution of International Human Rights: Visions Seen, 2nd ed. Philadelphia: University of Pennsylvania Press, 2003. 397 pp. $29.95. Journal of Cold War Studies, 9(4):174–176, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223544;   http://muse.jhu.edu/article/223544/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/174/696771/
jcws.2007.9.4.174.pdf.

Haynes:1999:HWS

[Hay99] John Earl Haynes. The Haunted Wood: Soviet Espionage in America — The Stalin Era (review). Journal of Cold War Studies, 1 (2):121–123, Spring 1999. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/9100;
 https://direct.mit.edu/jcws/article-pdf/1/2/121/695141/
jcws.1999.1.2.121.pdf.

Haynes:2000:CWD

[Hay00] John Earl Haynes. The Cold War debate continues: a traditionalist view of historical writing on domestic Communism and anti-Communism. Journal of Cold War Studies, 2(1):76–115, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9128;   http:/
/muse.jhu.edu/article/9128/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/1/76/695185/15203970051032381.pdf.

Haynes:2003:CSH

[Hay03] John Earl Haynes. Chasing Spies: How the FBI Failed in Counterintelligence but Promoted the Politics of McCarthyism in the Cold War Years (review). Journal of Cold War Studies, 5(2):117–119, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/40760;   http://
muse.jhu.edu/article/40760/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/2/117/700410/jcws.2003.5.2.117.pdf.

Haynes:2006:KOR

[Hay06] John Earl Haynes. Kathryn S. Olmsted, Red Spy Queen: a Biography of Elizabeth Bentley, and Lauren Kessler, Clever Girl: Elizabeth Bentley’s Life in and Out of Espionage. Journal of Cold War Studies, 8(2):148–150, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196279;   http://muse.jhu.edu/article/196279/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/148/700530/
jcws.2006.8.2.148.pdf;   https://direct.mit.edu/jcws/article-
pdf/8/2/148/700538/jcws.2006.8.2.148a.pdf.

Haynes:2011:NWO

[Hay11] John Earl Haynes. Nigel West and Oleg Tsarev, eds., Triplex: Secrets from the Cambridge Spies. New Haven: Yale University Press, 2009. 363 pp. Journal of Cold War Studies, 13(3):229–230, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/447481;   http:
//muse.jhu.edu/article/447481/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/3/229/700684/jcws_r_00114.pdf.

Haynes:2013:JLC

[Hay13] John Earl Haynes. Jennifer Luff, Commonsense Anticommunism: Labor and Civil Liberties between the World Wars. Chapel Hill: University of North Carolina Press, 2012. xii, 288 pp. Journal of Cold War Studies, 15(2):140–141, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/15/2/140/698169/jcws_r_00343.pdf.

Haynes:2015:VPB

[Hay15a] John Earl Haynes. A Very Principled Boy: The Life of Duncan Lee, Red Spy and Cold Warrior by Mark A. Bradley (review). Journal of Cold War Studies, 17(2):150–152, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/582442;   http://muse.jhu.edu/article/582442/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/2/150/
698698/jcws_r_00540.pdf.

Haynes:2015:RAC

[Hay15b] John Earl Haynes. Red Apple: Communism and McCarthyism in Cold War New York by Phillip Deery (review). Journal of Cold War Studies, 17(3):238–240, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595083;   http://muse.jhu.edu/article/595083/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/238/
1102472/jcws_r_00572.pdf.

Haynes:2016:FCH

[Hay16] John Earl Haynes. Frank Close, Half-Life: The Divided Life of Bruno Pontecorvo, Physicist or Spy. New York: Basic Books, 2015. xix + 378 pp. Journal of Cold War Studies, 18(3):233–236, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/632304;   http:
//muse.jhu.edu/article/632304/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/3/233/699065/jcws_r_00670.pdf.

Hayden:2017:MBV

[Hay17] Robert M. Hayden. Max Bergholz, Violence as a Creative Force: Identity, Nationalism and Memory in a Balkan Community. Journal of Cold War Studies, 19(4):251–254, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683681;   http://muse.jhu.edu/article/683681/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/251/699443/

jcws_r_00781.pdf. See erratum [Ano18j].

Heilbrunn:2006:CWT

[Hei06] Jacob Heilbrunn. Cold War Triumphalism: The Misuse of History after the Fall of Communism (review). Journal of Cold War Studies, 8(1):150–151, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/192210;   http://muse.jhu.edu/article/192210/pdf;

 https://direct.mit.edu/jcws/article-pdf/8/1/150/696349/
jcws.2006.8.1.150.pdf.

Heiss:2015:ERC

[Hei15] Mary Ann Heiss. Exposing “red Colonialism”: U.S. propaganda at the United Nations, 1953–1963. Journal of Cold War Studies, 17(3):82–115, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/595072;   http://muse.jhu.edu/article/595072/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/3/82/1102405/
jcws_a_00562.pdf.

Heikkila:2016:KKL

[Hei16] Pauli Heikkilä. Katalin Kádár Lynn, ed., The Inauguration of Organized Political Warfare: Cold War Organizations Sponsored by the National Committee for a Free Europe/Free Europe Committee. Saint Helena, CA: Helena History Press, 2013. 604 pp. $75.00/ EUR 57.00/£47.00. Journal of Cold War Studies, 18(4):235–237, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/645925;   http:
//muse.jhu.edu/article/645925/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/4/235/700152/jcws_r_00697.pdf.

Helleiner:2007:FGG

[Hel07] Eric Helleiner. Francis Gavin, Gold, Dollars and Power: The Politics of International Monetary Relations, 1958–1971. Chapel Hill: University of North Carolina Press, 2004. 263 pp. $45.00. Journal of Cold War Studies, 9(2):170–171, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214738;   http://muse.jhu.edu/article/214738/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/170/696445/
jcws.2007.9.2.170.pdf.

Hellendoorn:2022:CTA

[Hel22] Elmar Hellendoorn. Containing technology and Allies alike: The Cold War, intra-NATO relations, and the U.S. centrifuge classification initiative, 1958–1962. Journal of Cold War Studies, 24(4):59–111, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/4/
59/2062825/jcws_a_01091.pdf.

Hendrickson:2009:AWC

[Hen09] Ryan C. Hendrickson. Andreas Wenger, Christian Nuenlist, and Anna Locher, eds., Transforming NATO in the Cold War: Challenges beyond Deterrence in the 1960s. Journal of Cold War Studies, 11(1): 157–159, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/259191;
http://muse.jhu.edu/article/259191/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/1/157/1468512/jcws.2009.11.1.
157.pdf.

Hershberg:1999:DSA

[Her99] James G. Hershberg. Dr. Strangelove’s America: Society and Culture in the Atomic Age (review). Journal of Cold War Studies, 1 (2):123–125, Spring 1999. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/9101;
 https://direct.mit.edu/jcws/article-pdf/1/2/123/695144/
jcws.1999.1.2.123.pdf.

Herrera:2002:RSD

[Her02] Yoshiko M. Herrera. Russia’s Stillborn Democracy? From Gorbachev to Yeltsin (review). Journal of Cold War Studies, 4(3):155–157, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9277;   http:/
/muse.jhu.edu/article/9277/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/155/700262/jcws.2002.4.3.155.pdf.

Herken:2003:ASA

[Her03a] Gregg Herken. American Science in an Age of Anxiety: Scientists, Anticommunism, and the Cold War (review). Journal of Cold War Studies, 5(1):167–169, Winter 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/38253;   http://muse.jhu.edu/
article/38253/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/1/167/695776/jcws.2003.5.1.167.pdf.

Hershberg:2003:PPB

[Her03b] James G. Hershberg. Peace probes and the bombing pause: Hungarian and Polish diplomacy during the Vietnam War, December 1965– January 1966. Journal of Cold War Studies, 5(2):32–67, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40761;   http://muse.jhu.edu/
article/40761/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/32/700412/152039703763336462.pdf.

Herring:2004:PIS

[Her04a] George C. Herring. The Pueblo Incident: a Spy Ship and the Failure of American Foreign Policy (review). Journal of Cold War Studies, 6(3):168–170, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/171161;   http://muse.jhu.edu/article/171161/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/3/168/696095/
jcws.2004.6.3.168.pdf.

Hershberg:2004:USBa

[Her04b] James G. Hershberg. The United States, Brazil, and the Cuban Missile Crisis, 1962 (Part 1). Journal of Cold War Studies, 6(2):3–20, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/53674;   http://
muse.jhu.edu/article/53674/pdf;   https://direct.mit.edu/
jcws/article-pdf/6/2/3/695965/152039704773254740.pdf.

Hershberg:2004:USBb

[Her04c] James G. Hershberg. The United States, Brazil, and the Cuban Missile Crisis, 1962 (Part 2). Journal of Cold War Studies, 6(3):5–67, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171162;   http:
//muse.jhu.edu/article/171162/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/3/5/696015/1520397041447364.pdf.

Herspring:2005:SHC

[Her05] Dale R. Herspring. The Soviet High Command: a Military– Political History, 1918–1941 (review). Journal of Cold War Studies, 7(1):203, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/179562;   http://muse.jhu.edu/article/179562/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/1/203/700480/
jcws.2005.7.1.203.pdf.

Herken:2007:JBO

[Her07a] Gregg Herken. Jeremy Bernstein, Oppenheimer: Portrait of an Enigma. Journal of Cold War Studies, 9(1):98–99, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/210081;   http:
//muse.jhu.edu/article/210081/pdf;   http://muse.jhu.edu/
journals/journal_of_cold_war_studies/summary/v009/9.1herken.
html;   https://www.math.utah.edu/pub/bibnet/authors/o/oppenheimer-

j-robert.bib;   https://direct.mit.edu/jcws/article-pdf/
9/1/98/696602/jcws.2007.9.1.98.pdf.

Herken:2007:TCR

[Her07b] Gregg Herken. Thomas C. Reed, At the Abyss: an Insider’s History of the Cold War. New York: Ballantine Books, 2004. 368 pp. $25.95. Journal of Cold War Studies, 9(2):148–150, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/214739;   http:
//muse.jhu.edu/article/214739/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/2/148/696421/jcws.2007.9.2.148.pdf.

Herspring:2008:RJS

[Her08] Dale R. Herspring. Rudolf J. Schlaffer, Der Wehrbeauftragte 1951 bis 1985: Aus Sorge um den Soldaten. Journal of Cold War Studies, 10(3):200–201, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241531;   http://muse.jhu.edu/article/241531/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/200/697159/
jcws.2008.10.3.200.pdf.

Herken:2009:TES

[Her09] Gregg Herken. Target Enormoz: Soviet nuclear espionage on the West Coast of the United States, 1942–1950. Journal of Cold War Studies, 11(3):68–90, Summer 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/268391;   http://muse.jhu.edu/article/268391/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/3/68/697254/
jcws.2009.11.3.68.pdf.

Herd:2011:CAE

[Her11a] Alexander W. G. Herd. A “common Appreciation”: Eisenhower, Canada, and continental air defense, 1953–1954. Journal of Cold War Studies, 13(3):4–26, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447463;   http://muse.jhu.edu/article/447463/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/4/700692/jcws_
a_00140.pdf.

Herring:2011:MPB

[Her11b] George C. Herring. Mark Philip Bradley, Vietnam at War. Oxford, UK: Oxford University Press, 2009. 233 pp. $29.95. Journal of Cold War Studies, 13(1):226–228, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/424248;   http://muse.jhu.edu/article/424248/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/1/226/697707/
jcws_r_00078.pdf.

Herf:2012:AOA

[Her12] Jeffrey Herf. Alessandro Orsini, Anatomy of the Red Brigades: The Religious Mind-Set of Modern Terrorists, trans. from Italian by Sarah J. Nodes. Ithaca, NY: Cornell University Press, 2011. 317 pp. Journal of Cold War Studies, 14(1):137–139, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/467697;   http://muse.jhu.edu/article/467697/
pdf;   https://direct.mit.edu/jcws/article-pdf/14/1/137/
697913/jcws_r_00208.pdf.

Hershberg:2013:RGF

[Her13] James G. Hershberg. Reflections on George F. Kennan: an American life. Journal of Cold War Studies, 15(4):155–160, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/534428;   http:
//muse.jhu.edu/article/534428/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/4/155/698331/jcws_a_00399.pdf.

Herf:2014:WIE

[Her14] Jeffrey Herf. “At war with Israel”: East Germany’s key role in Soviet policy in the Middle East. Journal of Cold War Studies, 16(3):129–163, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/556078;   http://muse.jhu.edu/article/556078/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/3/129/698473/
jcws_a_00450.pdf.

Herd:2016:CJB

[Her16] Alexander W. G. Herd. Christopher J. Bright, Continental Defense in the Eisenhower Era: Nuclear Antiaircraft Arms and the Cold War. New York: Palgrave Macmillan, 2010. 280 pp. $100.00. Journal of Cold War Studies, 18(3):202–204, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/632290;   http://muse.jhu.edu/article/632290/
pdf;   https://direct.mit.edu/jcws/article-pdf/18/3/202/
699053/jcws_r_00664.pdf.

Herd:2017:RTL

[Her17] Alexander W. G. Herd. Robert Teigrob, Living with War: Twentieth-Century Conflict in Canadian and American History and Memory. Journal of Cold War Studies, 19(4):273–275, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/683691;   http:
//muse.jhu.edu/article/683691/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/4/273/699469/jcws_r_00791.pdf.

Hershberg:2019:MHJ

[Her19] James G. Hershberg. Man of the Hour: James B. Conant, Warrior Scientist by Jennet Conant (review). Journal of Cold War Studies, 21(1):186–189, Winter 2019. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
723558;   https://direct.mit.edu/jcws/article-pdf/21/1/186/
699742/jcws_r_00858.pdf.

Herken:2020:KYW

[Her20a] Gregg Herken. Ken Young and Warner R. Schilling, Superbomb: Organizational Conflict and the Development of the Hydrogen Bomb. Journal of Cold War Studies, 22(4):229–230, Fall 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/22/4/229/1860626/jcws_r_
00971.pdf;   https://muse.jhu.edu/article/776964.

Hershberg:2020:SBR

[Her20b] James G. Hershberg. Soviet-Brazilian relations and the Cuban Missile Crisis. Journal of Cold War Studies, 22(1):175–209, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/1/175/699825/jcws_a_00930.pdf;   https://muse.jhu.edu/
article/749654.

Herf:2021:USD

[Her21] Jeffrey Herf. The U.S. State Department’s opposition to Zionist aspirations during the Early Cold War: George F. Kennan and George C. Marshall in 1947–1948. Journal of Cold War Studies, 23(4):153–180, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/4/153/1970101/jcws_a_01042.pdf;
https://muse.jhu.edu/article/835880.

Herran:2023:BON

[Her23] Néstor Herran. The “Bounties of our new servant”: Isotopes, industry, and economy before and after Atoms for Peace. Journal of Cold War Studies, 25(3):14–34, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/3/14/2158672/jcws_a_01157.pdf.

Hess:2002:ATK

[Hes02] Gary R. Hess. American Tragedy; Kennedy, Johnson, and the Origins of the Vietnam War (review). Journal of Cold War Studies, 4(3):119– 121, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9278;   http:/
/muse.jhu.edu/article/9278/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/119/700220/152039702320201094.pdf.

Hess:2012:CGC

[Hes12] Gary R. Hess. Christopher E. Goscha and Christian F. Ostermann, eds., Connecting Histories: Decolonization and the Cold War in Southeast Asia, 1945–1962. Washington, DC: Woodrow Wilson Center Press, 2009. 431 pp. Journal of Cold War Studies, 14(3):194–196, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/490180;   http://muse.jhu.edu/article/490180/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/3/194/698053/
jcws_r_00253.pdf.

Hessler:2013:DFH

[Hes13] Julie Hessler. Donald Filtzer, The Hazards of Urban Life in Late Stalinist Russia: Health, Hygiene, and Living Standards, 1943–1953. New York: Cambridge University Press, 2010. xxx + 379 pp. Journal of Cold War Studies, 15(3):218–220, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/522733;   http://muse.jhu.edu/article/522733/
pdf;   https://direct.mit.edu/jcws/article-pdf/15/3/218/
698357/jcws_r_00388.pdf.

Hess:2021:LAA

[Hes21] Gary R. Hess. LBJ: Architect of American Ambition by Randall B. Woods, Cambridge, MA: Harvard University Press, 2006. 1007 pp. (review). Journal of Cold War Studies, 23(1):231–233, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/23/1/231/1911826/
jcws_r_00988.pdf;   https://muse.jhu.edu/article/788396.

Hezel:2002:NSS

[Hez02] Francis X. Hezel. National Security and Self-determination: United States Policy in Micronesia (1961–1972) (review). Journal of Cold War Studies, 4(2):148–150, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9251;   http://muse.jhu.edu/article/9251/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/2/148/695609/
jcws.2002.4.2.148.pdf.

Henrich-Franke:2014:CCR

[HF14] Christian Henrich-Franke. Cross-curtain radio cooperation and new international alignments during the Cold War. Journal of Cold War Studies, 16(4):110–132, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566654;   http://muse.jhu.edu/article/566654/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/110/698556/
jcws_a_00517.pdf.

Hershberg:2023:MCW

[HGP^{+}23] James G. Hershberg, David Greenberg, Barbara A. Perry, Luther Spoehr, and Fredrik Logevall. The making of a Cold War President. Journal of Cold War Studies, 25(3):188–209, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/3/
188/2158678/jcws_c_01146.pdf.

Hisano:2013:ROK

[His13] Ai Hisano. Ruth Oldenziel and Karin Zachmann, eds., Cold War Kitchen: Americanization, Technology, and European Users. Cambridge, MA: MIT Press, 2009. 415 pp. $36.00. Journal of Cold War Studies, 15(1):166–168, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/505759;   http://muse.jhu.edu/article/505759/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/1/166/700010/
jcws_r_00320.pdf.

Hitchcock:2003:RFG

[Hit03] William I. Hitchcock. Response to “France and the German Question, 1945–1955,” by Michael Creswell and Marc Trachtenberg. Journal of Cold War Studies, 5(3):33–36, Summer 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45197;   http://muse.jhu.edu/
article/45197/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/33/695935/jcws.2003.5.3.33.pdf. See [CT03a].

Hitchens:2016:ITC

[Hit16] Ralph M. Hitchens. Ingo Trauschweizer, The Cold War U.S. Army. Lawrence, KS: University of Kansas Press, 2008. 416 pp. $39.95. Journal of Cold War Studies, 18(4):264–267, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/645938;   http:
//muse.jhu.edu/article/645938/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/4/264/700208/jcws_r_00713.pdf.

Hixson:2005:TGP

[Hix05] Walter L. Hixson. The Tragedy of Great Power Politics (review). Journal of Cold War Studies, 7(3):149–151, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/185106;   http:
//muse.jhu.edu/article/185106/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/3/149/1102553/jcws.2005.7.3.149.pdf.

Haynes:2009:AVN

[HK09] John Earl Haynes and Harvey Klehr. Alexander Vassiliev’s notebooks and the documentation of Soviet intelligence activities in the United States during the Stalin Era. Journal of Cold War Studies, 11(3): 6–25, Summer 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/268389;
http://muse.jhu.edu/article/268389/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/3/6/697250/jcws.2009.11.3.6.
pdf.

Haynes:2020:FWA

[HK20] John Earl Haynes and Harvey Klehr. Framing William Albertson: The FBI’s “Solo” operation and the Cold War. Journal of Cold War Studies, 22(3):63–85, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/22/3/63/1860590/jcws_a_
00951.pdf;   https://muse.jhu.edu/article/762892.

Hochschild:2016:LRS

[Hoc16] Jennifer L. Hochschild. Lawrence R. Samuel, The American Dream: a Cultural History. Syracuse, NY: Syracuse University Press. 2012. 241 pp. $24.95. Journal of Cold War Studies, 18(2):197–199, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/620559;   http:
//muse.jhu.edu/article/620559/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/2/197/698889/jcws_r_00630.pdf.

Hoffmann:2000:CM

[Hof00] Stanley Hoffmann. Comment on Moravcsik. Journal of Cold War Studies, 2(3):69–73, Fall 2000. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/9155;
 http://muse.jhu.edu/article/9155/pdf;   https://direct.
mit.edu/jcws/article-pdf/2/3/69/695278/15203970051032200.

pdf. See [Mor00b, Mor00c].

Hoffmann:2006:RLB

[Hof06] Steven A. Hoffmann. Rethinking the linkage between Tibet and the China-India border conflict: a realist approach. Journal of Cold War Studies, 8(3):165–194, Summer 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/201235;   http://muse.jhu.edu/article/201235/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/3/165/696413/
jcws.2006.8.3.165.pdf.

Hoffman:2020:UMM

[Hof20] Colonel Jon T. Hoffman. Underdogs: The Making of the Modern Marine Corps by Aaron B. O’Connell (review). Journal of Cold War Studies, 22(1):268–270, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/22/1/268/699850/jcws_r_
00919.pdf;   https://muse.jhu.edu/article/749636.

Holland:1999:LCS

[Hol99] Max Holland. A Luce connection: Senator Keating, William Pawley, and the Cuban Missile Crisis. Journal of Cold War Studies, 1(3):139–167, Fall 1999. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/9114;
 https://direct.mit.edu/jcws/article-pdf/1/3/139/695164/
152039799316976832.pdf.

Holland:2005:PSU

[Hol05a] Max Holland. Private sources of U.S. foreign policy: William Pawley and the 1954 coup d’État in Guatemala. Journal of Cold War Studies, 7(4):36–73, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/188265;

http://muse.jhu.edu/article/188265/pdf;   https://direct.
mit.edu/jcws/article-pdf/7/4/36/696245/1520397055012442.
pdf.

Hollander:2005:ESP

[Hol05b] Paul Hollander. Enemies of the State: Personal Stories from the Gulag (review). Journal of Cold War Studies, 7(1):206–208, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/179563;   http:
//muse.jhu.edu/article/179563/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/1/206/700496/jcws.2005.7.1.206.pdf.

Holub:2006:RDA

[Hol06] Renate Holub. Richard Drake, Apostles and Agitators: Italy’s Marxist Revolutionary Tradition. Journal of Cold War Studies, 8(2):174–175, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196280;   http://muse.jhu.edu/article/196280/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/174/700562/
jcws.2006.8.2.174.pdf.

Hollander:2008:DES

[Hol08] Paul Hollander. David Everitt, A Shadow of Red: Communism and the Blacklist in Radio and Television. Journal of Cold War Studies, 10(2):160–162, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/237160;   http://muse.jhu.edu/article/237160/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/2/160/697007/
jcws.2008.10.2.160.pdf.

Holland:2009:FSE

[Hol09] Max Holland. I. F. Stone: encounters with Soviet intelligence. Journal of Cold War Studies, 11(3):144–205, Summer 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/268393;   http:
//muse.jhu.edu/article/268393/pdf;   https://direct.mit.
edu/jcws/article-pdf/11/3/144/697244/jcws.2009.11.3.144.
pdf.

Hollander:2010:JRE

[Hol10] Paul Hollander. John Rodden, ed., The Cambridge Companion to George Orwell. Journal of Cold War Studies, 12(1):171–173, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382501;
http://muse.jhu.edu/article/382501/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/1/171/1627415/jcws.2010.12.1.
171.pdf.

Hollander:2012:CSG

[Hol12] Paul Hollander. Commentary on Stalin’s genocides. Journal of Cold War Studies, 14(3):153–155, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/490177;   http://muse.jhu.edu/article/490177/

pdf. See [GRS^{+}12].

Holland:2015:AHE

[Hol15] Max Holland. Andrew Hoberek, ed., The Cambridge Companion to John F. Kennedy. New York: Cambridge University Press, 2015. 268 pp. Journal of Cold War Studies, 17(4):204–206, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/606040;   http:
//muse.jhu.edu/article/606040/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/4/204/698826/jcws_r_00608.pdf.

Holloway:2021:CWW

[Hol21] David Holloway. The Cold War: A World History by Odd Arne Westad (review). Journal of Cold War Studies, 23(3):232– 235, Summer 2021. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL https://direct.mit.edu/jcws/
article-pdf/23/3/232/1955775/jcws_r_01030.pdf;   https://
muse.jhu.edu/article/801910.

Hopmann:2003:AIR

[Hop03] Terrence Hopmann. Adapting international relations theory to the end of the Cold War. Journal of Cold War Studies, 5(3):96–101, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/45198;   http://
muse.jhu.edu/article/45198/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/3/96/695812/152039703322286791.pdf.

Horne:2007:JRH

[Hor07] Gerald Horne. Jonathan Rosenberg, How Far the Promised Land? World Affairs and the American Civil Rights Movement from the First World War to Vietnam. Cambridge, MA:

Harvard University Press, 2006. Journal of Cold War Studies, 9(4):142–143, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223545;   http://muse.jhu.edu/article/223545/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/142/696737/
jcws.2007.9.4.142.pdf.

Houghton:2015:SBP

[Hou15] David Patrick Houghton. Spies and boats and planes: an examination of U.S. decision-making during the Pueblo hostage crisis of 1968. Journal of Cold War Studies, 17(4):4–40, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/606019;   http:
//muse.jhu.edu/article/606019/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/4/4/698746/jcws_a_00595.pdf.

House:2016:MJL

[Hou16] Jonathan House. Michael Jones, Leningrad: State of Siege. New York: Basic Books, 2008. 322 pp. $27.95. Journal of Cold War Studies, 18(4):223–225, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645920;   http://muse.jhu.edu/article/645920/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/223/700100/
jcws_r_00692.pdf.

Howlett:2014:BNW

[How14] Charles F. Howlett. A Band of Noble Women: Racial Politics in the Women’s Peace Movement by Melinda Plastas (review). Journal of Cold War Studies, 16(4):224–226, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566659;   http://muse.jhu.edu/article/566659/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/224/698586/
jcws_r_00479.pdf.

Howard:2022:BRF

[How22] M. T. Howard. Book review: Fighting for Time: Rhodesia’s Military and Zimbabwe’s Independence by Charles D. Melson. Journal of Cold War Studies, 24(4):218–220, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/24/4/218/2062874/jcws_r_01108.pdf.

Harrison:2017:PUM

[HP17] Mark Harrison and Viktor Pál. Perspectives on unified military industries of the Soviet Bloc. Journal of Cold War Studies, 19(3):225–230, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668165;   http://muse.jhu.edu/article/668165/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/225/699204/
jcws_c_00761.pdf.

Hager:2015:USN

[HS15] Robert P. Hager, Jr. and Robert S. Snyder. The United States and Nicaragua: Understanding the breakdown in relations. Journal of Cold War Studies, 17(2):3–35, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/582433;   http://muse.jhu.edu/article/582433/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/2/3/698684/jcws_
a_00546.pdf.

Hershberg:2023:RCM

[HSB^{+}23] James G. Hershberg, Svetlana Savranskaya, Thomas Blanton, Don Munton, and Serhii Plokhy. Reexamining the Cuban Missile Crisis six decades later. Journal of Cold War Studies, 25(4):191–221, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/4/
191/2204997/jcws_c_01176.pdf.

Hurtado-Torres:2019:CMG

[HT19] Sebastián Hurtado-Torres. The Chilean moment in the global Cold War: International reactions to Salvador Allende’s victory in the Presidential election of 1970. Journal of Cold War Studies, 21(3):26–55, Summer 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/731282;   https:
//direct.mit.edu/jcws/article-pdf/21/3/26/699636/jcws_a_
00892.pdf.

Herring:2015:PPD

[HTP^{+}15] George C. Herring, Wallace Thies, Merle L. Pribbenow, Sophie Quinn-Judge, Jessica M. Chapman, Michael R. Adamson, and James G. Hershberg. Peace Proposals, Diplomacy, and War: Was an Opportunity Lost for an Early Settlement in Vietnam? Journal of Cold War Studies, 17(1):153–180, Winter 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/581234;   http://muse.jhu.edu/article/581234/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/1/153/
700782/jcws_c_00534.pdf.

Hughes:2008:BTA

[Hug08] Geraint Hughes. Britain, the Transatlantic Alliance, and the Arab– Israeli war of 1973. Journal of Cold War Studies, 10(2):3–40, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/237150;   http:
//muse.jhu.edu/article/237150/pdf;   https://direct.mit.
edu/jcws/article-pdf/10/2/3/697009/jcws.2008.10.2.3.pdf.

Huhle:2020:AMK

[Huh20] Teresa Huhle. Albert Manke and Katěrina Březinová, eds., Kleinstaaten und sekundäre Akteure im Kalten Krieg: Politische, wirtschaftliche, militärische und kulturelle Wechselbeziehungen zwischen Europa und Lateinamerika. Journal of Cold War Studies, 22(3):266–268, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/3/266/1860636/jcws_r_00961.pdf;
https://muse.jhu.edu/article/762905.

Hurrell:2005:GCR

[Hur05] Andrew Hurrell. Global Community: The Role of International Organizations in the Making of the Contemporary World (review). Journal of Cold War Studies, 7(1):186–189, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/179564;   http:
//muse.jhu.edu/article/179564/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/1/186/700486/jcws.2005.7.1.186.pdf.

Hurrell:2010:EHR

[Hur10] Andrew Hurrell. Emma Haddad, The Refugee in International Society. Between Sovereigns. Journal of Cold War Studies, 12(2):127– 128, Spring 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382209;
http://muse.jhu.edu/article/382209/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/2/127/697497/jcws.2010.12.2.
127.pdf.

Husain:2000:CWH

[Hus00] Aiyaz Husain. Cold War in the High Himalayas: The USA, China, and South Asia in the 1950s (review). Journal of Cold War Studies, 2(3):145–147, Fall 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9156;   http://muse.jhu.edu/article/9156/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/3/145/695294/
jcws.2000.2.3.145.pdf.

Husain:2005:GDV

[Hus05] Aiyaz Husain. Grand Designs and Visions of Unity: The Atlantic Powers and the Reorganization of Western Europe, 1955– 1963 (review). Journal of Cold War Studies, 7(4):169–170, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/188266;   http:
//muse.jhu.edu/article/188266/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/4/169/696325/jcws.2005.7.4.169.pdf.

Husain:2013:NFC

[Hus13] Aiyaz Husain. Niall Ferguson, Charles S. Maier, Erez Manela, and Daniel J. Sargent, eds., The Shock of the Global: The 1970s in Perspective. Cambridge, MA: Harvard University Press, 2010. 434 pp. $29.00. Journal of Cold War Studies, 15(1):171–172, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/505756;   http:
//muse.jhu.edu/article/505756/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/1/171/700018/jcws_r_00323.pdf.

Hutchings:2001:CWN

[Hut01] Robert Hutchings. From the Cold War to a New Era: The United States and the Soviet Union, 1983–1991 (review). Journal of Cold War Studies, 3(3):97–100, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9203;   http://muse.jhu.edu/article/
9203/pdf;   https://direct.mit.edu/jcws/article-pdf/3/3/
97/695472/jcws.2001.3.3.97.pdf.

Hutchings:2019:DDU

[Hut19] Robert Hutchings. Dealing with Dictators: The United States, Hungary, and East Central Europe, 1943–1989 by László Borhi (review). Journal of Cold War Studies, 21(1):214–217, Winter 2019. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/723547;   https://direct.mit.
edu/jcws/article-pdf/21/1/214/699800/jcws_r_00868.pdf.

Hutchinson:2020:JCC

[Hut20] George B. Hutchinson. Jean-Christophe Cloutier, Shadow Archives: The Lifecycles of African American Literature. Journal of Cold War Studies, 22(3):250–253, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/3/250/1860604/jcws_r_00954.pdf;
https://muse.jhu.edu/article/762898.

Iacob:2022:PNH

[Iac22] Bogdan C. Iacob. Peripheral Nerve: Health and Medicine in Cold War Latin America ed. by Anne-Emanuelle Birn and Raúl Necochea López (review). Journal of Cold War Studies, 24(1):258–261, Winter 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/24/1/258/1980885/
jcws_r_01064.pdf;   https://muse.jhu.edu/article/843086.

Iandolo:2014:IPS

[Ian14] Alessandro Iandolo. Imbalance of power: The Soviet Union and the Congo crisis, 1960–1961. Journal of Cold War Studies, 16(2): 32–55, Spring 2014. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/547323;
http://muse.jhu.edu/article/547323/pdf;   https://direct.
mit.edu/jcws/article-pdf/16/2/32/698414/jcws_a_00449.pdf.

Iatrides:2005:RSD

[Iat05] John O. Iatrides. Revolution or self-defense? Communist goals, strategy, and tactics in the Greek Civil War. Journal of Cold War Studies, 7(3):3–33, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/185107;
http://muse.jhu.edu/article/185107/pdf;   https://direct.
mit.edu/jcws/article-pdf/7/3/3/1102561/1520397054377179.
pdf.

Iatrides:2018:AJM

[Iat18] John O. Iatrides. Assassination and judicial misconduct in Cold War Greece: The Polk/Staktopoulos case in retrospect. Journal of Cold War Studies, 20(4):65–126, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718200;   https://direct.mit.
edu/jcws/article-pdf/20/4/65/1969636/jcws_a_00823.pdf.

Immerman:2016:SLC

[Imm16] Richard H. Immerman. Stephen Long, The CIA and the Soviet Bloc: Political Warfare, the Origins of the CIA and Countering Communism in Europe. London: I. B. Tauris, 2014. 336 pp. £68.00. Journal of Cold War Studies, 18(2):207–209, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/620553;   http://muse.jhu.edu/article/620553/
pdf;   https://direct.mit.edu/jcws/article-pdf/18/2/207/
698917/jcws_r_00642.pdf.

Ingimundarson:2004:SSF

[Ing04a] Valur Ingimundarson. Stalin and the Soviet–Finnish War, 1939–1940 (review). Journal of Cold War Studies, 6(2):86–88, Spring 2004. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/53675;   http://muse.jhu.edu/
article/53675/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/2/86/695997/jcws.2004.6.2.86.pdf.

Ingimundarson:2004:IAA

[Ing04b] Valur Ingimundarson. Immunizing against the American other: Racism, nationalism, and gender in U.S.–Icelandic military relations during the Cold War. Journal of Cold War Studies, 6(4): 65–88, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/174708;
http://muse.jhu.edu/article/174708/pdf;   https://direct.
mit.edu/jcws/article-pdf/6/4/65/696060/1520397042350892.
pdf.

Ingimundarson:2017:PVA

[Ing17] Valur Ingimundarson. Poul Villaume, Ann-Marie Ekengren, and Rasmus Mariager, eds., Northern Europe in the Cold War, 1965–1990: East-West Interactions of Trade, Culture, and Security. Journal of Cold War Studies, 19(4):283–286, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683696;   http://muse.jhu.edu/article/683696/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/283/699471/
jcws_r_00797.pdf.

Iriye:2002:HSF

[Iri02] Akira Iriye. Henry L. Stimson: The First Wise Man (review). Journal of Cold War Studies, 4(4):106–108, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9299;   http://muse.jhu.edu/article/9299/pdf;

 https://direct.mit.edu/jcws/article-pdf/4/4/106/695674/
jcws.2002.4.4.106.pdf.

Iriye:2011:KHC

[Iri11] Akira Iriye. Kuan-Hsing Chen, Asia as Method: Toward Deimperialization. Durham, NC: Duke University Press, 2010. 321 pp. Journal of Cold War Studies, 13(3):220–221, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/447476;   http://muse.jhu.edu/article/447476/
pdf;   https://direct.mit.edu/jcws/article-pdf/13/3/220/
700738/jcws_r_00156.pdf.

Iriye:2014:NCH

[Iri14] Akira Iriye. The New Cambridge History of American Foreign Relations, Vol. 4: Challenges to American Primacy, 1945 to the Present by Warren Cohen (review). Journal of Cold War Studies, 16(4):259–260, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566675;   http://muse.jhu.edu/article/566675/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/259/698647/
jcws_r_00494.pdf.

Jacobs:2016:EIX

[Jac16] Justin M. Jacobs. Exile Island: Xinjiang refugees and the “one China” policy in nationalist Taiwan, 1949–1971. Journal of Cold War Studies, 18(1):188–218, Winter 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/612011;   http://muse.jhu.edu/article/612011/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/1/188/698785/
jcws_a_00624.pdf.

Jackson:2017:STL

[Jac17] Galen E. Jackson. Strategy and two-level games: U.S. domestic politics and the road to a separate peace, 1977–1978. Journal of Cold War Studies, 19(3):160–195, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668161;   http://muse.jhu.edu/article/668161/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/160/699247/
jcws_a_00757.pdf.

Jacobs:2023:BRS

[Jac23] Justin M. Jacobs. Book review: Soviet Policy in Xinjiang: Stalin and the National Movement in Eastern Turkistan by Jamil Hasanli. Journal of Cold War Studies, 25(3):224–227, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/3/
224/2158674/jcws_r_01166.pdf.

Jakovina:2016:NMH

[Jak16] Tvrtko Jakovina. Nataša Mišković, Harald Fischer-Tiné, and Nada Boškovska, eds., The Non-Aligned Movement and the Cold War: Delhi–Bandung–Belgrade. London: Routledge, 2014. 232 pp. Journal of Cold War Studies, 18(3):208–211, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/632293;   http:
//muse.jhu.edu/article/632293/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/3/208/699029/jcws_r_00645.pdf.

Janos:2008:JRG

[Jan08] Andrew C. Janos. János Rainer and György Péteri, eds., Muddling through the Long Sixties: Ideas and Everyday Life in High Politics and the Lower Classes of Communist Hungary. Journal of Cold War Studies, 10(4):170–171, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/252665;   http://muse.jhu.edu/article/252665/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/4/170/697061/
jcws.2008.10.4.170.pdf.

Jarzabek:2016:IGQ

[Jar16] Wanda Jarzabek. The impact of the German question on Polish attitudes toward CSCE, 1964–1975. Journal of Cold War Studies, 18(3):139–157, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632281;   http://muse.jhu.edu/article/632281/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/139/699033/
jcws_a_00655.pdf.

Jensen-Eriksen:2019:NFT

[JE19] Niklas Jensen-Eriksen. The Northern Front in the technological Cold War: Finland and East–West trade in the 1970s and 1980s. Journal of Cold War Studies, 21(4):150–174, Fall 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/21/4/150/699738/jcws_a_
00909.pdf;   https://muse.jhu.edu/article/736845.

Jenkins:2003:AMB

[Jen03] Philip Jenkins. The Age of McCarthyism: a Brief History With Documents (review). Journal of Cold War Studies, 5(3):120–122, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/45199;   http://
muse.jhu.edu/article/45199/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/3/120/695839/jcws.2003.5.3.120.pdf.

Jensen:2017:JRH

[Jen17] Steven L. B. Jensen. Joe Renouard, Human Rights in American Foreign Policy: From the 1960s to the Soviet Collapse. Philadelphia: University of Pennsylvania Press, 2016. 324 pp. $69.95. Journal of Cold War Studies, 19(3):247–249, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668174;   http://muse.jhu.edu/article/668174/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/247/699275/
jcws_r_00734.pdf.

Jenks:2021:USH

[Jen21] Andrew Jenks. U.S.–Soviet handshakes in space and the Cold War imaginary. Journal of Cold War Studies, 23(2):100–132, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/23/2/100/1922980/
jcws_a_01010.pdf;   https://muse.jhu.edu/article/794728.

Jervis:2000:PTE

[Jer00] Robert Jervis. Power Ties: Economic Interdependence, Balancing, and War (review). Journal of Cold War Studies, 2(1):135–137, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9129;   http:/
/muse.jhu.edu/article/9129/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/1/135/695201/jcws.2000.2.1.135.pdf.

Jervis:2001:WCW

[Jer01] Robert Jervis. Was the Cold War a security dilemma? Journal of Cold War Studies, 3(1):36–60, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9175;   http://muse.jhu.edu/article/
9175/pdf;   https://direct.mit.edu/jcws/article-pdf/3/1/
36/695402/15203970151032146.pdf.

Jervis:2006:CSP

[Jer06] Robert L. Jervis. Containment strategies in perspective. Journal of Cold War Studies, 8(4):92–97, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/204621;   http://muse.jhu.edu/article/204621/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/4/92/700829/jcws.
2006.8.4.92.pdf.

Jervis:2008:CWA

[Jer08] Robert L. Jervis. The Cold War after Stalin’s Death: a Missed Opportunity for Peace? (review). Journal of Cold War Studies, 10(2): 131–133, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/237155;
http://muse.jhu.edu/article/237155/pdf.

Jervis:2011:DLM

[Jer11] Robert Jervis. Derek Leebaert, Magic and Mayhem: The Delusions of American Foreign Policy from Korea to Afghanistan. New York: Simon & Schuster, 2010. 336 pp. $26.00. Journal of Cold War Studies, 13(3):222–223, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447477;   http://muse.jhu.edu/article/447477/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/222/700704/
jcws_r_00110.pdf.

Jervis:2013:CDP

[Jer13] Robert Jervis. Carmel Davis, Power, Threat, or Military Capabilities. Lanham, MD: University Press of America, 2011. 140 pp. $24.95. Journal of Cold War Studies, 15(3):181–183, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/522714;   http:
//muse.jhu.edu/article/522714/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/3/181/698317/jcws_r_00375.pdf.

Jervis:2016:CDP

[Jer16a] Robert Jervis. Carmel Davis, Power, Threat, or Military Capabilities. Lanham, MD: University Press of America, 2011. 140 pp. $24.95. Journal of Cold War Studies, 18(2):189–191, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/620556;   http:
//muse.jhu.edu/article/620556/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/2/189/698862/jcws_r_00626.pdf.

Jervis:2016:MHG

[Jer16b] Robert Jervis. Michael Herman and Gwilym Hughes, eds., Intelligence in the Cold War: What Difference Did It Make? New York: Routledge, 2013. 150 pp. $145.00. Journal of Cold War Studies, 18(3):190–192, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632284;   http://muse.jhu.edu/article/632284/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/190/699045/
jcws_r_00658.pdf.

Jervis:2022:MFS

[Jer22a] Robert Jervis. The many faces of SALT. Journal of Cold War Studies, 24(4):198–214, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/24/4/198/2062819/jcws_a_01105.pdf.

Jervis:2022:RDH

[Jer22b] Robert Jervis. Restricted data: The history of nuclear secrecy in the United States. Journal of Cold War Studies, 24(2):159–162, Spring 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/2/
159/2013022/jcws_r_01079.pdf.

Jaroszynska-Kirchmann:2007:TPE

[JK07] Anna D. Jaroszyńska-Kirchmann. Tadeusz Piotrowski, ed., The Polish Deportees of World War II: Recollections of Removal to the Soviet Union and Dispersal throughout the World. Journal of Cold War Studies, 9(1):155–157, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210082;   http://muse.jhu.edu/article/210082/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/155/696550/
jcws.2007.9.1.155.pdf.

Johnson:2012:RSC

[JKM12] A. Ross Johnson, Michael Kraus, and Vojtech Mastny. Reassessing the Soviet–Czechoslovak Crisis of 1968. Journal of Cold War Studies, 14(4):216–223, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500334;   http://muse.jhu.edu/article/500334/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/216/698120/
jcws_r_00281.pdf.

Jervis:2017:NWC

[JLBK17] Robert Jervis, Mark Atwood Lawrence, William Burr, and Jeffrey P. Kimball. Nuclear Weapons, Coercive Diplomacy, and the Vietnam War: Perspectives on Nixon’s Nuclear Spector. Journal of Cold War Studies, 19(4):192–210, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683670;   http://muse.jhu.edu/article/683670/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/192/699386/
jcws_c_00770.pdf.

Jiang:2023:EED

[JM23] Huajie Jiang and Kazushi Minami. The eyes and ears of the dragon: Open-source intelligence and Chinese foreign policy during the Cold War. Journal of Cold War Studies, 25(2):41–63, Spring 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/2/
41/2140956/jcws_a_01141.pdf.

Jervis:2008:PCW

[JMG08] Robert L. Jervis, Thomas Maddux, and Bernd Greiner. Perspectives on the Cold War after Stalin’s death: a missed opportunity for peace? Journal of Cold War Studies, 10(2):131– 138, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/237155;
http://muse.jhu.edu/article/237155/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/2/131/696987/jcws.2008.10.2.
131.pdf.

Jockel:2007:ESN

[Joc07] Joseph T. Jockel. Erika Simpson, NATO and the Bomb: Canadian Defenders Confront Critics. Journal of Cold War Studies, 9(1):130–132, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210083;   http://muse.jhu.edu/article/210083/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/130/696536/
jcws.2007.9.1.130.pdf.

Johnson:2001:CCW

[Joh01] Robert David Johnson. Congress and the Cold War. Journal of Cold War Studies, 3(2):76–100, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9188;   http://muse.jhu.edu/article/

9188/pdf;   https://direct.mit.edu/jcws/article-pdf/3/2/
76/695378/152039701300373899.pdf.

Johns:2003:NPA

[Joh03] Andrew L. Johns. Northern Passage: American Vietnam War Resisters in Canada (review). Journal of Cold War Studies, 5(2):86–89, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/40762;   http://
muse.jhu.edu/article/40762/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/2/86/700454/jcws.2003.5.2.86.pdf.

Johnson:2005:FPC

[Joh05] Robert David Johnson. Foreign Policy and Congress: an International Relations Perspective (review). Journal of Cold War Studies, 7(3):162–163, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185108;   http://muse.jhu.edu/article/185108/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/162/1102559/
jcws.2005.7.3.162.pdf.

Johnson:2006:JPC

[Joh06] A. Ross Johnson. John P. C. Matthews, Tinderbox: East-Central Europe in the Spring, Summer, and Early Fall of 1956. Journal of Cold War Studies, 8(4):147–148, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/204622;   http://muse.jhu.edu/article/204622/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/4/147/700793/
jcws.2006.8.4.147.pdf.

Johns:2007:JAS

[Joh07a] Andrew L. Johns. The Johnson Administration, the Shah of Iran, and the changing pattern of U.S.–Iranian relations, 1965–1967: “Tired of being treated like a schoolboy”. Journal of Cold War Studies, 9(2): 64–94, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/214740;
http://muse.jhu.edu/article/214740/pdf;   https://direct.
mit.edu/jcws/article-pdf/9/2/64/696473/jcws.2007.9.2.64.
pdf.

Johns:2007:SYC

[Joh07b] Andrew L. Johns. Salim Yaqub, Containing Arab Nationalism: The Eisenhower Doctrine and the Middle East. Chapel Hill: University of North Carolina Press, 2004. 377 pp. $22.50. Journal of Cold War Studies, 9(4):147–149, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223546;   http://muse.jhu.edu/article/223546/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/147/696739/
jcws.2007.9.4.147.pdf.

Johnson:2007:MKB

[Joh07c] Loch K. Johnson. Michael K. Bohn, The Achille Lauro Hijacking: Lessons in the Politics and Prejudice of Terrorism. Washington, DC: Brassey’s, 2004. 235 pp. $26.95. Journal of Cold War Studies, 9(4):145–147, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223547;   http://muse.jhu.edu/article/223547/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/145/696729/
jcws.2007.9.4.145.pdf.

Johnson:2007:WJD

[Joh07d] Loch K. Johnson. William J. Daugherty, Executive Secrets: Covert Action & the Presidency. Journal of Cold War Studies, 9(1):106–109, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210084;   http://muse.jhu.edu/article/210084/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/106/696471/
jcws.2007.9.1.106.pdf.

Johns:2011:WTF

[Joh11a] Andrew L. Johns. W. Taylor Fain, American Ascendance and British Retreat in the Persian Gulf Region. New York: Palgrave MacMillan, 2008. 283 pp. Journal of Cold War Studies, 13(4):217–218, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457200;   http:
//muse.jhu.edu/article/457200/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/4/217/697893/jcws_r_00179.pdf.

Johnson:2011:BSF

[Joh11b] Loch K. Johnson. Bayard Stockton, Flawed Patriot: The Rise and Fall of CIA Legend Bill Harvey. Dulles, VA: Potomac Books, 2006. 368 pp. $28.95. Journal of Cold War Studies, 13(4):215–217, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457199;   http:
//muse.jhu.edu/article/457199/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/4/215/697849/jcws_r_00178.pdf.

Johnson:2013:SDB

[Joh13a] Ian Johnson. Sohail Daulatzai, Black Star Crescent Moon: The Muslim International and Black Freedom beyond America. Minneapolis: University of Minnesota Press, 2012. 257 pp. $22.50 paperback, $67.50 cloth. Journal of Cold War Studies, 15(2):141–143, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/509710;   http:
//muse.jhu.edu/article/509710/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/2/141/698193/jcws_r_00345.pdf.

Johnson:2013:JAC

[Joh13b] Loch K. Johnson. James Angleton and the Church Committee. Journal of Cold War Studies, 15(4):128–147, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.

jhu.edu/article/534418;   http://muse.jhu.edu/article/534418/
pdf;   https://direct.mit.edu/jcws/article-pdf/15/4/128/
698315/jcws_a_00397.pdf.

Johnson:2014:LHS

[Joh14a] Oliver Johnson. Lewis H. Siegelbaum, Cars for Comrades: The Life of the Soviet Automobile. Ithaca, NY: Cornell University Press, 2008. 309 pp. Journal of Cold War Studies, 16(1):237–239, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543744;   http:
//muse.jhu.edu/article/543744/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/1/237/698952/jcws_r_00422.pdf.

Johnson:2014:ERC

[Joh14b] Robert David Johnson. Eric R. Crouse, An American Stand: Senator Margaret Chase Smith and the Communist Menace, 1948–1972. Lanham, MD: Lexington Books, 2010. 183 pp. Journal of Cold War Studies, 16(1):220–222, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/543736;   http://muse.jhu.edu/article/543736/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/1/220/698954/
jcws_r_00414.pdf.

Johnson:2014:TWD

[Joh14c] Robert David Johnson. Thomas W. Devine, Henry Wallace’s 1948 Presidential Campaign and the Future of Postwar Liberalism. Chapel Hill: University of North Carolina Press, 2013. 408 pp. Journal of Cold War Studies, 16(2):130–132, Spring 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/547329;   http://muse.jhu.edu/article/547329/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/2/130/698439/
jcws_r_00456.pdf.

Johnson:2017:LGP

[Joh17] A. Ross Johnson. Lechosław Gawlikowski, Pracownicy Radia Wolna Europa: Biografie zwykłe i niezwykłe. Journal of Cold War Studies, 19(4):254–256, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683682;   http://muse.jhu.edu/article/683682/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/254/699453/
jcws_r_00782.pdf.

Johnson:2020:WPR

[Joh20a] A. Ross Johnson. The Warsaw Pact Reconsidered: International Relations in Eastern Europe, 1955–1969 by Laurien Crump (review). Journal of Cold War Studies, 22(1):249–252, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/22/1/249/699869/jcws_r_
00935.pdf;   https://muse.jhu.edu/article/749646.

Johnson:2020:SDB

[Joh20b] A. Ross Johnson. Siobhan Doucette, Books Are Weapons: The Polish Opposition Press and the Overthrow of Communism. Journal of Cold War Studies, 22(3):253–256, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/3/253/1860606/jcws_r_00955.pdf;
https://muse.jhu.edu/article/762899.

Johnson:2021:RPR

[Joh21] A. Ross Johnson. Rozgłósnia Polska Radia Wolna Europa w latach 1950–1975 [Polish Broadcasting Station of Radio Free Europe in 1950–1975], by Rafał Habielski and Paweł Machcewicz, Wrocław: Ossolineum, 2018. 416 pp. 36.00 Polish złoty. (review). Journal of Cold War Studies, 23(1):256–259, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/1/256/1911788/jcws_r_01004.pdf;
https://muse.jhu.edu/article/788390.

Johnston:2022:BRG

[Joh22] Rosamund Johnston. Book review: Guns, Guerrillas, and the Great Leader: North Korea and the Third World by Benjamin R. Young. Journal of Cold War Studies, 24(4):229–231, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/4/
229/2062890/jcws_r_01112.pdf.

Jolluck:2007:MJC

[Jol07] Katherine R. Jolluck. Marek Jan Chodakiewicz, Between the Nazis and Soviets: Occupation Politics in Poland, 1939–1947. Journal of Cold War Studies, 9(3):201–203, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/219457;   http://muse.jhu.edu/article/219457/
pdf;   https://direct.mit.edu/jcws/article-pdf/9/3/201/696662/
jcws.2007.9.3.201.pdf.

Jones:2003:AVN

[Jon03] Matthew Jones. American Visions of the Netherlands East Indies/ Indonesia: U.S. Foreign Policy and Indonesian Nationalism, 1920– 1949 (review). Journal of Cold War Studies, 5(3):128–130, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/45200;   http://
muse.jhu.edu/article/45200/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/3/128/695901/jcws.2003.5.3.128.pdf.

Jones:2004:USP

[Jon04] Matthew Jones. United States Policy towards Indonesia in the Truman and Eisenhower Years (review). Journal of Cold War Studies, 6(4):163–166, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/174709;   http://muse.jhu.edu/article/174709/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/4/163/696125/
jcws.2004.6.4.163.pdf.

Jonas:2006:WHZ

[Jon06] Manfred Jonas. Winfried Heinemann, Vom Zusammenwachsen des Bündnisses: Die Funktionsweise der NATO in ausgewählten Krisenfällen 1951–1956. Journal of Cold War Studies, 8(2):164–166, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196281;   http://muse.jhu.edu/article/196281/pdf;

 https://direct.mit.edu/jcws/article-pdf/8/2/164/700536/
jcws.2006.8.2.164.pdf.

Jones:2007:CKC

[Jon07] Matthew Jones. Christina Klein, Cold War Orientalism: Asia in the Middlebrow Imagination, 1945–1961. Berkeley: University of California Press, 2003. xiv + 316 pp. Journal of Cold War Studies, 9(2):161–163, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214741;   http://muse.jhu.edu/article/214741/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/161/696441/
jcws.2007.9.2.161.pdf.

Jones:2008:TCU

[Jon08] Matthew Jones. Targeting China: U.S. nuclear planning and “massive retaliation” in East Asia, 1953–1955. Journal of Cold War Studies, 10(4):37–65, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/252656;   http://muse.jhu.edu/article/252656/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/4/37/697075/
jcws.2008.10.4.37.pdf.

Jones:2014:JWE

[Jon14] Matthew Jones. James Waite, The End of the First Indochina War: a Global History. Abingdon, UK: Routledge, 2012. x + 299 pp. Journal of Cold War Studies, 16(1):231–233, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/543741;   http://muse.jhu.edu/article/543741/
pdf;   https://direct.mit.edu/jcws/article-pdf/16/1/231/
698956/jcws_r_00419.pdf.

Jones:2019:PSC

[Jon19] Matthew Jones. Prelude to the Skybolt Crisis: The Kennedy Administration’s approach to British and French strategic nuclear policies in 1962. Journal of Cold War Studies, 21(2):58–109, Spring 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/726150;   https://direct.
mit.edu/jcws/article-pdf/21/2/58/699577/jcws_a_00839.pdf.

Josephson:2020:AKC

[Jos20] Paul Josephson. Alison Kraft and Carola Sachse, eds., Science, (Anti-)Communism and Diplomacy: The Pugwash Conferences on Science and World Affairs in the Early Cold War. Journal of Cold War Studies, 22(4):243–245, Fall 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/4/243/1860660/jcws_r_00978.pdf;
https://muse.jhu.edu/article/776965.

Josephson:2023:APL

[Jos23] Paul R. Josephson. Atoms for Peace in the 1950s: Lessons from the spread of nuclear technology in the Early Cold War. Journal of Cold War Studies, 25(3):6–13, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/3/6/2372950/jcws_a_01156.pdf.

Kachurin:2002:RRP

[Kac02] Pamela Kachurin. The ROCI road to peace: Robert rauschenberg, perestroika, and the end of the Cold War. Journal of Cold War Studies, 4(1):27–43, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9228;   http://muse.jhu.edu/article/
9228/pdf;   https://direct.mit.edu/jcws/article-pdf/4/1/
27/695512/152039702753344816.pdf.

Kagan:2005:DUH

[Kag05] Donald Kagan. The Dawn of Universal History: Selected Essays from a Witness to the Twentieth Century (review). Journal of Cold War Studies, 7(3):152–153, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185109;   http://muse.jhu.edu/article/185109/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/152/1102563/
jcws.2005.7.3.152.pdf.

Kalinovsky:2009:DMS

[Kal09] Artemy Kalinovsky. Decision-making and the Soviet war in Afghanistan: From intervention to withdrawal. Journal of Cold War Studies, 11(4):46–73, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/366246;   http://muse.jhu.edu/article/366246/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/4/46/1627643/
jcws.2009.11.4.46.pdf.

Kalicki:2012:CPT

[Kal12] Jan H. Kalicki. Christopher P. Twomey, The Military Lens: Doctrinal Difference and Deterrence Failure in Sino–American Relations. Ithaca, NY: Cornell University Press, 2010. 252 pp. $35.00. Journal of Cold War Studies, 14(2):160–162, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/478783;   http://muse.jhu.edu/article/478783/
pdf;   https://direct.mit.edu/jcws/article-pdf/14/2/160/
697993/jcws_r_00228.pdf.

Kalinovsky:2014:SRB

[Kal14] Artemy M. Kalinovsky. Strange Rebels: 1979 and the Birth of the 21st Century by Christian Caryl (review). Journal of Cold War Studies, 16(4):248–250, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566669;   http://muse.jhu.edu/article/566669/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/248/698625/
jcws_r_00489.pdf.

Kalinovsky:2016:DIT

[Kal16] Artemy M. Kalinovsky. Daniel Immerwahr, Thinking Small: The United States and the Lure of Community Development. Cambridge, MA: Harvard University Press, 2015. 303 pp. Journal of Cold War Studies, 18(3):204–206, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632291;   http://muse.jhu.edu/article/632291/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/204/698973/
jcws_r_00643.pdf.

Kalinovsky:2018:NJC

[Kal18] Artemy Kalinovsky. Nathan J. Citino, Envisioning the Arab Future: Modernization in U.S.–Arab Relations, 1945–1967. Journal of Cold War Studies, 20(4):262–264, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718221;   https://direct.mit.
edu/jcws/article-pdf/20/4/262/1969622/jcws_r_00847.pdf.

Kalicki:2022:BRF

[Kal22] Jan H. Kalicki. Book review: FREEZE! The Grassroots Movement to Halt the Arms Race and End the Cold War by Henry Richard Maar III. Journal of Cold War Studies, 24(4):217–218, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/4/
217/2062876/jcws_r_01107.pdf.

Kampmark:2013:CJL

[Kam13] Binoy Kampmark. Commentary on John Lewis Gaddis, George F. Kennan: an American life. Journal of Cold War Studies, 15(4):216–224, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/534422;   http://muse.jhu.edu/article/534422/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/4/216/698339/

jcws_a_00405.pdf. See [Kra13b, Gad13].

Kami:2017:LDW

[Kam17] Hideaki Kami. The limits of dialogue: Washington, Havana, and Miami, 1977–1980. Journal of Cold War Studies, 19(3):4– 41, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/668157;
http://muse.jhu.edu/article/668157/pdf;   https://direct.
mit.edu/jcws/article-pdf/19/3/4/699212/jcws_a_00753.pdf.

Kanet:2010:CSK

[Kan10] Roger E. Kanet. Christian Saehrendt, Kunst als Botschafter einer künstlichen Nation: Studien zur Rolle der bildenden Kunst in der auswärtigen Kulturpolitik der DDR. (German) [Art as the Ambassador of an Artificial Nation: Studies on the Role of the Visual Arts in the Foreign Culture Policy of the GDR]. Journal of Cold War Studies, 12(3):141–142, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392787;   http://muse.jhu.edu/article/392787/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/141/697602/
jcws_r_00014.pdf.

Kanet:2012:AKS

[Kan12a] Roger E. Kanet. Artemy Kalinovsky and Sergey Radchenko, eds., The End of the Cold War and the Third World: New Perspectives on Regional Conflict. New York: Routledge, 2011. xii + 315 pp. $138.00. Journal of Cold War Studies, 14(4):248–250, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/500347;   http:
//muse.jhu.edu/article/500347/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/4/248/698183/jcws_r_00294.pdf.

Kanet:2012:SMD

[Kan12b] Roger E. Kanet. Sergey Mazov, A Distant Front in the Cold War: The USSR in West Africa and the Congo, 1956–1964. Washington, DC: Woodrow Wilson Press, 2010. xiv, 334 pp. $55.00. Journal of Cold War Studies, 14(3):218–220, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/490192;   http://muse.jhu.edu/article/490192/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/3/218/698086/
jcws_r_00265.pdf.

Kanet:2014:WSV

[Kan14] Roger E. Kanet. Washingtons Söldner: Verdeckte US-Interventionen im Kalten Krieg und ihre Folgen by Klaas Voß (review). Journal of Cold War Studies, 16(4):252–253, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566671;   http://muse.jhu.edu/article/566671/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/252/698662/
jcws_r_00509.pdf.

Kanet:2015:LNB

[Kan15] Roger E. Kanet. Lise Namikas, Battleground Africa: Cold War in the Congo, 1960–1965. Washington, DC: Woodrow Wilson Center Press, 2013. xiv + 352 pp. $60.00 cloth, $29.95 paper. Journal of Cold War Studies, 17(4):214–215, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606045;   http://muse.jhu.edu/article/606045/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/214/698821/
jcws_r_00613.pdf.

Kaplan:2004:NED

[Kap04] Lawrence S. Kaplan. NATO Enlargement during the Cold War: Strategy and System in the Western Alliance (review). Journal of Cold War Studies, 6(2):75–77, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/53676;   http://muse.jhu.edu/
article/53676/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/2/75/695989/jcws.2004.6.2.75.pdf.

Kaple:2010:SSS

[Kap10] Deborah Kaple. The Sino–Soviet Split: Cold War in the Communist World (review). Journal of Cold War Studies, 12(1):143–146, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/382489;   http://
muse.jhu.edu/article/382489/pdf.

Kaple:2012:C

[Kap12] Deborah Kaple. Commentary. Journal of Cold War Studies, 14(1): 103–106, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/467683;
http://muse.jhu.edu/article/467683/pdf.

Kaplan:2014:GAR

[Kap14] Lawrence S. Kaplan. Greek–American Relations from Monroe to Truman by Angelo Repousis (review). Journal of Cold War Studies, 16(4):287–289, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566688;   http://muse.jhu.edu/article/566688/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/287/698680/
jcws_r_00507.pdf.

Kaple:2016:ACS

[Kap16] Deborah Kaple. Agents of change: Soviet advisers and high Stalinist management in China, 1949–1960. Journal of Cold War Studies, 18 (1):5–30, Winter 2016. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/612004;
http://muse.jhu.edu/article/612004/pdf;   https://direct.
mit.edu/jcws/article-pdf/18/1/5/698755/jcws_a_00617.pdf.

Kaple:2017:MHO

[Kap17a] Deborah Kaple. Mark Harrison, One Day We Will Live without Fear: Everyday Lives under the Police State. Stanford, CA: Hoover Institution Press, 2016. 280 pp. $24.95. Journal of Cold War Studies, 19(2): 217–220, Spring 2017. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://www.mitpressjournals.org/doi/
full/10.1162/JCWS_r_00726;   https://direct.mit.edu/jcws/
article-pdf/19/2/217/699230/jcws_r_00726.pdf.

Kapstein:2017:SFC

[Kap17b] Ethan B. Kapstein. Success and failure in counterinsurgency campaigns: Lessons from the Cold War. Journal of Cold War Studies, 19(1):125–159, Winter 2017. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://www.mitpressjournals.org/
doi/full/10.1162/JCWS_a_00720;   https://direct.mit.edu/
jcws/article-pdf/19/1/125/699104/jcws_a_00720.pdf.

Kapstein:2020:PEI

[Kap20] Ethan B. Kapstein. Private enterprise, international development, and the Cold War. Journal of Cold War Studies, 22(4):113–145, Fall 2020.

CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/22/4/113/1860613/
jcws_a_00967.pdf;   https://muse.jhu.edu/article/776961.

Karsh:2004:AWM

[Kar04] Efraim Karsh. Arabs at War: Military Effectiveness, 1948–1991 (review). Journal of Cold War Studies, 6(3):144–145, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171163;   http:
//muse.jhu.edu/article/171163/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/3/144/696037/jcws.2004.6.3.144.pdf.

Karner:2014:WSB

[Kar14a] Stefan Karner. Wieviel Sicherheit braucht der Friede? Zivile und militärische Näherungen zur österreichischen Sicherheitsstrategie edited by Thomas Roithner, Johann Frank, and Eva Huber (review). Journal of Cold War Studies, 16(4):235–237, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/566663;   http:
//muse.jhu.edu/article/566663/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/4/235/698582/jcws_r_00483.pdf.

Karpozilos:2014:DGC

[Kar14b] Kostis Karpozilos. The defeated of the Greek civil war: From fighters to political refugees in the Cold War. Journal of Cold War Studies, 16 (3):62–87, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/556076;
http://muse.jhu.edu/article/556076/pdf;   https://direct.
mit.edu/jcws/article-pdf/16/3/62/698493/jcws_a_00471.pdf.

Karcic:2015:DBP

[Kar15a] Hamza Karčić. Džemal Bijedić: Politička biografija by Husnija Kamberović (review). Journal of Cold War Studies, 17(3):270–271, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595099;   http:
//muse.jhu.edu/article/595099/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/270/1102418/jcws_r_00511.pdf.

Karlsgodt:2015:HMN

[Kar15b] Elizabeth Campbell Karlsgodt. Holocaust Monuments and National Memory Cultures in France and Germany since 1989: The Origins and Political Function of the Vél d’Hiv’ in Paris and the Holocaust Monument in Berlin by Peter Carrier (review). Journal of Cold War Studies, 17(3):260–262, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595094;   http://muse.jhu.edu/article/595094/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/260/
1102484/jcws_r_00592.pdf.

Karner:2015:VKI

[Kar15c] Stefan Karner. Vasilii S. Khristoforov, Istoriya strany v dokumentakh arkhivov FSB Rossii: Sbornik stat’ei i materialov [The country’s history in documents from the archives of Russia’s FSB: a collection of articles and materials]. Journal of Cold War Studies, 17(3):253–255, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595091;   http:
//muse.jhu.edu/article/595091/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/253/1102490/jcws_r_00580.pdf.

Kassianova:2016:NPM

[Kas16] Alla Kassianova. Norman Polmar and Michael White, Project Azorian: The CIA and the Raising of K-129. Annapolis: Naval Institute Press, 2010. 238 pp. Journal of Cold War Studies, 18(3):236–238, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632305;   http://muse.jhu.edu/article/632305/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/236/699076/
jcws_r_00671.pdf.

Kaufman:2002:AAM

[Kau02] Robert G. Kaufman. Annelise Anderson, and Martin Anderson, eds., Reagan: In His Own Hand: The Writings of Ronald Reagan That Reveal His Revolutionary Vision for America (review). Journal of Cold War Studies, 4(4):112–114, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9300;   http://muse.jhu.edu/article/9300/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/4/112/695692/
jcws.2002.4.4.112.pdf.

Kaufman:2003:AEB

[Kau03] Victor S. Kaufman. Acheson and Empire: The British Accent in American Foreign Policy (review). Journal of Cold War Studies, 5(2):89–91, Spring 2003. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
40763;   http://muse.jhu.edu/article/40763/pdf;   https:/

/direct.mit.edu/jcws/article-pdf/5/2/89/700466/jcws.2003.
5.2.89.pdf.

Kaufman:2009:DCW

[Kau09] Robert Kaufman. David C. Wills, The First War on Terrorism: Counter-Terrorism Policy during the Reagan Administration. Journal of Cold War Studies, 11(2):164–166, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/265861;   http://muse.jhu.edu/article/265861/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/2/164/
697209/jcws.2009.11.2.164.pdf.

Kazmierczak:2009:CNW

[Kaź09] Janusz Kaźmierczak. The community that never was: The European Defense Community and its image in Polish visual propaganda of the 1950s. Journal of Cold War Studies, 11(4):118– 141, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/366248;
http://muse.jhu.edu/article/366248/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/4/118/1627619/jcws.2009.11.4.
118.pdf.

Kazmierczak:2019:BIO

[Kaź19] Janusz Kaźmierczak. Bojkot igrzysk olimpijskich jako instrument polityki miedzynarodowej w latach 1976–1988 by Michał Słoniewski (review). Journal of Cold War Studies, 21(2):201–203, Spring 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/726142;   https://direct.mit.
edu/jcws/article-pdf/21/2/201/699549/jcws_r_00887.pdf.

Kraus:2017:CWE

[KCG^{+}17] Michael Kraus, Anna M. Cienciala, Margaret K. Gnoinska, Douglas Selvage, Molly Pucci, Erik Kulavig, Constantine Pleshakov, A. Ross Johnson, Mark Kramer, and Vít Smetana. The Cold War and East– Central Europe, 1945–1989. Journal of Cold War Studies, 19(2): 158–214, Spring 2017. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://www.mitpressjournals.org/doi/
full/10.1162/JCWS_c_00723;   https://direct.mit.edu/jcws/
article-pdf/19/2/158/699210/jcws_c_00723.pdf.

Keeler:2000:RAM

[Kee00] John T. S. Keeler. A response to Andrew Moravcsik. Journal of Cold War Studies, 2(3):74–76, Fall 2000. CODEN ????

ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9157;   http://muse.jhu.edu/article/
9157/pdf;   https://direct.mit.edu/jcws/article-pdf/2/3/
74/695274/15203970051032219.pdf. See [Mor00b, Mor00c].

Kelley:2002:MIH

[Kel02] Col. Lawrence G. Kelley. Military Intelligence: a History (review). Journal of Cold War Studies, 4(1):104–106, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9229;   http://muse.jhu.edu/article/
9229/pdf;   https://direct.mit.edu/jcws/article-pdf/4/1/
104/695516/jcws.2002.4.1.104.pdf.

Keller:2019:RWT

[Kel19] Renata Keller. The revolution will be teletyped: Cuba’s Prensa Latina news agency and the Cold War contest over information. Journal of Cold War Studies, 21(3):88–113, Summer 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/731284;   https://direct.
mit.edu/jcws/article-pdf/21/3/88/699624/jcws_a_00895.pdf.

Kenney:2003:RNE

[Ken03] Padraic Kenney. Redrawing Nations: Ethnic Cleansing in East– Central Europe, 1944–1948 (review). Journal of Cold War Studies, 5(3):105–107, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/45201;   http://muse.jhu.edu/article/45201/pdf.

Kennedy:2005:SSA

[Ken05] David M. Kennedy. Surprise, Security, and the American Experience (review). Journal of Cold War Studies, 7(1):182–184, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/179565;
http://muse.jhu.edu/article/179565/pdf;   https://direct.
mit.edu/jcws/article-pdf/7/1/182/700464/1520397053326211.
pdf.

Kenney:2007:ROD

[Ken07] Padraic Kenney. Robin Okey, The Demise of Communist East Europe: 1989 in Context. Journal of Cold War Studies, 9 (3):160–161, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219458;   http://muse.jhu.edu/article/219458/pdf;

 https://direct.mit.edu/jcws/article-pdf/9/3/160/696642/
jcws.2007.9.3.160.pdf.

Kenez:2013:MPC

[Ken13] Peter Kenez. Mária Palasik, Chess Game for Democracy: Hungary between East and West: 1944–1947. Montreal: McGill–Queen’s University Press, 2011. vii + 230 pp. $32.95. Journal of Cold War Studies, 15(3):214–216, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522731;   http://muse.jhu.edu/article/522731/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/214/698353/
jcws_r_00386.pdf.

Kern:2015:FSB

[Ker15] Gary Kern. Father, son, and the Bomb. Journal of Cold War Studies, 17(4):166–174, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606025;   http://muse.jhu.edu/article/606025/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/166/698744/
jcws_a_00601.pdf.

Kessler:2007:KCB

[Kes07] Gijs Kessler. Karel C. Berkhoff, Harvest of Despair: Life and Death in Ukraine under Nazi Rule. Journal of Cold War Studies, 9(1):153–155, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210085;   http://muse.jhu.edu/article/210085/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/153/696542/
jcws.2007.9.1.153.pdf.

Ketchian:2014:JEC

[Ket14] Sonia I. Ketchian. Jacob Edmond, A Common Strangeness: Contemporary Poetry, Cross-Cultural Encounter, Comparative Literature. New York: Fordham University Press, 2012. xv + 272 pp. $20.00. Journal of Cold War Studies, 16(1):242–244, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543746;   http:
//muse.jhu.edu/article/543746/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/1/242/698979/jcws_r_00437.pdf.

Keys:2016:RAN

[Key16] Barbara Keys. Roham Alvandi, Nixon, Kissinger, and the Shah: The United States and Iran in the Cold War. Oxford, UK: Oxford University Press, 2014. 255 pp. $55.00. Journal of Cold War Studies, 18(3):240–241, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632307;   http://muse.jhu.edu/article/632307/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/240/699080/
jcws_r_00673.pdf.

Kagan-Guthrie:2011:SOE

[KG11] Zachary Kagan-Guthrie. Sue Onslow, ed., Cold War in Southern Africa: White Power, Black Liberation. New York: Routledge, 2009. 253 pp. $125.00. Journal of Cold War Studies, 13(3):244–246, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447489;   http://muse.jhu.edu/article/447489/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/244/700734/
jcws_r_00122.pdf.

Klehr:2023:FUB

[KH23] Harvey Klehr and John Earl Haynes. The first U.S.-based Soviet nuclear spy: The saga of Clarence Hiskey and Arthur Adams. Journal of Cold War Studies, 25(4):53–69, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/4/53/2204977/jcws_a_01170.pdf.

Khoo:2010:BRE

[Kho10] Nicholas Khoo. Breaking the ring of encirclement: The Sino–Soviet Rift and Chinese policy toward Vietnam, 1964–1968. Journal of Cold War Studies, 12(1):3–42, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/382485;   http://muse.jhu.edu/article/382485/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/1/3/1627467/
jcws.2010.12.1.3.pdf.

Khoo:2014:INM

[Kho14] Nicholas Khoo. Ian Nish and Mark Allen, eds., War, Conflict and Security in Japan and Asia–Pacific, 1941–52: The Writings of Louis Allen. Journal of Cold War Studies, 16(3):220–221, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/556085;   http:
//muse.jhu.edu/article/556085/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/3/220/698517/jcws_r_00461.pdf.

Kimball:2007:GPP

[Kim07] Jeffrey P. Kimball. Gareth Porter, Perils of Dominance: Imbalance of Power and the Road to War in Vietnam. Berkeley: University of California Press, 2005. 403 pp. $27.50. Journal of Cold War Studies, 9(2):164–166, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214743;   http://muse.jhu.edu/article/214743/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/164/696451/
jcws.2007.9.2.164.pdf.

Kimball:2011:MPY

[Kim11] Warren F. Kimball. S. M. Plokhy, Yalta: The Price of Peace. New York: Viking, 2010. 451 pp. $29.95. Journal of Cold War Studies, 13(3):205–207, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447469;   http://muse.jhu.edu/article/447469/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/205/700726/
jcws_r_00148.pdf.

Kim:2016:LRS

[Kim16] Inhan Kim. Land reform in south Korea under the U.S. military occupation, 1945–1948. Journal of Cold War Studies, 18(2):97– 129, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/620548;
http://muse.jhu.edu/article/620548/pdf;   https://direct.
mit.edu/jcws/article-pdf/18/2/97/698844/jcws_a_00639.pdf.

Kipp:2003:MSW

[Kip03] Jacob W. Kipp. Making Sense of War: The Second World War and the Fall of the Bolshevik Revolution (review). Journal of Cold War Studies, 5(2):122–124, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40764;   http://muse.jhu.edu/
article/40764/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/122/700447/jcws.2003.5.2.122.pdf.

Kipp:2015:RBN

[Kip15] Jacob W. Kipp. Richard Bidlack and Nikita Lomagin, eds., The Leningrad Blockade, 1941–1944: a New Documentary History from the Soviet Archives, trans. of documents by Marian Schwartz. New Haven: Yale University Press, 2014. xxix + 486 pp. $75.00. Journal of Cold War Studies, 17(4):195–197, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606035;   http://muse.jhu.edu/article/606035/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/195/698773/
jcws_r_00603.pdf.

Kirby:2008:MTC

[Kir08] Dianne Kirby. Merrilyn Thomas, Communing with the Enemy: Covert Operations, Christianity and Cold War Politics in Britain and the GDR. Journal of Cold War Studies, 10(4):151– 153, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/252659;
http://muse.jhu.edu/article/252659/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/4/151/697041/jcws.2008.10.4.
151.pdf.

Kirby:2009:MAM

[Kir09] Dianne Kirby. Michael Alexander, Managing the Cold War: a View from the Frontline. Journal of Cold War Studies, 11(2):156– 157, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/265865;
http://muse.jhu.edu/article/265865/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/2/156/697195/jcws.2009.11.2.
156.pdf.

Kirby:2016:DCW

[Kir16] Dianne Kirby. Daniel C. Williamson, Separate Agendas: Churchill, Eisenhower, and Anglo-American Relations, 1953–1955. New York: Rowman & Littlefield, 2006. 145 pp. Journal of Cold War Studies, 18(3):242–243, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632308;   http://muse.jhu.edu/article/632308/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/242/699085/
jcws_r_00674.pdf.

Kisatsky:2006:JHC

[Kis06] Deborah Kisatsky. James H. Critchfield, Partners at the Creation: The Men behind Postwar Germany’s Defense and Intelligence Establishments. Journal of Cold War Studies, 8(2):154–156, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/196282;   http:
//muse.jhu.edu/article/196282/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/2/154/700548/jcws.2006.8.2.154.pdf.

Klehr:2004:DDS

[Kle04] Harvey Klehr. The Dancer Defects: The Struggle for Cultural Supremacy during the Cold War (review). Journal of Cold War Studies, 6(4):151–153, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/174710;   http://muse.jhu.edu/article/174710/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/4/151/696087/
jcws.2004.6.4.151.pdf.

Klehr:2008:SUE

[Kle08] Harvey Klehr. Steven Usdin, Engineering Communism: How Two Americans Spied for Stalin and Founded the Soviet Silicon Valley. Journal of Cold War Studies, 10(2):155–157, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/237161;   http:
//muse.jhu.edu/article/237161/pdf;   https://direct.mit.
edu/jcws/article-pdf/10/2/155/697005/jcws.2008.10.2.155.
pdf.

Klehr:2011:RRG

[Kle11] Harvey Klehr. R. S. Rose and Gordon Scott, Johnny: a Spy’s Life. University Park, PA: Pennsylvania State University Press, 2010. 512 pp. $45.00. Journal of Cold War Studies, 13(3):227–229, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/447480;   http:
//muse.jhu.edu/article/447480/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/3/227/700708/jcws_r_00113.pdf.

Klehr:2016:CBI

[Kle16] Harvey Klehr. Colin Burke, Information and Intrigue: From Index Cards to Dewey Decimals to Alger Hiss. Cambridge, MA: MIT Press, 2014. 370 pp. Journal of Cold War Studies, 18(4):233–235, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/645924;   http:
//muse.jhu.edu/article/645924/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/4/233/700122/jcws_r_00696.pdf.

Klehr:2022:TAF

[Kle22] Harvey Klehr. Teaching Anticommunism: Fred Schwarz and American Postwar Conservatism by Hubert Villeneuve (review). Journal of Cold War Studies, 24(1):252–254, Winter 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/24/1/252/1980869/jcws_r_
01062.pdf;   https://muse.jhu.edu/article/843084.

Klehr:2023:KMC

[Kle23] Harvey Klehr. Book review: KGB Man: The Cold War’s Most Notorious Soviet Agent and the First to Be Exchanged at the Bridge of Spies by Cecil Kuhne. Journal of Cold War Studies, 25(2):257–259, Spring 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/2/
257/2140958/jcws_r_01151.pdf.

Kramer:2014:CEV

[KLHK14] Mark Kramer, David Lowenthal, John Earl Haynes, and Harvey Klehr. Correspondence: Exchange on Vassiliev notebooks and Alger Hiss. Journal of Cold War Studies, 16(1):200–209, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543730;   http:
//muse.jhu.edu/article/543730/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/1/200/698947/jcws_c_00436.pdf.

Klinghoffer:2002:ISR

[Kli02] Arthur Jay Klinghoffer. Israeli–Soviet Relations 1953–1967: From Confrontation to Disruption (review). Journal of Cold War Studies, 4(2):131–133, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9252;   http://muse.jhu.edu/article/9252/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/2/131/695577/
jcws.2002.4.2.131.pdf.

Klinghoffer:2005:SFK

[Kli05] Arthur Jay Klinghoffer. Support Any Friend: Kennedy’s Middle East and the Making of the U.S.–Israel Alliance (review). Journal of Cold War Studies, 7(4):155–157, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188267;   http://muse.jhu.edu/article/188267/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/155/696321/
jcws.2005.7.4.155.pdf.

Klimentov:2022:CMU

[Kli22] Vassily A. Klimentov. “Communist Muslims”: The USSR and the People’s Democratic Party of Afghanistan’s conversion to Islam, 1978– 1988. Journal of Cold War Studies, 24(1):4–38, Winter 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/24/1/4/1980879/jcws_a_
01055.pdf;   https://muse.jhu.edu/article/843076.

King:2022:JSC

[KM22] Amy King and Sherzod Muminov. “Japan still has cadres remaining”: Japanese in the USSR and Mainland China, 1945– 1956. Journal of Cold War Studies, 24(3):200–230, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/3/
200/2042024/jcws_a_01093.pdf.

Knaus:2002:CCW

[Kna02] John Kenneth Knaus. The Clandestine Cold War in Asia, 1945–65: Western Intelligence, Propaganda, and Special Operations (review). Journal of Cold War Studies, 4(1):92–95, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9230;   http:/
/muse.jhu.edu/article/9230/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/1/92/695555/jcws.2002.4.1.92.pdf.

Knaus:2003:OPC

[Kna03] John Kenneth Knaus. Official policies and covert programs: The U.S. State Department, the CIA, and the Tibetan Resistance. Journal of Cold War Studies, 5(3):54–79, Summer 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45202;   http://muse.jhu.edu/
article/45202/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/54/695814/152039703322286773.pdf.

Knapp:2005:BCM

[Kna05] Wilfrid Knapp. Britain and the Conflict in the Middle East, 1964– 1967: The Coming of the Six-Day War (review). Journal of Cold War Studies, 7(4):158–159, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188268;   http://muse.jhu.edu/article/188268/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/158/696289/
jcws.2005.7.4.158.pdf.

Knight:2003:KPC

[Kni03] Amy W. Knight. The KGB, perestroika, and the collapse of the Soviet Union. Journal of Cold War Studies, 5(1):67–93, Winter 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/38254;   http://muse.jhu.edu/
article/38254/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/1/67/695767/152039703320996722.pdf.

Knopf:2006:TWC

[Kno06] Jeffrey W. Knopf. Timothy W. Crawford, Pivotal Deterrence: Third-Party Statecraft and the Pursuit of Peace. Journal of Cold War Studies, 8(2):126–128, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196283;   http://muse.jhu.edu/article/196283/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/126/700506/
jcws.2006.8.2.126.pdf.

Knopf:2020:ALA

[Kno20] Jeffrey W. Knopf. Alexander Lanoszka, Atomic Assurance: The Alliance Politics of Nuclear Proliferation. Journal of Cold War Studies, 22(4):239–242, Fall 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/4/239/1860642/jcws_r_00976.pdf;
https://muse.jhu.edu/article/776966.

Kraft:2018:SIP

[KNS18] Alison Kraft, Holger Nehring, and Carola Sachse. Special issue: The Pugwash Conferences and the global Cold War: scientists, transnational networks, and the complexity of nuclear histories: Introduction. Journal of Cold War Studies, 20(1):4–30, Winter 2018. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/691016;
http://muse.jhu.edu/article/691016/pdf;   https://direct.
mit.edu/jcws/article-pdf/20/1/4/699355/jcws_e_00799.pdf.

Knusel:2020:SCI

[Knü20] Ariane Knüsel. Swiss counter intelligence and Chinese espionage during the Cold War. Journal of Cold War Studies, 22(3):4–31, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/3/4/1860587/jcws_a_00949.pdf;   https://muse.jhu.edu/
article/762890.

Kober:2006:GPI

[Kob06] Avi Kober. Great-power involvement and Israeli battlefield success in the Arab–Israeli wars, 1948–1982. Journal of Cold War Studies, 8 (1):20–48, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/192211;
http://muse.jhu.edu/article/192211/pdf;   https://direct.
mit.edu/jcws/article-pdf/8/1/20/696263/152039706775212058.
pdf.

Koch:2010:JMB

[Koc10] James V. Koch. James MacGregor Burns. Running Alone: Presidential Leadership from JFK to Bush II: Why It Has Failed and How We Can Fix It. Journal of Cold War Studies, 12(3):131–132, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/392784;   http:
//muse.jhu.edu/article/392784/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/3/131/697616/jcws_r_00011.pdf.

Kolodziej:2012:CGG

[Kol12] Edward A. Kolodziej. Commentary on globalizing de Gaulle. Journal of Cold War Studies, 14(1):78–80, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/467680;   http://muse.jhu.edu/article/467680/
pdf;   https://direct.mit.edu/jcws/article-pdf/14/1/78/697839/
jcws_a_00193.pdf. See [Loc05].

Kolodziej:2015:BMR

[Kol15] Edward Kolodziej. Benjamin M. Rowland, ed., Charles De Gaulle’s Legacy of Ideas. Lanham, MD: Lexington Books, 2011. Journal of Cold War Studies, 17(4):193–195, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606034;   http://muse.jhu.edu/article/606034/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/193/698792/
jcws_r_00602.pdf.

Komine:2014:WRJ

[Kom14] Yukinori Komine. Whither a “resurgent Japan”: The Nixon doctrine and Japan’s defense buildup, 1969–1976. Journal of Cold War Studies, 16(3):88–128, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/556077;   http://muse.jhu.edu/article/556077/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/3/88/698469/
jcws_a_00447.pdf.

Komska:2018:RRB

[Kom18] Yuliya Komska. RFE/RL broadcasting and West German society: Caught between nature protection activism and anti-Americanism. Journal of Cold War Studies, 20(3):180–206, Summer 2018. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/704574;   https://direct.mit.
edu/jcws/article-pdf/20/3/180/699622/jcws_a_00824.pdf.

Kopstein:2016:JHU

[Kop16] Jeffrey Kopstein. Jeffrey Herf, Undeclared Wars With Israel: East Germany and the West German Far Left, 1967–1989. New York: Cambridge University Press, 2016. 493 pp. $29.99. Journal of Cold War Studies, 18(4):217–219, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645917;   http://muse.jhu.edu/article/645917/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/217/700114/
jcws_r_00689.pdf.

Korbonski:2006:APS

[Kor06] Andrzej Korbonski. Andrzej Paczkowski, The Spring Will Be Ours — Poland and the Poles from Occupation to Freedom, trans. by Jane Cave. Journal of Cold War Studies, 8(4):160–161, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/204620;   http:
//muse.jhu.edu/article/204620/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/4/160/700799/jcws.2006.8.4.160.pdf.

Korbonski:2007:TCC

[Kor07] Andrzej Korbonski. Triggering Communism’s Collapse: Perceptions and Power in Poland’s Transition (review). Journal of Cold War Studies, 9(2):136–139, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214744;   http://muse.jhu.edu/article/214744/pdf.

Kort:2010:MHR

[Kor10] Michael Kort. Max Hastings, Retribution: The Battle for Japan, 1944–1945. Journal of Cold War Studies, 12(3):162–164, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/392798;   http:
//muse.jhu.edu/article/392798/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/3/162/697648/jcws_r_00025.pdf.

Kosicki:2015:THM

[Kos15] Piotr H. Kosicki. The Trouble with History: Morality, Revolution, and Counterrevolution by Adam Michnik (review). Journal of Cold War Studies, 17(3):291–293, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595110;   http://muse.jhu.edu/article/595110/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/291/
1102468/jcws_r_00556.pdf.

Kott:2017:VDM

[Kot17] Matthew Kott. Violeta Davoliūtė, The Making and Breaking of Soviet Lithuania: Memory and Modernity in the Wake of War. New York: Routledge, 2013. 211 pp. Journal of Cold War Studies, 19(3):241–243, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668171;   http://muse.jhu.edu/article/668171/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/241/699283/
jcws_r_00731.pdf.

Kourkouvelas:2012:DNS

[Kou12] Lykourgos Kourkouvelas. Denuclearization on NATO’s Southern Front: Allied reactions to Soviet proposals, 1957–1963. Journal of Cold War Studies, 14(4):197–215, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500333;   http://muse.jhu.edu/article/500333/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/197/698110/
jcws_a_00280.pdf.

Koumas:2017:CWD

[Kou17] Manolis Koumas. Cold War dilemmas, superpower influence, and regional interests: Greece and the Palestinian question, 1947–1949. Journal of Cold War Studies, 19(1):99–124, Winter 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http:

//www.mitpressjournals.org/doi/full/10.1162/JCWS_a_00719;
 https://direct.mit.edu/jcws/article-pdf/19/1/99/699112/
jcws_a_00719.pdf.

Kozovoi:2014:DVS

[Koz14] Andrei Kozovoi. Dissonant voices: Soviet youth mobilization and the Cuban Missile Crisis. Journal of Cold War Studies, 16(3):29– 61, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/556075;

http://muse.jhu.edu/article/556075/pdf;   https://direct.
mit.edu/jcws/article-pdf/16/3/29/698465/jcws_a_00470.pdf.

Kramer:1999:EPSb

[Kra99a] Mark Kramer. The early post-Stalin succession struggle and upheavals in East–Central Europe: Inernal-external linkages in Soviet policy making (Part 2). Journal of Cold War Studies, 1(2):3–38, Spring 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/9103;   https://direct.
mit.edu/jcws/article-pdf/1/2/3/695129/15203979952559504.
pdf.

Kramer:1999:EPSa

[Kra99b] Mark Kramer. The early post-Stalin succession struggle and upheavals in East–Central Europe: Internal–extrenal linkages in Soviet policy making (Part 1). Journal of Cold War Studies, 1(1):3–55, Winter 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/9093;   https://direct.
mit.edu/jcws/article-pdf/1/1/3/695107/15203970152521881.
pdf.

Kramer:1999:EPSc

[Kra99c] Mark Kramer. The early post-Stalin succession struggle and upheavals in East central Europe: Internal-external linkages in Soviet policy making (Part 3). Journal of Cold War Studies, 1(3):3–66, Fall 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/9116;   https://direct.
mit.edu/jcws/article-pdf/1/3/3/695156/152039799316976805.
pdf.

Kramer:1999:ENa

[Kra99d] Mark Kramer. Editor’s note. Journal of Cold War Studies, 1(1):1–2, Winter 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/9091;   https://direct.mit.edu/jcws/article-pdf/
1/1/1/695105/15203970152521872.pdf.

Kramer:1999:ENb

[Kra99e] Mark Kramer. Editor’s note. Journal of Cold War Studies, 1(2):1–2, Spring 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/9102;   https://direct.mit.edu/jcws/article-pdf/
1/2/1/695135/15203979952559496.pdf.

Kramer:1999:ENc

[Kra99f] Mark Kramer. Editor’s note. Journal of Cold War Studies, 1(3):1–2, Fall 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/9115;   https://direct.mit.edu/jcws/article-pdf/
1/3/1/695146/152039799316976797.pdf.

Kramer:2000:MNA

[Kra00] Mark Kramer. Memorial notice: Adam Bruno Ulam (1922–2000). Journal of Cold War Studies, 2(2):130–132, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9142;   http://muse.jhu.edu/article/
9142/pdf;   https://direct.mit.edu/jcws/article-pdf/2/2/
130/695233/15203970051032345.pdf.

Kramer:2001:SKC

[Kra01] Mark Kramer. Die Sowjetunion und das kommunistische China 1945– 1950: Der beschwerliche Weg zum Bündnis (review). Journal of Cold War Studies, 3(2):107–111, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9190;   http://muse.jhu.edu/article/9190/pdf;
 https://direct.mit.edu/jcws/article-pdf/3/2/107/695390/
jcws.2001.3.2.107.pdf.

Kramer:2003:CEE

[Kra03a] Mark Kramer. The collapse of East European Communism and the repercussions within the Soviet Union (Part I). Journal of Cold War Studies, 5(4):178–256, Fall 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/47931;   http://muse.jhu.edu/
article/47931/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/4/178/700384/152039703322483783.pdf.

Kramer:2003:ENa

[Kra03b] Mark Kramer. Editor’s note. Journal of Cold War Studies, 5(1):1–2, Winter 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/38251;   http://
muse.jhu.edu/article/38251/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/1/1/695763/152039703320996696.pdf.

Kramer:2003:ENb

[Kra03c] Mark Kramer. Editor’s note. Journal of Cold War Studies, 5(2):1–2, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/40765;   http://
muse.jhu.edu/article/40765/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/2/1/700392/152039703763336444.pdf.

Kramer:2003:ENc

[Kra03d] Mark Kramer. Editor’s note. Journal of Cold War Studies, 5(3):1–4, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/45203;   http://
muse.jhu.edu/article/45203/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/3/1/695806/152039703322286737.pdf.

Kramer:2003:ENd

[Kra03e] Mark Kramer. Editor’s note. Journal of Cold War Studies, 5(4):1–2, Fall 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/47929;   http://
muse.jhu.edu/article/47929/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/4/1/700370/152039703322483738.pdf.

Kramer:2003:SICb

[Kra03f] Mark Kramer. Special issue: The collapse of the Soviet Union (Part 2): Introduction. Journal of Cold War Studies, 5(4):3–42, Fall 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/47930;   http://muse.jhu.edu/
article/47930/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/4/3/700378/152039703322483747.pdf.

Kramer:2003:SICa

[Kra03g] Mark Kramer. Special issue: The collapse of the Soviet Union (Part I): Introduction. Journal of Cold War Studies, 5(1):3–16, Winter 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/38255;   http://muse.jhu.edu/
article/38255/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/1/3/695761/152039703320996704.pdf.

Kramer:2004:WHS

[Kra04a] Mark Kramer. What Happened to the Soviet Union? How and Why American Sovietologists Were Caught by Surprise (review). Journal of Cold War Studies, 6(3):177–179, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171164;   http:
//muse.jhu.edu/article/171164/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/3/177/696121/jcws.2004.6.3.177.pdf.

Kramer:2004:CEE

[Kra04b] Mark Kramer. The collapse of East European Communism and the repercussions within the Soviet Union (Part 2). Journal of Cold War Studies, 6(4):3–64, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/174712;   http://muse.jhu.edu/article/174712/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/4/3/696079/1520397042350955.
pdf.

Kramer:2004:ENa

[Kra04c] Mark Kramer. Editor’s note. Journal of Cold War Studies, 6(2):1–2, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/53677;   http://
muse.jhu.edu/article/53677/pdf;   https://direct.mit.edu/
jcws/article-pdf/6/2/1/695953/152039704773254731.pdf.

Kramer:2004:ENb

[Kra04d] Mark Kramer. Editor’s note. Journal of Cold War Studies, 6(3):1–3, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171159;   http:
//muse.jhu.edu/article/171159/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/3/1/696031/1520397041447382.pdf.

Kramer:2004:ENc

[Kra04e] Mark Kramer. Editor’s note. Journal of Cold War Studies, 6(4):1–2, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/174711;   http:
//muse.jhu.edu/article/174711/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/4/1/696073/1520397042350928.pdf.

Kramer:2005:CEE

[Kra05a] Mark Kramer. The collapse of East European Communism and the repercussions within the Soviet Union (Part 3). Journal of Cold War Studies, 7(1):3–96, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/179566;   http://muse.jhu.edu/article/179566/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/1/3/700438/1520397053326185.
pdf.

Kramer:2005:ENa

[Kra05b] Mark Kramer. Editor’s note. Journal of Cold War Studies, 7(1):1–2, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/179558;   http:
//muse.jhu.edu/article/179558/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/1/1/700424/1520397053326149.pdf.

Kramer:2005:ENb

[Kra05c] Mark Kramer. Editor’s note. Journal of Cold War Studies, 7(2):1–2, Spring 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/181901;   http:
//muse.jhu.edu/article/181901/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/2/1/696200/1520397053630592.pdf.

Kramer:2005:ENc

[Kra05d] Mark Kramer. Editor’s note. Journal of Cold War Studies, 7(3): 1–2, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/185110;
http://muse.jhu.edu/article/185110/pdf;   https://direct.
mit.edu/jcws/article-pdf/7/3/1/1102569/1520397054377214.
pdf.

Kramer:2005:ENd

[Kra05e] Mark Kramer. Editor’s note. Journal of Cold War Studies, 7(4):1–2, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/188269;   http:
//muse.jhu.edu/article/188269/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/4/1/696273/1520397055012433.pdf.

Kramer:2006:SHH

[Kra06a] Mark Kramer. Stalin and His Hangmen: The Tyrant and Those Who Killed for Him (review). Journal of Cold War Studies, 8(1):122–127, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/192213;   http://muse.jhu.edu/article/192213/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/1/122/696297/
jcws.2006.8.1.122.pdf.

Kramer:2006:ENa

[Kra06b] Mark Kramer. Editor’s note. Journal of Cold War Studies, 8(1): 1–2, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/192212;
http://muse.jhu.edu/article/192212/pdf;   https://direct.
mit.edu/jcws/article-pdf/8/1/1/696242/152039706775212012.
pdf.

Kramer:2006:ENb

[Kra06c] Mark Kramer. Editor’s note. Journal of Cold War Studies, 8(2):1–2, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/196284;   http:
//muse.jhu.edu/article/196284/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/2/1/700504/jcws.2006.8.2.1.pdf.

Kramer:2006:ENc

[Kra06d] Mark Kramer. Editor’s note. Journal of Cold War Studies, 8(3):1–4, Summer 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/201237;   http:
//muse.jhu.edu/article/201237/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/3/1/696409/jcws.2006.8.3.1.pdf.

Kramer:2006:ENd

[Kra06e] Mark Kramer. Editor’s note. Journal of Cold War Studies, 8(4):1–2, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/204623;   http:
//muse.jhu.edu/article/204623/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/4/1/700801/jcws.2006.8.4.1.pdf.

Kramer:2006:GPR

[Kra06f] Mark Kramer. Great-power rivalries, Tibetan guerrilla resistance, and the Cold War in South Asia: Introduction. Journal of Cold War Studies, 8(3):5–14, Summer 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/201238;   http://muse.jhu.edu/article/201238/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/3/5/696417/jcws.
2006.8.3.5.pdf.

Kramer:2007:ENa

[Kra07a] Mark Kramer. Editor’s note. Journal of Cold War Studies, 9(1):1–2, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/210086;   http:
//muse.jhu.edu/article/210086/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/1/1/696469/jcws.2007.9.1.1.pdf.

Kramer:2007:ENb

[Kra07b] Mark Kramer. Editor’s note. Journal of Cold War Studies, 9(2):1–4, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/214745;   http:
//muse.jhu.edu/article/214745/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/2/1/696429/jcws.2007.9.2.1.pdf.

Kramer:2007:ENc

[Kra07c] Mark Kramer. Editor’s note. Journal of Cold War Studies, 9(3):1–4, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/219451;   http:
//muse.jhu.edu/article/219451/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/3/1/696632/jcws.2007.9.3.1.pdf.

Kramer:2007:ENd

[Kra07d] Mark Kramer. Editor’s note. Journal of Cold War Studies, 9(4):1–5, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/223548;   http:
//muse.jhu.edu/article/223548/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/4/1/696725/jcws.2007.9.4.1.pdf.

Kramer:2008:EN

[Kra08a] Mark Kramer. Editor’s note. Journal of Cold War Studies, 10(3): 1–2, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/241495;
http://muse.jhu.edu/article/241495/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/3/1/697081/jcws.2008.10.3.1.
pdf.

Krammer:2008:I

[Kra08b] Mark Krammer. Introduction. Journal of Cold War Studies, 10(2): 97–98, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/237153;
http://muse.jhu.edu/article/237153/pdf.

Kramer:2013:DRM

[Kra13a] Mark Kramer. The dynamics of 1989: Reassessing a momentous year. Journal of Cold War Studies, 15(4):148–152, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/534419;   http:
//muse.jhu.edu/article/534419/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/4/148/698319/jcws_a_00398.pdf.

Kramer:2013:FGF

[Kra13b] Mark Kramer. Forum: George F. Kennan and the Cold War: Perspectives on John Gaddis’s biography. Journal of Cold War Studies, 15(4):153–154, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/534420;   http://muse.jhu.edu/article/534420/pdf;

 https://direct.mit.edu/jcws/article-pdf/15/4/153/698367/

jcws_e_00427.pdf. See [Kur13, Kam13, Ste13, May13].

Kraft:2018:DSE

[Kra18] Alison Kraft. Dissenting scientists in early Cold War Britain: The “fallout” controversy and the origins of Pugwash, 1954–1957. Journal of Cold War Studies, 20(1):58–100, Winter 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/691018;   http://muse.jhu.edu/article/691018/
pdf;   https://direct.mit.edu/jcws/article-pdf/20/1/58/699323/
jcws_a_00801.pdf.

Kramer:2022:DSU

[Kra22] Mark Kramer. The dissolution of the Soviet Union: A case study of discontinuous change. Journal of Cold War Studies, 24(1): 188–218, Winter 2022. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL https://direct.mit.edu/jcws/
article-pdf/24/1/188/1980883/jcws_a_01059.pdf;   https://
muse.jhu.edu/article/843081.

Krementsov:2007:SBU

[Kre07] Nikolai L. Krementsov. In the shadow of the bomb: U.S.–Soviet biomedical relations in the early Cold War, 1944–1948. Journal of Cold War Studies, 9(4):41–67, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223549;   http://muse.jhu.edu/article/223549/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/41/696763/jcws.
2007.9.4.41.pdf.

Krenn:2012:RLKb

[Kre12a] Michael L. Krenn. Richard Lentz and Karla K. Gower, The Opinions of Mankind: Racial Issues, Press, and Propaganda in the Cold War. Columbia, MO: University of Missouri Press, 2010. x + 349 pp. $39.95. Journal of Cold War Studies, 14(3):190–192, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/490178;   http:
//muse.jhu.edu/article/490178/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/3/190/698051/jcws_r_00251.pdf.

Krenn:2012:RLKa

[Kre12b] Michael L. Krenn. Richard Lentz and Karla K. Gower, The Opinions of Mankind: Racial Issues, Press, and Propaganda in the Cold War. Columbia: University of Missouri Press, 2010. x + 349 pp. $39.95. Journal of Cold War Studies, 14(1):149–151, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/467702;   http:
//muse.jhu.edu/article/467702/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/1/149/697931/jcws_r_00213.pdf.

Krikorian:2004:CSS

[Kri04] Robert Owen Krikorian. Contending with Stalinism: Soviet Power and Popular Resistance in the 1930s (review). Journal of Cold War Studies, 6(4):177–179, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/174713;   http://muse.jhu.edu/article/174713/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/4/177/696133/
jcws.2004.6.4.177.pdf.

Krikorian:2005:RCS

[Kri05] Robert Owen Krikorian. Russia’s Carnival: The Smells, Sights, and Sounds of Transition (review). Journal of Cold War Studies, 7(3):190–192, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185111;   http://muse.jhu.edu/article/185111/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/190/1102603/
jcws.2005.7.3.190.pdf.

Kroll:2008:JKF

[Kro08] Heidi Kroll. János Kornai, By Force of Thought: Irregular Memoirs of an Intellectual Journey. Journal of Cold War Studies, 10(1):154–156, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231576;   http://muse.jhu.edu/article/231576/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/154/700083/
jcws.2008.10.1.154.pdf.

Kroll:2012:DPJ

[Kro12] Heidi Kroll. Dana Polan, Julia Child’s “The French Chef.” Durham, NC: Duke University Press, 2011. 312 pp. $23.95. Journal of Cold War Studies, 14(4):224–225, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500335;   http://muse.jhu.edu/article/500335/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/224/698128/
jcws_r_00282.pdf.

Krupar:2004:BWM

[Kru04a] Jason N. Krupar. For Better or for Worse: The Marriage of Science and Government in the United States (review). Journal of Cold War Studies, 6(1):101–103, Winter 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/51867;   http://muse.jhu.edu/
article/51867/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/1/101/695899/jcws.2004.6.1.101.pdf.

Krupar:2004:SBS

[Kru04b] Jason N. Krupar. Scientists, Business, and the State, 1890–1960 (review). Journal of Cold War Studies, 6(3):130–133, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171165;   http:
//muse.jhu.edu/article/171165/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/3/130/696025/jcws.2004.6.3.130.pdf.

Krupar:2005:MWA

[Kru05] Jason N. Krupar. Mao’s War against Nature: Politics and the Environment in Revolutionary China (review). Journal of Cold War Studies, 7(4):174–176, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188270;   http://muse.jhu.edu/article/188270/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/174/696351/
jcws.2005.7.4.174.pdf.

Krugler:2007:WPD

[Kru07a] David F. Krugler. Wilson P. Dizard, Jr., Inventing Public Diplomacy: The Story of the U.S. Information Agency. Boulder, CO: Lynne Rienner Publishers, 2004. 255 pp. $49.95. Journal of Cold War Studies, 9(2):158–161, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214746;   http://muse.jhu.edu/article/214746/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/158/696447/
jcws.2007.9.2.158.pdf.

Krupar:2007:JCE

[Kru07b] Jason N. Krupar. Jennet Conant, 109 East Palace: Robert Oppenheimer and the Secret City of Los Alamos. New York: Simon and Schuster, 2005. 425 pp. Journal of Cold War Studies, 9(4):143–145, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/223550;   http:

//muse.jhu.edu/article/223550/pdf;   http://muse.jhu.edu/
journals/journal_of_cold_war_studies/v009/9.4krupar.html;
 https://direct.mit.edu/jcws/article-pdf/9/4/143/696717/
jcws.2007.9.4.143.pdf.

Krugler:2008:KOT

[Kru08] David F. Krugler. Kenneth Osgood, Total Cold War: Eisenhower’s Secret Propaganda Battle at Home and Abroad. Journal of Cold War Studies, 10(3):149–153, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241508;   http://muse.jhu.edu/article/241508/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/149/697115/
jcws.2008.10.3.149.pdf.

Khoo:2018:UAE

[KSR18] Nicholas Khoo, Balázs Szalontai, and Sergey Radchenko. The USSR, Asia, and the end of the Cold War: Perspectives on unwanted visionaries. Journal of Cold War Studies, 20(4):212–226, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718204;   https://direct.mit.
edu/jcws/article-pdf/20/4/212/1969650/jcws_c_00842.pdf.

Kubik:2007:HME

[Kub07] Jan Kubik. Historical memory and the end of Communism. Journal of Cold War Studies, 9(2):127–133, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/214747;   http://muse.jhu.edu/article/214747/
pdf;   https://direct.mit.edu/jcws/article-pdf/9/2/127/696427/
jcws.2007.9.2.127.pdf.

Kube:2020:HKP

[Kub20] Sven Kube. Helma Kaldewey, A People’s Music: Jazz in East Germany, 1945–1990. Journal of Cold War Studies, 22(3):258– 259, Summer 2020. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL https://direct.mit.edu/jcws/
article-pdf/22/3/258/1860620/jcws_r_00957.pdf;   https://
muse.jhu.edu/article/762901.

Kuklick:2011:DCA

[Kuk11] Bruce Kuklick. David C. Atkinson, In Theory and in Practice: Harvard’s Center for International Affairs, 1958–1983. Cambridge, MA: Harvard University Press, 2007. 248 pp. Journal of Cold War Studies, 13(1):225–226, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/424247;   http://muse.jhu.edu/article/424247/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/1/225/697722/
jcws_r_00077.pdf.

Kuklick:2013:DFH

[Kuk13] Bruce Kuklick. Daniel F. Harrington, Berlin on the Brink: The Blockade, the Airlift, and the Early Cold War. Lexington, KY: University Press of Kentucky, 2012. 414 pp. $90.00. Journal of Cold War Studies, 15(1):164–166, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/505760;   http://muse.jhu.edu/article/505760/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/1/164/700008/
jcws_r_00319.pdf.

Kuklick:2021:AEM

[Kuk21] Bruce Kuklick. Armed with Expertise: The Militarization of American Social Research during the Cold War by Joy Rohde (review). Journal of Cold War Studies, 23(2):174–176, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/23/2/174/1919944/jcws_r_
01021.pdf;   https://muse.jhu.edu/article/794721.

Kulavig:2010:JAG

[Kul10] Erik Kulavig. J. Arch Getty and Oleg V. Naumov, Yezhov: The Rise of Stalin’s “Iron Fist”. Journal of Cold War Studies, 12(1): 202–204, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382502;
http://muse.jhu.edu/article/382502/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/1/202/1627453/jcws.2010.12.1.
202.pdf.

Kulavig:2013:JSM

[Kul13] Erik Kulavig. Jeremy Smith and Melanie Ilic, eds., Khrushchev in the Kremlin: Policy and Government in the Soviet Union 1953–1964. London: Routledge, 2011. 249 pp. $150.00. Journal of Cold War Studies, 15(1):178–180, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/505752;   http://muse.jhu.edu/article/505752/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/1/178/700024/
jcws_r_00327.pdf.

Kurilla:2005:FCW

[Kur05a] Ivan I. Kurilla. The First Cold War: The Legacy of Woodrow Wilson in U.S.–Soviet Relations (review). Journal of Cold War Studies, 7(3):192–195, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185112;   http://muse.jhu.edu/article/185112/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/192/1102577/
jcws.2005.7.3.192.pdf.

Kuromiya:2005:MWF

[Kur05b] Hiroaki Kuromiya. Making War, Forging Revolution: Russia’s Continuum of Crisis, 1914–1921 (review). Journal of Cold War Studies, 7(4):191–192, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188271;   http://muse.jhu.edu/article/188271/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/191/696341/
jcws.2005.7.4.191.pdf.

Kuromiya:2010:SCS

[Kur10] Hiroaki Kuromiya. Stefan Creuzberger, Stalin: Machtpolitiker und Ideologe. Journal of Cold War Studies, 12(4):202–203, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/403811;   http:
//muse.jhu.edu/article/403811/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/4/202/697564/jcws_r_00063.pdf.

Kuromiya:2011:AAE

[Kur11] Hiroaki Kuromiya. Anne Applebaum, ed., Gulag Voices: an Anthology. New Haven: Yale University Press, 2011. 195 pp. $25.00. Journal of Cold War Studies, 13(4):229–230, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/457206;   http://muse.jhu.edu/article/457206/
pdf;   https://direct.mit.edu/jcws/article-pdf/13/4/229/
697911/jcws_r_00186.pdf.

Kurilla:2013:AJL

[Kur13] Ivan Kurilla. An assessment of John Lewis Gaddis’s George F. Kennan: an American life. Journal of Cold War Studies, 15(4):189–195, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/534416;   http:
//muse.jhu.edu/article/534416/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/4/189/698343/jcws_a_00403.pdf. See

[Kra13b].

Kurilla:2014:AER

[Kur14] Ivan Kurilla. Americans Experience Russia: Encountering the Enigma, 1917 to the Present edited by Choi Chatterjee and Beth Holmgren (review). Journal of Cold War Studies, 16(4):277–279, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/566683;   http:
//muse.jhu.edu/article/566683/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/4/277/698659/jcws_r_00502.pdf.

Kuromiya:2015:SVP

[Kur15] Hiroaki Kuromiya. Stalin, Vol. 1: Paradoxes of Power, 1878–1928 by Stephen Kotkin (review). Journal of Cold War Studies, 17(3):245–247, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595087;   http:
//muse.jhu.edu/article/595087/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/245/1102478/jcws_r_00576.pdf.

Kurosaki:2018:JSC

[Kur18] Akira Kurosaki. Japanese scientists’ critique of nuclear deterrence theory and its influence on Pugwash, 1954–1964. Journal of Cold War Studies, 20(1):101–139, Winter 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/691019;   http://muse.jhu.edu/article/691019/pdf;
 https://direct.mit.edu/jcws/article-pdf/20/1/101/699311/
jcws_a_00802.pdf.

Kuromiya:2021:AES

[Kur21] Hiroaki Kuromiya. Aufstieg zur Energiemacht: Der sowjetische Weg ins Erdölzeitalter: 1930-er bis 1950-er Jahre by Felix Rehschuh (review). Journal of Cold War Studies, 23(2):171–172, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/23/2/171/1919935/
jcws_r_01019.pdf;   https://muse.jhu.edu/article/794722.

Kushner:2002:EAA

[Kus02] Marilyn S. Kushner. Exhibiting art at the American national exhibition in Moscow, 1959: Domestic politics and cultural diplomacy. Journal of Cold War Studies, 4(1):6–26, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9232;   http://muse.jhu.edu/article/
9232/pdf;   https://direct.mit.edu/jcws/article-pdf/4/1/
6/695497/152039702753344807.pdf.

Kusa:2003:RNE

[Kus03] Dagmar Kusa. Redrawing Nations: Ethnic Cleansing in East–Central Europe, 1944–1948 (review). Journal of Cold War Studies, 5(3):111– 114, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/45204;   http:
//muse.jhu.edu/article/45204/pdf.

Kuzmarov:2016:TCF

[Kuz16] Jeremy Kuzmarov. Thomas C. Field, Jr., From Development to Dictatorship: Bolivia and the Alliance for Progress. Ithaca, NY: Cornell University Press, 2014. 196 pp. $45.00. Journal of Cold War Studies, 18(4):245–247, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645929;   http://muse.jhu.edu/article/645929/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/245/700181/
jcws_r_00702.pdf.

Kuzio:2019:NAE

[Kuz19] Taras Kuzio. The Near Abroad: Eastern Europe and Soviet Patriotism in Ukraine, 1956–1985 by Zbigniew Wojnowski (review). Journal of Cold War Studies, 21(1):200–203, Winter 2019. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/723542;   https://direct.mit.
edu/jcws/article-pdf/21/1/200/699783/jcws_r_00863.pdf.

Kemp-Welch:2007:TCC

[KW07] A. Kemp-Welch. Triggering Communism’s Collapse: Perceptions and Power in Poland’s Transition (review). Journal of Cold War Studies, 9(2):134–136, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214742;   http://muse.jhu.edu/article/214742/pdf.

Kimball:2009:ENA

[KW09] Jeffrey P. Kimball and James J. Wirtz. Exchange on the Nixon Administration and the Vietnam War. Journal of Cold War Studies, 11 (4):142–144, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/366249;
http://muse.jhu.edu/article/366249/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/4/142/1627615/jcws.2009.11.4.
142.pdf.

Kemp-Welch:2022:PSM

[KW22] Anthony Kemp-Welch. Poland’s solidarity movement and the global politics of human rights. Journal of Cold War Studies, 24(3):258– 260, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/24/3/258/2042015/jcws_r_01102.pdf.

Kemp-Welch:2007:PTC

[KWKS07] Tony Kemp-Welch, Andrzej Korbonski, and Michael Szporer. Perspectives on triggering communism’s collapse. Journal of Cold War Studies, 9(2):134–143, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/9/2/134/696437/jcws.2007.9.2.134.pdf.

Laird:2008:JNR

[Lai08] Robbin F. Laird. Julie Newton, Russia, France, and the Idea of Europe. Journal of Cold War Studies, 10(1):138–141, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/231577;
http://muse.jhu.edu/article/231577/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/1/138/700069/jcws.2008.10.1.
138.pdf.

Lak:2011:PHC

[Lak11] Martijn Lak. Peter L. Hahn, Caught in the Middle East: U.S. Policy toward the Arab-Israeli Conflict, 1945–1961. Chapel Hill: University of North Carolina Press, 2006. 398 pp. $19.95, paper. Journal of Cold War Studies, 13(1):236–238, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/424253;   http://muse.jhu.edu/article/424253/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/1/236/697741/
jcws_r_00084.pdf.

Lankov:2001:DNC

[Lan01] Andrei N. Lankov. The demise of non-Communist parties in North Korea (1945–1960). Journal of Cold War Studies, 3(1):103–125, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9177;   http:/
/muse.jhu.edu/article/9177/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/1/103/695404/15203970151032164.pdf.

Langguth:2010:KCS

[Lan10] A. J. Langguth. Kathryn C. Statler, Replacing France: The Origins of American Intervention in Vietnam. Journal of Cold War Studies, 12(1):178–179, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/382503;   http://muse.jhu.edu/article/382503/
pdf;   https://direct.mit.edu/jcws/article-pdf/12/1/178/
1627413/jcws.2010.12.1.178.pdf.

Lankov:2020:TBN

[Lan20] Andrei Lankov. Trouble brewing: The North Korean famine of 1954–1955 and Soviet attitudes toward North Korea. Journal of Cold War Studies, 22(2):3–25, Spring 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.

mit.edu/jcws/article-pdf/22/2/3/1860544/jcws_a_00938.pdf;
 https://muse.jhu.edu/article/756823.

Larres:2002:GPT

[Lar02] Klaus Larres. The German Problem Transformed: Institutions, Politics, and Foreign Policy, 1945–1995 (review). Journal of Cold War Studies, 4(2):113–118, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9253;   http://muse.jhu.edu/article/9253/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/2/113/695522/
jcws.2002.4.2.113.pdf.

Large:2007:HRH

[Lar07a] David Clay Large. Helmut R. Hammerich, Jeder für sich und Amerika gegen alle? Die Lastenteilung der NATO am Bespiel des Temporary Council Committee 1949 bis 1954. Journal of Cold War Studies, 9(1):128–130, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210087;   http://muse.jhu.edu/article/210087/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/128/696523/
jcws.2007.9.1.128.pdf.

Larsen:2007:MDS

[Lar07b] Jeffrey Arthur Larsen. Mira Duric, The Strategic Defence Initiative: US Policy and the Soviet Union. Journal of Cold War Studies, 9(3):176–178, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219459;   http://muse.jhu.edu/article/219459/pdf;

 https://direct.mit.edu/jcws/article-pdf/9/3/176/696656/
jcws.2007.9.3.176.pdf.

Laron:2010:SBT

[Lar10] Guy Laron. Stepping back from the Third World: Soviet policy toward the United Arab Republic, 1965–1967. Journal of Cold War Studies, 12(4):99–118, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/403798;
http://muse.jhu.edu/article/403798/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/4/99/697532/jcws_a_00049.pdf.

Larson:2011:OCT

[Lar11] Deborah Welch Larson. The origins of commitment: Truman and West Berlin. Journal of Cold War Studies, 13(1):180–212, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/424244;   http:
//muse.jhu.edu/article/424244/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/1/180/697736/jcws_a_00074.pdf.

Laron:2017:PMC

[Lar17] Guy Laron. Philip Muehlenbeck, Czechoslovakia in Africa, 1945–1968. Journal of Cold War Studies, 19(4):258–259, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/683684;   http:
//muse.jhu.edu/article/683684/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/4/258/699449/jcws_r_00784.pdf.

Larson:2022:BRW

[Lar22] Deborah Welch Larson. Book review: The War of Nerves: Inside the Cold War Mind by Martin Sixsmith. Journal of Cold War Studies, 24 (4):233–236, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/24/4/233/2062886/jcws_r_01114.pdf.

Lasoen:2019:DSA

[Las19] Kenneth Lasoen. DDR Spionage: Von Albanien bis Grossbritannien ed. by Helmut Müller-Enbergs and Thomas Wegener Friis (review). Journal of Cold War Studies, 21(1):208–210, Winter 2019. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/723545;   https://direct.mit.
edu/jcws/article-pdf/21/1/208/699802/jcws_r_00866.pdf.

Latham:2008:KCS

[Lat08] Michael E. Latham. Kathryn C. Statler and Andrew L. Johns, eds., The Eisenhower Administration, the Third World, and the Globalization of the Cold War. Journal of Cold War Studies, 10(1):173–175, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231578;   http://muse.jhu.edu/article/231578/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/173/700162/
jcws.2008.10.1.173.pdf.

Latham:2014:PMR

[Lat14] Michael E. Latham. Peter Mandler, Return from the Natives: How Margaret Mead Won the Second World War and Lost the Cold War. New Haven: Yale University Press, 2013. 366 pp. $40.00. Journal of Cold War Studies, 16(1):253–255, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/543751;   http://muse.jhu.edu/article/543751/
pdf;   https://direct.mit.edu/jcws/article-pdf/16/1/253/
699000/jcws_r_00442.pdf.

Latham:2016:HKO

[Lat16a] Michael E. Latham. Heonik Kwan, The Other Cold War. New York: Columbia University Press, 2010. 232 pp. $50.00. Journal of Cold War Studies, 18(3):192–194, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632285;   http://muse.jhu.edu/article/632285/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/192/699057/
jcws_r_00659.pdf.

Latham:2016:JGT

[Lat16b] Michael E. Latham. John A. Gronbeck-Tedesco, Cuba, the United States, and the Cultures of the Transnational Left, 1930–1975. New York: Cambridge University Press, 2015. 294 pp. $99.99. Journal of Cold War Studies, 18(4):208–210, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645913;   http://muse.jhu.edu/article/645913/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/208/700039/
jcws_r_00685.pdf.

Laufer:2000:YBC

[Lau00] Jochen Laufer. From Yalta to Berlin: The Cold War Struggle over Germany (review). Journal of Cold War Studies, 2(2):118–120, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9143;   http:/
/muse.jhu.edu/article/9143/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/2/118/695265/jwcs.2000.2.2.118.pdf.

Lawrence:2001:CVW

[Law01] Mark A. Lawrence. China and the Vietnam Wars, 1950–1975 (review). Journal of Cold War Studies, 3(3):120–122, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9205;   http://muse.jhu.edu/article/
9205/pdf;   https://direct.mit.edu/jcws/article-pdf/3/3/
120/695484/jcws.2001.3.3.120.pdf.

Lawrence:2005:BPW

[Law05] Mark Atwood Lawrence. A Bitter Peace: Washington, Hanoi, and the Making of the Paris Agreement, and: a Companion to the Vietnam War (review). Journal of Cold War Studies, 7(3):173–175, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185113;   http://muse.jhu.edu/article/185113/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/173/1102585/
jcws.2005.7.3.173.pdf.

Lawrence:2007:CGI

[Law07] Mark Atwood Lawrence. Christoph Giebel, Imagined Ancestries of Vietnamese Communism: Ton Duc Thang and the Politics of History and Memory. Journal of Cold War Studies, 9 (3):210–212, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219460;   http://muse.jhu.edu/article/219460/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/210/696682/
jcws.2007.9.3.210.pdf.

Lawrence:2008:RJT

[Law08] Mark Atwood Lawrence. Robert J. Topmiller, The Lotus Unleashed: The Buddhist Peace Movement in South Vietnam, 1964–1966. Journal of Cold War Studies, 10(4):154–155, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/252662;   http://muse.jhu.edu/article/252662/
pdf;   https://direct.mit.edu/jcws/article-pdf/10/4/154/
697053/jcws.2008.10.4.154.pdf.

Lawrence:2010:GRH

[Law10] Mark Atwood Lawrence. Gary R. Hess, Vietnam: Explaining America’s Lost War. Journal of Cold War Studies, 12(3):160–162, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/392797;   http:
//muse.jhu.edu/article/392797/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/3/160/697690/jcws_r_00024.pdf.

Lawrence:2011:WFU

[Law11] Mark Atwood Lawrence. War without Fronts: The USA in Vietnam (review). Journal of Cold War Studies, 13(3):196–199, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/447462;   http://
muse.jhu.edu/article/447462/pdf.

Lawrence:2012:DHV

[Law12] Mark Atwood Lawrence. David Hunt, Vietnam’s Southern Revolution: From Peasant Insurrection to Total War. Amherst: University of Massachusetts Press, 2008. 272 pp. $28.95. Journal of Cold War Studies, 14(1):121–123, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467689;   http://muse.jhu.edu/article/467689/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/121/697870/
jcws_r_00200.pdf.

Lawrence:2022:BQN

[Law22] Mark Atwood Lawrence. Beyond the quagmire: New interpretations of the Vietnam War. Journal of Cold War Studies, 24(2):166–168, Spring 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/2/
166/2013010/jcws_r_01082.pdf.

Lazarev:2007:DEM

[Laz07] Valery Lazarev. David Engerman, Modernization from the Other Shore: American Intellectuals and the Romance of Russian Development. Journal of Cold War Studies, 9(1):160–162, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/210088;   http:
//muse.jhu.edu/article/210088/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/1/160/696556/jcws.2007.9.1.160.pdf.

Leary:2001:RCC

[Lea01] William M. Leary. Raiders of the China Coast: CIA Covert Operations During the Korean War (review). Journal of Cold War Studies, 3(1):143–144, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9178;   http://muse.jhu.edu/article/9178/pdf;
 https://direct.mit.edu/jcws/article-pdf/3/1/143/695422/
jcws.2001.3.1.143.pdf.

Lebow:2000:WWP

[Leb00] Richard Ned Lebow. Ways of War and Peace (review). Journal of Cold War Studies, 2(1):138–139, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9131;   http://muse.jhu.edu/article/9131/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/1/138/695205/
jcws.2000.2.1.138.pdf.

Lebow:2003:BAC

[Leb03] Richard Ned Lebow. Before and After the Cold War: Using Past Forecasts to Predict the Future (review). Journal of Cold War Studies, 5(2):103–104, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40766;   http://muse.jhu.edu/
article/40766/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/103/700404/jcws.2003.5.2.103.pdf.

Lebow:2006:PIC

[Leb06] Richard Ned Lebow. The Purpose of Intervention: Changing Beliefs about the Use of Force (review). Journal of Cold War Studies, 8(1):148–149, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/192214;   http://muse.jhu.edu/article/192214/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/1/148/696333/
jcws.2006.8.1.148.pdf.

Lee:2006:CAN

[Lee06a] Jongsoo Lee. Charles Armstrong, The North Korean Revolution, 1945–1950. Journal of Cold War Studies, 8(4):151–153, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/204624;   http:
//muse.jhu.edu/article/204624/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/4/151/700795/jcws.2006.8.4.151.pdf.

Leebaert:2006:WWL

[Lee06b] Derek Leebaert. Winning the World: Lessons for America’s Future from the Cold War (review). Journal of Cold War Studies, 8(1):136–138, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/192215;   http://muse.jhu.edu/article/192215/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/1/136/696301/
jcws.2006.8.1.136.pdf.

Lee:2012:CSB

[Lee12] Jongsoo James Lee. Chris Springer, with Balázs Szalontai, North Korea Caught in Time: Images of War and Reconstruction. Reading, UK: Garnet Publishing, 2010. 176 pp. $49.95. Journal of Cold War Studies, 14(2):162–164, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/478784;   http://muse.jhu.edu/article/478784/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/2/162/697995/
jcws_r_00229.pdf.

Lees:2017:ILD

[Lee17] Lorraine M. Lees. Ivan Laković and Dmitar Tasić, The Tito-Stalin Split and Yugoslavia’s Military Opening toward the West, 1950–1954: In NATO’s Backyard. Lanham, MD: Lexington Books, 2016. 285 pp. $95.00. Journal of Cold War Studies, 19(3):262–264, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/668181;   http:
//muse.jhu.edu/article/668181/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/3/262/699316/jcws_r_00749.pdf.

Lee:2020:MHG

[Lee20] Erika Lee. Madeline Y. Hsu, The Good Immigrants: How the Yellow Peril Became the Model Minority. Journal of Cold War Studies, 22(2):171–173, Spring 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/2/171/1860554/jcws_r_00885.pdf;
https://muse.jhu.edu/article/756831.

Lee:2021:ATW

[Lee21] Peter W. Lee. Against Their Will: The Secret History of Medical Experimentation on Children in Cold War America by Allen M. Hornblum, Judith L. Newman and Gregory J. Dober, New York: Palgrave Macmillan, 2013. 266 pp. $27.00. (review). Journal of Cold War Studies, 23(1):233–235, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/1/233/1911818/jcws_r_00989.pdf;
https://muse.jhu.edu/article/788391.

Lees:2023:BRU

[Lee23] Lorraine M. Lees. Book review: Unmaking Détente: Yugoslavia, the United States, and the Global Cold War, 1968–1980 by Milorad Lazic. Journal of Cold War Studies, 25(3):218–220, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/3/
218/2158658/jcws_r_01153.pdf.

Lefebvre:2002:CEM

[Lef02] Stéphane Lefebvre. Contre-espionnage: Mémoires d’un patron de la DST [Counterintelligence: Memoirs of a former head of the DST]. Journal of Cold War Studies, 4(3):134–137, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9280;   http:/
/muse.jhu.edu/article/9280/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/134/700252/jcws.2002.4.3.134.pdf.

Leff:2003:RNE

[Lef03] Carol Skalnik Leff. Redrawing Nations: Ethnic Cleansing in East– Central Europe, 1944–1948 (review). Journal of Cold War Studies, 5(3):109–111, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/45205;   http://muse.jhu.edu/article/45205/pdf.

Leffler:2012:CCN

[Lef12] Melvyn P. Leffler. Curt Cardwell, NSC 68 and the Political Economy of the Early Cold War. New York: Cambridge University Press, 2011. 298 pp. Journal of Cold War Studies, 14(3):192–194, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/490179;   http:
//muse.jhu.edu/article/490179/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/3/192/698037/jcws_r_00252.pdf.

Leffler:2021:CWA

[Lef21] Melvyn P. Leffler. Cold Wars: Asia, the Middle East, Europe by Lorenz M. Lüthi (review). Journal of Cold War Studies, 23(2):162–164, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/2/162/1919921/jcws_r_01012.pdf;
https://muse.jhu.edu/article/794723.

Leitenberg:2008:JBT

[Lei08] Milton Leitenberg. Jonathan B. Tucker, War of Nerves: Chemical Warfare from World War I to al-Qaeda. Journal of Cold War Studies, 10(1):116–119, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231579;   http://muse.jhu.edu/article/231579/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/116/700065/
jcws.2008.10.1.116.pdf.

Leitenberg:2018:HOI

[Lei18] Milton Leitenberg. The hazards of operations involving nuclear weapons during the Cold War. Journal of Cold War Studies, 20(3): 207–249, Summer 2018. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/704575;
 https://direct.mit.edu/jcws/article-pdf/20/3/207/699656/
jcws_a_00826.pdf.

Lembcke:2014:VWA

[Lem14] Jerry Lembcke. The Vietnam War in American Memory: Veterans, Memorials, and the Politics of Healing by Patrick Hagopian (review). Journal of Cold War Studies, 16(4):275–277, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/566682;   http:
//muse.jhu.edu/article/566682/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/4/275/698643/jcws_r_00501.pdf.

Lenoe:2012:JPS

[Len12] Matthew Lenoe. Jan Plamper, The Stalin Cult: a Study in the Alchemy of Power. New Haven: Yale University Press, 2012. 310 pp. Journal of Cold War Studies, 14(4):255–257, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/500350;   http:
//muse.jhu.edu/article/500350/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/4/255/698167/jcws_r_00297.pdf.

Lenoe:2018:JRL

[Len18] Matthew Lenoe. Joshua Rubenstein, The Last Days of Stalin. Journal of Cold War Studies, 20(4):256–258, Fall 2018. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718219;   https://direct.mit.
edu/jcws/article-pdf/20/4/256/1969616/jcws_r_00845.pdf.

Lerner:2002:UFP

[Ler02] Mitchell Lerner. US Foreign Policy in World History (review). Journal of Cold War Studies, 4(4):104–106, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9302;   http://muse.jhu.edu/article/9302/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/4/104/695678/
15203970260209536.pdf.

Lerner:2004:DMN

[Ler04] Mitchell B. Lerner. A dangerous miscalculation: New evidence from Communist-bloc archives about North Korea and the crisis of 1968. Journal of Cold War Studies, 6(1):3–21, Winter 2004. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/51868;   http://muse.jhu.edu/
article/51868/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/1/3/695865/152039704772741579.pdf.

Lerner:2006:RMF

[Ler06] Mitchell B. Lerner. Richard A. Mobley, Flash Point North Korea: The Pueblo and EC-121 Crises. Journal of Cold War Studies, 8(2):136–137, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196285;   http://muse.jhu.edu/article/196285/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/136/700512/
jcws.2006.8.2.136.pdf.

Lerner:2015:BCW

[Ler15] Mitchell Lerner. Beyond the Cold War: Lyndon Johnson and the New Global Challenges of the 1960s edited by Francis J. Gavin and Mark Atwood Lawrence (review). Journal of Cold War Studies, 17(3):281–282, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595104;   http://muse.jhu.edu/article/595104/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/281/
1102424/jcws_r_00524.pdf.

Leustean:2013:RCD

[Leu13] Lucian N. Leustean. Roman Catholicism, diplomacy, and the European Communities, 1958–1964. Journal of Cold War Studies, 15(1):

53–77, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/505773;
http://muse.jhu.edu/article/505773/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/1/53/699974/jcws_a_00308.pdf.

Levine:2004:DEC

[Lev04] Steven I. Levine. Decisive Encounters: The Chinese Civil War, 1946–1950 (review). Journal of Cold War Studies, 6(3):162–164, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/6/
3/162/696099/jcws.2004.6.3.162.pdf.

Levine:2007:HYL

[Lev07] Steven I. Levine. Hua-yu Li, Mao and the Economic Stalinization of China, 1948–1953. Journal of Cold War Studies, 9(3):203–205, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219461;   http://muse.jhu.edu/article/219461/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/203/696670/
jcws.2007.9.3.203.pdf.

Levine:2008:C

[Lev08] Steven I. Levine. Commentary. Journal of Cold War Studies, 10(2): 102–107, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/237153;
http://muse.jhu.edu/article/237153/pdf.

Levering:2010:MPL

[Lev10a] Ralph B. Levering. Melvyn P. Leffler, For the Soul of Mankind: The United States, the Soviet Union, and the Cold War. Journal of Cold War Studies, 12(3):184–186, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392806;   http://muse.jhu.edu/article/392806/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/184/697658/
jcws_r_00034.pdf.

Levine:2010:SSS

[Lev10b] Steven I. Levine. The Sino–Soviet Split: Cold War in the Communist World (review). Journal of Cold War Studies, 12(1):128–133, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/382489;   http://
muse.jhu.edu/article/382489/pdf.

Lewis:2007:JWB

[Lew07] George Lewis. Jeff Woods, Black Struggle, Red Scare: Segregation and Anti-Communism in the South, 1948–1968. Journal of Cold War Studies, 9(1):101–103, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210089;   http://muse.jhu.edu/article/210089/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/101/696488/
jcws.2007.9.1.101.pdf.

Lewkowicz:2012:PLE

[Lew12] Nicolas Lewkowicz. N. Piers Ludlow, ed., European Integration and the Cold War: Ostpolitik-Westpolitik, 1965–1973. London: Routledge, 2009. xii + 194 pp. £21.15. Journal of Cold War Studies, 14(1):135–136, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467696;   http://muse.jhu.edu/article/467696/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/135/697921/
jcws_r_00207.pdf.

Li:2018:BBB

[Li18] Hongshan Li. Building a black bridge: China’s interaction with African–American activists during the Cold War. Journal of Cold War Studies, 20(3):114–152, Summer 2018. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/704572;   https://direct.mit.
edu/jcws/article-pdf/20/3/114/699616/jcws_a_00814.pdf.

Libbey:2001:AMU

[Lib01] James K. Libbey. Anatomy of Mistrust: U.S.–Soviet Relations During the Cold War (review). Journal of Cold War Studies, 3(3):104– 106, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9206;   http:/
/muse.jhu.edu/article/9206/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/104/695462/jcws.2001.3.3.104.pdf.

Liber:2007:MLS

[Lib07] George O. Liber. Moshe Lewin, The Soviet Century, ed. by Gregory Elliot. Journal of Cold War Studies, 9(3):197–199, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/219462;   http:
//muse.jhu.edu/article/219462/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/3/197/696684/jcws.2007.9.3.197.pdf.

Liberman:2009:KNP

[Lib09] Peter Liberman. Kevin Narizny, The Political Economy of Grand Strategy. Journal of Cold War Studies, 11(1):152–154, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/259189;   http:
//muse.jhu.edu/article/259189/pdf;   https://direct.mit.
edu/jcws/article-pdf/11/1/152/1468518/jcws.2009.11.1.152.
pdf.

Lichtman:2020:JPE

[Lic20] Robert M. Lichtman. John P. Enyeart, Death to Fascism: Louis Adamic’s Fight for Democracy. Journal of Cold War Studies, 22(2):175–177, Spring 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/2/175/1860564/jcws_r_00944.pdf;
https://muse.jhu.edu/article/756818.

Lieshout:2008:WLE

[Lie08] Robert H. Lieshout. Wilfried Loth, ed., La Gouvernance Supranationale dans la Construction Européenne [Supranational Governance during the Construction of Europe]. Journal of Cold War Studies, 10(3):158–160, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241511;   http://muse.jhu.edu/article/241511/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/158/697099/
jcws.2008.10.3.158.pdf.

Lieshout:2012:SRE

[Lie12] Robert H. Lieshout. Sebastian Rosato, Europe United: Power Politics and the Making of the European Community. Ithaca, NY: Cornell University Press, 2011. 265 pp. $35.00. Journal of Cold War Studies, 14(4):234–237, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500341;   http://muse.jhu.edu/article/500341/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/234/698149/
jcws_r_00288.pdf.

Lieberman:2021:ASS

[Lie21] David Isadore Lieberman. Arnold Schoenberg’s A Survivor from Warsaw in Postwar Europe, by Joy H. Calico, Berkeley: University of California Press, 2014. 254 pp. $60.00 (review). Journal of Cold War Studies, 23(1):254–256, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/1/254/1911824/jcws_r_01003.pdf;
https://muse.jhu.edu/article/788388.

Lin:2009:CEV

[Lin09] Mao Lin. China and the escalation of the Vietnam War: The first years of the Johnson Administration. Journal of Cold War Studies, 11(2):35–69, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/265886;   http://muse.jhu.edu/article/265886/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/2/35/697211/
jcws.2009.11.2.35.pdf.

Lin:2011:JGM

[Lin11] Mao Lin. Jean A. Garrison, Making China Policy: From Nixon to G. W. Bush. Boulder, CO: Lynne Rienner Publishers, 2005. 255 pp. $58.00 hardcover, $22.50 paper and Yukinori Komine, Secrecy in U.S. Foreign Policy: Nixon, Kissinger and the Rapprochement with China. Aldershot, UK: Ashgate Publishing, 2008. 287 pp. £52.25. Journal of Cold War Studies, 13(4):246–248, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457214;   http://muse.jhu.edu/article/457214/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/246/697800/
jcws_r_00126.pdf.

Lin:2015:VVW

[Lin15] Mao Lin. Voices from the Vietnam War: Stories from American, Asian, and Russian Veterans by Xiaobing Li (review). Journal of Cold War Studies, 17(3):244–245, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595086;   http://muse.jhu.edu/article/595086/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/244/
1102464/jcws_r_00575.pdf.

Little:2010:USK

[Lit10] Douglas Little. The United States and the Kurds: a Cold War story. Journal of Cold War Studies, 12(4):63–98, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/403797;   http:
//muse.jhu.edu/article/403797/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/4/63/697548/jcws_r_00048.pdf.

Little:2011:JPS

[Lit11a] Douglas Little. John Prados. Safe for Democracy: The Secret Wars of the CIA. Chicago: Ivan Dee Publishers, 2006. 696 pp. $35.00. Journal of Cold War Studies, 13(3):207–210, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/447470;   http:
//muse.jhu.edu/article/447470/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/3/207/700736/jcws_r_00149.pdf.

Litwak:2011:RJW

[Lit11b] Robert S. Litwak. Robert Jervis, Why Intelligence Fails: Lessons from the Iranian Revolution and the Iraq War. Ithaca, NY: Cornell University Press, 2010. 248 pp. $27.95. Journal of Cold War Studies, 13(1):223–225, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/424246;   http://muse.jhu.edu/article/424246/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/1/223/697701/
jcws_r_00076.pdf.

Livingston:2007:WSE

[Liv07] Jeffery C. Livingston. William Stueck, ed., The Korean War in World History. University Press of Kentucky, 2004. ix + 203 pp. $35.00. Journal of Cold War Studies, 9(2):168–169, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/214748;   http:
//muse.jhu.edu/article/214748/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/2/168/696453/jcws.2007.9.2.168.pdf.

Locher:2005:KGW

[Loc05] Anna Locher. Kennedy, de Gaulle, and Western Europe (review). Journal of Cold War Studies, 7(4):164–166, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.

jhu.edu/article/188272;   http://muse.jhu.edu/article/188272/
pdf;   https://direct.mit.edu/jcws/article-pdf/7/4/164/696295/

jcws.2005.7.4.164.pdf. See comments [Wal12c, Kol12].

Loedel:2010:JRZ

[Loe10] Peter Loedel. Jonathan R. Zatlin, The Currency of Socialism: Money and Political Culture in East Germany. Journal of Cold War Studies, 12(1):194–195, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/382504;   http://muse.jhu.edu/article/382504/

pdf;   https://direct.mit.edu/jcws/article-pdf/12/1/194/
1627447/jcws.2010.12.1.194.pdf.

Logevall:2001:BOS

[Log01] Fredrik Logevall. Bringing in the “other side”: New scholarship on the Vietnam Wars. Journal of Cold War Studies, 3(3):77– 93, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9207;   http:/
/muse.jhu.edu/article/9207/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/77/695432/152039701750419529.pdf.

Lomperis:2009:MAL

[Lom09] Timothy J. Lomperis. Mark Atwood Lawrence, Assuming the Burden: Europe and the American Commitment to War in Vietnam. Journal of Cold War Studies, 11(2):162–164, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/265862;   http://muse.jhu.edu/article/265862/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/2/162/
697207/jcws.2009.11.2.162.pdf.

Long:2011:QAC

[Lon11] Austin Long. The quiet Americans?: CIA, NSA, and counterinsurgency. Journal of Cold War Studies, 13(3):149–184, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/447465;   http:
//muse.jhu.edu/article/447465/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/3/149/700716/jcws_r_00146.pdf.

Lorke:2017:DSD

[Lor17] Christoph Lorke. Depictions of social dissent in East German television detective series, 1970–1989. Journal of Cold War Studies, 19(4):168–191, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683669;   http://muse.jhu.edu/article/683669/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/168/699364/
jcws_a_00768.pdf.

Loureiro:2016:WML

[Lou16] Felipe P. Loureiro. William M. LeoGrande and Peter Kornbluh, Back Channel to Cuba: The Hidden History of Negotiations between Washington and Havana. Chapel Hill: University of North Carolina Press, 2014. 524 pp. $25.00. Journal of Cold War Studies, 18(4):210–212, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645914;   http://muse.jhu.edu/article/645914/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/210/700059/
jcws_r_00686.pdf.

Loureiro:2023:MAP

[Lou23] Felipe P. Loureiro. Making the alliance for progress serve the few: U.S. economic aid to Cold War Brazil (1961–1964). Journal of Cold War Studies, 25(1):168–207, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/1/168/2073514/jcws_a_01125.pdf.

Loveman:2010:GJD

[Lov10] Brian Loveman. Gilbert Joseph and Daniela Spenser, eds., In from the Cold: Latin America’s New Encounter with the Cold War. Journal of Cold War Studies, 12(3):190–193, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/392809;   http://muse.jhu.edu/article/392809/
pdf;   https://direct.mit.edu/jcws/article-pdf/12/3/190/
697696/jcws_r_00037.pdf.

Loveman:2022:BAR

[Lov22] Brian Loveman. Beatriz Allende: A revolutionary life in Cold War Latin America. Journal of Cold War Studies, 24(3):233– 238, Summer 2022. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL https://direct.mit.edu/jcws/
article-pdf/24/3/233/2042001/jcws_r_01095.pdf;   https://
muse.jhu.edu/article/864238.

Lowen:2004:SCC

[Low04] Rebecca S. Lowen. Scientists in the Classroom: The Cold War Reconstruction of American Science Education (review). Journal of Cold War Studies, 6(4):155–157, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/174715;   http://muse.jhu.edu/article/174715/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/4/155/696097/
jcws.2004.6.4.155.pdf.

Lowenthal:2006:JPL

[Low06] Mark M. Lowenthal. John Prados, Lost Crusader: The Secret Wars of CIA Director William Colby, and: Rhodri Jeffreys-Jones, Cloak and Dollar: The History of American Secret Intelligence. Journal of Cold War Studies, 8(2):130–132, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196286;   http://muse.jhu.edu/article/196286/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/130/700510/
jcws.2006.8.2.130a.pdf;   https://direct.mit.edu/jcws/article-
pdf/8/2/130/700526/jcws.2006.8.2.130.pdf.

Lowenthal:2008:AMI

[Low08] Abraham F. Lowenthal. Alan McPherson, Intimate Ties, Bitter Struggles: The United States and Latin America since 1945. Journal of Cold War Studies, 10(1):119–121, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231580;   http://muse.jhu.edu/article/231580/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/119/700045/
jcws.2008.10.1.119.pdf.

Lieshout:2004:GMC

[LSvdV04] Robert S. Lieshout, Mathieu L. L. Segers, and Johanna Maria van der Vleuten. De Gaulle, Moravcsik, and the choice for Europe: Soft sources, weak evidence. Journal of Cold War Studies, 6(4): 89–139, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/174714;
http://muse.jhu.edu/article/174714/pdf;   https://direct.
mit.edu/jcws/article-pdf/6/4/89/696077/1520397042350900.

pdf. See [Mor00b, Mor00c].

Lubenow:2011:KCB

[Lub11] Gerald C. Lubenow. Kenneth C. Burt, The Search for a Civic Voice: California Latino Politics. Claremont, CA: Regina Books, 2007. 438 pp. $24.95. Journal of Cold War Studies, 13(1):248–250, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/424259;   http:
//muse.jhu.edu/article/424259/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/1/248/697769/jcws_r_00090.pdf.

Lucas:2002:DVU

[Luc02] Scott Lucas. Dueling Visions: U.S. Strategy toward Eastern Europe under Eisenhower (review). Journal of Cold War Studies, 4(3):132– 134, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9281;   http:/
/muse.jhu.edu/article/9281/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/132/700244/jcws.2002.4.3.132.pdf.

Luck:2007:NMY

[Luc07] Edward C. Luck. S. Neil MacFarlane and Yuen Foong Khong, Human Security and the UN: a Critical History. Bloomington: Indiana University Press, 2006. 346 pp. $85.00. Journal of Cold War Studies, 9(2):146–148, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214749;   http://muse.jhu.edu/article/214749/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/146/696433/
jcws.2007.9.2.146.pdf.

Ludlow:2013:RYE

[Lud13] N. Piers Ludlow. The real years of Europe?: U.S.–West European relations during the Ford Administration. Journal of Cold War Studies, 15(3):136–161, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522712;   http://muse.jhu.edu/article/522712/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/136/698276/
jcws_a_00373.pdf.

Lukes:2001:CPP

[Luk01] Igor Lukes. Changing patterns of power in Cold War politics: The mysterious case of Vladimír Komárek. Journal of Cold War Studies, 3(1):61–102, Winter 2001. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/9179;
 http://muse.jhu.edu/article/9179/pdf;   https://direct.
mit.edu/jcws/article-pdf/3/1/61/695408/15203970151032155.
pdf.

Lukes:2007:CSS

[Luk07] Igor Lukes. The Czechoslovak Special Services and their American adversary during the Cold War. Journal of Cold War Studies, 9(1):3–28, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/210091;   http:
//muse.jhu.edu/article/210091/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/1/3/696597/jcws.2007.9.1.3.pdf.

Lukes:2016:ECW

[Luk16] Igor Lukes. On the edge of the Cold War: a reply. Journal of Cold War Studies, 18(3):181–189, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632283;   http://muse.jhu.edu/article/632283/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/181/698969/
jcws_a_00657.pdf.

Lukoff:2017:GHG

[Luk17] Lee Lukoff. Gregg Herken, The Georgetown Set: Friends and Rivals in Cold War Washington. Journal of Cold War Studies, 19(4):277–279, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683693;   http://muse.jhu.edu/article/683693/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/277/699467/
jcws_r_00793.pdf.

Lukasiewicz:2022:SES

[Łuk22] Sławomir Łukasiewicz. Security empire: The secret police in Communist Eastern Europe. Journal of Cold War Studies, 24(3):252–254, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/3/
252/2041987/jcws_r_01099.pdf.

Lukasiewicz:2023:SPS

[Łuk23] Sławomir Łukasiewicz. A shadow party system: The political activities of Cold War Polish exiles. Journal of Cold War Studies, 25(1):46–74, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/1/
46/2073511/jcws_a_01121.pdf.

Lumbers:2004:IVJ

[Lum04] Michael Lumbers. The irony of Vietnam: The Johnson Administration’s tentative bridge building to China, 1965–1966. Journal of Cold War Studies, 6(3):68–114, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/171167;   http://muse.jhu.edu/article/171167/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/3/68/696023/1520397041447373.
pdf.

Lundy:2004:AIC

[Lun04] Derek Lundy. America and the Intellectual Cold Wars in Europe: Shepard Stone between Philanthropy, Academy, and Diplomacy (review). Journal of Cold War Studies, 6(3):152–153, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171168;   http:
//muse.jhu.edu/article/171168/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/3/152/696069/jcws.2004.6.3.152.pdf.

Lunak:2010:ZZR

[Luň10] Petr Luňák. Zbynek Zeman and Rainer Karlsch, Uranium Matters: Central European Uranium in International Politics, 1900– 1960. Journal of Cold War Studies, 12(3):149–150, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/392791;   http:
//muse.jhu.edu/article/392791/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/3/149/697604/jcws_r_00018.pdf.

Luthi:2006:IVG

[Lüt06] Lorenz M. Lüthi. Ilya V. Gaiduk, Confronting Vietnam: Soviet Policy toward the Indochina Conflict, 1945–1963. Journal of Cold War Studies, 8(4):164–165, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/204625;   http://muse.jhu.edu/article/204625/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/4/164/700809/
jcws.2006.8.4.164.pdf.

Luthi:2008:VWC

[Lüt08] Lorenz Lüthi. The Vietnam War and China’s third-line defense planning before the Cultural Revolution, 1964–1966. Journal of Cold War Studies, 10(1):26–51, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231581;   http://muse.jhu.edu/article/231581/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/26/700117/
jcws.2008.10.1.26.pdf.

Luthi:2009:BBB

[Lüt09] Lorenz M. Lüthi. Beyond betrayal: Beijing, Moscow, and the Paris negotiations, 1971–1973. Journal of Cold War Studies, 11(1):57–107, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/259185;   http://muse.jhu.edu/article/259185/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/1/57/1468569/
jcws.2009.11.1.57.pdf.

Luthi:2010:RC

[Lüt10] Lorenz Lüthi. Reply to the commentaries. Journal of Cold War Studies, 12(1):161–165, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/382489;   http://muse.jhu.edu/article/382489/pdf.

Luthi:2012:PRE

[Lüt12] Lorenz M. Lüthi. Priscilla Roberts, ed., Behind the Bamboo Curtain: China, Vietnam, and the World beyond Asia. Stanford, CA: Stanford University Press, 2007. 577 pp. $65.00. Journal of Cold War Studies, 14(1):126–128, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467692;   http://muse.jhu.edu/article/467692/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/126/697882/
jcws_r_00203.pdf.

Luthi:2014:RIR

[Lüt14] Lorenz M. Lüthi. Rearranging international relations?: How Mao’s China and de Gaulle’s France recognized each other in 1963–1964. Journal of Cold War Studies, 16(1):111–145, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543727;   http:
//muse.jhu.edu/article/543727/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/1/111/698934/jcws_a_00433.pdf.

Luthi:2016:NAM

[Lüt16] Lorenz M. Lüthi. The non-aligned movement and the Cold War, 1961–1973. Journal of Cold War Studies, 18(4):98–147, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/645910;
http://muse.jhu.edu/article/645910/pdf;   https://direct.
mit.edu/jcws/article-pdf/18/4/98/700035/jcws_a_00682.pdf.

Luthi:2022:BRS

[Lüt22a] Lorenz M. Lüthi. Book review: The Sino-Soviet Alliance: An International History by Austin Jersild. Journal of Cold War Studies, 24 (4):238–241, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/24/4/238/2062870/jcws_r_01116.pdf.

Lutsch:2022:WGN

[Lut22b] Andreas Lutsch. West Germany and NATO’s nuclear force posture in the early 1960s (part 1). Journal of Cold War Studies, 24(4):4–58, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/4/
4/2062829/jcws_a_01089.pdf.

Li:2014:JLM

[LX14] Danhui Li and Yafeng Xia. Jockeying for leadership: Mao and the Sino–Soviet split, October 1961–July 1964. Journal of Cold War Studies, 16(1):24–60, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/543724;   http://muse.jhu.edu/article/543724/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/1/24/698924/
jcws_a_00430.pdf.

Liang:2023:CWH

[LX23] Zhi Liang and Yafeng Xia. Cold War history studies in China in the 21st Century: The state of the field. Journal of Cold War Studies, 25 (2):11–40, Spring 2023. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/25/2/11/2140988/jcws_a_01140.pdf.

Lynn:2016:GNA

[Lyn16] Denise Lynn. Gendered narratives in anti-Stalinism and anti-Communism during the Cold War: The case of Juliet Poyntz. Journal of Cold War Studies, 18(1):31–59, Winter 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/612005;   http://muse.jhu.edu/article/612005/
pdf;   https://direct.mit.edu/jcws/article-pdf/18/1/31/698759/
jcws_a_00618.pdf.

Macqueen:2006:BDP

[Mac06] Norrie Macqueen. Belated decolonization and UN politics against the backdrop of the Cold War: Portugal, Britain, and Guinea-Bissau’s proclamation of independence, 1973–1974. Journal of Cold War Studies, 8(4):29–56, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/204626;   http://muse.jhu.edu/article/204626/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/4/29/700825/jcws.
2006.8.4.29.pdf.

Macekura:2019:IDP

[Mac19] Stephen Macekura. International Development: A Postwar History by Corinna R. Unger (review). Journal of Cold War Studies, 21(2): 206–207, Spring 2019. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/726144;
 https://direct.mit.edu/jcws/article-pdf/21/2/206/699589/
jcws_r_00889.pdf.

Machcewicz:2020:JES

[Mac20a] Paweł Machcewicz. Jörg Echternkamp and Stephan Jaeger, eds., Views of Violence: Representing the Second World War in German and European Museums and Memorials. Journal of Cold War Studies, 22(4):237–239, Fall 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/4/237/1860640/jcws_r_00975.pdf;
https://muse.jhu.edu/article/776967.

Macrakis:2020:VGA

[Mac20b] Kristie Macrakis. Valentina Glajar, Alison Lewis, and Corina Petrescu, eds., Cold War Spy Stories from Eastern Europe. Journal of Cold War Studies, 22(4):242–243, Fall 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/4/242/1860653/jcws_r_00977.pdf;
https://muse.jhu.edu/article/776968.

Maddux:2002:EDR

[Mad02] Thomas R. Maddux. Eisenhower Decides to Run: Presidential Politics and Cold War Strategy (review). Journal of Cold War Studies, 4(2):139–141, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9254;   http://muse.jhu.edu/article/9254/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/2/139/695595/
jcws.2002.4.2.139.pdf.

Maddux:2004:ASV

[Mad04] Thomas R. Maddux. Another Such Victory: President Truman and the Cold War, 1945–1953 (review). Journal of Cold War Studies, 6(1):96–99, Winter 2004. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
51869;   http://muse.jhu.edu/article/51869/pdf;   https:/
/direct.mit.edu/jcws/article-pdf/6/1/96/695907/jcws.2004.
6.1.96.pdf.

Maddux:2007:JBL

[Mad07] Thomas R. Maddux. Jonathan Bell, The Liberal State on Trial: The Cold War and American Politics in the Truman Years. Journal of Cold War Studies, 9(1):99–101, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/211267;   http://muse.jhu.edu/article/211267/
pdf;   https://direct.mit.edu/jcws/article-pdf/9/1/99/696614/
jcws.2007.9.1.99.pdf.

Maddux:2008:CWA

[Mad08a] Thomas Maddux. The Cold War after Stalin’s Death: a Missed Opportunity for Peace? (review). Journal of Cold War Studies, 10(2): 134–136, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/237155;
http://muse.jhu.edu/article/237155/pdf.

Maddux:2008:MJW

[Mad08b] Thomas R. Maddux. Mark J. White, Against the President: Dissent and Decision-Making in the White House — A Historical Perspective. Journal of Cold War Studies, 10(4):147–149, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/252657;
http://muse.jhu.edu/article/252657/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/4/147/697043/jcws.2008.10.4.
147.pdf.

Maddox:2009:MKC

[Mad09] Robert James Maddox. Michael Kort, The Columbia Guide to Hiroshima and the Bomb. Journal of Cold War Studies, 11(4):147– 149, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/366258;
http://muse.jhu.edu/article/366258/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/4/147/1627613/jcws.2009.11.4.
147.pdf.

Maddrell:2013:EDC

[Mad13] Paul Maddrell. The economic dimension of Cold War intelligencegathering: The West’s spies in the GDR’s economy. Journal of Cold War Studies, 15(3):76–107, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522710;   http://muse.jhu.edu/article/522710/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/76/698280/
jcws_a_00371.pdf.

Maddox:2015:WMM

[Mad15] Robert James Maddox. Wilson Miscamble, The Most Controversial Decision: Truman, the Atomic Bombs, and the Defeat of Japan. New York: Cambridge University Press, 2011. 174 pp. Journal of Cold War Studies, 17(4):206–207, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606041;   http://muse.jhu.edu/article/606041/pdf;

 https://direct.mit.edu/jcws/article-pdf/17/4/206/698800/
jcws_r_00609.pdf.

Maggs:2010:PKG

[Mag10] Peter B. Maggs. Patricia Kennedy Grimstead, F. J. Hoogewoud, and Eric Ketelaar, eds., Returned from Russia: Nazi Archival Plunder in Western Europe and Recent Restitution Issues. Journal of Cold War Studies, 12(1):200–202, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/382505;   http://muse.jhu.edu/article/382505/
pdf;   https://direct.mit.edu/jcws/article-pdf/12/1/200/
1627445/jcws.2010.12.1.200.pdf.

Maggor:2014:JAR

[Mag14a] Noam Maggor. Joyce Appleby, The Relentless Revolution: a History of Capitalism. New York: W. W. Norton, 2011. 494 pp. Journal of Cold War Studies, 16(1):214–216, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/543733;   http://muse.jhu.edu/article/543733/
pdf;   https://direct.mit.edu/jcws/article-pdf/16/1/214/
698920/jcws_r_00411.pdf.

Maggs:2014:TBW

[Mag14b] Peter Maggs. Tatiana Borisova and William Simons, eds., The Legal Dimension in Cold War Interactions: Some Notes from the Field. Leiden: Martinus Nijhoff Publishers, 2012. xxiv + 185 pp. Journal of Cold War Studies, 16(1):244–245, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/543747;   http://muse.jhu.edu/article/543747/
pdf;   https://direct.mit.edu/jcws/article-pdf/16/1/244/
698971/jcws_r_00438.pdf.

Maier:2005:MPD

[Mai05] Charles S. Maier. The Marshall Plan and the division of Europe. Journal of Cold War Studies, 7(1):168–174, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/179567;   http:
//muse.jhu.edu/article/179567/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/1/168/700434/1520397053326167.pdf.

Makko:2012:SEC

[Mak12] Aryo Makko. Sweden, Europe, and the Cold War: a reappraisal. Journal of Cold War Studies, 14(2):68–97, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/478776;   http:
//muse.jhu.edu/article/478776/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/2/68/697975/jcws_a_00221.pdf.

Makinen:2017:WAO

[Mak17] Marvin W. Makinen. The Wallenberg Affair and the onset of the Cold War. Journal of Cold War Studies, 19(3):215–224, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/668164;   http:
//muse.jhu.edu/article/668164/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/3/215/699343/jcws_r_00760.pdf.

Malloy:2015:CCS

[Mal15] Sean L. Malloy. Campbell Craig and Sergey Radchenko, The Atomic Bomb and the Origins of the Cold War. New Haven: Yale University Press, 2008. 232 pp. $27.00. Journal of Cold War Studies, 17(4):208–210, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606042;   http://muse.jhu.edu/article/606042/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/208/698816/
jcws_r_00610.pdf.

Manela:2010:KAT

[Man10] Erez Manela. Kristin L. Ahlberg, Transplanting the Great Society: Lyndon Johnson and Food for Peace. Journal of Cold War Studies, 12(3):129–131, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392783;   http://muse.jhu.edu/article/392783/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/129/697606/
jcws_r_00010.pdf.

Manela:2013:MLR

[Man13] Erez Manela. Michael E. Latham, The Right Kind of Revolution: Modernization, Development, and U.S. Foreign Policy from the Cold War to the Present. Ithaca, NY: Cornell University Press, 2011. 256 pp. Journal of Cold War Studies, 15(1):162–164, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/505761;   http:
//muse.jhu.edu/article/505761/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/1/162/699998/jcws_r_00318.pdf.

Manchester:2021:CDC

[Man21] Margaret Murányi Manchester. The corporate dimension of the Cold War in Hungary: ITT and the Vogeler/Sanders Case reconsidered. Journal of Cold War Studies, 23(2):41–74, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/23/2/41/1922987/jcws_a_
00983.pdf;   https://muse.jhu.edu/article/794726.

Marples:2001:RAR

[Mar01] David R. Marples. Red Atom: Russia’s Nuclear Power Program from Stalin to Today (review). Journal of Cold War Studies, 3(3):110– 112, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9208;   http:/
/muse.jhu.edu/article/9208/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/110/695460/jcws.2001.3.3.110.pdf.

Marchio:2002:PCG

[Mar02] James Marchio. The planning coordination group: Bureaucratic casualty in the Cold War campaign to exploit Soviet-bloc vulnerabilities 1. Journal of Cold War Studies, 4(4):3–28, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9303;   http:/
/muse.jhu.edu/article/9303/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/4/3/695688/15203970260209491.pdf.

Marsh:2005:CCR

[Mar05] Steve Marsh. Continuity and change: Reinterpreting the policies of the Truman and Eisenhower Administrations toward Iran, 1950–1954. Journal of Cold War Studies, 7(3):79–123, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/185114;
http://muse.jhu.edu/article/185114/pdf;   https://direct.
mit.edu/jcws/article-pdf/7/3/79/1102557/1520397054377205.
pdf.

Mark:2006:EOW

[Mar06a] Eduard Maximilian Mark. Experiment in Occupation: Witness to the Turnabout–Anti-Nazi War to Cold War, 1944–1946 (review). Journal of Cold War Studies, 8(1):140–142, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/192216;   http:
//muse.jhu.edu/article/192216/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/1/140/696319/jcws.2006.8.1.140.pdf.

Martin:2006:REN

[Mar06b] Terry Martin. E. A. Rees, ed., The Nature of Stalin’s Dictatorship: The Politburo, 1924–1953. Journal of Cold War Studies, 8(4):165–167, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/204628;   http://muse.jhu.edu/article/204628/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/4/165/700819/
jcws.2006.8.4.165.pdf.

Mark:2007:MCB

[Mar07a] Eduard Maximilian Mark. Michael Cassella-Blackburn, The Donkey, the Carrot, and the Club: William C. Bullitt and Soviet-American Relations, 1917–1948. Journal of Cold War Studies, 9(3):170–172, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219463;   http://muse.jhu.edu/article/219463/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/170/696628/
jcws.2007.9.3.170.pdf.

Marling:2007:DMB

[Mar07b] Karal Ann Marling. Donna M. Binkiewicz, Federalizing the Muse: United States Arts Policy and the National Endowment for the Arts, 1965–1980. Chapel Hill: University of North Carolina Press, 2004. 295 pp. Journal of Cold War Studies, 9 (2):156–158, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214750;   http://muse.jhu.edu/article/214750/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/156/696439/
jcws.2007.9.2.156.pdf.

Marquardt:2007:TSC

[Mar07c] James J. Marquardt. Transparency and security competition: Open Skies and America’s Cold War statecraft, 1948–1960. Journal of Cold War Studies, 9(1):55–87, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210092;   http://muse.jhu.edu/article/210092/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/55/696609/jcws.
2007.9.1.55.pdf.

Martin:2007:CEC

[Mar07d] Jessica Martin. Christopher Endy, Cold War Holidays: American Tourism in France. Chapel Hill: University of North Carolina Press, 2004. 286 pp. $19.95. Journal of Cold War Studies, 9(2):175–177, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214751;   http://muse.jhu.edu/article/214751/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/175/696463/
jcws.2007.9.2.175.pdf.

Mark:2008:GEW

[Mar08a] Eduard Mark. G. Edward White, Alger Hiss’s Looking-Glass War: The Covert Life of a Soviet Spy. Journal of Cold War Studies, 10(2):157–160, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/237162;   http://muse.jhu.edu/article/237162/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/2/157/697003/
jcws.2008.10.2.157.pdf.

Marling:2008:MKF

[Mar08b] Karal Ann Marling. Michael L. Krenn, Fall-Out Shelters for the Human Spirit: American Art and the Cold War. Journal of Cold War Studies, 10(1):168–170, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231582;   http://muse.jhu.edu/article/231582/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/168/700103/
jcws.2008.10.1.168.pdf.

Martin:2008:PCC

[Mar08c] Garret Martin. Playing the China card?: Revisiting France’s recognition of Communist China, 1963–1964. Journal of Cold War Studies, 10(1):52–80, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231583;   http://muse.jhu.edu/article/231583/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/52/700170/
jcws.2008.10.1.52.pdf.

Mark:2009:RAH

[Mar09] Eduard Mark. In Re Alger Hiss: a final verdict from the archives of the KGB. Journal of Cold War Studies, 11(3):26–67, Summer 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/268390;   http:
//muse.jhu.edu/article/268390/pdf;   https://direct.mit.
edu/jcws/article-pdf/11/3/26/697242/jcws.2009.11.3.26.pdf.

Mark:2010:GGS

[Mar10] Eduard Mark. Gabriel Gorodetsky, Stafford Cripps in Moscow, 1940–1942: Diaries and Papers. Journal of Cold War Studies, 12(3):168–171, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392800;   http://muse.jhu.edu/article/392800/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/168/697684/
jcws_r_00027.pdf.

Marshall:2012:PSB

[Mar12] Barbara Marshall. Paul Steege, Black Market, Cold War: Everyday Life in Berlin 1946–1949. New York: Cambridge University Press, 2007. 348 pp. $80.00. Journal of Cold War Studies, 14(1):141–143, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467699;   http://muse.jhu.edu/article/467699/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/141/697915/
jcws_r_00210.pdf.

Marantzidis:2013:GCW

[Mar13a] Nikos Marantzidis. The Greek Civil War (1944–1949) and the international Communist system. Journal of Cold War Studies, 15 (4):25–54, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/534424;
http://muse.jhu.edu/article/534424/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/4/25/698307/jcws_a_00394.pdf.

Martin:2013:LBD

[Mar13b] Garret J. Martin. Laurence Badel, Diplomatie et grands contrats: L’état français et les marchés extérieurs au XXe siècle. Paris: Publications de la Sorbonne, 2010. 512 pp. EUR 40. Journal of Cold War Studies, 15(3):202–204, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522725;   http://muse.jhu.edu/article/522725/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/202/698299/
jcws_r_00352.pdf.

Martin:2015:BCA

[Mar15] Kevin W. Martin. “Behind Cinerama’s aluminum Curtain”: Cold War spectacle and propaganda at the First Damascus International Exposition. Journal of Cold War Studies, 17(4):59– 85, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/606021;

http://muse.jhu.edu/article/606021/pdf;   https://direct.
mit.edu/jcws/article-pdf/17/4/59/698733/jcws_a_00597.pdf.

Martin:2016:NMI

[Mar16] Garret J. Martin. Nicolas Mary, Les Intellectuels et les figures politiques charismatiques: De Gaulle, Mendès France, Mitterrand. Paris: Les Indes Savantes, 2013. 467 pp. EUR 36.00. Journal of Cold War Studies, 18(4):263–264, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645937;   http://muse.jhu.edu/article/645937/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/263/700202/
jcws_r_00711.pdf.

Mariager:2018:DCW

[Mar18] Rasmus Mariager. Danish cold war historiography. Journal of Cold War Studies, 20(4):180–211, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718203;   https://direct.mit.
edu/jcws/article-pdf/20/4/180/1969648/jcws_a_00825.pdf.

Martin:2020:SMJ

[Mar20] Garret J. Martin. Sean McLaughlin, JFK and de Gaulle: How America and France Failed in Vietnam, 1961–1963. Journal of Cold War Studies, 22(4):230–232, Fall 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/4/230/1860622/jcws_r_00972.pdf;
https://muse.jhu.edu/article/776969.

Martin:2021:SMD

[Mar21] Barbara Martin. The Sakharov–Medvedev debate on détente and human rights: From the Jackson–Vanik Amendment to the Helsinki Accords. Journal of Cold War Studies, 23(3):138–174, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/23/3/138/1955783/
jcws_a_01009.pdf;   https://muse.jhu.edu/article/801907.

Marku:2022:SAA

[Mar22] Ylber Marku. Shifting alliances: Albania in the Early Cold War. Journal of Cold War Studies, 24(3):80–115, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/24/3/80/2041999/jcws_a_
01074.pdf.

Marantzidis:2023:SBD

[Mar23a] Nikos Marantzidis. Book review: Storms over the Balkans during the Second World War by Alfred J. Rieber. Journal of Cold War Studies, 25(4):257–259, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/25/4/257/2204987/jcws_r_01181.pdf.

Marshall:2023:UCW

[Mar23b] Jonathan Marshall. U.S. Cold War policy and the Italian far-right: The Nixon administration, Republican Party operatives, and the Borghese coup plot of 1970. Journal of Cold War Studies, 25(1): 138–167, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/25/1/138/2073499/jcws_a_01124.pdf.

Mastny:2000:CWS

[Mas00] Vojtech Mastny. Cold War Statesmen Confront the Bomb: Nuclear Diplomacy since 1945 (review). Journal of Cold War Studies, 2(2):125–127, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9144;   http://muse.jhu.edu/article/9144/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/2/125/695257/
jwcs.2000.2.2.125.pdf.

Mastny:2002:NHC

[Mas02] Vojtech Mastny. The new history of Cold War alliances. Journal of Cold War Studies, 4(2):55–84, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9255;   http://muse.jhu.edu/article/
9255/pdf;   https://direct.mit.edu/jcws/article-pdf/4/2/
55/695585/152039702753649647.pdf.

Mastny:2008:NTB

[Mas08] Vojtech Mastny. The 1963 nuclear test ban treaty: a missed opportunity for détente? Journal of Cold War Studies, 10(1):3– 25, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/231584;
http://muse.jhu.edu/article/231584/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/1/3/700142/jcws.2008.10.1.3.
pdf.

Mastny:2009:HAW

[Mas09] Vojtech Mastny. How able was “Able Archer”?: Nuclear trigger and intelligence in perspective. Journal of Cold War Studies, 11(1):108– 123, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/259186;
http://muse.jhu.edu/article/259186/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/1/108/1468502/jcws.2009.11.1.
108.pdf.

Mastny:2010:OBG

[Mas10a] Vojtech Mastny. Oliver Bange and Gottfried Niedhart, eds., Helsinki 1975 and the Transformation of Europe. Journal of Cold War Studies, 12(3):197–199, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392812;   http://muse.jhu.edu/article/392812/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/197/697688/
jcws_r_00040.pdf.

Mastny:2010:SUP

[Mas10b] Vojtech Mastny. The Soviet Union’s partnership with India. Journal of Cold War Studies, 12(3):50–90, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/392778;   http://muse.jhu.edu/article/392778/
pdf;   https://direct.mit.edu/jcws/article-pdf/12/3/50/697593/
jcws_a_00006.pdf.

Mason:2011:JJC

[Mas11a] Dwight N. Mason. Joseph T. Jockel, Canada in NORAD 1957– 2007: a History. Montreal: McGill–Queens University Press, 2007. 225 pp. $34.95. Journal of Cold War Studies, 13(4):219–221, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457201;   http:
//muse.jhu.edu/article/457201/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/4/219/697841/jcws_r_00180.pdf.

Mastny:2011:FRS

[Mas11b] Vojtech Mastny. Federico Romero, Storia della guerra fredda: L’ultimo conflitto per l’Europa [A History of the Cold War: The Last Conflict over Europe]. Turin: Einaudi, 2009. 356 pp. Journal of Cold War Studies, 13(1):250–251, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/424260;   http://muse.jhu.edu/article/424260/pdf;

 https://direct.mit.edu/jcws/article-pdf/13/1/250/697765/
jcws_r_00091.pdf.

Mastny:2012:LNE

[Mas12] Vojtech Mastny. Leopoldo Nuti, ed., The Crisis of Détente in Europe: From Helsinki to Gorbachev, 1975–1985. London: Routledge, 2009. 285 pp. Journal of Cold War Studies, 14(1):147–149, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/467701;   http:
//muse.jhu.edu/article/467701/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/1/147/697929/jcws_r_00212.pdf.

Mastny:2016:BLE

[Mas16] Vojtech Mastny. Bernd Lemke, ed., Periphery or Contact Zone? The NATO Flanks 1961 to 2013. Freiburg im Breisgau: Rombach, 2015. 231 pp. Journal of Cold War Studies, 18(3):214–216, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/632296;   http:
//muse.jhu.edu/article/632296/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/3/214/699024/jcws_r_00648.pdf.

Masso:2019:KJS

[Mas19a] Iivi Anna Masso. Kremlin jalanjäljet: Suomettuminen ja vuoden 2002 vakoilukohun tausta ed. by Jarmo Korhonen and Alpo Rusi (review). Journal of Cold War Studies, 21(1):203–206, Winter 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/723543;   https://direct.mit.
edu/jcws/article-pdf/21/1/203/699791/jcws_r_00864.pdf.

Mastny:2019:SIB

[Mas19b] Vojtech Mastny. Sowjetisch-indische Beziehungen 1941–1966: Imperiale Agenda und nationale Identität in der Ära von Dekolonisierung und Kaltem Krieg by Andreas Hilger (review). Journal of Cold War Studies, 21(3):260–261, Summer 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http:

//muse.jhu.edu/article/731290;   https://direct.mit.edu/
jcws/article-pdf/21/3/260/699667/jcws_r_00901.pdf.

Mastny:2021:EEE

[Mas21a] Vojtech Mastny. Engaging the Evil Empire: Washington, Moscow, and the Beginning of the End of the Cold War by Simon Miles (review). Journal of Cold War Studies, 23(4):246–247, Fall 2021. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/23/4/246/1970099/
jcws_r_01046.pdf;   https://muse.jhu.edu/article/835883.

Mastny:2021:NAS

[Mas21b] Vojtech Mastny. NATO after Sixty Years: A Stable Crisis ed. by James Sperling and S. Victor Papacosma, Kent, OH: Kent State University Press, 2012. 281 pp. and: NATO before the Korean War: April 1949–June 1950 by Lawrence S. Kaplan, Kent, OH: Kent State University Press, 2013. 216 pp. (review). Journal of Cold War Studies, 23(1):237–241, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/1/237/1911810/jcws_r_00991.pdf;
https://muse.jhu.edu/article/788395.

Mastny:2023:BRW

[Mas23] Vojtech Mastny. Book reviews: Would You Believe /the Helsinki Accords Changed the World? Advancing Global Human Rights and, for Decades, Security in Europe by Peter L. W. Osnos with Holly Cartner; and Defrosting the Cold War and Beyond: an Introduction to the Helsinki Process, 1954–2022 by Richard Davy. Journal of Cold War Studies, 25(3):210–212, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/3/210/2213294/jcws_r_01163.pdf.

Matlock:2004:WIC

[Mat04a] Jack F. Matlock, Jr. Western Intelligence and the Collapse of the Soviet Union, 1980–1990: Ten Years That Did Not Shake the World (review). Journal of Cold War Studies, 6(2):99–101, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/53678;   http://muse.jhu.edu/
article/53678/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/2/99/696011/jcws.2004.6.2.99.pdf.

Matray:2004:TWK

[Mat04b] James Irving Matray. Their War for Korea: American, Asian, and European Combatants and Civilians, 1945–1953 (review). Journal of Cold War Studies, 6(4):168–170, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/174716;   http://muse.jhu.edu/article/174716/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/4/168/696115/
jcws.2004.6.4.168.pdf.

Matlock:2008:ABS

[Mat08] Jack F. Matlock, Jr. Archie Brown, Seven Years That Changed the World: Perestroika in Perspective. Journal of Cold War Studies, 10(3):205–207, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241534;   http://muse.jhu.edu/article/241534/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/205/697135/
jcws.2008.10.3.205.pdf.

Matlock:2010:EIR

[Mat10a] Jack F. Matlock, Jr. Empathy in international relations: a commentary. Journal of Cold War Studies, 12(2):88–91, Spring 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/12/2/
89/697509/jcws.2010.12.2.89.pdf.

Matlock:2010:MAA

[Mat10b] Jack F. Matlock, Jr. Martin Anderson and Annelise Anderson, Reagan’s Secret War: The Untold Story of His Fight to Save the World from Nuclear Disaster. Journal of Cold War Studies, 12(4):186–188, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/403803;   http://muse.jhu.edu/article/403803/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/4/186/697544/
jcws_r_00055.pdf.

Matz:2012:SUS

[Mat12] Johan Matz. Sweden, the United States, and Raoul Wallenberg’s mission to Hungary in 1944. Journal of Cold War Studies, 14(3): 97–148, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/490176;
http://muse.jhu.edu/article/490176/pdf;   https://direct.
mit.edu/jcws/article-pdf/14/3/97/698041/jcws_a_00249.pdf.

Matthews:2015:KAI

[Mat15] Weldon C. Matthews. The Kennedy Administration, the International Federation of Petroleum Workers, and Iraqi labor under the Ba’thist regime. Journal of Cold War Studies, 17(1):97– 128, Winter 2015. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/581232;
http://muse.jhu.edu/article/581232/pdf;   https://direct.
mit.edu/jcws/article-pdf/17/1/97/700778/jcws_a_00532.pdf.

Matesan:2016:RSU

[Mat16a] Ioana E. Matesan. Randall Schweller, Unanswered Threats: Political Constraints on the Balance of Power. Princeton, NJ: Princeton University Press, 2006. 200 pp. $32.95. Journal of Cold War Studies, 18(4):261–263, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645936;   http://muse.jhu.edu/article/645936/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/261/700187/
jcws_r_00710.pdf.

Matray:2016:LBT

[Mat16b] James I. Matray. Larry Blomstedt, Truman, Congress, and Korea: The Politics of America’s First Undeclared War. Lexington, KY: University Press of Kentucky, 2016. xvi + 305 pp. $50.00. Journal of Cold War Studies, 18(4):212–214, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645915;   http://muse.jhu.edu/article/645915/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/212/700063/
jcws_r_00687.pdf.

Matray:2020:DCC

[Mat20a] James I. Matray. David Cheng Chang, The Hijacked War: The Story of Chinese POWs in the Korean War. Journal of Cold War Studies, 22(3):268–271, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/3/268/1860638/jcws_r_00962.pdf;
https://muse.jhu.edu/article/762906.

Matthiesen:2020:CWC

[Mat20b] Toby Matthiesen. The Cold War and the Communist Party of Saudi Arabia, 1975–1991. Journal of Cold War Studies, 22(3):32–62, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/3/32/1860592/jcws_a_00950.pdf;   https://muse.jhu.edu/
article/762891.

Mavrogordatos:2018:RVK

[Mav18] George Th. Mavrogordatos. Robert V. Keeley, The Colonels’ Coup and the American Embassy: a Diplomat’s View of the Breakdown of Democracy in Cold War Greece. Journal of Cold War Studies, 20 (4):243–245, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/718213;

https://direct.mit.edu/jcws/article-pdf/20/4/243/1969640/
jcws_r_00833.pdf.

Mawdsley:2008:OVK

[Maw08] Evan Mawdsley. Oleg V. Khlevniuk, The History of the Gulag: From Collectivization to the Great Terror. Journal of Cold War Studies, 10(3):181–183, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241522;   http://muse.jhu.edu/article/241522/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/181/697129/
jcws.2008.10.3.181.pdf.

Mayers:2013:GKC

[May13] David Mayers. Gaddis, Kennan, and the Cold War: an assessment of the biography. Journal of Cold War Studies, 15(4):161–169, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/534412;   http:
//muse.jhu.edu/article/534412/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/4/161/698323/jcws_a_00400.pdf. See

[Kra13b].

Maier:2021:AST

[MBR^{+}21] Charles S. Maier, Günter Bischof, Peter Ruggenthaler, Gerald Stourzh, and Wolfgang Mueller. The Austrian State Treaty and the Cold War: Contending perspectives. Journal of Cold War Studies, 23(4):211–245, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/4/211/1970091/jcws_c_01045.pdf;
https://muse.jhu.edu/article/835882.

Musoni-Chikede:2023:BRA

[MC23] Fungisai Musoni-Chikede. Book review: African Students in East Germany 1949–1975 by Sara Pugach. Journal of Cold War Studies, 25(4):253–255, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/25/4/253/2204979/jcws_r_01179.pdf.

McAllister:2007:LCG

[McA07] James McAllister. Lloyd C. Gardner and Ted Gittinger, eds., The Search for Peace in Vietnam 1964–1968. College Station: Texas A&M University Press, 2004. 406 pp. Journal of Cold War Studies, 9(2):166–168, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214752;   http://muse.jhu.edu/article/214752/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/166/696443/
jcws.2007.9.2.166.pdf.

McAllister:2008:LCG

[McA08] James McAllister. Lloyd C. Gardner and Ted Gittinger, eds., The Search for Peace in Vietnam 1964–1968. Journal of Cold War Studies, 10(1):121–122, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231585;   http://muse.jhu.edu/article/231585/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/121/700057/
jcws.2008.10.1.121.pdf.

McAllister:2014:CWC

[McA14] James McAllister. The Cambodian Wars: Clashing Armies and CIA Covert Operations by Kenneth Conboy (review). Journal of Cold War Studies, 16(4):250–251, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566670;   http://muse.jhu.edu/article/566670/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/250/698635/
jcws_r_00490.pdf.

McAdams:2022:BCR

[McA22] A. James McAdams. Between containment and rollback: The United States and the Cold War in Germany. Journal of Cold War Studies, 24(2):168–170, Spring 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/24/2/168/2013014/jcws_r_01083.pdf.

McBrady:2015:CPR

[McB15] Jared McBrady. The challenge of peace: Ronald Reagan, John Paul II, and the American bishops. Journal of Cold War Studies, 17(1):129–152, Winter 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/581233;   http://muse.jhu.edu/article/581233/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/1/129/700780/
jcws_a_00533.pdf.

McCormick:2014:FTI

[McC14] Evan McCormick. Freedom tide?: Ideology, politics, and the origins of democracy promotion in U.S. central America policy, 1980–1984. Journal of Cold War Studies, 16(4):60–109, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/566653;   http:
//muse.jhu.edu/article/566653/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/4/60/698627/jcws_a_00516.pdf.

McCauley:2020:KAU

[McC20] Martin McCauley. Kari Alenius, Unselfishly for Peace and Justice — And against Evil: The Rhetoric of the Great Powers in the UN Security Council, 1946–1956. Journal of Cold War Studies, 22(2):181–182, Spring 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/2/181/1860578/jcws_r_00946.pdf;
https://muse.jhu.edu/article/756820.

McDermott:2002:ACF

[McD02] Rose McDermott. Arms control and the first Reagan Administration: Belief-systems and policy choices. Journal of Cold War Studies, 4(4): 29–59, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9304;   http:/
/muse.jhu.edu/article/9304/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/4/29/695670/15203970260209509.pdf.

McDermott:2006:YGO

[McD06] Kevin McDermott. Yoram Gorlizki and Oleg Khlevniuk, Cold Peace: Stalin and the Soviet Ruling Circle, 1945–1953. Journal of Cold War Studies, 8(4):168–169, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/204627;   http://muse.jhu.edu/article/204627/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/4/168/700813/
jcws.2006.8.4.168.pdf.

McEnaney:2003:ONU

[McE03] Laura McEnaney. One Nation Underground: The Fallout Shelter in American Culture (review). Journal of Cold War Studies, 5(3):124– 126, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/45206;
 http://muse.jhu.edu/article/45206/pdf;   https://direct.
mit.edu/jcws/article-pdf/5/3/124/695915/jcws.2003.5.3.124.
pdf.

McGinn:1999:PCI

[McG99] John G. McGinn. The politics of collective inaction: NATO’s response to the Prague Spring. Journal of Cold War Studies, 1 (3):111–138, Fall 1999. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/9117;
 https://direct.mit.edu/jcws/article-pdf/1/3/111/695152/
152039799316976823.pdf.

McGranahan:2006:TCW

[McG06] Carole McGranahan. Tibet’s Cold War: The CIA and the Chushi Gangdrug Resistance, 1956–1974. Journal of Cold War Studies, 8(3):102–130, Summer 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/201239;   http://muse.jhu.edu/article/201239/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/3/102/696407/
jcws.2006.8.3.102.pdf.

McKenzie:2008:GCW

[McK08] Francine McKenzie. GATT and the Cold War: Accession debates, institutional development, and the Western Alliance, 1947– 1959. Journal of Cold War Studies, 10(3):78–109, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/241498;   http:
//muse.jhu.edu/article/241498/pdf;   https://direct.mit.
edu/jcws/article-pdf/10/3/78/697183/jcws.2008.10.3.78.pdf.

McKenzie:2015:IER

[McK15a] Francine McKenzie. International Economic Relations Since 1945 by Catherine R. Schenk (review). Journal of Cold War Studies, 17(2):143–145, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/582439;   http://muse.jhu.edu/article/582439/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/2/143/698690/
jcws_r_00537.pdf.

McKercher:2015:HFH

[McK15b] Asa McKercher. A helpful fixer in a hard place: Canadian mediation in the U.S. confrontation with Cuba. Journal of Cold War Studies, 17 (3):4–35, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/595070;
http://muse.jhu.edu/article/595070/pdf;   https://direct.
mit.edu/jcws/article-pdf/17/3/4/1102416/jcws_a_00551.pdf.

McMahon:2005:AVW

[McM05] Robert J. McMahon. America, the Vietnam War, and the World: Comparative and International Perspectives (review). Journal of Cold War Studies, 7(4):147–148, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188273;   http://muse.jhu.edu/article/188273/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/147/696299/
jcws.2005.7.4.147.pdf.

McMahon:2006:UPT

[McM06] Robert J. McMahon. U.S. policy toward South Asia and Tibet during the early Cold War. Journal of Cold War Studies, 8(3):131–144, Summer 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/201240;   http://muse.jhu.edu/article/201240/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/3/131/696411/
jcws.2006.8.3.131.pdf.

McMeekin:2008:SGP

[McM08] Sean McMeekin. Stanley G. Payne, The Spanish Civil War, the Soviet Union, and Communism. Journal of Cold War Studies, 10(1):125–127, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231586;   http://muse.jhu.edu/article/231586/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/125/700041/
jcws.2008.10.1.125.pdf.

McMahon:2010:MAL

[McM10] Robert J. McMahon. Mark Atwood Lawrence, The Vietnam War: a Concise International History. Journal of Cold War Studies, 12(3):159–160, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392796;   http://muse.jhu.edu/article/392796/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/159/697680/
jcws_r_00023.pdf.

McMillan:2011:JBM

[McM11] Priscilla Johnson McMillan. Jay Bergman, Meeting the Demands of Reason: The Life and Thought of Andrei Sakharov. Ithaca: Cornell University Press, 2009. 454 pp. Journal of Cold War Studies, 13(4):269–271, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457224;   http://muse.jhu.edu/article/457224/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/269/697817/
jcws_r_00136.pdf.

McMahon:2012:MJA

[McM12a] Robert J. McMahon. Matthew Jones, After Hiroshima: The United States, Race and Nuclear Weapons in Asia, 1945–1965. New York: Cambridge University Press, 2010. 502 pp. $110.00. Journal of Cold War Studies, 14(1):123–124, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467690;   http://muse.jhu.edu/article/467690/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/123/697907/
jcws_r_00201.pdf.

McMillan:2012:VZZ

[McM12b] Priscilla Johnson McMillan. Vladislav Zubok, Zhivago’s Children: The Last Russian Intelligentsia. Cambridge, MA: Belknap Press of Harvard University Press, 2009. $35.00. Journal of Cold War Studies, 14(1):118–120, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467687;   http://muse.jhu.edu/article/467687/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/118/697861/
jcws_r_00160.pdf.

McMahon:2014:OHC

[McM14] Robert J. McMahon. The Oxford Handbook of the Cold War edited by Richard H. Immerman and Petra Goedde (review). Journal of Cold War Studies, 16(4):253–255, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566672;   http://muse.jhu.edu/article/566672/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/253/698645/
jcws_r_00491.pdf.

McMahon:2016:SCZ

[McM16] Robert J. McMahon. Samantha Christiansen and Zachary A. Scarlett, eds., The Third World in the Global 1960s. New York: Berghahn, 2013. 223 pp. Journal of Cold War Studies, 18(2):188–189, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/620551;   http:
//muse.jhu.edu/article/620551/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/2/188/698846/jcws_r_00625.pdf.

McMahon:2021:TGE

[McM21a] Robert J. McMahon. Tonkin Gulf and the Escalation of the Vietnam War by Edwin E. Moïse (review). Journal of Cold War Studies, 23(2):177–179, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/2/177/1919933/jcws_r_01023.pdf;
https://muse.jhu.edu/article/794715.

McMillan:2021:SUS

[McM21b] Priscilla McMillan. Scientists under Surveillance: The FBI Files ed. by JPat Brown, B. C. D. Lipton and Michael Morisy (review). Journal of Cold War Studies, 23(2):167–168, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/23/2/167/1922978/jcws_r_
01014.pdf;   https://muse.jhu.edu/article/794716.

McMahon:2022:EAU

[McM22] Robert J. McMahon. The end of ambition: The United States and the Third World in the Vietnam Era. Journal of Cold War Studies, 24(3): 231–233, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/24/3/231/2041997/jcws_r_01094.pdf.

McPherson:2019:TCM

[McP19] Alan McPherson. Thomas C. Mann: President Johnson, the Cold War, and the Restructuring of Latin American Foreign Policy by Thomas Tunstall Allcock (review). Journal of Cold War Studies, 21 (3):262–264, Summer 2019. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
731291;   https://direct.mit.edu/jcws/article-pdf/21/3/262/
699653/jcws_r_00902.pdf.

Mehilli:2011:DSA

[Mëh11a] Elidor Mëhilli. Defying de-Stalinization: Albania’s 1956. Journal of Cold War Studies, 13(4):4–56, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457192;   http://muse.jhu.edu/article/457192/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/4/697788/jcws_
a_00169.pdf.

Mehta:2011:CCT

[Meh11b] Coleman Mehta. The CIA confronts the Tito–Stalin split, 1948– 1951. Journal of Cold War Studies, 13(1):101–145, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/424242;   http:

//muse.jhu.edu/article/424242/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/1/101/697767/jcws_a_00070.pdf.

Meijer:2015:BCS

[Mei15] Hugo Meijer. Balancing conflicting security interests: U.S. defense exports to China in the last decade of the Cold War. Journal of Cold War Studies, 17(1):4–40, Winter 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/581229;   http://muse.jhu.edu/article/581229/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/1/4/700776/jcws_
a_00529.pdf.

Melander:2001:NKC

[Mel01] Erik Melander. The Nagorno–Karabakh conflict revisited: Was the war inevitable? Journal of Cold War Studies, 3(2):48–75, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9192;   http:/
/muse.jhu.edu/article/9192/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/2/48/695372/152039701300373880.pdf.

Merten:2013:ATR

[Mer13] Ulrich Merten. Agnes Toth, Rückkehr nach Ungarn 1946–1950: Erlebnisberichte ungarndeutscher Vertriebener. Munich: Oldenburg Verlag, 2012. 389 pp. Journal of Cold War Studies, 15(3):209–211, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/522729;   http:
//muse.jhu.edu/article/522729/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/3/209/698349/jcws_r_00384.pdf.

Meyer:2011:BZE

[Mey11] David S. Meyer. Benjamin Ziemann, ed., Peace Movements in Western Europe, Japan and the USA during the Cold War. Essen: Klartext Verlag, 2007. 286 pp. 29.90. Journal of Cold War Studies, 13(4):237–239, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457210;   http://muse.jhu.edu/article/457210/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/237/697923/
jcws_r_00189.pdf.

Mickiewicz:2008:TCW

[Mic08] Ellen Mickiewicz. Thomas C. Wolfe, Governing Soviet Journalism: The Press and the Socialist Person after Stalin. Journal of Cold War Studies, 10(1):147–149, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231587;   http://muse.jhu.edu/article/231587/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/147/700077/
jcws.2008.10.1.147.pdf.

Mickiewicz:2011:EEE

[Mic11] Ellen Mickiewicz. Efficacy and evidence: Evaluating U.S. goals at the American national exhibition in Moscow, 1959. Journal of Cold War Studies, 13(4):138–171, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457196;   http://muse.jhu.edu/article/457196/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/138/697780/
jcws_a_00171.pdf.

Michaels:2017:WPC

[Mic17] Jeffrey H. Michaels. Waging “protracted conflict” behind the scenes: The Cold War activism of Frank R. Barnett. Journal of Cold War Studies, 19(1):70–98, Winter 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http:

//www.mitpressjournals.org/doi/full/10.1162/JCWS_a_00718;
 https://direct.mit.edu/jcws/article-pdf/19/1/70/699100/
jcws_a_00718.pdf.

Michelson:2021:LMR

[Mic21] Paul E. Michelson. The Language of the Moldovans: Romania, Russia, and Identity in an Ex-Soviet Republic by Matthew H. Ciscel (review). Journal of Cold War Studies, 23(3):246–248, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/23/3/246/1955773/
jcws_r_01035.pdf;   https://muse.jhu.edu/article/801915.

Middeke:2000:AAN

[Mid00] Michael Middeke. Anglo–American nuclear weapons cooperation after the Nassau Conference: The British policy of interdependence. Journal of Cold War Studies, 2(2):69–96, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9145;   http:/
/muse.jhu.edu/article/9145/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/2/69/695238/15203970051032318.pdf.

Middeke:2002:JFK

[Mid02] Michael Middeke. John F. Kennedy and Europe (review). Journal of Cold War Studies, 4(1):110–113, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9233;   http://muse.jhu.edu/article/9233/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/1/110/695548/
jcws.2002.4.1.110.pdf.

Midtgaard:2011:NSC

[Mid11] Kristine Midtgaard. National security and the choice of international humanitarian aid: Denmark and the Korean war, 1950–1953. Journal of Cold War Studies, 13(2):148–174, Spring 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/435669;   http://muse.jhu.edu/article/435669/
pdf;   https://direct.mit.edu/jcws/article-pdf/13/2/148/
700666/jcws_a_00107.pdf.

Mikkonen:2012:EES

[Mik12] Simo Mikkonen. Exploiting the exiles: Soviet Émigrés in U.S. Cold War strategy. Journal of Cold War Studies, 14(2):98–127, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/478777;
http://muse.jhu.edu/article/478777/pdf;   https://direct.
mit.edu/jcws/article-pdf/14/2/98/697979/jcws_a_00222.pdf.

Milward:2000:CAA

[Mil00] Alan S. Milward. A comment on the article by Andrew Moravcsik. Journal of Cold War Studies, 2(3):77–80, Fall 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9159;   http://muse.jhu.edu/article/
9159/pdf;   https://direct.mit.edu/jcws/article-pdf/2/3/
77/695276/15203970051032228.pdf. See [Mor00b, Mor00c].

Millett:2001:DSU

[Mil01] Allan R. Millett. In the Devil’s Shadow: U.N. Special Operations During the Korean War (review). Journal of Cold War Studies, 3(3): 122–124, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9209;   http:/
/muse.jhu.edu/article/9209/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/122/695474/jcws.2001.3.3.122.pdf.

Miller:2002:CRS

[Mil02] Steven E. Miller. Critical Reflections on Security and Change (review). Journal of Cold War Studies, 4(2):108–110, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9256;   http:/
/muse.jhu.edu/article/9256/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/2/108/695526/jcws.2002.4.2.108.pdf.

Milward:2004:QSE

[Mil04a] Alan S. Milward. A Question of Self-Esteem: The United States and the Cold War Choices in France and Italy, 1944–1958 (review). Journal of Cold War Studies, 6(3):156–158, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171169;   http:
//muse.jhu.edu/article/171169/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/3/156/696064/jcws.2004.6.3.156.pdf.

Milward:2004:GDV

[Mil04b] Alan S. Milward. Grand Designs and Visions of Unity: The Atlantic Powers and the Reorganization of Western Europe, 1955–1963 (review). Journal of Cold War Studies, 6(2):88–89, Spring 2004. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/53679;   http://muse.jhu.edu/
article/53679/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/2/88/695999/jcws.2004.6.2.88.pdf.

Milward:2006:RLM

[Mil06] Alan S. Milward. R. Laurence Moore and Maurizio Vaudagna, eds., The American Century in Europe. Journal of Cold War Studies, 8(2):170–172, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196287;   http://muse.jhu.edu/article/196287/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/170/700542/
jcws.2006.8.2.170.pdf.

Miller:2013:YMR

[Mil13] Jamie Miller. Yes, Minister: Reassessing South Africa’s intervention in the Angolan Civil War, 1975–1976. Journal of Cold War Studies, 15 (3):4–33, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/522707;
http://muse.jhu.edu/article/522707/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/3/4/698286/jcws_a_00368.pdf.

Millett:2014:CFI

[Mil14] Allan R. Millett. I Cannot Forget: Imprisoned in Korea, Accused at Home by Johnny Moore and Judith Fenner Gentry (review). Journal of Cold War Studies, 16(4):239–241, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/566665;   http://muse.jhu.edu/article/566665/
pdf;   https://direct.mit.edu/jcws/article-pdf/16/4/239/
698590/jcws_r_00485.pdf.

Millett:2015:NRS

[Mil15a] Allan R. Millett. Name, Rank, and Serial Number: Exploiting Korean War POWs at Home and Abroad by Charles S. Young (review). Journal of Cold War Studies, 17(2):156–158, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/582445;   http:
//muse.jhu.edu/article/582445/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/2/156/698704/jcws_r_00543.pdf.

Millett:2015:HMC

[Mil15b] Allan R. Millett. Hajimu Masuda, Cold War Crucible: The Korean Conflict and the Postwar World. Cambridge, MA: Harvard University Press, 2015. 388 pp. $39.95. Journal of Cold War Studies, 17(4):215–217, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606046;   http://muse.jhu.edu/article/606046/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/215/698834/
jcws_r_00614.pdf.

Milder:2017:GDC

[Mil17a] Stephen Milder. Gerrit Dworok and Christoph Weissmann, eds., 1968 und die 68er: Ereignisse, Wirkungen, und Kontroversen in der Bundesrepublik. Vienna: Böhlau Verlag, 2013. 227 pp. EUR 29.90. Journal of Cold War Studies, 19(3):252–253, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/668176;   http:
//muse.jhu.edu/article/668176/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/3/252/699292/jcws_r_00736.pdf.

Millett:2017:SRT

[Mil17b] Allan R. Millett. Stephen R. Taaffe, MacArthur’s Korean War Generals. Manhattan, KS: University of Kansas Press, 2016. 216 pp. $34.95. Journal of Cold War Studies, 19(3):236–239, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/668169;   http:
//muse.jhu.edu/article/668169/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/3/236/699264/jcws_r_00729.pdf.

Miles:2020:WSW

[Mil20] Simon Miles. The war scare that wasn’t: Able Archer 83 and the myths of the Second Cold War. Journal of Cold War Studies, 22(3): 86–118, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/22/3/86/1860596/jcws_a_00952.pdf;   https://muse.jhu.
edu/article/762893.

Millett:2021:PTC

[Mil21] Allan R. Millett. Passing the Test: Combat in Korea: April–June 1951 ed. by William T. Bowers and John T. Greenwood, Lexington, KY: University Press of Kentucky, 2011. 488 pp. $40.00. (review). Journal of Cold War Studies, 23(1):252–254, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/23/1/252/1911806/jcws_r_
01000.pdf;   https://muse.jhu.edu/article/788387.

Miles:2022:SCW

[Mil22] Simon Miles. The Second Cold War: Carter, Reagan, and the politics of foreign policy. Journal of Cold War Studies, 24(2):164–166, Spring 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/2/
164/2013024/jcws_r_01081.pdf.

Minkova:2019:SPP

[Min19] Kristina Minkova. Sovetskaya politika po rasshireniyu yuzhnykh granits: Stalin i azerbaidzhanskaya karta v bor’be za neft’ (1939–1945) by Jamil Hasanli (review). Journal of Cold War Studies, 21(1):196–200, Winter 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/723541;   https:
//direct.mit.edu/jcws/article-pdf/21/1/196/699789/jcws_r_
00862.pdf.

Miscamble:2023:BRK

[Mis23a] Wilson D. Miscamble. Book review: Kennan: a Life between Worlds by Frank Costigliola. Journal of Cold War Studies, 25(2):248–250, Spring 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/2/
248/2140962/jcws_r_01147.pdf.

Miscamble:2023:BRR

[Mis23b] Wilson D. Miscamble. Book review: Road to Surrender: Three Men and the Countdown to the End of World War II by Evan Thomas. Journal of Cold War Studies, 25(4):250–253, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/25/4/250/2204975/jcws_r_
01178.pdf.

Mitrovich:2004:IGK

[Mit04] Gregory Mitrovich. Interviews with George Kennan (review). Journal of Cold War Studies, 6(3):127–128, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/171170;   http://muse.jhu.edu/article/171170/
pdf;   https://direct.mit.edu/jcws/article-pdf/6/3/127/696029/
jcws.2004.6.3.127.pdf.

Mitrovich:2007:RPO

[Mit07] Gregory Mitrovich. Raymond P. Ojserkis, Beginnings of the Cold War Arms Race: The Truman Administration and the U.S. Arms Build-Up. Journal of Cold War Studies, 9(1):120–122, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/210093;   http:
//muse.jhu.edu/article/210093/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/1/120/696465/jcws.2007.9.1.120.pdf.

Modarressi:2017:PPU

[Mod17] Matin Modarressi. Philatelic propaganda: U.S. postage stamps during the Cold War. Journal of Cold War Studies, 19(3):196–201, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/668162;   http:
//muse.jhu.edu/article/668162/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/3/196/699232/jcws_a_00758.pdf.

Moeller:2003:KKP

[Moe03] Robert G. Moeller. Kalter Krieg und Propaganda: Die USA, der Kampf um die Weltmeinung und die ideelle Westbindung der Bundesrepublik Deutschland, 1945–1955, and: Cold War Politics in Postwar Germany (review). Journal of Cold War Studies, 5(2):93–96, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/40767;   http://

muse.jhu.edu/article/40767/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/2/93/700432/jcws.2003.5.2.93.pdf.

Moeller:2007:MDH

[Moe07] Robert G. Moeller. Matthew D. Hockenos, A Church Divided: German Protestants Confront the Nazi Past. Journal of Cold War Studies, 9(3):155–157, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219464;   http://muse.jhu.edu/article/219464/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/155/696648/
jcws.2007.9.3.155.pdf.

Moe:2018:KSG

[Moe18] Richard Moe. Kathryn Smith, The Gatekeeper: Missy LeHand, FDR, and the Untold Story of the Partnership That Defined a Presidency. Journal of Cold War Studies, 20(4):239–241, Fall 2018. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718211;   https://direct.mit.
edu/jcws/article-pdf/20/4/239/1969663/jcws_r_00830.pdf.

Moise:2003:REB

[Moi03] Edwin E. Moise. Review essay: Better late than never? The delayed debate over the costs of Vietnam. Journal of Cold War Studies, 5(2): 73–77, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/40768;
 http://muse.jhu.edu/article/40768/pdf;   https://direct.
mit.edu/jcws/article-pdf/5/2/73/700394/152039703763336480.
pdf.

Moise:2010:MPB

[Moï10] Edwin Moïse. Mark Philip Bradley and Marilyn B. Young, eds., Making Sense of the Vietnam Wars: Local, National, and Transnational Perspectives. Journal of Cold War Studies, 12(4):197–199, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/403809;   http:
//muse.jhu.edu/article/403809/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/4/197/697578/jcws_r_00061.pdf.

Moise:2011:WFU

[Moï11] Edwin Moïse. War without Fronts: The USA in Vietnam (review). Journal of Cold War Studies, 13(3):190–196, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/447462;   http://
muse.jhu.edu/article/447462/pdf.

Moise:2014:GJV

[Moi14] Edwin E. Moise. George J. Veith, Black April: The Fall of South Vietnam, 1973–75. New York: Encounter Books, 2012. xxvi + 587 pp. $29.95. Journal of Cold War Studies, 16(1):251–253, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543750;   http:
//muse.jhu.edu/article/543750/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/1/251/698984/jcws_r_00441.pdf.

Moise:2015:WSV

[Moi15] Edwin Moise. Why South Vietnam Fell by Anthony James Joes (review). Journal of Cold War Studies, 17(3):295–297, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595112;   http:
//muse.jhu.edu/article/595112/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/295/1102446/jcws_r_00558.pdf.

Moll:2016:GBJ

[Mol16] Martin Moll. Günter Bischof, Jason Dawsey, and Bernhard Fetz, eds. The Life and Work of Günther Anders: Émigré, Iconoclast, Philosopher, Man of Letters. Innsbruck: StudienVerlag, 2014. 202 pp. EUR 29.90. Journal of Cold War Studies, 18(4):243–245, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/645928;   http:
//muse.jhu.edu/article/645928/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/4/243/700164/jcws_r_00701.pdf.

Moore:2023:BRH

[Moo23] Brenda L. Moore. Book review: Her Cold War: Women in the U.S. Military 1945–1980 by Tanya L. Roth. Journal of Cold War Studies, 25(1):224–227, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/1/224/2073757/jcws_r_01133.pdf.

Moravcsik:2000:BGG

[Mor00a] Andrew Moravcsik. Beyond grain and grandeur: an answer to critics and an agenda for future research. Journal of Cold War Studies, 2(3): 117–142, Fall 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9161;   http:/
/muse.jhu.edu/article/9161/pdf;   https://direct.mit.edu/

jcws/article-pdf/2/3/117/695286/15203970051032264.pdf. See

[Mor00b, Mor00c].

Moravcsik:2000:GBGa

[Mor00b] Andrew Moravcsik. De Gaulle between grain and grandeur: The political economy of French EC policy, 1958–1970 (Part 1). Journal of Cold War Studies, 2(2):3–43, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9146;   http://muse.jhu.edu/article/
9146/pdf;   https://direct.mit.edu/jcws/article-pdf/2/2/
3/695228/15203970051032291.pdf. See [Mor00c], comments [Hof00,

Kee00, Mil00, Tra00], and reply [Mor00a].

Moravcsik:2000:GBGb

[Mor00c] Andrew Moravcsik. De Gaulle between grain and grandeur: The political economy of French EC policy, 1958–1970 (Part 2). Journal of Cold War Studies, 2(3):4–68, Fall 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9160;   http://muse.jhu.edu/article/
9160/pdf;   https://direct.mit.edu/jcws/article-pdf/2/3/
4/695280/15203970051032192.pdf. See [Mor00c], comments [Hof00,

Kee00, Mil00, Tra00, LSvdV04, Mor12], and reply [Mor00a].

Moran:2006:PAB

[Mor06] Daniel Moran. Paul Aussaresses, The Battle of the Casbah: Terrorism and Counter-Terrorism in Algeria, 1955–1957. Journal of Cold War Studies, 8(2):172–174, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196288;   http://muse.jhu.edu/article/196288/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/172/700552/
jcws.2006.8.2.172.pdf.

Moravcsik:2012:CGE

[Mor12] Andrew Moravcsik. Charles de Gaulle and Europe: The new revisionism. Journal of Cold War Studies, 14(1):53–77, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/467679;   http:
//muse.jhu.edu/article/467679/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/1/53/697821/jcws_a_00192.pdf. See

[Mor00b, Mor00c].

Moran:2013:IFP

[Mor13] Christopher Moran. Ian Fleming and the public profile of the CIA. Journal of Cold War Studies, 15(1):119–146, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/505771;   http:
//muse.jhu.edu/article/505771/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/1/119/699970/jcws_a_00310.pdf.

Moullec:2002:RTS

[Mou02] Gaël-Georges Moullec. The Road to Terror: Stalin and the Self-Destruction of the Bolsheviks, 1932–1939 (review). Journal of Cold War Studies, 4(3):160–161, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9282;   http://muse.jhu.edu/article/9282/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/3/160/700266/
jcws.2002.4.3.160.pdf.

Moullec:2003:NTS

[Mou03] Gaël-Georges Moullec. A Normal Totalitarian Society: How the Soviet Union Functioned and How It Collapsed (review). Journal of Cold War Studies, 5(3):154–156, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45207;   http://muse.jhu.edu/
article/45207/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/154/695931/jcws.2003.5.3.154.pdf.

Moyar:2011:HSA

[Moy11] Mark Moyar. Herbert Y. Schandler, America in Vietnam: The War That Couldn’t Be Won. Lanham, MD: Rowman & Littlefield, 2009. Journal of Cold War Studies, 13(1):231–234, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/424251;   http:
//muse.jhu.edu/article/424251/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/1/231/697711/jcws_r_00082.pdf.

MacFarquhar:2008:RC

[MS08] Roderick MacFarquhar and Michael Schoenhals. Response to the commentaries. Journal of Cold War Studies, 10(2):126–130, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/237153;   http://
muse.jhu.edu/article/237153/pdf.

Mastny:2021:SFE

[MSPN21] Vojtech Mastny, Vít Smetana, Vladimir Pechatnov, and Norman M. Naimark. Stalin and the fate of Europe after 1945: Contending perspectives. Journal of Cold War Studies, 23(3):208–231, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/23/3/208/1955779/
jcws_c_01028.pdf;   https://muse.jhu.edu/article/801909.

Mueller:2005:GWR

[Mue05] John E. Mueller. The Gulf War of 1991 Reconsidered (review). Journal of Cold War Studies, 7(1):191–193, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/179568;   http:
//muse.jhu.edu/article/179568/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/1/191/700484/jcws.2005.7.1.191.pdf.

Mueller:2010:HMA

[Mue10] Wolfgang Mueller. Horst Möller and Aleksandr Tschubarjan, eds., Mitteilungen der Gemeinsamen Kommission für die Erforschung der jüngeren Geschichte der deutsch-russischen Beziehungen/Soobshcheniya Sovmestnoi komissii po izucheniyu noveishei istorii rossiiskogermanskikh otnoshenii, Vol. 2. Journal of Cold War Studies, 12(3):144–147, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392789;   http://muse.jhu.edu/article/392789/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/144/697634/
jcws_r_00016.pdf.

Mueller:2011:RRD

[Mue11] Wolfgang Mueller. Recognition in return for Détente?: Brezhnev, the EEC, and the Moscow treaty with West Germany, 1970–1973. Journal of Cold War Studies, 13(4):79–100, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457194;   http:
//muse.jhu.edu/article/457194/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/4/79/697792/jcws_a_00167.pdf.

Mueller:2013:RFS

[Mue13] Tim B. Mueller. The Rockefeller Foundation, the social sciences, and the humanities in the Cold War. Journal of Cold War Studies, 15(3):108–135, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522711;   http://muse.jhu.edu/article/522711/pdf;

 https://direct.mit.edu/jcws/article-pdf/15/3/108/698290/
jcws_a_00372.pdf.

Mueller:2016:UPN

[Mue16] Wolfgang Mueller. The USSR and permanent neutrality in the Cold War. Journal of Cold War Studies, 18(4):148–179, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/645911;   http:
//muse.jhu.edu/article/645911/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/4/148/700071/jcws_a_00683.pdf.

Muehlenbeck:2017:JMA

[Mue17] Philip E. Muehlenbeck. Jamie Miller, An African Volk: The Apartheid Regime and Its Search for Survival. Journal of Cold War Studies, 19(4):279–281, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683694;   http://muse.jhu.edu/article/683694/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/279/699478/
jcws_r_00794.pdf.

Mulhern:2023:BRC

[Mul23] Matt Mulhern. Book review: Cold War Liberation: The Soviet Union and the Collapse of the Portuguese Empire in Africa, 1961–1975 by Natalia Telepneva. Journal of Cold War Studies, 25(1):219–222, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/1/
219/2073759/jcws_r_01131.pdf.

Munteanu:2010:WLB

[Mun10] Mircea Munteanu. When the levee breaks: The impact of the Sino–Soviet split and the invasion of Czechoslovakia on Romanian– Soviet relations, 1967–1970. Journal of Cold War Studies, 12(1): 43–61, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382486;
http://muse.jhu.edu/article/382486/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/1/43/1627461/jcws.2010.12.1.
43.pdf.

Munteanu:2012:HFR

[Mun12] Mircea Munteanu. Over the hills and far away: Romania’s attempts to mediate the start of U.S.–North Vietnamese negotiations, 1967–1968. Journal of Cold War Studies, 14(3):64–96, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/490175;
http://muse.jhu.edu/article/490175/pdf;   https://direct.
mit.edu/jcws/article-pdf/14/3/64/698047/jcws_a_00248.pdf.

Munton:2021:SGT

[Mun21] Don Munton. The Silent Guns of Two Octobers: Kennedy and Khrushchev Play the Double Game by Theodore Voorhees (review). Journal of Cold War Studies, 23(4):248–251, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/23/4/248/1970077/jcws_r_
01047.pdf;   https://muse.jhu.edu/article/835884.

Murray:2002:AAS

[Mur02] Donette Murray. Anglo–American Strategic Relations and the French Problem, 1960–1963: a Troubled Partnership (review). Journal of Cold War Studies, 4(2):118–120, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9257;   http://muse.jhu.edu/article/9257/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/2/118/695552/
jcws.2002.4.2.118.pdf.

Murphy:2004:BGP

[Mur04] Philip Murphy. British Government Policy and Decolonisation 1945– 1963: Scrutinising the Official Mind (review). Journal of Cold War Studies, 6(3):154–155, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/171171;   http://muse.jhu.edu/article/171171/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/3/154/696091/
jcws.2004.6.3.154.pdf.

Muravchik:2005:JEH

[Mur05] Joshua Muravchik. John Earl Haynes and Harvey Klehr, In Denial: Historians, Communism and Espionage (review). Journal of Cold War Studies, 7(1):213–215, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/179569;   http://muse.jhu.edu/article/179569/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/1/213/700502/
jcws.2005.7.1.213.pdf.

Murray:2007:MFH

[Mur07] Donette Murray. Michael F. Hopkins, Michael D. Kandiah, and Gillian Staerck, eds., Cold War Britain, 1945–1964: New Perspectives. Journal of Cold War Studies, 9(1):132–133, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/210094;   http:
//muse.jhu.edu/article/210094/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/1/132/696540/jcws.2007.9.1.132.pdf.

Mylonas:2012:CMA

[Myl12] Harris Mylonas. Carole McGranahan, Arrested Histories: Tibet, the CIA, and Memories of a Forgotten War. Durham, NC: Duke University Press, 2010. Journal of Cold War Studies, 14(2):154–156, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/478780;   http:
//muse.jhu.edu/article/478780/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/2/154/697989/jcws_r_00225.pdf.

Mylonas:2015:DDR

[Myl15] Harris Mylonas. Diplomacy and Displacement: Reconsidering the Turco–Greek Exchange of Populations, 1922–1934 by Onur Yildirim (review). Journal of Cold War Studies, 17(3):268–270, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595098;   http:
//muse.jhu.edu/article/595098/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/268/1102409/jcws_r_00510.pdf.

Nagy:2017:SHA

[Nag17] András Nagy. Shattered hopes amid violent repression: The Hungarian Revolution and the United Nations (part 1). Journal of Cold War Studies, 19(4):42–73, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683665;   http://muse.jhu.edu/article/683665/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/42/699398/
jcws_a_00764.pdf.

Nagy:2018:SHA

[Nag18] András Nagy. Shattered hopes amid violent repression: The 1956 Hungarian Revolution and the United Nations (part 2). Journal of Cold War Studies, 20(4):127–153, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718201;   https://direct.mit.
edu/jcws/article-pdf/20/4/127/1969620/jcws_a_00840.pdf.

Naimark:2002:DSO

[Nai02] Norman M. Naimark. Denazification in Soviet-Occupied Germany: Brandenburg 1945–1948 (review). Journal of Cold War Studies, 4(3):140–142, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9283;   http://muse.jhu.edu/article/9283/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/3/140/700246/
jcws.2002.4.3.140.pdf.

Naimark:2004:GSE

[Nai04] Norman M. Naimark. Gleichschaltung unter Stalin? Die Entwicklung der Parteien in ostlichen Europa, 1944–1949 (review). Journal of Cold War Studies, 6(4):170–172, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/174717;   http://muse.jhu.edu/article/174717/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/4/170/696103/
jcws.2004.6.4.170.pdf.

Naimark:2008:JZE

[Nai08a] Norman Naimark. Jürgen Zarusky, ed., Stalin und die Deutschen: Neue Beiträge der Forschung. Journal of Cold War Studies, 10(1): 156–159, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/231588;
http://muse.jhu.edu/article/231588/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/1/156/700110/jcws.2008.10.1.
156.pdf.

Naimark:2008:SCM

[Nai08b] Norman M. Naimark. Stefan Creuzberger and Manfred Görtemaker, eds., Gleichschaltung unter Stalin? Die Entwicklung der Parteien in östlichen Europa, 1944–1949. Journal of Cold War Studies, 10(3):203–205, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241533;   http://muse.jhu.edu/article/241533/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/203/697121/
jcws.2008.10.3.203.pdf.

Naimark:2010:BAP

[Nai10a] Norman M. Naimark. Balazs Apor, Peter Apor, and E. A. Rees, eds., The Sovietization of Eastern Europe: New Perspectives on the Postwar Period. Journal of Cold War Studies, 12(2):126– 127, Spring 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382208;

http://muse.jhu.edu/article/382208/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/2/126/697493/jcws.2010.12.2.
126.pdf.

Naimark:2010:WGD

[Nai10b] Norman M. Naimark. Wladimir Gelfand, Deutschland-Tagebuch 1945–1946: Aufzeichnungen eines Rotarmisten, trans. by Anja Lutter and Harmut Schröder, ed. by Elke Scherstjanoi. Journal of Cold War Studies, 12(1):191–193, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/382506;   http://muse.jhu.edu/article/382506/
pdf;   https://direct.mit.edu/jcws/article-pdf/12/1/191/
1627439/jcws.2010.12.1.191.pdf.

Naimark:2012:RC

[Nai12] Norman Naimark. Reply to the commentaries. Journal of Cold War Studies, 14(3):183–189, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/490177;   http://muse.jhu.edu/article/490177/

pdf. See [GRS^{+}12].

Naimark:2016:AJR

[Nai16] Norman Naimark. Alfred J. Rieber, Stalin and the Struggle for Supremacy in Eurasia. New York: Cambridge University Press, 2015. 420 pp. Journal of Cold War Studies, 18(3):219–221, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/632299;   http:
//muse.jhu.edu/article/632299/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/3/219/699049/jcws_r_00666.pdf.

Naimark:2021:KVV

[Nai21] Norman M. Naimark. Komandirovka vo Vlast’ by A. V. Minzhurenko (review). Journal of Cold War Studies, 23(4):253–255, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/23/4/253/1970105/
jcws_r_01049.pdf;   https://muse.jhu.edu/article/835886.

Naimark:2022:CWW

[Nai22] Norman M. Naimark. The Cold War: A World History by Odd Arne Westad (review). Journal of Cold War Studies, 24(1): 254–258, Winter 2022. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL https://direct.mit.edu/jcws/

article-pdf/24/1/254/1980863/jcws_r_01063.pdf;   https://
muse.jhu.edu/article/843085.

Namikas:2013:FRV

[Nam13] Lise Namikas. Frank R. Villafana, Cold War in the Congo: The Confrontation of Cuban Military Forces, 1960–1967. New Brunswick, NJ: Transaction Publishers, 2009. 224 pp. $29.95. Journal of Cold War Studies, 15(3):207–209, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522728;   http://muse.jhu.edu/article/522728/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/207/698345/
jcws_r_00383.pdf.

Namikas:2016:EGB

[Nam16] Lise Namikas. Emmanuel Gerard and Bruce Kuklick, Death in the Congo: Murdering Patrice Lumumba. Cambridge, MA: Harvard University Press, 2015. 276 pp. $29.95. Journal of Cold War Studies, 18(3):212–214, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632295;   http://muse.jhu.edu/article/632295/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/212/699003/
jcws_r_00647.pdf.

Nation:2004:EWG

[Nat04] R. Craig Nation. Enemies within the Gates? The Comintern and the Stalinist Repression, 1934–1939 (review). Journal of Cold War Studies, 6(4):174–177, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/174718;   http://muse.jhu.edu/article/174718/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/4/174/696123/
jcws.2004.6.4.174.pdf.

Nau:2004:BEA

[Nau04] Henry R. Nau. Between Empire and Alliance: America and Europe during the Cold War (review). Journal of Cold War Studies, 6(4):149–151, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/174719;   http://muse.jhu.edu/article/174719/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/4/149/696066/
jcws.2004.6.4.149.pdf.

Nehring:2011:WKC

[Neh11] Holger Nehring. Wolfram Kaiser, Christian Democracy and the Origins of European Union. New York: Cambridge University Press, 2007. 388 pp. £55.00. Journal of Cold War Studies, 13(4):242–244, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457212;   http:
//muse.jhu.edu/article/457212/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/4/242/697782/jcws_r_00124.pdf.

Nehring:2021:ASM

[Neh21] Christopher Nehring. Active and sharp measures: Cooperation between the Soviet KGB and Bulgarian State Security. Journal of Cold War Studies, 23(4):3–33, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.

mit.edu/jcws/article-pdf/23/4/3/1970095/jcws_a_01038.pdf;
 https://muse.jhu.edu/article/835876.

Nehring:2022:BSS

[Neh22] Christopher Nehring. Bulgaria as the sixteenth Soviet republic?: Todor Zhivkov’s proposals to join the USSR. Journal of Cold War Studies, 24(2):29–45, Spring 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/24/2/29/2013020/jcws_a_01076.pdf.

Neidhart:2000:IRR

[Nei00] Christoph Neidhart. Interpreting the Russian Revolution: The Language and Symbols of 1917 (review). Journal of Cold War Studies, 2(2):122–125, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9147;   http://muse.jhu.edu/article/9147/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/2/122/695269/
jwcs.2000.2.2.122.pdf.

Neidhart:2007:ZAC

[Nei07] Christoph Neidhart. Zlatko Anguelov, Communism and the Remorse of an Innocent Victimizer. Journal of Cold War Studies, 9(3):186–187, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219465;   http://muse.jhu.edu/article/219465/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/186/696664/
jcws.2007.9.3.186.pdf.

Neidhart:2008:FCC

[Nei08] Christoph Neidhart. Frederick C. Corney, Telling October: Memory and the Making of the Bolshevik Revolution. Journal of Cold War Studies, 10(3):193–195, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241528;   http://muse.jhu.edu/article/241528/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/193/697171/
jcws.2008.10.3.193.pdf.

Neidhart:2010:AYE

[Nei10a] Christoph Neidhart. Alexei Yurchak, Everything Was Forever, Until It Was No More: The Last Soviet Generation. Journal of Cold War Studies, 12(1):204–206, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/382507;   http://muse.jhu.edu/article/382507/
pdf;   https://direct.mit.edu/jcws/article-pdf/12/1/204/
1627441/jcws.2010.12.1.204.pdf.

Neier:2010:JWG

[Nei10b] Aryeh Neier. Jeffrey N. Wasserstrom, Greg Grandin, Lynn Hunt, and Marilyn B. Young, eds., Human Rights and Revolutions, 2nd ed. Journal of Cold War Studies, 12(4):182–184, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/403801;   http:
//muse.jhu.edu/article/403801/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/4/182/697542/jcws_r_00053.pdf.

Neier:2014:RFU

[Nei14] Aryeh Neier. Rosa Freedman, The United Nations Human Rights Council: a Critique and Early Assessment. New York: Routledge, 2012. 330 pp. $145.00. Journal of Cold War Studies, 16(1):212–214, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543732;   http:
//muse.jhu.edu/article/543732/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/1/212/698878/jcws_r_00410.pdf.

Neu:2003:REE

[Neu03] Charles E. Neu. Review essay: Efforts to make sense of the Vietnam War. Journal of Cold War Studies, 5(2):68–72, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40769;   http://muse.jhu.edu/
article/40769/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/68/700374/152039703763336471.pdf.

Neufeld:2011:WBD

[Neu11] Michael J. Neufeld. Wayne Biddle, Dark Side of the Moon: Wernher von Braun, the Third Reich, and the Space Race. New York: W. W. Norton, 2009. xix + 220 pp. $25.95. Journal of Cold War Studies, 13(4):223–224, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457203;   http://muse.jhu.edu/article/457203/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/223/697853/
jcws_r_00182.pdf.

Neuburger:2023:BRS

[Neu23] Mary Neuburger. Book review: Sugarland: The Transformation of the Countryside in Communist Albania by Artan Hoxha. Journal of Cold War Studies, 25(4):267–269, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/4/267/2205004/jcws_r_01185.pdf.

Newton:2007:GKD

[New07] Verne W. Newton. Gary Kern, A Death in Washington: Walter G. Krivitsky and the Stalin Terror. Journal of Cold War Studies, 9(1):109–111, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210095;   http://muse.jhu.edu/article/210095/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/109/696483/
jcws.2007.9.1.109.pdf.

Nichols:2012:DFS

[Nic12] Thomas M. Nichols. David F. Schmitz, Brent Scowcroft: Internationalism and Post-Vietnam War American Foreign Policy. Lanham, MD: Rowman & Littlefield, 2011. 214 pp. Journal of Cold War Studies, 14(4):227–228, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500337;   http://muse.jhu.edu/article/500337/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/227/698133/
jcws_r_00284.pdf.

Niedhart:2003:DDE

[Nie03] Gottfried Niedhart. Dealing with the Devil: East Germany, Detente, and Ostpolitik, 1969–1973 (review). Journal of Cold War Studies, 5(2):96–98, Spring 2003. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
40770;   http://muse.jhu.edu/article/40770/pdf;   https:/

/direct.mit.edu/jcws/article-pdf/5/2/96/700460/jcws.2003.
5.2.96.pdf.

Niedhart:2008:AHE

[Nie08] Gottfried Niedhart. Arne Hofmann, The Emergence of Détente in Europe: Brandt, Kennedy and the Formation of Ostpolitik. Journal of Cold War Studies, 10(4):168–170, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/252667;   http://muse.jhu.edu/article/252667/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/4/168/697057/
jcws.2008.10.4.168.pdf.

Niedhart:2010:GRH

[Nie10a] Gottfried Niedhart. Gerd-Rainer Horn, The Spirit of ’68: Rebellion in Western Europe and North America, 1956–1976. Journal of Cold War Studies, 12(1):183–184, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/382508;   http://muse.jhu.edu/article/382508/
pdf;   https://direct.mit.edu/jcws/article-pdf/12/1/183/
1627435/jcws.2010.12.1.183.pdf.

Nielsen:2010:GBP

[Nie10b] Claus Kjersgaard Nielsen. Gary Baines and Peter Vale, eds., Beyond the Border War: New Perspectives on Southern Africa’s Late-Cold War Conflicts. Journal of Cold War Studies, 12(1):173– 175, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382509;
http://muse.jhu.edu/article/382509/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/1/173/1627421/jcws.2010.12.1.
173.pdf.

Niebuhr:2011:NYA

[Nie11] Robert Niebuhr. Nonalignment as Yugoslavia’s answer to bloc politics. Journal of Cold War Studies, 13(1):146–179, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/424243;   http:
//muse.jhu.edu/article/424243/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/1/146/697719/jcws_a_00072.pdf.

Niedhart:2012:PVO

[Nie12] Gottfried Niedhart. Poul Villaume and Odd Arne Westad, eds., Perforating the Iron Curtain: European Détente, Transatlantic Relations, and the Cold War, 1965–1985. Copenhagen: Museum Tusculum Press, 2010. 272 pp. Journal of Cold War Studies, 14(2):175–177, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/478790;   http://muse.jhu.edu/article/478790/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/2/175/698001/
jcws_r_00233.pdf.

Niedhart:2016:ICG

[Nie16a] Gottfried Niedhart. Introduction: CSCE, the German Question, and the Eastern Bloc. Journal of Cold War Studies, 18(3):3–13, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/632277;
http://muse.jhu.edu/article/632277/pdf;   https://direct.
mit.edu/jcws/article-pdf/18/3/3/699018/jcws_e_00651.pdf.

Niedhart:2016:OTT

[Nie16b] Gottfried Niedhart. Ostpolitik: Transformation through communication and the quest for peaceful change. Journal of Cold War Studies, 18(3):14–59, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632278;   http://muse.jhu.edu/article/632278/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/14/699012/
jcws_a_00652.pdf.

Nielsen:2021:IES

[Nie21] Christian Axboe Nielsen. Imprisoning “Enemies of the state” in a Communist dictatorship: The record of Tito’s Yugoslavia, 1945–1953. Journal of Cold War Studies, 23(4):124–152, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/23/4/124/1970089/jcws_a_
01041.pdf;   https://muse.jhu.edu/article/835879.

Nilsson:2019:USN

[Nil19] Mikael Nilsson. The United States and neutral countries in Europe, 1945–1991. Journal of Cold War Studies, 21(4):208–230, Fall 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
21/4/208/699728/jcws_a_00912.pdf;   https://muse.jhu.edu/
article/736843.

Ninkovich:2003:CWC

[Nin03] Frank A. Ninkovich. Cold War Constructions: The Political Culture of United States Imperialism, 1945–1966, and: To Lead the Free World: American Nationalism and the Cultural Roots of the Cold War (review). Journal of Cold War Studies, 5(2):78–81, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40771;   http://muse.jhu.edu/
article/40771/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/78/700400/152039703763336499.pdf.

Njolstad:2002:CAI

[Njø02] Olav Njølstad. The Carter Administration and Italy: Keeping the Communists out of power without interfering. Journal of Cold War Studies, 4(3):56–94, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9284;   http://muse.jhu.edu/article/
9284/pdf;   https://direct.mit.edu/jcws/article-pdf/4/3/
56/700226/152039702320201076.pdf.

Nuenlist:2012:RC

[NLM12] Christian Nuenlist, Anna Locher, and Garret Martin. Reply to the commentaries. Journal of Cold War Studies, 14(1):93–95, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/467682;
http://muse.jhu.edu/article/467682/pdf;   https://direct.
mit.edu/jcws/article-pdf/14/1/93/697837/jcws_a_00195.pdf.

Nordlander:2000:SYS

[Nor00] David J. Nordlander. The Stalin Years: The Soviet Union, 1929– 1953 (review). Journal of Cold War Studies, 2(1):143–145, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9132;   http:/
/muse.jhu.edu/article/9132/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/1/143/695214/jcws.2000.2.1.143.pdf.

Norris:2005:HSH

[Nor05] Robert S. Norris. Hanford Site Historic District: History of the Plutonium Production Facilities, 1943–1990, and: Hanford: a Conversation about Nuclear Waste and Cleanup (review). Journal of Cold War Studies, 7(3):164–165, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185115;   http://muse.jhu.edu/article/185115/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/164/1102565/
jcws.2005.7.3.164.pdf.

Norris:2007:HGP

[Nor07] Robert S. Norris. Hugh Gusterson, People of the Bomb: Portraits of America’s Nuclear Complex. Minneapolis: University of Minnesota Press, 2004. 312 pp. $19.95. Journal of Cold War Studies, 9(2):152–154, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214753;   http://muse.jhu.edu/article/214753/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/152/696423/
jcws.2007.9.2.152.pdf.

Nordell:2008:MWL

[Nor08a] John R. Nordell, Jr. Martin Windrow, The Last Valley: Dien Bien Phu and the French Defeat in Vietnam. Journal of Cold War Studies, 10(1):123–124, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231589;   http://muse.jhu.edu/article/231589/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/123/700087/
jcws.2008.10.1.123.pdf.

Norris:2008:GJD

[Nor08b] Robert S. Norris. Gerard J. DeGroot, The Bomb: a Life. Journal of Cold War Studies, 10(4):142–144, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/252658;   http://muse.jhu.edu/article/252658/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/4/142/697051/
jcws.2008.10.4.142.pdf.

Norris:2009:JHI

[Nor09] Robert S. Norris. Jon Hunner, Inventing Los Alamos: The Growth of an Atomic Community. Journal of Cold War Studies, 11(4): 149–151, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/366259;
http://muse.jhu.edu/article/366259/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/4/149/1627617/jcws.2009.11.4.
149.pdf.

Norris:2010:MGS

[Nor10] Robert S. Norris. Michael S. Goodman, Spying on the Nuclear Bear: Anglo-American Intelligence and the Soviet Bomb. Journal of Cold War Studies, 12(3):199–201, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392813;   http://muse.jhu.edu/article/392813/pdf;

 https://direct.mit.edu/jcws/article-pdf/12/3/199/697698/
jcws_r_00041.pdf.

Norris:2015:DBS

[Nor15] Robert S. Norris. David Burke, The Spy Who Came in from the Co-op: Melita Norwood and the Ending of Cold War Espionage. Woodbridge, UK: The Boydell Press, 2008. ix + 209 pp. $34.95. Journal of Cold War Studies, 17(4):175–178, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606026;   http://muse.jhu.edu/article/606026/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/175/698714/
jcws_r_00581.pdf.

Norton:2020:AMB

[Nor20] Richard J. Norton. Andrew Marble, Boy on the Bridge: The Story of John Shalikashvili’s American Success. Journal of Cold War Studies, 22(3):264–266, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/3/264/1860624/jcws_r_00960.pdf;
https://muse.jhu.edu/article/762904.

Nuenlist:2005:BWC

[Nue05] Christian Nuenlist. The Berlin Wall Crisis: Perspectives on Cold War Alliances (review). Journal of Cold War Studies, 7(4):162–164, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188274;   http://muse.jhu.edu/article/188274/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/162/696317/
jcws.2005.7.4.162.pdf.

Nuenlist:2006:CMK

[Nue06] Christian Nuenlist. Christof Münger, Kennedy, die Berliner Mauer und die Kubakrise: Die westliche Allianz in der Zerreissprobe, 1961–63. Journal of Cold War Studies, 8(2):159–161, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/196289;   http:
//muse.jhu.edu/article/196289/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/2/159/700540/jcws.2006.8.2.159.pdf.

Nuenlist:2007:AWD

[Nue07] Christian Nuenlist. Andreas W. Daum, Kennedy in Berlin: Politik, Kultur und Emotionen im Kalten Krieg. Journal of Cold War Studies, 9(1):124–126, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210096;   http://muse.jhu.edu/article/210096/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/124/696494/
jcws.2007.9.1.124.pdf.

Nunan:2022:DGR

[Nun22] Timothy Nunan. “Doomed to good relations”: The USSR, the Islamic Republic of Iran, and anti-imperialism in the 1980s. Journal of Cold War Studies, 24(1):39–77, Winter 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/24/1/39/1980873/jcws_a_
01056.pdf;   https://muse.jhu.edu/article/843077.

Nuti:2002:MIA

[Nut02a] Leopoldo Nuti. Modernization as Ideology. American Social Science and “Nation Building” in the Kennedy Era (review). Journal of Cold War Studies, 4(1):99–101, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9234;   http://muse.jhu.edu/article/
9234/pdf;   https://direct.mit.edu/jcws/article-pdf/4/1/
99/695593/jcws.2002.4.1.99.pdf.

Nuti:2002:ICW

[Nut02b] Leopoldo Nuti. Italy and the Cold War: General introduction. Journal of Cold War Studies, 4(3):3–4, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9285;   http://muse.jhu.edu/article/
9285/pdf;   https://direct.mit.edu/jcws/article-pdf/4/3/
3/700224/152039702320201049.pdf.

Nuti:2002:USI

[Nut02c] Leopoldo Nuti. The United States, Italy, and the opening to the left, 1953–1963. Journal of Cold War Studies, 4(3):36–55, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9286;   http:/
/muse.jhu.edu/article/9286/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/36/700216/152039702320201067.pdf.

Nuti:2010:DBG

[Nut10] Leopoldo Nuti. Duccio Basosi, Il governo del dollaro. Interdipendenza economica e potere statunitense negli anni di Richard Nixon (1969–1973). Journal of Cold War Studies, 12(1):179–181, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382510;
http://muse.jhu.edu/article/382510/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/1/179/1627423/jcws.2010.12.1.
179.pdf.

Nuti:2011:DBG

[Nut11] Leopoldo Nuti. Duccio Basosi, Il governo del dollaro: Interdipendenza economica e potere statunitense negli anni di Richard Nixon (1969–1973). Firenze: Edizioni Polistampa, 2006. 250 pp. EUR 16.00. Journal of Cold War Studies, 13(1):242–244, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/424256;   http:
//muse.jhu.edu/article/424256/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/1/242/697727/jcws_r_00087.pdf.

Oren:2020:RTP

[OB20] Eitan Oren and Matthew Brummer. Reexamining threat perception in Early Cold War Japan. Journal of Cold War Studies, 22(4):71– 112, Fall 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/4/71/1860594/jcws_a_00948.pdf;   https://muse.jhu.edu/
article/776960.

ODriscoll:2009:ECD

[O’D09] Mervyn O’Driscoll. Explosive challenge: Diplomatic triangles, the United Nations, and the problem of French nuclear testing, 1959–1960. Journal of Cold War Studies, 11(1):28–56, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/259184;
http://muse.jhu.edu/article/259184/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/1/28/1468529/jcws.2009.11.1.
28.pdf.

Odom:2002:VPK

[Odo02] William E. Odom. Voenno-promyshlennyi kompleks SSSR v gody kholodnoi voiny. (Vtoraya polovina 40-kh — nachalo 60-kh godov) (review). Journal of Cold War Studies, 4(4):122–125, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9305;   http://muse.jhu.edu/article/
9305/pdf;   https://direct.mit.edu/jcws/article-pdf/4/4/
122/695702/jcws.2002.4.4.122.pdf.

Odom:2006:CWO

[Odo06] William E. Odom. The Cold War origins of the U.S. central command. Journal of Cold War Studies, 8(2):52–82, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/196290;   http:
//muse.jhu.edu/article/196290/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/2/52/700566/jcws.2006.8.2.52.pdf.

Olesen:2011:TTD

[Ole11] Thorsten Borring Olesen. Tango for Thule: The dilemmas and limits of the “neither confirm nor deny” doctrine in Danish–American relations, 1957–1968. Journal of Cold War Studies, 13(2):116–147, Spring 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/435668;   http:
//muse.jhu.edu/article/435668/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/2/116/700664/jcws_a_00108.pdf.

Olesen:2018:BBH

[Ole18] Mikkel Runge Olesen. To balance or not to balance: How Denmark almost stayed out of NATO, 1948–1949. Journal of Cold War Studies, 20(2):63–98, Spring 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http:

//muse.jhu.edu/article/697061;   https://direct.mit.edu/
jcws/article-pdf/20/2/63/699431/jcws_a_00818.pdf.

Oliver:2011:HKA

[Oli11] Kendrick Oliver. Heonik Kwon, After the Massacre: Commemoration and Consolation in Ha My and My Lai. Berkeley: University of California Press, 2006. 217 pp. $19.95. Journal of Cold War Studies, 13(1):230–231, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/424250;   http://muse.jhu.edu/article/424250/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/1/230/697705/
jcws_r_00081.pdf.

Olsakova:2018:PEE

[Olš18] Doubravka Olšáková. Pugwash in Eastern Europe: The limits of international cooperation under Soviet control in the 1950s and 1960s. Journal of Cold War Studies, 20(1):210–240, Winter 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/691022;   http:
//muse.jhu.edu/article/691022/pdf;   https://direct.mit.
edu/jcws/article-pdf/20/1/210/699340/jcws_a_00805.pdf.

OMalley:2021:SRC

[O’M21] Alanna O’Malley. The Simba Rebellion, the Cold War, and the Stanleyville hostages in the Congo. Journal of Cold War Studies, 23(2): 75–99, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/23/2/75/1922982/jcws_a_00985.pdf;   https://muse.jhu.
edu/article/794727.

Onuf:2014:MDD

[Onu14] Nicholas Onuf. Michael Dillon, Deconstructing International Politics. London: Routledge, 2013. 213 pp. $42.94 paper, $130.00 hardcover. Journal of Cold War Studies, 16(1):210–212, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543731;   http:
//muse.jhu.edu/article/543731/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/1/210/698938/jcws_r_00409.pdf.

OSullivan:2017:CDK

[O’S17] Dónal O’Sullivan. Christian Domnitz, Kooperation und Kontrolle: Die Arbeit der Stasi-Operativgruppen im sozialistischen Ausland. Journal of Cold War Studies, 19(4):260–261, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/683685;   http://muse.jhu.edu/article/683685/
pdf;   https://direct.mit.edu/jcws/article-pdf/19/4/260/
699459/jcws_r_00785.pdf.

Osgood:2002:HMU

[Osg02] Kenneth A. Osgood. Hearts and minds: The unconventional Cold War. Journal of Cold War Studies, 4(2):85–107, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9258;   http:/
/muse.jhu.edu/article/9258/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/2/85/695531/152039702753649656.pdf.

Osgood:2004:WCW

[Osg04] Kenneth A. Osgood. War and Cold War in American Foreign Policy, 1942-62 (review). Journal of Cold War Studies, 6(2):73–75, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/53680;   http://
muse.jhu.edu/article/53680/pdf;   https://direct.mit.edu/
jcws/article-pdf/6/2/73/695969/jcws.2004.6.2.73.pdf.

Owen:2007:RMB

[Owe07] Roger Owen. Robert McNamara, Britain, Nasser and the Balance of Power in the Middle East, 1952–1967. London: Frank Cass, 2003. xvii + 308 pp. Journal of Cold War Studies, 9(4):151–153, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223551;   http://muse.jhu.edu/article/223551/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/151/696755/
jcws.2007.9.4.151.pdf.

Oyen:2014:TRA

[Oye14] Meredith Oyen. “Thunder without Rain”: ARCI, the Far East Refugee Program, and the U.S. response to Hong Kong refugees. Journal of Cold War Studies, 16(4):189–221, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/566657;   http:
//muse.jhu.edu/article/566657/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/4/189/698641/jcws_a_00520.pdf.

PallA:2023:CNC

[PA23] GÃ‘!bor PallÃ^{3} and Matthew Adamson. Constructing nuclear culture under Soviet-style Communism: The Hungarian experience. Journal of Cold War Studies, 25(3):89–111, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/3/
89/2158682/jcws_a_01160.pdf.

Paczkowski:2006:SLP

[Pac06] Andrzej Paczkowski. A Secret Life: The Polish Officer, His Covert Mission and the Price He Paid to Save His Country (review). Journal of Cold War Studies, 8(1):120–122, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/192217;   http:
//muse.jhu.edu/article/192217/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/1/120/696255/152039706775212030.pdf.

Paine:2014:SVE

[Pai14] S. C. M. Paine. State Violence in East Asia edited by N. Ganesan and Sung Chull Kim (review). Journal of Cold War Studies, 16(4):243–244, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566667;   http://muse.jhu.edu/article/566667/pdf;

 https://direct.mit.edu/jcws/article-pdf/16/4/243/698619/
jcws_r_00486.pdf.

Paine:2016:BKM

[Pai16] S. C. M. Paine. Barak Kushner, Men to Devils, Devils to Men: Japanese War Crimes and Chinese Justice. Cambridge, MA: Harvard University Press, 2015. 403 pp. $45.00. Journal of Cold War Studies, 18(2):191–194, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/620554;   http://muse.jhu.edu/article/620554/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/2/191/698880/
jcws_r_00627.pdf.

Paine:2017:MOD

[Pai17] S. C. M. Paine. Meredith Oyen, The Diplomacy of Migration: Transnational Lives and the Making of U.S.-Chinese Relations in the Cold War. Journal of Cold War Studies, 19(4):244–245, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/683677;   http:
//muse.jhu.edu/article/683677/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/4/244/699419/jcws_r_00777.pdf.

Panofsky:2001:ASL

[Pan01] Wolfgang K. H. Panofsky. Atomic Spaces: Living on the Manhattan Project (review). Journal of Cold War Studies, 3(1):130–132, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9180;   http:/
/muse.jhu.edu/article/9180/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/1/130/695406/jcws.2001.3.1.130.pdf.

Papacosma:2023:BRG

[Pap23] S. Victor Papacosma. Book review: The Greek Military Dictatorship: Revisiting a Troubled Past, 1967–1974 by Othon Anastasakis and Katerina Lagos, eds. Journal of Cold War Studies, 25(1):214–217, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/1/
214/2073755/jcws_r_01129.pdf.

Parrott:2009:VMZ

[Par09] Bruce Parrott. Vladislav M. Zubok, A Failed Empire: The Soviet Union in the Cold War from Stalin to Gorbachev. Journal of Cold War Studies, 11(1):170–172, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/259197;   http://muse.jhu.edu/article/259197/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/1/170/
1468553/jcws.2009.11.1.170.pdf.

Paris:2010:KMS

[Par10] Timothy J. Paris. Karl E. Meyer and Shareen Blair Brysac, Kingmakers: The Invention of the Modern Middle East. Journal of Cold War Studies, 12(3):193–195, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392810;   http://muse.jhu.edu/article/392810/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/193/697662/
jcws_r_00038.pdf.

Parrott:2013:DCE

[Par13] Bruce Parrott. David C. Engerman, Know Your Enemy: The Rise and Fall of America’s Soviet Experts. New York: Oxford University Press, 2009. Journal of Cold War Studies, 15(2):159–161, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/509718;   http:
//muse.jhu.edu/article/509718/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/2/159/698201/jcws_r_00359.pdf.

Parrott:2016:PMB

[Par16] R. Joseph Parrott. Philip Muehlenbeck, Betting on the Africans: John F. Kennedy’s Courting of African Nationalist Leaders. New York: Oxford University Press, 2012. 360 pp. $27.95. Journal of Cold War Studies, 18(4):221–223, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645919;   http://muse.jhu.edu/article/645919/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/221/700085/
jcws_r_00691.pdf.

Parr:2019:VVV

[Par19] Helen Parr. Visions, Votes and Vetoes: The Empty Chair Crisis and the Luxembourg Compromise Thirty Years On ed. by Jean-Marie Palayret, Helen Wallace, and Pascaline Winand (review). Journal of Cold War Studies, 21(2):203–205, Spring 2019. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/726143;   https://direct.mit.
edu/jcws/article-pdf/21/2/203/699585/jcws_r_00888.pdf.

Parrott:2022:LCW

[Par22] Bruce Parrott. Lessons of the Cold War: Getting it wrong. Journal of Cold War Studies, 24(1):219–249, Winter 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/24/1/219/1982915/jcws_a_
01060.pdf;   https://muse.jhu.edu/article/843082.

Pastor:2003:OOB

[Pas03] Robert A. Pastor. Our Own Backyard: The United States and Central America, 1977–1992 (review). Journal of Cold War Studies, 5(3): 126–128, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/45208;
 http://muse.jhu.edu/article/45208/pdf;   https://direct.
mit.edu/jcws/article-pdf/5/3/126/695858/jcws.2003.5.3.126.
pdf.

Pastor:2010:MEO

[Pas10] Robert A. Pastor. Missing each other: a comment on “when empathy failed”. Journal of Cold War Studies, 12(2):92–94, Spring 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/382203;   http:
//muse.jhu.edu/article/382203/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/2/92/697511/jcws.2010.12.2.92.pdf.

Pavalko:2019:LPM

[Pav19] Nathan Pavalko. Last of the President’s Men by Bob Woodward (review). Journal of Cold War Studies, 21(1):193–195, Winter 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/723539;   https://direct.mit.
edu/jcws/article-pdf/21/1/193/699785/jcws_r_00860.pdf.

Payne:2021:AAF

[Pay21] Stanley G. Payne. Arguing Americanism: Franco Lobbyists, Roosevelt’s Foreign Policy, and the Spanish Civil War by Michael E. Chapman (review). Journal of Cold War Studies, 23(4): 255–257, Fall 2021. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL https://direct.mit.edu/jcws/
article-pdf/23/4/255/1970087/jcws_r_01050.pdf;   https://
muse.jhu.edu/article/835887.

Pechatnov:2008:GRS

[Pec08] Vladimir Pechatnov. Geoffrey Roberts, Stalin’s Wars: From World War to Cold War, 1939–1953. Journal of Cold War Studies, 10(3):179–181, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241521;   http://muse.jhu.edu/article/241521/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/179/697119/
jcws.2008.10.3.179.pdf.

Pechatnov:2013:DOD

[Pec13a] Vladimir Pechatnov. Donal O’Sullivan, Dealing with the Devil: Anglo-Soviet Intelligence Cooperation during the Second World War. New York: Peter Lang Publishing, 2010. xi + 338 pp. Journal of Cold War Studies, 15(3):191–193, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522719;   http://muse.jhu.edu/article/522719/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/191/698325/
jcws_r_00380.pdf.

Pechatnov:2013:GAT

[Pec13b] Vladimir O. Pechatnov. Gaddis’s achievement: Taking the measure of Kennan. Journal of Cold War Studies, 15(4):183–188, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/534421;   http:
//muse.jhu.edu/article/534421/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/4/183/698335/jcws_a_00402.pdf.

Pedersen:2007:RBC

[Ped07] Vernon L. Pedersen. R. Bruce Craig, Treasonable Doubt: The Harry Dexter White Spy Case. Journal of Cold War Studies, 9(1):111–113, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210097;   http://muse.jhu.edu/article/210097/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/111/696509/
jcws.2007.9.1.111.pdf.

Pedemonte:2019:CNS

[Ped19] Rafael Pedemonte. A case of “new Soviet internationalism”: Relations between the USSR and Chile’s Christian Democratic government, 1964–1970. Journal of Cold War Studies, 21(3):4–25, Summer 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/731281;   https://direct.
mit.edu/jcws/article-pdf/21/3/4/699620/jcws_a_00894.pdf.

Pee:2023:BRF

[Pee23] Robert Pee. Book review: Freedom on the Offensive: Human Rights, Democracy Promotion, and US Interventionism in the Late Cold War by William Michael Schmidli. Journal of Cold War Studies, 25(2): 250–252, Spring 2023. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/25/2/250/2140944/jcws_r_01148.pdf.

Pembleton:2016:IGS

[Pem16] Matthew R. Pembleton. Imagining a global sovereignty: U.S. counternarcotic operations in Istanbul during the early Cold War and the origins of the foreign “War on Drugs”. Journal of Cold War Studies, 18 (2):28–63, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/620546;
http://muse.jhu.edu/article/620546/pdf;   https://direct.
mit.edu/jcws/article-pdf/18/2/28/698842/jcws_a_00637.pdf.

Peng:2006:SGZ

[Pen06] Tao Peng. Shu Guang Zhang, Economic Cold War: America’s Embargo against China and the Sino-Soviet Alliance, 1949–1963. Journal of Cold War Studies, 8(2):144–146, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/196264;   http://muse.jhu.edu/article/196264/
pdf;   https://direct.mit.edu/jcws/article-pdf/8/2/144/700516/
jcws.2006.8.2.144.pdf.

Pendas:2018:GBE

[Pen18] Devin Pendas. Gerhard Besier, ed., 20 Jahre neue Bundesrepublik: Kontinuitäten und Diskontinuitäten. Journal of Cold War Studies, 20 (4):272–274, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/718226;
https://direct.mit.edu/jcws/article-pdf/20/4/272/1969644/
jcws_r_00852.pdf.

Pereira:2004:RAE

[Per04] Marina V. Pereira. Russian–American Economic Relations, 1763– 1999 (review). Journal of Cold War Studies, 6(3):179–181, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/171172;   http:
//muse.jhu.edu/article/171172/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/3/179/696139/jcws.2004.6.3.179.pdf.

Perovic:2007:TSS

[Per07] Jeronim Perović. The Tito–Stalin split: a reassessment in light of new evidence. Journal of Cold War Studies, 9(2):32–63, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/214754;   http:
//muse.jhu.edu/article/214754/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/2/32/696477/jcws.2007.9.2.32.pdf.

Petrescu:2008:RIR

[Pet08] Dragoş Petrescu. Radu Ioanid, The Ransom of the Jews: The Story of the Extraordinary Secret Bargain Between Romania and Israel. Journal of Cold War Studies, 10(1):127–128, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/231590;   http:
//muse.jhu.edu/article/231590/pdf;   https://direct.mit.
edu/jcws/article-pdf/10/1/127/700055/jcws.2008.10.1.127.
pdf.

Peters:2011:SPR

[Pet11a] Rita P. Peters. Silvio Pons and Robert Service, eds., A Dictionary of 20th-Century Communism. Princeton, NJ: Princeton University Press, 2010. 944 pp. $99.50. Journal of Cold War Studies, 13(4):226–229, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457205;   http://muse.jhu.edu/article/457205/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/226/697885/
jcws_r_00185.pdf.

Petersen:2011:STG

[Pet11b] Nikolaj Petersen. SAC at Thule: Greenland in the U.S. polar strategy. Journal of Cold War Studies, 13(2):90–115, Spring 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/435667;   http:
//muse.jhu.edu/article/435667/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/2/90/700672/jcws_a_00138.pdf.

Peteri:2013:DKE

[Pét13] György Péteri. Daniela Koleva, ed., Negotiating Normality: Everyday Lives in Socialist Institutions. New Brunswick, NJ: Transaction Publishers, 2012. 236 pp. Journal of Cold War Studies, 15(2):152–153, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/509715;   http://muse.jhu.edu/article/509715/pdf;

 https://direct.mit.edu/jcws/article-pdf/15/2/152/698181/
jcws_r_00356.pdf.

Petrovic:2018:SRA

[Pet18] Vladimir Petrović. Svetozar Rajak et al., eds., The Balkans in the Cold War. Journal of Cold War Studies, 20(4):258–262, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718220;   https://direct.mit.
edu/jcws/article-pdf/20/4/258/1969624/jcws_r_00846.pdf.

Peterson:2020:CWB

[Pet20] Christian Philip Peterson. Changing the world from “Below”: U.S. peace activists and the transnational struggle for peace and détente in the 1980s. Journal of Cold War Studies, 22(3):180–224, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/22/3/180/1860585/
jcws_a_00926.pdf;   https://muse.jhu.edu/article/762895.

Pfaltzgraff:2002:NCW

[Pfa02] Robert L. Pfaltzgraff, Jr. Norstad: Cold War NATO Supreme Commander: Airman, Strategist, Diplomat (review). Journal of Cold War Studies, 4(3):142–144, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9288;   http://muse.jhu.edu/article/9288/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/3/142/700256/
jcws.2002.4.3.142.pdf.

Pietrobon:2016:RNC

[Pie16] Allen Pietrobon. The role of Norman cousins and track II diplomacy in the breakthrough to the 1963 limited test ban treaty. Journal of Cold War Studies, 18(1):60–79, Winter 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/612006;   http://muse.jhu.edu/article/612006/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/1/60/698749/
jcws_a_00619.pdf.

Piffer:2015:OSS

[Pif15] Tommaso Piffer. Office of strategic services versus special operations executive: Competition for the Italian resistance, 1943–1945. Journal of Cold War Studies, 17(4):41–58, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606020;   http://muse.jhu.edu/article/606020/pdf;

 https://direct.mit.edu/jcws/article-pdf/17/4/41/698757/
jcws_a_00596.pdf.

Pike:2002:BEG

[Pik02] David Pike. Building the East German Myth: Historical Mythology and Youth Propaganda in the German Democratic Republic, 1945–1989 (review). Journal of Cold War Studies, 4(1):118–119, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9235;   http:/
/muse.jhu.edu/article/9235/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/1/118/695540/jcws.2002.4.1.118.pdf.

Pipes:2003:C

[Pip03] Richard Pipes. Correspondence. Journal of Cold War Studies, 5(3): 161, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/45209;
 http://muse.jhu.edu/article/45209/pdf;   https://direct.
mit.edu/jcws/article-pdf/5/3/161/695937/jcws.2003.5.3.161a.
pdf.

Pipes:2011:PSL

[Pip11] Richard Pipes. Polish sovietology in the lead-up to the Cold War. Journal of Cold War Studies, 13(2):175–193, Spring 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/435670;   http:
//muse.jhu.edu/article/435670/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/2/175/700668/jcws_a_00103.pdf.

Pipes:2015:EPE

[Pip15] Richard Pipes. Exiled in Palestine: The Emigration of Zionist Convicts from the Soviet Union, 1924–1934 by Ziva Galili and Boris Morozov (review). Journal of Cold War Studies, 17(3):247–248, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595088;   http:
//muse.jhu.edu/article/595088/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/247/1102488/jcws_r_00577.pdf.

Pitt:2023:BRP

[Pit23] Jorden Pitt. Book review: Pulp Vietnam: War and Gender in Cold War Men’s Adventure Magazines by Gregory A. Daddis. Journal of Cold War Studies, 25(1):212–214, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/1/
212/2073753/jcws_r_01128.pdf.

Pee:2022:RDP

[PL22] Robert Pee and Scott Lucas. Reevaluating democracy promotion: The Reagan Administration, allied authoritarian states, and regime change. Journal of Cold War Studies, 24(3):151–199, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/3/
151/2042013/jcws_a_01090.pdf.

Pocock:2003:MNA

[Poc03] Chris Pocock. By Any Means Necessary: America’s Secret Air War in the Cold War (review). Journal of Cold War Studies, 5(3):117–118, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/45210;   http://
muse.jhu.edu/article/45210/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/3/117/695818/jcws.2003.5.3.117.pdf.

Podvig:2014:SMS

[Pod14] Pavel Podvig. Sergo Mikoyan, The Soviet Cuban Missile Crisis: Castro, Mikoyan, Kennedy, Khrushchev, and the Missiles of November, ed. by Svetlana Savranskaya. Washington, DC: Woodrow Wilson Center Press, 2012. 589 pp. $24.95. Journal of Cold War Studies, 16(1):239–241, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/543745;   http://muse.jhu.edu/article/543745/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/1/239/698962/
jcws_r_00423.pdf.

Polonsky:2005:CMP

[Pol05] Antony Polonsky. Contested Memories: Poles and Jews during the Holocaust and Its Aftermath (review). Journal of Cold War Studies, 7(3):176–177, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185116;   http://muse.jhu.edu/article/185116/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/176/1102595/
jcws.2005.7.3.176.pdf.

Polmar:2006:GWW

[Pol06] Norman Polmar. Gary E. Weir and Walter J. Boyne, Rising Tide: The Untold Story of the Russian Submarines That Fought the Cold War. Journal of Cold War Studies, 8(2):134–135, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/196292;   http:
//muse.jhu.edu/article/196292/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/2/134/700524/jcws.2006.8.2.134.pdf.

Pollack:2007:EFC

[Pol07a] Jonathan D. Pollack. Evan A. Feigenbaum, China’s Techno-Warriors: National Security and Strategic Competition from the Nuclear to the Information Age. Journal of Cold War Studies, 9(1):139–141, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210098;   http://muse.jhu.edu/article/210098/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/139/696548/
jcws.2007.9.1.139.pdf.

Polmar:2007:MJB

[Pol07b] Norman Polmar. Martin J. Bollinger, Stalin’s Slave Ships: Kolyma, the Gulag Fleet, and the Role of the West. Journal of Cold War Studies, 9(3):180–182, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219466;   http://muse.jhu.edu/article/219466/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/180/696678/
jcws.2007.9.3.180.pdf.

Polonsky:2010:PRP

[Pol10] Antony Polonsky. Phillip T. Rutherford, Prelude to the Final Solution. The Nazi Program for Deporting Ethnic Poles, 1939–1941. Journal of Cold War Studies, 12(1):185–189, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/382511;   http://muse.jhu.edu/article/382511/
pdf;   https://direct.mit.edu/jcws/article-pdf/12/1/185/
1627443/jcws.2010.12.1.185.pdf.

Polonsky:2013:BZS

[Pol13] Antony Polonsky. Bożena Szaynok, Poland-Israel 1944–1968: In the Shadow of the Past and of the Soviet Union. Warsaw: Institute of National Remembrance, 2012. 501 pp. Journal of Cold War Studies, 15(3):212–214, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522730;   http://muse.jhu.edu/article/522730/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/212/698361/
jcws_r_00385.pdf.

Pons:2001:DSL

[Pon01a] Silvio Pons. Dimitrov and Stalin 1934–1943. Letters from the Soviet Archives (review). Journal of Cold War Studies, 3(3):116– 117, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9210;   http:/
/muse.jhu.edu/article/9210/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/116/695468/jcws.2001.3.3.116.pdf.

Pons:2001:STO

[Pon01b] Silvio Pons. Stalin, Togliatti, and the origins of the Cold War in Europe. Journal of Cold War Studies, 3(2):3–27, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9193;   http:/
/muse.jhu.edu/article/9193/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/2/3/695376/152039701300373862.pdf.

Pons:2013:MDF

[Pon13] Silvio Pons. Michael David-Fox, Showcasing the Great Experiment: Cultural Diplomacy and Western Visitors to the Soviet Union 1921– 1941. New York: Oxford University Press, 2012. xii + 396 pp. Journal of Cold War Studies, 15(2):156–159, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/509717;   http://muse.jhu.edu/article/509717/
pdf;   https://direct.mit.edu/jcws/article-pdf/15/2/156/
698185/jcws_r_00358.pdf.

Port:2008:MLD

[Por08] Andrew I. Port. Mark Landsman, Dictatorship and Demand: The Politics of Consumerism in East Germany. Journal of Cold War Studies, 10(1):161–163, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231591;   http://muse.jhu.edu/article/231591/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/161/700091/
jcws.2008.10.1.161.pdf.

Port:2009:GWS

[Por09] Andrew I. Port. Gerhard Wettig, Stalin and the Cold War in Europe: The Emergence and Development of East-West Conflict, 1939–1953. Journal of Cold War Studies, 11(1):161–164, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/259193;

http://muse.jhu.edu/article/259193/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/1/161/1468516/jcws.2009.11.1.
161.pdf.

Posner:2007:MHK

[Pos07] Gerald L. Posner. Max Holland, The Kennedy Assassination Tapes: The White House Conversations of Lyndon B. Johnson Regarding the Assassination, the Warren Commission, and the Aftermath. New York: Alfred A. Knopf, 2004. 453 pp. $26.95. Journal of Cold War Studies, 9(2):154–156, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214755;   http://muse.jhu.edu/article/214755/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/154/696435/
jcws.2007.9.2.154.pdf.

Pospisil:2019:ISA

[Pos19] Filip Pospíšil. Inspiration, subversion, and appropriation: The effects of Radio Free Europe music broadcasting. Journal of Cold War Studies, 21(4):124–149, Fall 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/21/4/124/699689/jcws_a_00908.pdf;
https://muse.jhu.edu/article/736846.

Powell:2001:SHS

[Pow01] David Powell. Storming the Heavens: The Soviet League of the Militant Godless (review). Journal of Cold War Studies, 3(2):113–116, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9194;   http:/
/muse.jhu.edu/article/9194/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/2/113/695392/jcws.2001.3.2.113.pdf.

Powell:2016:CWF

[Pow16] Nathaniel K. Powell. The “Cuba of the West”?: France’s Cold War in Zaïre, 1977–1978. Journal of Cold War Studies, 18(2):64– 96, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/620547;
http://muse.jhu.edu/article/620547/pdf;   https://direct.
mit.edu/jcws/article-pdf/18/2/64/698840/jcws_a_00638.pdf.

Pulcini:2021:OPP

[PR21] Giordana Pulcini and Or Rabinowitz. An ounce of prevention — a pound of cure?: The Reagan Administration’s nonproliferation policy and the Osirak Raid. Journal of Cold War Studies, 23(2):

4–40, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/23/2/4/1922976/jcws_a_01007.pdf;   https://muse.jhu.
edu/article/794725.

Pravda:1999:SMC

[Pra99] Alex Pravda. Sarah E. Mendelson, Changing Course: Ideas, Politics, and the Soviet Withdrawal from Afghanistan. Princeton, NJ: Princeton University Press, 1998. xiii + 140 pp. Journal of Cold War Studies, 1(3):182–184, Fall 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9118;   https://direct.mit.edu/jcws/
article-pdf/1/3/182/695178/jcws.1999.1.3.182.pdf.

Prados:2007:FGP

[Pra07] John Prados. Francis Gary Powers with Curt Gentry, Operation Overflight: a Memoir of the U-2 Incident. Journal of Cold War Studies, 9(3):175–176, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219467;   http://muse.jhu.edu/article/219467/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/175/696646/
jcws.2007.9.3.175.pdf.

Prados:2008:CPJ

[Pra08a] John Prados. Christopher A. Preble, John F. Kennedy and the Missile Gap. Journal of Cold War Studies, 10(1):141–143, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/231592;
http://muse.jhu.edu/article/231592/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/1/141/700047/jcws.2008.10.1.
141.pdf.

Prados:2008:HLG

[Pra08b] John Prados. Harold Lloyd Goodall, Jr., A Need to Know: The Clandestine History of a CIA Family. Journal of Cold War Studies, 10(2):152–153, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/237163;   http://muse.jhu.edu/article/237163/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/2/152/697001/
jcws.2008.10.2.152.pdf.

Prados:2014:BJF

[Pra14] John Prados. The Brothers: John Foster Dulles, Allen Dulles, and Their Secret Cold War by Stephen Kinzer (review). Journal of Cold War Studies, 16(4):255–256, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566673;   http://muse.jhu.edu/article/566673/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/255/698651/
jcws_r_00492.pdf.

Prados:2015:ECWa

[Pra15a] John Prados. Eisenhower and the Cold War Arms Race: “Open Skies” and the Military–Industrial Complex by Helen Bury (review). Journal of Cold War Studies, 17(2):142–143, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/582438;   http://muse.jhu.edu/article/582438/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/2/142/
698692/jcws_r_00536.pdf.

Prados:2015:ECWb

[Pra15b] John Prados. Eisenhower and the Cold War Arms Race: “Open Skies” and the Military–Industrial Complex by Helen Bury (review). Journal of Cold War Studies, 17(3):232–233, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595080;   http:
//muse.jhu.edu/article/595080/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/232/1102470/jcws_r_00569.pdf.

Prados:2017:RMD

[Pra17] John Prados. Robert M. Dienesch, Eyeing the Red Storm: Eisenhower and the First Attempt to Build a Spy Satellite. Lincoln: University of Nebraska Press, 2016. 296 pp. $34.95. Journal of Cold War Studies, 19(3):260–261, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668180;   http://muse.jhu.edu/article/668180/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/260/699288/
jcws_r_00747.pdf.

Pravda:2019:GHL

[Pra19] Alex Pravda. Gorbachev: His Life and Times by William Taubman (review). Journal of Cold War Studies, 21(1):217–221, Winter 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/723548;   https://direct.mit.
edu/jcws/article-pdf/21/1/217/699796/jcws_r_00869.pdf.

Prados:2020:SWI

[Pra20] John Prados. Silent Warriors, Incredible Courage: The Declassified Stories of Cold War Reconnaissance Flights and the Men Who Flew Them by Wolfgang W. E. Samuel (review). Journal of Cold War Studies, 22(1):245–246, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/22/1/245/699874/jcws_r_
00933.pdf;   https://muse.jhu.edu/article/749648.

Prados:2021:KRM

[Pra21] John Prados. Kevin Ruane and Matthew Jones, Anthony Eden, Anglo–American Relations and the 1954 Indochina Crisis. Journal of Cold War Studies, 23(2):176–177, Spring 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/23/2/176/1919929/jcws_r_
01022.pdf;   https://muse.jhu.edu/article/794717.

Preston:2011:DAJ

[Pre11a] Andrew Preston. David L. Anderson and John Ernst, eds., The War That Never Ends: New Perspectives on the Vietnam War. Lexington, KY: University Press of Kentucky, 2007. Journal of Cold War Studies, 13(1):228–230, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/424249;   http://muse.jhu.edu/article/424249/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/1/228/697715/
jcws_r_00079.pdf.

Preston:2011:JMC

[Pre11b] Andrew Preston. James M. Carter, Inventing Vietnam: The United States and State Building, 1954–1968. New York: Cambridge University Press, 2008. 278 pp. Journal of Cold War Studies, 13(3):212–214, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447472;   http://muse.jhu.edu/article/447472/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/212/700730/
jcws_r_00152.pdf.

Preston:2017:GDW

[Pre17] Andrew Preston. Gregory A. Daddis, Westmoreland’s War: Reassessing American Strategy in Vietnam. New York: Oxford University Press, 2014. xxv, 250 pp. $36.95. Journal of Cold War Studies, 19(3):256–258, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668178;   http://muse.jhu.edu/article/668178/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/256/699296/
jcws_r_00745.pdf.

Pruessen:2004:NEA

[Pru04] Ronald W. Pruessen. No Exit: America and the German Problem, 1943–1954 (review). Journal of Cold War Studies, 6 (3):150–151, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/171173;   http://muse.jhu.edu/article/171173/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/3/150/696041/
jcws.2004.6.3.150.pdf.

Patti:2020:WNA

[PS20] Carlo Patti and Matias Spektor. “We are not a nonproliferation agency”: Henry Kissinger’s failed attempt to accommodate nuclear Brazil, 1974–1977. Journal of Cold War Studies, 22(2):58–93, Spring 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/2/58/1860550/jcws_a_00940.pdf;   https://muse.jhu.edu/
article/756825.

Pucci:2021:ASP

[Puc21] Molly Pucci. The anatomy of Stalinist police states: Blanket surveillance networks in Communist Czechoslovakia and Poland, 1949–1952. Journal of Cold War Studies, 23(4):95–123, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/23/4/95/1970083/jcws_a_
01040.pdf;   https://muse.jhu.edu/article/835878.

Puddington:2012:QOH

[Pud12] Arch Puddington. Quenby Olsted Hughes, “In the Interest of Democracy”: The Rise and Fall of the Early Cold War Alliance between the American Federation of Labor and the Central Intelligence Agency. Bern: Peter Lang, 2011. 204 pp. $55.95. Journal of Cold War Studies, 14(3):228–230, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/490196;   http://muse.jhu.edu/article/490196/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/3/228/698093/
jcws_r_00269.pdf.

Painter:2005:SGM

[PW05] David S. Painter and Edmund A. Walsh. Staging Growth: Modernization, Development, and the Global Cold War (review). Journal of Cold War Studies, 7(4):137–138, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188275;   http://muse.jhu.edu/article/188275/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/137/696269/
1520397055012424.pdf.

Qureshi:2012:EDU

[QG12] Lubna Qureshi and Kristian Gustafson. Exchange: Debating U.S. involvement in Chile in the 1970s. Journal of Cold War Studies, 14(1):114–117, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467686;   http://muse.jhu.edu/article/467686/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/114/697888/
jcws_c_00199.pdf.

Quester:2016:CWE

[QMLL16] George H. Quester, Patrick M. Morgan, Jeffrey S. Lantis, and Elli Lieberman. Cold War-era deterrence and international relations in the Middle East. Journal of Cold War Studies, 18(4):180–207, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/645912;   http:
//muse.jhu.edu/article/645912/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/4/180/700093/jcws_c_00684.pdf.

Quandt:2004:SDW

[Qua04] William B. Quandt. Six Days of War: June 1967 and the Making of the Modern Middle East (review). Journal of Cold War Studies, 6(3):145–148, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/171174;   http://muse.jhu.edu/article/171174/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/3/145/696047/
jcws.2004.6.3.145.pdf.

Quandt:2009:NJA

[Qua09] William B. Quandt. Nigel J. Ashton, ed., The Cold War in the Middle East: Regional Conflict and the Superpowers, 1967–73. Journal of Cold War Studies, 11(1):159–161, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/259192;   http://muse.jhu.edu/article/259192/

pdf;   https://direct.mit.edu/jcws/article-pdf/11/1/159/
1468498/jcws.2009.11.1.159.pdf.

Quandt:2013:WRL

[Qua13] William B. Quandt. William Roger Louis and Avi Shlaim, eds., The 1967 Arab-Israeli War: Origins and Consequences. New York: Cambridge University Press, 2012. 325 pp. Journal of Cold War Studies, 15(2):165–167, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/509721;   http://muse.jhu.edu/article/509721/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/2/165/698205/
jcws_r_00362.pdf.

Quandt:2014:OWP

[Qua14] William B. Quandt. The October 1973 War: Politics, Diplomacy, Legacy edited by Asaf Siniver (review). Journal of Cold War Studies, 16(4):256–258, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566674;   http://muse.jhu.edu/article/566674/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/256/698637/
jcws_r_00493.pdf.

Quandt:2021:ISA

[Qua21] William B. Quandt. Imperfect Strangers: Americans, Arabs, and U.S.–Middle East Relations in the 1970s by Salim Yaqub, Ithaca, NY: Cornell University Press, 2016. 455 pp. (review). Journal of Cold War Studies, 23(1):243–246, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/1/243/1911790/jcws_r_00993.pdf;
https://muse.jhu.edu/article/788397.

Quandt:2023:BRL

[Qua23] William B. Quandt. Book review: A Lost Peace: Great Power Politics and the Arab–Israeli Dispute, 1967–1979 by Galen Jackson. Journal of Cold War Studies, 25(3):220–224, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/25/3/220/2158670/jcws_r_
01165.pdf.

Quester:2002:INB

[Que02] George H. Quester. India’s Nuclear Bomb: The Impact on Nuclear Proliferation (review). Journal of Cold War Studies, 4(2):127–128, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9259;   http:/
/muse.jhu.edu/article/9259/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/2/127/695581/jcws.2002.4.2.127.pdf.

Rabe:2013:CPC

[Rab13] Stephen G. Rabe. Colin A. Palmer, Cheddi Jagan and the Politics of Power: British Guiana’s Struggle for Independence. Chapel Hill: University of North Carolina Press, 2010. 363 pp. $39.95. Journal of Cold War Studies, 15(3):200–202, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522724;   http://muse.jhu.edu/article/522724/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/200/698309/
jcws_r_00351.pdf.

Radchenko:2006:MPS

[Rad06] Sergey S. Radchenko. Mongolian politics in the shadow of the Cold War: The 1964 coup attempt and the Sino–Soviet split. Journal of Cold War Studies, 8(1):95–119, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/192218;   http://muse.jhu.edu/article/192218/
pdf;   https://direct.mit.edu/jcws/article-pdf/8/1/95/696271/
152039706775212021.pdf.

Radchenko:2007:SSR

[Rad07] Sergey Radchenko. Sino–Soviet relations and the emergence of the Chinese Communist regime, 1946–1950: New documents, old story. Journal of Cold War Studies, 9(4):115–124, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/223552;   http:
//muse.jhu.edu/article/223552/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/4/115/696721/jcws.2007.9.4.115.pdf.

Radchenko:2011:ABR

[Rad11] Sergey Radchenko. Archie Brown, The Rise and Fall of Communism. New York: HarperCollins, 2009. 736 pp. $35.99. Journal of Cold War Studies, 13(1):256–258, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/424263;   http://muse.jhu.edu/article/424263/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/1/256/697773/
jcws_r_00094.pdf.

Radchenko:2012:PSB

[Rad12a] Sergey Radchenko. Pavel Stroilov, Behind the Desert Storm: a Secret Archive Stolen from the Kremlin That Sheds New Light on the Arab Revolutions in the Middle East. Chicago: Price World Publishing, 2011. Journal of Cold War Studies, 14(4):266–268, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/500354;   http:
//muse.jhu.edu/article/500354/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/4/266/698195/jcws_r_00301.pdf.

Radchenko:2012:RC

[Rad12b] Sergey Radchenko. Reply to the commentaries. Journal of Cold War Studies, 14(1):107–110, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467684;   http://muse.jhu.edu/article/467684/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/107/697833/
jcws_a_00197.pdf.

Radchenko:2014:SRA

[Rad14] Sergey Radchenko. Socialist Revolutions in Asia: The Social History of Mongolia in the Twentieth Century by Irina Morozova (review). Journal of Cold War Studies, 16(3):224–225, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/556087;   http:
//muse.jhu.edu/article/556087/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/3/224/698529/jcws_r_00463.pdf.

Radchenko:2015:MRS

[Rad15] Sergey Radchenko. Mao: The Real Story by Alexander V. Pantsov (review). Journal of Cold War Studies, 17(2):158–160, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/582446;   http:
//muse.jhu.edu/article/582446/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/2/158/698709/jcws_r_00544.pdf.

Radchenko:2016:ZXD

[Rad16] Sergey Radchenko. Zhang Xiaoming, Deng Xiaoping’s Long War: The Military Conflict between China and Vietnam, 1979–1991. Chapel Hill: University of North Carolina Press, 2015. 296 pp. $34.95. Journal of Cold War Studies, 18(3):198–200, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/632288;   http:

//muse.jhu.edu/article/632288/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/3/198/699061/jcws_r_00662.pdf.

Radchenko:2017:LCP

[Rad17] Sergey Radchenko. Lost chance for peace: The 1945 CCP– Kuomintang Peace Talks revisited. Journal of Cold War Studies, 19 (2):84–114, Spring 2017. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://www.mitpressjournals.org/doi/
full/10.1162/JCWS_a_00742;   https://direct.mit.edu/jcws/
article-pdf/19/2/84/699197/jcws_a_00742.pdf.

Raim:2011:PGA

[Rai11] Edith Raim. Philipp Gassert and Alan E. Steinweis, eds., Coping with the Nazi Past: West German Debates on Nazism and Generational Conflict, 1955–1975. New York: Berghahn Books, 2006. viii 1 339 pp. $34.95. Journal of Cold War Studies, 13(4):255–257, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457218;   http:
//muse.jhu.edu/article/457218/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/4/255/697802/jcws_r_00130.pdf.

Rajak:2014:NBC

[Raj14] Svetozar Rajak. No bargaining chips, no spheres of interest: The Yugoslav origins of Cold War non-alignment. Journal of Cold War Studies, 16(1):146–179, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/543728;   http://muse.jhu.edu/article/543728/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/1/146/698896/
jcws_a_00434.pdf.

Rakove:2015:FLJ

[Rak15] Robert B. Rakove. Frank Leith Jones, Blowtorch: Robert Komer, Vietnam, and American Cold War Strategy. Annapolis: Naval Institute Press, 2013. 416 pp. $52.95. Journal of Cold War Studies, 17(4):188–191, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606032;   http://muse.jhu.edu/article/606032/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/188/698767/
jcws_r_00588.pdf.

Ramet:2001:JRR

[Ram01] Sabrina P. Ramet. Jazz, Rock, and Rebels: Cold War Politics and American Culture in a Divided Germany (review). Journal of Cold War Studies, 3(1):137–138, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9181;   http://muse.jhu.edu/article/9181/pdf;
 https://direct.mit.edu/jcws/article-pdf/3/1/137/695438/
jcws.2001.3.1.137.pdf.

Ramcharan:2007:SGR

[Ram07] Lily Ramcharan. Stephen G. Rabe, U.S. Intervention in British Guiana: a Cold War Story. Chapel Hill: The University of North Carolina Press, 2005. 240 pp. Journal of Cold War Studies, 9(4):153–155, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223553;   http://muse.jhu.edu/article/223553/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/153/696765/
jcws.2007.9.4.153.pdf.

Ramet:2008:MCM

[Ram08] Sabrina P. Ramet. Mark Carroll, Music and Ideology in Cold War Europe. Journal of Cold War Studies, 10(3):178–179, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/241520;
http://muse.jhu.edu/article/241520/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/3/178/697151/jcws.2008.10.3.
178.pdf.

Ramet:2011:MFS

[Ram11a] Sabrina P. Ramet. Mark Fenemore, Sex, Thugs and Rock ‘n’ Roll: Teenage Rebels in Cold-War East Germany. New York and Oxford: Berghahn Books, 2007. 277 pp. $90.00/£45.00. Journal of Cold War Studies, 13(4):279–280, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457227;   http://muse.jhu.edu/article/457227/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/279/697835/
jcws_r_00159.pdf.

Ramsden:2011:BLC

[Ram11b] John Ramsden. Brian Lavery, Churchill Goes to War: Winston’s Wartime Journeys. Annapolis, MD: Naval Institute Press, 2007. 392 pp. $34.95. Journal of Cold War Studies, 13(3):236–238, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/447485;   http:
//muse.jhu.edu/article/447485/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/3/236/700722/jcws_r_00118.pdf.

Randall:2002:SWL

[Ran02] Amy E. Randall. Stalinism as a Way of Life: a Narrative in Documents (review). Journal of Cold War Studies, 4(2):129–131, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9260;   http:/
/muse.jhu.edu/article/9260/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/2/129/695599/jcws.2002.4.2.129.pdf.

Randall:2006:DHE

[Ran06] Amy E. Randall. David L. Hoffman, ed., Stalinism: The Essential Readings. Journal of Cold War Studies, 8(4):172–175, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/204630;   http:
//muse.jhu.edu/article/204630/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/4/172/700815/jcws.2006.8.4.172.pdf.

Randall:2010:FLB

[Ran10] Amy E. Randall. Frances Lee Bernstein, The Dictatorship of Sex: Lifestyle Advice for the Soviet Masses. Journal of Cold War Studies, 12(4):203–205, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/403812;   http://muse.jhu.edu/article/403812/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/4/203/697560/
jcws_r_00064.pdf.

Randall:2012:MBS

[Ran12] Amy E. Randall. Mark B. Smith, Property of Communists: The Urban Housing Program from Stalin to Khrushchev. DeKalb, IL: University of Northern Illinois Press, 2010. xii + 240 pp. $40.00. Journal of Cold War Studies, 14(4):260–263, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500352;   http://muse.jhu.edu/article/500352/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/260/698171/
jcws_r_00299.pdf.

Rath:2021:MCW

[Rat21] Thomas Rath. Mexico’s Cold War: Cuba, the United States, and the Legacy of the Mexican Revolution by Renata Keller, New York: Cambridge University Press, 2015 (review). Journal of Cold War Studies, 23(1):250–252, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/1/250/1911808/jcws_r_00998.pdf;
https://muse.jhu.edu/article/788392.

Raymond:2020:SCT

[Ray20] Gregory V. Raymond. Strategic culture and Thailand’s response to Vietnam’s occupation of Cambodia, 1979–1989: A Cold War epilogue. Journal of Cold War Studies, 22(1):4–45, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/22/1/
4/699808/jcws_a_00924.pdf;   https://muse.jhu.edu/article/
749658.

Reardon:2006:SPB

[Rea06] Lawrence C. Reardon. Steven E. Phillips, Between Assimilation and Independence: The Taiwanese Encounter Nationalist China, 1945–1950. Journal of Cold War Studies, 8(4):153–155, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/204631;   http:
//muse.jhu.edu/article/204631/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/4/153/700803/jcws.2006.8.4.153.pdf.

Readman:2011:CCI

[Rea11] Kristina Spohr Readman. Conflict and cooperation in intra-alliance nuclear politics: Western Europe, the United States, and the genesis of NATO’s dual-track decision, 1977–1979. Journal of Cold War Studies, 13(2):39–89, Spring 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/435666;   http://muse.jhu.edu/article/435666/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/2/39/700670/
jcws_a_00137.pdf.

Reckard:2017:JMA

[Rec17] Bryan R. Reckard. Joyce Mao, Asia First: China and the Making of Modern American Conservatism. Journal of Cold War Studies, 19(4):245–247, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683678;   http://muse.jhu.edu/article/683678/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/245/699436/
jcws_r_00778.pdf.

Rees:2016:MDK

[Ree16] E. A. Rees. Miriam Dobson, Khrushchev’s Cold Summer: Gulag Returnees, Crime and the Fate of Reform after Stalin. Ithaca, NY: Cornell University Press, 2011. viii + 264 pp. $24.95 paper. Journal of Cold War Studies, 18(2):201–203, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/620561;   http://muse.jhu.edu/article/620561/
pdf;   https://direct.mit.edu/jcws/article-pdf/18/2/201/
698905/jcws_r_00632.pdf.

Reilly:2016:SGZ

[Rei16] James Reilly. Shu Guang Zhang, Beijing’s Economic Statecraft during the Cold War: 1949–1991. Washington, DC: Woodrow Wilson Center Press, 2014. 496 pp. $65.00. Journal of Cold War Studies, 18(4):259–260, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645935;   http://muse.jhu.edu/article/645935/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/259/700192/
jcws_r_00709.pdf.

Reis:2019:DCW

[Rei19] Bruno C. Reis. Decentering the Cold War in Southern Africa: The Portuguese policy of decolonization and détente in Angola and Mozambique (1974–1984). Journal of Cold War Studies, 21 (1):3–51, Winter 2019. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
723551;   https://direct.mit.edu/jcws/article-pdf/21/1/3/
699781/jcws_a_00873.pdf.

Rempe:2008:AJJ

[Rem08] Dennis M. Rempe. Anthony James Joes, America and Guerrilla Warfare; and Anthony James Joes, Resisting Rebellion: The History and Politics of Counterinsurgency. Journal of Cold War Studies, 10(3):129–133, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241500;   http://muse.jhu.edu/article/241500/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/129/697105/
jcws.2008.10.3.129.pdf.

Reppert:2003:RSN

[Rep03] John C. Reppert. Russian Strategic Nuclear Forces (review). Journal of Cold War Studies, 5(2):124–125, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40772;   http://muse.jhu.edu/
article/40772/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/124/700414/jcws.2003.5.2.124.pdf.

Reppy:2012:TGM

[Rep12] Judith Reppy. Thomas G.Mahnken, Technology and the American Way of War since 1945. New York: Columbia University Press, 2008. 244 pp. Journal of Cold War Studies, 14(3):215–216, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/490190;   http:
//muse.jhu.edu/article/490190/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/3/215/698073/jcws_r_00263.pdf.

Reynolds:2002:WSC

[Rey02] David Reynolds. Whitehall and the Suez Crisis (review). Journal of Cold War Studies, 4(1):108–110, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9236;   http://muse.jhu.edu/article/9236/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/1/108/695520/
jcws.2002.4.1.108.pdf.

Reynolds:2023:BRM

[Rey23] David Reynolds. Book review: Mountbatten, Cold War and Empire, 1945–79 by Adrian Smith. Journal of Cold War Studies, 25(4):255– 257, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/4/
255/2204983/jcws_r_01180.pdf.

Robb:2015:ATD

[RG15] Thomas K. Robb and David James Gill. The ANZUS treaty during the Cold War: a reinterpretation of U.S. diplomacy in the Southwest Pacific. Journal of Cold War Studies, 17(4):109–157, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/606023;   http:
//muse.jhu.edu/article/606023/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/4/109/698790/jcws_a_00599.pdf.

Riabov:2017:GAE

[Ria17] Oleg Riabov. Gendering the American enemy in early Cold War Soviet films (1946–1953). Journal of Cold War Studies, 19(1):193–219, Winter 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://www.mitpressjournals.org/doi/full/10.
1162/JCWS_a_00722;   https://direct.mit.edu/jcws/article-
pdf/19/1/193/699110/jcws_a_00722.pdf.

Richmond:2004:CIH

[Ric04] Yale Richmond. Culture and International History (review). Journal of Cold War Studies, 6(4):153–155, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/174720;   http://muse.jhu.edu/article/174720/
pdf;   https://direct.mit.edu/jcws/article-pdf/6/4/153/696083/
jcws.2004.6.4.153.pdf.

Richelson:2007:CPT

[Ric07] Jeffrey T. Richelson. Curtis Peebles, Twilight Warriors: Covert Air Operations against the USSR. Journal of Cold War Studies, 9(1):113–114, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210099;   http://muse.jhu.edu/article/210099/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/113/696457/
jcws.2007.9.1.113.pdf.

Richelson:2009:ATQ

[Ric09] Jeffrey T. Richelson. Athan Theoharis, The Quest for Absolute Security: The Failed Relations among U.S. Intelligence Agencies. Journal of Cold War Studies, 11(4):153–155, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/366260;   http://muse.jhu.edu/article/366260/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/4/153/
1627629/jcws.2009.11.4.153.pdf.

Richelson:2010:MHJ

[Ric10] Jeffrey T. Richelson. Michael Holzman, James Jesus Angleton, the CIA, and the Craft of Counterintelligence. Journal of Cold War Studies, 12(4):193–195, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/403807;   http://muse.jhu.edu/article/403807/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/4/193/697558/
jcws_r_00059.pdf.

Richmond:2012:AFS

[Ric12] Yale Richmond. Ali Fisher and Scott Lucas, eds., Trials of Engagement: The Future of U.S. Public Diplomacy. Leiden: Martinus Nijhoff Publishers, 2011. 320 pp. Journal of Cold War Studies, 14(2):165–166, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/478785;   http://muse.jhu.edu/article/478785/pdf;

 https://direct.mit.edu/jcws/article-pdf/14/2/165/697997/
jcws_r_00230.pdf.

Richelson:2015:SIC

[Ric15] Jeffrey T. Richelson. Strategic Intelligence in the Cold War and Beyond by Jefferson Adams (review). Journal of Cold War Studies, 17(3):286–287, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595107;   http://muse.jhu.edu/article/595107/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/286/
1102432/jcws_r_00553.pdf.

Richardson:2022:BRR

[Ric22] Peter Richardson. The Beatles and the 1960s: Reception, revolution and social change. Journal of Cold War Studies, 24(3):238–240, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/3/
238/2042005/jcws_r_01096.pdf.

Rider:2016:CTS

[Rid16] Toby C. Rider. A campaign of truth: The State Department, propaganda, and the Olympic Games, 1950–1952. Journal of Cold War Studies, 18(2):4–27, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/620545;   http://muse.jhu.edu/article/620545/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/2/4/698836/jcws_
a_00636.pdf.

Rieber:2008:MKV

[Rie08] Alfred J. Rieber. Manfred Kittel, Vertreibung der Vertriebenen? Der historische deutsche Osten in der Erinnerungskultur der Bundesrepublik (1961–1982). Journal of Cold War Studies, 10(3):202–203, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241532;   http://muse.jhu.edu/article/241532/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/202/697175/
jcws.2008.10.3.202.pdf.

Rieber:2015:MSC

[Rie15] Alfred J. Rieber. Molotov: Stalin’s Cold Warrior by Geoffrey Roberts (review). Journal of Cold War Studies, 17(3):249–250, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595089;   http:

//muse.jhu.edu/article/595089/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/249/1102492/jcws_r_00578.pdf.

Ringstad:2012:EAC

[Rin12] Arnold Ringstad. The evolution of American civil defense film rhetoric. Journal of Cold War Studies, 14(4):93–121, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/500330;   http:
//muse.jhu.edu/article/500330/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/4/93/698118/jcws_a_00277.pdf.

Riste:2014:SBC

[Ris14] Olav Riste. “Stay Behind”: a clandestine Cold War phenomenon. Journal of Cold War Studies, 16(4):35–59, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.

jhu.edu/article/566652;   http://muse.jhu.edu/article/566652/
pdf;   https://direct.mit.edu/jcws/article-pdf/16/4/35/698595/
jcws_a_00515.pdf.

Riste:2015:SNU

[Ris15] Olav Riste. Second to None: U.S. Intelligence Activities in Northern Europe 1943–1946 by Peer Henrik Hansen (review). Journal of Cold War Studies, 17(3):258–260, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595093;   http://muse.jhu.edu/article/595093/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/258/
1102498/jcws_r_00591.pdf.

Roberts:2010:FMK

[RLV^{+}10] Priscilla Roberts, Steven I. Levine, Péter Vámos, Deborah Kaple, Jeremy Friedman, et al. Forum: Mao, Khrushchev, and China’s split with the USSR: Perspectives on the Sino–Soviet split. Journal of Cold War Studies, 12(1):120–165, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/12/1/120/1627427/jcws.2010.12.1.120.pdf.

Roberts:2002:LLP

[Rob02a] Geoffrey Roberts. Litvinov’s lost peace, 1941–1946. Journal of Cold War Studies, 4(2):23–54, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9261;   http://muse.jhu.edu/article/
9261/pdf;   https://direct.mit.edu/jcws/article-pdf/4/2/
23/695545/152039702753649638.pdf.

Roberts:2002:SPN

[Rob02b] Geoffrey Roberts. Stalin, the pact with Nazi Germany, and the origins of postwar Soviet diplomatic historiography. Journal of Cold War Studies, 4(4):93–103, Fall 2002. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/9306;
 http://muse.jhu.edu/article/9306/pdf;   https://direct.
mit.edu/jcws/article-pdf/4/4/93/695662/15203970260209527.
pdf.

Roberts:2004:GSM

[Rob04] Cynthia Roberts. German and Soviet military doctrinal innovation before World War II. Journal of Cold War Studies, 6(4): 140–144, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/174721;
http://muse.jhu.edu/article/174721/pdf;   https://direct.
mit.edu/jcws/article-pdf/6/4/140/696054/1520397042350946.
pdf.

Roberts:2007:STY

[Rob07] Geoffrey Roberts. Stalin at the Tehran, Yalta, and Potsdam conferences. Journal of Cold War Studies, 9(4):6–40, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/223554;   http:
//muse.jhu.edu/article/223554/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/4/6/696773/jcws.2007.9.4.6.pdf.

Roberts:2008:VPS

[Rob08] Geoffrey Roberts. Vladimir Pechatnov, Stalin, Ruzvel’t, Trumen: SSSR i SShA v 1940-kh gg. Journal of Cold War Studies, 10(1):159–161, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231593;   http://muse.jhu.edu/article/231593/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/159/700149/
jcws.2008.10.1.159.pdf.

Roberts:2010:SSS

[Rob10] Priscilla Roberts. The Sino–Soviet Split: Cold War in the Communist World (review). Journal of Cold War Studies, 12(1):120–128, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/382489;   http://
muse.jhu.edu/article/382489/pdf.

Roberts:2014:MRC

[Rob14] Sean P. Roberts. Markku Ruotsila, Churchill and Finland: a Study in Anti-Communism and Geopolitics. New York: Routledge, 2005. 199 pp.; and Craig Gerrard, The Foreign Office and Finland 1938–1940: Diplomatic Sideshow. London: Frank Cass, 2005. 189 pp. Journal of Cold War Studies, 16(1):235–237, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/543743;   http://muse.jhu.edu/article/543743/
pdf;   https://direct.mit.edu/jcws/article-pdf/16/1/235/
698967/jcws_r_00421.pdf.

Roblou:2020:JRS

[Rob20] Yann Roblou. J. Richard Stevens, Captain America, Masculinity, and Violence: The Evolution of a National Icon. Journal of Cold War Studies, 22(2):173–175, Spring 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/2/173/1860556/jcws_r_00886.pdf;
https://muse.jhu.edu/article/756832.

Rogers:2015:ACW

[Rog15] James I. Rogers. The American Culture of War: The History of U.S. Military Force from World War II to Operation Enduring Freedom by Adrian R. Lewis (review). Journal of Cold War Studies, 17(3):231–232, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595079;   http://muse.jhu.edu/article/595079/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/231/
1102462/jcws_r_00568.pdf.

Rohde:2017:SBS

[Roh17] Joy Rohde. Sarah Bridger, Scientists at War: The Ethics of Cold War Weapons Research. Cambridge, MA: Harvard University Press, 2015. 350 pp. $45.00. Journal of Cold War Studies, 19(3):254–256, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668177;   http://muse.jhu.edu/article/668177/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/254/699329/
jcws_r_00737.pdf.

Rohrlich:2018:ACI

[Röh18] Elisabeth Röhrlich. An attitude of caution: The IAEA, the UN, and the 1958 Pugwash Conference in Austria. Journal of Cold War Studies, 20(1):31–57, Winter 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/691017;   http://muse.jhu.edu/article/691017/pdf;
 https://direct.mit.edu/jcws/article-pdf/20/1/31/699294/
jcws_a_00800.pdf.

Romero:2017:LCS

[Rom17] Federico Romero. Leonardo Campus, I sei giorni che sconvolsero il mondo: La crisi dei missili di Cuba e le sue percezioni internazionali [Six days that shook the world: The Cuban missile crisis and its international perception]. Florence: Le Monnier, 2014. 541 pp. EUR 28.00. Journal of Cold War Studies, 19(3):245–247, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668173;   http://muse.jhu.edu/article/668173/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/245/699303/
jcws_r_00733.pdf.

Rosenberg:2000:RUF

[Ros00] Jonathan Rosenberg. Race and U.S. Foreign Policy During the Cold War (review). Journal of Cold War Studies, 2(2):112–113, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9148;   http:/
/muse.jhu.edu/article/9148/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/2/112/695251/jwcs.2000.2.2.112.pdf.

Rosenberg:2004:CWC

[Ros04] Jonathan Seth Rosenberg. The Cold War and the Color Line: American Race Relations in the Global Arena (review). Journal of Cold War Studies, 6(2):69–71, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/53681;   http://muse.jhu.edu/
article/53681/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/2/69/695967/152039704773254786.pdf.

Rosenberg:2005:WFR

[Ros05] Jonathan Rosenberg. Window on Freedom: Race, Civil Rights, and Foreign Affairs, 1945–1988 (review). Journal of Cold War Studies, 7(4):139–140, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188276;   http://muse.jhu.edu/article/188276/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/139/696234/
jcws.2005.7.4.139.pdf.

Ross:2006:SHU

[Ros06] George Ross. Stanley Henig, The Uniting of Europe: From Consolidation to Enlargement, 2nd ed. Journal of Cold War Studies, 8(2):168–170, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196293;   http://muse.jhu.edu/article/196293/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/168/700556/
jcws.2006.8.2.168.pdf.

Ross:2008:PRE

[Ros08] Robert Ross. Priscilla Roberts, ed., Behind the Bamboo Curtain: China, Vietnam, and the World beyond Asia. Journal of Cold War Studies, 10(2):143–146, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/237164;   http://muse.jhu.edu/article/237164/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/2/143/696985/
jcws.2008.10.2.143.pdf.

Rossabi:2009:XLR

[Ros09] Morris Rossabi. Xiaoyuan Liu, Reins of Liberation: an Entangled History of Mongolian Independence, Chinese Territoriality, and Great Power Hegemony, 1911–1950. Journal of Cold War Studies, 11(1): 164–165, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/259194;
http://muse.jhu.edu/article/259194/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/1/164/1468526/jcws.2009.11.1.
164.pdf.

Rose:2010:TCD

[Ros10a] Kenneth D. Rose. Tracy C. Davis, Stages of Emergency: Cold War Nuclear Civil Defense. Journal of Cold War Studies, 12(3):182–184, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/392805;   http:
//muse.jhu.edu/article/392805/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/3/182/697682/jcws_r_00032.pdf.

Rosendorf:2010:DJL

[Ros10b] Neal M. Rosendorf. Daniel J. Leab, Orwell Subverted: The CIA and the Filming of Animal Farm. Journal of Cold War Studies, 12(3):179–182, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392804;   http://muse.jhu.edu/article/392804/pdf;

 https://direct.mit.edu/jcws/article-pdf/12/3/179/697692/
jcws_r_00033.pdf.

Rossman:2012:CSG

[Ros12] Jeffrey J. Rossman. Commentary on Stalin’s genocides. Journal of Cold War Studies, 14(3):179–183, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/490177;   http://muse.jhu.edu/article/490177/

pdf. See [GRS^{+}12].

Rosen:2017:DHP

[Ros17] Lawrence Rosen. David H. Price, Cold War Anthropology: The CIA, the Pentagon, and the Growth of Dual Use Anthropology. Durham, NC: Duke University Press, 2016. 452 pp. $29.95. Journal of Cold War Studies, 19(3):239–241, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668170;   http://muse.jhu.edu/article/668170/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/239/699252/
jcws_r_00730.pdf.

Roth:2021:FWA

[Rot21] Sarah Roth. The Free World: Art and Thought in the Cold War by Louis Menand (review). Journal of Cold War Studies, 23(4):263–265, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/4/263/1970097/jcws_r_01053.pdf;
https://muse.jhu.edu/article/835890.

Rozman:2003:DNV

[Roz03] Gilbert Rozman. Distant Neighbors, Vol 1: Japanese–Russian Relations under Brezhnev and Andropov, and: Distant Neighbors, Vol. 2: Japanese–Russian Relations under Gorbachev and Yeltsin (reviews). Journal of Cold War Studies, 5(2):98–101, Spring 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40773;   http://muse.jhu.edu/
article/40773/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/98/700441/jcws.2003.5.2.98.pdf.

Rozman:2011:DDE

[Roz11] Gilbert Rozman. Donald E. Davis and Eugene P. Trani, Distorted Mirrors: Americans and Their Relations with Russia and China in the Twentieth Century. Columbia: University of Missouri Press, 2009. 461 pp. Journal of Cold War Studies, 13(4):249–250, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457215;   http:
//muse.jhu.edu/article/457215/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/4/249/697778/jcws_r_00127.pdf.

Radchenko:2013:GOF

[RT13] Sergey Radchenko and Lisbeth Tarlow. Gorbachev, Ozawa, and the failed back-channel negotiations of 1989–1990. Journal of Cold War Studies, 15(2):104–130, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/509705;   http://muse.jhu.edu/article/509705/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/2/104/698161/
jcws_a_00339.pdf.

Radchenko:2019:CSA

[RTYD19] Sergey Radchenko, Joseph Torigian, Radoslav Yordanov, and Frank Dikötter. Chinese society amid Mao’s Great Proletarian Cultural Revolution: The roots and nature of the tragedy. Journal of Cold War Studies, 21(2):174–196, Spring 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/726153;   https://direct.mit.
edu/jcws/article-pdf/21/2/174/699566/jcws_c_00880.pdf.

Ruane:2004:BRE

[Rua04] Kevin Ruane. Britain’s Retreat from East of Suez: The Choice between Europe and the World? (review). Journal of Cold War Studies, 6(4):161–163, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/174722;   http://muse.jhu.edu/article/174722/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/4/161/696107/
jcws.2004.6.4.161.pdf.

Rubenstein:2002:IEB

[Rub02] Joshua Rubenstein. Ilya Ehrenburg — between East and West. Journal of Cold War Studies, 4(1):44–65, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9237;   http://muse.jhu.edu/article/
9237/pdf;   https://direct.mit.edu/jcws/article-pdf/4/1/
44/695502/152039702753344825.pdf.

Rubenstein:2006:ALS

[Rub06] Joshua Rubenstein. Arno Lustiger, Stalin and the Jews: The Red Book: The Tragedy of the Soviet Jews and the Jewish Anti-Fascist Committee, trans. by Mary Beth Friedrich and Todd Bludeau. Journal of Cold War Studies, 8(4):169–172, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/204629;   http:
//muse.jhu.edu/article/204629/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/4/169/700823/jcws.2006.8.4.169.pdf.

Rubenstein:2010:VZZ

[Rub10] Joshua Rubenstein. Vladislav Zubok, Zhivago’s Children: The Last Russian Intelligentsia. Journal of Cold War Studies, 12(3):171–173, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/392801;   http:
//muse.jhu.edu/article/392801/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/3/171/697672/jcws_r_00026.pdf.

Rubinstein:2011:SSH

[Rub11] Murray A. Rubinstein. Shih-Shan Henry Tsai, Lee Teng-hui and Taiwan’s Quest for Identity (New York: Palgrave Macmillan, 2005); and Richard C. Kagan, Taiwan’s Statesman: Lee Teng-hui and Democracy in Asia. Annapolis, MD: Naval Institute Press, 2007. 231 pp. Journal of Cold War Studies, 13(4):250–252, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/457216;   http://muse.jhu.edu/article/457216/
pdf;   https://direct.mit.edu/jcws/article-pdf/13/4/250/
697786/jcws_r_00128.pdf.

Rubenstein:2012:CSG

[Rub12] Joshua Rubenstein. Commentary on Stalin’s genocides. Journal of Cold War Studies, 14(3):151–152, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/490177;   http://muse.jhu.edu/article/490177/

pdf. See [GRS^{+}12].

Rubenstein:2020:DWC

[Rub20] Joshua Rubenstein. Duncan White, Cold Warriors: Writers Who Waged the Literary Cold War. Journal of Cold War Studies, 22(3):248–250, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/3/248/1860611/jcws_r_00953.pdf;
https://muse.jhu.edu/article/762897.

Ruggenthaler:2011:SNG

[Rug11] Peter Ruggenthaler. The 1952 Stalin note on German unification: The ongoing debate. Journal of Cold War Studies, 13(4):172–212, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457197;   http:
//muse.jhu.edu/article/457197/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/4/172/697784/jcws_a_00145.pdf.

Ruggenthaler:2016:AHE

[Rug16a] Peter Ruggenthaler. Andreas Hilger, ed., Diplomatie für die deutsche Einheit: Dokumente des Auswärtigen Amts zu den deutsch-sowjetischen Beziehungen 1989/90. Munich: Oldenbourg, 2011. Journal of Cold War Studies, 18(2):203–205, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/620557;   http:
//muse.jhu.edu/article/620557/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/2/203/698910/jcws_r_00633.pdf.

Ruggenthaler:2016:FCC

[Rug16b] Peter Ruggenthaler. Franz Cede and Christian Prosl, Anspruch und Wirklichkeit: Österreichs Außenpolitik seit 1945. Innsbruck: Studien-Verlag, 2015. 168 pp. Journal of Cold War Studies, 18(4):251–254, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/645932;   http:
//muse.jhu.edu/article/645932/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/4/251/700176/jcws_r_00705.pdf.

Ruggenthaler:2016:GWS

[Rug16c] Peter Ruggenthaler. Gerhard Wettig, Die Stalin-Note: Historische Kontroverse im Spiegel der Quellen. Berlin: be.bra, 2015. 303 pp. Journal of Cold War Studies, 18(3):218–219, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/632298;   http:
//muse.jhu.edu/article/632298/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/3/218/699063/jcws_r_00665.pdf.

Ruggenthaler:2017:MGE

[Rug17] Peter Ruggenthaler. Maximilian Graf, ed., Österreich und die DDR 1949–1990: Politik und Wirtschaft im Schatten der deutschen Teilung. Journal of Cold War Studies, 19(4):264–266, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/683687;   http:

//muse.jhu.edu/article/683687/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/4/264/699451/jcws_r_00787.pdf.

Ruggenthaler:2018:EAA

[Rug18] Peter Ruggenthaler. Die Einheit: Das Auswärtige Amt, das DDR-Außenministerium und der Zwei-plus-Vier-Prozess ed. by Horst Möller, et al. (review). Journal of Cold War Studies, 20(2):157–159, Spring 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/697065;   https:
//direct.mit.edu/jcws/article-pdf/20/2/157/699447/jcws_r_
00811.pdf.

Russell:2004:SWP

[Rus04a] James A. Russell. The Savage Wars of Peace: Small Wars and the Rise of American Power (review). Journal of Cold War Studies, 6(3): 124–126, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/171175;
http://muse.jhu.edu/article/171175/pdf;   https://direct.
mit.edu/jcws/article-pdf/6/3/124/696021/1520397041447337.
pdf.

Russett:2004:OEU

[Rus04b] Bruce M. Russett. Our Enemies and US: America’s Rivalries and the Making of Political Science (review). Journal of Cold War Studies, 6(4):147–149, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/174723;   http://muse.jhu.edu/article/174723/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/4/147/696071/
jcws.2004.6.4.147.pdf.

Russell:2005:WIC

[Rus05] James A. Russell. Western Intelligence and the Collapse of the Soviet Union 1980–1990 (review). Journal of Cold War Studies, 7(4):187–188, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188277;   http://muse.jhu.edu/article/188277/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/187/696345/
jcws.2005.7.4.187.pdf.

Russell:2007:SBM

[Rus07] James A. Russell. Stephen Biddle, Military Power: Explaining Victory and Defeat in Modern Battle. Princeton, NJ: Princeton University Press, 2004. 337 pp. Journal of Cold War Studies, 9(4):125–127, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223555;   http://muse.jhu.edu/article/223555/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/125/696731/
jcws.2007.9.4.125.pdf.

Rusi:2016:MES

[Rus16] Alpo Rusi. Mary Elise Sarotte, The Collapse: The Accidental Opening of the Berlin Wall. New York: Basic Books, 2014. 291 pp. $27.99. Journal of Cold War Studies, 18(3):206–208, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/632292;   http:
//muse.jhu.edu/article/632292/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/3/206/699008/jcws_r_00644.pdf.

Rutland:2006:ABL

[Rut06] Peter Rutland. Archie Brown and Lilia Shevtsova, eds., Gorbachev, Yeltsin, Putin: Political Leadership in Russia’s Transition. Journal of Cold War Studies, 8(4):177–178, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/204632;   http://muse.jhu.edu/article/204632/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/4/177/700817/
jcws.2006.8.4.177.pdf.

Ruthchild:2010:HGA

[Rut10] Rochelle Goldberg Ruthchild. Helena Goscilo and Andrea Lanoux, eds., Gender and National Identity in Twentieth-Century Russian Culture. Journal of Cold War Studies, 12(1):206–208, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/382496;
http://muse.jhu.edu/article/382496/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/1/206/1627451/jcws.2010.12.1.
206.pdf.

Rutland:2011:TJC

[Rut11] Peter Rutland. Timothy J. Colton, Yeltsin: a Life. New York: Basic Books, 2008. 640 pp. $35.00. Journal of Cold War Studies, 13(4):230–233, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457207;   http://muse.jhu.edu/article/457207/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/230/697904/
jcws_r_00187.pdf.

Ruthchild:2012:ARS

[Rut12a] Rochelle Goldberg Ruthchild. Amy E. Randall, The Soviet Dream World of Retail Trade and Consumption in the 1930s. New York: Palgrave Macmillan, 2008. xi + 251 pp. $89.00. Journal of Cold War Studies, 14(2):168–170, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/478787;   http://muse.jhu.edu/article/478787/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/2/168/698003/
jcws_r_00239.pdf.

Ruthchild:2012:MPF

[Rut12b] Rochelle Goldberg Ruthchild. My Perestroika: a Film by Robin Hessman, Red Square Productions, 2010. DVD, 87 min. Journal of Cold War Studies, 14(3):222–225, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/490194;   http://muse.jhu.edu/article/490194/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/3/222/698090/
jcws_r_00267.pdf.

Ryan:2008:JBI

[Rya08a] David Ryan. Jim Brown, Impact Zone: The Battle of the DMZ in Vietnam, 1967–1968. Journal of Cold War Studies, 10(2):146– 147, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/237165;
http://muse.jhu.edu/article/237165/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/2/146/696995/jcws.2008.10.2.
146.pdf.

Ryan:2008:LGS

[Rya08b] David Ryan. Lesley Gill, The School of the Americas: Military Training and Political Violence in the Americas. Journal of Cold War Studies, 10(3):143–144, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241505;   http://muse.jhu.edu/article/241505/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/143/697083/
jcws.2008.10.3.143.pdf.

Ryan:2010:BBC

[Rya10] James G. Ryan. Brian E. Birdnow, Communism, Anti-Communism, and the Federal Courts in Missouri, 1952–1958: The Trial of the St. Louis Five. Journal of Cold War Studies, 12(4):190–191, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/403805;   http:
//muse.jhu.edu/article/403805/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/4/190/697566/jcws_r_00057.pdf.

Ryder:2014:PPC

[Ryd14] J. Luke Ryder. Priest, Politician, Collaborator: Jozef Tiso and the Making of Fascist Slovakia by James Mace Ward (review). Journal of Cold War Studies, 16(3):215–217, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/556083;   http://muse.jhu.edu/article/556083/
pdf;   https://direct.mit.edu/jcws/article-pdf/16/3/215/
698521/jcws_r_00459.pdf.

Sachse:2018:MPS

[Sac18] Carola Sachse. The Max Planck Society and Pugwash during the Cold War: an uneasy relationship. Journal of Cold War Studies, 20(1):170–209, Winter 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/691021;   http://muse.jhu.edu/article/691021/pdf;
 https://direct.mit.edu/jcws/article-pdf/20/1/170/699348/
jcws_a_00804.pdf.

Saeki:2007:NSA

[Sae07] Chizuru Saeki. Naoko Shibusawa, America’s Geisha Ally: Reimagining the Japanese Enemy. Cambridge, MA: Harvard University Press, 2006. 397 pp. Journal of Cold War Studies, 9(4):155–158, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/223556;   http:
//muse.jhu.edu/article/223556/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/4/155/696747/jcws.2007.9.4.155.pdf.

Saikal:2002:SDV

[Sai02] Amin Saikal. Superpowers Defeated: Vietnam and Afghanistan Compared (review). Journal of Cold War Studies, 4(2):133–135, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9262;   http:/
/muse.jhu.edu/article/9262/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/2/133/695597/jcws.2002.4.2.133.pdf.

Sakwa:2006:YMH

[Sak06] Richard Sakwa. Yoshiko M. Herrera, Imagined Economies: The Sources of Russian Regionalism. Journal of Cold War Studies, 8(4):175–176, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/204633;   http://muse.jhu.edu/article/204633/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/4/175/700831/
jcws.2006.8.4.175.pdf.

Salmon:2006:JAB

[Sal06] Patrick Salmon. Juhana Aunesluoma, Britain, Sweden and the Cold War, 1945–54: Understanding Neutrality. Journal of Cold War Studies, 8(2):166–168, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196294;   http://muse.jhu.edu/article/196294/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/166/700546/
jcws.2006.8.2.166.pdf.

Sandrow:2016:RLY

[San16] Nahma Sandrow. Rina Lapidus, Young Jewish Poets Who Fell as Soviet Soldiers in the Second World War. New York: Routledge, 2014. 240 pp. $145.00. Journal of Cold War Studies, 18(4):254–257, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/645933;   http:
//muse.jhu.edu/article/645933/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/4/254/700158/jcws_r_00706.pdf.

ndergaard:2023:CFW

[SÃņ23] Rasmus Sinding SÃņdergaard. The Committee for the Free World and the defense of democracy. Journal of Cold War Studies, 25(4): 70–100, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/25/4/70/2204993/jcws_a_01171.pdf.

Sapolsky:2003:SCW

[Sap03] Harvey M. Sapolsky. Science, Cold War, and the American State: Lloyd V. Berkner and the Balance of Professional Ideals (review). Journal of Cold War Studies, 5(3):118–120, Summer 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45211;   http://muse.jhu.edu/
article/45211/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/118/695829/jcws.2003.5.3.118.pdf.

Sarantakes:2005:CWP

[Sar05a] Nicholas Evan Sarantakes. Cold war pop culture and the image of U.S. foreign policy: The perspective of the original star trek series. Journal of Cold War Studies, 7(4):74–103, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/188278;   http:
//muse.jhu.edu/article/188278/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/4/74/696259/1520397055012488.pdf.

Sarotte:2005:LST

[Sar05b] Mary Elise Sarotte. Licensed to Spy: With the Top Secret Military Liaison Mission in East Germany, and: Germany’s Cold War: The Global Campaign to Isolate East Germany, 1949– 1969 (review). Journal of Cold War Studies, 7(4):177–178, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/188279;   http:
//muse.jhu.edu/article/188279/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/4/177/696335/jcws.2005.7.4.177.pdf.

Sarotte:2011:FBM

[Sar11] Mary Sarotte. Frédéric Bozo, Mitterrand, the End of the Cold War, and German Unification, trans. by Susan Emanuel. New York: Berghahn, 2009. Journal of Cold War Studies, 13(4):257–258, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457219;   http:
//muse.jhu.edu/article/457219/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/4/257/697813/jcws_r_00131.pdf.

Sargent:2016:PLE

[Sar16] Daniel Sargent. Piers Ludlow, ed., European Integration and the Cold War: Ostpolitik-Westpolitik, 1965–1973. New York: Routledge, 2007. 194 pp. $150.00. Journal of Cold War Studies, 18(4):241–243, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/645927;   http:
//muse.jhu.edu/article/645927/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/4/241/700129/jcws_r_00700.pdf.

Sarkar:2019:UPC

[Sar19] Jayita Sarkar. U.S. policy to curb West European nuclear exports, 1974–1978. Journal of Cold War Studies, 21(2):110–149, Spring 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/726151;   https://direct.mit.
edu/jcws/article-pdf/21/2/110/699557/jcws_a_00877.pdf.

Sarantakes:2023:OCW

[Sar23] Nicholas Evan Sarantakes. The Olympics and the Cold War: a historiography. Journal of Cold War Studies, 25(4):127–158, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/4/
127/2204985/jcws_a_01173.pdf.

Sassoon:2014:EGM

[Sas14a] Joseph Sassoon. The East German Ministry for state security and Iraq, 1968–1989. Journal of Cold War Studies, 16(1):4–23, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/543723;
http://muse.jhu.edu/article/543723/pdf;   https://direct.
mit.edu/jcws/article-pdf/16/1/4/698902/jcws_a_00429.pdf.

Sassoon:2014:NAB

[Sas14b] Joseph Sassoon. Nigel Ashton and Bryan Gibson, eds., The Iran–Iraq War: New International Perspectives. New York: Routledge, 2013. 245 pp. $135.00. Journal of Cold War Studies, 16(1):216–218, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543734;   http:
//muse.jhu.edu/article/543734/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/1/216/698930/jcws_r_00412.pdf.

Saul:2012:LDJ

[Sau12] Scott Saul. Lisa E. Davenport, Jazz Diplomacy: Promoting America in the Cold War Era. Jackson: University Press of Mississippi, 2009. 219 pp. $50.00. Journal of Cold War Studies, 14(3):198–199, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/490182;   http://muse.jhu.edu/article/490182/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/3/198/698071/
jcws_r_00255.pdf.

Savranskaya:2009:ABE

[Sav09] Svetlana Savranskaya. Archie Brown, ed., The Demise of Marxism– Leninism in Russia. Journal of Cold War Studies, 11(1):172– 174, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/259198;
http://muse.jhu.edu/article/259198/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/1/172/1468536/jcws.2009.11.1.
172.pdf.

Savranskaya:2014:DTU

[Sav14] Svetlana Savranskaya. Divided Together: The United States and the Soviet Union in the United Nations, 1945–1965 by Ilya V. Gaiduk (review). Journal of Cold War Studies, 16(4):279–281, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/566684;   http:
//muse.jhu.edu/article/566684/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/4/279/698672/jcws_r_00503.pdf.

Sayle:2017:FBF

[Say17] Timothy Andrews Sayle. Frédéric Bozo, French Foreign Policy since 1945: an Introduction. Journal of Cold War Studies, 19(4):239–241, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683675;   http://muse.jhu.edu/article/683675/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/239/699404/
jcws_r_00775.pdf.

Sbardellati:2005:PDP

[Sba05] John Sbardellati. Power to Destroy: The Political Uses of the IRS from Kennedy to Nixon (review). Journal of Cold War Studies, 7(3):158–159, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185117;   http://muse.jhu.edu/article/185117/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/158/1102567/
jcws.2005.7.3.158.pdf.

Schaller:2000:EDJ

[Sch00a] Michael Schaller. Embracing Defeat: Japan in the Wake of World War II (review). Journal of Cold War Studies, 2(2):108–109, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9149;   http:/
/muse.jhu.edu/article/9149/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/2/108/695243/15203970051032336.pdf.

Schoenfeld:2000:RTS

[Sch00b] Gabriel Schoenfeld. The Road to Terror: Stalin and the Self-Destruction of the Bolsheviks, 1932–1939 (review). Journal of Cold War Studies, 2(3):143–145, Fall 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9162;   http://muse.jhu.edu/article/9162/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/3/143/695311/
jcws.2000.2.3.143.pdf.

Schwartz:2001:CMC

[Sch01] Stephen I. Schwartz. Critical Masses: Citizens, Nuclear Weapons Production, and Environmental Destruction in the United States and Russia (review). Journal of Cold War Studies, 3(3):107–109, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9211;   http:/
/muse.jhu.edu/article/9211/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/107/695464/jcws.2001.3.3.107.pdf.

Schmemann:2002:FCA

[Sch02a] Serge Schmemann. Failed Crusade: America and the Tragedy of Post-Communist Russia (review). Journal of Cold War Studies, 4(3):153– 155, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9289;   http:/
/muse.jhu.edu/article/9289/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/153/700258/jcws.2002.4.3.153.pdf.

Schrecker:2002:IIB

[Sch02b] Ellen Schrecker. Imagining Internationalism in British and American Labor, 1939–49 (review). Journal of Cold War Studies, 4(1):106– 108, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9238;   http:/
/muse.jhu.edu/article/9238/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/1/106/695534/jcws.2002.4.1.106.pdf.

Schweller:2002:WPA

[Sch02c] Randall L. Schweller. From War to Peace: Altered Strategic Landscapes in the Twentieth Century (review). Journal of Cold War Studies, 4(4):133–135, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9307;   http://muse.jhu.edu/article/9307/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/4/133/695712/
jcws.2002.4.4.133.pdf.

Schroeder:2003:RTH

[Sch03] Paul W. Schroeder. Rethinking Theory and History in the Cold War: The State, Military Power and Social Revolution (review). Journal of Cold War Studies, 5(2):82–84, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40774;   http://muse.jhu.edu/
article/40774/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/82/700436/jcws.2003.5.2.82.pdf.

Schwartz:2005:AAK

[Sch05] Thomas Alan Schwartz. The Ambivalent Alliance: Konrad Adenauer, The CDU/CSU and the West, 1949–1966 (review). Journal of Cold War Studies, 7(3):171–173, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185118;   http://muse.jhu.edu/article/185118/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/171/1102575/
jcws.2005.7.3.171.pdf.

Schaefer:2008:TLE

[Sch08] Bernd Schaefer. Thomas Lindenberger, ed., Massenmedien im Kalten Krieg: Akteure, Bilder, Resonanzen [Mass Media in Cold War: Actors, Images, Resonances]. Journal of Cold War Studies, 10(3):166–169, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241515;   http://muse.jhu.edu/article/241515/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/166/697113/
jcws.2008.10.3.166.pdf.

Schaller:2009:AAP

[Sch09] Michael Schaller. Alan Armstrong, Preemptive Strike: The Secret Plan That Would Have Prevented the Attack on Pearl Harbor. Journal of Cold War Studies, 11(2):143–144, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/265872;   http://muse.jhu.edu/article/265872/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/2/143/
697189/jcws.2009.11.2.143.pdf.

Schaefer:2010:SCK

[Sch10] Bernd Schaefer. Stefan Creuzberger, Kampf für die Einheit: Das gesamtdeutsche Ministerium und die politische Kultur des Kalten Krieges 1949–1969. Journal of Cold War Studies, 12(4):199–201, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/403810;   http:
//muse.jhu.edu/article/403810/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/4/199/697554/jcws_r_00062.pdf.

Schneider:2011:JKM

[Sch11] Eric C. Schneider. Jeremy Kuzmarov, The Myth of the Addicted Army: Vietnam and the Modern War on Drugs. Amherst: University of Massachusetts Press, 2009. 303 pp. $26.95 paper. Journal of Cold War Studies, 13(3):214–215, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447473;   http://muse.jhu.edu/article/447473/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/214/700740/
jcws_r_00153.pdf.

Schaller:2013:JME

[Sch13] Michael Schaller. James I. Matray, ed., Northeast Asia and the Legacy of Harry S. Truman: Japan, China, and the Two Koreas. Kirksville, MO: Truman State University Press, 2012. 362 pp. $34.95. Journal of Cold War Studies, 15(1):168–169, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/505758;   http://muse.jhu.edu/article/505758/
pdf;   https://direct.mit.edu/jcws/article-pdf/15/1/168/
700016/jcws_r_00321.pdf.

Schmidl:2016:LNC

[Sch16] Erwin A. Schmidl. Lukewarm neutrality in a Cold War?: The case of Austria. Journal of Cold War Studies, 18(4):36–50, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/645907;   http:
//muse.jhu.edu/article/645907/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/4/36/700029/jcws_a_00679.pdf.

Schumacher:2017:GMT

[Sch17] Frank Schumacher. Gregory M. Tomlin, Murrow’s Cold War: Public Diplomacy for the Kennedy Administration. Journal of Cold War Studies, 19(4):241–243, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683676;   http://muse.jhu.edu/article/683676/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/241/699417/
jcws_r_00776.pdf.

Schmidli:2019:SMH

[Sch19] William Michael Schmidli. From Selma to Moscow: How Human Rights Activists Transformed U.S. Foreign Policy by Sarah B. Snyder (review). Journal of Cold War Studies, 21(2):197–199, Spring 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/726154;   https://direct.mit.
edu/jcws/article-pdf/21/2/197/699553/jcws_r_00881.pdf.

Schroeder:2020:AJW

[Sch20] James Schroeder. Audra J. Wolfe, Freedom’s Laboratory: The Cold War Struggle for the Soul of Science. Journal of Cold War Studies, 22(4):234–237, Fall 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/4/234/1860647/jcws_r_00974.pdf;
https://muse.jhu.edu/article/776970.

Schmidt:2022:TSR

[Sch22a] Elizabeth Schmidt. Thomas Sankara: A revolutionary in Cold War Africa. Journal of Cold War Studies, 24(3):240–242, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/3/
240/2041991/jcws_r_01066.pdf.

Schretter:2022:BRS

[Sch22b] Lukas Schretter. Book review: Scars of War: The Politics of Paternity and Responsibility for the Amerasians of Vietnam by Sabrina Thomas. Journal of Cold War Studies, 24(4):226–228, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/4/
226/2062884/jcws_r_01111.pdf.

Scionti:2020:AAC

[Sci20] Andrea Scionti. “I am afraid Americans cannot understand”: The Congress for Cultural Freedom in France and Italy, 1950–1957. Journal of Cold War Studies, 22(1):89–124, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/22/1/89/699813/jcws_a_
00927.pdf;   https://muse.jhu.edu/article/749656.

Scott:2009:MDO

[Sco09] Len Scott. Michael Dobbs, One Minute to Midnight: Kennedy, Khrushchev, and Castro on the Brink of Nuclear War. Journal of Cold War Studies, 11(4):151–153, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/366261;   http://muse.jhu.edu/article/366261/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/4/151/
1627623/jcws.2009.11.4.151.pdf.

Scott:2017:BUH

[Sco17] Ian Scott. Bryn Upton, Hollywood and the End of the Cold War: Signs of Cinematic Change. London: Rowman & Littlefield, 2014. 195 pp.

$70.00 (£44.95) hardcover; $69.99 (£44.95) eBook. Journal of Cold War Studies, 19(3):231–233, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668166;   http://muse.jhu.edu/article/668166/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/231/699206/
jcws_r_00725.pdf.

Selvage:2002:CWS

[Sel02] Douglas Selvage. A Cold War in the Soviet Bloc: Polish–East German Relations, 1945–1962 (review). Journal of Cold War Studies, 4(4): 128–130, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9308;   http:/
/muse.jhu.edu/article/9308/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/4/128/695706/jcws.2002.4.4.128.pdf.

Selvage:2019:ODE

[Sel19] Douglas Selvage. Operation “Denver”: The East German Ministry of State Security and the KGB’s AIDS disinformation campaign, 1985–1986 (part 1). Journal of Cold War Studies, 21(4):71– 123, Fall 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
21/4/71/699714/jcws_a_00907.pdf;   https://muse.jhu.edu/
article/736847.

Selvage:2021:HMS

[Sel21a] Douglas Selvage. From Helsinki to “Mars”: Soviet-bloc active measures and the struggle over détente in Europe, 1975–1983. Journal of Cold War Studies, 23(4):34–94, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/23/4/34/1970103/jcws_a_
01039.pdf;   https://muse.jhu.edu/article/835877.

Selvage:2021:ODE

[Sel21b] Douglas Selvage. Operation “Denver”: The East German Ministry for State Security and the KGB’s AIDS disinformation campaign, 1986–1989 (part 2). Journal of Cold War Studies, 23(3):4–80, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
23/3/4/1955787/jcws_a_01024.pdf;   https://muse.jhu.edu/
article/801905.

Senn:2016:SFC

[Sen16] Alfred Erich Senn. Stephen F. Cohen, Soviet Fates and Lost Alternatives: From Stalinism to the New Cold War. New York: Columbia University Press, 2009. xiv + 308 pp. $28.50. Journal of Cold War Studies, 18(3):238–240, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632306;   http://muse.jhu.edu/article/632306/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/238/699067/
jcws_r_00672.pdf.

Sepp:2001:UNM

[Sep01] Kalev Sepp. The U.S. Naval Mission to Haiti, 1959–1963 (review). Journal of Cold War Studies, 3(1):140–143, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9182;   http://muse.jhu.edu/article/
9182/pdf;   https://direct.mit.edu/jcws/article-pdf/3/1/
140/695430/jcws.2001.3.1.140.pdf.

Sergi:2001:GUU

[Ser01] Bruno S. Sergi. German Unification and the Union of Europe: The Domestic Politics of Integration Policy (review). Journal of Cold War Studies, 3(1):133–135, Winter 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9183;   http://muse.jhu.edu/article/9183/pdf;
 https://direct.mit.edu/jcws/article-pdf/3/1/133/695434/
jcws.2001.3.1.133.pdf.

Serfaty:2013:LKE

[Ser13] Simon Serfaty. Ludger Kuhnhardt, ed., Crises in European Integration: Challenges and Responses, 1945–2005. New York: Berghahn Books, 2009. Journal of Cold War Studies, 15(2):148–150, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/509713;   http:
//muse.jhu.edu/article/509713/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/2/148/698177/jcws_r_00354.pdf.

Sewell:2023:BRC

[Sew23] Kenneth Sewell. Book review: The CIA’s Greatest Covert Operation: Inside the Daring Mission to Recover a Nuclear-Armed Soviet Sub by David H. Sharp. Journal of Cold War Studies, 25(4):259–261, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/4/
259/2204991/jcws_r_01182.pdf.

Stuart-Fox:2011:LCG

[SF11] Martin Stuart-Fox. Laurent Cesari, Les grandes puissances et le Laos, 1954–1964. Arras, France: Artois Presses Université, 2007. 374 pp. EUR 22.00. Journal of Cold War Studies, 13(3):240–242, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/447487;   http:
//muse.jhu.edu/article/447487/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/3/240/700714/jcws_r_00120.pdf.

Sfikas:2001:WPS

[Sfi01] Thanasis D. Sfikas. War and peace in the strategy of the Communist Party of Greece, 1945–1949. Journal of Cold War Studies, 3(3): 5–30, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9212;   http:/
/muse.jhu.edu/article/9212/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/5/695440/152039701750419493.pdf.

Sfikas:2002:AWW

[Sfi02] Thanasis D. Sfikas. After the War Was Over: Reconstructing the Family, Nation, and State in Greece, 1943–1960 (review). Journal of Cold War Studies, 4(3):137–140, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9290;   http://muse.jhu.edu/article/9290/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/3/137/700240/
jcws.2002.4.3.137.pdf.

Shaw:2001:PCW

[Sha01] Tony Shaw. The politics of Cold War culture. Journal of Cold War Studies, 3(3):59–76, Fall 2001. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/9213;
 http://muse.jhu.edu/article/9213/pdf;   https://direct.
mit.edu/jcws/article-pdf/3/3/59/695444/152039701750419510.
pdf.

Shaw:2002:MMM

[Sha02] Tony Shaw. Martyrs, miracles, and Martians: Religion and Cold War cinematic propaganda in the 1950s. Journal of Cold War Studies, 4(2): 3–22, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9263;   http:/
/muse.jhu.edu/article/9263/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/2/3/695565/152039702753649629.pdf.

Shaw:2006:RCW

[Sha06] Tony Shaw. Rethinking Cold War Culture (review). Journal of Cold War Studies, 8(1):144–146, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/192219;   http://muse.jhu.edu/article/192219/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/1/144/696327/
jcws.2006.8.1.144.pdf.

Shannon:2007:DLA

[Sha07a] Vaughn P. Shannon. Douglas Little, American Orientalism: The United States and the Middle East since 1945. Chapel Hill: The University of North Carolina Press, 2004. 424 pp. $19.95. Journal of Cold War Studies, 9(4):149–151, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223557;   http://muse.jhu.edu/article/223557/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/149/696745/
jcws.2007.9.4.149.pdf.

Shaw:2007:GSS

[Sha07b] Tony Shaw. Giles Scott-Smith and Hans Krabbendam, eds., The Cultural Cold War in Western Europe, 1945–1960. Journal of Cold War Studies, 9(1):105–106, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210100;   http://muse.jhu.edu/article/210100/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/105/696461/
jcws.2007.9.1.105.pdf.

Shaw:2008:ADB

[Sha08a] Tony Shaw. Andrew Defty, Britain, America and Anti-Communist Propaganda 1945–53: The Information Research Department; and James R. Vaughan, The Failure of American and British Propaganda in the Middle East, 1945–1957: Unconquerable Minds. Journal of Cold War Studies, 10(1):171–173, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/231594;   http://muse.jhu.edu/article/231594/
pdf;   https://direct.mit.edu/jcws/article-pdf/10/1/171/
700131/jcws.2008.10.1.171.pdf.

Shaw:2008:DSB

[Sha08b] Tony Shaw. David Seed, Brainwashing: The Fictions of Mind Control — A Study of Novels and Films since World War II. Journal of Cold War Studies, 10(3):185–187, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241524;   http://muse.jhu.edu/article/241524/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/185/697155/
jcws.2008.10.3.185.pdf.

Share:2011:STY

[Sha11] Michael Share. Steve Tsang Yui-Sang. The Cold War’s Odd Couple: The Unintended Partnership between the Republic of China and the UK, 1950–1958. London: I. B. Tauris, 2006. 269 pp. $84.95. Journal of Cold War Studies, 13(4):272–273, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457225;   http://muse.jhu.edu/article/457225/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/272/697827/
jcws_r_00157.pdf.

Shaw:2012:NNP

[Sha12] Tony Shaw. Nightmare on Nevsky Prospekt: The blue bird as a curious instance of U.S.–Soviet film collaboration during the Cold War. Journal of Cold War Studies, 14(1):3–33, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/467676;   http:
//muse.jhu.edu/article/467676/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/1/3/697825/jcws_a_00174.pdf.

Shackelford:2018:EKK

[Sha18] Philip Shackelford. Edward Kaplan, To Kill Nations: American Strategy in the Air-Atomic Age and the Rise of Mutually Assured Destruction. Journal of Cold War Studies, 20(4):227–229, Fall 2018. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718205;   https://direct.mit.
edu/jcws/article-pdf/20/4/227/1969608/jcws_r_00807.pdf.

Shackelford:2020:SSW

[Sha20] Philip C. Shackelford. Stalin’s Secret Weapon: The Origins of Soviet Biological Warfare by Anthony Rimmington (review). Journal of Cold War Studies, 22(1):252–253, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/22/1/252/699879/jcws_r_
00936.pdf;   https://muse.jhu.edu/article/749645.

Sheetz:2003:FGQ

[She03] Mark S. Sheetz. France and the German question: Avant-garde or rearguard: Comment on Creswell and Trachtenberg. Journal of Cold War Studies, 5(3):37–45, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45212;   http://muse.jhu.edu/
article/45212/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/37/695943/jcws.2003.5.3.37.pdf. See [CT03a].

Sheng:2006:MTK

[She06] Michael M. Sheng. Mao, Tibet, and the Korean war. Journal of Cold War Studies, 8(3):15–33, Summer 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/201241;   http://muse.jhu.edu/article/201241/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/3/15/696400/jcws.
2006.8.3.15.pdf.

Sherry:2007:MJH

[She07] Michael S. Sherry. Michael J. Hogan and Thomas G. Paterson, eds., Exploring the History of American Foreign Relations. Journal of Cold War Studies, 9(1):96–98, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210101;   http://muse.jhu.edu/article/210101/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/96/696592/jcws.
2007.9.1.96.pdf.

Sheehan:2012:MWW

[She12a] James J. Sheehan. Manfred Wilke, Der Weg zur Mauer: Stationen der Teilungsgeschichte. Berlin: Ch. Links Verlag, 2011. 472 pp. EUR 39.90. Journal of Cold War Studies, 14(3):200–201, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/490183;   http:
//muse.jhu.edu/article/490183/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/3/200/698061/jcws_r_00256.pdf.

Sheng:2012:C

[She12b] Michael Sheng. Commentary. Journal of Cold War Studies, 14(1): 96–99, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/467683;
http://muse.jhu.edu/article/467683/pdf.

Sherry:2012:JMC

[She12c] Michael Sherry. Joseph Maiolo, Cry Havoc: How the Arms Race Drove the World to War, 1931–1941. New York: Basic Books, 2010. 473 pp. $35.00. Journal of Cold War Studies, 14(1):133–134, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/467695;   http:
//muse.jhu.edu/article/467695/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/1/133/697900/jcws_r_00206.pdf.

Sherry:2013:JMC

[She13] Michael Sherry. Joseph Maiolo, Cry Havoc: How the Arms Race Drove the World to War, 1931–1941. New York: Basic Books, 2010. 473 pp. $35.00. Journal of Cold War Studies, 15(1):155–157, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/505764;   http:
//muse.jhu.edu/article/505764/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/1/155/700000/jcws_r_00315.pdf.

Sheetz:2014:BEI

[She14] Mark Sheetz. Building a European Identity: France, the United States, and the Oil Shock, 1973–1974 by Aurélie Elisa Gfeller (review). Journal of Cold War Studies, 16(4):285–287, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/566687;   http:
//muse.jhu.edu/article/566687/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/4/285/698664/jcws_r_00506.pdf.

She:2020:CWC

[She20] Gangzheng She. The Cold War and Chinese policy toward the Arab– Israeli conflict, 1963–1975. Journal of Cold War Studies, 22(1):125– 174, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/1/125/699806/jcws_a_00928.pdf;   https://muse.jhu.edu/
article/749655.

Shinji:2013:SRP

[Shi13] Yokote Shinji. Soviet repatriation policy, U.S. occupation authorities, and Japan’s entry into the Cold War. Journal of Cold War Studies, 15 (2):30–50, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/509702;
http://muse.jhu.edu/article/509702/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/2/30/698143/jcws_a_00336.pdf.

Shlapentokh:2006:GRM

[Shl06] Vladimir Shlapentokh. Gábor T. Rittersporn, Malte Rolf, Jan C. Behrends, eds., Spharen von Offentlichkeit in Gesellschaften sowietischen Typs Zwischen partei-staatlicher Selbstinszenierung und kirchlichen Gegenwelten : Public Spheres in Soviet-Type Societies: Between the Great Show of the Party-State and Religious Counter-Cultures. Journal of Cold War Studies, 8(4):161–163, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/204634;   http:
//muse.jhu.edu/article/204634/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/4/161/700807/jcws.2006.8.4.161.pdf.

Shlapentokh:2009:GPE

[Shl09] Vladimir Shlapentokh. György Péteri, ed., Nylon Curtain: Transnational and Trans-Systemic Tendencies in the Cultural Life of State-Socialist Russia and East-Central Europe. Journal of Cold War Studies, 11(1):175–176, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/259199;   http://muse.jhu.edu/article/259199/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/1/175/
1468550/jcws.2009.11.1.175.pdf.

Shmelev:2023:BRC

[Shm23] Anatol Shmelev. Book review: Cold War Radio: The Russian Broadcasts of the Voice of America and Radio Free Europe/Radio Liberty by Mark Pomar and Under the Radar: Tracking Western Radio Listeners in the Soviet Union by R. Eugene Parta. Journal of Cold War Studies, 25(4):261–265, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/4/261/2204995/jcws_r_01183.pdf.

Shore:2015:PAD

[Sho15] Zachary Shore. Provoking America: Le Duan and the origins of the Vietnam War. Journal of Cold War Studies, 17(4):86– 108, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/606022;
http://muse.jhu.edu/article/606022/pdf;   https://direct.
mit.edu/jcws/article-pdf/17/4/86/698761/jcws_a_00598.pdf.

Sibley:2011:JEH

[Sib11] Katherine Sibley. John Earl Haynes and Harvey Klehr, Early Cold War Spies: The Espionage Trials That Shaped American Politics.

New York: Cambridge University Press, 2006. 251 pp. Journal of Cold War Studies, 13(1):240–242, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/424255;   http://muse.jhu.edu/article/424255/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/1/240/697761/
jcws_r_00086.pdf.

Siekmeier:2006:RKO

[Sie06] James F. Siekmeier. Robert Kirkland, Observing our Hermanos de Armas: U.S. Military Attachés in Guatemala, Cuba, and Bolivia, 1950–1964. Journal of Cold War Studies, 8(2):132–133, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/196291;   http:
//muse.jhu.edu/article/196291/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/2/132/700518/jcws.2006.8.2.132.pdf.

Siljak:1999:TRM

[Sil99] Ana Siljak. Time and Revolution: Marxism and the Design of Soviet Institutions (review). Journal of Cold War Studies, 1(1):119–122, Winter 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9094;   https://
direct.mit.edu/jcws/article-pdf/1/1/119/695111/jcws.1999.
1.1.119.pdf.

Siljak:2004:PPI

[Sil04] Ana Siljak. Power and Persuasion: Ideology and Rhetoric in Communist Yugoslavia, 1944–1953 (review). Journal of Cold War Studies, 6(2):82–84, Spring 2004. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
53682;   http://muse.jhu.edu/article/53682/pdf;   https:/
/direct.mit.edu/jcws/article-pdf/6/2/82/695995/jcws.2004.
6.2.82.pdf.

Silins:2005:INH

[Sil05] Ints Silins. Imagining the Nation: History, Modernity, and Revolution in Latvia (review). Journal of Cold War Studies, 7 (1):208–211, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/179570;   http://muse.jhu.edu/article/179570/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/1/208/700500/
jcws.2005.7.1.208.pdf.

Silva:2011:CHP

[Sil11] Patricio Silva. Carlos Huneeus, The Pinochet Regime. Boulder, CO: Lynne Rienner Publishers, 2007. 559 pp. $69.95. Journal of Cold War Studies, 13(3):242–244, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447488;   http://muse.jhu.edu/article/447488/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/242/700724/
jcws_r_00121.pdf.

Simons:2004:USP

[Sim04] Thomas W. Simons, Jr. The United States and Pakistan, 1947– 2000: Disenchanted Allies (review). Journal of Cold War Studies, 6(4):166–168, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/174724;   http://muse.jhu.edu/article/174724/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/4/166/696093/
jcws.2004.6.4.166.pdf.

Simons:2010:WPF

[Sim10] Thomas W. Simons, Jr. When policy failed: a view from the middle distance. Journal of Cold War Studies, 12(2):95–101, Spring 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/382204;   http:
//muse.jhu.edu/article/382204/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/2/95/697513/jcws.2010.12.2.95.pdf.

Simons:2014:UFP

[Sim14] Thomas W. Simons. US Foreign Policy and Democracy Promotion: From Theodore Roosevelt to Barack Obama edited by Michael Cox, Timothy J. Lynch, and Nicolas Bouchet (review). Journal of Cold War Studies, 16(4):289–291, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566689;   http://muse.jhu.edu/article/566689/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/289/698682/
jcws_r_00508.pdf.

Simons:2016:APR

[Sim16] Thomas W. Simons, Jr. Andrzej Paczkowski, Revolution and Counterrevolution in Poland, 1980–1989, trans. by Christine Manetti. Rochester, NY: University of Rochester Press, 2015. 387 pp. $99.00. Journal of Cold War Studies, 18(4):237–241, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/645926;   http:
//muse.jhu.edu/article/645926/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/4/237/700172/jcws_r_00698.pdf.

Simons:2017:LSW

[Sim17] Thomas W. Simons, Jr. Louis Sell, From Washington to Moscow: US–Soviet Relations and the Collapse of the USSR. Journal of Cold War Studies, 19(4):249–251, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683680;   http://muse.jhu.edu/article/683680/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/249/699423/
jcws_r_00780.pdf.

Simons:2023:RCR

[Sim23] Jr. Thomas W. Simons. Reykjavik up close: Reagan and Gorbachev, October 1986 and after. Journal of Cold War Studies, 25(4):159–190, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/4/
159/2204981/jcws_a_01174.pdf.

Sinitsky:2018:JFW

[Sin18] Julia Sinitsky. Jeremy Friedman, Winning the Third World: Sino-American Rivalry during the Cold War? Shadow Cold War: The Sino-Soviet Competition for the Third World. Journal of Cold War Studies, 20(4):245–248, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718214;   https://direct.mit.
edu/jcws/article-pdf/20/4/245/1969646/jcws_r_00834.pdf.

Szalontai:2023:MBB

[SJ23] Balázs Szalontai and Yoo Jinil. Maneuvering between Baghdad and Tehran: North Korea’s relations with Iraq and Iran during the Cold War. Journal of Cold War Studies, 25(2):179–247, Spring 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/2/
179/2140970/jcws_a_01119.pdf.

Sanz-Lafuente:2023:GLD

[SL23] Gloria Sanz-Lafuente. From global to local: The development of heavy water in international nuclear programs (1945–1970). Journal of Cold War Studies, 25(3):35–67, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/3/35/2158686/jcws_a_01158.pdf.

Stelzl-Marx:2012:DSA

[SM12] Barbara Stelzl-Marx. Death to spies!: Austrian informants for Western intelligence services and Soviet capital punishment during the occupation of Austria. Journal of Cold War Studies, 14(4):167–196, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500332;   http://muse.jhu.edu/article/500332/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/167/698116/
jcws_a_00279.pdf.

Smetana:2013:GFK

[Sme13] Vít Smetana. George F. Kennan and the division of Europe. Journal of Cold War Studies, 15(4):225–232, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/534413;   http://muse.jhu.edu/article/534413/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/4/225/698351/
jcws_a_00406.pdf.

Smetana:2015:ULC

[Sme15] Vít Smetana. The U.S. “Loss” of Czechoslovakia: On the edge of historical truth. Journal of Cold War Studies, 17(3):220–226, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595075;   http:
//muse.jhu.edu/article/595075/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/220/1102442/jcws_a_00565.pdf.

Smith:2003:FNA

[Smi03] Tony Smith. The French North African Crisis: Colonial Breakdown and Anglo–French Relations, 1945–62 (review). Journal of Cold War Studies, 5(2):91–92, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40775;   http://muse.jhu.edu/
article/40775/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/91/700416/jcws.2003.5.2.91.pdf.

Smith:2007:PJH

[Smi07] Tony Smith. Patrick J. Haney and Walt Vanderbush, The Cuban Embargo: The Domestic Politics of an American Foreign Policy. Pittsburgh: University of Pittsburgh Press, 2005. 222 pp. Journal of Cold War Studies, 9(4):140–142, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223558;   http://muse.jhu.edu/article/223558/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/140/696727/
jcws.2007.9.4.140.pdf.

Smith:2008:FFC

[Smi08] Marcia S. Smith. Francis French and Colin Burgess, Into that Silent Sea: Trailblazers of the Space Era, 1961–1965. Journal of Cold War Studies, 10(3):209–211, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241536;   http://muse.jhu.edu/article/241536/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/209/697163/
jcws.2008.10.3.209.pdf.

Smith:2011:HGE

[Smi11] Marcia S. Smith. Von Hardesty and Gene Eisman, Epic Rivalry: The Inside Story of the Soviet and American Space Race. Washington, DC: National Geographic, 2007. 275 pp. $16.95. Journal of Cold War Studies, 13(4):221–222, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457202;   http://muse.jhu.edu/article/457202/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/221/697857/
jcws_r_00181.pdf.

Smidrkalova:2023:ASB

[Šmi23] Michaela Šmidrkalová. Atoms for socialism: The birth of a Czechoslovak–Soviet nuclear utopia. Journal of Cold War Studies, 25(3):112–141, Summer 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/3/112/2158684/jcws_a_01161.pdf.

Snitar:2022:TSM

[Sni22] Corina Snitar. Trăgători şi Mistificatori: Contrarevoluţia Securită ţii în decembrie 1989 [(Romanian). Shooters and Mysticists: The Security Counter-Revolution in December 1989], by Andrei Ursu and Roland O. Thomasson. Journal of Cold War Studies, 24(2):170–172, Spring 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/2/
170/2013018/jcws_r_01084.pdf.

Snyder:1999:RUQ

[Sny99] Timothy Snyder. “To resolve the Ukrainian question once and for all”: The ethnic cleansing of Ukrainians in Poland, 1943–1947. Journal of Cold War Studies, 1(2):86–120, Spring 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9104;   https://direct.mit.edu/jcws/
article-pdf/1/2/86/695137/15203979952559531.pdf.

Snyder:2003:RNE

[Sny03a] Timothy Snyder. Redrawing Nations: Ethnic Cleansing in East– Central Europe, 1944–1948 (review). Journal of Cold War Studies, 5(3):102–105, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/45213;   http://muse.jhu.edu/article/45213/pdf.

Snyder:2003:TWE

[Sny03b] Timothy Snyder. Trophies of War and Empire: The Archival Heritage of Ukraine, World War II, and the International Politics of Restitution (review). Journal of Cold War Studies, 5(2):101–103, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/40776;   http://
muse.jhu.edu/article/40776/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/2/101/700396/jcws.2003.5.2.101.pdf.

Snyder:2006:YIE

[Sny06] Timothy Snyder. Yaroslav Isaevych, ed., Volyn’ i Kholmshchyna 1938–1947 rr.: Pols’ko-ukrains’ke protystoyannya ta ioho vidlunnya: Doslidzhennya, dokumenty, spohady. Journal of Cold War Studies, 8(4):157–160, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/204635;   http://muse.jhu.edu/article/204635/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/4/157/700811/
jcws.2006.8.4.157.pdf.

Snyder:2007:KBB

[Sny07] Timothy Snyder. Kate Brown, A Biography of No Place: From Ethnic Borderland to Soviet Heartland. Journal of Cold War Studies, 9(1):152–153, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210102;   http://muse.jhu.edu/article/210102/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/152/696534/
jcws.2007.9.1.152.pdf.

Snyder:2008:GHF

[Sny08] David J. Snyder. Gerald Horne, The Final Victim of the Blacklist: John Howard Lawson, Dean of the Hollywood Ten. Journal of Cold War Studies, 10(3):134–139, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241502;   http://muse.jhu.edu/article/241502/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/134/697087/
jcws.2008.10.3.134.pdf.

Snyder:2009:AJK

[Sny09] Timothy Snyder. Alex J. Kay, Exploitation, Resettlement, Mass Murder: Political and Economic Planning for German Occupation Policy in the Soviet Union, 1940–1941. Journal of Cold War Studies, 11(1): 166–167, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/259195;
http://muse.jhu.edu/article/259195/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/1/166/1468544/jcws.2009.11.1.
166.pdf.

Snyder:2011:MES

[Sny11] Sarah B. Snyder. Mary Elise Sarotte, 1989: The Struggle to Create Post-Cold War Europe. Princeton, NJ: Princeton University Press, 2009. 344 pp. $29.95. Journal of Cold War Studies, 13(1):258–260, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/424264;   http://muse.jhu.edu/article/424264/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/1/258/697759/
jcws_r_00095.pdf.

Snyder:2013:DRL

[Sny13] David J. Snyder. Domesticity, rearmament, and the limits of U.S. public diplomacy in the Netherlands during the early Cold War. Journal of Cold War Studies, 15(3):47–75, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/522709;
http://muse.jhu.edu/article/522709/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/3/47/698284/jcws_a_00370.pdf.

Snyder:2017:FCM

[Sny17] Sarah B. Snyder. Frank Costigliola and Michael J. Hogan, eds., Explaining the History of American Foreign Relations, 3rd ed. New York: Cambridge University Press, 2016. 389 pp. Journal of Cold War Studies, 19(3):233–234, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668167;   http://muse.jhu.edu/article/668167/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/233/699218/
jcws_r_00727.pdf.

Soares:2006:SIH

[Soa06] John A. Soares, Jr. Strategy, ideology, and human rights: Jimmy Carter confronts the left in Central America, 1979–1981. Journal of Cold War Studies, 8(4):57–91, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/204636;   http://muse.jhu.edu/article/204636/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/4/57/700827/jcws.
2006.8.4.57.pdf.

Soares:2009:SWD

[Soa09] John A. Soares, Jr. Stephen Wagg and David L. Andrews, eds., East Plays West: Sport and the Cold War. Journal of Cold War Studies, 11(2):160–162, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/265863;   http://muse.jhu.edu/article/265863/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/2/160/697199/
jcws.2009.11.2.160.pdf.

Soares:2013:NES

[Soa13] John Soares. Nicholas Evan Sarantakes, Dropping the Torch: Jimmy Carter, the Olympic Boycott and the Cold War. New York: Cambridge University Press, 2011. xvi + 340 pp. Journal of Cold War Studies, 15(3):194–196, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522721;   http://muse.jhu.edu/article/522721/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/194/698292/
jcws_r_00348.pdf.

Soares:2020:DAW

[Soa20a] John Soares. Defending the American Way of Life: Sport, Culture and the Cold War ed. by Toby C. Rider and Kevin Witherspoon (review). Journal of Cold War Studies, 22(1):247–249, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/22/1/247/699860/jcws_r_
00934.pdf;   https://muse.jhu.edu/article/749647.

Soares:2020:PAS

[Soa20b] John Soares. Pan-Asian Sports and the Emergence of Modern Asia, 1913–1974 by Stefan Huebner (review). Journal of Cold War Studies, 22(1):264–265, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:

//direct.mit.edu/jcws/article-pdf/22/1/264/699833/jcws_r_
00917.pdf;   https://muse.jhu.edu/article/749638.

Soares:2022:BRE

[Soa22] John Soares. Book review: The Ends of Modernization: Nicaragua and the United States in the Cold War Era by David Johnson Lee. Journal of Cold War Studies, 24(4):221–223, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/24/4/221/2062882/jcws_r_
01109.pdf.

Soares:2023:BRD

[Soa23] John Soares. Book review: Degrees of Difficulty: How Women’s Gymnastics Rose to Prominence and Fell from Grace by Georgia Cervin. Journal of Cold War Studies, 25(1):237–239, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/1/
237/2073751/jcws_r_01137.pdf.

Sobers:2019:III

[Sob19] Candace Sobers. Independence, intervention, and internationalism: Angola and the international system, 1974–1975. Journal of Cold War Studies, 21(1):97–124, Winter 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/723554;   https://direct.mit.
edu/jcws/article-pdf/21/1/97/699734/jcws_a_00854.pdf.

Sokov:1999:EVS

[Sok99] Nikolai N. Sokov. Etiudy o vremeni (Sketches about time) (review). Journal of Cold War Studies, 1(2):127–128, Spring 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9105;   https://direct.mit.edu/jcws/
article-pdf/1/2/127/695154/jcws.1999.1.2.127.pdf.

Solchanyk:2013:WJR

[Sol13] Roman Solchanyk. William Jay Risch, The Ukrainian West: Culture and the Fate of Empire in Soviet Lviv. Cambridge, MA: Harvard University Press, 2011. 360 pp. $49.95. Journal of Cold War Studies, 15(1):176–178, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/505753;   http://muse.jhu.edu/article/505753/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/1/176/700014/
jcws_r_00326.pdf.

Song:2002:OCR

[Son02] Yongyi Song. The Origins of the Cultural Revolution, Vol. 3: The Coming of the Cataclysm, 1961–1966 (review). Journal of Cold War Studies, 4(2):125–127, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9264;   http://muse.jhu.edu/article/9264/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/2/125/695591/
jcws.2002.4.2.125.pdf.

Sondergaard:2019:DPN

[Søn19] Rasmus Sinding Søndergaard. Democracy Promotion, National Security and Strategy: Foreign Policy under the Reagan Administration by Robert Pee (review). Journal of Cold War Studies, 21(2):199–201, Spring 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/726141;   https:
//direct.mit.edu/jcws/article-pdf/21/2/199/699545/jcws_r_
00882.pdf.

Sonmez:2022:AMM

[Sön22] Erdem Sönmez. America and the making of modern Turkey: Science, culture and political alliances. Journal of Cold War Studies, 24(3): 242–244, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/24/3/242/2041995/jcws_r_01067.pdf.

Souchen:2021:MWU

[Sou21] Alex Souchen. Military Waste: The Unexpected Consequences of Permanent War Readiness by Joshua O. Reno, Berkeley: University of California Press, 2019. 288 pp. $34.95. (review). Journal of Cold War Studies, 23(1):235–237, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/1/235/1911816/jcws_r_00990.pdf;
https://muse.jhu.edu/article/788393.

Spark:2014:CHP

[Spa14] Clare L. Spark. The Collaboration: Hollywood’s Pact with Hitler by Ben Urwand (review). Journal of Cold War Studies, 16(4):226–228, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/566660;   http:
//muse.jhu.edu/article/566660/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/4/226/698604/jcws_r_00480.pdf.

Sperling:2001:UFT

[Spe01] Valerie Sperling. Unarmed Forces: The Transnational Movement to End the Cold War (review). Journal of Cold War Studies, 3(3):100– 102, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9214;   http:/
/muse.jhu.edu/article/9214/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/100/695436/jcws.2001.3.3.100.pdf.

Spector:2013:PDN

[Spe13] Ronald H. Spector. Phat Diem: nationalism, religion, and identity in the Franco–Viet Minh War. Journal of Cold War Studies, 15(3): 34–46, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/522708;
http://muse.jhu.edu/article/522708/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/3/34/698278/jcws_a_00369.pdf.

Sperling:2014:JRB

[Spe14] Valerie Sperling. Joanna Regulska and Bonnie G. Smith, eds., Women and Gender in Postwar Europe: From Cold War to European Union. New York: Routledge, 2012. 243 pp. $43.95. Journal of Cold War Studies, 16(1):233–235, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/543742;   http://muse.jhu.edu/article/543742/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/1/233/698965/
jcws_r_00420.pdf.

Sperling:2015:WGP

[Spe15] Valerie Sperling. Women and Gender in Postwar Europe: From Cold War to European Union edited by Joanna Regulska and Bonnie G. Smith (review). Journal of Cold War Studies, 17(3):256–258, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595092;   http:
//muse.jhu.edu/article/595092/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/256/1102496/jcws_r_00590.pdf.

Spelling:2018:WJC

[Spe18] Alex Spelling. The Wilson–Johnson Correspondence, 1964–69 ed. by Simon C. Smith (review). Journal of Cold War Studies, 20(2):

155–157, Spring 2018. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/697064;
 https://direct.mit.edu/jcws/article-pdf/20/2/155/699427/
jcws_r_00806.pdf.

Spohr:2012:PPS

[Spo12] Kristina Spohr. Precluded or precedent-setting?: The “NATO Enlargement Question” in the triangular Bonn–Washington–Moscow diplomacy of 1990–1991. Journal of Cold War Studies, 14(4):4– 54, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/500328;
http://muse.jhu.edu/article/500328/pdf;   https://direct.
mit.edu/jcws/article-pdf/14/4/4/698114/jcws_a_00275.pdf.

Springhall:2007:CBT

[Spr07] John Springhall. Christopher Bayly and Tim Harper, Forgotten Armies: The Fall of British Asia, 1941–1945. Cambridge, MA: The Belknap Press of Harvard University Press, 2005. 555 pp. $29.95. Journal of Cold War Studies, 9(4):172–174, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/223559;   http:
//muse.jhu.edu/article/223559/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/4/172/696761/jcws.2007.9.4.172.pdf.

Sprigge:2017:EKC

[Spr17] Martha Sprigge. Elaine Kelly, Composing the Canon in the German Democratic Republic: Narratives of Nineteenth-Century Music. Journal of Cold War Studies, 19(4):262–264, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.

jhu.edu/article/683686;   http://muse.jhu.edu/article/683686/
pdf;   https://direct.mit.edu/jcws/article-pdf/19/4/262/
699441/jcws_r_00786.pdf.

Scott-Smith:2014:FEU

[SS14] Giles Scott-Smith. The free Europe University in Strasbourg: U.S. state-private networks and academic “Rollback”. Journal of Cold War Studies, 16(2):77–107, Spring 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/547325;   http://muse.jhu.edu/article/547325/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/2/77/698418/
jcws_a_00452.pdf.

Sanchez-Sibony:2018:EGG

[SS18a] Oscar Sanchez-Sibony. Economic growth in the governance of the Cold War divide: Mikoyan’s encounter with Japan, Summer 1961. Journal of Cold War Studies, 20(2):129–154, Spring 2018. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/697063;   https://direct.mit.
edu/jcws/article-pdf/20/2/129/699406/jcws_a_00741.pdf.

Scott-Smith:2018:MPC

[SS18b] Giles Scott-Smith. Michael Patrick Cullinane, David Ryan, eds., U.S. Foreign Policy and the Other. Journal of Cold War Studies, 20(4): 236–237, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/718209;
https://direct.mit.edu/jcws/article-pdf/20/4/236/1969618/
jcws_r_00827.pdf.

Stan:2017:DLD

[ST17] Marius Stan and Vladimir Tismaneanu. The death of a Leninist dictator: “The Memory of Comrade Gheorghe Gheorghiu-Dej Forever Alive in the Heart of the Party, of the Working Class, of the People”. Journal of Cold War Studies, 19(3):202–214, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/668163;   http:
//muse.jhu.edu/article/668163/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/3/202/699269/jcws_a_00759.pdf.

Stafford:2000:BSP

[Sta00] David Stafford. Britain’s Secret Propaganda War (review). Journal of Cold War Studies, 2(3):147–148, Fall 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9164;   http://muse.jhu.edu/article/9164/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/3/147/695299/
jcws.2000.2.3.147.pdf.

Statiev:2007:TSS

[Sta07] Alexander Statiev. Timothy Snyder, Sketches from a Secret War: a Polish Artist’s Mission to Liberate Soviet Ukraine. Journal of Cold War Studies, 9(3):165–168, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219468;   http://muse.jhu.edu/article/219468/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/165/696636/
jcws.2007.9.3.165.pdf.

Stan:2013:SBA

[Sta13] Lavinia Stan. Suzanne Bardgett et al., eds., Justice, Politics and Memory in Europe after the Second World War. Vol. 2: Landscapes after Battle. London: Vallentine Mitchell, 2011. 360 pp. $74.94. Journal of Cold War Studies, 15(1):172–174, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/505755;   http:
//muse.jhu.edu/article/505755/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/1/172/700004/jcws_r_00324.pdf.

Stapleton:2015:KSE

[Sta15] Bradford Ian Stapleton. The Korea syndrome: an examination of war-weariness theory. Journal of Cold War Studies, 17(3):36–81, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595071;   http:
//muse.jhu.edu/article/595071/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/36/1102434/jcws_a_00561.pdf.

Stanciu:2016:FER

[Sta16] Cezar Stanciu. Fragile equilibrium: Romania and the Vietnam War in the context of the Sino–Soviet split, 1966. Journal of Cold War Studies, 18(1):161–187, Winter 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/612010;   http://muse.jhu.edu/article/612010/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/1/161/698775/
jcws_a_00623.pdf.

Stent:1999:IOG

[Ste99] Angela E. Stent. Imperial Overstretch: Germany in Soviet Policy from Stalin to Gorbachev (review). Journal of Cold War Studies, 1 (1):117–119, Winter 1999. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/9095;
 https://direct.mit.edu/jcws/article-pdf/1/1/117/695113/
jcws.1999.1.1.117.pdf.

Steel:2002:SDC

[Ste02] Brent S. Steel. Smokestack Diplomacy: Cooperation and Conflict in East–West Environmental Politics (review). Journal of Cold War Studies, 4(3):146–148, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9291;   http://muse.jhu.edu/article/9291/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/3/146/700250/
jcws.2002.4.3.146.pdf.

Stent:2005:SCI

[Ste05] Angela E. Stent. Social Construction of International Politics: Identities and Foreign Policies, Moscow, 1955 and 1999 (review). Journal of Cold War Studies, 7(1):184–186, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/179571;   http:
//muse.jhu.edu/article/179571/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/1/184/700478/jcws.2005.7.1.184.pdf.

Stern:2008:JVB

[Ste08] Jacques Stern. J. V. Boone, A Brief History of Cryptology. Journal of Cold War Studies, 10(4):144–145, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/252664;   http://muse.jhu.edu/article/252664/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/4/144/697037/
jcws.2008.10.4.144.pdf.

Steininger:2010:HHD

[Ste10] Rolf Steininger. Hannfried von Hindenburg, Demonstrating Reconciliation: State and Society in West German Foreign Policy toward Israel, 1952–1965. Journal of Cold War Studies, 12(3):186–188, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/392807;   http:
//muse.jhu.edu/article/392807/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/3/186/697676/jcws_r_00035.pdf.

Stephanson:2013:GKD

[Ste13] Anders Stephanson. Gaddis’s Kennan: a different Kennan? Journal of Cold War Studies, 15(4):233–240, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/534414;   http://muse.jhu.edu/article/534414/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/4/233/698355/
jcws_a_00407.pdf. See [Kra13b].

Stevenson:2014:TMD

[Ste14] Michael D. Stevenson. “Tossing a match into dry hay”: Nuclear weapons and the crisis in U.S.–Canadian relations, 1962–1963. Journal of Cold War Studies, 16(4):5–34, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566651;   http://muse.jhu.edu/article/566651/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/5/698564/jcws_
a_00514.pdf.

Stent:2015:ZSS

[Ste15] Angela Stent. Zbig: The Strategy and Statecraft of Zbigniew Brzezinski edited by Charles Gati (review). Journal of Cold War Studies, 17(2):140–142, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/582437;   http://muse.jhu.edu/article/582437/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/2/140/698688/
jcws_r_00535.pdf.

Stent:2018:SCT

[Ste18] Angela Stent. Samuel Charap and Timothy J. Colton, Everyone Loses: The Ukraine Crisis and the Ruinous Contest for Post-Soviet Eurasia. Journal of Cold War Studies, 20(4):234–235, Fall 2018. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718208;   https://direct.mit.
edu/jcws/article-pdf/20/4/234/1969612/jcws_r_00810.pdf.

Steininger:2020:UNO

[Ste20] Rolf Steininger. Die UdSSR und der Nahe Osten: Zionismus, ägyptischer Antikolonialismus und sowjetische Außenpolitik bis 1956 by Wiebke Bachmann (review). Journal of Cold War Studies, 22(1): 272–274, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/22/1/272/699846/jcws_r_00921.pdf;   https://muse.jhu.
edu/article/749640.

Stiles:2006:FBA

[Sti06] David Stiles. A fusion bomb over Andalucía: U.S. information policy and the 1966 Palomares Incident. Journal of Cold War Studies, 8 (1):49–67, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/192220;
http://muse.jhu.edu/article/192220/pdf;   https://direct.
mit.edu/jcws/article-pdf/8/1/49/696283/152039706775212067.
pdf.

Stiffler:2010:SSS

[Sti10] Douglas A. Stiffler. The Sino–Soviet Split: Cold War in the Communist World (review). Journal of Cold War Studies, 12(1):151–160, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/382489;   http://
muse.jhu.edu/article/382489/pdf.

Storm:2006:AHA

[Sto06] Carl Storm. Andrew Hemingway, Artists on the Left; American Artists and the Communist Movement 1926–1956. Journal of Cold War Studies, 8(2):150–151, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196295;   http://muse.jhu.edu/article/196295/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/150/700544/
jcws.2006.8.2.150.pdf.

Stokes:2007:GRH

[Sto07] Gale Stokes. Gerd-Rainer Horn and Padraic Kenney, eds., Transnational Moments of Change: Europe 1945, 1968, 1989. Journal of Cold War Studies, 9(1):150–152, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210103;   http://muse.jhu.edu/article/210103/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/150/696554/
jcws.2007.9.1.150.pdf.

Stone:2008:CII

[Sto08] David R. Stone. CMEA’s international investment bank and the crisis of developed socialism. Journal of Cold War Studies, 10(3):48–77, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241497;   http://muse.jhu.edu/article/241497/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/48/697179/
jcws.2008.10.3.48.pdf.

Stolarik:2011:PHS

[Sto11] M. Mark Stolarik. Paul Hacker, Slovakia on the Road to Independence: an American Diplomat’s Eyewitness Account. University Park, PA: The Pennsylvania State University Press, 2010. 240 pp. $65.00. Journal of Cold War Studies, 13(4):266–267, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457222;   http:
//muse.jhu.edu/article/457222/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/4/266/697831/jcws_r_00134.pdf.

Storkmann:2014:EGM

[Sto14] Klaus Storkmann. East German military aid to the Sandinista government of Nicaragua, 1979–1990. Journal of Cold War Studies, 16 (2):56–76, Spring 2014. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/547324;

http://muse.jhu.edu/article/547324/pdf;   https://direct.
mit.edu/jcws/article-pdf/16/2/56/698405/jcws_a_00451.pdf.

Stocker:2015:ARZ

[Sto15] James Stocker. Accepting regional zero: Nuclear weapon free zones, U.S. nonproliferation policy and global security, 1957–1968. Journal of Cold War Studies, 17(2):36–72, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/582434;   http://muse.jhu.edu/article/582434/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/2/36/698670/
jcws_a_00547.pdf.

Strub:2009:DKJ

[Str09] Whitney Strub. David K. Johnson, The Lavender Scare: The Cold War Persecution of Gays and Lesbians in the Federal Government. Journal of Cold War Studies, 11(2):158–160, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/265864;   http:
//muse.jhu.edu/article/265864/pdf;   https://direct.mit.
edu/jcws/article-pdf/11/2/158/697205/jcws.2009.11.2.158.
pdf.

Stromquist:2012:AJK

[Str12] Nelly P. Stromquist. Andrew J. Kirkendall, Paulo Freire & the Cold War Politics of Literacy. Chapel Hill: University of North Carolina Press, 2010. 170 pp. Journal of Cold War Studies, 14(3):210–213, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/490188;   http://muse.jhu.edu/article/490188/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/3/210/698067/
jcws_r_00261.pdf.

Stueck:2015:ARC

[Stu15] William Stueck. The Ashgate Research Companion to the Korean War edited by James I. Matray and Donald W. Boose, Jr. (review). Journal of Cold War Studies, 17(3):240–242, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595084;   http:
//muse.jhu.edu/article/595084/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/240/1102482/jcws_r_00573.pdf.

Stueck:2019:IRK

[Stu19] William Stueck. The Interrogation Rooms of the Korean War: The Untold Story by Monica Kim (review). Journal of Cold War Studies, 21(3):258–260, Summer 2019. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
731292;   https://direct.mit.edu/jcws/article-pdf/21/3/258/
699645/jcws_r_00900.pdf.

Suh:2008:ARM

[Suh08] Dae-Sook Suh. Allan R. Millett, The War for Korea, 1945– 1950: a House Burning. Journal of Cold War Studies, 10(2):149– 152, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/237166;
http://muse.jhu.edu/article/237166/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/2/149/696997/jcws.2008.10.2.
149.pdf.

Sullivan:2022:BRE

[Sul22] Neil J. Sullivan. Book review: The Essential Writings of Vannevar Bush edited by G. Pascal Zachary. Journal of Cold War Studies, 24 (4):236–238, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/24/4/236/2062888/jcws_r_01115.pdf.

Suri:2002:EEC

[Sur02] Jeremi Suri. Explaining the end of the Cold War: a new historical consensus? Journal of Cold War Studies, 4(4):60–92, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9309;   http:/
/muse.jhu.edu/article/9309/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/4/60/695682/15203970260209518.pdf.

Suri:2013:GTM

[Sur13] Jeremi Suri. Gil Troy, Moynihan’s Moment: America’s Fight against Zionism as Racism. New York: Oxford University Press, 2013. 336 pp. $29.95. Journal of Cold War Studies, 15(1):157–159, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/505763;   http:
//muse.jhu.edu/article/505763/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/1/157/699988/jcws_r_00316.pdf.

Svik:2016:CFW

[Svi16] Peter Svik. The Czechoslovak factor in Western Alliance building, 1945–1948. Journal of Cold War Studies, 18(1):133–160, Winter 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/612009;   http:
//muse.jhu.edu/article/612009/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/1/133/698771/jcws_a_00622.pdf.

Shen:2009:HCD

[SX09] Zhihua Shen and Yafeng Xia. Hidden currents during the honeymoon: Mao, Khrushchev, and the 1957 Moscow conference. Journal of Cold War Studies, 11(4):74–117, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/366247;   http://muse.jhu.edu/article/366247/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/4/74/1627645/
jcws.2009.11.4.74.pdf.

Shen:2014:CNK

[SX14] Zhihua Shen and Yafeng Xia. Chinese–North Korean relations and Chinese policy toward Korean cross-border migration, 1950–1962. Journal of Cold War Studies, 16(4):133–158, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/566655;   http:
//muse.jhu.edu/article/566655/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/4/133/698576/jcws_a_00518.pdf.

Shaw:2017:CWS

[SY17] Tony Shaw and Denise J. Youngblood. Cold War sport, film, and propaganda: A comparative analysis of the superpowers. Journal of Cold War Studies, 19(1):160–192, Winter 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http:

//www.mitpressjournals.org/doi/full/10.1162/JCWS_a_00721;
 https://direct.mit.edu/jcws/article-pdf/19/1/160/699108/
jcws_a_00721.pdf.

Szabo:2007:NWW

[Sza07] Stephen F. Szabo. Norbert Wiggershaus and Winfried Heinemann, eds., Nationale Aussen- und Bündnispolitk der NATO-Mitgliedstaaten. Journal of Cold War Studies, 9(1):126–128, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/210104;   http:
//muse.jhu.edu/article/210104/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/1/126/696518/jcws.2007.9.1.126.pdf.

Szabo:2010:ASE

[Sza10a] Stephen F. Szabo. Alexander Stephan, ed., Americanization and Anti-Americanism: The German Encounter with American Culture after 1945. Journal of Cold War Studies, 12(3):188–189, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/392808;   http:
//muse.jhu.edu/article/392808/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/3/188/697686/jcws_r_00036.pdf.

Szalontai:2010:JVY

[Sza10b] Bálazs Szalontai. James V. Young, Eye on Korea: an Insider Account of Korean-American Relations. Journal of Cold War Studies, 12(3):155–157, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392794;   http://muse.jhu.edu/article/392794/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/155/697639/
jcws_r_00021.pdf.

Szalontai:2011:TBS

[Sza11] Balázs Szalontai. Thomas A. Bass, The Spy Who Loved Us: The Vietnam War and Pham Xuan An’s Dangerous Game. New York: PublicAffairs, 2009. 297 pp. Journal of Cold War Studies, 13(1):244–246, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/424257;   http://muse.jhu.edu/article/424257/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/1/244/697754/
jcws_r_00088.pdf.

Szabo:2012:JJA

[Sza12a] Stephen F. Szabo. Jeffrey J. Anderson and Eric Langenbacher, eds., From the Bonn to the Berlin Republic: Germany at the Twentieth Anniversary of Unification. New York: Berghahn Books, 2010. 424 pp. Journal of Cold War Studies, 14(3):201–203, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/490184;   http:
//muse.jhu.edu/article/490184/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/3/201/698069/jcws_r_00257.pdf.

Szalontai:2012:SVN

[Sza12b] Balázs Szalontai. In the shadow of Vietnam: a new look at North Korea’s militant strategy, 1962–1970. Journal of Cold War Studies, 14(4):122–166, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500331;   http://muse.jhu.edu/article/500331/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/122/698108/
jcws_a_00278.pdf.

Szalontai:2014:MCM

[Sza14] Balázs Szalontai. Modern China–Myanmar Relations: Dilemmas of Mutual Dependence by David I. Steinberg and Hongwei Fan (review). Journal of Cold War Studies, 16(4):245–248, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/566668;   http:
//muse.jhu.edu/article/566668/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/4/245/698633/jcws_r_00487.pdf.

Szalontai:2018:SLG

[Sza18] Balázs Szalontai. The “Sole Legal Government of Vietnam”: The Bao Dai factor and Soviet attitudes toward Vietnam, 1947– 1950. Journal of Cold War Studies, 20(3):3–56, Summer 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/704570;   https://direct.
mit.edu/jcws/article-pdf/20/3/3/699633/jcws_a_00813.pdf.

Szabo:2021:OEG

[Sza21] Stephen F. Szabo. Ostpolitik, 1969–1974: European and Global Responses ed. by Carole Fink and Bernd Schaffer, New York: Cambridge University Press, 2009. 289 pp. (review). Journal of Cold War Studies, 23(1):241–243, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/1/241/1911800/jcws_r_00992.pdf;
https://muse.jhu.edu/article/788389.

Szemere:2019:EPM

[Sze19] Anna Szemere. East Punk Memories dir. by Lucile Chaufour (review). Journal of Cold War Studies, 21(2):207–209, Spring 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/726145;   https://direct.mit.
edu/jcws/article-pdf/21/2/207/699573/jcws_r_00890.pdf.

Sheng:2012:PSR

[SZK12] Michael Sheng, Qiang Zhai, and Deborah Kaple. Perspectives on Sergey Radchenko’s Two Suns in the Heavens. Journal of Cold War Studies, 14(1):96–106, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467683;   http://muse.jhu.edu/article/467683/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/96/697815/
jcws_a_00196.pdf.

Szoke:2010:DRS

[Szo10] Zoltán Szok̋e. Delusion or reality?: Secret Hungarian diplomacy during the Vietnam War. Journal of Cold War Studies, 12(4):119–180, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/403799;   http://muse.jhu.edu/article/403799/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/4/119/697550/
jcws_a_00050.pdf.

Szporer:2007:TCC

[Szp07a] Michael Szporer. Triggering Communism’s Collapse: Perceptions and Power in Poland’s Transition (review). Journal of Cold War Studies, 9(2):139–143, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214756;   http://muse.jhu.edu/article/214756/pdf.

Szporer:2007:SFP

[Szp07b] Michael Szporer. The security forces and Polish Communism: Reclaiming history from myth. Journal of Cold War Studies, 9(1): 88–95, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/210105;
http://muse.jhu.edu/article/210105/pdf;   https://direct.
mit.edu/jcws/article-pdf/9/1/88/696578/jcws.2007.9.1.88.
pdf.

Szporer:2009:SCP

[Szp09] Michael Szporer. Sławomir Cenckiewicz and Piotr Gontarczyk, SB a Lech Wałesa: Przyczynek do biografii [The SB and Lech Wałesa: a Contribution toward a Biography]. Journal of Cold War Studies, 11(2):119–121, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/265882;   http://muse.jhu.edu/article/265882/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/2/119/697157/
jcws.2009.11.2.119.pdf.

Szporer:2010:MRC

[Szp10] Michael Szporer. Managing religion in Communist-era Poland: Catholic priests versus the secret police. Journal of Cold War Studies, 12(3):115–120, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392780;   http://muse.jhu.edu/article/392780/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/115/697629/
jcws_a_00004.pdf.

Szporer:2011:AWL

[Szp11] Michael Szporer. Anna Walentynowicz and the legacy of solidarity in Poland. Journal of Cold War Studies, 13(1):213–222, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/424245;   http:
//muse.jhu.edu/article/424245/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/1/213/697713/jcws_a_00075.pdf.

Szporluk:2012:CSG

[Szp12] Roman Szporluk. Commentary on Stalin’s genocides. Journal of Cold War Studies, 14(3):175–179, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/490177;   http://muse.jhu.edu/article/490177/

pdf. See [GRS^{+}12].

Szporer:2014:LWS

[Szp14] Michael Szporer. Lech Wałesa and the solidarity era: The myth revived? Journal of Cold War Studies, 16(3):205–212, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/556081;   http:
//muse.jhu.edu/article/556081/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/3/205/698507/jcws_a_00474.pdf.

Taagepera:2013:TTL

[Taa13] Rein Taagepera. Tina Tamman, The Last Ambassador: August Torma, Soldier, Diplomat, Spy. Amsterdam: Rodopi B. V., 2011. xvii + 251 pp. Journal of Cold War Studies, 15(2):163–165, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/509720;   http:
//muse.jhu.edu/article/509720/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/2/163/698215/jcws_r_00361.pdf.

Tagirova:2022:CMR

[Tag22] Alsu Tagirova. From crisis management to realignment of forces: The diplomatic “Geometry” of the 1969–1978 Sino–Soviet Border Talks. Journal of Cold War Studies, 24(1):116–154, Winter 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/24/1/116/1982401/jcws_a_
01027.pdf;   https://muse.jhu.edu/article/843079.

Taliaferro:1999:SSC

[Tal99] Jeffrey W. Taliaferro. Statecraft and Security: The Cold War and Beyond (review). Journal of Cold War Studies, 1(2):125–127, Spring 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9106;   https://
direct.mit.edu/jcws/article-pdf/1/2/125/695150/jcws.1999.
1.2.125.pdf.

Tal:2008:OSP

[Tal08] David Tal. From the Open Skies Proposal of 1955 to the Norstad Plan of 1960: a plan too far. Journal of Cold War Studies, 10(4):66–93, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/252655;   http://muse.jhu.edu/article/252655/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/4/66/697077/
jcws.2008.10.4.66.pdf.

Talbot:2011:JCS

[Tal11] Ian Talbot. Joya Chatterji, Spoils of Partition: Bengal and India 1947–1967. Cambridge, UK: Cambridge University Press, 2007. 332 pp. Journal of Cold War Studies, 13(4):244–246, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457213;   http:
//muse.jhu.edu/article/457213/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/4/244/697790/jcws_r_00125.pdf.

Tannenwald:2005:IEA

[Tan05] Nina Tannenwald. Ideas and explanation: Advancing the theoretical agenda. Journal of Cold War Studies, 7(2):13–42, Spring 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/181903;   http:
//muse.jhu.edu/article/181903/pdf;   https://direct.mit.
edu/jcws/article-pdf/7/2/13/696220/1520397053630619.pdf.

Taubman:2002:OHG

[Tau02] William Taubman. “One Hell of a Gamble”: Khrushchev, Castro and Kennedy, 1958–1964 (review). Journal of Cold War Studies, 4(1):123– 124, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9239;   http:/
/muse.jhu.edu/article/9239/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/1/123/695562/jcws.2002.4.1.123.pdf.

Taylor:2000:CSM

[Tay00] Brian D. Taylor. The Collapse of the Soviet Military (review). Journal of Cold War Studies, 2(1):132–134, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9133;   http://muse.jhu.edu/article/9133/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/1/132/695203/
jcws.2000.2.1.132.pdf.

Taylor:2003:SMD

[Tay03] Brian D. Taylor. The Soviet military and the disintegration of the USSR. Journal of Cold War Studies, 5(1):17–66, Winter 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/38256;   http://muse.jhu.edu/
article/38256/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/1/17/695765/152039703320996713.pdf.

Taylor:2015:OHC

[Tay15] Moe Taylor. “One hand can’t clap”: Guyana and North Korea, 1974–1985. Journal of Cold War Studies, 17(1):41–63, Winter 2015. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/581230;
http://muse.jhu.edu/article/581230/pdf;   https://direct.
mit.edu/jcws/article-pdf/17/1/41/700772/jcws_a_00530.pdf.

Taylor:2019:SAC

[Tay19] Keith W. Taylor. Southeast Asia’s Cold War: An Interpretive History by Ang Cheng Guan (review). Journal of Cold War Studies, 21(1): 195–196, Winter 2019. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/723540;
 https://direct.mit.edu/jcws/article-pdf/21/1/195/699779/
jcws_r_00861.pdf.

Teigrob:2011:PME

[Tei11] Robert Teigrob. Penny Marie Von Eschen, Satchmo Blows Up the World: Jazz Ambassadors Play the Cold War. Cambridge, MA:

Harvard University Press, 2004. 329 pp. $18.95. Journal of Cold War Studies, 13(4):235–237, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457209;   http://muse.jhu.edu/article/457209/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/4/235/697917/
jcws_r_00188.pdf.

Telepneva:2018:SGR

[Tel18] Natalia Telepneva. Saving Ghana’s revolution: The demise of Kwame Nkrumah and the evolution of Soviet policy in Africa, 1966–1972. Journal of Cold War Studies, 20(4):4–25, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/718198;   https://direct.
mit.edu/jcws/article-pdf/20/4/4/1969642/jcws_a_00838.pdf.

Terrill:2005:MRR

[Ter05] Ross Terrill. Mao: a Reinterpretation (review). Journal of Cold War Studies, 7(1):211–213, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/179572;   http://muse.jhu.edu/article/179572/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/1/211/700498/
jcws.2005.7.1.211.pdf.

Tuma:2023:REC

[TGS^{+}23] Oldřich Tůma, Venelin I. Ganev, Marius Stan, Piotr Kosicki, and Kyrill Kunakhovich. Reassessing the end of the Cold War and its implications. Journal of Cold War Studies, 25(4):222–248, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/4/
222/2205006/jcws_c_01175.pdf.

Thayer:2006:KBN

[Tha06] Carlyle A. Thayer. Kim N. B. Ninh, A World Transformed: The Politics of Culture in Revolutionary Vietnam, 1945–1965. Journal of Cold War Studies, 8(2):177–179, Spring 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/196296;   http://muse.jhu.edu/article/196296/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/2/177/700564/
jcws.2006.8.2.177.pdf.

Theoharis:2015:FBE

[The15] Athan Theoharis. F. B. Eyes: How J. Edgar Hoover’s Ghostreaders Framed African American Literature by William J. Maxwell (review). Journal of Cold War Studies, 17(3):236–237, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595082;   http:
//muse.jhu.edu/article/595082/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/236/1102460/jcws_r_00571.pdf.

Thies:2014:LKN

[Thi14] Wallace J. Thies. Lawrence A. Kaplan, NATO before the Korean War: April 1949–June 1950. Kent, OH: Kent State University Press, 2013. 216 pp. $60.00. Journal of Cold War Studies, 16(2):128, Spring 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/547327;   http://muse.jhu.edu/article/547327/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/2/128/698432/
jcws_r_00454.pdf.

Thomas:2005:HRI

[Tho05] Daniel C. Thomas. Human rights ideas, the demise of Communism, and the end of the Cold War. Journal of Cold War Studies, 7(2): 110–141, Spring 2005. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/181904;
http://muse.jhu.edu/article/181904/pdf;   https://direct.
mit.edu/jcws/article-pdf/7/2/110/696214/1520397053630600.
pdf.

Thompson:2008:SCR

[Tho08] Ewa Thompson. Simon Cosgrove, Russian Nationalism and the Politics of Soviet Literature: The Case of Nash sovremennik, 1981–91. Journal of Cold War Studies, 10(1):149–151, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/231595;   http:
//muse.jhu.edu/article/231595/pdf;   https://direct.mit.
edu/jcws/article-pdf/10/1/149/700096/jcws.2008.10.1.149.
pdf.

Thorpe:2011:GHH

[Tho11] Andrew Thorpe. Geraint Hughes, Harold Wilson’s Cold War: The Labour Government and East-West Relations, 1964–1970. London: Royal Historical Society Studies in History, 2009. 202 pp. £50.00. Journal of Cold War Studies, 13(3):238–240, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/447486;   http:

//muse.jhu.edu/article/447486/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/3/238/700732/jcws_r_00119.pdf.

Thompson:2021:NDU

[Tho21] Sue Thompson. The Nixon doctrine and U.S. policy on regional cooperation in Southeast Asia after the Second World War. Journal of Cold War Studies, 23(1):126–162, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/1/126/1911804/jcws_a_00965.pdf;
https://muse.jhu.edu/article/788382.

Tierney:2007:PHR

[Tie07] Dominic Tierney. “Pearl Harbor in reverse”: Moral analogies in the Cuban Missile Crisis. Journal of Cold War Studies, 9(3):49– 77, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/219469;
http://muse.jhu.edu/article/219469/pdf;   https://direct.
mit.edu/jcws/article-pdf/9/3/49/696680/jcws.2007.9.3.49.
pdf.

Tismaneanu:2015:UEC

[Tis15] Vladimir Tismaneanu. An Uncanny Era: Conversations between Václav Havel and Adam Michnik transedited by Elzbieta Matynia (review). Journal of Cold War Studies, 17(3):289–291, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595109;   http:
//muse.jhu.edu/article/595109/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/289/1102458/jcws_r_00555.pdf.

Tuminez:2021:FIA

[TJ21] Astrid S. Tuminez and Andrew Paull Jensen. Freedom Incorporated: Anticommunism and Philippine Independence in the Age of Decolonization by Colleen Woods (review). Journal of Cold War Studies, 23(3):252–255, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/3/252/1955769/jcws_r_01016.pdf;
https://muse.jhu.edu/article/801918.

Lin:2016:SGC

[tL16] Hsiao ting Lin. Stephen G. Craft, American Justice in Taiwan: The 1957 Riots and Cold War Foreign Policy. Lexington, KY: University Press of Kentucky, 2016. 267 pp. $45.00. Journal of Cold War Studies, 18(4):219–220, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645918;   http://muse.jhu.edu/article/645918/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/219/700067/
jcws_r_00690.pdf.

Todd:2017:PTA

[Tod17] Molly Todd. The paradox of trans-American solidarity: Gender, race, and representation in the Guatemalan refugee camps of Mexico, 1980–1990. Journal of Cold War Studies, 19(4):74– 112, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/683666;
http://muse.jhu.edu/article/683666/pdf;   https://direct.
mit.edu/jcws/article-pdf/19/4/74/699362/jcws_a_00765.pdf.

Toft:2005:WRR

[Tof05] Monica Duffy Toft. War and Reconciliation: Reason and Emotion in Conflict Resolution (review). Journal of Cold War Studies, 7(3):160–162, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185119;   http://muse.jhu.edu/article/185119/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/160/1102547/
jcws.2005.7.3.160.pdf.

Tolnay:2003:SRH

[Tol03] Adam Tolnay. Seeing Red: Hungarian Intellectuals in Exile and the Challenge of Communism (review). Journal of Cold War Studies, 5 (3):145–147, Summer 2003. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic).  URL http://muse.jhu.edu/article/
45214;   http://muse.jhu.edu/article/45214/pdf;   https:/
/direct.mit.edu/jcws/article-pdf/5/3/145/695925/jcws.2003.
5.3.145.pdf.

Tolvaisas:2010:CWB

[Tol10] Tomas Tolvaisas. Cold war “Bridge–Building”: U.S. exchange exhibits and their reception in the Soviet Union, 1959–1967. Journal of Cold War Studies, 12(4):3–31, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/403795;   http://muse.jhu.edu/article/403795/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/4/3/697536/jcws_
a_00068.pdf.

Tompson:2004:KMH

[Tom04] William J. Tompson. Khrushchev: The Man and His Era (review). Journal of Cold War Studies, 6(4):172–174, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/174725;   http:
//muse.jhu.edu/article/174725/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/4/172/696119/jcws.2004.6.4.172.pdf.

Torigian:2022:YDK

[Tor22] Joseph Torigian. “You don’t know Khrushchev well”: The ouster of the Soviet leader as a challenge to recent scholarship on authoritarian politics. Journal of Cold War Studies, 24(1):78–115, Winter 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
24/1/78/1980877/jcws_a_01043.pdf;   https://muse.jhu.edu/
article/843078.

Trachtenberg:2000:GME

[Tra00] Marc Trachtenberg. De Gaulle, Moravcsik, and Europe. Journal of Cold War Studies, 2(3):101–116, Fall 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9165;   http://muse.jhu.edu/article/9165/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/3/101/695284/
15203970051032255.pdf. See [Mor00b, Mor00c].

Trachtenberg:2002:KWB

[Tra02] Marc Trachtenberg. Kennedy’s Wars: Berlin, Cuba, Laos, and Vietnam (review). Journal of Cold War Studies, 4(2):143–145, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9265;   http:/
/muse.jhu.edu/article/9265/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/2/143/695603/jcws.2002.4.2.143.pdf.

Trachtenberg:2005:MPT

[Tra05] Marc Trachtenberg. The Marshall Plan as tragedy. Journal of Cold War Studies, 7(1):135–140, Winter 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/179573;   http://muse.jhu.edu/article/179573/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/1/135/700472/
1520397053326220.pdf.

Trachtenberg:2008:USE

[Tra08] Marc Trachtenberg. The United States and Eastern Europe in 1945: a reassessment. Journal of Cold War Studies, 10(4):94–132, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/252651;   http:
//muse.jhu.edu/article/252651/pdf;   https://direct.mit.
edu/jcws/article-pdf/10/4/94/697073/jcws.2008.10.4.94.pdf.

Trapeznik:2009:AMD

[Tra09] Alexander Trapeznik. “Agents of Moscow” at the dawn of the Cold War: The Comintern and the Communist Party of New Zealand. Journal of Cold War Studies, 11(1):124–149, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/259187;   http:
//muse.jhu.edu/article/259187/pdf;   https://direct.mit.
edu/jcws/article-pdf/11/1/124/1468522/jcws.2009.11.1.124.
pdf.

Trachtenberg:2011:FFU

[Tra11] Marc Trachtenberg. The French factor in U.S. foreign policy during the Nixon–Pompidou period, 1969–1974. Journal of Cold War Studies, 13(1):4–59, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/424240;   http://muse.jhu.edu/article/424240/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/1/4/697717/jcws_
a_00073.pdf.

Trachtenberg:2012:GP

[Tra12] Marc Trachtenberg. The de Gaulle problem. Journal of Cold War Studies, 14(1):81–92, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467681;   http://muse.jhu.edu/article/467681/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/81/697819/
jcws_a_00194.pdf.

Trachtenberg:2022:USS

[Tra22] Marc Trachtenberg. The United States and strategic arms limitation during the Nixon–Kissinger period: Building a stable international system? Journal of Cold War Studies, 24(4):157–197, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/4/
157/2062823/jcws_a_01104.pdf.

Tromly:2016:MMN

[Tro16a] Benjamin Tromly. The making of a myth: The National Labor Alliance, Russian Émigrés, and Cold War intelligence activities. Journal of Cold War Studies, 18(1):80–111, Winter 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/612007;   http:
//muse.jhu.edu/article/612007/pdf;   https://direct.mit.
edu/jcws/article-pdf/18/1/80/698763/jcws_a_00620.pdf.

Tromly:2016:RHP

[Tro16b] Benjamin Tromly. Robert Hornsby, Protest, Reform and Repression in Khrushchev’s Soviet Union. Cambridge, UK: Cambridge University Press, 2013. 313 pp. $99.00. Journal of Cold War Studies, 18(4):267–268, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645939;   http://muse.jhu.edu/article/645939/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/267/700197/
jcws_r_00714.pdf.

Tsipis:2008:DRR

[Tsi08] Kosta Tsipis. David Rowe and Robert Schulmann, Einstein on Politics. Journal of Cold War Studies, 10(3):140–143, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/241504;
http://muse.jhu.edu/article/241504/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/3/140/697109/jcws.2008.10.3.
140.pdf.

Tumarkin:2022:WSP

[TSW^{+}22] Nina Tumarkin, Ronald Grigor Suny, Elizabeth Wood, Stephen E. Hanson, and Andrea Graziosi. The weight of the Soviet past in post-1991 Russia: Alternative perspectives. Journal of Cold War Studies, 24(2):132–158, Spring 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/24/2/132/2013034/jcws_c_01078.pdf.

Tsypkin:2008:LDN

[Tsy08] Mikhail Tsypkin. Leonid Dubonosov, Nelegal za okeanom [The Illegal Agent Overseas]. Journal of Cold War Studies, 10(1):144– 147, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/231596;

http://muse.jhu.edu/article/231596/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/1/144/700081/jcws.2008.10.1.
144.pdf.

Tudda:2005:RST

[Tud05] Christopher J. Tudda. “Reenacting the story of Tantalus”: Eisenhower, Dulles, and the failed rhetoric of liberation. Journal of Cold War Studies, 7(4):3–35, Fall 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/188280;   http://muse.jhu.edu/article/188280/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/4/3/696238/1520397055012479.
pdf.

Tuminez:2003:NEP

[Tum03] Astrid S. Tuminez. Nationalism, ethnic pressures, and the breakup of the Soviet Union. Journal of Cold War Studies, 5(4):81–136, Fall 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/47932;   http://muse.jhu.edu/
article/47932/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/4/81/700376/152039703322483765.pdf.

Tumarkin:2014:EMC

[Tum14] Nina Tumarkin. A European Memory? Contested Histories and Politics of Remembrance edited by Małgorzata Pakier and Bo Strath (review). Journal of Cold War Studies, 16(3):213–215, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/556082;   http:
//muse.jhu.edu/article/556082/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/3/213/698515/jcws_r_00458.pdf.

Tumarkin:2016:KSS

[Tum16] Nina Tumarkin. Konstantin Sheiko and Stephen Brown, History as Therapy: Alternative History and Nationalist Imaginings in Russia, 1991–2014. Stuttgart: Ibidem Verlag, 2014. 232 pp. Journal of Cold War Studies, 18(2):205–207, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/620552;   http://muse.jhu.edu/article/620552/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/2/205/698926/
jcws_r_00634.pdf.

Turner:2001:WVI

[Tur01] Robert F. Turner. Why Vietnam Invaded Cambodia: Political Culture and the Causes of War (review). Journal of Cold War Studies, 3(3):

117–120, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9215;   http:/
/muse.jhu.edu/article/9215/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/117/695476/jcws.2001.3.3.117.pdf.

Tannenwald:2005:IRI

[TW05] Nina Tannenwald and William Curti Wohlforth. Introduction: The role of ideas and the end of the Cold War. Journal of Cold War Studies, 7(2):3–12, Spring 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/181902;   http://muse.jhu.edu/article/181902/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/2/3/696204/1520397053630574.
pdf.

Uelzmann:2020:BDS

[Uel20] Jan Uelzmann. Building domestic support for West Germany’s integration into NATO, 1953–1955. Journal of Cold War Studies, 22(2):133–162, Spring 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/2/133/1860546/jcws_a_00941.pdf;
https://muse.jhu.edu/article/756827.

Ulam:1999:FUM

[Ula99] Adam Bruno Ulam. A few unresolved mysteries about Stalin and the Cold War in Europe: a modest agenda for research. Journal of Cold War Studies, 1(1):110–116, Winter 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9096;   https://direct.mit.edu/jcws/
article-pdf/1/1/110/695109/15203970152521917.pdf.

Ulianova:2021:CCD

[US21] Olga Ulianova and Alessandro Santoni. The Chilean Christian Democratic Party, the U.S. government, and European politics during Pinochet’s military regime (1973–1988). Journal of Cold War Studies, 23(1):163–195, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/1/163/1911822/jcws_a_00982.pdf;
https://muse.jhu.edu/article/788383.

Usdin:2009:RRR

[Usd09] Steven T. Usdin. The Rosenberg ring revealed: Industrial-scale conventional and nuclear espionage. Journal of Cold War Studies, 11(3):91–143, Summer 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/268392;   http://muse.jhu.edu/article/268392/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/3/91/697252/
jcws.2009.11.3.91.pdf.

Usdin:2011:AML

[Usd11] Steven Usdin. Andrew Meier, The Lost Spy: an American in Stalin’s Secret Service. New York: W. W. Norton, 2008. 402 pp. $25.95. Journal of Cold War Studies, 13(1):238–240, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/424254;   http:
//muse.jhu.edu/article/424254/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/1/238/697731/jcws_r_00085.pdf.

Usdin:2016:LCE

[Usd16] Steven Usdin. Lori Clune, Executing the Rosenbergs: Death & Diplomacy in a Cold War World. New York: Oxford University Press, 2016. 261 pp. $29.95. Journal of Cold War Studies, 18(4):215–216, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645916;   http://muse.jhu.edu/article/645916/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/215/700079/
jcws_r_00688.pdf.

Usdin:2020:HKM

[Usd20] Steven Usdin. Harvey Klehr, The Millionaire Was a Soviet Mole: The Twisted Life of David Karr. Journal of Cold War Studies, 22(2):163–167, Spring 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/22/2/163/1860560/jcws_r_00942.pdf;
https://muse.jhu.edu/article/756828.

Usdin:2022:FCL

[Usd22] Steven Usdin. The first counterspy: Larry Haas, Bell aircraft, and the FBI’s attempt to capture a Soviet mole by Kay Haas and Walter w. pickut. Journal of Cold War Studies, 24(4):215–217, Fall 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/4/
215/2062872/jcws_r_01106.pdf.

Valiyev:2012:JHS

[VA12] Anar Valiyev and Natavan Aghayeva. Jamil Hasanli, Stalin and the Turkish Crisis of the Cold War, 1945–1953. Lanham, MD: Lexington Books, 2011. 419 pp. $72.00 cloth. Journal of Cold War Studies, 14(3):216–218, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/490191;   http://muse.jhu.edu/article/490191/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/3/216/698075/
jcws_r_00264.pdf.

Vaillant:2005:TSP

[Vai05] Janet G. Vaillant. The Teachers of Stalinism: Policy, Practice and Power in Soviet Schools of the 1930s (review). Journal of Cold War Studies, 7(3):188–190, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185120;   http://muse.jhu.edu/article/185120/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/188/1102591/
jcws.2005.7.3.188.pdf.

Vajda:2020:CBE

[Vaj20] Barnabás Vajda. Csaba Békés, Enyhülés és emancipáció: Magyarország, a szovjet blokk és a nemzetközi politika, 1944–1991. Journal of Cold War Studies, 22(4):233–234, Fall 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/22/4/233/1860630/jcws_r_
00973.pdf;   https://muse.jhu.edu/article/776971.

Valiyev:2008:JHD

[Val08] Anar Valiyev. Jamil Hasanli, At the Dawn of the Cold War: The Soviet-American Crisis over Iranian Azerbaijan, 1941–1946. Journal of Cold War Studies, 10(3):174–176, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.

jhu.edu/article/241518;   http://muse.jhu.edu/article/241518/
pdf;   https://direct.mit.edu/jcws/article-pdf/10/3/174/
697127/jcws.2008.10.3.174.pdf.

Valkov:2014:RNS

[Val14] Nikolay Valkov. Rusya–NATO i sredata na sigurnost sled Studenata voina: Chast 1 — 1989–1999 by Nadia Boyadjieva (review). Journal of Cold War Studies, 16(4):283–285, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/566686;   http://muse.jhu.edu/article/566686/
pdf;   https://direct.mit.edu/jcws/article-pdf/16/4/283/
698653/jcws_r_00505.pdf.

Vamos:2010:SSS

[Vám10] Péter Vámos. The Sino–Soviet Split: Cold War in the Communist World (review). Journal of Cold War Studies, 12(1):133–143, Winter 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/382489;   http://
muse.jhu.edu/article/382489/pdf.

Vamos:2022:FAS

[Vám22] Péter Vámos. Friendly assistance and self-reliance: The Hungarian Geophysical Expedition in China, 1956–1962. Journal of Cold War Studies, 24(3):116–150, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/24/3/116/2042009/jcws_a_01075.pdf.

VanOudenaren:1999:FGS

[Van99a] John Van Oudenaren. Francesca Gori and Silvio Pons, eds., The Soviet Union and Europe in the Cold War, 1943–53. Journal of Cold War Studies, 1(3):180–182, Fall 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/9119.

VanOudenaren:1999:SMC

[Van99b] John Van Oudenaren. Sarah E. Mendelson, Changing Course: Ideas, Politics, and the Soviet Withdrawal from Afghanistan. Journal of Cold War Studies, 1(3):180–182, Fall 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/1/3/180/695170/jcws.1999.1.3.180.pdf.

Vanke:2000:RG

[Van00] Jeffrey Vanke. Reconstructing De Gaulle. Journal of Cold War Studies, 2(3):87–100, Fall 2000. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/9163;
 http://muse.jhu.edu/article/9163/pdf;   https://direct.
mit.edu/jcws/article-pdf/2/3/87/695282/15203970051032246.
pdf.

VanOudenaren:2002:CWR

[Van02a] John Van Oudenaren. Cold War Respite: The Geneva Summit of 1955 (review). Journal of Cold War Studies, 4(3):144–146, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9287;   http:/
/muse.jhu.edu/article/9287/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/144/700242/jcws.2002.4.3.144.pdf.

Vanke:2002:PYR

[Van02b] Jeffrey Vanke. The Pompidou Years, 1969–1974 (review). Journal of Cold War Studies, 4(1):113–115, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9240;   http://muse.jhu.edu/article/9240/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/1/113/695518/
jcws.2002.4.1.113.pdf.

Vanke:2004:GMD

[Van04] Jeffrey Vanke. Georges Marchais and the decline of French Communism. Journal of Cold War Studies, 6(1):90–94, Winter 2004. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/51870;   http://muse.jhu.edu/
article/51870/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/1/90/695880/152039704772741605.pdf.

VanAlstein:2011:EEP

[Van11] Maarten Van Alstein. From enigma to enemy: Paul-Henri Spaak, the Belgian diplomatic elite, and the Soviet Union, 1944–1945. Journal of Cold War Studies, 13(3):126–148, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/447461;   http://muse.jhu.edu/article/447461/
pdf;   https://direct.mit.edu/jcws/article-pdf/13/3/126/
700710/jcws_a_00144.pdf.

Varallyay:2009:RGG

[Vár09] Julius Várallyay. Roger Gough, A Good Comrade: János Kádár, Communism and Hungary. Journal of Cold War Studies, 11(1):167– 169, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/259196;
http://muse.jhu.edu/article/259196/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/1/167/1468532/jcws.2009.11.1.
167.pdf.

Vaughan:2015:CWO

[Vau15] Rachel J. Vaughan. The Cold War and the 1984 Olympic Games: a Soviet–American Surrogate War by Philip A. D’Agati (review). Journal of Cold War Studies, 17(3):271–273, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595100;   http:
//muse.jhu.edu/article/595100/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/271/1102426/jcws_r_00512.pdf.

vanCreveld:2013:JFP

[vC13] Martin van Creveld. John France, Perilous Glory: The Rise of Western Military Power. New Haven: Yale University Press, 2011. 438 pp. Journal of Cold War Studies, 15(2):134–138, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/509707;   http:
//muse.jhu.edu/article/509707/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/2/134/698147/jcws_r_00341.pdf.

vanDijk:2002:BEF

[vD02] Ruud van Dijk. Bereitschaft zu Einheit in Freiheit? Die sowjetische Deutschland-Politik 1945–1955 (review). Journal of Cold War Studies, 4(2):120–125, Spring 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9246;   http://muse.jhu.edu/article/9246/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/2/120/695570/
jcws.2002.4.2.120.pdf.

vanDijk:2003:WUD

[vD03] Ruud van Dijk. Walter Ulbricht: Eine deutsche Biografie (review). Journal of Cold War Studies, 5(3):147–150, Summer 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45215;   http://muse.jhu.edu/
article/45215/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/147/695933/jcws.2003.5.3.147.pdf.

vanDijk:2005:SNM

[vD05] Ruud van Dijk. Die Stalin-Note vom 10. Marz 1952: Neue Quellen und Analysen (review). Journal of Cold War Studies, 7(3):180–182, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185098;   http://muse.jhu.edu/article/185098/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/180/1102587/
jcws.2005.7.3.180.pdf.

vanDijk:2012:MDN

[vD12] Ruud van Dijk. Mike Dennis and Norman LaPorte, State and Minorities in Communist East Germany. New York: Berghahn Books, 2011. xvii + 236 pp. Journal of Cold War Studies, 14(4):244–246, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/500345;   http:
//muse.jhu.edu/article/500345/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/4/244/698159/jcws_r_00292.pdf.

vanDijk:2021:INR

[vD21] Ruud van Dijk. Interrogation Nation: Refugees and Spies in Cold War Germany by Keith R. Allen, Lanham, MD: Rowman & Littlefield, 2017. xxxii + 276 pp. (review). Journal of Cold War Studies, 23(1):246–248, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/1/246/1911798/jcws_r_00995.pdf;
https://muse.jhu.edu/article/788385.

vanDijk:2007:PRP

[vDG07] Ruud van Dijk and Peter E. Grieder. Perspectives on resistance with the people. Journal of Cold War Studies, 9(3): 144–154, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219470;   http://muse.jhu.edu/article/219470/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/144/696638/
jcws.2007.9.3.144.pdf.

vanDijk:2008:FPW

[vDO08] Ruud van Dijk and Arnold A. Offner. Forum: Perspectives on [Wilson D. Miscamble, C.S.C.,] From Roosevelt to Truman: Potsdam, Hiroshima, and the Cold War. Journal of Cold War Studies, 10(4):133–141, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/252653;   http://muse.jhu.edu/article/252653/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/4/133/697039/
jcws.2008.10.4.133.pdf.

Veith:2013:TPM

[Vei13] George J. Veith. Thomas P. McKenna, Kontum: The Battle to Save South Vietnam. Lexington, KY: University Press of Kentucky, 2011. 376 pp. Journal of Cold War Studies, 15(1):147–149, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/505768;   http:
//muse.jhu.edu/article/505768/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/1/147/699980/jcws_r_00311.pdf.

Veith:2014:HRV

[Vei14a] George J. Veith. Hanoi’s Road to the Vietnam War, 1954– 1965 by Pierre Asselin (review). Journal of Cold War Studies, 16(3):221–223, Summer 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/556086;   http://muse.jhu.edu/article/556086/pdf;

 https://direct.mit.edu/jcws/article-pdf/16/3/221/698527/
jcws_r_00462.pdf.

Veith:2014:PWM

[Vei14b] George J. Veith. The Pro-War Movement: Domestic Support for the Vietnam War and the Making of Modern American Conservatism by Sandra Scanlon (review). Journal of Cold War Studies, 16(4):273–275, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566681;   http://muse.jhu.edu/article/566681/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/273/698668/
jcws_r_00500.pdf.

Verovsek:2018:SME

[Ver18] Peter J. Verovšek. Screening migrants in the Early Cold War: The geopolitics of U.S. immigration policy. Journal of Cold War Studies, 20(4):154–179, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/718202;
https://direct.mit.edu/jcws/article-pdf/20/4/154/1969654/
jcws_a_00841.pdf.

vonHagen:2009:WLN

[vH09] Mark L. von Hagen. Wendy Lower, Nazi Empire-Building and the Holocaust in Ukraine. Journal of Cold War Studies, 11(1):177– 178, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/259200;
http://muse.jhu.edu/article/259200/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/1/177/1468540/jcws.2009.11.1.
177.pdf.

Vickers:2008:HWB

[Vic08] Rhiannon Vickers. Harold Wilson, the British labour party, and the war in Vietnam. Journal of Cold War Studies, 10 (2):41–70, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/237151;   http://muse.jhu.edu/article/237151/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/2/41/697011/
jcws.2008.10.2.41.pdf.

Villaume:2016:HGB

[Vil16] Poul Villaume. Henrik G. Bastiansen and Rolf Werenskjold, eds., The Nordic Media and the Cold War. Gothenburg: NORDI-COM, 2015. 366 pp. EUR 32.00. Journal of Cold War Studies, 18(3):200–201, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632289;   http://muse.jhu.edu/article/632289/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/3/200/699041/
jcws_r_00663.pdf.

Vitalis:2004:EAF

[Vit04] Robert Vitalis. Egypt and American Foreign Assistance 1952– 1956: Hopes Dashed (review). Journal of Cold War Studies, 6(3):148–150, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/171176;   http://muse.jhu.edu/article/171176/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/3/148/696081/
jcws.2004.6.3.148.pdf.

VonBencke:2003:SCO

[Von03] Matthew J. Von Bencke. Star-Crossed Orbits: Inside the U.S.– Russian Space Alliance (review). Journal of Cold War Studies, 5(3): 137–139, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/45182;
 http://muse.jhu.edu/article/45182/pdf;   https://direct.
mit.edu/jcws/article-pdf/5/3/137/695923/jcws.2003.5.3.137.
pdf.

Vu:2019:SWR

[Vu19] Tuong Vu. In the service of world revolution: Vietnamese Communists’ radical ambitions through the three Indochina wars. Journal of Cold War Studies, 21(4):4–30, Fall 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/21/4/4/699691/jcws_a_00905.pdf;   https:
//muse.jhu.edu/article/736849.

Vujacic:2002:YHT

[Vuj02] Veljko Vujačić. Yugoslavia as History: Twice There Was a Country, 2nd edition, New York: Cambridge University Press, 2000. 487 pp. $24.95 (review). Journal of Cold War Studies, 4(1):120–122, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9241;   http:/
/muse.jhu.edu/article/9241/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/1/120/695573/jcws.2002.4.1.120.pdf.

vanWyk:2019:ABR

[vW19] Anna-Mart van Wyk. Apartheid’s bomb and regional liberation: Cold War perspectives. Journal of Cold War Studies, 21(1):151–165, Winter 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/723556;   https:
//direct.mit.edu/jcws/article-pdf/21/1/151/699709/jcws_a_
00855.pdf.

Wallander:2003:WPD

[Wal03] Celeste A. Wallander. Western policy and the demise of the Soviet Union. Journal of Cold War Studies, 5(4):137–177, Fall 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/47933;   http://muse.jhu.edu/
article/47933/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/4/137/700382/152039703322483774.pdf.

Wall:2007:CEC

[Wal07] Irwin Wall. Christopher Endy, Cold War Holidays: American Tourism in France. Journal of Cold War Studies, 9 (1):122–124, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210106;   http://muse.jhu.edu/article/210106/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/122/696475/
jcws.2007.9.1.122.pdf.

Wallander:2008:HHG

[Wal08] Celeste A. Wallander. Helga Haftendorn, Georges-Henri Soutou, Stephen F. Szabo, and Samuel F. Wells, Jr., eds., The Strategic Triangle: France, Germany, and the United States in the Shaping of the New Europe. Journal of Cold War Studies, 10(3):198– 199, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/241530;
http://muse.jhu.edu/article/241530/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/3/198/697143/jcws.2008.10.3.
198.pdf.

Walk:2012:DHE

[Wal12a] Robert D. Walk. D. Hank Ellison, Chemical Warfare during the Vietnam War: Riot Control Agents in Combat. New York: Routledge, 2011. 202 pp. Journal of Cold War Studies, 14(4):233–234, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/500340;   http:

//muse.jhu.edu/article/500340/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/4/233/698130/jcws_r_00287.pdf.

Walker:2012:JCF

[Wal12b] Breck Walker. Jonathan Colman, The Foreign Policy of Lyndon B. Johnson: The United States and the World, 1963–1969. Edinburgh: Edinburgh University Press, 2010. 231 pp. $100.00. Journal of Cold War Studies, 14(4):229–230, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500338;   http://muse.jhu.edu/article/500338/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/229/698126/
jcws_r_00285.pdf.

Wall:2012:CGG

[Wal12c] Irwin Wall. Commentary on globalizing de Gaulle. Journal of Cold War Studies, 14(1):48–52, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467678;   http://muse.jhu.edu/article/467678/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/48/697875/
jcws_a_00191.pdf. See [Loc05].

Wallace:2013:CRL

[Wal13a] James C. Wallace. Contained?: The religious life of George F. Kennan and its influence. Journal of Cold War Studies, 15(4):196–215, Fall 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/534417;   http:
//muse.jhu.edu/article/534417/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/4/196/698347/jcws_a_00404.pdf.

Wallace:2013:RWC

[Wal13b] James C. Wallace. A religious war?: The Cold War and religion. Journal of Cold War Studies, 15(3):162–180, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/522713;   http:
//muse.jhu.edu/article/522713/pdf;   https://direct.mit.
edu/jcws/article-pdf/15/3/162/698282/jcws_a_00374.pdf.

Wall:2017:PLR

[Wal17a] Sir Stephen Wall. N. Piers Ludlow, Roy Jenkins and the European Commission Presidency, 1976–1980: At the Heart of Europe. Journal of Cold War Studies, 19(4):281–283, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/683695;   http://muse.jhu.edu/article/683695/

pdf;   https://direct.mit.edu/jcws/article-pdf/19/4/281/
699465/jcws_r_00795.pdf.

Wallace:2017:MRF

[Wal17b] James C. Wallace. Markku Ruotsila, Fighting Fundamentalist: Carl McIntire and the Politicization of American Fundamentalism. Journal of Cold War Studies, 19(4):266–268, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/683688;   http://muse.jhu.edu/article/683688/
pdf;   https://direct.mit.edu/jcws/article-pdf/19/4/266/
699457/jcws_r_00788.pdf.

Walder:2021:JFT

[Wal21a] Andrew G. Walder. June Fourth: The Tiananmen Protests and Beijing Massacre of 1989 by Jeremy Brown (review). Journal of Cold War Studies, 23(4):257–260, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/4/257/1970107/jcws_r_01051.pdf;
https://muse.jhu.edu/article/835888.

Walker:2021:ATM

[Wal21b] J. Samuel Walker. Atomic Testing in Mississippi: Project Dribble and the Quest for Nuclear Weapons Treaty Verification in the Cold War Era by David Allen Burke (review). Journal of Cold War Studies, 23(4):251–252, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/4/251/1970081/jcws_r_01048.pdf;
https://muse.jhu.edu/article/835885.

Wandycz:2009:TSR

[Wan09] Piotr S. Wandycz. Timothy Snyder, The Red Prince: The Secret Lives of a Habsburg Archduke. Journal of Cold War Studies, 11(2):117–119, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/265883;   http://muse.jhu.edu/article/265883/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/2/117/697139/
jcws.2009.11.2.117.pdf.

Wang:2016:YME

[Wan16] Zuoyue Wang. Yanek Mieczkowski, Eisenhower’s Sputnik Moment: The Race for Space and World Prestige. Ithaca, NY: Cornell University Press, 2013. 358 pp. Journal of Cold War Studies, 18(4):257–259, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/645934;   http://muse.jhu.edu/article/645934/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/4/257/700174/
jcws_r_00707.pdf.

Wang:2017:NIG

[Wan17] Tao Wang. Neutralizing Indochina: The 1954 Geneva Conference and China’s efforts to isolate the United States. Journal of Cold War Studies, 19(2):3–42, Spring 2017. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://www.mitpressjournals.org/
doi/full/10.1162/JCWS_a_00739;   https://direct.mit.edu/
jcws/article-pdf/19/2/3/699201/jcws_a_00739.pdf.

Warner:2001:MCM

[War01] Michael Warner. Many Are the Crimes: McCarthyism in America (review). Journal of Cold War Studies, 3(2):101–103, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9195;   http:/
/muse.jhu.edu/article/9195/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/2/101/695382/jcws.2001.3.2.101.pdf.

Ward:2006:DRS

[War06] James Mace Ward. Das Dritte Reich und die Slowakei 1939– 1945: Politischer Alltag zwischen Kooperation und Eigensinn (review). Journal of Cold War Studies, 8(1):132–134, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/192221;   http:
//muse.jhu.edu/article/192221/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/1/132/696291/jcws.2006.8.1.132.pdf.

Ware:2008:LEH

[War08] Susan Ware. Linda Eisenmann, Higher Education for Women in Postwar America, 1945–1965. Journal of Cold War Studies, 10(4):145–147, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/252660;   http://muse.jhu.edu/article/252660/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/4/145/697047/
jcws.2008.10.4.145.pdf.

Watry:2017:MDI

[Wat17] David M. Watry. Michael Doran, Ike’s Gamble: America’s Rise to Dominance in the Middle East. Journal of Cold War Studies, 19(4):270–272, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683690;   http://muse.jhu.edu/article/683690/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/270/699463/
jcws_r_00790.pdf.

Watry:2018:WHA

[Wat18] David M. Watry. William I. Hitchcock, The Age of Eisenhower: America and the World in the 1950s. Journal of Cold War Studies, 20 (4):266–268, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/718223;
https://direct.mit.edu/jcws/article-pdf/20/4/266/1969632/
jcws_r_00849.pdf.

Weathersby:2012:ARM

[Wea12] Kathryn Weathersby. Allan R. Millett, The War for Korea, 1950– 1951: They Came from the North. Lawrence, KS: University Press of Kansas, 2010. 644 pp. $45.00. Journal of Cold War Studies, 14(2):158–160, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/478782;   http://muse.jhu.edu/article/478782/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/2/158/697991/
jcws_r_00227.pdf.

Weathersby:2015:KWS

[Wea15] Kathryn Weathersby. The Korean War at Sixty: New Approaches to the Study of the Korean War edited by Steven Casey (review). Journal of Cold War Studies, 17(3):242–243, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595085;   http:
//muse.jhu.edu/article/595085/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/242/1102486/jcws_r_00574.pdf.

Weeks:2018:FLF

[Wee18] Gregory Weeks. Fernando López, The Feathers of Condor: Transnational State Terrorism, Exiles and Civilian Anticommunism in South America. Journal of Cold War Studies, 20(4):264–266, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718222;   https://direct.mit.
edu/jcws/article-pdf/20/4/264/1969661/jcws_r_00848.pdf.

Weiss:2001:CWU

[Wei01] Erik D. Weiss. Cold war under the ice: The Army’s bid for a longrange nuclear role, 1959–1963. Journal of Cold War Studies, 3(3):

31–58, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9216;   http:/
/muse.jhu.edu/article/9216/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/31/695424/152039701750419501.pdf.

Weiner:2002:APR

[Wei02] Robert Weiner. Ana Pauker: The Rise and Fall of a Jewish Communist (review). Journal of Cold War Studies, 4(4):126–127, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9311;   http:/
/muse.jhu.edu/article/9311/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/4/126/695719/jcws.2002.4.4.126.pdf.

Weiner:2004:IRR

[Wei04a] Robert Weiner. Israeli–Romanian Relations at the End of the Ceausescu Era: As Observed by Israel’s Ambassador to Romania, 1985–89 (review). Journal of Cold War Studies, 6(2):102–104, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/53683;   http://muse.jhu.edu/
article/53683/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/2/102/695959/jcws.2004.6.2.102.pdf.

Weinthal:2004:ISU

[Wei04b] Erika Weinthal. Islam and the Soviet Union: From the Second World War to Gorbachev (review). Journal of Cold War Studies, 6(2):104– 105, Spring 2004. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/53684;
 http://muse.jhu.edu/article/53684/pdf;   https://direct.
mit.edu/jcws/article-pdf/6/2/104/695947/jcws.2004.6.2.104.
pdf.

Weiss:2012:RJC

[Wei12] Andrea Weiss. Robert J. Corber, Cold War Femme: Lesbianism, National Identity, and Hollywood Cinema. Durham, NC: Duke University Press, 2011. 225 pp. $23.95. Journal of Cold War Studies, 14(3):196–197, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/490181;   http://muse.jhu.edu/article/490181/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/3/196/698057/
jcws_r_00254.pdf.

Welch:2000:PTI

[Wel00] David A. Welch. Political Theory and International Relations (review). Journal of Cold War Studies, 2(2):113–115, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9150;   http:/
/muse.jhu.edu/article/9150/pdf;   https://direct.mit.edu/
jcws/article-pdf/2/2/113/695261/jwcs.2000.2.2.113.pdf.

Welch:2006:AAH

[Wel06] David A. Welch. Awaiting Armageddon: How Americans Faced the Cuban Missile Crisis (review). Journal of Cold War Studies, 8(1):146–147, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/192222;   http://muse.jhu.edu/article/192222/pdf;
 https://direct.mit.edu/jcws/article-pdf/8/1/146/696353/
jcws.2006.8.1.146.pdf.

Wemheuer:2021:ADI

[Wem21] Felix Wemheuer. Agents of Disorder: Inside China’s Cultural Revolution by Andrew Walder, and: A Decade of Upheaval: The Cultural Revolution in Rural China by Dong Guoqiang and Andrew Walder (review). Journal of Cold War Studies, 23(4):260–263, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/23/4/260/1970093/
jcws_r_01052.pdf;   https://muse.jhu.edu/article/835889.

Wenger:2004:CON

[Wen04] Andreas Wenger. Crisis and opportunity: NATO’s transformation and the multilateralization of detente, 1966–1968. Journal of Cold War Studies, 6(1):22–74, Winter 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/51871;   http://muse.jhu.edu/
article/51871/pdf;   https://direct.mit.edu/jcws/article-
pdf/6/1/22/695873/152039704772741588.pdf.

Wend:2009:AGR

[Wen09] Henry Burke Wend. Armin Grünbacher, Reconstruction and Cold War in Germany: The Kreditanstalt für Wiederaufbau (1948– 1961); David Monod, Settling Scores: German Music, Denazification, and the Americans, 1945–1953. Journal of Cold War Studies, 11(2):126–129, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/265879;   http://muse.jhu.edu/article/265879/pdf;

 https://direct.mit.edu/jcws/article-pdf/11/2/126/697149/
jcws.2009.11.2.126.pdf.

Westad:2000:IPK

[Wes00] Odd Arne Westad. Irresolute Princes: Kremlin Decision Making in Middle East Crises, 1967–1973 (review). Journal of Cold War Studies, 2(1):127–128, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9134;   http://muse.jhu.edu/article/9134/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/1/127/695195/
jcws.2000.2.1.127.pdf.

Wettig:2001:ADE

[Wet01] Gerhard Wettig. American Diplomacy and the End of the Cold War: an Insider’s Account of U.S. Policy in Europe, 1989– 1992 (review). Journal of Cold War Studies, 3(3):102–104, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9217;   http:/
/muse.jhu.edu/article/9217/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/102/695456/jcws.2001.3.3.102.pdf.

Wettig:2009:DVM

[Wet09] Gerhard Wettig. Damian van Melis and Henrik Bispinck, eds., “Republikflucht”: Flucht und Abwanderung aus der SBZ/DDR 1945 bis 1961 (German) [“Flight from the Republic”: Flight and Emigration from the Soviet Zone/German Democratic Republic, 1945–1961]. Journal of Cold War Studies, 11(2):130–131, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/265878;
http://muse.jhu.edu/article/265878/pdf;   https://direct.
mit.edu/jcws/article-pdf/11/2/130/697165/jcws.2009.11.2.
130.pdf.

Wettig:2010:CBP

[Wet10] Gerhard Wettig. Chad Bryant, Prague in Black. Nazi Rule and Czech Nationalism. Journal of Cold War Studies, 12(3):142–144, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/392788;   http:
//muse.jhu.edu/article/392788/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/3/142/697614/jcws_r_00015.pdf.

Wettig:2011:TDWa

[Wet11a] Gerhard Wettig. Torsten Diedrich, Winfried Heinemann, and Christian F. Ostermann, eds., Der Warschauer Pakt: Von der Gründung bis zum Zusammenbruch 1955 bis 1991. Berlin: Ch. Links Verlag 2009. 368 pp. Journal of Cold War Studies, 13(1):254–256, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/424262;   http:
//muse.jhu.edu/article/424262/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/1/254/697750/jcws_r_00093.pdf.

Wettig:2011:TDWb

[Wet11b] Gerhard Wettig. Torsten Diedrich, Winfried Heinemann, and Christian F. Ostermann, eds., Der Warschauer Pakt: Von der Gründung bis zum Zusammenbruch 1955 bis 1991. Berlin: Ch. Links Verlag, 2009. 368 pp. Journal of Cold War Studies, 13(4):253–255, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/457217;   http:
//muse.jhu.edu/article/457217/pdf;   https://direct.mit.
edu/jcws/article-pdf/13/4/253/697798/jcws_r_00129.pdf.

Wettig:2017:ASD

[Wet17] Gerhard Wettig. Arvid Schors, Doppelter Boden: Die SALT-Verhandlungen 1963–1979. Journal of Cold War Studies, 19(4):237– 239, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/683674;   http:
//muse.jhu.edu/article/683674/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/4/237/699381/jcws_r_00774.pdf.

Wegener-Friis:2012:GBF

[WF12] Thomas Wegener-Friis. Gary Bruce, The Firm: The Inside Story of the Stasi. New York: Oxford University Press, 2010. 239 pp. Journal of Cold War Studies, 14(4):242–244, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500344;   http://muse.jhu.edu/article/500344/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/242/698145/
jcws_r_00291.pdf.

Whelan:2006:IMP

[Whe06] Bernadette Whelan. Ireland, the Marshall Plan, and U.S. Cold War concerns. Journal of Cold War Studies, 8(1):68–94, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/192223;   http:

//muse.jhu.edu/article/192223/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/1/68/696287/152039706775212076.pdf.

Whitfield:2002:BCC

[Whi02a] Stephen J. Whitfield. British Cinema and the Cold War: The State, Propaganda and Consensus (review). Journal of Cold War Studies, 4(4):120–122, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9310;   http://muse.jhu.edu/article/9310/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/4/120/695728/
jcws.2002.4.4.120.pdf.

Whitfield:2002:MCW

[Whi02b] Stephen J. Whitfield. The Making of the Cold War Enemy: Culture and Politics in the Military–Intellectual Complex (review). Journal of Cold War Studies, 4(3):127–129, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9292;   http://muse.jhu.edu/article/9292/pdf;
 https://direct.mit.edu/jcws/article-pdf/4/3/127/700238/
jcws.2002.4.3.127.pdf.

Whitfield:2003:PAC

[Whi03a] Stephen J. Whitfield. The Politics of Apolitical Culture: The Congress for Cultural Freedom, the CIA and Post-War American Hegemony (review). Journal of Cold War Studies, 5(2):113–115, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40777;   http://muse.jhu.edu/
article/40777/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/113/700398/jcws.2003.5.2.113.pdf.

Whiting:2003:WLI

[Whi03b] Allen Suess Whiting. The Wizards of Langley: Inside the CIA’s Directorate of Science and Technology (review). Journal of Cold War Studies, 5(2):111–113, Spring 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40778;   http://muse.jhu.edu/
article/40778/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/111/700386/jcws.2003.5.2.111.pdf.

Whitfield:2004:DDA

[Whi04] Stephen J. Whitfield. Dictators, Democracy, and American Public Culture: Envisioning the Totalitarian Enemy, 1920s-1950s (review). Journal of Cold War Studies, 6(4):157–159, Fall 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/174726;   http:
//muse.jhu.edu/article/174726/pdf;   https://direct.mit.
edu/jcws/article-pdf/6/4/157/696089/jcws.2004.6.4.157.pdf.

Whitfield:2006:RCW

[Whi06] Stephen J. Whitfield. Religion and the Cold War (review). Journal of Cold War Studies, 8(1):138–140, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/192224;   http://muse.jhu.edu/article/192224/
pdf;   https://direct.mit.edu/jcws/article-pdf/8/1/138/696309/
jcws.2006.8.1.138.pdf.

White:2008:C

[Whi08a] Lynn White. Commentary. Journal of Cold War Studies, 10(2): 98–102, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/237153;
http://muse.jhu.edu/article/237153/pdf.

Whitfield:2008:TSH

[Whi08b] Stephen J. Whitfield. Tony Shaw, Hollywood’s Cold War. Journal of Cold War Studies, 10(3):145–146, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/241506;   http://muse.jhu.edu/article/241506/
pdf;   https://direct.mit.edu/jcws/article-pdf/10/3/145/
697091/jcws.2008.10.3.145.pdf.

Whitfield:2010:MKC

[Whi10] Stephen J. Whitfield. Michael Kimmage, The Conservative Turn: Lionel Trilling, Whittaker Chambers, and the Lessons of Anti-Communism. Journal of Cold War Studies, 12(4):191–193, Fall 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/403806;   http:
//muse.jhu.edu/article/403806/pdf;   https://direct.mit.
edu/jcws/article-pdf/12/4/191/697552/jcws_r_00058.pdf.

Whitfield:2011:AJF

[Whi11a] Stephen J. Whitfield. Andrew J. Falk, Upstaging the Cold War: American Dissent and Cultural Diplomacy, 1940–1960. Amherst: University of Massachusetts Press, 2010. 264 pp. $34.95. Journal of Cold War Studies, 13(4):233–235, Fall 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/457208;   http://muse.jhu.edu/article/457208/pdf;

 https://direct.mit.edu/jcws/article-pdf/13/4/233/697811/
jcws_r_00151.pdf.

Whitfield:2011:AWJ

[Whi11b] Stephen J. Whitfield. Aviva Weingarten, Jewish Organisations’ Response to Communism and to Senator McCarthy. London: Vallentine Mitchell, 2008. 164 pp. $75.00 cloth, $35.00 paper. Journal of Cold War Studies, 13(3):223–225, Summer 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/447478;   http://muse.jhu.edu/article/447478/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/3/223/700688/
jcws_r_00111.pdf.

Whitfield:2012:RGL

[Whi12] Stephen J. Whitfield. Robert Genter, Late Modernism: Art, Culture, and Politics in Cold War America. Philadelphia: University of Pennsylvania Press, 2010. 375 pp. $49.95. Journal of Cold War Studies, 14(4):225–227, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500336;   http://muse.jhu.edu/article/500336/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/225/698124/
jcws_r_00283.pdf.

Whitfield:2013:JSJ

[Whi13] Stephen J. Whitfield. John Sbardellati, J. Edgar Hoover Goes to the Movies: The FBI and the Origins of Hollywood’s Cold War. Ithaca, NY: Cornell University Press, 2012. 256 pp. $27.95. Oliver Boyd-Barrett, David Herrera, and Jim Baumann, Hollywood and the CIA: Cinema, Defense, and Subversion. London: Routledge, 2011. 208 pp. $138.00. Journal of Cold War Studies, 15(1):159–162, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/505762;   http://muse.jhu.edu/article/505762/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/1/159/700002/
jcws_r_00317.pdf.

Whitfield:2014:AVM

[Whi14a] Stephen J. Whitfield. Annette Vowinckel, Marcus M. Payk, and Thomas Lindenberger, eds., Cold War Cultures: Perspectives on Eastern and Western Societies New York: Berghahn Books, 2012. 385 pp. $95.00. Journal of Cold War Studies, 16(1):249–251, Winter 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/543749;   http:

//muse.jhu.edu/article/543749/pdf;   https://direct.mit.
edu/jcws/article-pdf/16/1/249/698993/jcws_r_00440.pdf.

Whitfield:2014:HEE

[Whi14b] Stephen J. Whitfield. Hollywood Exiles in Europe: The Blacklist and Cold War Film Culture by Rebecca Prime (review). Journal of Cold War Studies, 16(4):222–224, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566658;   http://muse.jhu.edu/article/566658/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/222/698560/
jcws_r_00478.pdf.

Whitfield:2015:GBC

[Whi15] Stephen J. Whitfield. Greg Barnhisel, Cold War Modernists: Art, Literature, and American Cultural Diplomacy. New York: Columbia University Press, 2015. 322 pp. $40.00. Journal of Cold War Studies, 17(4):210–212, Fall 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/606043;   http://muse.jhu.edu/article/606043/pdf;
 https://direct.mit.edu/jcws/article-pdf/17/4/210/698832/
jcws_r_00611.pdf.

White:2016:PMB

[Whi16] George White, Jr. Philip E. Muehlenbeck, Betting on the Africans: John F. Kennedy’s Courting of African Nationalist Leaders. New York: Oxford University Press, 2012. 333 pp. $55.00. Journal of Cold War Studies, 18(2):195–197, Spring 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/620558;   http://muse.jhu.edu/article/620558/pdf;
 https://direct.mit.edu/jcws/article-pdf/18/2/195/698855/
jcws_r_00629.pdf.

Widmaier:2005:GDP

[Wid05] Wesley W. Widmaier. Gold, Dollars, and Power: The Politics of International Monetary Relations 1958–1971 (review). Journal of Cold War Studies, 7(3):153–155, Summer 2005. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/185121;   http://muse.jhu.edu/article/185121/pdf;
 https://direct.mit.edu/jcws/article-pdf/7/3/153/1102545/
jcws.2005.7.3.153.pdf.

Wiest:2007:HBS

[Wie07] Andrew A. Wiest. Howard B. Schaffer, Ellsworth Bunker: Global Troubleshooter, Vietnam Hawk. Chapel Hill: The University of North Carolina Press, 2003. 380 pp. $34.95. Journal of Cold War Studies, 9(4):176–178, Fall 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/223560;   http://muse.jhu.edu/article/223560/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/4/176/696753/
jcws.2007.9.4.176.pdf.

Williams:2000:RAS

[Wil00] Warren W. Williams. The road to the Austrian state treaty. Journal of Cold War Studies, 2(2):97–107, Spring 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9151;   http://muse.jhu.edu/article/
9151/pdf;   https://direct.mit.edu/jcws/article-pdf/2/2/
97/695222/15203970051032327.pdf.

Williams:2002:GBV

[Wil02] Warren W. Williams. Green Berets in the Vanguard: Inside Special Forces, 1953–1963 (review). Journal of Cold War Studies, 4(4): 110–112, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9312;   http:/
/muse.jhu.edu/article/9312/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/4/110/695686/jcws.2002.4.4.110.pdf.

Williams:2003:BOA

[Wil03] Warren W. Williams. Britain and the Occupation of Austria, 1943– 1945 (review). Journal of Cold War Studies, 5(3):131–132, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/45216;   http://
muse.jhu.edu/article/45216/pdf;   https://direct.mit.edu/
jcws/article-pdf/5/3/131/695921/jcws.2003.5.3.131.pdf.

Williams:2007:FAC

[Wil07] Warren C. Williams. Flashpoint Austria: The Communistinspired strikes of 1950. Journal of Cold War Studies, 9(3): 115–136, Summer 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/219471;   http://muse.jhu.edu/article/219471/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/3/115/696644/
jcws.2007.9.3.115.pdf.

Williams:2008:WMS

[Wil08a] Warren W. Williams. Wolfgang Mueller, Die sowjetische Besatzung in Österreich 1945–1955 und ihre politische Mission. (German) [The Soviet Occupying Forces in Austria, 1945–1955, and Their Political Mission]. Journal of Cold War Studies, 10(3):196– 198, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/241529;
http://muse.jhu.edu/article/241529/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/3/196/697125/jcws.2008.10.3.
196.pdf.

Williams:2008:LRB

[Wil08b] Warren Wellde Williams. Lawrence R. Bailey and Ron Martz, Solitary Survivor: The First American POW in Southeast Asia. Journal of Cold War Studies, 10(4):156–158, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/252663;   http://muse.jhu.edu/article/252663/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/4/156/697049/
jcws.2008.10.4.156.pdf.

Williams:2012:RER

[Wil12] Kieran Williams. Review essay: The Russian view(s) of the Prague Spring. Journal of Cold War Studies, 14(2):128–141, Spring 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/478778;   http:
//muse.jhu.edu/article/478778/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/2/128/697981/jcws_a_00223.pdf.

Williams:2017:MVP

[Wil17] Kieran Williams. Miroslav Vaněk and Pavel Mücke, Velvet Revolutions: an Oral History of Czech Society. Journal of Cold War Studies, 19(4):256–258, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683683;   http://muse.jhu.edu/article/683683/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/256/699429/
jcws_r_00783.pdf.

Williams:2023:WFE

[Wil23] Brad Williams. Why the five eyes? Power and identity in the formation of a multilateral intelligence grouping. Journal of Cold War Studies, 25(1):101–137, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/1/101/2073765/jcws_a_01123.pdf.

Winkler:2007:SCC

[Win07] Allan M. Winkler. Stanley Corkin, Cowboys as Cold Warriors: The Western and U.S. History. Journal of Cold War Studies, 9(1):103–105, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210107;   http://muse.jhu.edu/article/210107/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/103/696505/
jcws.2007.9.1.103.pdf.

Winkler:2011:HLI

[Win11] David F. Winkler. Harold LeeWise, Inside the Danger Zone: The U.S. Military in the Persian Gulf, 1987–1988. Annapolis: Naval Institute Press, 2007. 272 pp. $36.95. Journal of Cold War Studies, 13(1):234–236, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/424252;   http://muse.jhu.edu/article/424252/pdf;
 https://direct.mit.edu/jcws/article-pdf/13/1/234/697703/
jcws_r_00083.pdf.

Winkler:2015:DPM

[Win15] Allan M. Winkler. Daisy Petals and Mushroom Clouds: LBJ, Barry Goldwater, and the Ad That Changed American Politics by Robert Mann (review). Journal of Cold War Studies, 17(3):262–264, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/595095;   http:
//muse.jhu.edu/article/595095/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/3/262/1102494/jcws_r_00593.pdf.

Winger:2017:NDU

[Win17] Gregory Winger. The Nixon Doctrine and U.S. relations with the Republic of Afghanistan, 1973–1978: Stuck in the middle with Daoud. Journal of Cold War Studies, 19(4):4–41, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/683664;   http:
//muse.jhu.edu/article/683664/pdf;   https://direct.mit.
edu/jcws/article-pdf/19/4/4/699372/jcws_a_00763.pdf.

Wirtz:2003:SLW

[Wir03] James J. Wirtz. Strategy: The Logic of War and Peace (review). Journal of Cold War Studies, 5(3):115–116, Summer 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/45217;   http://muse.jhu.edu/
article/45217/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/3/115/695808/jcws.2003.5.3.115.pdf.

Wirtz:2006:HDL

[Wir06] James J. Wirtz. How Democracies Lose Small Wars (review). Journal of Cold War Studies, 8(1):134–136, Winter 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/192225;   http://muse.jhu.edu/article/192225/
pdf;   https://direct.mit.edu/jcws/article-pdf/8/1/134/696313/
jcws.2006.8.1.134.pdf.

Wirtz:2007:JKV

[Wir07a] James J. Wirtz. Jeffrey Kimball, The Vietnam War Files: Uncovering the Secret History of Nixon-Era Strategy. Journal of Cold War Studies, 9(1):117–118, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210108;   http://muse.jhu.edu/article/210108/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/1/117/696500/
jcws.2007.9.1.117.pdf.

Wirtz:2007:RMC

[Wir07b] James J. Wirtz. Robert M. Citino, Blitzkrieg to Desert Storm: The Evolution of Operational Warfare. Lawrence, KS: University Press of Kansas, 2004. 424 pp. $39.95. Journal of Cold War Studies, 9(2):150–152, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/214757;   http://muse.jhu.edu/article/214757/pdf;
 https://direct.mit.edu/jcws/article-pdf/9/2/150/696431/
jcws.2007.9.2.150.pdf.

Wirtz:2013:TCM

[Wir13] James J. Wirtz. Tony Cash and Mike Gerrard, The Coder Special Archive: The Untold Story of Navy National Servicemen Learning and Using Russian during the Cold War. Kingston-upon-Thames, UK: Hodgson Press, 2012. 438 pp. £12.99. Journal of Cold War Studies, 15(3):189–190, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522718;   http://muse.jhu.edu/article/522718/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/189/698337/
jcws_r_00379.pdf.

Wishnick:2000:BAR

[Wis00] Elizabeth Wishnick. Brothers in Arms: The Rise and Fall of the Sino–Soviet Alliance, 1945–1963 (review). Journal of Cold War Studies, 2(1):128–129, Winter 2000. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/9135;   http://muse.jhu.edu/article/9135/pdf;
 https://direct.mit.edu/jcws/article-pdf/2/1/128/695197/
jcws.2000.2.1.128.pdf.

White:2008:FMC

[WLX^{+}08] Lynn White, Steven I. Levine, Yafeng Xia, Joseph W. Esherick, David E. Apter, and other. Forum: Mao and the cultural revolution in China: Commentaries on Mao’s Last Revolution. Journal of Cold War Studies, 10(2):97–130, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/237153;   http://muse.jhu.edu/article/237153/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/2/97/697015/
jcws.2008.10.2.97.pdf.

Wohlforth:1999:CIS

[Woh99] William Curti Wohlforth. A certain idea of science: How international relations theory avoids reviewing the Cold War. Journal of Cold War Studies, 1(2):39–60, Spring 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9107;   https://direct.mit.edu/jcws/
article-pdf/1/2/39/695133/15203979952559513.pdf.

Wohlforth:2005:ECW

[Woh05] William Curti Wohlforth. The end of the Cold War as a hard case for ideas. Journal of Cold War Studies, 7(2):165–173, Spring 2005. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/181905;
http://muse.jhu.edu/article/181905/pdf;   https://direct.
mit.edu/jcws/article-pdf/7/2/165/696212/1520397053630583.
pdf.

Wolff:2013:JSP

[Wol13] David Wolff. Japan and Stalin’s policy toward Northeast Asia after World War II. Journal of Cold War Studies, 15(2):4–29, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/509701;
http://muse.jhu.edu/article/509701/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/2/4/698151/jcws_a_00335.pdf.

Wong:2004:BPM

[Won04] Roy Bin Wong. Between Politics and Markets: Firms, Competition, and Institutional Change in Post-Mao China (review). Journal of Cold War Studies, 6(3):166–168, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/171177;   http://muse.jhu.edu/article/171177/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/3/166/696127/
jcws.2004.6.3.166.pdf.

Woodward:2007:MHM

[Woo07] Susan L. Woodward. Munevera Hadžišehović, A Muslim Woman in Tito’s Yugoslavia. Journal of Cold War Studies, 9(1):148–150, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/210109;   http:
//muse.jhu.edu/article/210109/pdf;   https://direct.mit.
edu/jcws/article-pdf/9/1/148/696546/jcws.2007.9.1.148.pdf.

Wooden:2012:JAY

[Woo12] Amanda E. Wooden. Julian Agyeman and Yelena Ogneva-Himmelberger, eds., Environmental Justice and Sustainability in the Former Soviet Union. Cambridge, MA: MIT Press, 2009. 296 pp. $25.00. Journal of Cold War Studies, 14(3):225–227, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/490195;   http:
//muse.jhu.edu/article/490195/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/3/225/698080/jcws_r_00268.pdf.

Woodwell:2013:PME

[Woo13] Douglas Woodwell. Philip E. Muehlenbeck, ed., Race, Ethnicity, and the Cold War: a Global Perspective. Nashville: Vanderbilt University Press. 324 pp. $27.95. Journal of Cold War Studies, 15(3):198–199, Summer 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/522723;   http://muse.jhu.edu/article/522723/pdf;
 https://direct.mit.edu/jcws/article-pdf/15/3/198/698294/
jcws_r_00350.pdf.

Wyman:2003:BSU

[Wym03] Francis Wyman. Britain, the Soviet Union and Russia (review). Journal of Cold War Studies, 5(2):119–122, Spring 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/40779;   http://muse.jhu.edu/
article/40779/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/2/119/700430/jcws.2003.5.2.119.pdf.

Wyss:2012:MNT

[Wys12] Marco Wyss. Mikael Nilsson, Tools of Hegemony: Military Technology and Swedish-American Security Relations, 1945–1962. Stockholm: Santérus Academic Press, 2007, 486 pp. 315 Krona. Journal of Cold War Studies, 14(3):208–209, Summer 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/490187;   http://muse.jhu.edu/article/490187/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/3/208/698059/
jcws_r_00260.pdf.

Wyss:2018:CWN

[Wys18] Marco Wyss. The challenge of Western neutralism during the Cold War: Britain and the buildup of a Nigerian Air Force. Journal of Cold War Studies, 20(2):99–128, Spring 2018. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic).

URL http://muse.jhu.edu/article/697062;   https://direct.
mit.edu/jcws/article-pdf/20/2/99/699411/jcws_a_00817.pdf.

Xia:2006:CEP

[Xia06] Yafeng Xia. China’s elite politics and Sino–American rapprochement, January 1969–February 1972. Journal of Cold War Studies, 8(4):3–28, Fall 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/204637;   http:
//muse.jhu.edu/article/204637/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/4/3/700821/jcws.2006.8.4.3.pdf.

Xia:2008:C

[Xia08a] Yafeng Xia. Commentary. Journal of Cold War Studies, 10(2):108– 116, Spring 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/237153;
http://muse.jhu.edu/article/237153/pdf.

Xia:2008:SCW

[Xia08b] Yafeng Xia. The study of Cold War international history in China: a review of the last twenty years. Journal of Cold War Studies, 10(1):81–115, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231597;   http://muse.jhu.edu/article/231597/pdf;

 https://direct.mit.edu/jcws/article-pdf/10/1/81/700166/
jcws.2008.10.1.81.pdf.

Xia:2008:WCK

[Xia08c] Yafeng Xia. William C. Kirby, Robert S. Ross, and Gong Li, eds., Normalization of U.S.–China Relations: an International History. Journal of Cold War Studies, 10(4):161–163, Fall 2008. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/252668;
http://muse.jhu.edu/article/252668/pdf;   https://direct.
mit.edu/jcws/article-pdf/10/4/161/697055/jcws.2008.10.4.
161.pdf.

Xia:2014:MCS

[Xia14] Yafeng Xia. Mao’s China and the Sino–Soviet Split: Ideological Dilemma by Mingjiang Li (review). Journal of Cold War Studies, 16(4):260–263, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566676;   http://muse.jhu.edu/article/566676/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/260/698655/
jcws_r_00495.pdf.

Xia:2023:BRS

[Xia23] Yafeng Xia. Book review: The Sino-Soviet Alliance: An International History by Austin Jersild. Journal of Cold War Studies, 25(1):233– 237, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/25/1/233/2073747/jcws_r_01136.pdf.

Xia:2023:CCW

[XL23] Yafeng Xia and Zhi Liang. China and the Cold War: Introduction. Journal of Cold War Studies, 25(2):5–10, Spring 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/2/
5/2140964/jcws_e_01139.pdf.

Yankova:2008:DPE

[Yan08] Gergana Yankova. Daniel Philpott, ed., The Politics of Past Evil: Religion, Reconciliation, and the Dilemmas of Transitional Justice. Journal of Cold War Studies, 10(3):191–193, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/241527;   http:
//muse.jhu.edu/article/241527/pdf;   https://direct.mit.

edu/jcws/article-pdf/10/3/191/697141/jcws.2008.10.3.191.
pdf.

Yaqub:2002:JFD

[Yaq02] Salim Yaqub. John Foster Dulles: Piety, Pragmatism, and Power in U.S. Foreign Policy (review). Journal of Cold War Studies, 4(3):123– 125, Summer 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9293;   http:/
/muse.jhu.edu/article/9293/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/3/123/700222/jcws.2002.4.3.123.pdf.

Yarrow:2009:SNV

[Yar09] Andrew L. Yarrow. Selling a new vision of America to the world: Changing messages in early U.S. Cold War print propaganda. Journal of Cold War Studies, 11(4):3–45, Fall 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/366245;   http://muse.jhu.edu/article/366245/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/4/3/1627639/
jcws.2009.11.4.3.pdf.

Li:2001:DCT

[yL01a] Hua yu Li. Daughter of China: a True Story of Love and Betrayal (review). Journal of Cold War Studies, 3(3):124–126, Fall 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9196;   http:/
/muse.jhu.edu/article/9196/pdf;   https://direct.mit.edu/
jcws/article-pdf/3/3/124/695480/jcws.2001.3.3.124.pdf.

Li:2001:PSC

[yL01b] Hua yu Li. The political stalinization of China: The establishment of one-party constitutionalism, 1948–1954. Journal of Cold War Studies, 3(2):28–47, Spring 2001. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/9191;   http://muse.jhu.edu/article/
9191/pdf;   https://direct.mit.edu/jcws/article-pdf/3/2/
28/695374/152039701300373871.pdf.

Li:2007:SZS

[yL07] Hua yu Li. Shen Zhihua, Sulian zhuanjia zai Zhongguo (1948– 1960) [Soviet Experts in China: 1948–1960]. Journal of Cold War Studies, 9(1):142–144, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/210090;   http://muse.jhu.edu/article/210090/pdf;

 https://direct.mit.edu/jcws/article-pdf/9/1/142/696558/
jcws.2007.9.1.142.pdf.

Li:2009:RCC

[yL09] Hua yu Li. Reactions of Chinese citizens to the death of Stalin: Internal Communist Party reports. Journal of Cold War Studies, 11(2):70–88, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/265885;   http://muse.jhu.edu/article/265885/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/2/70/697217/
jcws.2009.11.2.70.pdf.

Li:2012:YCC

[yL12] Hua yu Li. Yinghong Cheng, Creating the “New Man”: From Enlightenment Ideals to Socialist Realities. Honolulu: University of Hawai’i Press, 2009. 265 pp. $60.00. Journal of Cold War Studies, 14(1):128–129, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467693;   http://muse.jhu.edu/article/467693/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/1/128/697895/
jcws_r_00204.pdf.

Yong:2008:LCE

[Yon08] Tan Tai Yong. Lionel Carter, ed., Punjab Politics, 1 January 1944– 3 March 1947: Last Years of the Ministries Governors’ Fortnightly Reports and Other Key Documents. Journal of Cold War Studies, 10(3):162–164, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241513;   http://muse.jhu.edu/article/241513/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/162/697095/
jcws.2008.10.3.162.pdf.

Yordanov:2015:FIA

[Yor15] Radoslav A. Yordanov. Foreign Intervention in Africa: From the Cold War to the War on Terror by Elizabeth Schmidt, and: Buried in the Sands of the Ogaden: The United States, the Horn of Africa, and the Demise of Détente by Louise Woodroofe (reviews). Journal of Cold War Studies, 17(3):297–301, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595113;   http://muse.jhu.edu/article/595113/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/297/
1102466/jcws_r_00559.pdf.

Yordanov:2017:MOI

[Yor17a] Radoslav Yordanov. Mezhdunarodni otnosheniya [International Relations] by Nadia Boyadjieva (review). Journal of Cold War Studies, 19(4):234–236, Fall 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/683673;   http://muse.jhu.edu/article/683673/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/4/234/699377/
jcws_r_00773.pdf.

Yordanov:2017:PRC

[Yor17b] Radoslav Yordanov. Peter Ruggenthaler, The Concept of Neutrality in Stalin’s Foreign Policy, 1945–1953. Lanham, MD: Lexington Books, 2015. 442 pp. $120.00. Journal of Cold War Studies, 19(3):264–267, Summer 2017. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/668182;   http://muse.jhu.edu/article/668182/pdf;
 https://direct.mit.edu/jcws/article-pdf/19/3/264/699322/
jcws_r_00750.pdf.

Yordanov:2018:JJB

[Yor18] Radoslav Yordanov. Jeffrey James Byrne, Mecca of Revolution: Algeria, Decolonization, and the Third World Order. Journal of Cold War Studies, 20(4):229–232, Fall 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/718206;   https://direct.mit.
edu/jcws/article-pdf/20/4/229/1969628/jcws_r_00808.pdf.

Yordanov:2019:CCE

[Yor19a] Radoslav Yordanov. Comrades of Color: East Germany in the Cold War World ed. by Quinn Slobodian (review). Journal of Cold War Studies, 21(1):211–213, Winter 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/723546;   https://direct.mit.
edu/jcws/article-pdf/21/1/211/699787/jcws_r_00867.pdf.

Yordanov:2019:WPC

[Yor19b] Radoslav A. Yordanov. Warsaw pact countries’ involvement in Chile from Frei to Pinochet, 1964–1973. Journal of Cold War Studies, 21 (3):56–87, Summer 2019. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/731283;
 https://direct.mit.edu/jcws/article-pdf/21/3/56/699614/
jcws_a_00893.pdf.

Yordanov:2020:BCW

[Yor20a] Radoslav Yordanov. The Balkans in the Cold War ed. by Svetozar Rajak et al. (review). Journal of Cold War Studies, 22(1):253–255, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL https://direct.mit.edu/jcws/article-pdf/
22/1/253/699817/jcws_r_00913.pdf;   https://muse.jhu.edu/
article/749644.

Yordanov:2020:LJE

[Yor20b] Radoslav Yordanov. Leslie James and Elisabeth Leake, eds., Decolonization and the Cold War: Negotiating Independence. Journal of Cold War Studies, 22(3):260–262, Summer 2020. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/22/3/260/1860628/jcws_r_
00958.pdf;   https://muse.jhu.edu/article/762902.

Yordanov:2021:GUS

[Yor21] Radoslav Yordanov. Gangsterismo: The United States, Cuba and the Mafia, 1933 to 1966 by Jack Colhoun, and: Celia Sánchez Manduley: The Life and Legacy of a Cuban Revolutionary by Tiffany A. Sippial (reviews). Journal of Cold War Studies, 23(3):239–243, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

https://direct.mit.edu/jcws/article-pdf/23/3/239/1955767/
jcws_r_01033.pdf;   https://muse.jhu.edu/article/801913.

Yordanov:2023:BRE

[Yor23a] Radoslav Yordanov. Book review: Eisenhower & Cambodia: Diplomacy, Covert Action and the Origins of the Second Indochina War by William J. Rust. Journal of Cold War Studies, 25(1):230–233, Winter 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/1/
230/2073743/jcws_r_01135.pdf.

Yordanov:2023:LMC

[Yor23b] Radoslav Yordanov. The long misunderstanding: Cuba’s economic ties with the Soviet Bloc. Journal of Cold War Studies, 25(4):24–52, Fall 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/4/
24/2207540/jcws_a_01169.pdf.

Yossef:2015:KSM

[Yos15] Amr Yossef. Kalemat as-Ser: Mudhakkirat Mohamed Hosni Mubarak: Yuniu 1967–October 1973 by Mohamed Hosni Mubarak (review). Journal of Cold War Studies, 17(2):152–154, Spring 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/582443;   http:
//muse.jhu.edu/article/582443/pdf;   https://direct.mit.
edu/jcws/article-pdf/17/2/152/698700/jcws_r_00541.pdf.

Young:2007:MSR

[You07] Ken Young. A most special relationship: The origins of Anglo– American nuclear strike planning. Journal of Cold War Studies, 9 (2):5–31, Spring 2007. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/214758;
http://muse.jhu.edu/article/214758/pdf;   https://direct.
mit.edu/jcws/article-pdf/9/2/5/696498/jcws.2007.9.2.5.pdf.

Young:2012:JKA

[You12a] Crawford Young. John Kent, America, the UN and Decolonisation: Cold War Conflict in the Congo. London: Routledge, 2010. 244 pp. Journal of Cold War Studies, 14(4):246–248, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/500346;   http:
//muse.jhu.edu/article/500346/pdf;   https://direct.mit.
edu/jcws/article-pdf/14/4/246/698163/jcws_r_00293.pdf.

Young:2012:AJV

[You12b] Marilyn B. Young. Andrew L. Johns, Vietnam’s Second Front: Domestic Politics, the Republican Party, and the War. Lexington, KY: University Press of Kentucky, 2012. x + 434 pp. Journal of Cold War Studies, 14(4):231–232, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/500339;   http://muse.jhu.edu/article/500339/pdf;
 https://direct.mit.edu/jcws/article-pdf/14/4/231/698137/
jcws_r_00286.pdf.

Young:2013:RN

[You13a] Ken Young. Revisiting NSC 68. Journal of Cold War Studies, 15 (1):3–33, Winter 2013. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/505769;
http://muse.jhu.edu/article/505769/pdf;   https://direct.
mit.edu/jcws/article-pdf/15/1/3/699966/jcws_a_00306.pdf.

Youngblood:2013:JMS

[You13b] Denise J. Youngblood. Jamie Miller, Soviet Cinema: Politics and Persuasion under Stalin. London: I. B. Tauris, 2010. xv + 224 pp. Journal of Cold War Studies, 15(2):161–163, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/509719;   http://muse.jhu.edu/article/509719/
pdf;   https://direct.mit.edu/jcws/article-pdf/15/2/161/
698210/jcws_r_00360.pdf.

Young:2022:SEC

[You22] Glennys J. Young. Spain and the Early Cold War: The “Isolation paradigm” revisited. Journal of Cold War Studies, 24(3):43–79, Summer 2022. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/24/3/
43/2042020/jcws_a_01088.pdf.

You:2023:HBA

[You23] Lan You. Hanoi’s balancing act: The Vietnamese Communists and the Sino-Soviet Split, 1960–1965. Journal of Cold War Studies, 25 (2):64–92, Spring 2023. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/25/2/64/2140950/jcws_a_01142.pdf.

Zabecki:2016:DKE

[Zab16] David T. Zabecki. Dieter Krüger, ed., Schlachtfeld Fulda Gap. Fulda, Germany: Parzellers Buchverlag, 2014. 313 pp. EUR 17.95. Journal of Cold War Studies, 18(3):222–225, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/632300;   http://muse.jhu.edu/article/632300/
pdf;   https://direct.mit.edu/jcws/article-pdf/18/3/222/
699059/jcws_r_00667.pdf.

Zake:2010:SCA

[Zak10] Ieva Zake. Soviet campaigns against “capitalist ideological subversives” during the Cold War: The Latvian experience. Journal of Cold War Studies, 12(3):91–114, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/392779;   http://muse.jhu.edu/article/392779/pdf;
 https://direct.mit.edu/jcws/article-pdf/12/3/91/697608/
jcws_a_00007.pdf.

Zake:2018:SIF

[Zak18] Ieva Zake. Soviet Inturist and foreign travel to the Latvian SSR in the post-Stalin era: A case of ethnic tourism. Journal of Cold War Studies, 20(2):38–62, Spring 2018. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/697060;   https://direct.mit.
edu/jcws/article-pdf/20/2/38/699401/jcws_a_00816.pdf.

Zaloga:2004:RWY

[Zal04] Steven J. Zaloga. Red Wings over the Yalu: China, the Soviet Union, and the Air War in Korea (review). Journal of Cold War Studies, 6(3):164–166, Summer 2004. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/171178;   http://muse.jhu.edu/article/171178/pdf;
 https://direct.mit.edu/jcws/article-pdf/6/3/164/696109/
jcws.2004.6.3.164.pdf.

Zaretsky:2021:BBP

[Zar21] Natasha Zaretsky. Battling Bella: The Protest Politics of Bella Abzug by Leandra Zarnow, Cambridge, MA: Harvard University Press, 2019. 441 pp. $35.00 (review). Journal of Cold War Studies, 23(1):229–231, Winter 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/1/229/1911820/jcws_r_00987.pdf;
https://muse.jhu.edu/article/788398.

Zivotic:2016:RBY

[ŽČ16] Aleksandar Životić and Jovan Čavoški. On the road to Belgrade: Yugoslavia, Third World neutrals, and the evolution of global nonalignment, 1954–1961. Journal of Cold War Studies, 18(4):79– 97, Fall 2016. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/645909;
http://muse.jhu.edu/article/645909/pdf;   https://direct.
mit.edu/jcws/article-pdf/18/4/79/700075/jcws_a_00681.pdf.

Zeitlin:2008:VGI

[Zei08] Jonathan Zeitlin. Victoria de Grazia, Irresistible Empire: America’s Advance through Twentieth-Century Europe. Journal of Cold War Studies, 10(3):189–191, Summer 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/241526;   http://muse.jhu.edu/article/241526/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/3/189/697131/
jcws.2008.10.3.189.pdf.

Zhai:2002:CCC

[Zha02] Qiang Zhai. Contending with Contradictions: China’s Policy toward Soviet Eastern Europe and the Origins of the Sino–Soviet Split, 1953–1960 (review). Journal of Cold War Studies, 4(4):132– 133, Fall 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9313;   http:/
/muse.jhu.edu/article/9313/pdf;   https://direct.mit.edu/
jcws/article-pdf/4/4/132/695698/jcws.2002.4.4.132.pdf.

Zhai:2006:TCB

[Zha06] Qiang Zhai. Tibet and Chinese-British-American relations in the early 1950s. Journal of Cold War Studies, 8(3):34–53, Summer 2006. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/201242;   http:
//muse.jhu.edu/article/201242/pdf;   https://direct.mit.
edu/jcws/article-pdf/8/3/34/696415/jcws.2006.8.3.34.pdf.

Zhai:2009:PBC

[Zha09] Qiang Zhai. Pierre Brocheux, Ho Chi Minh: a Biography. Journal of Cold War Studies, 11(2):139–141, Spring 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/265874;   http://muse.jhu.edu/article/265874/
pdf;   https://direct.mit.edu/jcws/article-pdf/11/2/139/
697187/jcws.2009.11.2.139.pdf.

Zhang:2010:DXC

[Zha10] Xiaoming Zhang. Deng Xiaoping and China’s decision to go to war with Vietnam. Journal of Cold War Studies, 12(3):3–29, Summer 2010. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/392776;
http://muse.jhu.edu/article/392776/pdf;   https://direct.
mit.edu/jcws/article-pdf/12/3/3/697598/jcws_a_00001.pdf.

Zhai:2012:C

[Zha12a] Qiang Zhai. Commentary. Journal of Cold War Studies, 14(1):99– 103, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL http://muse.jhu.edu/article/467683;
http://muse.jhu.edu/article/467683/pdf.

Zhai:2012:NKC

[Zha12b] Qiang Zhai. Nicholas Khoo, Collateral Damage: Sino-Soviet Rivalry and the Termination of the Sino-Vietnamese Alliance. New York: Columbia University Press, 2011. x + 267 pp. Journal of Cold War Studies, 14(1):125–126, Winter 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/467691;   http://muse.jhu.edu/article/467691/pdf;

 https://direct.mit.edu/jcws/article-pdf/14/1/125/697851/
jcws_r_00202.pdf.

Zhai:2015:BAC

[Zha15] Qiang Zhai. Brothers in Arms: Chinese Aid to the Khmer Rouge, 1975–1979 by Andrew Mertha (review). Journal of Cold War Studies, 17(3):275–277, Summer 2015. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/595102;   http://muse.jhu.edu/article/595102/
pdf;   https://direct.mit.edu/jcws/article-pdf/17/3/275/
1102412/jcws_r_00522.pdf.

Zhang:2023:SVM

[Zha23a] Yang Zhang. Strategic vigilance: Mao’s “Anti-Peaceful evolution” strategy and China’s policy toward the United States, 1959–1976. Journal of Cold War Studies, 25(2):93–111, Spring 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/25/2/
93/2140966/jcws_a_01143.pdf.

Zhao:2023:LCN

[Zha23b] Xuegong Zhao. The limits of confrontation: Nuclear weapons, the 1958 Taiwan Strait Crisis, and China–U.S. relations. Journal of Cold War Studies, 25(2):112–149, Spring 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.edu/

jcws/article-pdf/25/2/112/2140960/jcws_a_01144.pdf.

Zhihua:2000:SSR

[Zhi00] Shen Zhihua. Sino–Soviet relations and the origins of the Korean war: Stalin’s strategic goals in the Far East. Journal of Cold War Studies, 2(2):44–68, Spring 2000. CODEN ???? ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/9152;
 http://muse.jhu.edu/article/9152/pdf;   https://direct.
mit.edu/jcws/article-pdf/2/2/44/695216/15203970051032309.
pdf.

Zhuk:2016:KTV

[Zhu16] Sergei I. Zhuk. Kiril Tomoff, Virtuosi Abroad: Soviet Music and Imperial Competition during the Early Cold War, 1945–1958. Ithaca, NY: Cornell University Press, 2015. xi + 262 pp. $27.95. Journal of Cold War Studies, 18(3):229–231, Summer 2016. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/632302;   http://muse.jhu.edu/article/632302/pdf;

 https://direct.mit.edu/jcws/article-pdf/18/3/229/699055/
jcws_r_00669.pdf.

Zhuk:2020:SKB

[Zhu20] Sergei I. Zhuk. Swans of the Kremlin: Ballet and Power in Soviet Russia by Christina Ezrahi (review). Journal of Cold War Studies, 22(1): 262–263, Winter 2020. CODEN ???? ISSN 1520-3972 (print), 1531-

3298 (electronic). URL https://direct.mit.edu/jcws/article-
pdf/22/1/262/699841/jcws_r_00916.pdf;   https://muse.jhu.
edu/article/749639.

Zierler:2008:ANC

[Zie08] David Zierler. Arvid Nelson, Cold War Ecology: Forests, Farms, and People in the East German Landscape, 1945–1989. Journal of Cold War Studies, 10(1):151–153, Winter 2008. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/231598;   http://muse.jhu.edu/article/231598/pdf;
 https://direct.mit.edu/jcws/article-pdf/10/1/151/700138/
jcws.2008.10.1.151.pdf.

Ziegler:2021:PPP

[Zie21] Charles E. Ziegler. Project Plowshare: The Peaceful Use of Nuclear Explosives in Cold War America by Scott Kaufman (review). Journal of Cold War Studies, 23(3):237–239, Summer 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/23/3/237/1955763/jcws_r_
01032.pdf;   https://muse.jhu.edu/article/801912.

Zimmermann:2009:IPC

[Zim09] Hubert Zimmermann. The improbable permanence of a commitment: America’s troop presence in Europe during the Cold War. Journal of Cold War Studies, 11(1):3–27, Winter 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/259183;   http://muse.jhu.edu/article/259183/pdf;
 https://direct.mit.edu/jcws/article-pdf/11/1/3/1468559/
jcws.2009.11.1.3.pdf.

Zinoman:2011:NVG

[Zin11] Peter Zinoman. Nhân Văn–Giai Ph´ âm and Vietnamese “Reform Communism” in the 1950s: a revisionist interpretation. Journal of Cold War Studies, 13(1):60–100, Winter 2011. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/424241;   http://muse.jhu.edu/article/424241/

pdf;   https://direct.mit.edu/jcws/article-pdf/13/1/60/697709/
jcws_a_00071.pdf.

Zlotnik:2003:YGP

[Zlo03] Marc D. Zlotnik. Yeltsin and Gorbachev: The politics of confrontation. Journal of Cold War Studies, 5(1):128–164, Winter 2003. CO-DEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/38257;   http://muse.jhu.edu/
article/38257/pdf;   https://direct.mit.edu/jcws/article-
pdf/5/1/128/695790/152039703320996740.pdf.

Zhang:2023:QPU

[ZP23] Weizhen Zhang and Tao Peng. The Qingdao Pattern and U.S.– Chinese Crisis management: The KMT, the CCP, and the U.S. Marines in Qingdao during the Chinese Civil War (1945–1949). Journal of Cold War Studies, 25(2):150–178, Spring 2023. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https:/

/direct.mit.edu/jcws/article-pdf/25/2/150/2140954/jcws_a_
01145.pdf.

Zelikow:2021:ECW

[ZRS^{+}21] Philip Zelikow, Condoleezza Rice, Kristina Spohr, James Goldgeier, and Vojtech Mastny. Ending the Cold War and entering a new era: Perspectives on To Build a Better World. Journal of Cold War Studies, 23(4):181–210, Fall 2021. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL https://direct.mit.

edu/jcws/article-pdf/23/4/181/1970079/jcws_c_01029.pdf;
https://muse.jhu.edu/article/835881.

Zubok:1999:AKS

[Zub99] Vladislav M. Zubok. Andrei A. Kokoshin, Soviet Strategic Thought, 1917–91. Journal of Cold War Studies, 1(2):129–131, Spring 1999. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL https://direct.mit.edu/jcws/article-pdf/1/2/
132/695139/15203979952559559.pdf.

Zubok:2013:LTU

[Zub13] Vladislav Zubok. Lost in a triangle: U.S.–Soviet back-channel documents on the Japan factor in tripartite diplomacy, 1969–1972. Journal of Cold War Studies, 15(2):51–71, Spring 2013. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/509703;   http://muse.jhu.edu/article/509703/

pdf;   https://direct.mit.edu/jcws/article-pdf/15/2/51/698157/
jcws_a_00337.pdf.

Zubok:2014:RNM

[Zub14] Vladislav Zubok. The Readers of Novyi Mir: Coming to Terms with the Stalinist Past by Denis Kozlov (review). Journal of Cold War Studies, 16(4):281–283, Fall 2014. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.edu/

article/566685;   http://muse.jhu.edu/article/566685/pdf;
 https://direct.mit.edu/jcws/article-pdf/16/4/281/698676/
jcws_r_00504.pdf.

Zuk:2019:SWI

[Żuk19] Piotr Żuk. The secret war of intelligence: The mysterious mission of “Jack Strong” and its impact on the Cold War in the 1970s and 1980s. Journal of Cold War Studies, 21(1):166–185, Winter 2019. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL

http://muse.jhu.edu/article/723557;   https://direct.mit.
edu/jcws/article-pdf/21/1/166/699719/jcws_a_00872.pdf.